A Complete Bibliography of Publications in Journal of the ACM

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

10 November 2025 Version 1.226

Title word cross-reference

#NFA [ACJR21].

(2 + ϵ) [BYCHS17]. (Δ + 1) [Bar16, BEG22, HSS18]. (s, t) [ATSWZ00]. 1 [HV95, MM24]. 1 − 1/e^{2} [JL23]. 1/2N^{3} [Yen72]. 2 [AF67, BKSW24, Cha10, CL16, DG16, GR79, HMZ13, HdLT01, Lar76]. 3 [Bul06, Cha10, CNP85, KT17, KPS19, SS13, Shi10]. 3/2 [TV19]. 4/3 [AB17, CG93]. 5 [Wat61]. 6 [Wat61]. 8 [Wat61]. = [JJUW11]. ^{1} [MT88]. A [Bui79, SB73]. A^{∗} [MP10]. AC^{0} [Bra10]. C [Sla72a]. ∩ [AR05, FGHS23]. Δ [Bar16]. E [GNRS92]. ℓ_{1} [BC05]. ℓ_{p} [AKPS24]. ϵ [AKN^{+}08, GW21, RA20]. ∃R [AKM25]. ∃R [AAM22]. f(z) [RS63]. f(z) = 0 [War57]. g [HI92]. H [DFHT05]. J [CT81]. K [OOW99, AMR11, BBMN15, BG00, BP03a, CK95a, CeLT00, CGH^{+}05, DKK24, FGYZ21, Gui78, GHLL22, HSTV14, JV01, KLM19, KP95, Lin18, MLS76, Mol18, ORSS12, PPSZ05, YR78, KIM81]. k + 1 [YR78]. K_{m} [vdLL86]. L [Bui79, San65]. l_{1} [KV15, Abd71]. L_{p} [Lee80]. Λ [Gri68]. ln n [Fei98]. log^{∗} n [CGH^{+}05]. log^{c} n [BR91]. m [AM78b, DG62].

M′ = ∨_{i=1}M^{i} [Sim67, Sim65]. R^{3} [dMRST20]. m × n [NJ57]. n [CK95a, Mey84, MO68, Mou65a, Mow67, Was81, And66, BJL^{+}82, HP59, Kac79, Mey88, WW73, Yen72]. n + 1 [Mou65a]. n/2/F/\overline{F} [KS75]. n^{1/5} [KT17]. N^{3} [Yen72]. NC [BR91]. O(1) [MMR15, FKS84]. O(ϵ^{2}) [Van85]. O(g) [HI92]. O(log(n)^{4/3}) [ATSWZ00]. O(log^{2} n) [BHR19]. O(log Δ/ϵ log log Δ) [BYCHS17]. O(log n) [Bra87]. O(sort(n)) [BM11]. O(n) [FG05, CFM10, Upf92]. O(n^{2}(m + n log n) log n) [GT88a]. O(n^{2}) [DI05, KS11, MMR15, PSSZ13]. O(n^{3}) [GHL^{+}87]. O(n^{3} log n) [GHL^{+}87]. O(n log n) [BK09, FG05, HvdH22, Hwa79]. ω [BGB12, EKS20, CH84]. Ω(n^{k}) [AIK84]. ω_{n} [She57b]. P [All89, SG76]. P = 2 [CM64]. ∂^{2}u/∂x^{2} + ∂^{2}u/∂y^{2} + (K/y)(∂u/∂y) = 0 [Ehr59]. Q [GLH57]. R [Sla76]. R^{3} [DG98]. s [AKS15, TV19]. T [NOT06, AKS15, TV19]. t < n/3 [MMR15]. T_{0} [Das77]. Θ(log n) [FPPR22]. W [Lee61]. X + Y [JK78]. X_{t} = X_{t−3p} ⊕ X_{t−3q} [Fus90]. Z [Dix73, Wan93].

-Approximation [BYCHS17, RA20]. -ary [Gui78]. -Automata [EKS20, BGB12]. -Biclique [Lin18]. -Bit [MM24]. -Body [CK95a]. -center [CGH^{+}05]. -CNFs [FPPR22]. -Colorable [KT17]. -Coloring [Bar16, HSS18, BEG22]. -Colourings [Mol18]. -Complete [SG76, AAM22, AKM25]. -Complexes [GR79]. -cut [GHLL22]. -D [Cha10, Shi10]. -Dimensional [Mey84, Mey88]. -Domination [Sla76]. -Edge [CNP85, HdLT01]. -element [Bul06]. -Evaluation [Was81]. -Free [Gri68]. -Groups [Lar76]. -hard [CGH^{+}05]. -Languages [San65]. -Linear [Sla72a]. -Machine [Lee61]. -Maximal [CT81]. -Means [AMR11, ORSS12]. -Median [JV01]. -metric [Lee80]. -Minimal [CT81]. -minor-free [DFHT05]. -Module [Wan93]. -Nearest-Neighbors [CK95a]. -nets [AKN^{+}08]. -Node [WW73, Yen72]. -path [TV19]. -Player [HMZ13]. -Process [BJL^{+}82]. -Regression [AKPS24]. -Resolution [Dix73]. -SAT [FGYZ21, PPSZ05]. -Sequence [CH84]. -Sequences [AM78b]. -Server [BBMN15, DG16, BG00, KP95]. -set [CeLT00]. -Stable [Bui79, SB73]. -State [AF67, Wat61]. -SUM [DKK24, KLM19]. -Symbol [Wat61]. -Uniform [All89]. -Variables [And66]. -Wise [BR91]. -Writer [HV95]. -XOR [DKK24].

0/1 [Sah75]. 0S [ER80a].

1 [Rid76]. 1103 [Bau56]. 1955 [Hou56c].

2-Approximation [AGH^{+}25]. 2.0 [tCL09]. 2006 [VV09]. 2007 [LV09]. 2314 [ADW68].

3-query [Yek08].

407 [Boy57].

50th [Hal03].

60 [Bot62, BG70, RR64]. 650 [Gor56, HK54].

701 [Bak56, BO56, GKR55]. 702 [BBR54].

= [FGHS23, Sha92, She92].

A* [Dav90, DP85, SW91]. Abacus [Li59]. Abbreviating [BF61]. Abduction [CDOY11, EG95]. abductive [EM07]. ability [NS06a]. Absence [KBI^{+}17]. Absolute [CK65, Gab16]. Abstract [Ehr82, FLO^{+}25, LR72, McL84, Mil73, Nou83, Ove66, BT07, CCM12, GRS00, NOT06, GB92, Hue80]. Abstraction [EPT18, HHHN14, Imi91, vW96, BN01, CGJ^{+}03, DW99, GS07, SP07]. Abstractions [ADH15]. Academy [Leb56]. ACC [Wil14]. Accelerated [HST76, ST72]. Accelerating [Fox70]. Acceleration [AP25, Tro84]. Accept [Ste67]. Acceptable [Sch68]. Acceptance [BR54, Gra55a, Hen85]. Accepted [Gin61b]. Access [BFGY19, Bsh93, ER64b, FKS84, GS69, GW85, GFL87, GS73, Low68, MY79, Mey85, Rei87, SS79, Sch88, Tob82, Tra63, AW08, BEP09, HHR97]. accesses [GHS09]. Accessing [BK71]. Account [Oet57]. Accounting [Tre54]. Accumulation [KL70]. Accuracy [KS75, Ral59, AC24]. accurate [XB06]. Achieve [Ral59]. achieving [DMR08]. Ackermann [Cha00b]. acknowledgment [DGS01]. ACM [Ano79, Gra77, Hou56c, Alt03, Ano68, BJP59, Sal70]. Acquisition [SH22]. Action [LS95]. Actions [Set82]. Activity [SD62]. Acyclic [BMT89, BFMY83, HT24, San88, DGP07, GLS01, Ott12]. Acyclicity [Fag83]. Adams [Alo60]. Adaptation [Pea68]. adapting [AF03]. Adaptive [Bla25, CK95b, DM25, GN94, HRS24, Hol62b, LR79, Lyn69, PGL76, Ric75, ŠVV09, CST^{+}18, KS11]. Addendum [CH88, EV62a]. Adding [AM09, MS16b, WL85]. Addition [BEF^{+}21, Win65a, YSS78]. Additional [Leu93, BB64]. Additions [Yen72]. Additive [AB17, BILW20, BSLRZ14, GSK59]. Address [Car57, Elg54, Eli74, Flo60, Hou56c, Hou57, IS56, Lip57, PK87, TTK66]. Addressable [Ros72]. Addressed [Kau71]. Addresses [BAG92]. Addressing [AGOW66, Tra63]. Addressless [Bli72]. Adjacency [ADK17, DEG^{+}21]. Adjusting [ST85]. admission [LNT97]. Admits [ACJR21]. Advanced [GM56]. Advances [DFN^{+}24, GHS75]. Advantage [YZ24]. Adversarial [BKR^{+}01, BKN14, ISSS22]. Adversarially [BEJWY22, HKM^{+}22]. Adversaries [HRS24, HT93]. adversary [KS11]. Adverse [Mad66]. Advice [LV21, CBFH^{+}97, PM06]. AdWords [GML15, MSVV07]. AFA [GG72]. AFL [GH70]. After [GM56]. Again [Szp90, de 83]. Against [LST25, BKN14]. Age [Got03]. Agents [CGM24, DHK19, GH23, HR25]. Aggregate [Klu82, RSG05, CNS07, HLNW01]. Aggregated [Ste83]. Aggregating [ACN08]. Aggregation [CS85, Kan83, Van78, Van85]. Aggregation/Disaggregation [CS85]. Aging [Via87]. Agreement [ADLS94, BCP20, DR85, DLP^{+}86, DRS90, HPZ24, KS16, CeLT00, KS11, LLR06]. Agreements [PSL80]. Ahead [San69]. AI [Red03]. Aid [KH64]. Aided [BT69, GR74]. Aids [Pra60]. AIMD [WSY^{+}19]. Aitken [GL64]. Aitken-Hermite [GL64]. Algebra [BHK90, GTWW77, Klu82, NB95, Sal66, SV85, Wil62, Yeh68, YR64, KJJ03, LW08]. Algebraic [ACPS93, BBKO21, BGKM23, BHR84a, BG15, Bra67, BIZ18, DSS22, ES17, FL18, GUK^{+}16, GMM81, GX22, Hem64, HM85a, HOPW23, IW91, Ja’83, KM20, KT78, Lan59, LST25, LKK83, LFKN92, LT93, Mar71, Pic67, vdLL86, BR96, IOT19, MP24]. Algebraic-geometric [GX22]. Algebras [AGS17, BK72, FZ77, GTWW77, Liu69, Mil79, Nip90, Ven87, WM74, BT95]. ALGOL [Bot62, BG70, GR62, GR63b, RR64, OR00]. ALGOL-Like [GR63b]. Algorithm [AM85, AKS15, AGG62, ACW91, Bak77, BMPT91, BBN12, BBMN15, BRK74, Bau85, BSP82, BCS92, BDS87, BBB^{+}21, BLS92, Bro71, BT71, BT79, CK70, CE92, CL16, CG86, CG70b, CK65, CF59, Deo76, DFK91, EHR84, ES17, Elk20a, Fei84, FP94, FW74, Gab76, Gab82, GS78a, GM79, Gal82, GT88a, Gal95, GNP^{+}93, Gib69, GJ80, GL88b, Goo11, GW76, GR79, GLW70, HW90, HPR15, Hib63, HW76, Hor75, HP76, HLS77, Hsu75, Hwa79, ILG87, Joh73, Jon62, KS78, Kam87, Kan80, KL86, KKT95, KW85, KIM81, Kri64, LH90, LP79, LL85, Leh74, LP73, LS77, LM86, LB79, Man75, Man79a, Man79b, Mar76, MP72, McC76, Mey84, Moo67, Mor68a, Mow67, MT69, MC72, Mul91, Mye92, NS80, OV20, ON68, Pag76, RA20, Rah70, Reg25, Rei86, RPE81, RW88]. Algorithm [RP89, Sag80, SK73, SD76, SS17, Sha67a, Sin86, Sla64, ST73, Sti81, Sto70, Sto73, STV20, Tar75, TS74, TSOA80, Ull76, VV63, Wag76, Wan74a, Wan23, Wat73, Wei72, Wil78, WW73, Yan85, Zad72, Zoh69, Zol78, dCS77, ATSWZ00, AMN^{+}98, BG00, BK09, CE00, Cha00b, CLL^{+}08, CHL01, CJK^{+}06, DPSV08, HP99, Hoc01, HRM02, IFF01, JSV04, KP96, LZ01, Mye99, OOW99, PPSZ05, PR02, ST04, SW97, Vaz12, Vem10]. Algorithmic [AC24, BG70, CG70a, CGS15, EK72, Gil66a, Lat81, Per67, RZ19, BLP09, CD08]. Algorithms [ADF^{+}16, AKPS24, ABC17, AMOT90, ACHS16, ABS15, Ata94, ALS94, AF03, ADS83, BM83, BCW85, Bak94, BDD^{+}24, BDB23, BEJWY22, BBBL95, BJM79, BP03b, BPP^{+}16, BGSS20, Blu71, Blu94, BK78, Bro82, BS72, CLTZ18, CCMS22, CH14, CF12, Cla95, Cof68, CaKMTM19, CH79, CM93, Col87, CDRS93, CD94, DJ81, DJSW19, DM81, Ehr73, Fai85, Fal62, FGR^{+}17, Flo67, FLPS16, FGLS19, Fre79, FT87, Fri73, GGS89, GT91, Gav83, Gil68, GS93, GS85, GL85, GK88, HKM^{+}22, Hei71, Hel89, HLMW92, Hir77, HS86, HS87, HS76, HT24, HD73, Iba77, IK75, IK77, IM83, IKM17, IK74, ICR87, JSS16, Joh77, Joh87, KU76, KKMO18, KR93, KZ93, KN81, KS74, KS75, KSS75, KSS77, Kun76, La 64, LR79, Len89, LL73, MF63, MB83, Meg83, Mey88, MP65]. Algorithms [MP66a, MW84, Mot94, NS82, NPRR18, OR90, Ove74, PKR84, Pag70, PKT90, Rac82, Rei72, RC86, Sah75, Sah76, Sch80, She67, ST72, SR71, Tar81a, Tv84, Tro84, TS95, ÜD90, Yao80, vdSS89, AT11, BSZ06, CKL13, DFHT05, DFSL14, DHW97, EGIN97, ENRS00, FK99, FGK09, FKV04, GKPS06, GW95, GMS10, GKPR07, Hal07, HK99, HdLT01, JV01, JMM^{+}03, JSMM03, KLSS05, KT02, LR99, Pel98, SCD02, Sho03, Spe97, ST04, XB06, GG78]. All-pairs [PSSZ13]. Allen [KJJ03, NB95]. Allocating [Ros74, Ros75]. Allocation [BCW85, BN79, BCS92, Bro82, CC76a, CC76b, DL56, DJSW19, GL71, HKMW66, KIM81, PGL76, Rob71, Rob74, TW88, BNBYF^{+}01]. Allocations [LS23]. Allow [RT88]. Allowing [BC74, Rem81]. Allows [DV14]. Almost [AM94, AGPT16, AZ08, CGL^{+}23, CKL^{+}07, CF12, CKL^{+}25, CJLW22, CADMD21, ER64a, Gab82, Gal82, GB74, GSV84, Kol04, LR19, Wil59b, BZ23, BLQT22, GS06]. Almost-Linear [CKL^{+}25, Gab82, GS06]. Almost-tight [AZ08]. ALOHA [Tob82]. Alpha [Pit06, Yer66]. Alpha-structural [Pit06]. Alphabets [GX22]. Alphameric [BO56]. Alphanumeric [Mor68a]. Alternating [AHK02, CH14, CD60, DG62, MT85]. Alternating-time [AHK02]. Alternation [CKS81, PZ19]. ambients [MN05]. Ambiguities [Gor63]. Ambiguity [Bel60, Can62, GU66a, HU66, Mau69, Ros70a]. Ambiquity [GU66b]. Among [FMI80, HU66, JMMW79, PF79, SY80, You77, CNS07]. Amortized [BHR19, Cha01]. Amphisbaenic [Nag59, Liv60, Nag61]. amplification [CKOR14, Din07]. Amplifiers [GGG^{+}17, HK55, HA56]. Amplifying [AK10]. Analog [BW57a, Edw54, Fis58, HS55, Mos55, SR60, Wad56, ZL58]. Analog-Digital [BW57a]. Analogs [Sha66]. Analogue [Ehr59, Fis56, Jac54, LWS55]. analyses [McA02]. Analysing [CG18]. Analysis [ADW68, Abe68, AMR11, AM70, Bak62, BB77, BN71, Bla25, BBH^{+}87, BKK14, BR15, BS89, Bur76, But78, CDOY11, CW65, Cof68, Cof69a, CG86, CPT94, CM57, CM67, Dad88, Dav17, Deb92, DFGN24, DI87, Doy65, EBA72, FHV91, Flo61a, Flo63, Fra69, Fra58b, FB75, GKL24, GG78, Gav71, Gil60, Gil65, Gla91, GLPR21, GW76, Gui78, HRS24, HP72b, Hem64, HC88, Hoc65, Hof84, Hou56a, Jef56, KU76, KT69, KR92, KTU94, Kou77, KSS75, Kri64, Lar83, Lat81, Lie59, MY80, Mil73, Moc71, MT19, Mor68b, Mot94, NBC72, Neg66, Oli67, OM78, One75, PKT90, Rah70, Ram66, Rao78, RL80, RL81, Ros79a, RT90, Sal73, San88, SS17, SAKS64, Smi84, ST73, Sti65, Sur87, Tv84, TTK66, Vit83, Vit87, Wal61, Wel66, Wil61, Win65b, WM74, You58]. Analysis [YL77, Zar59, ZS99, AW08, AS13, BG01, CLMW11, CCM12, DP01, DGS01, EKL10, FGK09, HMSS01, HHR97, LMW05, Man01, McA02, Pel98, RV07, SSS06, SRM97, SPK13, SSS96, ST04, Cof69b, KSS77]. Analyst [Wil71]. analytic [Drm03]. Analytical [Har64, YDL93]. Analyze [JM77]. Analyzed [Kle66, JMM^{+}03]. Analyzer [CF68, Les58, Wad54]. Analyzing [AB05, Gof71, Hae16]. Anarchy [GT23, Rou15]. Ancestry [FK16]. and/or [Hun84, MB85, SL71]. Annealing [SH88, ŠVV09]. anniversary [Hal03]. Announcement [Ano58a]. Anomalies [BK86, BK87]. Anonymous [AAHK94, ASW88, KM20, Tau22]. Answer [Nag61]. Answering [CM87, Coo64, Koc69, Win82, AL08]. Antiferromagnetic [GSV15]. Antonymy [LBB67]. Any [BM92, BHT18, LP15, Was80, KSS10, PW93]. Aperiodic [BMT92]. API [BFM^{+}19]. Applicable [Lie59]. Application [AM78b, BHK59, BG77, Bre70, CLRS86, Doy65, FM79, Gil63, GSV84, Jon62, Kob74a, Kob74b, Mos58, Sag80, Sch61, Sup83, YR64, Zaf70, JRS09, SS06b]. Applications [AEK^{+}12, BEG22, Bau58b, BKST78, BGKM23, BLMS97, CK95a, CW67, CK15, CDRS93, CGS15, EKLS20, FNFSZ23, FLPS16, FK16, GKL24, Haj83, Har86, Har64, Hib62, HL87, HMR83, HS74, Hue80, JP84, Kob74b, KM20, MSM85, MO68, OM78, Orm61, PV76, RC69, ST72, Tar79, TZ25, Val15, Ver95, Yeh68, YW73, ANP^{+}04, AZ05, AAF^{+}97, Bas99, BDG^{+}13, CKL13, FG99, FLMM09, GIS99, GKPS06, GK96, LNT97]. Applied [ABF17, BK77b, BFGL79, Jac54, RTW94]. Applying [Meg83]. Approach [AV89, AK16, BGP84, BL90, BBKO21, Bea72, Bea73a, Bea73b, BSLRZ14, Dav17, DST92a, DST92b, ER64b, FKRV15, FM70, FG79, GMM81, GT88b, GS93, GLJT23, GPV89, HS86, KR77, KMP^{+}12, KS63, Law64, MY80, Rot70, Sla72a, Tar81b, TKDB24, TG82, WYCF14, AT10, AL05, BNBYF^{+}01, BLR13, BFU01, Dar03, DI04, Drm03, FGK09, FGY00, GE01, KS96, KMPS09, KVW00, MP24, RV07]. Approaches [BM85, BK72]. Approaching [GKKS14, HS21, RZR24, TV19]. Approximability [CGSV24, DJKK08]. Approximate [AAGT15, Bab16, CLRS16, CMS20, Cry70, DLW25, DLP^{+}86, FNS21, Gel75, HS76, Hou58a, KMS98, KS75, KPW18, LPSP15, MS19, Moi19, OV23, PKT90, Sah75, TZ05, Wig83, AMS05, AMN^{+}98, AMM09, Cha00a, CGH^{+}05, Mye99, Tho04]. Approximated [PW93]. Approximately [BILW20, Hab70, LOS02]. Approximating [AHPSV97, AHPV04, BEG^{+}21, CDG^{+}20, CDPP14, CR86a, CvH65, DFK91, EKLS20, GJ12, LY94, MS11, RWZ23, SLM19, San96, Sig67, SL15, You56, BNBYF^{+}01, BEP09, CRS01, FGL^{+}96, Fei98, Kho05]. Approximation [Abd71, AHPW19, AGH^{+}25, Bak94, BYCHS17, BHM11, BKSW24, BC61, BLTY94, Blu94, BH65, BGRS13, Cha16, CLTZ18, CL16, CW17, CAFS21, Con57, Cru76, CF59, DDFS14, DJSW19, DP14, Dun74, Eda22, FLS18, Fre79, GLH57, GMS10, GKPR07, HM85b, HS86, HS87, IK75, JV01, KLSS05, KT02, Kob74a, Kob74b, Kur62, MV25, Mas56, MSU99, RA20, Rid76, SG76, Sha70a, She57a, Shi69, Sto64b, STV20, Aro98, ALM^{+}98, BBG13, Coh00, ENRS00, GKPS06, GW95, GGR98, GHS13, JSV04, JM24, KSS10, LR99, OR02, RS09a, SK96, SS06b]. Approximations [Ber69, CBHH61, CS84, Fra65, KV94b, Laa58b, Luk57, Mae60, Mae63, Min57, Spi61, Ste67, TZ25, dV59, AM07, FKV04, MP24]. Arbitrary [BIW63, DST92a, DST92b, Fra58a, Hoo66b, Nat67, AY13, BKL99, DLT07]. architecture [Val00]. Architectures [HRS88]. Arcs [Ros73b]. Area [BB95, BK81, BK82]. Area-Time [BK81, BK82]. Area-Universal [BB95]. Arguments [EFKP22, Ked79, RZR25]. Arising [Cry71]. Arithmetic [Alt88, AAB17, AM59, Ash64, BL19, BSZ65, Bre74, BFGL79, Cha90b, CGL^{+}22, CS83, EV57, Erc60, GL70, Hei71, KL73, Kan65, Mát90, MP76, Pap87, Par55, RR64, RZ65, Raz13, Rei87, SU70, Sho79, Win75, HJV01].

Arithmetical [Jef56, CM57]. Arithmetics [Wad60]. Armstrong [BDFS84]. Arrangement [Riv78, Yan85]. arrangements [ANP^{+}04]. Array [AK84, DS78, Fis65a, Fra76, Fre82a, Kau71, SJ80, KSB06]. Arrays [GE84, KA88, Ros74, Ros75, RS77, Wai67]. Arrivals [Adi73, Del70, SB82]. Arrow [Vaz12]. Art [AAM22]. Arthur [HM13]. Article [Tar17a, Tar17b, Tar17c, Tar17d, Tar17e, Tar18a, Tar18b, Tar18c, Tar18d, Tar20a, Tar21a, Tar21b, Tar21c, Tar21d, Tar22a, Tar22b, Via11b, Via12c, Via12d, Via13b, Via13d, Via14c, Via15b, Via15a, Via11a, Via12e, Via12a, Via12b, Via13c, Via13e, Via14b]. Articles [Ano15, Tar16a, Tar16b, Tar16c, Tar16d, Tar17f, Tar17g, Tar17h, Tar19a, Tar19b, Tar20c, Tar20b, Tar21e, Tar21f, Via11d, Via14a, Via14d, Via15c, WK15, Via10b, Via11c, Via13f]. ary [Gui78]. ASCII [Chu81, Sar83]. Aspects [FW74, HSS11, Hel61, KP90, KB81, Mit74, Sri65, Sta72]. Assembly [Sch72]. Assertions [Weg77]. Assessment [MH82]. Assign [GMB85]. Assignment [AP86, BMPT91, Cur63b, DS78, Kur62, Liu63, Sch72, TTK66, BP03b, HB02]. Assignments [TWS80]. Associated [DM62, ALS09, MPR98]. Association [Ano68, Ano74, San65, Sti61, Wil54]. Associative [Pap87, Sal63, Sti81]. Associative-Commutative [Sti81]. Associativity [Sla74]. Assumption [DG21]. Assumptions [GS22]. Asymmetric [CGH^{+}05, STV20, KLSS05]. asymmetry [Vöc03]. Asymptotic [BC61, FNS21, Har66, KT90, MM84, McK87, MM86, MPT96]. Asymptotically [GST16, Hor78, LT82, TRE73, CKOR14]. Asymptotics [MW84]. Asynchronous [AACH^{+}14, AFL83, ABND^{+}90, ADG15, BCP20, Bau78, Boj84, BFR^{+}22, BT85, BS89, EGM16, GH23, GGGK13, HYC16, JT75, Liu63, LM86, MMR15, TS95, ÜD90, AC08, DFSL14, HS99, MRR03]. Atomic [AAD^{+}93, AACHE15, AACHE18, FT22, GPV89, HV95, Her90, SAG94, AKMS11, FG10]. Attack [Cun85]. Attempts [BFGY19]. Attenuation [San88]. Attribute [Dil89, EF89, GVW15, Jaz81, Boy05, NV02, RS99]. Attribute-Based [GVW15]. attributes [RS99]. Auction [DHL^{+}20, JL23]. Auctions [CKS16, DV16, DRY16, DFN^{+}24, GML15, BLP09, GG13b, LOS02, LBNS09]. Augmentation [TZ25, KKPW25]. Augmented [Kau71]. Augmenting [BMPT91, GM80]. authenticated [KOY09, LLR06]. Author [Ano58b, Ano68, Ano95a, Ano97, Ano98a]. Authoritative [Kle99]. Autoepistemic [Got95b, MT91]. Automata [Aho69, ADO91, Arb61, Bar65, Bav68, BM70, Bro64, BW57b, BW57c, Bur70, Con91, CDPP23, DH94, EKS20, FMNP19, Fle62, FHO72, Fri62, Gin61b, GGH67a, GGH67b, GH70, GB69a, GB69b, GB69c, GB70, Har66, HS68, Har69, HH94, HU68a, KV94a, Kel61, KB67, KS66, Lee61, PP65, Pic67, RC69, Rit63, RS94c, SEY15, Tra66, Ukk82, Ull69, Val75, Wee65, Wee67, Win64, Yeh68, Yoe65, von66, ANTSV02, BGB12, BBB^{+}00, FPS98, KVW00, OG96, TS10]. automata-theoretic [KVW00]. Automated [BG01, BY96, GOBS69, MW86, Ove74, Sla74, Win82]. Automatic [BB77, BB63, BB64, Cha66b, CM04, DL56, Doy65, Edm69, Gia64, GKR55, Gri73, Hen56a, HW55, KU80, Koc69, Lan59, Mar61, McA93, Mel56, Mos55, Nev70, Nev74, O’C65, Oet57, Per54, Sal73, Sch61, Sco58, Sla67, Sla72b, Smi82, Swi57, YL57, Yer66, YS76, YSS78]. Automatically [Hob93]. Automating [AM20, HC88]. Automaton [Fle65, Oeh63, Wee62, BYG96]. Automorphism [Fle65, Wee65, Wee67]. Automorphisms [Bar65, GB69a, GB69b]. Availability [Her90, LM94b, dG89]. Average [AM85, ABD^{+}21, BKST78, CM89, HRST17, Mot94, RS66, Rid76, TS95, JLV00, Lue09, SSS06]. Average-Case [HRST17, Mot94, JLV00]. Averaging [CDPP14]. Avoidance [Min82]. Avoiding [CCSS14, Mor97]. Aware [DJW14, CMT20]. Axiom [Cla79, Sal66]. Axiomatic [Bea72, Bea73a, Bea73b, MH82, AT10, BT00]. Axiomatics [CK79]. Axiomatization [PT20, SU82, Sci82, Kif98]. Axiomatizations [CGH83]. Axiomatizing [Men79]. Axioms [Dix73, Wan93, DFPP02]. AYDAR [ZL58].

B [FG99, Yi12]. B-tree [FG99]. B-Trees [Yi12]. Backbone [GKL24]. Backing [BAP06]. Backlogs [Gav71]. Backoff [BFGY19, GGMM88]. Backtrack [Ber73, Fre82b, Fre85, GB65, KZ93]. Backtrack-Bounded [Fre85]. Backtrack-Free [Fre82b]. Backtracking [SD76]. Backus [Can62]. bad [KVV04, RT02]. Balance [CHT77]. Balanced [AGH00, DPRS89, LS23, Tar79]. Balancing [GJL21, OV23, Rab89, RY91, TT85, AAF^{+}97, BM96, KP02, Vöc03]. Ball [BBB^{+}21]. Balloon [GH70]. Ballot [RV78]. Banded [RMK79, Sav84]. Bandit [BEL21, JWW^{+}23, GMS10]. Bandits [BKS18, ISSS22, KSU19]. Bank [Flo64]. Bar [RBCC19]. bargaining [Vaz12]. Barrier [CRWW24, DI05, GR98, KS11]. Base [Kan65, Low68]. Based [AV89, Aki70, ABB^{+}17, Bak62, Bau85, BMR97, Bry91, CGOS13, CM74, DM87, DH86, EG95, Eld59, FL18, GVW15, Gri86, HMZ13, KLW95, KUB^{+}15, NO80, RS94c, Rob65, San65, San69, Sha67b, SV85, Smi70, TPH86, Wan80, Wes67, BG01, BP01, BEL21, CE00, CGL^{+}22, DG21, Kor08, MSU99, Mye99, NV02, Reg04, Vem10, HZ92]. Bases [Baa93, CKN18]. Basic [BCL19, Cur63a, Mat82]. Basis [Baa93, Gal82, MH89]. Batch [Del70, HB21, Wil91]. Batched [HY77]. Bateman [Tho57]. Baur [CGS15]. Bayesian [BGJ10, CKS16, Dar03, FKR19, San96, XZ25]. Be [BIM95, KT78, Yao81b, BDFG19, PW93, Was80, Wei66]. beacons [KSW09]. Becoming [Hal97]. Beek [ZY03]. been [Sho03]. before [BZ14]. Beginning [PTW88]. beginnings [Alt03]. Behavior [BC61, Lam82, TG82]. Behavioral [MN05, PS00]. Belief [CF20, CO17, DPW18a]. Bellman [FL18]. Bending [Was54]. Benefit [ST73]. Benefit-Cost [ST73]. benefits [AFH96]. Bernard [DNRS03]. Bernouilli [Jen73]. Bernoulli [Bau58a, DHKN21, Mou65b]. Bernstein [GR74]. Bernstein-Bézier [GR74]. Bessel [CU59, dWF58]. Best [BC61, CF59, DLN73, DP85, GLH57, HP59, KIM81]. Best-First [DP85]. Better [BKSW24, BK77a, JSV97, TWZ23, TZ25, YR78, GKPR07]. Better-than- [BKSW24]. Better-Than-2 [TZ25]. Between [Bea82, HU68b, HS78a, HS78b, LBB67, PS78, RV80, SDPF81, WW18, WM74, AM85, AGPV90, BPC^{+}12, BS88, BW94, KA88, PU89, SDPF87, ŠVV09, Tho07]. betweenness [DEP10]. Beyond [BZ14, BS85, CHO^{+}22, CO84, DEG^{+}21, GR98, LMW05, LW08, APR13]. Bézier [GR74]. Bi [CDOY11]. Bi-Abduction [CDOY11]. Bias [BHT18, FACP81, ACKM09]. Bibel [SW01]. Bibliographic [Sal63]. Bibliography [Hou56a]. Biclique [Lin18]. Biconnectivity [KV94b, HdLT01]. Bidding [AHC19]. bidirected [MMHW97]. Bidirectional [ON68, dCS77, de 83]. Bierstone [MC72]. Big [BK77a]. Biharmonic [Par61]. bijective [AS13]. Bilinear [Ja’80, BL04b]. bill [Hal00]. Billing [Wis54]. Bills [Sal60]. Bin [GR20, LL85, Yao80, CJK^{+}06, Sei02]. Bin-Packing [LL85]. Binary [AM78a, BC74, BK81, BK82, Bur76, Cha66a, Cha69a, CJLW22, Dev86, DH86, FM79, Fit60, GGMM88, Har74, HD64, KR89, KW89, LM94a, Lio71, MM24, MS74, MMR15, NW73, PRR74, Par55, Pen65, Pro80, RV78, Sla64, ST85, SF80, Spr92, Zer85, Zol78, AD01, Drm03, MR98, Ree03]. Binding [Dav17]. Binding-Time [Dav17]. Binomial [Wei62, You67]. Binomial-Weighted [You67]. Bipartite [BHR19, FHTZ22, IRT78, RA20]. Bird [Cha14]. Bisection [Wil78]. bisimilarity [JS08]. Bisimulation [BBK93, BIM95, vW96, BCG07, DV95, SP07]. Bisimulations [CH89]. Bistable [Pul55]. Bit [DMR08, MM24, KS11, Mye99]. bit-vector [Mye99]. Bitcoin [GKL24]. Bitmaps [Hob93]. bits [HN04]. Black [DHKN21, Kla85, KNY19]. Black-Box [KNY19, DHKN21]. blackbox [SS13]. Blank [ER64a]. Block [Cur63a, CV59, Sch72, Wat73]. Blocking [CPT94, DLT94, KR76]. Blocks [CCSS14]. Bodies [AGTG21, DFK91]. Body [CK95a]. Book [Ano60, Ham56b, Ham56c, Ham56d, Ham57a, Ham57b, Ham57c, Ham57d, Ham58a, Ham58b, Ham58c, Ham58d, Ham59a, Ham59b, Ham59c, Ham59d, Ham60]. Boolean [Sim67, AGG62, BBCP20, BP89, Bre70, BDG^{+}13, GKR16, HRST17, HP72a, Har63a, Hay76, Hir58, KV94a, KLV94, KB81, Law64, Lee02, MP65, MP66a, Nat67, Sim65, SV85, Svo75, War62, WM74, YCW99]. Boosting [FGMS05]. Both [KW66]. Bottleneck [GG78, HS86]. Bound [And65a, Bry82, Das75, ES86, FGR^{+}17, Fre81, GW85, HS21, Hel89, Iba77, KZ93, Kir81, Kla85, KS74, MST91, Mor73, Raz19, RS94b, Rub15, Rub22, Sch88, Smi84, Yao81a, CR12, JLV00, Ste03, SM02]. Boundaries [BN13]. Boundary [AT71, Ber54, BH65, CM60, CM61, Col58, CD60, Cry70, Cry71, Ehr58, Ehr59, HW67, MP59, Usm66, Was71, Yak76, Zaf69]. Boundary-Value [Was71]. Bounded [AM85, BHM11, BSV20, Con91, Coo71, DH94, DJM^{+}20, ES17, Fre85, GNP20, GX22, HV02, HH94, HU69, Iba78, Lov16, MS92, SS79, SL15, Sud75, TS74, BFU01, BL04b, DR03, DFHT05, DW99, FJWY25, MLS76, MP24]. bounded-genus [DFHT05]. Bounding [CM96, HKP84, ZS99]. Bounds [Abe80, AAHK94, AC81, AC23, AG17, AHU76, AACH^{+}14, AH92b, ABDH^{+}20, ADLS94, BBH^{+}21, BH89, BZ14, BBBL95, BN23, Car58, Cau58, CN95, CG67, Cha90a, Cha90b, CC16, CRTY23, Che19, CdDPL^{+}24, CADMD21, CHK82, CS84, CFG^{+}17, DKK24, DHK19, DR85, DCKT92, FL74, FMP^{+}15, FB72, FNS21, FG79, GR60, GH23, GLVV12, GYY80, GLNS25, Gup79, GHLL22, HHHN14, Hib61, HS78a, HS78b, JK78, Kar67, Ked79, Kis91, KMW16, Kun76, KMS^{+}25, LT82, LMNM03, LY89, LST25, LM94b, MV25, Mey85, MN24, MP75, NW73, Oma77, PKR84, Pet15, Pot17, Raz13, Rei80a, Rob74, RVW18, She14, Smi70, Ste83, SEY15, Tiw87, Tro84, Tsa74, Wil14, WC76, YY76, AS04, AC05, AK10, AAF^{+}01, Asp98, AC08, BBC^{+}01, BSSV03, BLR10, BL04b, CR12, CeLT00, CKL^{+}07, CD08]. bounds [EFR07, FG10, FLvMV05, FV02, GHS13, GHS09, HKM^{+}02, Kol04, LLW10, LNT97, Lue09, Niv10, Raz04, SS13]. Box [KNY19, DHKN21, ZS99]. BQP [RT22]. Bracketed [San72]. Branch [BDSV24, Hel89, Iba77, KZ93, KS74, Smi84]. Branch-and-Bound [Iba77, KZ93, KS74]. Branching [BIZ18, EH86, Weg88, vW96, DV95, KVW00]. branching-time [KVW00]. break [BK79]. Breaking [ADH15, BEPS16, CRWW24, KS11, DI05, DMR08]. Bridge [Ber63]. bridging [Zwi02]. Broadcast [AGPV90, BT85, FGL77, WXSD25, FKL98]. Broadcasting [CD21]. Brouwer [CD08]. Brownian [MV25]. Büchi [CH14]. Bucket [Tra63]. Buckets [BK77a, DR03]. Buddy [PS70]. Buffer [BGH^{+}95, GL71, Hof77]. Buffering [Fer60, MS59, Sch72]. buffers [BFU01]. Building [CJJK00, Cur63a]. Built [Sla72b, Wan93]. Built-in [Sla72b, Wan93]. Bulk [Adi73, Del70, HM72]. Bureau [Mey65]. Burroughs [Mie57]. Burrows [Man01]. Bursts [Das75]. Business [PM71]. Busy [BN71, Dad88, NBC72]. butterfly [BCH^{+}96]. butterfly-like [BCH^{+}96]. Buyer [BILW20]. BWT [GNP20]. BWT-Runs [GNP20]. Byzantine [BCP20, Bra87, DR85, DRS90, DW04, HPZ24, KS11, KS16, LSP82, Lam83, LR19, LLR06, MMR15, WXSD25].

C [Ano98b, OM78]. cabbage [HP99]. Cache [MT19, Rao78, FV02, SCD02]. cache-efficient [SCD02]. Caching [LV21]. Cake [HR25]. Cake-Cutting [HR25]. Calculating [BS74, CLW95, GK58, Leb56, dG89, dL89]. Calculation [AS54, Col71b, Flo60, IS56, von56]. Calculations [Cha66b, Sto64a]. Calculator [BT57, HK54, Mos54, Par55]. Calculi [Gai20, BG98, CHL96]. Calculus [ABF17, BBFD23, Bra78, Fra61, Fri63b, Klu82, Man69, MN24, Sla63, Sla70, van72, DKP07, OR00, PS00, PR99]. Call [ST80, LNT97]. Call-by-value [ST80]. Calls [AM82]. Can [EC64, FHV92, KT78, Was80, BRSU97]. Canceling [GT89]. candidates [CSL07]. Cannot [PW93, BDFG19]. Canonical [AM82, BD94, Cav70, CK76, Cur59a, Cur59b, GNP^{+}93, GH73]. Can’t [BIM95]. Capabilities [Tra63]. Capability [GGV^{+}24, WL85]. Capacitated [Gav83]. Capacities [TWS80]. Capacity [Che19, EW75, Kne93, Oma77, SM64]. Capturing [GUK^{+}16]. Card [Ber54, Tit55, Was54]. Card-Program [Tit55]. Cards [Leh57, Per54]. Care [Chu61]. Carlo [Ehr59, FKV04, HH57, Ham56a, HH63, Mos58, RTW94, TM66, TI70]. Carnap [Got95a]. Carr [GR60]. Carrier [Tob82]. Carroll [HHR97]. Carrying [Bli72]. Cartesian [Mor97]. Carvings [KV94b]. Cascade [Yoe65]. Case [Cha90a, CM89, FKS84, Fri63a, GW85, GS90, HRST17, Mot94, Ros83, Tv84, TSG85, DKP98, JLV00, LBNS09, NPRR18, SSS06]. Cash [Bui77]. Catastrophe [Nel87]. Categories [Ano68]. Categorizing [Lee61]. Catenation [DST94, KT99]. Cayley [Lar76]. Cech [BDK24]. Cell [Wol65, LMNM03]. Cell-Size [Wol65]. Cellular [Bur70, Muk68, Smi71b]. center [CGH^{+}05]. Centers [GM70]. Centext [LW93]. Centext-Free [LW93]. Central [Men65, PGL76, Pri76]. Centrality [RP89, DEP10]. century [Bro03]. Cerise [GGV^{+}24]. Certain [Boo78, Col58, Dub66, Gau61b, Hib62, Hof77, Kun76, LV57, Phi62, Ros68, Sla64, Usm66, Win75]. Certifiable [BCM^{+}21]. Certified [SFLB^{+}25]. CFI [GLNS25]. Chain [CG86, DJM17, Hof84, XB06]. Chained [Sta72]. chaining [BG98]. Chains [AAGT15, CS85, CI74b, RW23, Ste83, WSY^{+}19, AKN^{+}08, DMR08, EY09]. challenge [Hoa03]. challenges [Bro03, Cor03, Fei03]. Change [CG70a, EHR84, Wri75, Rag03]. Change-Making [CG70a, Wri75]. Changing [CK76, SR60]. Channel [BFGY19, BBBL95, FGL77, GK94, AZ05]. Channel-Access [BFGY19]. Channels [AAF^{+}94, Che19, GW85, GFL87, GS16, MS92]. Chaotic [LM86]. Character [Chu65, CC67, EC64]. Characteristic [BGKM23, DH56, Fra58b, Giv57]. Characteristics [Kro66, Low68]. Characterization [CM83, Cul79, Dem92, GS83, HS71, KS74, MST93, NSW24, RZ19, AS98, BCG07, BKL^{+}06]. Characterizations [Coo71, HU74]. Characterizing [AGHR17, HP07, Smi79]. Chart [Sco58]. Charting [GKS04]. Charts [BS72]. Chasing [AGTG21]. Chasm [GKKS14]. Cheating [Rab94]. Chebycheff [SS58]. Chebyshev [Dun74, Mou65a, She57a, Sto64b]. Check [BK95]. Checker [de 86]. Checking [BKTW22, BKV25, Cal61, CKN18, Gin67, Kob13, Mar59, Par55, SFLB^{+}25, SEY15, AS98, CGJ^{+}03, DNS05, KVW00, NR09, WB97]. Checkpoint [Gel79]. Cheeger [LGT14]. Chemical [CDRS23]. Chemistry [GV61, WJR57]. Chen [ZM22]. Chervonenkis [BEHW89]. Chess [KSU^{+}57]. chief [Jou03, Hal97]. Chinese [AB03, BGJ10, Li59]. Choice [DK65, LS24, LS23]. Choiceless [LS24]. Chow [DDFS14, Win71]. Christofides [AKS15]. Chromatic [Wan74a]. Church [DKRW15, MNO88, Ros73a, Set74, Set75, Sha88, Toy87, Yao03]. Circle [GSH89]. circles [ANP^{+}04]. Circuit [All89, BSKK^{+}16, BKK^{+}18, BEGY22, CRTY23, Cur61, Cur63a, DFHM24, GP18a, Joh93, Kle67a, Wil14, SM02]. Circuit-SAT [BSKK^{+}16]. Circuitry [ILP97]. Circuits [Chu61, CS88b, Cur59b, FU82, Gil65, GVW15, HRST17, IKL^{+}19, Lei57, LST25, LMN93, Liu63, RW92, RVW18, Rub74, AAF^{+}97, AACH12, Bra10, SS13]. Circular [dWF58, RS99]. Circularity [Dil89, Jaz81]. Circulation [Ros71]. Circulations [GT89]. Circumscription [KP90, RWM94]. clairvoyance [KP00]. Clairvoyant [IKM17]. Clamped [AC55]. Class [AM78b, Bak62, Cas75, CC95, DG84, Di 69b, DCKT92, ES86, GK88, HP66, Har64, Hir58, HD73, KS66, Mat81, Ove74, PTW88, SU82, She57a, Win65b]. Classes [ACJR21, Bar65, BCMP75, BY73, Boo78, CF12, Di 73, DSS22, ESvL80, EG82, ESY85, FMI80, Har63a, Hib61, La 64, LR72, LED76, LS86, MY16, Ros69, Rus82, SFM78, Sud75, Tor91, XDB83, Yu75, Har03]. Classical [GS57, SS17, Yao03, van72, NS06a]. Classification [BB63, BB64, Doy65, Mey65, San65, KT02]. Clauses [Fag82, Lew78, NM90]. Claw [FOS14]. Claw-Free [FOS14]. clearing [BSZ06]. Clinching [GML15]. Clique [ABD^{+}21, BBB^{+}21, DP22, Weg88]. Cliques [FGR^{+}17, MC72, FGL^{+}96]. Clock [DHSS95, LR19, ST87, DW04, LLW10]. Clocked [ILP97]. Clocks [GS21, LMS85, NT93a, HV02]. Closed [BCMP75, CM96, CLW95, GK58, KT90, MM84, McK87, MM86, RL80, RL81, Bas99]. Closeness [BFR^{+}13]. Closest [Val15, LMW02]. Closure [DSS22, JCG97, NO80, Rei80b, Rem81, DI05, TS10]. Closures [Gra55b, Kel73b, Sac85]. CLS [FGHS23]. Cluster [AM70]. Clustered [FT22, TTK66, YL77]. Clustering [AGPT16, BBG13, BBS90, CGOS13, CM96, CAFS21, CaKMTM19, GK68, Gui78, MB83, ACN08, KSS10, OR02]. clusterings [KVV04]. Clusters [Mou65b]. CNFs [FPPR22]. Co [CDPP23]. Co-lexicographically [CDPP23]. Coalesced [Vit83]. Coarse [CD94]. Coarse-Grained [CD94]. Code [AJ76, AJU77, Bea72, Bea73a, Bea73b, BG77, BL75, BS76, DR54, Ear74, GGV^{+}24, Gon81, GS16, Har76, Kor58, PGE75, SU70, BY96]. Coded [Mor68a]. Codes [AM80, BSCI^{+}23, CJLW22, Das75, DPW18b, GS57, GX22, HS21, Kle67a, KMRZS17, LH90, Mul78, Vit87, GS06, GUV09, KSY14, Reg09, Tar08, Yek08]. Coding [AYZ95, Bak56, BZ23, BKN14, BEGH17, Cor63, GKR55, Gor57, GM56, Hae16, HW55, KS63, Kor58, LZ25, Mar71, Mel56, ST73, ANTSV02, WR03]. coefficient [BL04b]. Coefficients [CD87a, McC73]. cognitive [Val00]. coherence [PD98]. Coin [BHT18, CK76, Asp98]. coin-flipping [Asp98]. Collapses [DV14]. Collapsing [Ko90]. Collecting [AGH^{+}25]. Collections [LRY80]. collision [AS04]. Color [AYZ95]. Color-Coding [AYZ95]. Colorable [KT17]. Coloring [BE11, Bar16, Blu94, GJ76a, HSS18, Hsu88, KT17, MB83, Wig83, BEG22, KMS98]. Colors [DDST16, KT17]. Colourings [Mol18]. Combination [Wes67]. Combinational [Alt88, Ric73a]. Combinations [Ehr73, MBM89, Sho84, Wei66, dWF58]. Combinatorial [AIK84, AK16, Bon70, BC73, CKS16, DEL80, DV16, DRY16, Ehr73, ET76, FMP^{+}15, Gus83, Hib62, Rei79, TNS82, WC74, ARV08, BLP09, BR96, BFG03, BM96, CR12, GG13b, IFF01, LOS02, LBNS09]. Combinatorics [BSLRZ14, Gri86]. Combined [BKK^{+}18, Wis54]. Combining [Ked79, PM06, RS83]. Commands [DS78]. Comments [BCPW74, Bui77, Wil71]. Commitments [LP15]. Commodity [Ita78, Ram80, CKL^{+}07, KRV09]. Common [AHU76, AJU77, AS97, Bro71, DST80, HM90, Hel89, Hir77, Kal88, MST91, MO68, NT93a, PS80, Lue09]. Communicating [Apt83, BZ83, BHR84b]. Communication [ABF17, AAF^{+}94, AGPV90, BKS17, BSLRZ14, BEGH17, CIV88, CYZG14, Chu81, DG16, FLMS93, FNFSZ23, GKR16, Jai15, KLPP15, Lam86a, Lov16, LT93, Mae77, Sar83, She14, Tiw87, TS95, Upf84, AAPS96, BPC^{+}12, BDHS12, FKL98, GJS12, JRS09]. Communications [MC87]. Commutative [Baa93, Kor89, Sti81]. Commutativity [Sla74]. Compact [Tho04]. Compactly [ABF^{+}02]. Comparative [BM83]. Comparing [Alt88]. Comparison [BK79, Gau59, Gau60, KS74, Mul59, PRR74, Sch72]. Comparisons [Yen72, FG05]. Compatibility [Gin61a, RG67]. Compatible [RBCC19]. CompCertTSO [SVN^{+}13]. Competitive [AGTG21, BBMN15, IKM17, LV21, BG00, CL11]. Compilation [FU82, Lib01, SK96]. Compiled [Dix73, GHG60]. Compiler [MO56, Moc56, SVN^{+}13, Sim63, Hoa03]. Compiling [GGH67b, HN84, Pet62]. Compiling-Parameter [Pet62]. Complementation [Rei84]. Complementedness [HU66]. Complements [ERR81]. Complete [Bak94, BBH^{+}87, CM87, CK79, ET76, GHKL11, GI79, Jaf90, Mic76, PS81b, Pie73, Rei86, Röc84, SU82, SG76, Sal66, Sch78, Sci82, Sla72a, SFLB^{+}25, Yen72, AAM22, AKM25, APR13, BT95, GRS00, GPP^{+}19, HHR97, SV03, TKB95]. Completely [Hab70]. Completeness [AB70, BBFD23, Bau85, BE97, CS71, Dav77, GJ78, HWW90, HL76, HR91, Kap68, LF88, SW82, VVAG97, WRC65, MS19, SW01]. Completion [GHS75, SSS06]. Complex [ABL80, BIW63, CC17, GB74, Lyn75, Pen65, WS76, Wil78, PD98]. Complexes [AKM25, GR79, BER06, DG98]. Complexities [Alt88, Elk20a, FZ77, HU68b, Iba72, MP75]. Complexity [AVV97, AP93, AHU76, All89, ABB^{+}17, ABDH^{+}20, AACHE15, AACHE18, Awe85, Bab16, BY73, BSLRZ14, BSKK^{+}16, BKK^{+}18, BP89, Blu67, BR23, BG15, BD87, Bor72b, Bor72a, BK81, BK82, Bsh93, CC17, CV17, CM85, CF12, CYZG14, CPY17, CST^{+}18, COW16, CS77, CR86b, CY95, CS99, Dil89, EG95, EF75, ESY85, FGHS23, FGHS25, FNFSZ23, FT88, Fre81, Fre82a, GJ76a, Gat83, Gen78, GL76, GS93, GW21, GHJ^{+}24, GK84, GLPR21, GHKL21, GP18a, GI81, Har68, Har69, HH71, HHPW14, HZ16, Hod70, Hun82a, Hun84, HK77, IKL^{+}19, Ja’84, Jai15, Jaz81, JS82, JM78, KB89, KKG^{+}22, KNY19, KMRZS17, KP96, KPP^{+}15, LR72, LS81b, LZ25, Lin18, LS86, LT93, Mai78, MSY81, MT85, Mar58, MM81, MHG^{+}88, MP65, MP66a, MT19, MSM85, Mor73, Mor75, MGLA00]. Complexity [OKSW94, Ove24, Pap76, Pap81, PY82, Pap84, PF79, Pip89, Raz16, Rot17, SLM19, SH88, SFM78, Sel72, SSSU83, SC85, SV85, Sud75, Sud78, TZ16, Tiw87, Tor91, TW79, TS95, Weg77, Weg88, Wil70, XDB83, Yan79, ZCLD25, AGHK09, BPC^{+}12, BG01, BR96, BK10, Bul13, BL04b, CSV09, Cha00b, CDT09, DMR08, FCFM00, FHS98, FKP13, GLS01, GKPS05, Gro07, Har03, HLP99, JLV00, JS96, KZ13, McA02, MPT96, MMHW97, Mul11, NR09, Raz03, SM02, tCL09]. compliance [LMW05]. Component [CH14, DST92a, DST92b, HW90, Sam81, CLMW11]. Components [Nat67]. Composable [Can20]. Composite [DI87, Ral59, SB73]. Composition [NSSV24, Weg69, LLR06]. Compositional [Bro97, CDOY11, VSC24]. Compound [von62]. comprehensive [JS96]. Compressed [MM24, FM05]. Compressing [FLMM09, GLNS25]. Compression [ABDH^{+}20, BZ23, BW94, GS85, Kor58, LWZ21, MY16, RPE81, SS82, Tar79, ABF^{+}02, BFG03, FGMS05, VK96]. Compromise [KG74]. Computability [AFR19, SS63, TNS82, ZW03, HS99]. Computable [Abe68, CG67, Ste83, BT95]. Computation [BKS15, Bau56, BW57a, Bro71, CM85, CDRS23, Col71a, Dal73, DVSC77, Dye68, EV62a, EV62b, Fik59, FLPS16, FJ60, Gab16, GS22, Gau61b, GKR15, GHKL11, HS55, Ja’80, Kam87, Kap68, KZ93, KMW16, LF80, Lar76, LV57, LY89, LT93, Mac55, MST93, Meg83, MY84, Mor75, Ove66, PS80, Roc94, RVW18, Smi71b, Yao82, AM07, BSSV03, DP01, GN08, Ind06, Val00]. Computation-Universal [Smi71b]. Computational [AVV97, Bor72a, Har68, HH71, HZ16, KS63, KP90, MW84, Par61, PV88, Pra60, RKN75, SK73, Sav72, Sri65, Sri76, TM73, Wil70, Fei03, HLP99, MMHW97, RS00, Bor72b]. Computationally [GS16]. Computations [Abe80, Ano57, BBE70, Bau85, BL79, BP89, Bli72, Fri62, Gen78, Giv58, GK84, HH63, Har68, Hem64, Hen56a, Hen56b, Hun84, JS82, KRR22, KMW67, KM20, MS19, MST91, ME67, Mát90, NJ57, Pyl64, Rei68, San96, Tsa74, Wei55, WJR57, BL99]. Compute [CAA74, Rei87]. Computed [KT78]. Computer [Adi69, AK70, Ano56a, Ano56b, Ano56c, Ano56d, Bau58b, BK72, Bla59, Blu58, BO56, Bra77, Bra67, BT69, BFGL79, Chu65, Cof69a, Cof69b, Den61, Den58, Den65, EV57, EV62a, EV62b, FH80, Flo60, Flo61a, Fri63a, Fri56, GP83, Gav67, Gel75, Gla57, Gol57a, Gol57b, Gol57c, Gol57d, GL70, GR74, Got54, Ham69, Hol62a, HR63b, HR63a, HY77, Isr57, KS78, Kam80, Kob74b, KSS75, KSS77, Lan59, Lat81, LS88, Lei57, LNSW59, LY61, Les58, LS71, Lie59, LR65, LKK83, LM94b, Mae77, Mam66, Mau66, MC74, MO56, Mil75, Min70, Mos58, Nel87, Neu54a, Neu54b, Neu54c, Neu54d, Neu55a, Neu55b, Neu55c, Neu55d, Par55, Per56, PPV60, Pri76, Ras70, RS66, RS67, Rob63, RY91, Row57, SM64, SL68, Sal72, SGRR75, SAKS64, TT85, Til58, Tiw87]. Computer [Wad56, Wal61, Wil62, Wun67, dG89, Bro03, Cor03, Val03]. Computer-Aided [BT69, GR74]. Computer-Feasible [Wal61]. Computers [ABV60, BC54, BR54, Coo71, CU59, Dub66, Gol82, Har64, HST76, HI59, Jac54, Lip57, Mul59, NS80, Rub55, Sch64, Sch66, Swi57, Wil68, Wir68a, Wir68b, CRS01, KP02, Lam03]. Computing [AGMT95, ASV06, Ano68, Ano74, ASW88, BNNS94, BIW63, BCL19, CKM^{+}14, Cha66a, Cha69a, Cha69b, CLY25, Cor63, CM57, DP60, DL19, DG98, DH59, Elg54, FGHS25, Fra65, Fre82a, Gal82, GF60, Gus83, HS74, HS75, JP84, Jef56, LW63a, LW63b, LS87, LM86, LM94b, OD67, PR08, Ric74, RZR25, Sho81, Wan57, Wil78, Wil54, Zad72, ALS09, BP03b, CDT09, EM07, FKP13, Hoa03]. Concave [EGGI92b, AC24, CdDPL^{+}24, VY11]. Concavity [DLT94]. Concentration [Men15]. Concept [Sti73, Svo75, WRCS67]. concepts [BGM^{+}98]. Concerning [And68a, Gli65, Iba72, Lio71, Men65, Pic67, PT57, Rob74]. Conclusions [GM56]. Concurrency [BC14, BBG89, DH94, HM85a, Her90, HH94, Klu83, Pap82, Pap85, SVN^{+}13, YDL93]. Concurrent [ACHS16, CRR18, CHL01, DNS04, LTV96, MM79, Pap79b, Pel87, SM90, VV12, AACH12, DW99, HV02]. Condensation [Mac57]. Condensers [Pul55, BKS^{+}10]. Condition [Esc60, FL74, Fre82b, Fre85, Loi69, CS99]. Conditional [AFR19, Bro82, DKK24, Ham56a, KMS^{+}25, Set78]. Conditioning [Osb60, MS19]. Conditions [AGHR17, Bon70, Bra77, Bri90, Chu61, CD60, Ern69, Gav71, Gla91, HW67, Mad66, Mit72, MRR03, SS24, Ats04]. Conductance [CGLP18]. Conference [Ano57, Giv58]. Configuration [GE84, KS78]. Configurations [Ehr73, SS13]. confinement [BN05]. Conflict [LR78]. Conflict-Free [LR78]. Conflicts [BCW85, GW85, GFL87]. Confluence [CHL96]. Confluent [Boo82, Hue80, CKL^{+}07]. Conformal [Rab66]. Congested [DP22]. Congestion [CL16, ARV08, AZ08, CN06]. Congruence [NO80]. Congruences [LV57]. Congruential [ADH63, DKRW15, HD64, Wes67]. Conjecture [CMS20, CR86a, CG93, GOS13, KV15, Kor89, KP95, KMR95, Was81, Zhu20, HN07, ZM22]. Conjunctive [AGK^{+}17, BFGP20, GKS06, GLVV12, Klu88, Mar13, ADK06, CNS07, DS12, GLS01]. Connected [AK84, DST92a, DST92b, HW90, KA88, Mae77, NS80, Sam81]. Connected-Component [DST92a, DST92b]. Connecting [LW57]. Connection [BE97, Joh93, KLM^{+}97, Kow75, NS82, GGR98, SW01, ŠVV09]. Connection-Graph [BE97]. Connections [Bib81, WL57]. Connective [Gin60]. Connectivity [BM70, ES17, KT19, LNP^{+}25, MP71b, Pot17, Ram66, Ros70b, Yan79, Yel75, ATSWZ00, BPC^{+}12, Gab06, HdLT01, Rei08]. coNP [AR05]. Conquer [DS13, ENRS00, FGK09, Rou01]. Consecutive [CCSS14, Gup79]. Consensus [BT85, CHT96, CMS89, DDS87, DLS88, FLP85, FNS21, JSS16, LR19, MMR15, NSW24, Ove24, AKMS11, Asp98, AC08, MRR03]. Consequences [LZ25]. Conservative [Nic93, KZ13]. Considerations [Cha69a, Ram66, She67]. Consistency [AK25, AGHR17, BK14a, Har60, Kra62, Mar59, SK80, SN04, ZY03, vD95, vD97, AL08]. Consistent [BDFG19, CK79, PW93, Ros71, FKL98, MP10]. Constant [AM94, BSKK^{+}16, BFGY19, BR71, BHT18, BEGH17, CDG^{+}20, CMS89, CADK^{+}24, Gal95, GR20, HMZ13, Hob89, Hof84, LP15, LMN93, RZR25, STV20, WXSD25, GMPS00, KLOW00]. Constant-factor [STV20]. Constant-Rate [BEGH17]. Constant-Round [LP15]. Constant-Service [Hof84]. Constant-Time [CMS89, Gal95]. Constants [CLW95, Dub66, Lam82, Pla04]. Constrained [FK85, GC65, PM71]. Constraint [BK14a, BBKO21, BAG14, BMR97, BMM18, BR23, CLRS16, Das75, EW75, LM94a, Mar13, vD97, Bas99, BDLW98, BK10, Bul06, Bul13, Gro07, ZY03, vD95]. Constraints [BLT93, GJ82, Gri86, IKM17, Llo82, MH89, Ram80, Shr70, CJJK00, DJKK08, FL02, JCG97, Pac10]. Construct [GGM86, Rem81, Zha21]. Constructing [HV95, JSS16, Kle67a, OG96, ST80, KLOW00]. Construction [CBHH61, CY82, CS77, Dil89, Jaz81, MF63, McC76, Sum77, Yu75, FCFM00, KSB06, Kel73b]. Constructions [GS16, Wil70, BKS^{+}10, NR04, Reg04]. Constructive [HSS11, RBCC19, MT10]. Constructs [Cla79]. Containing [Klu88]. Containment [HR77, MM81, MS04, tCL09]. Content [Eli74, Heh77, Kau71, Min70]. Content-Addressed [Kau71]. Contention [DHW97, GMPS00, AAF^{+}01, AF03, MPR98]. contention-resolution [AAF^{+}01]. Context [Aho68, BBK93, CMT20, CRGM78, Ell72, Gab16, GS63, GU66a, GU66b, GL76, Gor63, GH72, Gre65, Gre66, Gre69, Har67, HU66, Hib74, Jez19, KT69, Kor89, Mau69, MSW83, Par66, Pav72, Ros68, Ros70a, Ros67b, Sud75, Sud78, DR00, Lee02, MLS76]. Context-aware [CMT20]. Context-Free [Aho68, BBK93, CRGM78, Ell72, GS63, GU66b, GL76, Gor63, GH72, Gre65, Gre66, Gre69, Har67, HU66, KT69, Kor89, Mau69, MSW83, Par66, Pav72, Ros68, Ros70a, Ros67b, Sud75, Sud78, DR00, Lee02]. Context-Limited [Hib74]. contig [HRM02]. Continuation [ST80, Wan80]. Continuation-Based [Wan80]. Continued [Mac55, MB56, Mae60, Spi61]. Continuous [CDRS23, CMYZ12, GTWW77, GKpKL22, LZ25, Mon69, Sha66, She57a, Sig67]. Contour [Mor68b]. Contour-Line [Mor68b]. Contours [Mas56]. Contraction [CLY25, CM60, CM61, GGS89, MP59]. Contributions [Ano68, Ano74, AV82, PR99]. Control [ABV60, BB59, Chi86, CG57, FGL77, JT75, KL95, Laz84, LED76, Mos55, Pap82, Pap85, Par79, Pol64, TG82, YDL93, Par80]. Controllability [Mow70]. Controlled [Cha66b, AW08]. controlling [BAL^{+}21]. Controls [IL79]. Convergence [ADG15, Cha14, Dun65, Dun66, FG76, Hou56b, JM60, KNP^{+}24, Pag76, PKT90, Sta56, Tor64, TW79, ABDCBH97]. Convergent [Moo67, Was80]. Converges [MM24]. Conversational [IW71]. Conversion [Kan65, RS66, RS67]. Convex [AGTG21, BEL21, CK70, CD87b, CH79, DRY16, DFK91, GW23, HS90, Jia23, MSSV24, Sku01, Yao81a, AHPSV97, BS88, BV04, Cha01, Cha10, DPSV08, LZ01, Vaz12, vD95, EGGI92b]. Convexity [AEK^{+}12, ICR87]. convey [NS06a]. Convolution [Skl70, Tra69]. Cooperating [de 86]. Cooperative [Mel56, She59b]. Coordinates [ICR87]. Coordination [FKR19]. Copy [vO93]. Coq [SFLB^{+}25]. Core [ESY85, GN08]. Cores [BD87]. Corners [Laa58b]. Correct [GL70, LS95, SFLB^{+}25]. Correctable [KMRZS17]. Correcting [Das75]. Correction [And68a, Bri90, EV62a, LW75, Pap85, SDPF87, She65a, Tai79, WF74]. Corrections [MC72]. Correctness [AGK^{+}17, BGGR23, Cap02, HZ92, HLMW92, Kow79, KP80, Mat81, Nou83, RK23]. Corrector [Alo60, Cha62, CL62, CK65, GS64, Ham59e, HC63, Kar68, Kro66]. Correctors [KT67]. Correlated [FCMG22, FM79, PR08]. Correlation [Cov60, HS55, HH63]. Correlations [Val15]. Correspondence [Hun84]. Corresponds [BGP17]. Corrigenda [DST92a, MP66a]. Corrigendum [Ano79, Bor72b, BK82, FH98, Gau60, Hal99a, HS78a, Par80, RL81, Ros75, Sim67, Wak68a, Wee67, Wir68a]. coset [HMR^{+}10]. Cosine [DA68]. Cosmological [SM02]. Cost [BBS90, CYZG14, CKL^{+}25, Dav90, DGG^{+}25, EW75, EGGI92a, EGGI92b, GT88a, GT89, HMR83, IAWK83, Nic93, OV23, PGE75, Sch88, Sla64, ST73, Wil91, YZJ18, GKPR07, RS09b, CK95b]. Cost-Error [Dav90]. cost-sharing [RS09b]. Costs [AM80, CC76a, CC76b, BDHS12]. Count [AT71]. Countable [AP86]. Counter [BKK14, Dil89, GI82, SEY15]. Counter-Example [Dil89]. Counterexample [CGJ^{+}03, RWM94, SW01]. Counterexample-guided [CGJ^{+}03]. Counterexamples [Win82]. Counters [BE62, Bai66, Fit60, You58]. Counting [ACJR21, AHS94, BGL^{+}22, CC17, FGYZ21, FG79, KM20, Moi19, SV88, Tor91, Bul13, BDG^{+}13, DGP07, ŠVV09]. Coupling [Ram80]. Covariance [MV25, Pag76]. Cover [BYCHS17, BRK74, Fei98]. Coverability [KMS^{+}25]. Covering [BG77, Bre70, Fri73, GH72, Hen79, HM85b, Mic76, SK73]. Coverings [Nat67]. Covers [Mai80, Ott12]. CPU [BN79, TWS80]. Crash [BFR^{+}22, BN01, VJ00]. Crash-resilient [BFR^{+}22]. Crashes [FLMS93]. CRCW [BH89]. Creating [VVAG97]. Credit [Sal60]. Criteria [BCFC^{+}23, BT00]. Criterion [Ema63, Kna75, LY25, RK23, SS58, Wil59a]. Critical [Llo82]. Criticality [BBD^{+}15]. Critically [Par79, Par80]. Crout [GLW70]. Cryptanalysis [DKK24]. Cryptographic [BCM^{+}21, KV94a, Reg04]. cryptographically [AW08]. Cryptography [AAB17, May01, Reg09]. CS [McC03]. CSP [CC17, DJKK08, Zhu20]. CSPs [BDG^{+}13, CGSV24, KZ13, MP24, RWZ23, TZ16]. Cumulative [Wei62]. Current [CLS21]. Curve [Aki70, Bas58, VV63, Fer64]. Curve-Fitting [VV63]. Curves [DL19, GR74, Hob89, PB89, Ros73b, GK99]. Customer [Mit72, PTW88, Tre54]. Customers [AAI69, BCMP75, CM96, Kam82]. Cut [KV15, LP25, Mar76, Yan85, CK09, GW95, GHLL22, KS96, LR99, SW97]. Cuts [Joh87, Kar00]. Cutset [TSOA80]. Cutsets [CNP85, TSOA80]. Cutting [CADMD21, FPPR22, HR25]. Cycle [Cho80, Gib69, OM78, SD83]. Cycles [APY91, CM93, CGS15, GT89, Wei72, DL07]. Cyclic [Adi73, BKK84, Cho80, LS71, Wel66]. Cyclic-Queue [LS71]. Cyclically [WLS85]. Cyclicity [HT24]. cycling [VH06]. Cyclotomic [CDW21].

D [Cha10, Shi10]. D0L [CH84]. DAGs [DI05]. Danilewski [Han63]. Data [ADLM14, APR13, AF58, AS97, BDD^{+}24, BDSV24, BV84, Bla25, BJL^{+}82, Chu81, DL56, Ehr82, FP86, Fre83, GS85, Goo11, GK84, Gri86, Her90, HWB56, KU76, Kou77, KUB^{+}15, Kow79, LMV16, LZ25, LED76, LRY80, Lom62, Low68, McG59, Mor68b, Oet57, Opp80, Orm61, RG67, RPE81, Ros79a, Ros72, Ros79b, Sar83, Sch64, SN85, SS82, Tre54, ÜD90, VV63, Wal61, WS76, WL85, ZL58, AFN04, AZ05, AL05, AL08, AACH12, BKL^{+}06, BT07, BMSS09, Cha10, FKK11, FG99, FMW10, Fre99, GE01, GN08, GS10, GHS09, Ind06, PMH09, VK96]. Data-Independent [BDSV24]. Data-Oblivious [Goo11]. Database [AV89, BFMY83, BK86, BK87, CM87, CH91a, Fag82, Fag83, FH80, GMSV93, Gal82, GS83, GPV89, Had88, HY84, Jac82, Klu83, Kor83, Lie82, Mai80, MW84, Pap79b, Pap82, Pap85, RT90, Sac85, SDPF81, SDPF87, Sag91, Set82, SK80, VVAG97, Via87, Wil91, Yan82, YDL93, BLR13, MS96]. Databases [Abi85, BNNS94, Dea89, ES76, GL22, HN84, IL84, Lip81, MH89, Men79, Rei79, Rei80b, Rei86, SU82, TSG85, Bas99, GE01, Kif98, RSG05]. Dataflow [Deb92]. Datalog [Ros94, GK04, HMSS01, KNP^{+}24, GKpKL22, KKG^{+}22]. Davenport [Niv10, Pet15]. Davidon [Ste67]. Davis [NOT06]. Dead [Kou77]. Deadlines [GJ76b, Mar82, PTW88]. Deadlock [AH92a, CR95, DG84, IAWK83, Kam80, Min82, JSMM03]. Deadlock-Free [DG84, IAWK83, JSMM03]. Deadlock-Freedom [Kam80]. Dealer [Rab94]. Dealing [FCB08]. Debreu [Vaz12]. deBruijn [CDMP92]. decade [Via10a]. Decentralized [BFR^{+}22, Jaf83]. Dechter [ZY03]. Decidability [BBK93, BBFD23, BGG17, FL88, GHKL21, Hun82b, Hun84, OPSPW19, SJ80, Ven87, Kif98]. Decidable [AMP73, CH84, ER80a, GF80, Gur85, HU68a, KS66, SMK18, GG13a]. Deciding [CNS07, ES17, FG01, GKS17, IM83, Jez19, LS77, LB79, Sho81, Sho84, Van07]. Decimal [ADH63, Mos54]. Decimation [CO17]. Decision [BH89, BLQT22, Chu65, CC67, CH90, Di 69b, Di 73, EY15, Fri62, Fri63a, Fri63b, Iba78, Joy76, KLM19, LW93, LWZ21, MT85, Mat81, MSM85, Nau83, NO80, Pag74, RS66, RS67, Sho79, Weg88, BSSV03, BT00, DFPP02, MGLA00]. Decisions [BK77b, PM71, Sla64, KP02]. declarative [ANV13]. declustering [CC04]. decodable [Yek08]. Decoding [GX22, KSY14]. Decomposable [Dar01, Tal69, FG01]. Decomposing [GG13b, KLSS05]. Decomposition [BE97, CK95a, CH14, CDMP92, CPT94, CS84, Cur63a, Cur63b, FOS14, Fre91, GF67, Har74, MS89, VBNB17, WZC95, GR98]. Decompositions [BC69, CPSZ21, GLPR21, Yoe65, DL07, FFG02, GMS09, Kol04]. Decoupling [MS18]. Decremental [HKN18]. Dedicated [Mul11]. Deduction [AM90, AJM80, MW86]. Deductive [BNNS94, Coo64, QH68, Kif98]. Deep [DFN^{+}24]. Default [Got95b, TK91, FH01]. Defects [GE84]. defender [JS08]. Definability [GR63a, Ros68, Gro12]. Definable [BLSS03, Ros67a]. Defined [CF59, Opp80, Tor91]. definedness [Van07]. Defining [HHP93, VV12]. Definite [Di 69b, PP65, RC69, dS58, AT11]. Definition [LZ25, San88, GL70]. Definitions [AM82, Kra62, MH82]. deformations [BLR10]. degeneracy [AC05]. Degenerate [BGL^{+}22]. Degenerates [GP18b]. Degradable [LS87]. Degree [BSV20, Hoo66b, MN24, SL15, ACKM09, SSZ98]. Degrees [Cas71, CS68, Fag83, Hom87]. Delaunay [BM11, Rub15]. Delay [Haj83, OM78, OR90, Tob82, DGS01, GMPS00]. delaying [CW96]. delays [LMNM03]. Delegate [KRR22]. Delegating [GKR15]. Deleting [LR81]. Deletion [Che19, ES81, YSS78]. Deletion-type [Che19]. Deletions [HS21]. Demand [Gav71, HY75, IK74, Mit72]. Demand-Interrupting [HY75].

Demise [ZM22]. Demodulation [WRCS67]. Demodulators [Bla55]. Demonstration [BC54]. Denotation [BG70]. Denotational [LP14]. Denotations [Gab16]. Dense [ANTSV02, CD87a, DKK24, GKS17, MSW82, SSV22, JM24]. Density [BD87, Gau64, LO85]. Dependencies [BV84, BDFS84, CKV90, Fag82, Hon82, MSY81, Men79, SU82, Sag80, SDPF81, SW82, SDPF87, Sci82, Via87]. Dependency [ADS83, Cur62, Gal82, GZ82, GMV86, GJ82, Hul84]. Dependent [Del70, GKPS06, MM86, OR90, Rid76, Tow80, AZ05, WSY^{+}19]. depending [CS99]. Depth [BN23, CCL23, GKKS14, HRST17, LST25, LMN93, Nau83, RW92, Wil84, Hås21, SS13, SPK13]. depth- [SS13]. Depth-First [Wil84]. Dequantizing [CGL^{+}22]. deques [KT99]. Derandomization [CRTY23, ACR98]. Derivability [Coo66]. Derivation [Flo64, Hog81, MS74, Yel75]. Derivations [HZ92]. Derivatives [Brz64, ST74, She14, Ste67]. Descent [AP25, FGHS23, Zol78]. Describing [Cas71]. Description [GHKL21, Mel56, FMW10, MR10b]. Descriptive [BR23, Flo61b]. descriptor [FPS98]. descriptor-vector [FPS98]. Design [But78, Den65, DHL^{+}20, DHKN21, EKLS20, Fra61, Gil63, GR74, GM61, Huf57, IAWK83, IW71, KS78, LS11, LS88, MP72, Meg83, Orm61, OF61, Rou24, TS81, Vit87, XZ25, vdSS89, GKPR07]. Designations [WL57]. Designed [Cof68]. Designing [BK95, Lei57, JSMM03, LR99]. Desirability [BFMY83]. Desk [Par55]. Desk-Calculator [Par55]. Detect [Leh74]. Detecting [CSSS20, CM93, Com68, FGR^{+}17]. Detection [Chu81, Gor63, Gri73, HPZ24, Sar83, Wai67, Yak76, FG10]. Detector [CHT96]. detectors [CT96]. Determinacy [BGG17]. Determinant [HS75, TC21, Raz09]. Determination [Bro82, CF59, Dub66, Hei71, She57b]. Determining [AMS05, CH79, Mar71, Pra60, Was54]. Deterministic [BE11, Bar16, BB95, BSV20, CJLW22, Col71a, Elk20a, GHL^{+}87, GS90, Hae15, HH74, KT19, LP25, NB96, NV95, SMK18, Ste03, Sud78, Ukk82, Upf92, Val75, DR00, HdLT01, OG96]. Deterministically [GW85]. Developing [BD77]. Development [Li59, LR77]. Deviates [Mul59]. Device [Bau58b, BCM^{+}21, GS73, Men65, TWS80]. Devices [CCF88, CC76a, CC76b, Fra69, GS65, GM73, GM75, HS55, HY75, MS16a]. DFA [PW93]. Diagonal [Shr70, von62]. Diagonalization [GH59, PT57]. Diagonally [BS74]. Diagram [BGK^{+}22, Lom73]. Diagrams [AEK^{+}12, Koc69, Rey77, Lee80]. Diameter [CGS15, DLW25, YZJ18]. Dichotomy [Zhu20, Bul06, DS12]. Diclique [Har74]. Dictionary [Fre83, Sch63]. Difference [Con57, DLN73, Ehr59, Esc60, Laa58a, Par61, SY80, Set82, Ste67, Wad54, Zaf69, WR03]. Differences [BH65, Cry70, Cry71, Fis56]. Different [BM83, BCMP75, Gin61b, Kam82, MT19]. Differentiable [DFN^{+}24]. Differential [ABV60, BGP17, Bra67, BKT67, Bui79, But65, CBHH61, Cas75, Cas76, Cha62, CM60, CR56, Cra54, Cra57, Cry71, DG71, Fis56, Fro61, Gla65, Gra55b, Ham59e, HKM^{+}22, HW67, IOT19, JM60, Kar68, Kro67, Kro73, Les58, Lin69, MR59, MR60, MR62, MP59, Mor62, Pic64a, PT20, Rah69, Rah70, Usm66, Wad54, Wil91, You56, Zaf70, ABF^{+}02, Dar03]. Differential-algebraic [IOT19]. Differentially [CCMS22]. Differentiation [Blu66]. Diffie [DG21]. Diffusion [Cra55, Kob74a, Kob74b, Moc71, Sta56]. Digit [Fra76, Kan65, Mat82]. Digital [AT71, Alt62, ABV60, BC54, BW57a, Baz57, Blu59, BR54, CU59, CM57, Fra61, Gla57, HI59, HL66, Jef56, Lan59, Les58, MB56, Mas56, McL65, Mul59, Neg66, Neu54a, Neu54b, Neu54c, Neu54d, Neu55a, Neu55b, Neu55c, Neu55d, Par55, RP66, Ros70b, Ros73b, Rub55, SM64, Sch64, Sel72, SAKS64, SR71, SR60, Sti65, Swi57, Wad56, Wil62, Wun67, Ano56a, Ano56b, Ano56c, Ano56d, Gol57a, Gol57b, Gol57c, Gol57d]. Digitization [Mon70]. Digitized [Hue71, Mon69]. Digits [BFGL79, FG76]. Digraph [BG77, Hsu75, MT69]. digraphs [Tho04]. Dijkstra [Joh73]. Dimension [ADF^{+}16, AM85, AM94, BEHW89, Cla95, KPS19, Meg84, MP71a, ZTX21, BC05]. Dimensional [BEF^{+}21, Cal61, Com68, Fis65a, Mey84, Mey88, Ros83, YUKY17, BKL99, CGT10, GW21, Lee80]. Dimensionality [Ked79]. Dimensions [CD87b, Kac79, Mor59, RS94b, RKN75, AHPSV97, ABDCBH97, AMN^{+}98, ES09, Kol04, KSS10]. Diophantine [GI82, Mey85]. Direct [CG67, EC68, Hoc65, HJ61, Jai15, Low68, Mau69, Sto64b, Toy87, Wee65, Wee67, Wil61, Zol78, TKB95]. Directed [Ber73, CF68, Gav83, KBI^{+}17, KKPW25, KP80, LS68, OV23, Pot17, SD76, Wei72, AZ08, BK09, CLL^{+}08, CK09, DGP07, KLSS05]. Direction [CD60, DG62]. Directional [She14]. Directories [BDS87]. Directory [CCF88]. Dirichlet [Laa58b, You55]. Dis [EKLS20]. Disaggregation [CS85]. Disaggregations [Abi85]. Discipline [Gof71, MC74, TPH86, Noe79]. Disciplines [KK74, KM72a]. Discount [HMZ13]. Discovering [DSS22]. Discovery [EG82, GS10]. Discrepancy [Tez87, CC04]. Discrete [Abd71, BFR^{+}13, BMM18, Bro65, DS13, FM79, Fox70, Laa58b, Nic93, ÖWA91, Rot70, Sha66, Sur87, ÜD90, VV63, HN04]. Discretization [BDSV24]. Discrimination [LBB67]. discriminative [BB10]. Diseconomies [MS18]. Disequations [BB94]. Dishonest [WXSD25]. Disjoint [BG77, CL16, PS78, AZ06, HW04]. disjunctive [CJJK00]. Disk [ADW68, BBB^{+}21, CCF84, CCF88, Fra69, GM73, GM75, Wil77, AGL00, Kor08]. disk-based [Kor08]. Disks [NV95, One75]. Dispersal [Rab89]. dispersers [BKS^{+}10, SSZ98]. Display [CvH65]. Dissection [DGG^{+}25, AY13]. Dissemination [Sal71a]. Dissolution [MR93]. Distance [BEG^{+}21, CDG^{+}20, CGL^{+}23, CM74, GYY80, HPR15, Mon68, SD76, WW73, Yen72, BOG02, OR07, TZ05]. Distances [CADK^{+}24, Tho04]. Distinct [Rus82, dV71]. distinctness [AS04]. Distinguishes [Gin58]. Distortion [CRWW24, FHS13, OR07]. Distortions [Har57]. Distributed [Abe80, AFL83, AS97, BYCHS17, BE11, BEPS16, BEG22, BZ14, CRR18, CH89, CPSZ21, DM87, DDS87, Elk20b, Elk20a, FHV92, FLP85, GMB85, HM90, HSS18, JP84, LPSP15, LT93, NT93a, NSW24, PKR84, RY91, SNPT13, TT85, Tiw87, TS95, UW87, Asp98, CT96, CMYZ12, FKP13, FGY00, KP02, MRR03]. Distributing [MS16a]. Distribution [BAL^{+}21, BBS90, BT69, Cho80, FG76, FW74, GM79, Gil63, Gus83, HB21, Kne93, McK87, PK87, Pri76, SM81, She57b, ST72, VH06, ZZC^{+}25, BGM^{+}98]. Distribution-Free [FW74, BAL^{+}21, BGM^{+}98]. Distributions [BFR^{+}13, BKK84, CMT70, GKpKL22, Kob74a, Kob74b, MY84, RT88, SB82, Sev74, Tob82, dL89, ACKM09, Ind06]. Divergence [AH92a]. Diversity [RS94c]. Diversity-Based [RS94c]. Divide [DS13, ENRS00, Rou01]. Divide-and-conquer [ENRS00, Rou01]. Division [GMPZ17, Kam87]. Divisions [Ked79]. Divisor [Kam87, MST91]. Divisors [Bro71, Kal88]. Do [Rei87]. Dobrushin [LY25]. Document [BK77b, BB63, BB64, CJLW22, CM74, FKRV15, Sal63, Sal73, San65, NV02]. Documentation [Ano68]. documents [AW08]. Dodgson [HHR97]. Does [LaP93]. Domain [Dem92, GLJT23, Laa58b, Rei80b, Rey77, Sti65, BDG^{+}13]. Domain-Independent [Dem92]. Domain-theoretic [GLJT23]. Domains [LP14]. Dominance [Iba77]. Dominant [BS74]. Dominantly [Kon24]. Dominating [SLM19, AFN04]. Domination [Sla76]. Dominoes [Har86]. Don’t [Chu61]. Double [Ber63, Mac57]. Double-Dummy [Ber63]. Doubling [CAFS21, CGT10]. Doubly [Sta72]. Down [CF68, Col87, SMK18, TRA71, UV88, War57]. Down-Hill [War57]. Downhill [Cal59]. dozen [Gra03]. DPDAs [Oya87]. DPLL [NOT06]. Draw [Wei66]. Driven [Hof84]. Driving [Ram86]. Drum [CC76a, CC76b, Cof69a, Cof69b, FB75, HK54, Knu61, Tre54]. Drum-Like [CC76a, CC76b]. DTDs [BFG08, FL02]. Dual [AK16, BBN12, Blu58, FK85, Her90, HS87, DPSV08, JV01, JMM^{+}03]. duality [LPW02]. Due [EV57, Erc60, Mar82, SC80]. Dummy [Ber63]. duration [AHC19, HB02]. duty [LW08]. Duval [HN07]. Dwork [DNRS03]. Dynamic [AAJS17, Adi71a, AKMS11, AT07, BCW85, BN13, Bar16, BHK59, Bel60, Bel62, BN79, BKSW24, Bon70, Cha01, CH14, DFGN24, DHSS95, EGGI92a, EGGI92b, FN93, FMS97, Glu63, GMM81, GR17, Hel89, KM20, Lam82, MW78, MT93, ÖWA91, Pel87, Rob71, Rob74, RT90, SAKS64, Via87, Vit87, WL85, Yi12, YZJ18, vO93, BFU01, Cha10, DI04, DI05, EGIN97, GIS99, HK99, HdLT01, Mye99]. Dynamics [AAGT15, FCMG22, LY25]. DynFO [DKM^{+}18]. DYSEAC [Lei54]. Dyson [MV25].

E. [SW01]. Early [DRS90]. EASIAC [Per56]. Easy [Hae16, LR19, SV88]. Eavesdropping [FGY00]. Economical [McC76]. Economics [DFN^{+}24]. economies [EKLS20]. Economization [Ral63]. Economy [Log78]. Eder [SW01]. Edge [BP23, CL16, CNP85, ES81, FHTZ22, HW04, KT19, OR90, Pap76, Wag76, AZ06, HdLT01]. Edge-Deletion [ES81]. Edge-Disjoint [CL16, HW04, AZ06]. Edge-Faulty [BP23]. Edge-Length [OR90]. Edge-Sparse [Wag76]. Edge-Weighted [FHTZ22]. Edges [Hue71, Hue73, Hue74, MS16b]. Edit [BEG^{+}21, CDG^{+}20, CJLW22, OR07]. Editing [WC76]. Editor [GG61, Goe61, Lei91, Liv60, Nag61, WSD59, Ano56e, Erc60, Hal97]. Editor-in-Chief [Hal97]. Editorial [Hal99a, Hal99b, Hal00, Rag03, Sal71a, Sal72, Via10a, Via13a]. editors [Jou03]. editors-in-chief [Jou03]. Edmonds [Gab76, Zad72]. Edmonds-Karp [Zad72]. Effect [Att57, EW75, Pri76, Smi79, Svo75, Yan79]. Effective [Blu71, CGH83, Har86, YS76, YSS78]. Effectiveness [YL77, ORSS12]. Effects [PM71, RKN75]. Efficiency [AFL83, BN79, BBEG18, CN95, CB72, Dun74, EK72, EKLS20, GL85, HH57, HC63, Joa59, Lar76, Mus78, Pag70, PRR74, PU89, Sha70a, Tar75, Ull72, WRC65, Yer66, Zad72, Fre99]. Efficient [ADF^{+}16, ABNC^{+}96, ADO91, ACJR21, BCP20, BRK74, BCS92, BBH^{+}87, BF01, BKK14, CH14, CG86, CPT94, CG70b, Deb92, ES76, Eli74, ERS24, FK99, FHS13, FPS98, FLPS16, FLS18, Gab76, GGS89, GP83, GN08, HW90, HSTV14, HT74, ILG87, JL23, JM24, Joh77, KU88, KOY09, Kea98, KMP^{+}12, LW93, MSSV24, PGE75, PS80, Rab89, Reg25, Ric73b, Ric73a, Roc94, Sla64, Sto73, UV88, Upf84, WH73, BGM99, CBDF^{+}99, DHL^{+}20, DW99, Hoc01, LOS02, NR04, SCD02]. Efficiently [ABL17, Dea89, Got54]. Effort [GM79, She59b]. EFX [CGM24]. Egd [GW90]. Eigenvalue [BJ66, BLR10, Joh61, MV25]. Eigenvalue-Gap [MV25]. Eigenvalues [Cor63, EV62a, EV62b, Kah95, La 64, von56]. Eigenvectors [BS74, Cor63, CS84, EV62a, EV62b, La 64, von56]. Electing [FL87]. Election [CD21, GS21, KPP^{+}15]. elections [CSL07, HHR97]. Electrical [WL57]. Electron [LWS55]. Electronic [Baz57, EV57, Fri56, Leb56, McG59, PPV60, Wad54, Wad56, Wis54]. Electrostatic [Pel98]. element [AS04, Bul06]. Elementary [Bre76, CLL^{+}21, Wan74b, Wil59b]. Elements [Pul55, Ric73b, Ric73a, She65a, She65b, RS00]. Eliminating [Ram88]. Elimination [BK86, BIW63, FLSY74, Lov68, Lov69b, Lov69a, Mar76, Ple74, Ske79, BR96, Bas99]. Elliott [YL57]. Elliptic [Cry67, Cry71, DH59, JM60, LW63a, LW63b, GK99]. Elusive [SAG94]. Embeddability [AKM25, FT22, KV15, dMRST20]. Embedded [CM87, CADMD21, SW82]. Embedded-Complete [CM87]. Embedding [CFG^{+}17, DEL80, FL18, KSW09]. Embeddings [FHS13, HMR83, ARV09, CGT10, Ind06, OR07]. Empirical [GSK59, LBNS09, O’C65, PM71]. Emulations [KLM^{+}97, BCH^{+}96]. Encoding [Heh77, KR92, Ros79b, Sch63, ABF^{+}02]. Encryption [BV18, DG21, GVW15, Unr15]. End [BGP84, CH14]. End-Component [CH14]. End-to-End [BGP84]. Ended [Bai66]. Energy [EKLS20, Kis91, BKP07]. engine [BYG08]. Engineering [Baz57, BFM^{+}19, WJR57]. English [BF61, KS63, Koc69]. English-Like [Koc69]. Enhanced [GS21]. Enough [KPW18]. ensures [BN05]. Entrance [Bab75]. Entries [HS75, JSV04]. entropies [SU05]. Entropy [Gre73, VV17, CKOR14, JRS09]. Entry [Mac57, RS66, RS67]. Enumerable [Cul79, ERR81, ER80b]. Enumerate [TSOA80]. Enumerating [Kle60, SF80]. Enumeration [ACJR21, BB85, CPSZ21, Hay76, Mar76, SSV22]. Enumerations [You69]. Environment [AGOW66, ABND^{+}90, CE72, HM90, LL73, SH22, DKP98, Kle99]. Environments [CHJS15, GR17, PM06]. Envy [HR25]. Envy-Free [HR25]. Epistemic [EFW10]. Epsilon [Rub22]. Epsilon-nets [Rub22]. EPTAS [BBB^{+}21]. Equal [Gli65]. Equality [CS71, DH86, ER80b, Gin67, RWM94, Rei80b, Set78, Sla72b, Wan93]. Equality-Based [DH86]. Equality-Oriented [Wan93]. Equation [BJ66, BH65, CM60, Con57, Cra54, Cra55, Cra57, Cry67, Dou59, Esc60, Fis58, Her74, Hoc65, Laa58a, MP59, Mor62, Par61, PT20, Sig67, Sim65, Sim67, Sta56, SW68, You56, Hal07]. Equational [BD94, BT95, GNRS92, PS81b, Ver95]. Equations [Abd71, ABV60, And65b, Boj84, BGP17, Bra67, BKT67, Bui79, BB94, But65, CD87a, CBHH61, Cas75, Cas76, CG67, Cha62, CR56, Cry71, DLN73, DG71, Fis56, FK15, Fra58b, FR63, Fri73, Fro61, Gat83, Gau59, Gau60, Gla65, GLH57, Gra55b, GI82, GLW70, Ham59e, Has70, Hen70, HW67, IOT19, Jez16, JM60, Kac79, KK70, Kar68, Kro67, Kro73, Lan59, Leh61, Lin69, LKK83, McC73, Mey85, MR59, MR60, MR62, Moo67, MO68, Mou65a, Osb61, Phi62, Pic64a, Rah69, Rah70, RMK79, Ros67b, SK77, Sto73, Tho57, TW79, Two63, Usm66, Was81, Zaf70, Zel68, Zoh74, EY09, KMP00, Pla04]. Equidistributed [Hab70]. Equilibration [Bun71]. Equilibria [Bab16, IKM17, JWW^{+}23, KL95, CDT09, PR08]. Equilibrium [FCMG22, DPSV08, VY11, Kob74a]. Equipment [Jac54, Wis54]. Equiprobable [PGE75]. Equivalence [AV88, BBK93, Bar65, BHR84a, CH84, DLT94, ER80a, GW90, Gri68, GI81, HR77, IM83, IL83, Klu82, Lie82, Mar71, Oya87, RV80, SDPF81, SDPF87, SMK18, Tho07, Tsi70, Ukk82, Lib01, MS04, PS00]. Equivalences [SY80, WW18, CNS07]. Equivalent [ABL^{+}22, Hsu75, MT69, You77]. ERA [Bau56]. ERA-1103 [Bau56]. Erasure [SFLB^{+}25]. Ergodic [OD67]. ERMETH [Rut55]. Errata [Bea73a, Bea73b, Bui77, CC76a, Cof69b, GM75, GB69b, GB70, KSS77, Lov69b, Set75]. Erratum [AACHE18, HR63a, Hue74, LW63a, SW01, ZY03]. Error [And65a, Att57, BR71, BCLM56, Car58, Cas76, Cau58, CG67, CC67, Chu81, CADK^{+}24, CKL^{+}23, Dav90, Dou59, DW64, GR60, Hou58b, KL70, Laa58b, Sar83, Smi70, Ste83, Tsa74, Van85, Wee60, Wil61, Win64, Zaf70, Cha00a, CS99, MR10a, Pel98]. Errors [CM96, CJLW22, EV57, Erc60, Gra55b, KL73, LPR13, MB56, BKL99, Reg09]. Essays [Bur70]. Establishing [AB70, Ked79]. Estimate [BR71, Rob71]. Estimates [BFGL79, Cas76, FG76, Har66, Zaf70, CS99]. Estimating [FM79, GFL87, SGP11, VV17, AT11]. Estimation [Gau64, Igl76, Mos58, SW68, DLT07]. Estimators [VV17]. ETH [GLR^{+}25]. Euclid [Bro71, BT71]. Euclidean [Aro98, CRS01, ES15, HPR15, Mon68, Wan23, YUKY17]. Evaluating [Ata94, CS83, Wag97, CL11]. Evaluation [Ash64, BGKM23, BGG^{+}24, BN79, Boo65, Bre74, Bre76, Fro61, HK77, Isr57, Joh61, Kro64, Kun76, MB56, ME67, MP76, NT93b, Rei86, Sag91, SL68, ST74, Tit55, Tra69, Was81, Win75, dWF58, BP11, CW96, FFG02, GK96, GKPS05, KZ98]. Evaluator [Mie57]. Even [BBE^{+}21, CFGS22, JM77, RMK79]. Event [FM79, Nic93, ÖWA91, Sur87]. Events [BZ14, Brz67, BC69, CEW58, Mey69, Muk68, PP65, Sal66]. Every [Pet15]. Everywhere [GB74]. Evidence [PM71]. evolution [ALS09, FK99]. Evolvability [Val09]. Evolving [AGPT16]. Exact [CGL^{+}23, CG86, Elk20b, FLPS16, FGLS19, HHR97, HS75, HS76, JS82, KS75, McC73, SS24, Van78, FGK09]. Examination [Mou65b]. Example [Bau85, Dal73, Dil89, McK78]. Examples [CS88a, GJ78, PV88, Sum77, Urq87]. excellence [Rag03]. Exchange [ADLM14, CJLW22, DR85, HI59, AL08, APR13, FKK11, GN08, KOY09]. Excluded [FLS18, Gro12]. Exclusion [BJL^{+}82, DM25, Lam86a, Lam86b, Sta82]. Exclusive [KS83]. Exclusiveness [Svo75]. Execution [AC81, BL90, Leu93, Rem81, Str83]. Execution/Sleep [Leu93]. Exercising [Ric73b, Ric73a]. Exhibit [Gui78]. Existence [BD87, Bor72b, Bor72a, KL95, Wan74b, CKL^{+}07, PKT90]. Existential [EGG00, GKS04, GHKL21, Ven87]. Exists [CGM24]. Expander [ARV09, CPSZ21, Gab06]. expanders [GUV09]. Expanding [MS16a]. Expansion [AM82, HS56, Kah95, WH73, BDHS12]. Expansions [KT90, MM84, McK87, MM86]. expectation [HP07]. Expected [Bra87, BDF81, CC76a, CC76b, CR95, Gon81, KKMO18, KS16, LZBY25, MMR15, WXSD25, CFM10, GMPS00]. Experience [SK73]. Experiment [Gin58, Mar61]. Experimental [MF63, PM71]. Experiments [AF58, BF67, BB64, Coh55, Hib61, KSU^{+}57, Oli67, PT57, Rab66, Sim63, SB68, SD69, You56]. expert [CBFH^{+}97, PM06]. Experts [KSU19]. explanations [EM07]. Explicit [Cra55, FHS13, SSZ98, CHL96]. Exploiting [CD21]. Exploration [DHK19]. Exponent [AB17]. Exponential [BFGY19, Blu59, BKK84, BDF81, BCL19, CRTY23, COW23, Cho80, Dil89, FMP^{+}15, GKR16, GGMM88, Jaz81, KW66, LNT97, Luk57, Mac55, RS94b, Rot17, SD83, Van78, AT07, OPSPW19, PPSZ05]. Exponential-time [CRTY23, PPSZ05]. Exponentially [BHH23, DP22]. Exponentiating [FT88]. Exposing [RW88]. expressibility [BDG^{+}13, KMP00]. Expression [AJ76, BS72, Mye92, GK96]. Expressions [AJU77, AGG62, Bea72, Bea73a, Bea73b, Bre74, Bre70, Brz64, CS83, FU82, Gin67, Hen79, Klu83, Kun76, Mar71, MP76, OF61, PS80, RR64, SY80, SU70, Set78, Wag97, Win75, WH73, ACM02, BCG07, BYG96]. Expressive [CKP17, KKG^{+}22, BDLW98, GK04, tCL09]. Expressiveness [FLO83, NV02, CSV09]. Extended [CH91b, Cur63a, HN07]. Extendible [Ros74, Ros75, RS77]. extensible [SS06a]. Extension [Aho68, And66, Kei56, LW75, MC87, Rot17, Shr70]. extensions [HMSS01]. Extensive [FCMG22]. Extensive-form [FCMG22]. extent [AHPV04]. External [Goe61, GL88a, LRY80, RA76, FG99, GHS09]. Extra [Hor78]. Extracting [Edm69]. Extraction [FKRV15, CM04, GK04]. Extractors [Tre01, BKS^{+}10, GUV09, SU05]. Extrapolation [TM66]. Extrema [DS91].

Extremal [BLT93]. Extremely [SDS^{+}18]. exudation [CDE^{+}00].

F [Zoh69, LTT22]. Face [FLMS93]. Facility [ADW68, CPT94, GSV84, JV01, JMM^{+}03]. Fact [Coo64]. Factor [CDG^{+}20, CF20, CADK^{+}24, Flo64, HMZ13, MS21, Sti61, JMM^{+}03, STV20]. factor-revealing [JMM^{+}03]. Factored [HHHN14]. Factorial [Oli67]. Factories [DHKN21]. Factoring [DSS22, Reg25]. Factorization [BSV20, Coh55, CEG86, CR86b, CD94, Joh93, Mus75, Pag76]. Fails [CO17, KMR95]. Failure [CHT96, CMS89, HL87, CT96, FG10]. Failures [LF88, BN01, VJ00]. Fair [DM25, GM92, KPW18]. Fairest [GMPZ17]. Fairness [DHK^{+}17, GHKL11, Har86, VV12]. False [CMS20]. Families [FLPS16, GR62, GZ82, Hul84, MSW82]. Family [GJ82, HZ92, Ral59]. Fan [HKP84]. Fan-out [HKP84]. Fanout [Ben80, But78, Hay75, Hay76]. Fanout-Free [Ben80, But78, Hay76]. Fast [AM07, AKPS24, ALS94, BYG96, BGKM23, BGG^{+}24, BKN14, Bre76, BK78, BT79, CKL13, CKN18, DJSW19, Fei84, FGYZ21, FKV04, GM79, GS21, GW76, Hae15, HW90, Hoc65, IK75, Kam87, KL70, KW85, KT78, LH90, Lee02, Mey88, Mor73, Mul91, NO80, Pap79a, Pea68, Raz19, RZR25, SD76, Sch80, Tar81a, Tro84, XB06, BDHS12, Mye99]. Faster [AMOT90, AC24, BHH23, Col87, DP22, DGM94, GT91, HVL17, KS20, NSSV24, OV20]. fat [ES09]. Fault [CYZG14, DHSS95, DFSL14, DI87, JCT98, LS88, Rab89, BN01, VJ00]. fault-tolerance [BN01]. Fault-Tolerant [DHSS95, DI87, LS88, DFSL14, JCT98]. Faults [DLP^{+}86, LMS85, PSL80, DW04]. Faulty [Bar16, BP23, FLP85, AGMT95]. Fd [GW90]. Feasible [TC21, Wal61]. Feature [MF63]. Features [Swi57]. Fee [Rou24]. Feedback [AF77, CY82, CK68, CH86, CH88, Del70, FACP81, HM72, JWW^{+}23, LP73, MC74, Mow67, Sim84, SAKS64, WLS85, YLC76, CLL^{+}08]. Ferroelectric [Pul55]. ferromagnetic [GJ12]. FERUT [CG57, HW55]. few [CSL07, GHS09]. Fibonacci [FT87]. Field [Abe68, WJR57]. Fields [BGG^{+}24, CK95a, FT88, HVL17, HvdH22, Ja’80, KB89, AY13, Bas99, Hoc01, KT02, Pel98]. FIFO [MS92]. Fifth [Gru70, MR62]. Fifth-Order [MR62]. File [Blu58, Gau61a, GS69, Low68, TWS80, Win65b]. Files [BBH^{+}87, Eli74, GS69, Gli65, HY77, MY79, YL77]. Fill [CEG86]. Filon [Fli60]. Filtering [And65a]. Filters [Blu59, Gau64, HL66, Orm61, Sha66, Wad56]. Filtrations [BDK24]. FINAC [EV57]. Find [AI16, KKT95, Gab06]. Finder [de 86]. Finding [BRK74, CFM10, CT06, Fra58a, FGM88, GNP^{+}93, GT89, Hsu75, IK74, KLP75, La 64, LP79, Man75, MM75, MT69, PKR84, Pag74, PS78, RS94b, RS63, RS99, Sla64, Sla72a, Val15, Wei72, WW73, Yao81a, Yen72, CD08, FKV04, MP91]. Fine [DKK24, BGM99, MT88]. Fine-grained [DKK24, BGM99]. fingerprint [Tar08]. Finish [GS76]. Finite [AAI69, ADO91, Arb61, Bav68, BGG^{+}24, BH65, BZ83, Cha66a, Cha69a, CGSV24, Con57, Cry70, Cry71, CF59, DLN73, Das77, DA68, Dun74, DS92a, DS92b, FT88, Fri62, Gib69, HVL17, HvdH22, Ja’80, JM78, Kam82, KB89, Kar81, KV94a, Kel61, Kne93, KS66, Kon24, KR76, Liu69, MM81, MS21, MT88, Rit63, RS94c, Roc94, SB82, Sav72, TZ16, Wei72, Win82, Zaf69, ANTSV02, CSV09, GHL25, MGLA00, OG96, DH94]. finite-horizon [MGLA00]. Finite-Source [Kam82]. Finite-State [BZ83, Liu69, OG96]. Finite-Valued [TZ16]. Finitely [Hul84, MP24]. finiteness [Kif98]. First [AM90, BKMZ15, CBHH61, Che17, DP85, DH59, Gab16, GKS17, HN84, Her74, JL23, LW63a, LW63b, Man69, MG93, Phi62, Pic64b, Pic64a, PZ19, Rei80b, SK77, SW68, Two63, Wil84, van72, BG98, BLSS03, BKL^{+}06, DKT13, FG01]. First-Order [AM90, BKMZ15, Che17, Gab16, GKS17, HN84, Man69, PZ19, Rei80b, BG98, BLSS03, BKL^{+}06, DKT13, FG01]. Fishspear [FP94]. Fitting [Aki70, AF58, CADK^{+}24, Mae60, Mae63, SS58, VV63, JMM^{+}03]. Five [But67]. Fixed [AM94, CLY25, ER80b, EV62a, EV62b, Gro12, KLM^{+}97, LM86, Meg84, OD67, RW63, Wol65, AMN^{+}98, CLL^{+}08, CD08, DJKK08]. Fixed-Connection [KLM^{+}97]. fixed-parameter [CLL^{+}08]. Fixed-Plus-Variable [EV62a, EV62b]. Fixed-Point [OD67, Gro12]. fixed-value [DJKK08]. Fixpoint [AVV97, AM82, SN85]. Fliess [Kor89]. Flipping [BHT18, Asp98]. Floating [AM59, CO84, JMMW79, KL73, Mol67, Tsa74, Wad60]. Floating-Point [JMMW79, KL73, Tsa74, Wad60]. Flocking [Cha14]. Flow [BDF81, CKL^{+}25, CT81, DGG^{+}25, EK72, Ehr58, Fai67, GT88a, GT88b, GW76, HU74, Ita78, KU76, KKPW25, KR77, KS75, KL95, Kou77, Kri64, Laz84, MS11, Mil79, OV20, Rey77, Röc84, Ros79a, Sco58, SM90, SN85, Str83, BL04a, BK09, CK09, GR98, KP03, LR99]. Flow-augmentation [KKPW25]. flow-cut [CK09]. Flow-Shop [CT81, KS75]. Flowchart [Hun82a]. Flowcharts [BB85, Yel75]. Flowgraphs [Bak77, Mil79]. Flows [Joh87, LP25, LNP^{+}25, Zad72, ARV09, BLR10, CKL^{+}07, KRV09, MMHW97]. Flowshop [PK80]. Floyd [BÉ91]. Floyd-Hoare [BÉ91]. Fluid [Har57]. Flutter [Fra58b]. FO [BKTW22, SSV22]. Folk [Haj83]. Fooling [OST22]. fools [Bra10]. Footnote [Gol60]. Forbidden [Lev13]. Forcing [Gai20, JS08]. Ford [FL18, Man79a]. Ford-Johnson [Man79a]. Forest [AGH^{+}25, BHM11]. Forests [SV19]. Foreword [Ano15, KV04, Tar16a, Tar16b, Tar16c, Tar16d, Tar17a, Tar17b, Tar17f, Tar17c, Tar17d, Tar17g, Tar17h, Tar17e, Tar18a, Tar18b, Tar18c, Tar18d, Tar19a, Tar19b, Tar20c, Tar20a, Tar20b, Tar21a, Tar21e, Tar21b, Tar21c, Tar21f, Tar21d, Tar22a, Tar22b, Via11b, Via11d, Via12c, Via12d, Via13b, Via13d, Via14a, Via14c, Via14d, Via15b, Via15c, Via15a, WK15, Via10b, Via11c, Via12e, Via12a, Via12b, Via13c, Via13f, Via13e, Via14b]. Forget [NR94]. Forgetting [FKS95]. Fork [BMT89, DLT94, ZTX21]. Fork-Join [BMT89, DLT94]. Form [BKK84, CHT77, CM83, CM96, Cur59a, Cur59b, DHM65, FJ60, GR74, Gre65, MSW83, McK87, Min70, Pra60, SB82, Wat73, dL89, BP03b, Dar01, FCMG22, Noe79]. Formal [Apt83, BOG02, Blu66, BK78, Cha74, CS68, Cud70, FKRV15, HR77, LW57, McL84, MNO88, PD98, QH68, Ros69, Yu75, Zei72, HW08]. Formalism [SFR68]. Formalisms [Fis65b]. Formalization [Lom73, MP70a]. Formally [ABK16, Cha75]. Format [AB70, HY84, Lov69a, LR81, Rei71]. Formation [YUKY17]. Formed [Bur76]. Forms [Cav70, GL76, Ja’80, Ros67b, Wil59b, AL05]. Formula [CL62]. Formulae [KV94a]. Formulas [AHK93, BE97, BEGY22, CO17, Cra55, Cra57, Dem92, DH59, Di 69b, Di 73, HP66, KLV94, LW63b, Ral59, Raz13, Sho77, Sur83, Tra69, ANP07, Ats04, Raz09, LW63a]. Formulation [MTZ60]. Formulations [Gav83]. FORTRAN [Wei66, Fer60, GHG60]. Fortran-Compiled [GHG60]. found [Sho03]. Foundation [AAJS17, HW08, JSMM03]. Foundations [CCMS22, CMT20, CM78, KLW95, MS96]. Founded [vRS91]. Four [But67, GKKS14, HR25, Mye92, Kol04]. Fourier [BT96, CM67, Cru76, DA68, Hoc65, KL70, LMN93, Mor73, Pap79a, Pea68, Pea69b, Roc94]. Fourth [Con57, Cra54, Cra57, Ano79]. FP [HWW90]. FPRAS [ACJR21]. Fractional [GLPR21]. Fractions [Mac55, MB56, Mae60, Spi61]. Fragment [SDPF81, SDPF87, Ven87, MS04, Ott12]. fragments [tCL09]. Frame [KLW95]. Frame-Based [KLW95]. Framework [BEJWY22, BG15, BS89, CGL^{+}22, HHP93, HS19]. Frameworks [HSTV14]. Fraud [HPZ24]. Fredholm [Fis58, Has70, Her74, KK70, SK77, SW68, Two63]. Free [Aho68, ACHS16, ALS94, BBK93, Ben80, Bli72, But78, CRGM78, Cry70, DG84, Ell72, FOS14, FW74, Fre82b, GS63, GU66a, GU66b, GL76, GR74, Gor63, GH72, Gre65, Gre66, Gre69, Gri68, Har67, Hay76, HU66, HR25, Huf57, IAWK83, KT69, Kor89, LR78, LW93, LTV96, Mau69, MSW83, Mey69, MMR15, Nau89, Par66, Pav72, Ros68, Ros70a, Ros67b, Sta82, Sud75, Sud78, AGHK09, BAL^{+}21, BGM^{+}98, DFHT05, DR00, IKL^{+}19, Jay97, JCT98, JSMM03, Lee02, SS06a]. Freedom [Kam80]. Freezing [Mol18]. Frege [BBCP20, Hås21]. Frequency [FJ60]. Frequent [KMP^{+}12]. Frequentist [XZ25]. Freshman [Sla63]. Frobenius [BGK^{+}22, MS21]. Frontier [Che17, KZTH05, GKS04]. Full [AF77, Bra77, EPT18, Sci82]. Full-Text [AF77]. Fully [DST94, GNP20, GIS99, HKT^{+}20, DI05, HK99, HdLT01]. fully-dynamic [HdLT01]. Function [Ash64, BE62, CF59, FJ60, Gia64, GF60, Gre73, HP72a, HA56, LP15, Luk57, McK78, Mie57, MP65, MP66a, Nat67, Nau83, Nau89, Net59, Phu76, Pra60, She57a, She57b, Sho79, Weg69, dV71, FCB08, GJ12]. Function-Free [Nau89]. Functional [AMP73, ADS83, BDFS84, BV18, Cur59a, Cur59b, Fri63b, GNP20, GZ82, Hon82, Lom61, MSY81, MP70a, Via87, KT99, RV07]. Functional-Dependency [GZ82]. Functions [AM85, Alt88, ABB^{+}17, Bas58, Bav68, BAG14, Ben80, Ber69, BHT18, Blu67, Bol79, BM75, Bre76, Bsh93, BT96, CK15, CLW95, Chu65, CC67, CaKMTM19, CvH65, CU59, Dub66, DNRS03, EGGI92a, EGGI92b, Fai85, Fra58a, Fra65, GKR16, GB74, GGM86, GK58, Har63a, Hay75, Hay76, HL87, Hir58, Jia23, Klo61, Klu82, KSG68, KB81, KT78, Liu69, Mac55, Mar58, McK87, MV15, Pet62, PT57, Ral63, Ran58, RS83, Rei87, Rob74, Rus82, San96, She57a, SS63, Sti81, Sto64b, Svo75, TM66, TI70, Ull72, Wan74b, Wee66, Weg88, Zha21, dWF58, ADM^{+}99, BBB^{+}00, BDG^{+}13, CL11, IFF01, NR04, YCW99, ZW03]. Fundamental [Kel73a]. Further [Har63b, Pag74]. Fuzzy [AJM80, Lee72].

Gage [Per54]. Gallai [SS13]. Gallery [AAM22]. Game [AIK84, Ber63, Con91, Lai13, LT82, MT21, Nau83, Rei80a, San69, SD69, Tar83, KZ98, Vaz12]. Games [ABS15, AHC19, BGG17, CH14, CV17, Doy23, EY15, HMZ13, HM13, KV15, ARV08, FGY00, GS07, PR08]. Gamma [GF60]. Gap [Con72, KV15, MV25, Din07]. Gaps [BSCI^{+}23, Bor72b, Bor72a, CK09]. Gates [Ben80, But78, Pip89, Ric73b]. Gaussian [ABDH^{+}20, MV25, Mar76, Ske79]. Gaussians [LM23]. Gaver [BCPW74]. Gene [ZZ17, ALS09]. General [BJ66, BW57a, BGN^{+}22, BO56, Bre74, BM62, Cud70, DPW18a, DST92a, DST92b, DG57, GT91, GP83, Kar68, Lie59, LM23, Por66, RC69, SB82, Sur87, Tit55, Wil77, vRS91, ACR98, BFU01, BGM^{+}98, CW96, DR03, MT10, CI74a]. General-Purpose [BW57a, GP83]. Generalization [BM82, BDSV24, But67, GR60, HR78, Kra62, McG59, RT69]. generalizations [MMW07, Niv10]. Generalized [AK70, AS54, Bab75, Bai66, BIW63, BD87, BB61, CH86, CH88, CL62, Cur61, Cur63a, DLN73, DP85, Dye68, EKLS20, Erd66, GMS09, GLPR21, GS64, GJ82, Gri68, HZ80, LP73, MN24, NS82, Oli67, OV20, Pic64b, Pyl64, RP89, Yeh68, Yoe65, MSVV07, Noe79, SS06c, ZW03]. Generals [Bra87, LSP82, Lam83]. Generated [BR71, Hou58b]. Generating [BBK93, CLW95, Ehr73, HHHN14, Hob93, LY61, LR77, Mow67, Mul59, MC72, Zer85, BP03a]. Generation [AJ76, AJU77, ACJR21, AM78b, BL75, BS76, Cha72, CU59, Cov60, Den61, Fis65a, Gib69, GLW70, MB56, Mos54, Nat67, Nev70, PGE75, Pro80, RV78, SU70, Spr92, Van67, Win82, YCW99, BF01, DFSL14, Fus90]. Generative [Gor63, GKpKL22]. Generator [GSK59, Gre61, Kro64, Rot60, Wes67, SU05]. Generators [ADH63, Boy89, CH86, CH88, CM67, Ell72, Gil63, HD64, MM65, RT69, RW88, Smi71a, TRA71, Wes67, Ind06, Tre01]. Genus [FGM88, HI92, DFHT05, JM24]. Geometric [AKM25, ABC17, DL19, DS91, Hod70, RA20, Shi10, Aro98, ARV09, BFJ^{+}03, BGM^{+}98, CS11, GX22, Mul11, OR02, RV07]. Gerschgorin [Smi70]. Getting [Lam03]. GFSR [Tez87]. Given [BM92, Fit60, Kal88, RG67]. Givens [Cau58, CD94, Joh61]. Global [BNR13, BN13, GW76, KU76, Kou77, Wil78, AAPS96, DR00, vD95, vD97]. Globally [Was80]. go [Ram88, LS80]. Goal [Nev70, GJS12]. goal-oriented [GJS12]. Goals [LR81, Gra03]. Goes [HZ92]. Going [PZ19]. Golden [Got03]. Good [Cla79, Raz19, Tar75, KVV04]. Gossip [Hae15, Hae16, GGGK13, KKD04]. GPS [Ern69]. Gradient [FGHS23, KH64, Pyl64]. Gradients [JNG^{+}21]. Gradual [LTT22]. Graeffe [Bar60, Bau58a, Gra63, Wee60]. grain [Koc23]. Grained [CD94, BGM99, DKK24]. Grammar [CR86a, GG74, GL76, HS78a, HS78b, Hun82b, Lee02]. Grammars [ABL80, Aho68, Bea82, CF20, CS68, Cud70, Dil89, EF89, GH72, Gre65, Hib74, Hun84, Jaz81, KT69, MSW83, MP72, McN67, Mic76, MLS76, Pav72, Ros67b, Ros69, San72, Boy05, EY09, NV02, RS99]. Grammatical [HS71, KS63, MSW82]. grand [Fei03, Hoa03]. Graph [AV20, AGPT16, ACPS93, AK84, AM70, ADS83, BBE70, BDHS12, BLR14, BE97, Blu94, Che17, CDMP92, CG70b, CFG^{+}17, DM87, DHK19, Fre91, FGM88, GGS89, GT91, GJ76a, GMM81, GS93, HMR83, Hsu75, Ja’84, JRS^{+}14, KRV09, KV94b, KNY19, LaP93, Leh74, LW93, LB79, MB85, Mar76, ME67, MB83, MHG^{+}88, Mol18, MT69, Pav72, PS78, Pfa72, Sha70b, Sup83, TSOA80, Wan74a, Wei72, Wig83, ARV09, BK09, CFM10, EGIN97, FGY00, HMR^{+}10, HK99, KMS98, Kor08, SGP11, SW01]. Graph-Like [Che17]. Graph-Matching [GT91]. Graph-Representable [Sha70b]. Graph-Theoretic [GS93, FGY00]. Graphical [DG71, FN93, Moi19]. Graphics [CLRS86]. Graphs [ADK17, APY91, Bak94, BHM11, BGL^{+}22, Ber73, BJM79, BP15, BP23, BBB^{+}21, BGdM^{+}24, CGL^{+}23, CF20, CADMD21, CM93, CDRS93, CNP85, DLW25, DK65, DJM^{+}20, DEG^{+}21, EBA72, EPL72, FOS14, GSH89, Gab76, GHL^{+}87, Gib69, GKS17, GLNS25, GSV84, HI92, HU74, HKN18, Hsu87, Hsu88, IRT78, JMMW79, Kah95, KT17, Ker71, KPS19, Kow75, LMV16, MTTV97, MR87, Pfa75, PV76, Ram66, Ros72, SD76, SSV22, SL71, Sla76, Str83, TNS82, Wag76, WLS85, Wel66, ACKM09, ATSWZ00, BKS^{+}10, CGP02, DFHT05, DKT13, DGP07, Gab06, GKS04, Gro12, HW04, JM24, KKPW25, KK13, SPK13]. great [Bro03]. Greatest [Bro71, Kal88, MST91]. Greed [NV95]. Greedy [BW94, CK76, Fai85, JMM^{+}03, KR92, TWZ23, AAF^{+}01, HRM02]. Grid [CC16, FLS18]. Grid-Minor [CC16]. Grids [Hås21, FV02]. Gröbner [Baa93]. Ground [GNP^{+}93]. Group [Fle65, Glu63, Spi69, Tra66, Wee62, Wee65, Wee67]. Group-Testing [Glu63]. Group-Type [Tra66]. Groups [Bar65, Fle62, Lar76, Roc94, DG98, Ott12]. Grow [BC93]. Growth [Gof71, Ked79, Lam82]. Guarantee [Wig83]. Guaranteed [KS75]. Guaranteeing [KPW18]. Guarantees [BDSV24]. Guarded [BTS15, GHKL21, Ott12]. Guided [CO17, CGJ^{+}03].

half [Bro03]. half-century-old [Bro03]. Halfspaces [DDFS14, Vem10]. Halting [AF67, BM84, CF12, Ell72, Ull69]. Hamilton [EHR84, Rub74]. Hamiltonian [GHS75]. Hamiltonicity [CKN18]. Handling [BO56, GM70, GPV89, Lom61, Wal61]. Happened [BZ14]. Happened-before [BZ14]. Hard [AM20, ABD^{+}21, BHR84a, CS88a, CADMD21, ESY85, FPPR22, GSV84, JM77, LS80, LL73, MM24, Mey88, Urq87, CT06, CGH^{+}05, CSL07, MR08, dMRST20, HL13]. Hard-Core [ESY85]. Hard-Real-Time [LL73]. Harder [HS90]. Hardness [CHO^{+}22, DLW25, DMOZ22, Kho05, LY94, MS11, AZ06, AZ08, ALM^{+}98, CN06, CK09, FGL^{+}96, GMS09, LBNS09]. Hardware [Bry91, BRSU97]. harmonic [BG00]. Hash [BCFC^{+}23, Gon81, Mar71, Riv78, ADM^{+}99, SS06a]. Hash-Table [BCFC^{+}23]. Hashing [BCFC^{+}23, Bol79, FNSS92, GL88a, Gui78, KU88, Lar83, LRY80, PT12, RS77, Ull72, Vit83, Yao85]. haven’t [Sho03]. Having [FZ77, HC72, Klu82]. Hazard [Huf57, IKL^{+}19]. Hazard-Free [Huf57, IKL^{+}19]. Hazards [YR64]. Head [CCF84, GM73, GM75, One75]. Headquarters [Man03]. Heads [CCF84, JSV97, YR78]. heap [Cha00a]. Heaps [FT87, Fre99]. Heat [Dou59, Ehr58, Sig67]. Heavy [Gav71, Mit72, ZTX21]. Hedging [AP25]. Height [Dev86, Drm03, Ree03]. Hellman [DG21]. Help [EC64, LaP93, Rei87, BRSU97]. helps [Vöc03]. Henschen [Bri90]. Henschen-Naqvi [Bri90]. Herbrand [Lov72]. Hermite [GL64]. Hermitian [dS58]. Heuristic [BM85, GK88, IK77, MB85, Sim63, Sla63, SB68, dCS77, de 83, ZS99]. Heuristics [BM83, KR92, MP10]. Hidden [FSV06]. Hierarchical [CaKMTM19, Len89, Mit74, SK80]. Hierarchies [BY73, Dea89, Di 69a, ES86, LED76, MSW82, PZ19, RC70, TS81, BGJ10, Jay97]. Hierarchy [AH92b, BBS86, DV14, Gre69, HRST17, Ko90, WEM25]. High [AC24, BDD^{+}24, BC54, Bau58b, Den58, Fra61, Har86, KSY14, KMRZS17, Kro66, Leb56, MM75, Ral59, WYCF14, Yer66, AZ05, LW08, PSSZ13]. High-accuracy [AC24]. high-level [LW08]. High-Performing [BDD^{+}24]. High-Rate [KMRZS17, KSY14]. High-Scanning-Rate [Bau58b]. High-Security [WYCF14]. High-Speed [BC54, Den58, Fra61, Leb56, AZ05]. Higher [BDK24, Fis56, Fro61, Kob13, LGT14, NM90, PZ19, JR07]. Higher-Order [Kob13, LGT14, NM90, JR07]. Highly [Ott12, ASV06]. Highway [ADF^{+}16]. Hilbert [Att57, Baa93]. Hill [War57]. History [RW63, Ver95]. Hitchcock [DM62]. Hoare [BÉ91, CK79, Cla79, CGH83, Wan78]. Hole [CSSS20]. HOM [GG13a]. homogeneous [KP02, MP24]. Homology [BCL19, DG98]. Homomorphic [Cul79]. Homomorphism [Ros08, Gro07]. Homomorphisms [RWM94, ADK06, DGP07]. homotopic [DL07]. Homotopy [AGS17]. Homotopy-Initial [AGS17]. Honest [Rab94]. Hopcroft [Deo76]. horizon [MGLA00]. Horn [CH91b, EM07, Fag82, HW74, Lew78, NM90]. Huffman [LH90, Vit87]. hull [Cha01]. Hulls [GW23, Yao81a, Cha10]. Human [Nev74]. Hwang [Man79b]. Hwang-Lin [Man79b]. Hybrid [Gai20]. Hydrodynamic [Har57]. Hyperbolic [Mac55]. Hypercubes [BC93]. Hypergeometric [AS54]. Hypergraph [CLTZ18, Mar13, Ott12]. Hypergraphs [Fag83, Sac85]. hyperlinked [Kle99]. Hypertree [GLPR21, GMS09]. Hypotheses [CRTY23]. Hypothesis [GLR^{+}25].

I/O [CAA74]. IA [Ste57]. Ianov [Rut64]. IBM [ADW68, Bac54, Bak56, BBR54, Boy57, BO56, Gor56, GKR55, HK54, Wei66]. Iceberg [BCFC^{+}23]. Ideal [Coh55, CDW21, GP18a, LPR13, BEP09, Hal07]. Identical [Cha75, Ric73b]. Identification [Oet57, RC86]. Identifier [BG70].

Identifying [KMP^{+}12]. Identities [Sch80, TC21]. Identity [DG21, GP18a, AK98, AB03, SS13]. Identity-based [DG21]. II [BN13, BSZ65, BB64, BW57c, Cha90b, CM60, CI74b, Drm03, Dun66, DS92b, EGGI92b, HH94, Kel73b, Kob74b, Lam86b, Mae63, McD82, MR60, Mul91, MP71b, RS67]. III [Mae63]. Ill [MS19]. Ill-conditioning [MS19]. im [BGI^{+}12]. Image [DST92a, DST92b, HM85b, Ric74, WYCF14, Hoc01]. Images [Mon69, Yak76]. Imbedding [FGM88, Was71]. Immortality [Hoo66a]. Impact [FKS95, ARV08]. imperfect [GS07]. Implementation [AK70, BS59, BJL^{+}82, Ehr82, FLSY74, Gab76, GGS89, GL70, KU88, Kau71, BLP09, HJV01]. Implementations [HHPW14, Nou83, AD01, AGHK09, EFR07, FG10]. Implementing [FLMS93, FKL98, GH23]. Implication [CKV90]. Implicational [Hul84]. Implications [GJ78, MSY81]. Implicit [BKS15, But79, Cas75, Cas76, Con57, Cra55, Fre83, Jen73, AT11, Kor08]. Implies [BHT18, CFGS22]. importance [Coo03]. Impossibility [DV16, FLMS93, FLP85, BC05]. Impossible [BEGH17, Cla79]. Improved [AD01, BR15, BFGL79, DEL80, FT87, GW95, JSS16, Kro66, LPSP15, Lue09, Niv10, Rei80a, Rou01, Rub22, SS13, SV19, Sin86, VV17, dCS77, PPSZ05]. Improvement [Lov77]. Improvements [EK72, Man79b]. Improving [AKS15, BFG03, KMS^{+}25, Wig83]. in-depth [SPK13]. in-place [FG05]. Inapproximability [DFHM24, GSV15, GLR^{+}25, Hås01]. Inaugural [Car57]. Including [Sla72b]. Inclusion [CKV90, GF80]. Incompatibility [BBEG18]. Incomplete [DH59, IL84, LW63a, LW63b, Lip81, Wal61, BLPS10]. Incompleteness [Wan78]. inconsistent [ACN08]. Incorrectness [BGGR23]. Increase [JM60]. Increased [CK65, Klu83]. Increasing [HH57]. Increasingly [CKP17]. Incremental [BGSS20, MS89]. Incrementality [GM80]. Independence [BR91, CH91a, FE76, GL22, HU66, Ros68, BN05, BKS^{+}10, Bra10]. Independence-Reducible [CH91a]. Independent [AC81, AHPW19, BDSV24, BBH^{+}21, BF67, Blu67, Cha16, Dem92, Fra65, Gin61a, Hsu88, IK77, KW85, Sag91, Sah76, SC80, CDRS23]. Index [Ano58b, Ano95a, Ano95b, Ano97, Ano98a, Ano98c, CS77, Fra61, Fra76, GS69, GK68, HKMW66, IOT19, You63, BYG08]. Index-Digit [Fra76]. Indexability [Yi12, HKM^{+}02]. Indexed [Aho68]. Indexes [AG86]. Indexing [BK77b, CM78, FM05, MK60, Mar61, O’C64, SL68, YS76, YSS78, FLMM09, Shi10]. Indistinguishability [BV18]. Individual [BSV20]. Induced [ADK17]. Induction [RBCC19, Pit06]. Inductive [Kra62, Pit89, Weg77]. inefficiency [RS09b]. Inequalities [Klu88, LGT14, Sho81]. INF [Sho77]. Inference [Ang82, BD94, Chi86, KU80, KBI^{+}17, Kra62, McA93, MG93, Moi19, MR87, Nau89, Pit89, RS94c, SGRR75, Sla72a, BGJ10, DS12, Dar03, DR03]. Inferring [Boy89, Sag80]. Infinite [AHC19, Cha69b, CKP17, Gre69, GL22, Har86]. Infinite-duration [AHC19]. Infinite-State [CKP17]. Infinitesimal [Sur87]. Influence [Low68].

Information [Abe80, AGPV90, Bak62, Beu70, BO56, BT57, Cha74, Cha75, Coo64, Dal73, DFGN24, DR85, FKRV15, Fai67, FHS13, GKR16, Gla57, Glu63, GYY80, GM61, GM70, GPV89, Heh77, IL84, JP84, Lip81, MK60, Mor68a, PM71, Rab89, RZ19, Sal63, Sal71a, Sin86, SH22, Sti61, Was80, Wer79, Win65b, YLC76, YLS82, ACN08, AL05, BLPS10, CKGS98, CM04, GS07, GK04, Gra03, Lev13, MSS01]. information-technology [Gra03]. Information-Theoretic [Cha74, AL05]. Informational [GV61]. Informational-Logical [GV61]. Inherent [GU66b, HU66, Mau69]. Initial [AGS17, GTWW77, KK70, Man75, Usm66]. Initial-Value [KK70]. Initializing [CH86, CH88]. Input [Bar65, BF67, Cha70, CW17, Cof69a, Cof69b, Fer60, Gil66b, Gin61a, Hof84, HR63b, HR63a, MS59, Oeh63, RW63, Sch64, SM81, Win64, LMNM03, MRR03, DK59]. Input-Error-Limiting [Win64]. Input-Independent [BF67, Gin61a]. Input-Output [Fer60, Gil66b, HR63b, HR63a, MS59, DK59]. input-queued [LMNM03]. Input/Output [Cof69a, Cof69b]. inputs [ABF^{+}02]. Inquiry [Mar61]. Insertions [HS21]. Instability [Kro67, Mil75]. Instance [ABC17, OKSW94, de 86]. Instance-Optimal [ABC17]. instances [GS10]. Instants [SM81]. Institutions [GB92]. Instruction [CS83]. Instructions [Ano68, Mau66, MO56]. Insurance [Wis54]. Integer [BL19, BNNS94, Cla95, EC68, Fei84, FJWY25, Hei71, Kam89, Kan80, Kan83, LR77, MST91, MTZ60, Pap81, Phu76, SK73, Sri65, TS74, SS06b, Tho99]. Integral [And65b, Fis58, Gla65, Has70, Her74, KK70, MM84, Phi62, SK77, SW68, Tho57, Tra69, Two63, TC21]. Integrality [KV15]. Integrals [DH59, Gau61b, Har64, Kac79, LW63a, LW63b, dWF58]. Integrated [Bau56, FU82]. Integrating [Rah70]. Integration [And66, Bui79, But65, Car58, Cas75, Cas76, Ema63, Fli60, HH57, HP66, Kau71, Kro73, Lot56, Mor62, Pic64a, Rah69, RTW94, Sla63, Wil59a]. integrity [FL02]. Integro [Lin69]. Integro-Differential [Lin69]. intelligence [Fei03]. Interaction [DS92a, Yer65, JS96]. Interactive [BKN14, BEGH17, FGL^{+}96, GKR15, LFKN92, Bro97]. Interarrival [KW66]. Interconnection [Gol82, HP90, JT75]. Interdeparture [Tob82]. Interface [BER06, MT21]. Interleaved [BC73, Hof84]. Intermediate [RK23]. Internal [Flo61a, GL88a, HI59, IL79, Mac66, Goe61]. Interpolant [Sha67b]. Interpolation [Aki70, GL64, KP71, MT93, Sla70, TI70, Fer64]. Interpolator [Mos55]. Interpretation [FLO^{+}25, CCM12]. interpretations [GRS00]. interpreted [BL00]. Interprocess [Lam86a]. Interrupting [HY75]. Intersecting [CE92]. Intersection [BK86, BK87, CD87b, DL19, FCB08, ZS99]. Intersections [Com68, Vem10]. Interval [CRR18, Gel79, GL70, HJV01, LB79, NB95, AKN^{+}08, KJJ03]. Interval-Linearizability [CRR18]. Intervals [HS91]. Intractability [BBE^{+}21, MS92]. intractable [Lib01]. Intrinsic [Rou15]. Intrinsically [Dil89, Jaz81]. Introduced [CM96]. Introduction [LV09, MV05, Sal71b, SV06, VV09, Via11a]. Invariance [Pip89, PT20, HKM12]. Invariant [Doy62]. Invariants [BNR13, BN13, HOPW23, KBI^{+}17]. Inventory [CG57]. Inverse [Att57, Bas58, BIW63, DLN73, DHM65, DH56, Erd66, MP70b, Pyl64, Tit55, Cha00b]. Inverse-Ackermann [Cha00b]. Inverses [HC72, FN10]. Inversion [Cru76, DA68, HC72, Mar58, NBC72, Pea67, Pea69a, Roc94, Two63, Wat73, Wee66, Wil61, Zie68, Zoh69, von62]. Inverted [BBH^{+}87]. Invertible [Har63a]. Inverting [CLW95, FK99]. Investigation [SM64, Sri65, TW61]. Invited [Ano15, For16, Tar16a, Tar16b, Tar16c, Tar16d, Tar17a, Tar17b, Tar17f, Tar17c, Tar17d, Tar17g, Tar17h, Tar17e, Tar18a, Tar18b, Tar18c, Tar18d, Tar19a, Tar19b, Tar20c, Tar20a, Tar20b, Tar21a, Tar21e, Tar21b, Tar21c, Tar21f, Tar21d, Tar22a, Tar22b, Via10b, Via11b, Via11c, Via11d, Via12e, Via12a, Via12b, Via12c, Via12d, Via13b, Via13c, Via13d, Via13f, Via13e, Via14a, Via14c, Via14d, Via14b, Via15b, Via15c, Via15a, WK15, Via11a]. Involving [BL66, Ehr59, Har57, Rac82, dWF58]. IP [Adl05, BFM^{+}19, Sha92, She92]. IPMs [DGG^{+}25]. Irreducibility [Mus78]. Irreducible [HR63b, HR63a, Nat67]. Irredundant [BRK74]. Irrelevant [KW20]. Isolation [CG18, CFGS22]. Isomorphism [Ber73, BJM79, CG70b, Fle62, GHL^{+}87, KMR95, LB79, SD76, Ull76, HMR^{+}10]. Isoperimetry [BS23]. Item [GR20]. Itemsets [KMP^{+}12]. Iterated [Sch74]. Iteration [Bau58a, BÉ91, CV59, Gau59, Gau60, HMZ13, KY83, KT74, Man67b, TW79, Was80]. Iterations [Hou56b, Was81]. Iterative [And65b, Bau78, BR71, BGRS13, CS85, Fis65a, HST76, KU76, MM24, Mol67, Par61, She59a, Tor64, Wai67, von56, BEG22]. IV [BGdM^{+}24].

J [Nag61]. J. [Bui77]. JACM [Cof03, Fis03, Got03, Hal97, Hal03, Lei03, Mil03, Ros03, Sal69, Jou03, Via10a, Via11a, Via13a]. Jacobi [Cor63, GMvN59, Han62, PT57]. Java [MT21]. Job [FMI80, Leu93, MS11, Rus82]. Jobs [Bru76, CAA74, McK87]. Johnson [KN14, Man79a]. Join [BMT89, DLT94, MSY81, NPRR18, Sci82, ZTX21]. Joins [BC81, HT24, Mor97, Tay93]. Joint [dL89]. Jordan [Loi69]. Journal [Alt03, Ano68, Ano74, Sal70]. Junta [CST^{+}18]. Justification [Apt83].

Kanellakis [Ano98b]. Karp [Zad72]. Kedma [ACDF78]. Kernel [BEL21, Kac79, LP79]. Kernel-based [BEL21]. Kernelization [BFL^{+}16, KW20]. Kernels [KK70]. Key [Bol79, KOY09, McG59]. Keys [Bur76, MS16a, Riv78, TTK66, BF01]. Kikuchi [WEM25]. Killing [TW24]. Kilter [Sin86]. Kind [DH59, Has70, Her74, Phi62, Pic64b, Raz16, SK77, SW68, Two63]. Kinds [BM83, LW63a, LW63b]. Kinetic [Rub15]. Kinodynamic [DXCR93]. KKT [FGHS25]. Kleene [MS21]. Knapsack [HW76, HS74, IK75, Mey84, Sah75, Yao82, Zol78]. Knapsacks [BKS18, ISSS22]. knot [HLP99]. Know [FHV92]. Knowledge [CFGS22, DS92b, EFKP22, FHV91, FHV92, FH94, FH98, GMW91, GOS12, HM90, HZ92, HT93, NT93a, SK96, BBP95, BLMS97, DNS04, SV03, HZ92].

Knowledge-Based [HZ92]. Knuth [Pfa75]. Koch [RW23]. Kronecker [AK70, SPK13]. Kuratowski [Wil84]. Kutta [But67, But79, BL66, Car58, Cas75, Cas76, CR56, Fro61, GR60, Gru70].

labeled [FLMM09]. Labeling [ADK17, FK16, Sam81, ST72, KT02]. Labelling [DST92a, DST92b, DEG^{+}21]. Lagrangian [JV01]. Laguerre [Wee66]. LALR [SSSU83]. Lambda [OR00]. Lambda-calculus [OR00]. Lamport [BZ14]. Language [Bro58, CK79, Cla79, FGP64, Flo61b, GHG60, Gra62, Har67, Mau69, McL65, Nev70, Rem81, VK76, Wir68a, Wir68b, AK98, Cap02, DR00, Lai13]. Languages [Ang82, BBK93, Boo78, BM62, CB72, CDPP23, CRGM78, DKRW15, ER64b, ER80b, ESvL80, Fai85, Gil66a, GR62, GR63a, GS63, GR63b, GS65, GU66a, GU66b, GH68, GMW91, Gor63, GLJT23, Gre66, Gre69, HS71, HH74, HU66, HR77, Hun82a, KLW95, Klu82, KB67, Kor89, KY83, LW93, Lyn77, MNO88, MH82, NB96, Par66, Ros79a, Ros67a, Ros68, Ros70a, Ros69, San65, Sud75, Sud78, VVAG97, BP03a, BDLW98, BLSS03, BGM99, FMW10, GK04, KMP00, NV02, Van07, Ano79, Gra77]. Lanky [ES15]. Laplace [Bas58, CM60, Cru76, DA68, LZ01, MP59, Tit55, Wee66]. Laplacian [CLTZ18]. Large [CLRS16, CCL23, Dea89, ES76, Har57, Has70, Joh61, Kau71, KT90, Lie59, Pea69b, Pol64, SU82, Sch66, TW84, CM04, RV07]. Large-Scale [Kau71]. Largest [Gau61a]. Last [MB83]. Latency [Knu61]. Latent [Bak62, Win65b]. Lattice [AR05, Ber69, CRS01, Far63, Liu69, Reg04]. lattice-based [Reg04]. Lattices [CDW21, LPR13, Kho05, Reg09]. Law [BS23, Erd66, ACKM09]. Laws [HM85a]. Layout [Ram86]. Leader [CD21, FL87, GS21, KPP^{+}15]. Learn [BDD^{+}24, DKP98, HPRW96]. Learnability [ABL^{+}22, BEHW89, LMN93, ABDCBH97]. Learned [LV21]. Learning [AHK93, ABDH^{+}20, ABL17, BDSV24, BBB^{+}00, BLQT22, CGL^{+}22, DJW14, DHL^{+}20, DFN^{+}24, FCMG22, FKS95, GS92a, GR17, JWW^{+}23, JNG^{+}21, KV94a, KLV94, KR97, LM23, LPR13, Men15, Men19, PV88, Raz19, Val15, XZ25, ZZC^{+}25, ZCLD25, BB10, BKW03, BLR13, BGM^{+}98, CBDF^{+}99, CHW12, GGR98, Kea98, Reg09, Vem10]. Least [DLN73, GLH57, Hib61, Mad66, MP70b, Osb61, Ple74]. left [TKB95]. left-linear [TKB95]. Legality [BBE70]. Leman [BN23, GLNS25, KPS19]. Lemma [AI16, BM82, FGYZ21, GST16, GJL19, Moi19, AKS12, MT10, HSS11]. Lemmas [Pet76]. Length [BGP17, Cer58, Cha66a, Cha69a, Gin58, Gli65, Gon81, KW89, LH90, LW57, NW73, OR90, Pag74, RZR24, WW73, dL89, GS06, Lue09, Yek08]. Length-Limited [LH90]. Lengths [MM84, Wol65, Yen72, LMNM03]. Lenses [ANP^{+}04]. Less [KT17]. Letter [AM80, GG61, Goe61, Lei91, Liv60, Nag61, WSD59]. Letters [Erc60, PGE75]. Level [HLS77, HR63b, HR63a, ILP97, LW08]. Level-Clocked [ILP97]. Levels [Zar59]. Lewis [HHR97]. Lexicographic [AG86]. lexicographically [CDPP23]. Lie [BK72]. Life [Wis54]. Like [Che17, CC76a, CC76b, GR63b, Koc69, BCH^{+}96]. likelihood [CT06]. Limit [KKG^{+}22, Mon70, ZTX21]. Limitations [BRS22, Cha74, HMR^{+}10, KV94a, PV88, Sch64, SH22, SRM97]. Limited [AACHE15, AACHE18, Cof68, GL88a, Hib74, LH90, PK80, RS66, RS67]. Limited-Entry [RS66, RS67]. Limited-Use [AACHE15, AACHE18]. Limiting [Sch74, Win64]. Limits [BDSV24, NS06a]. Lin [Man79b]. Lindenstrauss [KN14]. Line [BB95, BLS92, CE92, ES81, GS78a, Gri73, HL76, IM83, IL83, JS82, Kal88, LL85, Leh74, Mae77, Man75, MS92, Mor68b, AAF^{+}97, CDRS93, DGS01, GJL21, Par61]. Linear [Abd71, AM94, ADM^{+}99, AB70, ADH15, AGTG21, ADO21, ADG15, ABL17, BAG14, BJM79, BLTY94, CD87a, CS71, CG67, CKL^{+}25, Cla95, CLS21, Coh64, Col58, CC95, Cra55, Cry67, DGG^{+}25, DP14, EC68, EH86, EKS20, ERS24, Fai85, FZ77, Fra60, Gab82, GS78a, Gal82, Gau59, Gau60, GHL25, Gib69, Gil65, GLH57, GL88b, GW76, Gre66, GR79, GS90, GLW70, HPR15, Har76, HKN18, HS90, HK77, Jon62, Kam87, KLM19, KKT95, KSB06, KT19, Kor08, KSS10, LS11, LL78, Lin69, LS77, LKK83, Lov72, LB79, Man75, McC73, MW78, Meg84, MY80, Mey85, Mey84, Mor73, Mor75, Mou65a, NB96, Osb61, Pag76, Pav72, PK87, Ple74, RW92, RV87, Rei71, RG67, RPE81, RMK79, Rot70, SK78, San96, Sho81, Shr70, SC85, Sla72a, Sri76, Sto73]. Linear [Sud75, TNS82, Tar75, Tho57, TW84, TS81, TM66, TSOA80, Two63, Was80, Wel66, Yan85, Zoh74, AC05, BNWN25, BP11, CAFS21, Coh00, FGMS05, GS06, HW04, Kar00, OR00, RA20, Raz09, Reg09, SS06b, Tho99, TKB95, VY11, EGGI92a]. Linear-Constraint [BAG14]. Linear-Time [DP14, GS78a, Gal82, GS90, KKT95, Man75, NB96, TNS82, Kor08, KSS10]. Linearizability [CRR18, RK23, VSC24]. Lines [HW67, Hue73, Hue74, KM74, Zaf70]. Linguistic [ACDF78, Zar59]. Linguistics [IW71]. link [HLP99]. linked [BAP06]. Lipschitz [Eda22]. LISP [BM75, Sum77]. List [BC74, Fox70, GHG60, GH68, GX22, LWZ21, AS13]. List-Processing [GHG60]. List-Storage [GH68]. Listings [CF12, Sco58]. Lists [DST94, MY79, BAP06, SS06a]. Literals [dV71]. Literature [Doy61]. Little [HZ92, KS20]. Live [BDFG19, Kou77]. Live-Dead [Kou77]. Lived [HHPW14]. Lively [Nag61]. LL [Bea82]. Lloyd [ORSS12]. Lloyd-type [ORSS12]. Load [Bra77, Rab89, RY91, TT85, AAF^{+}97, KP02, Vöc03]. Loading [AG73]. Local [AI16, AAPS96, Aki70, AGPT16, AF77, BNR13, BN13, BK14a, CHT77, FGYZ21, FGLS19, GST16, Gri86, GJL19, HSS11, Hue73, Hue74, KL73, KMW16, Moi19, OT95, RZR24, AKS12, FKP13, GHS13, MT10, ZY03, vD97]. Locality [BEPS16, CHO^{+}22]. Localization [ABL17, Gri86]. Locally [BEG22, GS06, KMRZS17, FG01, Yek08]. Locally-iterative [BEG22]. Locates [Hue71]. Locating [Gau61a]. Location [FH80, GSV84, New65, JV01, JMM^{+}03, KKD04]. Lock [ACHS16, SS06a]. Lock-Free [ACHS16, SS06a]. Locking [BS85, FHS13, KS83, Klu83, Kor83, MW84, RT90, TSG85, Yan82, Yan84]. Locks [KS83]. Log [AC24, CMS20, CdDPL^{+}24, Lyn77, HN04, Rei08]. Log-Approximate-Rank [CMS20]. Log-concave [AC24, CdDPL^{+}24]. log-space [Rei08]. Logarithmic [AZ06, BKV25, RV87, Sch88, Cha01, HdLT01, LP25, LNP^{+}25]. Logemann [NOT06]. Logic [AM90, AP93, AH94, AV82, BC14, BNR13, BBCP20, BFM^{+}19, BÉ91, BK14b, BGGR23, Bry91, BW57b, BW57c, CH91b, CC95, Dav17, Deb92, EG95, EH86, EKS20, ERS24, FMNP19, Gab16, GMSV93, Got95a, Got95b, HS91, Hog81, Jac82, Lee72, Lic23, Mam66, MT91, McD82, Net59, Nev74, Pel87, Rob65, Sag80, SDPF81, SDPF87, She65a, She65b, VK76, Win71, vRS91, AB05, AHK02, BMSS09, BLMS97, CW96, DFSL14, EGG00, GKS04, RS00, SSS96, SM02, TS10]. Logic-Based [EG95]. Logical [BHK59, CEW58, Eld59, GV61, Hod70, Hol62b, HKP84, HR63b, HR63a, KLW95, Moc56, SH21, TKDB24, UV88]. Logically [dV71]. Logics [AVV97, CF12, CGH83, CKP17, Gai20, GHKL21, HHP93, HLNW01, MST93, SC85, TK91, DV95, MR10b]. Logspace [ACJR21, LZ77]. Long [BCLM56, HZ92, HHPW14, Ram86]. Long-Lived [HHPW14]. Longest [AHU76, Gon81, Hir77, KR92, Lue09]. Look [Mac57, Rah69, San69, MMW07]. Look-Ahead [San69]. Look-Up [Mac57]. Lookup [Pap87]. Loop [Bli72, GK58, HS56, Hun82a, KK74, KM72b, Sho81, Swa80, vdLL86]. Loop-Free [Bli72]. Loopless [Ehr73]. Loops [BAG14]. looseness [ZY03, vD97]. Lopsided [KR89]. Loss [Sev74, CKOR14]. Lossless [GS07]. Lovász [AI16, AKS12, GJL19, HSS11, Moi19, MT10]. Love [GP18b]. Loveland [NOT06]. Low [AH92b, CGL^{+}22, CW17, DDFS14, FHS13, LO85, LST25, MV25, OR07, AM07, CGT10, FKV04, CK95b]. Low-Cost [CK95b]. Low-Density [LO85]. Low-Depth [LST25]. Low-Distortion [FHS13]. Low-Rank [CW17, MV25, CGL^{+}22, AM07, FKV04]. Low-Weight [DDFS14]. Lower [Abe80, AAHK94, AC23, AC05, AH92b, Asp98, BBH^{+}21, BN23, BL04b, Cha90a, Cha90b, CRTY23, CdDPL^{+}24, CADMD21, CFG^{+}17, FGR^{+}17, FMP^{+}15, Fre81, FG79, GH23, GHS13, GW85, GHS09, GLNS25, HHHN14, HS78b, JK78, Ked79, Kir81, Kis91, KMW16, Kun76, LY89, LST25, MST91, Mey85, MN24, Mor73, PKR84, Raz13, Raz19, RVW18, Sch88, She14, Sla70, Tiw87, Tro84, Wil14, YY76, Yao81a, AS04, AK10, BBC^{+}01, BSSV03, EFR07, FLvMV05, JLV00, Raz04, Ste03, SM02, HS78a]. LP [CLRS16, JMM^{+}03, MSU99]. LP-based [MSU99]. LR [Bea82, Mic76, MLS76]. LRU [OOW99]. LRU- [OOW99]. LSH [CK15]. LSH-Preserving [CK15].

M [DNRS03, Laz84, OM78, Rid76, Laz84, OM78, Rid76]. M/M/1 [Rid76]. M/M/C [OM78]. M/M/m [Laz84]. Machine [Baz57, Blu67, BCLM56, BS76, Bru76, CGL^{+}22, Cur65, Das77, Gau59, Gau60, GGV^{+}24, Gin58, Gin59, GH64, Har68, Hun84, JNG^{+}21, Leb56, Lee61, Leh61, LR65, LV21, Ove66, RW63, Rob65, Röc84, SM64, Swi57, Was54, Yer65, Zar59, ZL58, AAF^{+}97, BP01, BEP09, CN06, CHW12, GK96, BS59, Tre54]. Machine-Independent [Blu67]. Machine-Oriented [Rob65]. Machinery [Ano68, Ano74, Wil54]. Machines [AF67, AVV97, ADH63, Arb61, BF67, Bel60, Ber54, BZ83, BL75, Bsh93, Cas71, Coh64, Col71a, Coo71, Cur62, Cur63b, Elg54, ER64b, ER64a, ESvL80, FHV92, Far63, Fis65b, FKS95, Gil66b, GF67, Gin60, Gin61a, Gri68, GI82, GV61, Har63b, HS66, Hib61, HU69, HD64, Iba78, Kar67, Kle66, Liu69, Mey85, OF61, Rei87, RC69, Sav72, SS79, Sch88, Wan57, Wat61, BL04a, BEP09, KMPS09]. Macro [ESvL80]. Made [Hae16, BSW01]. Maehly [Dun65, Dun66]. Magic [DNRS03, Ros94, WZC95, SSS96]. Magnetic [HK54, Tre54]. Magnification [CHO^{+}22]. Magnified [Shr70]. Main [Smi79]. Maintain [Dea89]. Maintained [SD62]. Maintaining [Fre82a, MH89]. Maintenance [AS97, DS91, Gra55a, ACK08]. Majority [WXSD25]. Makanin [KP96]. Make [Set82]. Makespan [BDF81]. Making [CG70a, GRS00, Sla64, Wri75, MR93]. Malicious [GH23]. Malleable [DPW18b]. Man [Ham69, SM64, Yer65]. Man-Machine [SM64, Yer65]. manage [BKP07]. Management [ADLM14, HLMW92, Hof77, Hol62a, AAPS96, LMW05]. Managers [Jaf83]. Managing [GE01]. Manifolds [CGOS13]. manipulate [CSL07]. Manipulating [BK78]. Manipulation [Ata94, ADS83, FGP64, Flo61b, Ros73a]. Manual [Bla59]. Manuscript [Ano68]. Many [BCFC^{+}23, CS88a, DDST16, FMI80, GS92b, Mor59, Ram86, RW23, Wal88, HPRW96]. Many-Sorted [Wal88]. Map [CGP02]. Mapping [HRS88, Rab66, GS10]. Mappings [ADLM14, BSZ65, GH64, GS65, RZ65, FN10]. MapReduce [BEG^{+}21]. Maps [CKM^{+}14, CLY25, Doy61, Eda22, BL04b]. Market [DPSV08, VY11, BSZ06]. Markets [CPY17, JWW^{+}23]. marking [Adl05]. Markov [AAGT15, BN71, CS85, CDPP14, CI74b, DJM17, EY09, EY15, Hoc01, Hof84, KR14, KT02, MGLA00, RW88, Ste83, WSY^{+}19]. Markov-Chain-Driven [Hof84]. Markovian [MM84, MM86]. Marriage [ILG87, TTS21]. Massively [BHH23]. Master [Blu58, DS13, Rou01]. Mastermind [DDST16]. Match [FP86, KR92]. Matching [AV20, BHH23, BHR19, BKSW24, CD08, CP91, DP14, DGM94, FHTZ22, Gab76, GT91, Gal81, Gal95, GK88, HO82, HKT^{+}20, IRT78, JWW^{+}23, LPSP15, Mot94, Mye92, RA20, RR92, RW92, RPE81, Rot17, SH88, TWZ23, CFM10, MSVV07, Mye99]. Matchings [BBH^{+}21, CGS15, CKN18, DJM17]. Mathematical [Adi71b, Ber54, Dub66, ES76, Gia64, Gof71, Gri73, Kap68, Lom62, Mor68b, She67]. Mathematics [GOBS69]. Matings [And81, GNRS92]. Matrices [AK70, Bib81, BS74, BGdM^{+}24, Bun71, Cor63, CS84, DM62, DG57, EV62a, EV62b, FL74, GMvN59, GH59, Har60, HS75, HC72, IOT19, Joh61, La 64, Lio71, Loi69, MV25, Mar59, Net59, OD67, Ort60, Osb60, Pea69a, PT57, SD76, War62, Wat73, Zie68, dS58, von62, RV07]. Matrix [AY13, Ano57, Att57, Bau58a, BIW63, BR71, Ber58, CGL^{+}22, CLS21, Erd66, Gen78, Giv58, Has70, Hou56b, Hou58a, Hou58c, KS20, LM86, MP70b, MO68, Mor60, Oli67, OPSPW19, Pea67, Pul55, Ros67b, Sav84, SS17, Sim65, Sim67, WSD59, Wil59b, Wil61, WW18, Yel75, Zoh69, vdLL86, von56, AM07, AT11, BDHS12, CKL13, JSV04, Lee02, Zwi02]. Matrix-Algebraic [vdLL86]. Matrix-exponential [OPSPW19]. Matroid [BIKK18]. Max [BSP82, Bun71, DHK^{+}17, LNP^{+}25, PS81a, LR99, Chi86, DJKK08, RWZ23]. Max-CSPs [RWZ23]. max-flow [LR99]. Max-Flows [LNP^{+}25]. Max-Min [DHK^{+}17, PS81a]. Max-Norm [Bun71]. Maxima [BKST78, KLP75]. Maximal [BBH^{+}21, BHH23, BJM79, Cer58, CH14, CT81, KW85, Kel73a, Kel73b, Log79, NB95, Wer79, Zad72]. Maximin [KPW18]. Maximization [FNFSZ23, GW21, HP72a, OV20]. Maximizing [PT57]. Maximum [BBB^{+}21, Bry82, CKL^{+}25, DP14, FGM88, Gab76, GT88b, Hsu88, Joh87, KR77, LP25, New65, PKR84, SH88, SM90, Yan79, ANP07, BFJ^{+}03, BK09, CFM10, CT06, GW95]. Maximum-Finding [PKR84]. Maximum-Flow [GT88b]. Maximum-Genus [FGM88]. May [GS90]. Meager [LF88]. Mean [BR71, CV17, Dad88, FM79, PKT90, RL80, TSG85, RL81]. Mean-Payoff [CV17]. Mean-Value [RL80, RL81]. meaning [BG96]. Means [AMR11, CDOY11, HA56, SB82, AK10, ORSS12]. Measure [DI87, FGK09]. Measurement [FACP81, DKP07]. Measurements [GK58, MM24]. Measures [MY79, PF79, Pip89, dG89, AHPV04, FH01]. Measuring [CHJS15]. Mechanical [Ber63, BM84, EG82, Gor63, Lov68, Mey65, PPV60, Sha88, Wel66, Lov69b]. Mechanism [BILW20, DHKN21, LS11, Rou24]. Mechanisms [BKS15, DRY16, Kon24, Mur55, SAKS64, RS09b]. Mechanization [Bar60, Pie73]. Mechanized [O’C64]. Media [HWB56]. median [OR02, JV01]. Mediated [BKK^{+}18]. Mediators [GH23]. Meets [DGG^{+}25]. Membership [BMT92, OPSPW19]. Membrane [GS57]. Memoriam [Ano98b, DNRS03]. Memories [Fal62, Por66, Rao78]. Memory [AAD^{+}93, AG73, ABND95, AGHR17, BKV25, Bry82, BC73, CE72, CLRS86, EGM16, Flo64, FK85, Gau61a, Har67, Joa59, JM78, KS78, KU88, Kau71, Lie59, Log78, MSSV24, Mit74, Pul55, RC70, Raz19, SVN^{+}13, Smi79, Tau22, TG82, UW87, Wil87, WC74, Cap02, DHW97, DR00, FG99, GHS09, NR09, SN04]. Memory-Constrained [FK85]. MERCURY [BT69]. Merge [FB72, HHHN14, KG74]. Merge-and-Shrink [HHHN14]. Merging [AB70, And68b, BT79, Gli65, Goe61, HD73, Man79b, MBM89, Rei71, Sob62, YY76, HRM02, MPT96]. Merlin [HM13]. Mersenne [Fra61]. Mesh [AK84, KA88, MMHW97, Mor62, NS80]. Mesh-Connected [AK84, KA88, NS80]. Message [ABND95, DDWY93, Elk20a, Paw80, Pol64, Sch72, WH80, BPC^{+}12]. Message-Passing [ABND95]. Messages [MMR15, NS06a]. Meta [BFL^{+}16]. Metadata [ADLM14]. Metalgorithm [Ric75]. metatheory [PR99]. Method [Aki70, Alo60, ABV60, AMR11, BJ66, BIW63, Ber54, Ber58, Boj84, But65, But79, Cal59, CM74, CD60, Cor63, Cry71, CV59, DHM65, Dun65, Dun66, DG57, Ehr58, Fli60, Fra60, GC65, Glu63, Gof71, GMvN59, GF60, HH57, Han63, Has70, HHHN14, HW67, HI59, HP59, HR91, Jen73, Joh61, Joh93, Kei56, Kle67a, Leh61, Liu63, Lot56, Mar59, McL84, MN24, Mon68, Par55, Pet62, PT57, Pyl64, Rab66, RS63, Sal60, Shi69, Sho77, Shr70, Sig67, Ste67, SEY15, Sto64b, Tho57, TM66, TI70, Wal61, Wan93, War57, Was54, Wee60, WSD59, Win65b, YS76, Zaf70, ACR98, KT68]. Methodology [Bry91, Sum77, CGH04, LBNS09]. Methods [BK14a, Bau78, BNNS94, BF61, BEL21, Bui79, But67, BL66, Cas75, Cha62, Chu61, CPT94, DLN73, DG71, DG62, DW64, Dye68, Edm69, Fis58, Fra65, Fri73, Fro61, Gil60, GR74, Gor57, GS64, Gru70, Ham59e, Han62, HST76, Hen70, Kar68, KT67, KG74, Kro66, Kur62, Les58, Lin69, LFKN92, Mae60, Mae63, MB85, MM75, MR62, Mul59, O’C64, Par61, Sha67b, Sin86, SB73, Tor64, Wil61, Zel68, ORSS12]. Metric [BKMZ15, CRWW24, FL18, KSU19, JV01, KT02, Lee80]. Metrical [BLS92, BLR10]. Metrics [CAFS21, CADK^{+}24, KV15, CGT10, ENRS00]. MGS [Ple74]. Micro [Mer57]. Micro-Programming [Mer57]. Microcellular [Min67]. microprocessor [BY96]. Microprogramming [Gla56]. MIDAC [Coh55, Gra55a]. Middleweight [MT21]. Midpoint [GR73, Lot56]. Mild [LWZ21]. Mildly [CDW21]. Milne [Kei56]. MIMD [LKK83]. Mimimax [VV63]. Min [BSP82, DHK^{+}17, DGG^{+}25, GT88a, PS81a, Yan85, LR99, OR02, SU05, SW97, Chi86]. Min-Cost [DGG^{+}25, GT88a]. Min-Cut [Yan85, LR99, SW97]. min-entropies [SU05]. Min-Max [BSP82]. min-sum [OR02]. Mini [DR03]. Mini-buckets [DR03]. Minima [Ber69]. Minimal [BF67, BL79, Cha72, CT81, CR56, DM81, DDS87, EHR84, EW75, FM70, GS22, Gav83, Gin60, Hib61, Hom87, Hsu75, Hwa79, Jaf90, Nat67, Pag74, Pra60, dV71]. minimality [vD95]. Minimax [BDB23, CF59, Fra65, Sha70a]. Minimization [BHR84a, CMNP18, Law64, LY94, MP65, MP66a, Sch74, Ste67, AZ08, CN06, FFM08]. Minimize [BDF81, GS76, GS73, BL04a]. Minimizers [Jia23]. Minimizing [AGL00, CC76a, CC76b, CADK^{+}24, Jia23, KP03, Knu61, Nau89, IFF01]. Minimum [Bay72, CKL^{+}25, GT89, IAWK83, Joh87, KKT95, Kar00, LP25, LW57, Mai80, MT69, MR08, OR90, PW93, Sch63, Sev74, SL15, Sla64, Sup83, Wag97, Cha00b, CHL01, HdLT01, KS96, PR02, RS09a, RS66, RS67]. Minimum-Cost [CKL^{+}25, GT89, IAWK83, Sla64]. Minimum-weight [MR08]. Minor [CC16, DFHT05]. Minors [FLS18, Gro12]. Miss [FE76]. Missile [Orm61]. Mixed [ADU72, ADH63, BBD^{+}15, BCMP75, BNNS94, BH65, CM61, CD60, HD64, IOT19, KM72a, ASV06]. Mixed-Criticality [BBD^{+}15]. Mixtures [ABDH^{+}20, LM23]. ML [KTU94]. Mobile [ABF17, AP95, YUKY17, MN05]. Modal [Got95a, HS91, MST93, McD82, TK91, DP01]. Model [BBG89, BKTW22, Bra77, BK71, CM85, Cha90b, DM87, FHV91, FLSY74, Gia64, GH86, GN94, GB92, GHKL21, Gri86, GPV89, Her74, HY75, Hof77, HY84, Kle66, Kob13, KR76, KSS75, KSS77, LS71, Lov68, Lov69b, Lov69a, Mai80, Mor68b, Muk68, Par79, Par80, San88, Sch72, SEY15, TSG85, Wak69, Wil77, YLC76, YLS82, ALS09, BB10, BKL99, BKW03, CGJ^{+}03, GJ12, HKM^{+}02, KVW00, LPW02, MSS01]. Model-Based [Gri86]. Model-Theoretic [FHV91]. Modeled [BN71]. Modeling [Kob74b, Nel87, TG82, YDL93, Zei72, GE01, MMHW97]. Models [Adi71b, BBE70, Bay72, CMS89, CK68, CPT94, DVSC77, Gav67, GL71, Gel75, Gri73, HP90, HP72b, Hem64, KM72a, Lie82, Mae77, ME67, Mát90, MW84, Moi19, Tow80, Win82, LBNS09, PD98, RS00]. Modern [Bau58a, RVW18]. Modification [Fli60, GK59, KT68, Ste67]. Modified [BC74, But65, Gau61a, Joh61, KT67, Sha67b, Win65b, Zie68]. Modular [BSZ65, MS89, NSSV24, RZ65, Ros94, Sto70]. Modularity [APY91]. Modulators [Bla55]. Module [BHK90, Gus83, Wan93]. Modules [Mam66, SH21, de 86]. Modulo [NOT06]. Modulus [Leh57, Smi71a]. Moments [Alt62, MM84]. Monadic [AMP73, GK04, VBNB17]. Monitoring [BKMZ15, Swi57]. monitors [CSV09]. Monniaux [FLO^{+}25]. Monogenic [Hoo66b]. Monoids [AK25, BMT92, MT88]. Monotone [AG87, BT96, CPY17, FGLS19, Pot17, RW92, SEY15, Tal69, AACH12, EY09]. Monotonic [Lib01]. Monsters [ZM22]. Monte [Ehr59, FKV04, HH57, Ham56a, HH63, Mos58, RTW94, TM66, TI70]. Monte-Carlo [FKV04]. Moore [FL18]. Moore-Bellman-Ford [FL18]. Moran [GGG^{+}17]. Morphic [ERR81]. Morphisms [Zei72]. Morphological [Mey65]. Morphology [Fai67]. Morse [Bla59]. Moser [HS19]. Most [HL13, KPS19]. Motion [DXCR93, MV25, RS94a, BLMS97]. Motions [Rub15]. Motivation [LZ25, GJ78]. Movable [GM73, GM75]. Movable-Head [GM73, GM75]. moves [FG05]. Moving [Ehr58, One75, RS94a]. Moving-Head [One75]. Mr. [Nag61]. MST [Elk20a]. Mu [BBFD23]. Mu-Calculus [BBFD23]. Much [BDD^{+}24, HS90, BRSU97]. Muggles [GKR15]. Multi [AP25, Bab75, CYZG14, CPT94, DCKT92, GW21, Raz09, SKM18, ZZC^{+}25, EFR07, PR08]. Multi-Class [DCKT92]. Multi-dimensional [GW21]. Multi-Distribution [ZZC^{+}25]. Multi-Entrance [Bab75]. Multi-Facility [CPT94]. Multi-linear [Raz09]. Multi-Party [CYZG14]. multi-player [PR08]. Multi-Step [AP25]. Multi-Way [SKM18]. multi-writer [EFR07]. Multicategory [Ble66]. Multichain [RL80, RL81]. Multiclass [RTW94]. Multicommodity [LR99]. Multicounter [Iba78]. Multidimensional [CK95a, FP86, Har64, Wai67, Wil87, CC04]. Multifunctional [Cur59b]. multigraphs [KLSS05]. Multihead [FMR72, LS81a]. Multilayer [BBBL95]. Multilevel [AG73, BK79, Law64, RA76]. multimedia [MS96, MSS01]. Multimodule [LS87, WC74]. Multiobjective [BR15, MP10, SW91]. Multioperand [Pap87]. Multiparameter [Wei55]. Multiparty [BEGH17, GS22, HYC16, JS96]. Multiple [Bol79, Bre76, Chu81, CG86, Cur62, DR54, DHK19, ES86, Flo64, GW85, GFL87, HR63b, HR63a, KH64, LNSW59, Men65, Mor97, Mou65b, Sar83, Tay93, Tob82, Tra63]. Multiple-Bank [Flo64]. Multiple-Chain [CG86]. Multiple-Class [ES86]. Multiple-Gradient [KH64]. Multiple-Key [Bol79]. Multiple-Precision [Bre76]. Multiplication [Boy57, BK81, BK82, CLS21, Edw54, HVL17, HvdH22, Hor75, KB89, KS20, Spi69, Win67, BDHS12, Lee02, Zwi02]. Multiplications [Ked79, ST74, FPS98]. Multiplicative [FZ77, KB89, Smi71a]. Multiplicities [GFL87]. multiplicity [BBB^{+}00]. Multiplier [Fra61]. Multipoint [BGKM23, BGG^{+}24, KT74, Tor64, Was81]. Multiprocessing [CE72, CR95]. Multiprocessor [BL90, HY77, Lat81, MC70]. Multiprocessors [Bau78, CK95b]. Multiprogrammed [Bra77, Den65, LS71, Pri76, RC70]. Multiprogramming [AHY73, CW65, Gav67, Hel61, KSS75, KSS77, Lat81, LL73, Mit72]. Multiprogramming-Multiprocessor [Lat81]. Multipurpose [SB68]. Multiread [HV95]. Multiresource [Oma77]. Multisalesmen [BH74]. Multiserver [CI74a]. Multisets [JK78]. Multistage [HP90]. Multistep [But65, But67, Dye68, GS64, KT67, Lin69, Sha67b, SB73]. Multitape [GG72, HS66, KS66]. Multiterminal [GK94]. multithreaded [BL99]. Multivalued [HV95, Men79, Sag80, SW82]. Multivariable [Fer64, TM66, TI70]. Multivariate [BGKM23, BGG^{+}24, Col71b, Mus75]. Multiway [LGT14]. Multiwrite [Por66]. Mutual [BFG^{+}99, BJL^{+}82, DM25, Lam86a, Lam86b, RZ19, Sta82].

Names [ABF17, BF61, LY61]. Naming [BY73]. Naqvi [Bri90]. narrow [BSW01]. narrowing [AEH00]. Nash [Bab16, CDT09, Vaz12]. National [Mey65]. Natural [Cha69b, CHO^{+}22, MV15]. NC [AV20, CM93, MT88]. Near [AG17, ABDH^{+}20, BN23, BNWN25, CPSZ21, CC04, CAFS21, DJSW19, Elk20a, Fra65, GJ76a, HKN18, KLM19, KT19, LS11, RA20, Rub15, Sha70a, Cha01, Coh00, Kar00, ŠVV09]. Near-Linear [HKN18, KT19, BNWN25, CAFS21, RA20, Coh00, Kar00]. near-logarithmic [Cha01]. Near-Minimax [Fra65, Sha70a]. Near-Optimal [AG17, Elk20a, GJ76a, LS11, ABDH^{+}20, BN23, CPSZ21, CC04, KLM19, ŠVV09]. Near-Quadratic [Rub15]. Nearest [CK95a, MTTV97, AMN^{+}98, AMM09, Cha10]. Nearly [BBBL95, CS85, DDFS14, DGG^{+}25, DMOZ22, ES17]. Nearly-Linear [DGG^{+}25]. Necessary [Bon70, Esc60, Rob71]. Need [CCL23]. Needed [DDS87, GW85, AEH00, HPRW96]. Negation [BTS15, Ros94, Dar01, FCB08]. Negative [BNWN25, GT89, KV15, BG96]. Negative-Type [KV15]. Negative-Weight [BNWN25]. Neighbor [MTTV97, AMN^{+}98, AMM09, Cha10]. Neighborhood [Sup83]. Neighbors [CK95a]. Neighbours [BK14b]. Nested [Aho69, BBG89, DGG^{+}25, GPV89, AY13, BGJ10, TS10]. nesting [AM09]. Net [HPR15]. Nets [ADO91, Arb61, CEW58, CM57, CLL^{+}21, DG84, GK94, Gil60, Jef56, Koc23, LR78, MM81, Ric73b, Ric73a, SD62, AKN^{+}08, Rub22]. Network [Awe85, BPP^{+}16, Cun85, DPRS89, EK72, EKLS20, FT87, FK85, Hae16, IW71, KS78, Lam82, MC87, Mul78, NS82, Pol64, SM81, Sur83, Tow80, Wak68a, Wak68b, AAPS96, BPC^{+}12, GKPR07, Noe79]. Network-Oblivious [BPP^{+}16]. networking [ANV13]. Networks [AHS94, AEK^{+}12, BMT89, BM85, Bar16, BCMP75, BB95, BP89, BS89, But78, CHT77, CM83, CDRS23, CM96, CK95b, CLW95, Col87, CG86, CD21, Dad88, DLT94, DFGN24, DCKT92, ES86, FH80, Haj83, HP90, HKP84, Huf57, IS78, IW71, Joh77, Joh87, KT90, Kob74a, Kob74b, KLM^{+}97, KM20, MM84, McK87, MY84, MV15, MM86, Mow70, MP75, OR90, PKT90, Pip89, Pot17, RV87, RL80, RL81, RTW94, SB82, Tar72, Tiw87, TW88, Van78, Weg69, WW73, YY76, Yen72, YZJ18, ZTX21, dL89, ABNC^{+}96, AFGZ05, BCH^{+}96, BM96, Dar03, FPS98, JSMM03, KP02, KLOW00, MPT96, OG96, vD95]. Neumann [CM61, Muk68]. Neural [ADO91, Arb61, SD62, OG96]. neuroidal [Val00]. neuromata [ŠW98]. Never [EH86]. News [Alt57a, Alt57b, Alt57c, Alt57d, Ano54a, Ano54b, Ano54c, Ano54d, Ano55a, Ano55b, Ano55c, Ano55d, Ano56f, Ano56g, Ano56h, Ano56i]. Newsletter [Ano56a, Ano56b, Ano56c, Ano56d, Gol57a, Gol57b, Gol57c, Gol57d, Neu54a, Neu54b, Neu54c, Neu54d, Neu55a, Neu55b, Neu55c, Neu55d]. Newton [Boj84, SEY15, TW79, Zel68]. Newtonian [EKL10]. next [BK79, FMW10, Gra03, McC03]. Nicholas [Yer65]. No [DV14, FCMG22, FKR19, KRR22, Röc84, TSG85]. No-regret [FCMG22]. No-Signaling [KRR22]. No-Wait [Röc84]. No-Waiting [TSG85]. Node [WW73, Yen72]. Nodes [EBA72]. Noise [ABL17, BKW03, BKN14, BGM^{+}98, LS23, CBDF^{+}99, Kea98]. Noise-tolerant [BKW03, BGM^{+}98, Kea98]. Noisy [SH22]. Nominal [Gab16, LP14]. Non [CPY17, CST^{+}18, DPW18b, IKM17, MS92, Ukk82, HP07]. Non-adaptive [CST^{+}18]. Non-Clairvoyant [IKM17]. Non-FIFO [MS92]. Non-Malleable [DPW18b]. Non-Monotone [CPY17]. non-probabilistic [HP07]. Non-Real-Time [Ukk82]. Nonarithmetic [Lom62]. Nonbinary [ZZ17]. Nonclairvoyant [BL04a]. nonclairvoyantly [KP03]. Nonclausal [AM90]. Nonconstructive [FL88]. Nonconvex [JNG^{+}21]. Noncooperative [KL95]. Noncounting [CRGM78]. Nondeterminism [AP86, HM85a, XDB83]. Nondeterministic [Flo67, GH68, Gri68, GS90, HS71, Iba72, MT85, Mát90, SFM78, SN85, Ste03]. Nondistinct [Bur76]. Nonequilibrium [Kob74b]. Nonhomogeneous [WSY^{+}19]. Nonidentical [HS76, IK77]. Noninteractive [GOS12, BLR13]. Nonlinear [And65b, BJ66, BL19, Boj84, FR63, GS90, Kac79, KH64, Klo61, Mad66, Man67b, Moc71, Mow70, RR92, Sch88, Shr70, TI70, Was81, Zel68, EY09]. Nonmalleable [LP15]. Nonmonotonic [BNNS94, MST93, McD82, TK91, DFPP02, McD82]. Nonnegative [CS84, LM86, WW73, Yen72, JSV04]. Nonnegative-Distance [WW73, Yen72]. Nonnested [ESvL80]. Nonnormality [Loi69]. Nonnumerical [Bra67]. Nonoblivious [FNSS92]. nonparametric [BGJ10]. Nonpolynomial [Sha67b]. Nonpreemptive [CG93]. Nonpriority [Mit72]. Nonstandard [BB77]. Nonsymmetric [Hou58c, Mor60]. Nontrivial [DV14]. Nonuniform [GJ76b, PK87, Wil14]. Nonuniqueness [GSV15]. nonzeros [FJWY25]. Nordsieck [KT68]. Norm [Bun71, FHS13]. Normal [Coo66, FJ60, GH59, Gre65, HH63, Hoo66a, Hoo66b, Kro64, Mul59, Ros67b, Yas67, AL05, Dar01]. Normal-Form [Gre65]. Normalization [CLW95, ERS24, Lam82]. Normalized [CV59]. Norms [She59a]. Notation [RS83]. Note [Adi71b, Ano56e, AJM80, Cal59, Cas71, Chu81, Col58, Deo76, Dev86, Fik59, Gla56, GF60, HM72, Iba72, Joh73, Kam89, KP71, Kra62, Lot56, Mey69, MP70b, Mor73, Pic67, Pri76, Ros71, Ros70a, SB82, San72, Sar83, SF80, Sud75, TM66, Ull72, WW73]. Notes [Gre61, Lyn69]. Nothing [GMW91]. Notices [Alt57a, Alt57b, Alt57c, Alt57d, Ano54a, Ano54b, Ano54c, Ano54d, Ano55a, Ano55b, Ano55c, Ano55d, Ano56f, Ano56g, Ano56h, Ano56i]. Notion [MP71a, Vaz12]. Notions [GMV86]. Nowhere [GKS17, SSV22]. NP [AR05, AS98, AM20, Bak94, Coo03, GJ78, GPP^{+}19, GMW91, Got95a, GMS09, GI79, GSV84, HHR97, HL13, MS19, Mul11, MR08, Röc84, dMRST20]. NP-Complete [Bak94, GI79, Röc84, GPP^{+}19]. NP-Completeness [GJ78, MS19]. NP-Hard [HL13, AM20, GSV84, MR08, dMRST20]. NP-hardness [GMS09]. NQL [Sch78]. Null [Rei86]. Number [Abe68, AAI69, AM85, ADH63, AM78b, Ben80, BKST78, Boy89, CLRS86, CH86, CH88, CM67, DL19, DJM^{+}20, GR20, Gra63, GSK59, Gre61, GI79, Har63a, HD64, Kan65, Ked79, Kon24, Kro64, LV57, LP73, MM65, MM24, NR04, Rot60, SKM18, ST74, Smi71a, TRA71, Van67, Wan74a, Wes67, Fus90]. Number-Theoretic [GI79, NR04]. Numbers [Cer58, Cha69b, Cov60, FL74, Hab70, Hen56b, KR93, Loi69, MP66b, Mos54, Pen65, Pic64b, Sch68, Sto64a, Tez87, BT07]. Numeric [Chu81, Sar83]. Numerical [And65a, ABV60, Boo65, Bui79, But65, CBHH61, CW67, Cas75, Cas76, CM61, Col58, Cra54, Cru76, Cry67, Dou59, DA68, Ehr58, Ema63, Fli60, Fra58b, FR63, Gil60, Gla65, Gra55b, Her74, HW67, HJ61, Hou56a, Jon62, Kro67, Kro73, MY84, Mil73, Mil75, MR59, MR60, MR62, Moc71, Mor59, Mor62, Orm61, Phi62, Rab66, Rah69, Ske79, Spr92, Sta56, SW68, Two63, Usm66, Wee66, Wer79, Wil59a, Wil71]. Numerically [CLW95].

O [CAA74]. obfuscating [BGI^{+}12]. Obfuscation [BV18]. Object [AK98, KLW95, VVAG97, Yak76, BN05, BY96, FG10, ZS99]. Object-Creating [VVAG97]. Object-Oriented [KLW95, BN05]. Objective [CaKMTM19, Phu76]. Objectives [Doy23]. Objects [AI16, CRR18, CD87b, Com68, AGMT95, FKL98, JCT98]. Oblivious [AKL^{+}23, BPP^{+}16, CCMS22, Goo11, SDS^{+}18, GO96]. Observation [Ber75]. Observational [AGHR17]. Obstacles [RS94a, SR94]. obstruction [AGHK09]. obstruction-free [AGHK09]. Obtain [Cla79, Col87, KMS^{+}25]. Obtaining [Glu63, HR63b, HR63a, Mon68, Pet62]. Octave [Neg66]. Odd [CSSS20, RMK79]. Odd-Even [RMK79]. Off [AGPV90, BKT67, Dal73, Dou59, KT68, PU89, Pip78, Rei80a, BPC^{+}12, BSSV03, CS99, KL70]. Off-Step [BKT67, KT68]. Office [BB61]. Offs [DEL80, Ja’83, Sav84, Adl05, DI05, HMR83, LT82]. Ogden [BM82]. old [Bro03]. Omega [MC87]. On-Line [BB95, BLS92, ES81, GS78a, LL85, Man75, MS92, AAF^{+}97, CDRS93, DGS01].

Once [AHK93, BCFC^{+}23]. One [BHT18, BS76, CIV88, Coo66, FNFSZ23, Fis65a, FLP85, GGH67a, GH68, Ham69, HS71, Har68, HHPW14, KT74, LP15, Lip57, Neg66, Sel72, SL15, SEY15, Yer65]. One-Counter [SEY15]. One-Dimensional [Fis65a]. One-Normal [Coo66]. One-Pass [Sel72]. One-Plus-One [Lip57]. One-Point [KT74]. One-Register [BS76]. One-Shot [HHPW14]. One-Tape [Har68]. One-Third [Neg66]. One-Way [BHT18, CIV88, FNFSZ23, GGH67a, GH68, HS71, LP15]. Online [ABL^{+}22, AP95, BHR19, BSZ06, CS11, CvH65, DJSW19, DHL^{+}20, FHTZ22, HKT^{+}20, ZCLD25, MSVV07, NR09, Sei02]. ONR [Neu54a, Neu54b, Neu54c, Neu54d, Neu55a, Neu55b, Neu55c, Neu55d]. Ontology [BKK^{+}18]. Ontology-Mediated [BKK^{+}18]. Open [BCMP75, GS76, GK58, Win82, Red03]. Operating [SM64]. Operation [Cof69a, Cof69b, Heh77, MC74, Sch64, Wee62, HK99]. Operation-Preserving [Wee62]. Operational [HK55, RV80, RZ19]. Operations [Ata94, BK77b, Doy62, EV57, Erc60, GR63a, KW66, Oli67, Par55, RP66, Skl70, Wis54, AD01, Cha01, FV02]. Operator [Con72, Flo63, Hue71, Hue73, Hue74, Jon62, TW79]. Operators [Lom61, SY80, Tal69, HLNW01, RSG05]. optical [ABNC^{+}96]. Optimal [AAJS17, ABC17, AG17, ADU71, AJ76, ADK17, AGPT16, AKL^{+}23, ACW91, ACK08, BILW20, BZ23, BH89, BBBL95, BL79, BCH^{+}96, BGH^{+}95, Boj84, BLS92, BEGY22, BL75, BFGL79, CCF88, CGL^{+}23, CE92, CF12, CJLW22, CS88b, CD94, Cun85, DDFS14, DL07, DJSW19, DM81, DMOZ22, DRY16, DFN^{+}24, ES17, ES15, Elk20a, FGL77, FK16, FB72, Fri73, GNP20, Gal95, GJ76a, GS21, GL88b, GHLL22, GS16, GX22, GLW70, Hor78, HT24, HPZ24, HY77, IK74, ILG87, Ker71, KN81, KL95, KA88, KT74, LH90, LS11, Laz84, LP79, LS87, LS88, Leh74, Man79a, MM24, Mor62, NS80, NPRR18, PTW88, PGE75, RS66, RS67, Riv78, RY91, SKM18, SU70, SL71, SL15, ST87, Sta70, TT85, Tar08, Tar83, TW84, TWS80, TS81, TG82, VK96, Wil91, Yao85, ZZC^{+}25]. optimal [AGH00, AMN^{+}98, ABDH^{+}20, BN23, CKOR14, CPSZ21, Cha00a, CC04, CHL01, FGMS05, Hås01, KP02, KLM19, PR02, ŠVV09, NV95]. Optimality [Bol79, DP85, FB72, KMS^{+}25, Pap79a, Rei72, Tay93, Yi12, YW73, OOW99]. Optimally [FL74]. Optimisation [WSY^{+}19]. Optimization [AV88, AG73, BRS22, BFGP20, Bea72, Bea73a, Bea73b, BMR97, BG77, BR15, BEL21, Ear74, Fai85, FMP^{+}15, Fra69, FT87, GMSV93, HM13, HS90, JNG^{+}21, MS18, MSSV24, RC70, Rot70, Shi69, You77, BFG03, CHW12, SSS96]. Optimized [Den61, GC65, GG61]. Optimizing [BCFC^{+}23, Gor56, ILP97]. Optimum [CCF84, Chu65, Cra57, GM79, Gel79, KR89, MBM89, Wol65]. OptORAMa [AKL^{+}23]. OR-dispersers [SSZ98]. Oracle [BFM^{+}19, DHL^{+}20, KMR95, RT22, CGH04]. Oracle-efficient [DHL^{+}20]. Oracles [BBS86, CGL^{+}23, LS86, Tho04, TZ05]. ORAM [SDS^{+}18]. Orbit [COW16, Dye68, GHL25, KL86]. Orbit-finite [GHL25]. Order [AM90, AAJS17, BKMZ15, BDK24, CBHH61, Che17, CR56, Con57, Cra54, Cra57, Erd66, Fis56, FLO83, Fro61, Gab16, GW23, GKS17, Gru70, HN84, Kob13, Kro66, KT74, LGT14, Man69, MG93, MM75, MR62, Mow67, MP71b, NM90, Pet15, Pic64a, Pie73, PZ19, Rei80b, Wer79, van72, BG98, BLSS03, BKL^{+}06, DKT13, EGG00, FG01, GKS04, JR07]. Ordered [BG98, BGdM^{+}24, Fai85, MY79, AT07, BG01, SS06a]. Ordering [BZ14, CDPP23, EBA72, MB83, Rei71, Ros71, Sla72b, WH73]. Orderings [BD94, JR07]. Ordinal [BY73, Sri76]. Ordinary [BGP17, BKT67, Bui79, But65, Cas75, Cas76, Cha62, DG71, Gla65, Gra55b, Ham59e, Kro67, Kro73, MR62, Mor62, Pic64a, Rah69, Rah70, Usm66]. ORDVAC [You55]. Organization [CLRS86, EV62b, GS69, HY84, KMW67, Low68, Moc56, Pea69b, Win65b, EV62a]. Organizations [Gau59, Gau60, WC74]. Organizing [AM78a]. orientable [DL07]. Oriented [Gra62, KLW95, Nev74, Rob65, Wan93, BN05, GJS12]. Origin [DK65, Li59]. Orphan [HLMW92]. Orthogonal [AF58, Bro65, Cha90a, Cha90b, CEG86, Fre81, You67]. Orthographic [Wei66]. Orthologic [RS00]. Orthonormal [Rab66]. Orthonormalizing [DR54]. Oscillating [Sob62]. Other [BBE70, Boo82, Ehr73, GK84, JK78, VV17, Aro98, Gro07]. Out-of-Kilter [Sin86]. Outerplanar [BJM79]. Outline [Hol62b]. Outlines [Hob93]. Output [Cof69a, Cof69b, Fer60, Gil66b, Hof77, HR63b, HR63a, Leh74, MS59, SM81, DK59]. Output-Buffer [Hof77]. Overdetermined [Abd71, GLH57]. Overflow [EV57, Erc60]. Overhead [BKV25, LS71, RZR25]. Overlap [CAA74]. Overview [HH71]. owned [DR00]. Ownership [BN05].

P [Coo03, Mul11, SEY15]. P-Time [SEY15]. P. [ZY03]. Package [Wei66]. Packet [FGL77, Paw80, Tob82, Upf92, Adl05, AFGZ05, BFU01]. Packet-Routing [Upf92]. Packing [GR20, HM85b, LL85, RT88, Yao80, CJK^{+}06, GG13b, Sei02, XB06]. Packings [MTTV97]. PACT [Bak56, DL56, GM56, HS56, MO56, Moc56, Ste57]. Page [ADU71, FE76, SMC79, OOW99]. Paged [Cof69a, Cof69b]. Pagenumber [HI92]. PageRank [SGP11]. Paging [AS13, BBN12, FW74, IK74]. Pair [HB21, Liu69, Val15, Yeh68]. pairing [Fre99]. Pairs [Fre91, PS78, DI04, PSSZ13, Zwi02]. Pairwise [Cha16, KT02]. Pairwise-Independent [Cha16]. Palindrome [Man75]. Palstar [GS78a]. pants [DL07]. Paper [BCPW74, Bui77, CM57, For16, Gol60, HWB56, Nag61, dV59]. Paper-Tape [HWB56]. Papers [Gra77, O’C65, Ano79]. Parabolic [Con57, Cra54, Cra57, DG62, HW67, Zaf70]. Paradoxes [KP02]. Parallel [AP93, AM94, AGK^{+}17, ACW91, Ata94, BL90, BLT93, BMPT91, BW91, BKS17, BHH23, Bre74, BDFG19, CW67, CN95, CHK82, CR86b, CD94, DVSC77, DR00, Fal62, FT88, FKR19, FL18, GP83, Gal95, Gen78, Gil68, Gol82, GS85, GK84, HW90, HP90, HRS88, HST76, HT24, HK77, ICR87, JMG93, Joh87, KU88, KW85, KZ93, Kel73a, Kel73b, Kun76, LF80, LL78, LR79, LY89, Meg83, MC87, MP76, NS80, NS82, NT93b, Nic93, NV95, Pea67, Pea68, Rei87, Rei68, RVW18, SK78, SS79, Sch61, Sch66, SL71, SR71, Sto73, TNS82, Upf84, ÜD90, Wag97, Win75, Yao82, AGL00, BL04a, Cap02, CHL01, GK96, HHR97, KZ98, KMPS09, Spe97, YCW99]. Parallel-Connected [HW90]. Parallel-Correctness [AGK^{+}17]. Parallel-Search [Fal62]. Parallelism [BGSS20, Kel73a, Kel73b, Log79, BGM99]. Parallelizable [EFKP22]. Parallelize [LY25]. Parameter [Pet62, CLL^{+}08]. Parameterized [BBE^{+}21, EGM16, FLPS16, GLR^{+}25, Lin18, SLM19, DFHT05]. Parameters [DDFS14, Win71]. Parametric [Gus83, LR77]. Parametricity [OT95, SH21, TTS21]. Parametrization [Ehr82]. Paraphrase [SGRR75]. Parenthesis [Lyn77, McN67]. Paris [Ano98b]. Parities [Val15]. Parity [Cal61, Raz19, BKW03, KK13]. Parsers [GM80]. Parsing [ADU72, BW94, Ber75, CHK82, HH74, NB96, Sha70b, Lee02, NS06b]. Part [CDPP23, BNR13, BN13, BB64, BW57b, BW57c, CB72, Dun66, Lam86a, Lam86b, LZ25, Mae60, MR60]. Partial [AAJS17, Bra67, CN95, Cra54, Cra57, Cry71, DLS88, Esc60, Fis56, FP86, Fre82a, Gla65, HS19, HW67, JM60, LS81b, Sla72b, Zaf70]. Particularly [EV57]. Partition [EW75, Mul91, GJ12]. Partitioning [AM78b, AAI86, BSP82, Kir81, LGT14, Pea69a, PS81a, SKM18, Sto70, ARV09, BLR10, KRV09]. Partitions [HS74, Ker71]. Parts [Mae63]. Party [CYZG14, GHKL11]. Parvaresh [GUV09]. Pass [Sel72]. Passes [EF89]. Passing [ABND95, WH80]. Passive [dV71]. passwords [KOY09]. Past [NR94]. Path [ADF^{+}16, AMOT90, AKS15, BMPT91, BP23, Bra78, Gre73, HP59, Joh73, KW89, Llo82, MR87, NW73, OR90, SDS^{+}18, Tar79, Tar81a, Tar81b, Wag76, Wai67, WW18, Yel75, HRM02, JR07, TV19]. Path-Connectivity [Yel75]. path-merging [HRM02]. Paths [APY91, BNWN25, BG77, CL16, DP22, EHR84, Elk20b, Fre91, HKN18, Joh77, MP91, MR93, OV23, PS78, RS94b, Rub74, SV19, Shi80, SR94, Wan23, WW73, Yen72, AHPSV97, AMS05, AZ06, BS88, Coh00, DI04, PSSZ13, Tho99, Zwi02]. PATRICIA [Mor68a, Szp90]. Pattern [Alt62, AM82, Ble66, Doy62, DW64, DGM94, Gol82, HO82, Kel61, MF63, Mye92, RR92]. Pattern-of-Calls [AM82]. Patterns [BLR14, CMNP18]. Payment [BKS15, Sal60]. Payoff [CV17]. PCA [WEM25]. PCF [EPT18]. PCL [BDFG19]. PCP [Din07, MR10a]. PCPs [BSKK^{+}16, GS06]. PDAs [GF80]. Pebble [LT82, Rei80a]. Pebbles [DHK19, Kla85]. Peer [Kon24]. Pell [Hal07]. Percentiles [She57b]. Perceptrons [Kel61]. Perfect [AI16, AV20, CKN18, DJM17, GK88, Hsu87, Hsu88, RT88]. Perfectly [DDWY93]. Perform [Win65a, Win67]. Performability [dG89]. Performance [BW91, DI87, FW74, FB72, Gau59, Gau60, GM73, MY79, Nel87, NT93b, Rao78, Rus82, RT90, TSG85, TRA71, TPH86, Vit83, Weg76, Wig83, Wil77, AAF^{+}01, WR03, GM75]. Performing [BDD^{+}24]. Period [AM78b, BN71, NBC72, KLOW00]. Periodic [BB58, CM93, DPRS89, GF67, GB69c, GB70]. periodically [GE01]. Periodicity [HN07]. Periodification [KLOW00]. Periods [Dad88, Fit60]. permanent [JSV04, Raz09]. Permutation [BM92, EPL72, Fra76, KS74, MV15, NS82, Wak68a, Wak68b]. Permutations [Ehr73, HP99]. Permuting [GK84, PMH09]. Persistence [CGOS13]. Persistence-Based [CGOS13]. Persistent [DST94, LR78]. personalized [AT10]. Perspective [KSG68]. Persymmetric [HC72]. Perturbation [Gla91, Sur87]. Perturbations [MV25]. Perturbed [RW88]. Petri [CLL^{+}21, DG84, Koc23, LR78, MM81]. PH [RT22]. Phase [BS85, GS21, ILP97]. phenomena [MPR98]. Phenomenon [GR73]. Philadelphia [Hou56c]. Phrase [BM62, Gre65]. Physical [HRS88, BEP09]. Physics [Ber54, Yao03]. Pi [ABF17, PS00]. pi-calculus [PS00]. Picture [HP76, RP66, Sel72]. Pictures [AT71, Hue71, Ros70b, Ros73b, Sha70b, SR71]. piecewise [VY11]. piecewise-linear [VY11]. pigeonhole [Raz04]. PILOT [LNSW59]. PIR [DG16]. Pivot [Mul78]. PL360 [Wir68a, Wir68b]. Place [WSY^{+}19, FG05]. Place-dependent [WSY^{+}19]. Placement [CCF88, GS73, Mam66, MW78]. Planar [AV20, Bak94, BHM11, CGL^{+}23, CK95b, DGG^{+}25, DJM^{+}20, DEG^{+}21, Fre91, Hsu87, Hsu88, Joh87, KPS19, MP91, Mul91, PB89, BKL^{+}06, BK09, Cha01, Tho04]. Planar-Adaptive [CK95b]. Planarity [Deo76, FT22, GR79, HT74, Kna75, Len89, GIS99]. Plane [CE92, Hod70, Rub22, SR94, Wan23, Wei66, Wil78, YUKY17, HW04]. Planes [FPPR22]. Planning [DXCR93, RS94a, San69, BLMS97]. Plans [BJP59]. Planted [FGR^{+}17]. Plate [AC55, CD60, Was54]. Plausibility [FH01]. Player [HMZ13, PM71, CDT09, PR08]. Playing [Ber63, DDST16]. Pliability [RWZ23]. Plotting [Bas58]. PLS [FGHS23]. Pluralism [Smi82]. Plus [EV62a, EV62b, Lip57, Mac66]. PODS [LV09, VV09]. Point [Abd71, Alo60, AM59, ACW91, BG77, CK95a, CLY25, CO84, CF59, ER80b, HW67, JMMW79, KL73, Kei56, KT68, KT74, LM86, Mol67, OD67, RW23, Tsa74, Wad60, AF03, CD08, FKL98, Gro12]. Point-Disjoint [BG77]. point-to-point [FKL98]. Pointer [ABB^{+}17, GK96]. Pointers [BAG92]. Points [BG77, BKT67, BL66, GS69, JNG^{+}21, AHPV04]. Poisson [BH65, Hoc65, Laa58a, SB82]. Polarization [BGN^{+}22]. Polarized [CM74]. Poles [Sti65]. Policies [BJP59, Glu63, MT19, NT93b, PTW88, Yan82, LW08, MSU99]. Policy [BGH^{+}95, Leu93, One75, Rus82, Sal69]. Polling [KM74, Swa80]. Poly [HdLT01, LP25, LNP^{+}25]. Poly-logarithmic [HdLT01, LP25, LNP^{+}25]. Polyadic [GB69a, GB69b]. Polygon [ACW91, LP79]. Polygonal [AAI86, SR94]. Polygons [AHPW19]. Polyhedra [CH79, BS88]. Polyhedral [GML15, IKM17, AMS05]. Polylog [Coh00]. Polylog-time [Coh00]. Polylogarithmic [AACH12, AACHE15, AACHE18, BBMN15, BE11, BFGY19, BKSW24, Bra10, CL16, HK99, SSZ98]. Polylogarithmic-Competitive [BBMN15]. Polymorphic [JR07, Lai13, OR00, PS00]. Polynomial [AFN04, Aro98, BBS86, BZ23, BLQT22, BC61, BGP17, Bro71, CD87a, CC16, CLY25, CK09, CM93, Col67, Col71b, CKV90, CDW21, CF59, DV14, DFK91, ET76, FL88, FLS18, Fra65, Fri73, GSH89, GNP^{+}93, Gat83, GP18a, Hal07, HMZ13, HVL17, HvdH22, Hei71, Hen70, HW76, Hom87, HS75, Hor75, KB89, Kam89, Kan80, Kan83, KL86, KS16, Ko90, KMRZS17, KM20, Lad75, Leh61, LS80, Lic23, LS24, Mát90, McC73, Mey84, MN24, Moo67, Mou65b, Mus75, Mus78, OV20, OR02, PW93, Sch80, ST74, Shi80, Smi70, SEY15, VV63, Yan85, Gro12, HP99, IFF01, JSV04, JM24, Raz09, RS09a, ST04]. Polynomial-Space [LS80]. Polynomial-Time [BBS86, CM93, CKV90, DV14, FL88, FLS18, HW76, Kan83, KL86, Ko90, Mát90, AFN04, Hal07, OR02, JSV04, JM24]. Polynomiality [GR20]. Polynomials [AF58, BSV20, Boo65, COW23, FT88, Kal88, Rab66, RC86, Spi61, Wil78, Wil62, You67, BBC^{+}01]. Polytope [GML15, Rot17, AHPSV97, LZ01]. Polytopes [CK70, FMP^{+}15, OST22, HKM12]. Population [GS21]. Posets [FK16]. Positive [AG87, But78, CS84, AT11, Tho99]. possibility [BGI^{+}12]. Post [AF67, BB61, Hoo66a, Hoo66b, WM74, Yas67]. Posteriori [Tsa74]. Postman [Fre79]. Potential [CK95a, San96]. Potentiometers [HA56]. Potts [GJ12]. Power [ADO21, ABL17, BCP20, BK78, CIV88, DH94, DS92a, Gau64, Hen56a, HH94, Iba77, KRR22, KKG^{+}22, Min57, PT12, Smi82, Spi61, dV59, ACKM09, BDLW98, GK04, MSU99, SRM97]. power-law [ACKM09]. Powerful [Gur85, Jez16, KP00]. Powers [Nip90]. PPAD [FGHS23, DFHM24]. Practical [Fik59, HS87, Mad66, Mor68a, Sho79, Sta72]. Practically [ACHS16]. PRAM [BH89, CD94]. Pre [KNP^{+}24, Osb60]. Pre- [KNP^{+}24]. Pre-Conditioning [Osb60]. Precedence [ADU72, Col70, Flo63, GH73, Har60, KY83, Mar59, MP72, TPH86]. Precedence-Based [TPH86]. Precision [Bla55, Bre76, Cha66b, YS76, YLS82]. Precondition [KKMO18]. Preconditioning [SS17]. Preconditions [LZBY25]. Predicate [Hoa78, Man69, Sla70, VK76, van72]. Predicates [KS66]. Predicting [CEG86]. Prediction [BAL^{+}21, CW67, Kon24]. Predictor [Alo60, Cha62, CK65, GS64, Ham59e, HC63, Kar68, Kro66]. Predictor-Corrector [Alo60, Cha62, CK65, GS64, Ham59e, HC63, Kar68, Kro66]. Predictors [KT67]. Preemptive [BBD^{+}15, CG93, GS78b, GJ80, HLS77, LL78, Mar82, MC70]. Preference [CMT20]. prefetching [VK96]. Prefix [BP89, LF80, PGE75]. premises [BG96]. Premium [Wis54]. Preparation [Sco58]. Presburger [Sho77]. Presence [ADLS94, DLP^{+}86, DLS88, GGV^{+}24, GE84, LMS85, PSL80, RS94a, BFG08, CBDF^{+}99, DW04, FL02]. Presentations [Ros70a]. Preservation [GU66b, ADK06, Ros08]. Preserve [GR63a]. Preserved [Gin60]. Preserving [Bav68, CK15, KLM^{+}97, Rag03, Ram88, Wee62, Zei72]. Presidential [Car57, Hou57, Hou56c]. Pressburger [SJ80]. Prevention [BCLM56]. Price [GT23, JL23, Rou15]. priced [CL11]. Pricing [KUB^{+}15]. Primal [AK16, BBN12, Mul78, Nip90, ST73, DPSV08, JV01]. Primal-Dual [AK16, BBN12, JV01]. Primal-Simplex [Mul78]. Primality [AB03, GK99].

Prime [DP75, Smi71a]. Primes [DP75, Fis65a, Fra61, HS68]. primitive [AK98]. Primitives [Kor83, Sta82]. principal [CLMW11, Hal07]. Principle [Lee72, Rob65, HKM12, Raz04]. Principles [ADU71, Ano79, Gra77, XZ25, HJV01]. Priori [EC64]. Priorities [Adi69, RA76]. Priority [Adi71a, Bab75, DLT07, FP94, KK74, Leu93, NT93b, Sim84, Cha00a, MSU99, Tho07]. Privacy [CKOR14, DJW14, HKM^{+}22, BLR13, EFW10, FGY00, JRS09]. Private [ABL^{+}22, CKGS98, MV25]. Prize [AGH^{+}25]. Prize-Collecting [AGH^{+}25]. Probabilistic [AAHK94, AS98, BGB12, BZ23, BKK14, CHJS15, Con91, CM78, CY95, EPT18, Ell72, FKK11, GG78, GL22, IM83, Kar94, Ko90, KB81, MT85, MK60, MW84, NS06b, Pit89, Rac82, Sch80, SS24, SEY15, Adl05, CSV07, DS12, HP07, MPR98, RSG05, Tar08]. Probabilistically [Kam89]. Probabilities [AM80, CR86a, EC64, ME67, Wei62, ALS09]. Probability [AFR19, CC67, FH94, FH98, Gav67, GL71, HT93, OD67, She65a, She65b, Wil70, YW73, PSSZ13]. Probe [Gon81]. Probing [BBS90, MY80, PK87]. Problem [AF67, AAM22, AHU76, AMOT90, BGP84, BMPT91, BBMN15, BJ66, BMT92, Bel62, BH74, BBE^{+}21, BCS92, BEF^{+}21, BN62, BM84, Bre70, BC73, Can62, CG70a, CF12, COW16, CD60, Coo66, CNP85, CH84, Cur63b, CLL^{+}21, DDFS14, Den58, Di 69b, Di 73, Dil89, DST80, DM62, DG57, ER80a, Ehr58, Ell72, EG82, ET76, ES81, FOS14, Fei84, FLO^{+}25, FN93, Fre83, Fri63a, GM79, GG78, Gav83, GS57, Gil63, Giv57, GT88b, GHS75, GYY80, GF80, Gri68, GHLL22, GI79, GI81, Gus83, Har76, HRS88, HW76, Hir77, HP59, Hoo66a, HS74, Jaz81, Joh71, Kan80, KL86, KW85, KIM81, KR77, KS75, Kur62, LSP82, Lam83, Lam86a, Lam86b, Lin18, LR81, LW75, MM81, Men65, Mey84, Mic76, MP91, Moc71, MS16b, Net59, Oya87, Pag74, PW93]. Problem [PB89, QH68, Rei80a, RKN75, Sah75, San69, Sch74, SM90, Shi80, Spr92, STV20, Tai79, TW61, TTK66, TS74, Tsi70, Ukk82, Val15, WF74, WC74, WC76, Wri75, Yao82, You55, Zaf69, AZ06, BFJ^{+}03, BKW03, Bul13, CLL^{+}08, DGS01, GG13a, Hal07, KS96, Kho05, ORSS12, SV03, Sei02, SS06c, SM02]. Problem-Solving [EG82, QH68]. Problems [AIK84, AC23, ABS15, AK84, BIKK18, Bak94, BK14a, BH89, Ber54, Ber63, BHR84a, BMM18, BR23, BH65, BR54, BCLM56, Chi86, CM61, CADMD21, Col58, CC95, CKV90, Cry70, Cry71, Cud70, CFG^{+}17, DJSW19, DH56, Dun65, Dun66, EC68, EK72, Ehr59, EF75, Esc60, FMS97, Fre79, FG79, GT91, GMSV93, GL71, GR63b, GC65, GH72, Gur85, GSV84, HPR15, Har57, HH63, Har69, HZ80, HL13, HM85b, HS86, HS87, HJ61, Hou58a, Hsu88, HR77, HS78a, HS78b, Hun82b, Hun84, IK75, Iba78, IL83, IRT78, Ja’83, Ja’84, KLM19, Kan83, KV15, Kir81, KS74, KNY19, Kou77, Laa58b, LM94a, LO85, LR65, LR77, LZ77, LY94, Mai78, MS18, MT85, MS11, McC03, Mey88, MS21, MTZ60, Mot94, NJ57, PY82, Ram80, RWM94, RP89, RKN75]. Problems [Rot70, SG76, Sav84, Shi69, Sla63, ST72, TNS82, Tar81a, Tar81b, Usm66, Val75, Was71, Wil91, WW18, Yan79, AS04, AR05, Aro98, ALM^{+}98, BSSV03, BK10, Bul06, CK09, GW95, GG13b, Gro07, GMS10, HLP99, Hoc01, JV01, KT02, KPR04, KSS10, LMW02, Lib01, MGLA00, Red03, Ste03, Val03]. Procedural [Ros79a]. Procedure [Bar60, BV84, Ber73, Com68, CR56, DP60, FLSY74, Fri62, Fri63b, Gin67, GH59, KT69, Kow75, LR77, Lov69a, Mad66, Mat81, Men19, Mou65b, PPV60, Rub74, Sch61, Sha70a, Sho79, von56, NOT06]. Procedures [Aki70, And66, And65b, Blu71, Cas76, GR60, Har64, HC63, Joy76, JM60, LW93, LW57, Lom62, Lov72, Mac57, Mae60, Mos58, NO80, Sla64, Smi84, Sri76, Wun67]. Process [Bar60, BN71, BJL^{+}82, Car58, FLP85, GGG^{+}17, Gra63, Smi79, BGJ10, MGLA00]. Processes [Apt83, BBK93, Bau58a, BR71, BHR84b, CDPP14, EY15, GS92b, Koc23, MM79, She59a, BP01, CSV07]. Processing [BLT93, BW91, BKS17, Blu58, Bry82, CW67, CN95, CAA74, Fox70, GHG60, GM61, HB21, HM85b, Imi91, Kle66, Lom62, McG59, NT93b, Oet57, Orm61, Pap87, Pea67, Pea68, PGL76, RS66, RP66, SL68, SS79, Tre54, GHS09]. Processor [AK84, CMT70, CG93, FMI80, Gab82, GS78b, HY75, KM72a, Kne93, Llo82, Pri76, SC80]. Processor-Sharing [CMT70]. Processors [DJ81, FK85, GJ76b, Gen78, GK84, HS76, IK77, KA88, LL78, Pea69b]. Produced [Boy89, Two63]. Producing [MO56]. Product [BKK84, CHT77, CM83, CM96, DLN73, DHM65, HM13, Jai15, McK87, SB82, Wee65, Wee67, dL89, Noe79]. Product-Form [CM83, CM96, McK87, SB82, dL89]. Production [Man67a, Mel56]. Products [Erd66, Kam89, Mor97]. Program [AGOW66, Ber63, Bra78, Bra67, BM62, BGGR23, Cha75, CB72, ER64b, Fri63a, GGV^{+}24, Gor56, Gus83, HC88, HR63b, HR63a, HOPW23, Hun82a, Hun84, Kel73a, Kel73b, Log78, Log79, Lov77, Ram88, RD81, Rut64, Sch74, SFR68, Sla63, SB68, Smi82, SH21, Sum77, Tit55, TG82, Wan80, Weg76, DNS05, DPSV08, EKL10, Vaz12]. Programmability [Por66]. Programmed [Boy57, Ros69, MS59]. Programmers [Row57]. Programming [AM94, Ano79, AV82, BNNS94, BHK59, Bel60, Bel62, BS59, Bon70, CK79, Cla79, Cla95, CC95, CB72, DEL80, EC68, ER64b, EGGI92a, EGGI92b, Fei84, FN93, Fox70, GHL25, Glu63, GMM81, GB92, GB65, Gor57, GLJT23, Gra77, Gra62, Har76, Hel89, Hol62a, Hun82a, Kan80, Kan83, KH64, KB67, LS11, LL78, LR77, Lip57, McL65, Meg84, Mer57, MH82, MTZ60, Nev70, Pap81, Rem81, Rut55, Sch61, Sri65, Sri76, SR60, SS24, Til58, TS74, VK76, Wir68a, Wir68b, YL57, Yer66, GW95, KMS98, Lai13, Mye99, SS06b, Sku01, GK59]. Programs [AP93, ADH15, AK16, ADO21, BL90, BL79, BK95, BKK14, BIZ18, BD77, Cha66a, Cha69a, Cha69b, CLS21, Deb92, FGHS25, GMSV93, GJL21, GI81, Gur85, Heh77, HL76, Hog81, IM83, IL83, Isr57, JM77, JM78, Kal88, Kob13, Kow79, Lee61, LZBY25, Man69, MP70a, Mat81, Pag74, Phu76, RV80, RS66, RS67, Ros94, SD69, SJ80, Tsi70, Weg88, Wei66, You77, vRS91, AB05, BN05, BGI^{+}12, BV04, CW96, FJWY25, SS06b]. progress [Hal98]. Projection [Pyl64]. projections [McC03, Pac10]. Projects [DRY16]. Promise [BBKO21]. Pronounceable [LY61]. Proof [Apt83, ALM^{+}98, BD94, BV84, BG15, BM84, Cha70, CG93, Dav77, FLSY74, GMW91, GL70, GHJ^{+}24, GP18a, Haj83, Kow75, LFKN92, Mau69, MH89, PPV60, Raz16, Sha88, She92, SH21, Zhu20, HN07, LMW05, MT10, OOW99, Raz03]. proof-of-compliance [LMW05]. Proof-Relevant [SH21]. Proofs [BD94, BB77, CHO^{+}22, GMW91, GKR15, Hås21, HZ92, KRR22, MV15, Nou83, RZR24, TC21, AS98, BSW01, BY96, FGL^{+}96]. Propagated [Att57]. Propagation [BCLM56, CF20, CMT20, CO17, Gra55b, MMW07]. Proper [Di 73, Fik59]. Properly [BLQT22]. Properties [ABL80, AMP73, BBE70, BKMZ15, Bol79, CLTZ18, Cha62, Coh64, DLT94, FHV92, Far63, Gin60, GZ82, GKS17, Her90, Hib62, Hoa78, Hod70, Hue80, LR72, LR78, LW93, Man69, MP70a, Mar13, Mon70, MP71a, MP71b, PS70, RS83, Ros68, Sel72, VV17, WS76, BKL^{+}06, Cap02, CHL96, DKT13, FG01, JCG97]. Property [BB58, GGR98, KBI^{+}17, KNY19, Set74, Set75, Toy87, dS58, JRS09]. Property-Directed [KBI^{+}17]. Proposed [Fis58]. Propositional [BK14b, CH91b, HS91, Raz03, Raz16, Sag80, SDPF81, SDPF87, SC85, van72, EM07]. Prospects [GV61]. Protection [San88, GO96]. Protein [Sha67a, BER06, Shi10, XB06]. protein-protein [BER06]. Protocol [Bra87, GKL24, SDS^{+}18]. Protocols [AGPV90, BT85, CJLW22, CMS89, GS21, HZ92, KS83, MS92, MS16a, TL95, Aba99, AB05, AAF^{+}01, BOG02, KKD04, MPR98, PD98]. Provable [ADG15]. Provably [ADF^{+}16, BGM99, LS95, You77]. Provenance [GUK^{+}16]. Prover [RZR25, Win82, DNS05]. Provers [de 86]. Provide [Wil87]. providing [BN01]. Proving [And81, BM75, BBEG18, Cha70, FL88, GNRS92, Hen79, HR91, KBI^{+}17, Lov68, Lov69b, Lov69a, MN24, Nev74, Ove74, Pet76, Rob63, RZR25, Sho77, Sla67, SB68, Sla72b, Sla74, Sti73, Wan93, WS76, WRC65, WRCS67, Dix73]. Proximity [BSCI^{+}23]. Prune [HPR15]. Pseudo [BB58, Boy89, BL66, Cer58, Cov60, Gre61, Gru70, HP72a, Mos54, Per56, Rot60, Smi71a, Sto64a, TRA71, Van67, ANP^{+}04, NR04]. Pseudo-Boolean [HP72a]. pseudo-circles [ANP^{+}04]. Pseudo-Computer [Per56]. Pseudo-Random [BB58, Boy89, Cer58, Cov60, Gre61, Mos54, Rot60, Smi71a, Sto64a, TRA71, Van67, NR04]. Pseudo-Runge [BL66, Gru70]. Pseudorandom [AM78b, CH86, CH88, KR93, Kro64, LP73, MV15, Tez87, Ind06, SU05, Tre01]. Pseudorandomness [DMOZ22, IMZ19]. PSPACE [JJUW11, Pla04, Sha92, She92]. Psychological [Row57]. Public [DRY16, Tre54]. publication [Hal98, Hal02]. Publications [BJP59]. pulse [DFSL14]. Punch [Ber54]. Punched [Per54, Was54]. Punched-Card [Was54]. punctuality [AFH96]. Pure [DFHM24]. Pure-Circuit [DFHM24]. Purely [Cul79, KT99, Ven87]. Purpose [BW57a, DR54, GP83, ZL58]. Pushdown [CV17, Col71a, Coo71, Ukk82, Val75, HH94]. Putnam [NOT06]. PV [HZ80]. Pyramid [Kla85].

QCSP [ZM22]. QIP [JJUW11]. QR [CR86b]. Quadratic [AM85, FGHS25, FJ60, Phu76, Rub15, CDG^{+}20, Sku01]. Quadrature [Gat83, LR79, Lyn69, Mor59, Ral59, Ric75, Two63, Wer79]. Quadric [Wei66].

Quadrilateral [KP71]. Quadtrees [Sam81]. Qualitative [BGG17, DFPP02, BT00]. Quality [ABK16, Nau83]. Quanta [PGL76]. Quantification [DP60]. Quantified [BBCP20]. Quantifier [BN23, PZ19, BR96, Bas99, KK13]. Quantifiers [Tor91]. Quantifying [RS09b]. Quantile [Igl76]. Quantiles [CKL^{+}23]. Quantitative [LZ25, RK23]. Quantized [MP71a, MP71b]. Quantum [AS04, BBC^{+}01, BN79, BEG^{+}21, BCM^{+}21, Bro82, CGL^{+}22, CCL23, CFGS22, CDW21, HM13, HZ16, LZBY25, MS16a, Reg25, Unr15, YZ24, Zha21, ANTSV02, AKS12, Hal07, HMR^{+}10, JRS09, May01, NS06a, RS00, Sho03]. Quantumness [BCM^{+}21]. Quasi [Cry67, Mon68, Zel68, RS09a]. Quasi-Euclidean [Mon68]. Quasi-Linear [Cry67]. Quasi-Newton [Zel68]. quasi-polynomial [RS09a]. Quasicyclic [Han62]. Quasilinear [Sch78]. Queries [AGK^{+}17, AHK93, BFGP20, BC81, BKK^{+}18, CM87, CLY25, CY82, ES17, Fre81, GS92a, GUK^{+}16, GLVV12, HN84, Klu88, Mar13, Sag91, SSV22, ADK06, BL00, Cha10, CC04, Chi86, CNS07, DS12, FFM08, GLS01, GKS06, HPRW96, HKM^{+}02, Kea98, SSS96]. Query [ABB^{+}17, Bab16, BKS17, BSKK^{+}16, Che17, CST^{+}18, CdDPL^{+}24, FFG02, Imi91, Klu82, KMRZS17, KUB^{+}15, Rei86, Wil91, AK98, AL08, BDLW98, BLSS03, BKW03, GKPS05, Kif98, MR10a, NV02, Van07, Yek08, tCL09]. Query-Based [KUB^{+}15]. Querying [BLR14, LMV16]. Question [Coo64, Koc69, Coo03]. Question-Answering [Coo64, Koc69]. Questions [HU68a, Koc69, Rac82, Win82]. Queue [AAI69, Adi71a, AHY73, Bab75, Cof69a, Cof69b, Del70, DJM^{+}20, FP94, Kam82, Kne93, Kob74a, Laz84, LS71, MM84, Rid76, dL89, Cha00a, LMNM03]. Queue-Dependent [Del70]. Queue-Length [dL89]. Queue-Number [DJM^{+}20]. queued [LMNM03]. Queueing [ADW68, BMT89, Bra77, CHT77, CM83, CK68, DLT94, Del70, HP72b, IW71, Kle66, KM72a, KT90, Kob74a, Kob74b, KR76, Leu93, Mae77, MC74, OM78, One75, PKT90, Ras70, TPH86, Van78, ZTX21, Noe79]. Queues [Adi69, Adi73, BCMP75, BKK84, Cho80, CI74a, FACP81, HM72, HB21, IS78, JMG93, Oma77, PTW88, RA76, SD83, Sim84, Tar72, Tho07]. Queuing [BL90, CM96, CLW95, CG86, DCKT92, ES86, GT23, Gla91, GM92, Haj83, HP90, Hof84, Lam82, McK87, MY84, RL80, RL81, RTW94, SB82, SM81, Sur83, Tow80, TW88, dL89, BKR^{+}01]. Quotients [GS63].

R [Bui77]. R. [ZY03]. Race [BS89]. Rackoff [KMS^{+}25]. Radio [CD21]. Radius [LM86]. Radix [HI59, Mac66, Mat82]. RAM [AKL^{+}23, SDS^{+}18]. Ramakrishna [Mul11]. RAMs [DR00, GO96, PMH09]. Ramsey [BKS^{+}10, KNY19, MSM85]. Random [AI16, ADH63, AP86, BYG08, BC93, BB58, BBS90, Boy89, Bsh93, Cer58, CO17, CDRS93, Cov60, CM67, Di 69a, DFK91, ER64b, FPPR22, FJ60, Fus90, GS69, GW23, GGM86, GSK59, Gre61, HD64, KK13, KMR95, LWZ21, MM65, MY79, Mey85, Mol18, Mos54, Rei87, Rot60, SNPT13, SS79, Sch88, Smi71a, Smi84, Sto64a, TRA71, TRE73, Tra63, Van67, Wes67, Zha21, ANP07, Ats04, BV04, BEP09, CGH04, CFM10, GHS09, Hoc01, KT02, NR04, Ree03, Reg09, Vem10]. Random-Access [ER64b, MY79, Sch88, Tra63]. random-sampling-based [Vem10]. Randomization [ADG15, dG89, CSV09]. Randomized [AT11, BBN12, BDB23, BGSS20, Bra87, BGRS13, Goo11, HK99, KKMO18, KKT95, KR93, KZ93, MR98, MC87, TWZ23, Asp98, AC08, BSSV03, FHS98, VH06]. Randomness [BCM^{+}21, MS16a, GUV09]. Range [AC23, Cha90a, Cha90b, CK65, Fre81, Gra63, WL85, CC04, HKM^{+}02]. Ranges [MST93]. Rank [BSLRZ14, CMS20, CW17, Lic23, Lov16, MV25, Raz13, AM07, CKL13, CGL^{+}22, FKV04, SS13]. Ranking [BAG14, YW73, AT10, ACN08]. Raster [CLRS86]. Raster-Graphics [CLRS86]. Rasterizing [Hob89]. Rate [ADG15, Bau58b, BSKK^{+}16, BEGH17, CDRS23, GS16, GX22, Ked79, KMRZS17, Cha00a, KSY14]. Rate-independent [CDRS23]. Rates [Bry82, DW64, JM60, MM86, Sta56]. Ratio [AGTG21, FE76, CL11]. Rational [Bas58, GH23, Hen56b, Jia23, Kun76, Luk57, Mae60, Mae63, Ral63, Spi61, Sto64b, BT07, Vaz12]. Reach [ADLS94]. Reachability [BEF^{+}21, CLL^{+}21, DKM^{+}18, Tho04]. Reaching [DLP^{+}86, PSL80]. Reaction [CDRS23]. Reactive [VV12, PM06]. Read [AHK93, CCF84, HV07]. Read-Once [AHK93]. Read/Write [CCF84, HV07]. Real [BLT93, CW65, Col71a, ES17, EV62a, EV62b, FMNP19, Fis65a, FMR72, Gal81, GH68, GMvN59, Hei71, Hol62a, Isr57, La 64, LS81a, LL73, MS19, Man67a, MC70, Oya87, Ros67a, Ros68, Rub55, Sch64, TW61, Ukk82, Wan74b, Yak76, von56, Bas99, KT99]. Real-Time [BLT93, Col71a, Fis65a, FMR72, GH68, Isr57, LS81a, Man67a, MC70, Oya87, Ros67a, Ros68, Rub55, Sch64, TW61, FMNP19, KT99]. Realistic [CMS89]. Realization [CEW58, Gil66b, PPV60]. Really [AH94, ES15]. reason [KR97]. Reasonable [Doy65]. Reasoning [ABK16, BNR13, BN13, FH94, FH98, GS92b, GS93, KKMO18, KJJ03, LZBY25, NB95, Opp80, Wan93, BMSS09, DFPP02, FH01, HP07]. RECAL [CG86]. Recognition [Alt62, Ble66, Chu65, CC67, Doy62, DW64, EC64, GS78a, Gla57, Gre66, Gri86, Har67, HS68, Kel61, Lyn77, MF63, McA93, O’C65, DR00]. Recognizes [Hue73, Hue74]. Recognizing [GSH89, Hsu87, Joh71]. Recompression [Jez16]. Reconciliation [ZZ17]. Reconciling [MR10b]. Reconfiguration [LS87]. Reconstructing [BKL99, Sha67a]. reconstruction [CS11]. Recontamination [LaP93]. Record [CC76b, ES76, CC76a]. Recording [Per54]. Records [GS73, MW78, Wol65]. Recoverable [DM25, TL95]. Rectangle [GL88b, MP86]. rectangular [FV02, Zwi02]. Rectilinear [Hwa79, DKP98]. Recurrence [Cra55, Cra57, KMW67, Kar94]. Recurrences [DS13, HK77, Kun76, Rou01]. recurrent [OG96]. Recursion [IW91, JM78, Sch74, Fus90, Pit06, SP07]. Recursive [AM82, BL79, Blu67, BD77, Di 69b, EY09, EY15, Gau61b, GB74, HN84, HL66, LR72, Nau89, RV80, SS63, Tra69, JR07]. Recursively [Cul79, ERR81, ER80b, GR63b, Opp80]. Reduced [Col67, DG57, Has70]. Reducibilities [Hom87]. Reducibility [GB69c, GB70, JMMW79, Lad75, Lyn75, AK10]. Reducible [CH91a, HU74, WLS85]. Reduction [AAJS17, ACPS93, Bra67, CH90, Cur62, Gin59, Gra63, GH72, HWB56, IOT19, Lio71, LR81, RW88, Wei55, Wil59b, ZL58, AFN04, BC05]. Reductions [DHKN21, Hue80, HS78a, HS78b, PS81b, RMK79, Wil87, Lib01]. Redundancy [Sch63]. redux [Via13a]. Reed [BSCI^{+}23]. References [DS78]. Refinement [AGHR17, BN23, Mad66, Mol67, Bro97, CGJ^{+}03, MMHW97]. Refinements [GLNS25]. Refining [XDB83]. Refutation [CS71]. Refutational [HR91]. Refutations [HW74]. Regards [EV57]. Regenerative [IS78, IL79]. Regime [FGYZ21]. Region [AAI86, BNR13, GSV15, MP91]. Register [BP01, BS76, CH86, CH88, HKMW66, LP73, RT69, SAG94, You58]. Register-machine [BP01]. Registers [Mow67]. Regression [AKPS24, CW17, Shr70, Sri76, Wal61]. Regret [AG17, FCMG22]. Regular [ABD^{+}21, BLR14, BE97, Brz64, BC69, CDPP23, DKRW15, FU82, GH64, Gin67, HV95, Kah95, Mye92, OF61, Sal66, ACKM09, ACM02, BCG07, BYG96, BP03a, KLSS05]. Regularity [Val75]. Reinforcement [Cun85, ZCLD25]. Relabeling [ST72]. Related [ABS15, BC73, CM85, Di 73, Erd66, GR62, KLM19, Mot94, Val75, Wee65, Wee67, WC74, Fre99, Hoc01, SSS96]. Relating [BSLRZ14, DA68]. Relational [AV88, AV89, AVV97, ANV13, BK87, BDLW98, BL00, BC81, CP84, Fag83, Gal82, GH86, GPV89, IL84, Klu82, Mai80, MH89, Men79, Rei86, SU82, SY80, SDPF81, SDPF87, AL05, GE01, WR03]. Relations [AK25, BDFS84, Col70, FHS13, Gil66b, Gla65, Har74, HU68b, Iba77, KMW67, Kar94, MW86, McA93, NB95, PF79, SH21, BG98, BLSS03, KMP00, KJJ03]. Relationship [Bea82, BW94, LBB67]. Relationships [Dav90, WM74, KT02]. Relative [Ber75, CKL^{+}23, KMR95, LF88, Sup83, JRS09]. Relativizations [XDB83]. Relativized [Ko90, Rac82]. Relativizing [LS86]. Relaxation [JM60, JV01]. Relaxations [CLRS16, Sku01]. Relaxed [SVN^{+}13]. Relaxed-Memory [SVN^{+}13]. relaxing [AFH96]. Release [Mar82, Unr15]. Relevance [MK60, YLC76]. Relevant [SH21]. Reliability [DI87, Had88, Jac54, WB97]. Reliable [AAF^{+}94, FLMS93, BPC^{+}12, CT96]. Remainder [Col67]. remaindering [AB03]. Remark [Sch68, Yas67]. Remarks [CM57, Mor60, Nag61, dV59, Liv60]. Remote [Boy05, Gav71]. Removing [MS16b]. Renamable [Sla67]. Renaming [AACH^{+}14, ABND^{+}90, Lew78, CR12]. Renamings [Log78, Log79]. Rent [GMPZ17]. Repairable [LM94b, dG89]. Repeated [Gra55b, LR81]. Repeating [NR94]. Repetitive [RC69]. Replacement [ADU71, MT19, OOW99]. Replacements [BHR19]. Replete [LF88]. Replicated [BDS87, Her90]. Replication [Wak69]. Report [You56, Hal98, Man03]. Reporting [Cha90a]. Reports [BT69]. Representable [Sha70b]. Representation [BK72, Das77, ERR81, ER80b, Har74, HP90, KSG68, Mat82, Muk68, Spi61, BN05]. Representations [Boo78, DST92a, DST92b, GB69c, GB70, HR63b, HR63a, MM84]. Representative [FLPS16, KW20, Lib01]. represented [BBB^{+}00]. Representing [Pfa75]. Reproducing [von66]. Require [AIK84, GS90, RW92]. Required [Spi69, Win65a, Win67]. Requirement [RS67, Yan79]. Requirements [BJL^{+}82, CE72, GM70, Har67, Kar67]. Requires [BKV25, CLRS16, MSSV24, Raz19, Lee02]. Requiring [CR56]. Resampling [HS19]. Research [Baz57, BK77b, Min67, Gra03, Hoa03]. Resequencing [BGP84, JMG93]. Residue [Kam87, Pap87]. Residues [Sho81, Wil91]. Resilience [HPZ24]. Resilient [BEGY22, BFR^{+}22]. Resistance [Cha16]. Resolution [AB70, And68a, And68b, AM20, ABD^{+}21, BE97, CS88a, DH86, Dix73, GFL87, HR78, Joy76, Lee72, Man67b, MR87, Raz04, Rei71, Rob65, Sla67, Sla70, Urq87, AAF^{+}01, BG01, GMPS00, MPR98, SW01, BSW01]. Resolve [GW85]. Resolving [BCW85]. Resource [Dal73, DJSW19, Jaf83, KIM81, Llo82, SH22, AAPS96, BNBYF^{+}01, KKD04]. Resources [TW88]. Respect [Leh57]. Response [Bro82, HB21, IS78]. responsibilities [Hal00]. restaurant [BGJ10]. restless [GMS10]. Restorable [BP23]. Restricted [PY82]. Restriction [WL85]. Restrictions [LWZ21, Mey88]. Restructuring [MP76]. Result [SW82, Wan78, SW01, WB97]. result-checking [WB97]. Resultant [Bar60, CE00]. Resultants [Col71b]. Results [Ble66, CM85, Di 73, DV16, DSS22, GJ78, Gen78, GK84, Har63b, HS87, HU69, Jai15, JS82, Kap68, Kel73a, Ko90, Lio71, Pag74, Par55, Par61, PKT90, Rei71, TM73, Van67, AAF^{+}01, Bas99, CN06, Hås01, MMHW97]. Retiring [Hou57]. Retrieval [AF77, ACDF78, Bak62, BBH^{+}87, BK77b, CC76a, CC76b, Coo64, Eli74, EF75, FP86, Gup79, HSTV14, MK60, Sal63, Sal73, Sta72, Sti61, Win65b, YLC76, YL77, YLS82, CKGS98, MSS01]. Retrieve [Mor68a]. Retry [LS88]. revealing [JMM^{+}03]. revelation [LOS02]. Revenue [GW21]. Reversal [Iba78]. Reversal-Bounded [Iba78]. reversals [HP99]. Reverse [Erd66]. Reversibility [BM70, DLT94]. Reversible [Ang82]. Reviews [Ano60, Ham56b, Ham56c, Ham56d, Ham57a, Ham57b, Ham57c, Ham57d, Ham58a, Ham58b, Ham58c, Ham58d, Ham59a, Ham59b, Ham59c, Ham59d, Ham60]. Revised [DHM65]. Revision [DPW18a]. Revisited [EH86, KMS^{+}25, Min82, Szp90, ZZ17, CGH04]. Revisited* [BCP20]. Revisiting [ADG15]. Revocable [Unr15]. Rewrite [BGK^{+}22, GNP^{+}93, HWW90]. Rewriting [BGK^{+}22, DM87, Hue80, Toy87, WZC95, BT95, TKB95]. Rewritings [BKK^{+}18]. Richardson [Fra60]. Riemannian [CGOS13]. right [MLS76]. right-context [MLS76]. rights [Hal00]. Rigid [GNRS92]. Rigorous [BFM^{+}19, KMP^{+}12]. Ring [ASW88, Fit60, FL87, Wan93]. Rings [AAHK94, LPR13, DMR08]. Risk [BAL^{+}21]. Risk-controlling [BAL^{+}21]. RNA [Sha67a]. Road [Doy61]. Robin [Ras70]. Robots [Mur55, YUKY17]. Robust [ABDH^{+}20, BEJWY22, CLMW11, DFSL14, FK15, HKM^{+}22, Jay97, MS16a, Rab94]. Robustly [ABND95, LM23]. Robustness [BFGY19, BS23, FNFSZ23, Rou15, Sur83]. Role [RW88, Sal70]. Romberg [And66]. Room [KR76]. Root [Leh74, Lov16, Mou65b, Wee60]. Root-Squaring [Wee60]. Rooted [Gre73]. Roots [BJ66, Brz67, DSS22, MM75, Mou65b]. Rosser [MNO88, DKRW15, Ros73a, Set74, Set75, Sha88, Toy87]. Rotational [Hou58b]. Rotations [PT57, Zer85]. Round [Dou59, KL70, LP15, Ras70, SD83, WXSD25, BPC^{+}12, CS99, GS22]. Round-Off [Dou59, KL70, CS99]. Round-Robin [Ras70]. Rounding [BGRS13, DRY16, GKPS06]. Roundoff [BR71, KL73]. Rounds [BYCHS17, Bra87, HSS18]. Route [Pol64]. Routine [Bas58, Lyn69, YL57]. Routines [Eld59]. Routing [BB95, BBBL95, CK95b, DEP10, GK94, GN94, GL88b, Haj83, MP86, NS80, PU89, Tow80, Upf92, ABNC^{+}96, AGH00, AFGZ05, AAF^{+}97, BOG02, BRSU97, BFU01, HW04, JSMM03, RT02]. row [FJWY25, vD95]. row-convex [vD95]. RS [ER64a]. RS-Machines [ER64a]. RSA [BF01, HN04]. Rules [GNP^{+}93, GHKL21, HWW90, Nau89, Ral59, UV88, MR10b]. Rumor [CGLP18, Hae15]. run [WB97]. run-time [WB97]. Runge [But67, But79, BL66, Car58, Cas75, Cas76, Fro61, GR60, Gru70]. Running [Got54]. Runs [GNP20, TRA71, Wei55]. Runtime [BFR^{+}22]. Runtimes [KKMO18, LZBY25]. Russians [Mye92].

S [FHO72]. S-Semigroups [FHO72]. S.P.M. [Kle66]. Saddle [JNG^{+}21]. Safe [Yan82]. Salesman [Bel62, BH74, CNP85, FN93, GG78, MTZ60, MS16b, PB89, SV19, STV20, Aro98, BFJ^{+}03]. Same [CCSS14]. Sample [ABDH^{+}20, CBDF^{+}99, GW21, MY16, RC86, Wil91, ZCLD25]. Sample-efficient [CBDF^{+}99]. Samples [BRS22]. Samplesort [FM70]. Sampling [AG17, AC24, Bay72, CdDPL^{+}24, CGL^{+}22, Dav77, FGYZ21, FM70, GJL19, RV07, ACKM09, BYG08, CMYZ12, DLT07, ŠVV09, Vem10]. Sampling-based [CGL^{+}22]. SAT [BSKK^{+}16, FGYZ21, GST16, NOT06, PPSZ05]. Satellite [Dye68]. Satisfaction [BK14a, BBKO21, BMR97, BMM18, BR23, CLRS16, GMV86, Hon82, LS81b, Mar13, BK10, Bul06, Bul13, Gro07]. Satisfiability [DV14, FK15, HZ16, Pla04, Sch78, ANP07, BFG08, FLvMV05, GW95]. Savage [DFPP02]. scaffolding [HRM02]. Scalable [KS11]. Scalar [Was81]. Scale [ABDCBH97, DK65, EKLS20, Kau71, MS18, Pea69b]. Scale-sensitive [ABDCBH97]. Scaled [FL74]. Scaling [BFGY19, GT91, Gia64, Lam82, Ske79, ZTX21, BKP07]. Scan [One75, RR64]. Scanning [Bau58b]. scenario [CSV07]. Scene [CF20]. Schaefer [BP15]. Schedule [AP25]. Scheduled [Cof69a, Cof69b]. Schedules [AC81, CT81, DVSC77, Man67a]. Scheduling [AZ05, BLT93, BBD^{+}15, BL99, CG93, DJ81, FK85, Gab82, GJ76b, GS76, GS78b, GJ80, HS87, HS76, HLS77, IK77, IKM17, KM72a, KN81, KSS75, KSS77, LL78, Leu93, LL73, Llo82, Mar82, MS11, MC70, NJ57, PTW88, PK80, Ras70, Rei68, RY91, Sah76, SC80, Sev74, She67, AFGZ05, AAF^{+}97, BNBYF^{+}01, BL04a, BGM99, CN06, KMPS09, MSU99, SSS06, Sku01]. Schema [ADLM14, GS10, Hel89, FN10, JV01]. Schemas [AMP73, GS83]. Schemata [GG74, Kel73a, Kel73b, Log78, Log79, Rut64]. Schematic [San88]. Scheme [FK16, KU80, Lie59, Upf92, YW73, Zaf69, Coh00, DR03, JM24, KLOW00, RS09a, SS06b]. Schemes [AHPW19, ABDH^{+}20, BHM11, BFMY83, BK86, BK87, CM87, CH91a, CAFS21, Edw54, Fag83, FLS18, GH73, HM85b, Hun82a, Hun84, LRY80, MY80, Mon70, MY16, RS77, Sag91, San88, Upf84, Aro98, KSS10, OR02]. Schinzel [Niv10, Pet15]. Science [Ham69, Min70, Sal72, Bro03, Val03]. Sciences [Ano68, Leb56]. Scientific [Gof71, O’C65, Sal71a]. Scott [LP14]. Search [AM78a, BM83, BM85, BC74, Bur76, DP85, Dev86, DG71, EC68, Fal62, FGLS19, FKR19, Fre82b, Fre85, Fri73, GC65, Hen70, HJ61, KZ93, KNY19, LaP93, MB85, MT93, Mey84, MM75, Mil75, Nau83, Por66, Rub74, SD69, ST85, SMC79, Tar83, Vit83, Wei72, Wil87, Wil84, dCS77, de 83, AT07, BYG08, BFG^{+}99, Drm03, FG99, KZTH05, Kor08, MP10, MR98, Ree03]. Searchers [Doy61]. Searching [AC23, Bab63, BT57, Cha90a, Cha90b, Dav90, GNP20, GMM81, Gon81, Hib62, HY77, MHG^{+}88, Sta70, AMN^{+}98, AMM09, BYG96]. Second [DH59, Dun65, Dun66, FLO83, Has70, LW63a, LW63b, Pie73, EGG00, GKS04]. Second-Order [FLO83, Pie73, EGG00, GKS04]. Secondary [GS73]. Secrecy [Aba99, AB05]. Secretary [BIKK18]. Secrets [Rab94]. Section [Via14a, LV09, VV09, Via10b]. Secure [ABF17, DDWY93, GS22, GHKL11, KOY09]. Securely [MS16a]. Security [AW08, Can20, Chi86, LS77, Rab89, Rei79, WYCF14, Aba99, AB05, HN04, LMW05, LW08, May01]. seen [Gro07]. Segment [vO93]. Segmentation [Den65, ES76, HP76, KPR04, Hoc01]. Segmented [CAA74]. Segments [CE92]. Selected [Ja’84]. Selection [CM89, DM81, FG79, GM79, GS69, JK78, Kir81, Nev70, Row57, TWS80]. Self [AM78a, Cas71, DW04, JRS^{+}14, Kra62, LR19, ST85, von66, AK10]. Self-Adjusting [ST85]. Self-Consistency [Kra62]. Self-Describing [Cas71]. Self-Organizing [AM78a]. self-reducibility [AK10]. Self-Reproducing [von66]. Self-Stabilising [LR19]. Self-Stabilizing [JRS^{+}14, DW04]. selfish [RT02]. Semantic [BFGP20, Doy61, FCB08, GK68, HR91, KP80, MR87, RV80, Sla67]. Semantic-Syntax-Directed [KP80]. Semantics [BG70, Gab16, GTWW77, GHKL21, LP14, MT21, Rey77, ST80, SN85, VK76, vRS91, vW96, Lai13]. Semaphore [Sta82]. Semi [BC81, BN71, Cas76, DL56, Fri63b, GOBS69, dS58, AT11, BB10]. Semi-Automated [GOBS69]. Semi-Automatic [DL56]. Semi-Decision [Fri63b]. Semi-Definite [dS58, AT11]. Semi-Implicit [Cas76]. Semi-Joins [BC81]. Semi-Markov [BN71]. semi-supervised [BB10]. Semialgebraic [AC23, BCL19]. semicomputable [BT95]. Semidefinite [AK16, GW95, KMS98, Sku01]. Semidegenerate [KK70]. Semigroup [Oeh63]. Semigroups [FHO72, OPSPW19]. Semilinear [Ros70a]. Semiring [BMR97]. Semiring-Based [BMR97]. Semirings [JS82, KNP^{+}24]. Sense [Tob82]. Sensing [MM24]. sensitive [ABDCBH97]. Separable [HS90, VY11]. Separated [Sti65]. Separating [Ko90, Lic23, SFM78]. Separation [BK14b, CCF84, GKR16, Har03, HT24, RT22, LW08]. Separations [ABB^{+}17, GHJ^{+}24]. Separators [ABL17, MTTV97]. September [Hou56c]. Sequence [CH84, ER80a, FM79, Gon81, MS92, RT69, TL95, TRE73, Tro84]. Sequences [AG86, AM78b, Ata94, BB58, Boy89, Bro65, Bur76, Cer58, Cha66a, Cha69a, Col67, Hab70, Ort60, Pet15, RV78, Sha67a, AGH00, BP03a, KR14, Niv10]. Sequencing [Bru76, BDF81, Hel61, Sch61]. Sequential [Apt83, Bel60, BHR84b, Cur62, Cur63b, Far63, GS69, Gil65, Gil66b, GF67, Gin59, GC65, Har63b, Hib61, Kar67, Ker71, Kle66, Liu63, Mos58, Mow70, OF61, Ran58, RP66, XZ25]. sequentially [FKL98]. Serial [Cov60, Meg83]. Serializability [Pap79b, Set82, Yan84]. Series [AS54, BK78, CBHH61, Cru76, Hen56a, HY75, Min57, SL71, Spi61, TNS82, dV59, YCW99]. Series-Parallel [SL71, TNS82, YCW99]. Served [Men65]. Server [BBMN15, DG16, BG00, KP95, SS06c]. Servers [vdLL86]. Service [BGH^{+}95, BDF81, Del70, HM72, Hof84, KM72b, KW66, MM86, PTW88, Pri76, SB82, Sev74]. Service-Time [SB82]. Serving [Leu93]. Session [HYC16]. Set [Abd71, AHPW19, Ata94, BRK74, BKST78, BBE^{+}21, CBHH61, CF59, FOS14, GW90, Hsu88, KW85, Kir81, KLP75, Lew78, MS21, Pac10, Rei72, Rit63, SLM19, SK73, Tar75, Tv84, WH73, WRC65, Zoh74, AFN04, Bul06, CeLT00, CLL^{+}08, Fei98, FCB08, GK96, GG13b]. Set-Manipulation [Ata94]. set-theoretically [FCB08]. Sets [AAJS17, AGPT16, BBH^{+}21, Bar65, BAL^{+}21, Bra67, Bro64, BCL19, CK95a, Cha69b, CH91b, Che17, CS83, Cul79, Di 69a, ERR81, ES17, GNP^{+}93, Gin61b, GH64, GH86, HW74, KW20, Lyn75, Mar76, Mat82, MP66b, PS81b, Ros70a, Ros94, RW23, Sch68, Sla72b, WLS85, dV71, AT07, KSW09, SSS96, WR03, Zwi02]. Settling [CDT09, CST^{+}18, ZCLD25]. Several [Ber69, KSS75, KSS77, NT93b, She59a, Wil62]. Shape [CDOY11]. Share [LTV96, UW87, BS59, BB59, DK59, GK59, MS59, She59b]. Shared [AAD^{+}93, BJL^{+}82, CK68, ES76, EGM16, KU88, KS83, KM72a, KN81, Kne93, KW66, Ras70, Tau22, AGMT95, BF01, Cap02, DHW97, FKL98, HV07, JCT98, Kle67b, SN04]. Shared-Memory [EGM16, Cap02]. Shares [KPW18]. Sharing [Adi69, AAI69, Adi71a, Adi71b, AGOW66, ABND95, Bab75, BN71, Cof68, CMT70, FMI80, Fra69, HP72b, KM72a, NBC72, Rab94, She67, WYCF14, GKPR07, RS09b]. Sharp [Pet15]. Shellability [GPP^{+}19]. Shellsort [Goo11, JLV00]. Shift [CH86, CH88, LP73, Mow67, RT69, You58]. Shift-Register [RT69]. Shifting [BE62, BSP82, Jen73, PV76]. Shop [CT81, GS76, KS75, MS11, Röc84]. Short [BSW01, BEGY22, CDW21, ES15]. Short-Circuit [BEGY22]. Shortest [ADF^{+}16, AMOT90, BMPT91, BNWN25, BBE^{+}21, BLTY94, BP23, CGS15, DP22, Elk20b, Fre91, GYY80, HKN18, Joh73, Joh77, MP91, OR90, OV23, RS94b, SR94, Wag76, Wan23, WW73, Yen72, AHPSV97, AMS05, BS88, Coh00, DL07, DI04, Kho05, PSSZ13, Tho99, Zwi02]. Shortest-Path [OR90]. Shot [HHPW14]. Should [Yao81b]. Showing [Dil89, Jaz81]. Shrink [HHHN14]. Shrinkage [IMZ19]. Shuffles [RVW18]. Shunning [BCP20]. side [Gro07, XB06]. side-chain [XB06]. Sieving [Wun67]. Sifting [Mac66]. Sign [BM92]. Signaling [KRR22]. Signals [BGG17]. Signature [MMR15]. Signature-Free [MMR15]. signed [HP99]. Significant [Kan65, KMP^{+}12, Man79b]. Silence [GM20]. Silver [AP25, CDE^{+}00]. Similar [AC81]. Similarity [Doy62, Wil59b]. Similarity-Invariant [Doy62]. Simple [AP93, ACW91, BILW20, Boo78, CK79, CMS89, DW99, EF75, Elk20a, FCMG22, FLO83, Goo11, GI81, GS16, Hae15, Hib63, Jez16, JM77, Kna75, Koc69, KY83, LL85, Mau69, MP72, Par55, PT12, Rei87, Rid76, SU05, Smi71b, SDS^{+}18, Tsi70, Wei72, BSW01, SW97]. Simpler [Dil89, Jaz81, OV20, GKPR07]. Simplex [AM85, DHM65, Mul78, ON68, ST04]. simplicial [DG98]. Simplicity [Cha69b]. Simplification [Bre70, Cav70, Hir58, IL83]. Simplified [Lov69a, She92]. Simplifiers [Sla74]. Simplify [DNS05]. Simplifying [BN01, Chu61]. Simpson [Lyn69]. Simulated [SH88, ŠVV09]. Simulating [BKS^{+}10, BR91, CI74a, CI74b, Igl76, Les58, McL65, NT93a]. Simulation [ADO91, Bay72, BB61, Bry91, FMR72, FM79, GM61, HA56, IS78, IL79, Isr57, Jaf83, Mos58, PM71, Rub55, Wad56, Zei72, GO96]. Simulations [HS66, KA88, LS81a, Nic93]. Simulator [Cur65]. Simultaneous [CBHH61, Rei87]. Single [BLP09, BE62, Bai66, BNWN25, BCM^{+}21, Bru76, BJL^{+}82, Car58, DCKT92, Elg54, Fra65, HKN18, LY25, RR64, RS94b, AGL00, BL04a, BP03b, CKL^{+}07, KRV09, Tho99]. Single-Class [DCKT92]. single-commodity [CKL^{+}07]. Single-Ended [Bai66]. Single-Exponential [RS94b]. Single-Scan [RR64]. Single-Site [LY25]. Single-Source [BNWN25, HKN18, Tho99]. Single-Step [Car58]. Single-value [BLP09]. Singleton [HS21]. singly [BAP06]. singularities [Pel98]. SINR [AEK^{+}12]. Site [LY25]. situation [PR99]. Size [BP89, CCSS14, Cha75, FE76, GL76, GLVV12, Mor62, RV87, Rid76, Rob71, Wol65, Raz09]. Size-Time [BP89]. Sizing [AG73]. Skeletons [Mon68, Mon69]. Sketching [CGSV24]. Skewing [McK78]. Skip [JRS^{+}14, JRS^{+}14]. Sleep [Leu93]. Slepian [BZ23]. Slotted [Tob82]. Slowing [Col87]. SLR [MLS76]. Small [BGKM23, BIZ18, Cla95, Hås21, RC86, KSW09, SM02]. Small-depth [Hås21]. Smaller [AM85]. Smallest [Gin58, Man75, MB83]. Smallest-Last [MB83]. Smooth [Aki70, Eda22, ES17, MP24]. Smoothed [AMR11, BR15, DFGN24, HRS24, ST04]. SMT [BBEG18]. Snapshot [CG18]. Snapshots [AAD^{+}93, AACHE15, AACHE18, EFR07]. SNOBOL [FGP64]. Sockets [BFM^{+}19]. soft [Cha00a]. Software [GO96, McL84, WB97]. Sojourn [BKK84, Kne93, McK87, MY84]. Sojourn-Time [MY84]. Solitude [AAHK94]. Solomon [BSCI^{+}23]. Solution [BJ66, Boj84, BKT67, CG70a, CG67, CM60, CM61, Col58, Cry67, Cry70, Cur63b, Dou59, EC68, Fis56, Fis58, Fra60, Fra58b, FR63, Gau59, Gau60, Gla65, Gra55b, HST76, Her74, HW67, Hoc65, HP59, HJ61, Hou58a, Kac79, Kar68, Kro67, Laa58a, Lan59, McC73, MR59, MR60, MR62, MP59, Mou65a, Phi62, Phu76, RG67, RKN75, SB82, Shi80, Sta56, Sto73, SW68, TW61, TW84, Two63, Usm66, Was81, You56, Zoh74, vdLL86, CR86a]. Solutions [Abd71, BGP17, BH65, CK76, DLN73, DDFS14, Ehr59, FGHS25, FGYZ21, KIM81, Laa58b, Lam86b, LV57, MO68, Osb61, Pap84, Sig67, SL71, You55, Noe79, MW84]. Solvability [AF67, Coo66, Di 73, GH64, MRR03]. Solvable [BK14a, CC95, Fri63a, HZ80, LZ77]. Solve [BC81]. Solver [San69]. Solvers [ADG15, SK78, CCM12]. Solves [Sla63]. Solving [ABV60, ADH15, Ber54, BV04, BBEG18, BB94, CHT96, CLS21, CR56, CD60, Cry71, Ehr58, EG82, FOS14, Fri73, Gat83, GC65, GSV84, Has70, Hen70, LO85, Leh61, LKK83, MS18, Mey85, Moo67, NOT06, QH68, Ram80, Rot70, SK77, Tar81a, Tho57, War57, Was71, RWM94]. Some [AIK84, Adi71b, AM70, Ble66, Bui79, Cau58, Chu61, Coh55, DH56, EHR84, Ear74, EF75, Fei03, FL74, FW74, Fre79, Gen78, GR63b, Hås01, Har64, Hib62, Hoa78, HU69, IRT78, Ja’83, JS82, Kap68, Kar67, KP90, Lio71, Lov72, Mai78, Man67b, MS74, Mit74, NJ57, Par61, PS81b, PR99, Rei72, RT88, Rob74, Rut55, ST74, She67, SD69, Sri65, SR71, Tar83, TM73, Tsa74, Ukk82, Van67, Wil71, WM74, vdSS89, PM71, Sal71a]. Sometimes [EH86, Rei86]. Sort [GH86, NV95, RV87, Sob62, VH06]. Sorted [Wal88, Yao81b]. Sorting [Beu70, BN62, Col87, CS88b, DPRS89, Flo60, Flo61a, FM70, Fri56, Goe61, GK96, Goo11, Hib62, Hib63, HI59, Hor75, Hor78, IS56, Leh57, Liv60, Mac66, Man79a, MBM89, MP75, Nag59, Nag61, NS82, NV95, Tar72, FCFM00, FG05, HP99, KLOW00, Tho07]. sorting-complexity [FCFM00]. Sound [Rei86, SAKS64]. Soundness [Bau85, TKDB24]. Source [AAJS17, AFGZ05, BNWN25, HKN18, Kam82, Lov77, Sch64, Tho99]. Source-to-Source [Lov77]. Sources [Par79, Par80, Kle99]. Space [AMM09, BG15, Con91, DEL80, DM81, DG62, ET76, GNP20, GS21, GS90, HHPW14, HSTV14, HPV77, Hor78, Ja’83, LT82, LS80, LED76, Log79, Lyn77, McC76, Ove24, PU89, Pip78, Raz19, Rei80a, Sav84, YUKY17, ATSWZ00, BSSV03, CRS01, FHS98, FLvMV05, Rei08]. Space-Bounded [Con91]. Space-Economical [McC76]. Space-Efficient [HSTV14]. Space-Time [DEL80, Log79, Sav84, AMM09]. Spacefilling [PB89]. Spaces [HHHN14, KSU19, MP71a, MP71b, Smi71b]. span [VJ00]. Spanner [AB17]. Spanners [ES15, FKRV15]. Spanning [Hwa79, KKT95, PY82, SL15, Sup83, ACK08, Cha00b, CHL01, HdLT01, PR02]. SPARKs [EFKP22]. SPARQL [GUK^{+}16]. Sparse [AHPW19, BBS86, BSV20, BDK24, CEG86, EGGI92a, EGGI92b, FKS84, Gri86, GLW70, Joh77, Kam87, LS86, Lyn75, TY79, Wag76, CE00, CFM10, DKT13]. Sparser [KN14]. Sparsification [DV14, EGIN97, AY13]. Sparsity [CW17]. Spatial [CFGS22, KKD04, BKL^{+}06]. Special [MW86, ZL58, LV09, VV09]. Specifiable [Hul84]. Specific [Pet62]. Specification [AV89, BFM^{+}19, Ehr82, GB92, Hem64, McL84, HV07]. Specifications [Lei54, BT95, BG96]. specifying [LW08]. Spectra [AK70]. Spectral [CLTZ18, Gau64, LGT14, LM86, Neg66, She59a, BLR10, KVV04]. Spectrum [BT96]. Speech [SAKS64]. Speed [BKP07, BC54, Cha69b, Den58, Fra61, GFL87, KP00, Leb56, PT57, Rub15, TWS80, AZ05]. Speedcoding [Bac54]. Speeding [Blu71, EGIN97]. Speedup [CN95, CHK82, LR79]. Sphere [CKM^{+}14, MTTV97]. Sphere-Packings [MTTV97]. Spherical [CU59]. Spin [GSV15]. Split [SS06a]. Split-ordered [SS06a]. Spontaneous [CD21]. Sporadic [BBD^{+}15]. Spreading [CGLP18, DFGN24, Hae15, ENRS00]. Square [AC55, BR71]. Squares [DLN73, Mad66, MP70b, Osb61, Ple74, Rit63, CJK^{+}06]. Squaring [Wee60]. Sri [Mul11]. st [BK09]. st-flow [BK09]. Stabbing [AC23]. Stabilising [LR19]. Stability [BW91, Cha62, CL62, CK65, Ema63, Esc60, FGL77, GGMM88, Kar68, Kro66, MR59, MR60, ÖWA91, Rah70, Ske79, Wil59a, Wil59b, AAF^{+}01, BBG13]. Stabilizability [ÖWA91]. Stabilization [Man67a]. Stabilizing [JRS^{+}14, DW04]. Stable [Bui79, Con57, CI74a, CI74b, DG71, FOS14, Gil65, GHKL21, Ham59e, Hor78, Igl76, Ind06, ILG87, Mow67, SK78, SB73]. Stack [Aho69, BL75, ESvL80, GGH67a, GGH67b, HS71, Ull69]. Stacks [Tar72]. staged [DP01]. Stages [But67]. stall [AGL00]. Standard [BH74, Got95b]. Standardized [Gor57]. standards [BOG02, Mey65]. Star [Brz67, Mey69, MS21]. Star-Free [Mey69]. start [Via10a]. Starting [Alo60]. Starts [AC24]. Starvation [Sta82]. Starvation-Free [Sta82]. State [AF67, BZ83, CKP17, Cur63b, DS92a, DS92b, Gin60, HHHN14, Hib61, Liu63, Liu69, LM94b, MH89, MM86, Tow80, Wat61, CSV09, OG96, PD98]. State-Dependent [MM86, Tow80]. Statement [Lam86b, Rem81]. Statements [BK79, Eld59]. States [BEF^{+}21, Gin58, Gin59, Gin61a, HM13, RW88, SM81, ASV06, HMR^{+}10, NS06a]. Static [AG73, Bar16, Eli74, HMSS01, TT85, YR64, BP03b, CCM12, McA02]. Stationary [Was80]. Statistical [Doy65, DW64, Dub66, FGR^{+}17, GLJT23, Hem64, Her74, HJ61, HPZ24, LBB67, MP65, PS70, SW68, YLC76, van72, BKW03, Kea98, SV03, Cha69a, MP66a]. Statistically [KMP^{+}12]. statistics [FSV06, Hal02]. Steady [LM94b]. stealing [BL99]. Steiner [AGH^{+}25, BHM11, BGRS13, Cha72, TZ25]. stencil [FV02]. Step [AP25, AACHE15, AACHE18, BKT67, Car58, KT68, Pic64a]. Step-by-Step [Pic64a]. Steps [AM85, BKS17, BN23, Wag97]. Stepsize [AP25]. Stiff [Bui79, Cas75, Cas76]. Stiffly [DG71]. Stirling [Pic64b]. Stochastic [BGG17, Bru76, CI74a, CI74b, Dad88, Doy23, EY15, Gil63, HMZ13, HY75, Hof77, Igl76, Nel87, San69, Sch72, Shi69, EY09, FPS98, GE01, MSU99, SPK13, SS06b]. Stochasticity [JNG^{+}21]. Stock [Hal99a, Hal99b]. Stopping [DRS90]. Storage [ADW68, BCW85, Bau58b, CCF88, CM74, CC76a, CC76b, CR56, DL56, Eli74, Fra69, FM70, FB75, GL71, GH68, GL88a, GM73, GM75, GS73, Gup79, GM70, Hof84, HWB56, Kar67, MW78, Men65, PRR74, PK80, RS67, Rob71, Rob74, Ros74, Ros75, Smi79, Tra63, TS81, YW73, AKMS11]. Store [Col71a, Rob71]. Stored [ER64b]. Storing [FKS84, Sin86, TY79]. Straight [GJL21, HL76, IM83, IL83, JS82, Kal88]. Straight-Line [HL76, IM83, IL83, JS82, Kal88, GJL21]. Strain [Per54]. Strain-Gage [Per54]. Strand [Her74]. Strassen [CGS15]. Strategic [GT23]. Strategies [DP85, EG82, HWW90, HR91, Joy76, Mul78, Sch72, Tay93, Wan80, BLP09, CBDF^{+}99, CC04, NS06b, SSS96]. Strategy [ADU72, Ber63, FB72, HMZ13, LS87, WRC65, AEH00]. Strategyproofness [BBEG18]. Stratification [Ros94]. stream [Ind06]. Streaming [AD17, BEJWY22, CKL^{+}23, FNFSZ23, HKM^{+}22, PMH09]. streams [CMYZ12, SGP11]. Strict [Set82]. strictness [SRM97]. String [BGK^{+}22, CP91, FGP64, Gal81, Gal95, HSTV14, LW75, Man75, RPE81, SMK18, WF74, WC76, FG99, LMW02, Mye99]. String-Matching [Gal95]. String-to-String [LW75, WF74]. Strings [HS21, BLSS03, EGG00]. Strong [BE97, BGN^{+}22, GJ78, Jai15, CHL96, SW01]. Strongest [HOPW23]. Strongly [CM93, HMZ13, OV20, Wee65, Wee67, IFF01]. Structural [Gla91, Ros68, Pit06]. Structure [AGOW66, Bav68, BDFS84, BGK^{+}22, Bot62, BM62, CB72, EV62a, EV62b, Gre65, Har63b, Hay75, Lad75, Lom62, Ram88, WYCF14, Wee62, Wel66, YZ24, Zei72, ARV08, AM09, Cha10, FN10, FG99, HS99, MT88]. Structured [BB85, DEL80, GPV89, HL76, HY77, NV02]. Structures [BHR84a, Bru76, FLO83, Fre83, GUK^{+}16, JT75, Kow79, LED76, LRY80, Oeh63, Opp80, Pfa72, Ros79b, Sta70, WS76, WL85, vO93, AACH12, BL00, Fre99, FG01, MP24, Shi10]. Structuring [Bak77, Gil68]. Students [Til58]. Studies [LR65, Sal73, van72]. Study [BM83, BF61, CvH65, DVSC77, LWS55, MF63, O’C65, Ras70, Rei79, Ros83, RKN75, YR64, JS96, LBNS09]. Sturm [Ort60]. Sub [CDG^{+}20, KMRZS17]. Sub-Polynomial [KMRZS17]. Sub-quadratic [CDG^{+}20]. subalgebras [KJJ03]. Subclass [GF80, NB95, SW82]. subclasses [DKT13]. subconstant [MR10a]. Subcubic [WW18]. subdeterminants [FJWY25]. Subdivision [MP91, CE00]. subdivision-based [CE00]. Subexponential [ABS15, BBB^{+}21, DFHT05, Yek08]. Subexpression [DST80]. Subexpressions [AJU77, PS80]. Subgradients [Eda22]. Subgraph [Ull76, Yan79]. Subgraphs [BGL^{+}22, Wil84]. Subgroups [Cha16]. Subinterval [Ric74]. Subject [Ano95b, Ano98c, LS77, O’C65]. Sublinear [Bar16, BSKK^{+}16, CGL^{+}22, CHW12, Elk20b, GS90, KSY14]. sublinear-time [KSY14]. Sublogarithmic [HSS18]. Submodular [DV16, FNFSZ23, IFF01]. Subnetworks [Dad88]. Suboptimal [Glu63]. Subpolynomial [DG16]. Subproblem [de 86]. Subquadratic [BEG^{+}21, BBP95, Val15]. Subrecursive [CB72, Di 69a]. Subresultants [BT71, Col67]. Subroutine [Hen56b, Kro73]. Subroutines [Den61, GG61]. Subsampling [Bla25]. Subscript [DP75]. Subsequence [AHU76, Hir77]. Subsequences [Mai78, Lue09]. Subset [Dem92, IK75, LO85, SW82, DLT07]. Subsets [AHPW19, Dun74]. Substantially [GS90]. Substitution [MV15, Sti73, SS82]. Substitution-Permutation [MV15]. substitutions [CHL96]. substring [LMW02]. Subsumption [GL85]. subtyping [FCB08]. Success [Ern69]. Successful [McG59]. Succinct [EFKP22, RZR25].

Succinctness [BKK^{+}18]. Suffices [Bla25]. Sufficient [BDD^{+}24, Bon70, Das75, Ern69, Esc60, Fre82b, Fre85, Ats04]. Suffix [GNP20, McC76, NB96, FCFM00, KSB06, Shi10]. Sum [CCSS14, IK75, LO85, Toy87, OR02, Chi86, DKK24, KLM19, CJK^{+}06]. Sum-of-Squares [CJK^{+}06]. Summary [DHM65]. Summation [Kar81, KH64, RTW94]. Sums [Fre82a, DLT07, TKB95]. SUP [Sho77]. SUP-INF [Sho77]. super [Raz09]. super-polynomial [Raz09]. Superdeterministic [GF80]. Superfluous [Gin59]. Superlinear [MSSV24]. Supernormal [MSW83]. Supernormal-Form [MSW83]. Superpolynomial [LST25]. Supersequences [Mai78]. Superstrings [BLTY94]. supervised [BB10]. Supervisory [BB59]. Supplement [Neu54a, Neu54b, Neu54c, Neu54d, Neu55a, Neu55b, Neu55c, Neu55d]. Support [GM80, WRC65]. Surely [AM94]. surface [DL07]. Surfaces [GR74, New65, Wei66, AMS05, BER06]. Survey [Edw54, Fra65, Min67, MMW07]. SWAC [AF58, NJ57]. Swap [Ove24]. Swapping [Cof68]. Sweeps [EF89]. Switch [DJM17, HW04]. switches [LMNM03]. Switching [Chu61, FGL77, Har66, Hay75, Huf57, KR77, Kis91, KB81, Liu63, MP75, Paw80, Pot17, Ric73b, Ric73a]. Sylvester [SS13]. Symbol [Flo61b, Wat61]. Symbolic [AAGT15, BS59, Eld59, Gra62, GLW70, Ove66, Sla63, WL57, WH73, CGJ^{+}03, PD98]. Symbols [Sho79, BP03a]. Symbolwise [BW94]. Symmetric [ADO21, Boo65, Bun71, But78, Cor63, EV62a, EV62b, GMvN59, La 64, LS24, PP65, PT57, Rei84, Zie68, von56, AT11]. Symmetric-Definite [PP65]. Symmetry [BEPS16, DLT94, DMR08]. Symposium [Ano79, Gra77]. Synchronisation [LR19]. Synchronism [DDS87]. Synchronization [Awe85, CN95, DHSS95, Doy23, HS21, HZ80, Nic93, RC69, ST87, DW04, FHS98, LLW10]. Synchronized [NT93a]. Synchronizing [LMS85]. Synchronous [AFL83, DVSC77, FL87, YUKY17]. Synchrony [DLS88]. Synonymy [LBB67]. Synonymy/Antonymy [LBB67]. Syntactic [Flo63, GS83, WH80]. Syntactical [Dem92]. Syntax [CF68, Gil66a, KT69, KP80, LS68, MG93, MM79]. Syntax-Analysis [KT69]. Syntax-Directed [CF68, LS68]. Synthesis [Bau85, BHK59, DG84, Fit60, Gil65, HL66, MC74, Per67, Smi82]. Synthesizing [CHJS15, Weg77]. System [Apt83, AS97, Bac54, Bak56, Bau56, BW57a, BN71, Blu66, Blu58, BS59, BLS92, BO56, BB59, BT69, BD77, CCF84, CG57, Cof69a, Cof69b, CR95, Del70, DK59, GMB85, Gel75, GLH57, GK59, GP18a, GM61, Hof84, HW55, KK74, KM72b, KM74, Kor83, LTT22, Lat81, LS87, Lei54, LNSW59, LY61, LS71, Lom61, Mar58, Mel56, Mey65, MS59, NBC72, OM78, Par79, Par80, Paw80, Pen65, PS70, QH68, Rah70, RG67, RY91, SM64, SC80, SK78, She59b, SAKS64, Sto73, Swa80, Swi57, Two63, UW87, Wan78, Yer66, BG96, Cor03, HHR97]. Systematically [BF61]. Systems [Adi71b, AFL83, AG73, AF77, ACDF78, ABND95, BL90, BW91, BBD^{+}15, BBG89, BHK59, BZ14, BBCP20, BGH^{+}95, BEF^{+}21, BG15, Boo82, Bra77, BB61, Bry82, BC73, CD87a, Can62, Cas76, Cha74, CN95, CHJS15, Cla79, CM64, CK68, CMT70, Coo66, CKP17, Coo64, CI74a, CI74b, CH84, DM87, Den65, DI87, DG62, ER80a, EGM16, FHV92, FK15, Fra60, Fra69, FR63, Fri56, GT23, GSV15, Gav67, GH23, GS92b, Gla91, GMW91, GS78b, GLW70, Had88, HP90, HST76, HZ80, Hol62b, Hoo66a, Hoo66b, HR63b, HR63a, Hue80, HY77, IAWK83, Igl76, Kam80, Kar68, Kle60, Kle67b, KM72a, KN81, KSS75, KSS77, LS88, Leu93, LS71, Lom73, LM94b, LFKN92, Mae77, Man67b, ME67, McC73, McL65, MNO88, MS74, Mit74, Mit72, MC70, NT93a, NT93b, NSW24]. Systems [ÖWA91, Per67, Pri76, RC70, Ras70, Ros73a, SM64, Sal66, SK80, Sla72a, SEY15, Sti65, Sur87, TT85, Toy87, TW84, Ver95, VV12, Wil77, Yan82, Yas67, dG89, AGL00, AT10, BT95, Bro97, CT96, EY09, FGY00, HV02, MS96, MRR03, TKB95]. Systolic [vdSS89].

Table [BCFC^{+}23, FKS84, Pap87, Riv78, TY79, BFG03]. Tabled [CW96]. Tables [Lar76, Mac57, Pag74, PU89, Pic64b, Pic64a, RS66, RS67, Yao81b, SS06a, WR03]. Tabulation [PT12]. Tag [CM64]. Tagged [Yel75]. Tails [ZTX21]. takes [ST04]. Taking [BC93, Hal99a, Hal99b]. Tandem [HB21]. Tape [Bab63, Blu58, Bro64, ER64a, FMR72, GS65, Gli65, Har68, HS66, HWB56, HU68b, HU69, Iba72, LS81a, Sud75, Sud78]. Tape-Bounded [HU69, Sud75]. Tapes [Gin61b, JSV97]. Tardos [HS19]. Tarjan [Deo76]. Task [BBD^{+}15, BCS92, BLS92, Bru76, FK85, HB02, KSS75, KSS77, Man67a]. Task-Scheduling [KSS75, KSS77]. Tasks [AC81, BDF81, CRR18, DJ81, GJ76b, GS90, IK77, Kon24, MC70, Sah76, SC80]. Tausworthe [TRA71, TRE73]. Tautogies [BE97]. Taxonomic [MG93]. Taylor [CBHH61]. Tchebysheff [dV59]. TCP [BFM^{+}19, DGS01, LPW02]. TCP/IP [BFM^{+}19]. Technical [BT69]. Technique [And65a, AB70, Bay72, BC74, Bri90, Den58, GM56, Jez16, KH64, PRR74, Phi62, Sob62, EGIN97]. Techniques [AM70, Bab63, BB77, CS85, ES76, GOS12, Gui78, Hof77, Hol62a, Isr57, Kle60, Lip57, RR64, Rut55, Sal63, ST73, SR60, Usm66, Niv10]. technology [Gra03]. Telescoping [Mae60]. Teletype [Pol64]. temperature [BKP07]. Template [SU82]. Temporal [AM90, AH94, BKMZ15, BMM18, BR23, CKP17, Dav17, Dea89, EH86, EKS20, ERS24, NB95, SC85, AHK02, BK10, KJJ03]. Temporary [PK80]. Tensor [DLN73, HM13, HL13, Raz13, WEM25]. Tensor-Rank [Raz13]. Term [Hue80, Svo75, Toy87, Ven87, Yu75, YSS78, YLS82, BT95, TKB95]. Terminal [Gav71, Gin58, Hib61]. Terminals [LW57, Men65]. Terminating [Hun84]. Termination [AH92a, Bri90, TKB95, UV88]. Terms [Coo71, GK68, Kar81, Spi61]. Ternary [YR64]. Test [BFM^{+}19, BCM^{+}21, CG57, Eld59, GHL^{+}87, Isr57, Kro67, LR77, Mus78, SD76, SS13]. Test-Oracle [BFM^{+}19]. Testable [KMRZS17, GS06]. Testing [BFR^{+}13, Bra67, BR54, CST^{+}18, DKT13, Glu63, GR17, GP18a, HK55, HM13, Hon82, HT74, Kam80, KNY19, Kro73, Len89, MSY81, O’C64, RD81, Set74, SSSU83, AB03, AC05, CSV07, GIS99, GGR98, GK99, Set75]. Tests [Gra55a, GSK59, Row57, Set78, UV88]. tetrahedra [ES09]. Text [AF77, BW94, BBH^{+}87, Doy65, GNP20, KR92, SL68, Sal73, BYG96, FM05]. Textual [SS82, FGMS05]. TFNP [GHJ^{+}24]. th [GLH57, HP59]. Their [BK95, CK15, CLW95, CS84, Eda22, FT87, Gau59, Gau60, GMW91, GFL87, Iba78, KBI^{+}17, MM79, Nat67, Nip90, OF61, PV76, ANP^{+}04, Har64, LR99, Niv10, Nou83, O’C64]. Them [DA68, GMPZ17]. Theorem [And81, Baa93, BDB23, BP15, BDFG19, Cha70, CC16, DEL80, DS13, GR60, GNRS92, Gre65, HRST17, Haj83, Hen79, HL76, HR91, Kar68, KY83, KG74, Lov68, Lov69b, Lov69a, MSW83, MSM85, Nev74, Ove74, Pap82, Pap85, Pet76, Rob63, Sha88, Sla67, SB68, Sla72b, Sla74, Sti73, TW88, War62, Win82, WRC65, WRCS67, de 86, Bul06, DNS05, Din07, ZY03, CGS15, Dix73]. Theorem-Proving [HR91, Lov68, Lov69b, Lov69a, Nev74, Ove74, Rob63, SB68, Sla74, Sti73, Dix73]. Theoremhood [van72]. Theorems [BB77, BM75, ESY85, Ros73a, Sla70, Smi70, CKL^{+}07, LR99, Ros08, Rou01]. Theoretic [Cha74, CM78, FHV91, GS93, GI79, MH89, AL05, FGY00, GLJT23, KVW00, NR04]. Theoretical [AM70, EK72, Kle67b, KS74, MH82, Zad72, HS87]. theoretically [FCB08]. Theories [Baa93, BÉ91, CS71, CCM12, LS95, McD82, PS81b, Sho84, Sla72b, Sla74, BG98, NOT06]. Theory [AV82, ACPS93, AGS17, BL90, BK87, Blu67, BGK^{+}22, BHR84b, BT71, Cha75, CLRS86, Cur63a, Cur63b, DP60, Ehr82, Glu63, GB92, Had88, Har66, Har69, HH71, Hol62b, HY84, IW91, KK70, Kap68, Kle66, KG74, Lam86a, LZ25, MS19, Mau66, Nel87, Pic67, Pie73, RBCC19, Ras70, RZ19, ŠW98, SV85, Svo75, VSC24, Ven87, Ver95, Wan57, WH80, Yan82, Yeh68, You69, Zei72, von66, BLR13, BKR^{+}01, DFPP02, EM07, FKP13, GJS12, MN05, Mul11, Pel98, SK96, SCD02, SN04, ZW03]. Thermocouple [Per54]. Thesis [Yao03]. Thickenability [FT22]. Thin [ES15]. Thinning [SR71]. Third [CR56, Neg66, Zaf69, Gra77]. Third-Order [CR56]. Thompson [AG17]. Thoughts [Sal71a]. threads [CHL01]. Three [Alo60, BM85, Bro03, CCSS14, CGM24, CD87b, Com68, DV95, Kei56, Red03, RS94b, Röc84, Ros83, She57b, Val03, YUKY17, Zar59, AHPSV97, BKL99, BT00, ES09]. Three-Dimensional [Com68, Ros83, YUKY17, BKL99]. Three-Machine [Röc84]. Three-Point [Alo60, Kei56]. Threesomes [GP18b]. Threshold [Mol18, She65b, WYCF14, Win71, Fei98, She65a]. Thresholded [Skl70]. Thresholding [MM24]. Throughput [BFGY19, ZTX21]. Thue [Boo82, MNO88]. tick [DFSL14]. tick-generation [DFSL14]. Tiebreaking [BP23]. Tight [AB17, AAHK94, AACH^{+}14, AC08, CeLT00, CKL^{+}07, CADMD21, CFG^{+}17, DFHM24, DKK24, DHK19, FG10, FV02, FNS21, GST16, Kla85, LT82, LLW10, AZ08, Kol04]. tightness [vD97]. Time [AIK84, Adi69, AAI69, Adi71a, Adi71b, AM94, AGOW66, ADLS94, Bab75, BLT93, BBS86, BE11, Bar16, BZ23, BSSV03, BZ14, BNWN25, Ber75, BN71, BKSW24, BP89, BLQT22, BEG^{+}21, BGP17, BKK84, BK81, BK82, BDF81, BCL19, CDG^{+}20, CW65, CKL^{+}25, CMS89, Cho80, CW17, Cof68, CK68, CMT70, CAFS21, CM93, CLS21, Col71a, CR95, Coo71, CKV90, CDW21, Dav17, DEL80, Del70, DV14, DM81, DGG^{+}25, DP14, DFK91, Elk20b, Elk20a, EFR07, EH86, FL88, Fis65a, FMR72, Flo60, Flo64, FLS18, FLvMV05, Fra69, FKS84, GSH89, GS78a, Gal81, Gal82, Gal95, GNP^{+}93, GH68, GS93, GS76, GL88b, GW85, GR79, GS73, GS90, HS91, Hal98, HPR15, HB21, HvdH22, HP72b, HKN18, HW76, Hol62a, HU68b, HPV77, Hor78, Isr57, Ja’83, Kam87, Kan83, KL86]. Time [KKT95, KT19, KS16, Kle67b, KM72a, KN81, Kne93, Knu61, Ko90, KW66, Lad75, LT82, LS81a, Lic23, LS24, LED76, LS77, LL73, Log79, LB79, Man67a, Man75, Mát90, McK87, Meg84, MY84, MMR15, MC70, NBC72, NB96, OR90, Oya87, Paw80, Pip78, Pri76, RA20, Ras70, Raz19, RV87, RS66, Rei80a, Rid76, Ros67a, Ros68, Rub55, SM64, SB82, SH88, Sav72, Sav84, SS79, SD83, Sch64, SFM78, Sev74, She67, Spe97, Spi69, SEY15, Sti65, TNS82, Tob82, TW61, TSOA80, Ukk82, Val15, Wag97, Win65a, Win67, vW96, AFN04, AGL00, AHK02, AZ05, Aro98, AMM09, BL04a, BP11, BM11, Cha01, CFM10, CRTY23, Coh00, FGMS05, FMNP19, Gro12, Hal07, HK99, HW04, JSV04, JM24, KP03, KT99, Kar00, KSY14, Kor08]. time [KSS10, KVW00, OR02, PPSZ05, PSSZ13, RS09a, SSS06, ST04, Ste03, Tho99, WB97, Hal02]. Time-Bounded [Coo71]. Time-Dependent [OR90, Rid76]. Time-Domain [Sti65]. Time-Shared [CK68, KN81, KW66, Ras70, Kle67b]. Time-Sharing [Adi69, AAI69, Adi71a, Adi71b, AGOW66, Bab75, BN71, Cof68, Fra69, HP72b, NBC72, She67]. Time-Space [Ja’83, LT82, Pip78, Raz19, Rei80a, BSSV03, FLvMV05]. time-varying [AZ05]. Time-work [Spe97]. Timed [ACM02, FMNP19, Par79, Par80, Unr15]. Timed-Release [Unr15]. Times [AC81, BDF81, CAA74, IS78, KM74, KW66, Mar82, RKN75, SC80]. Timestamp [HHPW14, HV02]. timestamping [DW99]. Timing [ADLS94]. Toeplitz [HC72, Wat73, Zoh69, Zoh74]. Tolerance [CYZG14, Rab89, BN01]. Tolerant [DHSS95, DI87, LS88, BKW03, BGM^{+}98, DFSL14, JCT98, Kea98]. Tools [ACDF78, FL88, KW20]. Top [CF68, HSTV14, SMK18, UV88]. Top- [HSTV14]. Top-Down [CF68, SMK18, UV88]. topic [BGJ10]. Topics [Ear74]. Topological [BBFD23, Cry71, MP71a, MP71b, Nat67, NSW24, BKL^{+}06, HS99]. Topologies [Das77]. Topology [Cha72, KLPP15, TM73, CR12]. Total [CADK^{+}24, Col70, HKN18, LW57, NW73, Sev74, BL04a]. trace [AT11]. traceable [DW99]. traceback [Adl05]. Traced [BIM95]. traceroute [ACKM09]. Tracking [AP95]. Tractability [Che17, McA93, GKS04]. Tractable [BKTW22, Mar13, NB95, CJJK00, GMS09, KJJ03]. Trade [Adl05, AGPV90, Dal73, DEL80, DI05, HMR83, Ja’83, LT82, PU89, Pip78, Rei80a, Sav84, BPC^{+}12, BSSV03]. Trade-Off [AGPV90, Dal73, PU89, Pip78, Rei80a, BPC^{+}12, BSSV03]. Trade-Offs [DEL80, Ja’83, Sav84, Adl05, DI05, HMR83, LT82]. Tradeoff [Log79]. Tradeoffs [Raz16, AMM09, Spe97]. Trailing [BFGL79, FG76]. Trajectories [LWS55]. Transaction [AV89, BBG89, Rou24]. Transaction-Based [AV89]. Transactional [AGHR17]. Transactions [AV88, BDFG19].

Transcendental [Lan59]. Transcendentals [Dav77, RD81]. Transcode [HW55]. Transcription [Bla59]. Transducers [AD17, SMK18, ANV13]. Transducing [KR14]. Transduction [LS68]. Transfer [Abe80, GK58, HA56, JP84, Sal60]. Transferability [AGK^{+}17]. Transfinite [HR91]. Transform [Bas58, DA68, KB81, LMN93, Mor73, NBC72, Pap79a, Pea68, Tit55, LZ01, Man01]. Transformation [BMT92, BH74, BD77, DM25, Lov77, VVAG97, Wan80, Wee62]. Transformations [Har86, Wil59b]. Transformed [But79]. Transformer [She65a, She65b]. Transformers [Hoa78]. Transforming [HP99, MLS76]. Transforms [Cru76, DA68, KN14, KL70, Wee66]. Transient [RW88]. Transition [Bav68, Lom73, OD67, BG96]. Transition-Preserving [Bav68]. Transitions [Jun03]. Transitive [EPL72, TS10, BG98, DI05]. Translated [Yer65]. Translating [AGG62, Got95b]. Translation [BM62, Bro58, DK59, EKS20, Lyn77, Mey65, RR64, SFR68, Weg69, Zar59]. Translations [KP80]. Translator [Gra62]. Translator-Oriented [Gra62]. Transmission [DDWY93, MS92, TL95, BPC^{+}12]. Transmissions [CD21]. Transportation [Den58, DG57, Har76, Ram80, RKN75, ST73, TW61]. Transposing [Ber58, WSD59]. Trapdoor [BM92]. Trapezoids [AAI86]. Traveling [BH74, CNP85, FN93, GG78, MTZ60, MS16b, STV20, Aro98, BFJ^{+}03]. Travelling [Bel62, PB89]. Traversal [HP76]. Traversing [Pap76]. Treatment [Bel62, Cra54, Kle67b, BT00, BM96]. Tree [AD17, BSP82, BGRS13, CH90, CADK^{+}24, Cur61, Cur63a, CMNP18, Dav90, DGM94, FM70, FL18, GSV15, Gav83, GS83, HP76, HR91, HY77, Hwa79, McC76, MSM85, PY82, PS81a, Ros73a, SMK18, Sta70, Tai79, TZ25, ACK08, Cha00b, CT06, FCFM00, FG99, FFG02, FG01, HdLT01, PR02, Ree03, Shi10, TS10]. tree-decomposable [FG01]. tree-decompositions [FFG02]. Tree-Manipulation [Ros73a]. Tree-Searching [Dav90]. Tree-to-String [SMK18]. Tree-to-Tree [Tai79]. Trees [AJ76, AM78a, ADK17, BC93, BLQT22, Bur76, Cha72, Dev86, FK16, GNP20, GJ80, Got95a, Gre73, GSV84, Har86, Hen85, Hib62, HO82, JSS16, KLM19, KR89, KKT95, KW89, MT85, Nau83, NW73, Pfa75, Pro80, RR92, Ros79b, RP89, RV78, San69, SL15, SD69, ST85, Smi84, SF80, Spr92, ST72, Sta72, Sup83, Tar79, Tar83, Weg88, Wil87, Yan85, Yi12, Zer85, ZZ17, vO93, AT07, BMSS09, CHL01, Drm03, FLMM09, GKS06, KZ98, MR98, Van07]. Treewidth [BHM11, GLVV12]. Trench [Zoh69]. Triangle [CPSZ21, WW18]. Triangles [GP18b]. Triangular [Wil59b]. Triangularization [Hou58c, Mor60]. Triangulation [KSW09, MR08, RS09a]. Triangulations [Rub15, RW23, BM11]. Tridiagonal [HST76, Ort60, Sto73]. Tridiagonalization [Hou58b]. Trie [CS77]. Tries [Szp90, BYG96]. Trinomial [LV57]. Trip [SD83]. Triple [Elg54, von62]. Triple-Diagonal [von62]. Trivalent [GHL^{+}87]. True [BC14, GOS13]. Truly [BEG^{+}21, CDG^{+}20]. Truncation [Gra55b, Laa58b, Wee60]. trust [LMW05]. Truth [LOS02, Mie57, Nat67, Net59, Pra60]. Truthful [BKS15, DV16, Kon24, LS11]. Tseitin [Hås21]. Tshebysheff [Min57]. TSP [AKS15, KLSS05, TV19]. Tuned [Hob93]. Tunnel [HWB56]. Tuple [AG86]. Turing [Yao03, Arb61, Cas71, Cur65, Fis65b, Har68, HS66, HU69, Hun84, Kar67, LR65, RW63, Wan57, Wat61]. Turn [GN94, HMZ13]. Turn-Based [HMZ13]. turnip [HP99]. Twin [BKTW22, BGdM^{+}24]. Twin-Width [BGdM^{+}24, BKTW22]. Two [AM85, BEF^{+}21, BMSS09, BKT67, BS85, BL66, CCF84, Cal61, CD87b, Cof68, CG93, CH79, CP91, DW64, Fei84, Gab82, GJ76b, GS22, GR62, GS65, GH70, GI82, HP72b, HS66, Hib61, HW67, HW76, HR63b, HR63a, ILP97, Ita78, JSV97, Kan80, KS66, KSG68, La 64, Lee80, Lip57, MR10a, Par61, PS78, Ram80, Rei71, Sal66, Shi80, Wag97, Wes67, Win82, de 86, vdLL86, BS88, CDT09, FJWY25, SS06c, GHKL11]. Two-Commodity [Ita78, Ram80]. Two-Dimensional [BEF^{+}21, Cal61, Lee80]. Two-Level [HR63b, HR63a]. Two-line [Par61]. Two-Party [GHKL11]. Two-Phase [BS85, ILP97]. two-player [CDT09]. Two-Point [HW67]. Two-Processor [CG93, Gab82]. Two-query [MR10a]. Two-round [GS22]. two-server [SS06c]. Two-Tape [GS65, HS66]. Two-Variable [Fei84, Kan80, BMSS09]. Two-Way [CP91, GH70, GI82, KS66]. Typability [KTU94]. Type [AGS17, Bak56, BBR54, Bau58a, Bon70, FLO83, Fro61, HK54, KV15, Pie73, RBCC19, SFLB^{+}25, TKDB24, Tra66, BT07, Cha00b, Che19, ORSS12, SP07, Tre54]. Types [Boo82, Ehr82, Gin61b, GW23, GR20, HYC16, KU80, SH21, Wil87, AB05, FCB08]. typing [Aba99, GKPS05].

U.S.S.R. [Leb56]. Ultimate [PP65, RC69]. Ultimate-Definite [PP65]. Ultra [CGT10]. Ultra-low-dimensional [CGT10]. Ultraconcurrent [Rem81]. Unambiguity [GU66b]. Unambiguous [KT69]. Unary [CKV90]. Unbalanced [GUV09]. unbordered [HN07]. Unbounded [BKK14]. Uncertainty [ADLS94, FHS13, BLMS97]. Unconditional [May01]. Uncoupled [CS85, FCMG22]. Undecidability [BK14b, Har86, JS08, Wan74b]. Undecidable [GMSV93, Har69, HU68a]. understand [Lam03]. Understanding [LPW02, TWZ23]. Undirected [DLW25, DHK19, Gib69, HKN18, Rei08, Shi80, Tho99, Wel66, AZ06, ATSWZ00, Coh00]. undominated [BLP09]. Unequal [AM80, PGE75, AM80]. Unification [Baa93, GNRS92, Jaf90, Jez19, Nip90, Sti81, Wal88]. Unified [BS89, EKS20, HS86, Kir81, Tar81b, BNBYF^{+}01, KMPS09, SN04]. Uniform [All89, ACJR21, CM67, DSS22, Gin58, GS78b, GJL19, GPV89, KMW67, Lar83, MM65, SC80, TC21, Wag97, Wes67, Yao85, ABDCBH97]. Unifying [CRR18, Lov72]. Unimodal [New65]. Union [SY80, Tar75, Tv84, vO93, ES09, FCB08]. Union-Copy [vO93]. unions [ADK06, DS12]. Uniprocessor [BBD^{+}15]. Unique [ABS15, KV15, Pap84]. Uniqueness [PKT90]. Unit [Ano68, BBB^{+}21, Cha70, HW74, LM86, PGL76, Rub15]. Unitary [Hou58c, Mor60]. Units [Bli72, FMR72, FB75, LS81a]. Univalence [TTS21]. Univariate [CD87a]. Universal [ADK17, AAF^{+}01, BZ23, BB95, Baz57, BS23, CF68, FK16, Gol82, Gor57, KBI^{+}17, KPP^{+}15, Smi71b, Wat61, AD01, YCW99]. Universal-stability [AAF^{+}01]. Universality [CM64, MS21]. Universally [Can20]. Unknown [YZJ18, DKP98, HB02]. Unknowns [Mou65a]. unless [DV14]. Unnormalized [AM59, Ash64]. Unrecognizable [MP66b]. Unrelated [DJ81, LL78, KMPS09]. Unreliable [AAF^{+}94, CT96, Pip89]. Unrestricted [Men19]. unsatisfiability [Ats04]. Unseen [VV17]. Unsolvability [BM84, CS68, Di 69b, Gre66, Gri68]. Unsolvable [GR63b]. unstructured [ABF^{+}02]. Unsynchronised [WSY^{+}19]. Untrusted [GGV^{+}24, MS16a]. Up-and-Down [TRA71]. Up-to- [GW21]. Update [BKSW24, Hal02, HKN18, Smi79, AS13]. updated [GE01]. Updates [CP84, Pap79b, Set82]. Updating [DHM65, Sin86]. Upon [Bak62, Smi70]. Upper [Che19, CHK82, Hib61, Kis91, KMW16, NW73, RS94b, SEY15, CR12, Kol04]. Upward [DHK^{+}17]. URAL [Baz57]. Usage [Low68]. Use [AF58, AACHE15, AACHE18, Beu70, Bot62, Cur63b, Gra63, HW67, Huf57, LS88, MF63, Por66, RD81, Was54, CBFH^{+}97, DW99, FV02, LR99]. Used [Gra55a, Rah70, She65a, She65b, BY96]. Useful [Doy62]. Useless [Set82]. User [Fra61, AZ05]. user-dependent [AZ05]. Users [AP95]. Uses [DR54, FT87]. Using [AGPT16, AGHR17, BB77, BC81, BKT67, BJL^{+}82, Chu61, Cru76, Cry70, DLN73, Dea89, Dub66, GGS89, Gab06, GNRS92, Gau61a, Gla65, Glu63, GM56, GX22, HS87, Hoc65, HWB56, KT68, Kow75, LY61, Mac55, MS16a, Mon68, Mou65b, Oli67, Pea67, PGE75, PT57, Pul55, Pyl64, Sal63, Sam81, SD76, Sev74, She14, Tar72, Ver95, Was80, Wee66, Win82, Yen72, YSS78, YLS82, Zer85, dG89, FKL98, GW95, GK99, HV02, JV01, JMM^{+}03, KOY09, KRV09, KSW09, PD98, PMH09, Zwi02]. USSR [GV61]. Usually [GW76, ST04]. utilities [VY11]. Utility [CM78, Tre54]. Utility-Theoretic [CM78]. Utilization [GV61, Pri76].

Validation [BFM^{+}19]. Validity [GMW91]. Valuations [DV16]. Value [Ber54, BH65, CM61, Col58, Cry71, Dad88, DH56, Ehr59, Giv57, KK70, PKT90, RL80, RK23, TSG85, Usm66, Was71, Zaf69, BLP09, DJKK08, ST80, RL81]. Value-Vector [Giv57]. Valued [TZ16, KZ13]. Values [ABF17, Fik59, Pet62, Rei86]. Vanish [MR93]. Vapnik [BEHW89]. Vardy [GUV09]. Variable [BJL^{+}82, Cur62, EV62a, EV62b, Fei84, Fra65, Kan80, KU80, Liu63, TS74, BMSS09]. Variables [And66, Ber69, DG62, FJ60, HV95, HW76, HL76, KSG68, LTV96, OT95, Wil62, HV07]. Variance [Bay72, Hem64, Lie59]. Variant [Wan57]. variants [GMS09]. Variations [DST80]. Varieties [Nip90]. Various [Ben80, Wol65]. varying [AZ05]. VASS [KMS^{+}25]. VC [MY16]. Veblen [Kle60]. Veblen-Wedderburn [Kle60]. Vector [BBE^{+}21, BEF^{+}21, Giv57, OD67, Str83, BOG02, FPS98, HV02, Kho05, Mye99]. Vectors [BKST78, CDW21, KLP75, MRR03]. Vegas [Cla95, GHL^{+}87, LPW02]. Venture [Mel56]. Verifiable [YZ24]. Verification [AAHK94, AAGT15, BFR^{+}22, Bra78, Bry91, CY95, EGM16, GGV^{+}24, Sch80, SJ80, Win82, ALM^{+}98, BOG02, PD98]. Verified [SVN^{+}13]. Verifiers [DS92a, DS92b]. Verifying [BL19, CKP17, Kam89, Weg76, Hoa03]. Versus [AFL83, HPV77, AG87, BAG92, Coo03, EH86, Ste03, vD97]. Vertex [BYCHS17, BE11, LNP^{+}25, ME67, TW88, WLS85, CLL^{+}08, Gab06]. Vertex-Allocation [TW88]. vertical [Kol04]. Vertices [KW20, PS78]. Very [Lie59]. VI [Igl76]. Via [CKN18, AB03, AC24, ABDH^{+}20, BKK^{+}18, BLR10, BBEG18, BS23, BGRS13, CPSZ21, CF12, DPSV08, DRY16, ENRS00, FOS14, FFG02, FGLS19, GS92a, GKPR07, HKM^{+}22, Hof84, HPZ24, KB81, LS11, MS18, MMHW97, OV23, RPE81, SS82, VK96, WB97]. Vibration [AC55]. View [FL18, Ham69, Lov72, Yer65]. Viewpoint [Mit72]. Views [CP84, Wei66]. virtual [AAF^{+}97]. Visibility [ACW91]. VISION [Wei66]. Visits [EF89]. Visual [Hue73, Hue74]. VLSI [CM85, CS88b, CDMP92, GE84, HM85b, JP84, Ja’84, Kis91, Ram86, Ros83]. Volterra [Lin69]. Volume [DFK91, LZ01]. Volumes [CH79, You63]. Voronoi [Lee80]. Vortex [TW24]. Votes [GMB85]. Voting [BDS87, CRWW24, HHR97]. VPN [GOS13]. vs [CG93, Cra55, Elg54, KNY19, Mul11].

W. [SW01, Zoh69]. Wait [ACHS16, ALS94, LTV96, Röc84, Jay97, JCT98]. Wait-Free [ACHS16, ALS94, LTV96, Jay97, JCT98]. Waiting [CMT70, Flo64, KM74, KR76, Paw80, TSG85]. Waiting-Time [Flo64]. walking [TS10]. Walks [AI16, BC93, CDRS93, SNPT13, BV04]. Warm [AC24]. Way [BHT18, CIV88, CP91, FNFSZ23, GGH67a, GH68, GH70, GI82, HZ92, HS71, KS66, LP15, SKM18]. Weak [ADU72, AKN^{+}08, BCL19, GYY80, Lam83, Rub22, Sti73, CHL96, KOY09, Raz04]. Weakest [CHT96, KKMO18, LZBY25]. Weather [CW67]. Web [GK04]. websites [CM04]. Wedderburn [Kle60]. Weight [BNWN25, Das75, DDFS14, DP14, Gil63, MR08, RS09a]. Weighted [BBN12, BDS87, CDRS93, FOS14, FHTZ22, GK88, MP91, TZ25, You67, AMS05]. Weighting [YS76, YLS82]. Weights [CC17, Tho99]. Weisfeiler [BN23, GLNS25, KPS19]. Well [vRS91, Van07]. well-definedness [Van07]. Well-Founded [vRS91]. Westwater [Her74]. Wheeler [Man01]. Which [Cla79, ET76, GMPZ17, Gin58, GR63a, Hue71, Hue73, Hue74, Ral59]. while [Ram88]. White [Kla85, KNY19]. White-Box [KNY19]. Whole [Koc23]. Whole-grain [Koc23]. Whose [AM85]. Widely [Sti65, BY96]. Width [BGdM^{+}24, BIZ18, Hob89, BKTW22]. Wilf [Ema63]. Wind [HWB56]. Wire [LW57]. Wireless [AEK^{+}12, KLPP15, AZ05]. Wires [Ram86]. Wise [BR91]. Within [CDG^{+}20, SMC79, Wag97, CADK^{+}24, PW93, SL15]. without [AKMS11, ADH15, BLT93, Men15, NR94, Pel98, YZ24]. Witness [RZR24]. Witnessed [LS24]. Wolf [BZ23]. Word [FMS97, Gau61a, Jaf90, Jez16, LZ77, MF63, FSV06, KMP00, Pla04]. Words [BF61, KS63, LBB67, MS21, PGE75, PZ19, RW63, AM09, HN07]. Work [BK95, GS90, KLM^{+}97, Sav72, BL99, Coh00, KSB06, Spe97]. Work-Preserving [KLM^{+}97]. World [CFGS22, Yak76]. wormhole [JSMM03]. Worst [FKS84, GW85, NPRR18, Tv84]. Worst-Case [Tv84, NPRR18]. Write [CCF84, Smi79, HV07]. Writer [HV95, EFR07]. WRITES [Rei87].

XML [AW08, ADLM14, AL05, AL08, BLPS10, BMSS09, FL02, FK16, GKPS05]. XOR [DKK24]. XPath [BFG08, BP11, FFM08, GKPS05, MS04, tCL09, TS10]. XrML [HW08].

Yaw [ZL58]. years [McC03]. Yen [WW73]. Yield [GMW91]. Young [Til58].

Zero [CFGS22, GMW91, GOS12, Hei71, Joh71, Ros71, BBP95, DNS04, SV03, DS92b]. Zero-Knowledge [GMW91, GOS12, BBP95, DNS04]. Zeros [COW23, Fra58a, Klo61, RS63, Smi70, Wan74b, Wil78]. Zvegintzov [Yer65].

References

Applebaum:2017:AC

[AAB17] Benny Applebaum, Jonathan Avron, and Chris Brzuska. Arithmetic cryptography. Journal of the ACM, 64(2):10:1–10:??, June 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Aspnes:2012:PCD

[AACH12] James Aspnes, Hagit Attiya, and Keren Censor-Hillel. Polylogarithmic concurrent data structures from monotone circuits. Journal of the ACM, 59(1):2:1–2:24, February 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Alistarh:2014:TBA

[AACH^{+}14] Dan Alistarh, James Aspnes, Keren Censor-Hillel, Seth Gilbert, and Rachid Guerraoui. Tight bounds for asynchronous renaming. Journal of the ACM, 61(3):18:1–18:??, May 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Aspnes:2015:LUA

[AACHE15] James Aspnes, Hagit Attiya, Keren Censor-Hillel, and Faith Ellen. Limited-use atomic snapshots with polylogarithmic step complexity. Journal of the ACM, 62(1):3:1–3:??, February 2015.

CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Aspnes:2018:ELU

[AACHE18] James Aspnes, Hagit Attiya, Keren Censor-Hillel, and Faith Ellen. Erratum: Limited-use atomic snapshots with polylogarithmic step complexity. Journal of the ACM, 65(6):38:1– 38:??, November 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3231592.

Afek:1993:ASS

[AAD^{+}93] Yehuda Afek, Hagit Attiya, Danny Dolev, Eli Gafni, Michael Merritt, and Nir Shavit. Atomic snapshots of shared memory. Journal of the ACM, 40(4):873–890, September 1993. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
153741.html.

Afek:1994:RCU

[AAF^{+}94] Yehuda Afek, Hagit Attiya, Alan Fekete, Michael Fischer, Nancy Lynch, Yishay Mansour, Dai-Wei Wang, and Lenore Zuck. Reliable communication over unreliable channels. Journal of the ACM, 41(6):1267–1297, November 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/195651.html.

Aspnes:1997:LRV

[AAF^{+}97] James Aspnes, Yossi Azar, Amos Fiat, Serge Plotkin, and Orli Waarts. On-line routing of virtual circuits with applications to load balancing and machine scheduling. Journal of the ACM, 44 (3):486–504, May 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/citations/journals/jacm/1997-44-3/p486-aspnes/.

Andrews:2001:USR

[AAF^{+}01] Matthew Andrews, Baruch Awerbuch, Antonio Fernández, Tom Leighton, Zhiyong Liu, and Jon Kleinberg. Universal-stability results and performance bounds for greedy contention-resolution protocols. Journal of the ACM, 48(1):39–69, January 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL http://www.acm.org/pubs/articles/journals/
jacm/2001-48-1/p39-andrews/p39-andrews.pdf;   http:/

/www.acm.org/pubs/citations/journals/jacm/2001-48-1/
p39-andrews/.

Agrawal:2015:AVS

[AAGT15] Manindra Agrawal, S. Akshay, Blaise Genest, and P. S. Thiagarajan. Approximate verification of the symbolic dynamics of Markov chains. Journal of the ACM, 62(1):2:1–2:??, February 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Abrahamson:1994:TLB

[AAHK94] Karl Abrahamson, Andrew Adler, Lisa Higham, and David Kirkpatrick. Tight lower bounds for probabilistic solitude verification on anonymous rings. Journal of the ACM, 41(2):277–310, March 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/174656.html.

Adiri:1969:TSQ

[AAI69] I. Adiri and B. Avi-Itzhak. A time-sharing queue with a finite number of customers. Journal of the ACM, 16(2):315–323, April 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Asano:1986:PPR

[AAI86] Takao Asano, Tetsuo Asano, and Hiroshi Imai. Partitioning a polygonal region into trapezoids. Journal of the ACM, 33(2):290– 312, April 1986. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/5387.html.

Abdulla:2017:SSF

[AAJS17] Parosh Aziz Abdulla, Stavros Aronis, Bengt Jonsson, and Konstantinos Sagonas. Source sets: a foundation for optimal dynamic partial order reduction. Journal of the ACM, 64(4):25:1– 25:??, September 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Abrahamsen:2022:AGP

[AAM22] Mikkel Abrahamsen, Anna Adamaszek, and Tillmann Miltzow. The art gallery problem is ∃R-complete. Journal of the ACM, 69(1):4:1–4:70, February 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3486220.

Afek:1996:LMG

[AAPS96] Yehuda Afek, Baruch Awerbuch, Serge Plotkin, and Michael Saks. Local management of a global resource in a communication network. Journal of the ACM, 43(1):1–19, January 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
jacm/227596.html.

Anderson:1970:LFR

[AB70] Robert Anderson and W. W. Bledsoe. A linear format for resolution with merging and a new technique for establishing completeness. Journal of the ACM, 17(3):525–534, July 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Agrawal:2003:PIT

[AB03] Manindra Agrawal and Somenath Biswas. Primality and identity testing via Chinese remaindering. Journal of the ACM, 50(4): 429–443, July 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Abadi:2005:ASP

[AB05] Martín Abadi and Bruno Blanchet. Analyzing security protocols with secrecy types and logic programs. Journal of the ACM, 52 (1):102–146, January 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Abboud:2017:ASE

[AB17] Amir Abboud and Greg Bodwin. The 4/3 additive spanner exponent is tight. Journal of the ACM, 64(4):28:1–28:??, September 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Abadi:1999:STS

[Aba99] Martín Abadi. Secrecy by typing in security protocols. Journal of the ACM, 46(5):749–786, September 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ambainis:2017:SQC

[ABB^{+}17] Andris Ambainis, Kaspars Balodis, Aleksandrs Belovs, Troy Lee, Miklos Santha, and Juris Smotrovs. Separations in query complexity based on pointer functions. Journal of the ACM, 64(5): 32:1–32:??, October 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Afshani:2017:IOG

[ABC17] Peyman Afshani, Jérémy Barbay, and Timothy M. Chan. Instance-optimal geometric algorithms. Journal of the ACM, 64 (1):3:1–3:??, March 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Abdelmalek:1971:LAD

[Abd71] Nabih N. Abdelmalek. Linear L_{1} approximation for a discrete point set and L_{1} solutions of overdetermined linear equations. Journal of the ACM, 18(1):41–47, January 1971. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Atserias:2021:CHA

[ABD^{+}21] Albert Atserias, Ilario Bonacina, Susanna F. De Rezende, Massimo Lauria, Jakob Nordström, and Alexander Razborov. Clique is hard on average for regular resolution. Journal of the ACM, 68(4):23:1–23:26, August 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3449352.

Alon:1997:SSD

[ABDCBH97] Noga Alon, Shai Ben-David, Nicolò Cesa-Bianchi, and David Haussler. Scale-sensitive dimensions, uniform convergence, and learnability. Journal of the ACM, 44(4):615–631, July 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org:80/pubs/citations/
journals/jacm/1997-44-4/p615-alon/.

Ashtiani:2020:NOS

[ABDH^{+}20] Hassan Ashtiani, Shai Ben-David, Nicholas J. A. Harvey, Christopher Liaw, Abbas Mehrabian, and Yaniv Plan. Nearoptimal sample complexity bounds for robust learning of Gaussian mixtures via compression schemes. Journal of the ACM, 67(6):32:1–32:42, November 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/doi/10.1145/3417994.

Aberth:1968:ACN

[Abe68] Oliver Aberth. Analysis in the computable number field. Journal of the ACM, 15(2):275–299, April 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Abelson:1980:LBI

[Abe80] Harold Abelson. Lower bounds on information transfer in distributed computations. Journal of the ACM, 27(2):384–392, April 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ajtai:2002:CEU

[ABF^{+}02] Miklos Ajtai, Randal Burns, Ronald Fagin, Darrell D. E. Long, and Larry Stockmeyer. Compactly encoding unstructured inputs with differential compression. Journal of the ACM, 49(3):318– 367, May 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Abadi:2017:APC

[ABF17] Martín Abadi, Bruno Blanchet, and Cédric Fournet. The applied pi calculus: Mobile values, new names, and secure communication. Journal of the ACM, 65(1):1:1–1:??, December 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Abiteboul:1985:DD

[Abi85] Serge Abiteboul. Disaggregations in databases. Journal of the ACM, 32(1):79–101, January 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/2463.html.

Almagor:2016:FRA

[ABK16] Shaull Almagor, Udi Boker, and Orna Kupferman. Formally reasoning about quality. Journal of the ACM, 63(3):24:1–24:??, September 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Abramson:1980:CPG

[ABL80] Fred G. Abramson, Yuri Breitbart, and Forbes D. Lewis. Complex properties of grammars. Journal of the ACM, 27(3):484– 498, July 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Awasthi:2017:PLE

[ABL17] Pranjal Awasthi, Maria Florina Balcan, and Philip M. Long. The power of localization for efficiently learning linear separators with noise. Journal of the ACM, 63(6):50:1–50:??, February 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Alon:2022:POL

[ABL^{+}22] Noga Alon, Mark Bun, Roi Livni, Maryanthe Malliaris, and Shay Moran. Private and online learnability are equivalent. Journal of the ACM, 69(4):28:1–28:??, August 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3526074.

Aggarwal:1996:ERO

[ABNC^{+}96] Alok Aggarwal, Amotz Bar-Noy, Don Coppersmith, Rajiv Ramaswami, Baruch Schieber, and Madhu Sudan. Efficient routing in optical networks. Journal of the ACM, 43(6):973–1001, November 1996. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/jacm/235812.html.

Attiya:1990:RAE

[ABND^{+}90] Hagit Attiya, Amotz Bar-Noy, Danny Dolev, David Peleg, and Rüdiger Reishuk. Renaming in an asynchronous environment. Journal of the ACM, 37(3):524–548, July 1990. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
79158.html.

Attiya:1995:SMR

[ABND95] Hagit Attiya, Amotz Bar-Noy, and Danny Dolev. Sharing memory robustly in message-passing systems. Journal of the ACM, 42(1):124–142, January 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/200869.html.

Arora:2015:SAU

[ABS15] Sanjeev Arora, Boaz Barak, and David Steurer. Subexponential algorithms for unique games and related problems. Journal of the ACM, 62(5):42:1–42:??, November 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anderson:1960:NMS

[ABV60] W. H. Anderson, R. B. Ball, and J. R. Voss. A numerical method for solving control differential equations on digital computers. Journal of the ACM, 7(1):61–68, January 1960. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Abramowitz:1955:VSC

[AC55] Milton Abramowitz and William F. Cahill. On the vibration of a square clamped plate. Journal of the ACM, 2(3):162–168, July 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Achugbue:1981:BSI

[AC81] James O. Achugbue and Francis Y. Chin. Bounds on schedules for independent tasks with similar execution times. Journal of the ACM, 28(1):81–99, January 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ailon:2005:LBL

[AC05] Nir Ailon and Bernard Chazelle. Lower bounds for linear degeneracy testing. Journal of the ACM, 52(2):157–171, March 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Attiya:2008:TBA

[AC08] Hagit Attiya and Keren Censor. Tight bounds for asynchronous randomized consensus. Journal of the ACM, 55(5):20:1–20:26, October 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Afshani:2023:LBS

[AC23] Peyman Afshani and Pingan Cheng. Lower bounds for semialgebraic range searching and stabbing problems. Journal of the ACM, 70(2):16:1–16:??, April 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/doi/10.1145/3578574.

Altschuler:2024:FHA

[AC24] Jason M. Altschuler and Sinho Chewi. Faster high-accuracy logconcave sampling via algorithmic warm starts. Journal of the ACM, 71(3):24:1–24:??, June 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/doi/10.1145/3653446.

Attar:1978:KLT

[ACDF78] R. Attar, Y. Choueka, N. Dershowitz, and A. S. Fraenkel. Kedma — linguistic tools for retrieval systems. Journal of the ACM, 25 (1):52–66, January 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Alistarh:2016:LFC

[ACHS16] Dan Alistarh, Keren Censor-Hillel, and Nir Shavit. Are lockfree concurrent algorithms practically wait-free? Journal of the ACM, 63(4):31:1–31:??, November 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Arenas:2021:NAF

[ACJR21] Marcelo Arenas, Luis Alberto Croquevielle, Rajesh Jayaram, and Cristian Riveros. #NFA admits an FPRAS: Efficient enumeration, counting, and uniform generation for logspace classes. Journal of the ACM, 68(6):48:1–48:40, December 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3477045.

Awerbuch:2008:OMS

[ACK08] Baruch Awerbuch, Israel Cidon, and Shay Kutten. Optimal maintenance of a spanning tree. Journal of the ACM, 55(4): 18:1–18:45, September 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Achlioptas:2009:BTS

[ACKM09] Dimitris Achlioptas, Aaron Clauset, David Kempe, and Cristopher Moore. On the bias of traceroute sampling: Or, power-law degree distributions in regular graphs. Journal of the ACM, 56 (4):21:1–21:28, June 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Asarin:2002:TRE

[ACM02] Eugene Asarin, Paul Caspi, and Oded Maler. Timed regular expressions. Journal of the ACM, 49(2):172–206, March 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ailon:2008:AII

[ACN08] Nir Ailon, Moses Charikar, and Alantha Newman. Aggregating inconsistent information: Ranking and clustering. Journal of the ACM, 55(5):23:1–23:27, October 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Arnborg:1993:ATG

[ACPS93] Stefan Arnborg, Bruno Courcelle, Andrzej Proskurowski, and Detlef Seese. An algebraic theory of graph reduction. Journal of the ACM, 40(5):1134–1164, November 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/169807.html.

Andreev:1998:NGD

[ACR98] Alexander E. Andreev, Andrea E. F. Clementi, and José D. P. Rolim. A new general derandomization method. Journal of the ACM, 45(1):179–213, January 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:/

/www.acm.org:80/pubs/citations/journals/jacm/1998-45-
1/p179-andreev/.

Atallah:1991:OPA

[ACW91] Mikhail J. Atallah, Danny Z. Chen, and Hubert Wagener. An optimal parallel algorithm for the visibility of a simple polygon from a point. Journal of the ACM, 38(3):516–533, July 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/116827.html.

Attiya:2001:IIB

[AD01] Hagit Attiya and Eyal Dagan. Improved implementations of binary universal operations. Journal of the ACM, 48(5):1013–1037, September 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Alur:2017:STT

[AD17] Rajeev Alur and Loris D’Antoni. Streaming tree transducers. Journal of the ACM, 64(5):31:1–31:??, October 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Abraham:2016:HDP

[ADF^{+}16] Ittai Abraham, Daniel Delling, Amos Fiat, Andrew V. Goldberg, and Renato F. Werneck. Highway dimension and provably efficient shortest path algorithms. Journal of the ACM, 63(5): 41:1–41:??, December 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Avron:2015:RAL

[ADG15] Haim Avron, Alex Druinsky, and Anshul Gupta. Revisiting asynchronous linear solvers: Provable convergence rate through randomization. Journal of the ACM, 62(6):51:1–51:??, December 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Allard:1963:MCR

[ADH63] J. L. Allard, A. R. Dobell, and T. E. Hull. Mixed congruential random number generators for decimal machines. Journal of the ACM, 10:131–141, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anderson:2015:SLP

[ADH15] Matthew Anderson, Anuj Dawar, and Bjarki Holm. Solving linear programs without breaking abstractions. Journal of the ACM, 62(6):48:1–48:??, December 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Adiri:1969:CTS

[Adi69] Igal Adiri. Computer time-sharing queues with priorities. Journal of the ACM, 16(4):631–645, October 1969. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Adiri:1971:DTS

[Adi71a] Igal Adiri. A dynamic time-sharing priority queue. Journal of the ACM, 18(4):603–610, October 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Adiri:1971:NSM

[Adi71b] Igal Adiri. A note on some mathematical models of time-sharing systems. Journal of the ACM, 18(4):611–615, October 1971. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Adiri:1973:CQB

[Adi73] Igal Adiri. Cyclic queues with bulk arrivals. Journal of the ACM, 20(3):416–428, July 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Atserias:2006:PUH

[ADK06] Albert Atserias, Anuj Dawar, and Phokion G. Kolaitis. On preservation under homomorphisms and unions of conjunctive queries. Journal of the ACM, 53(2):208–237, March 2006. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Alstrup:2017:OIU

[ADK17] Stephen Alstrup, Søren Dahlgaard, and Mathias Bæk Tejs Knudsen. Optimal induced universal graphs and adjacency labeling for trees. Journal of the ACM, 64(4):27:1–27:??, September 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Adler:2005:TOP

[Adl05] Micah Adler. Trade-offs in probabilistic packet marking for IP traceback. Journal of the ACM, 52(2):217–244, March 2005. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Amano:2014:XSM

[ADLM14] Shun’ichi Amano, Claire David, Leonid Libkin, and Filip Murlak. XML schema mappings: Data exchange and metadata management. Journal of the ACM, 61(2):12:1–12:??, April 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Attiya:1994:BTR

[ADLS94] Hagit Attiya, Cynthia Dwork, Nancy Lynch, and Larry Stockmeyer. Bounds on the time to reach agreement in the presence of timing uncertainty. Journal of the ACM, 41(1):122–152, January 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/174649.html.

Alon:1999:LHF

[ADM^{+}99] Noga Alon, Martin Dietzfelbinger, Peter Bro Miltersen, Erez Petrank, and Gábor Tardos. Linear hash functions. Journal of the ACM, 46(5):667–683, September 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/articles/journals/jacm/
1999-46-5/p667-alon/p667-alon.pdf;   http://www.acm.
org/pubs/citations/journals/jacm/1999-46-5/p667-alon/

.

Alon:1991:ESF

[ADO91] Noga Alon, A. K. Dewdney, and Teunis J. Ott. Efficient simulation of finite automata by neural nets. Journal of the ACM, 38(2):495–514, April 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/103523.html.

Atserias:2021:PSL

[ADO21] Albert Atserias, Anuj Dawar, and Joanna Ochremiak. On the power of symmetric linear programs. Journal of the ACM, 68 (4):26:1–26:35, August 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3456297.

Ausiello:1983:GAF

[ADS83] Giorgio Ausiello, Alessandro D’Atri, and Domenico Saccà. Graph algorithms for functional dependency manipulation. Journal of the ACM, 30(4):752–766, October 1983. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Aho:1971:POP

[ADU71] Alfred V. Aho, Peter J. Denning, and Jeffrey D. Ullman. Principles of optimal page replacement. Journal of the ACM, 18 (1):80–93, January 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Aho:1972:WMS

[ADU72] A. V. Aho, P. J. Denning, and J. D. Ullman. Weak and mixed strategy precedence parsing. Journal of the ACM, 19(2):225– 243, April 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Abate:1968:QAI

[ADW68] Joseph Abate, Harvey Dubner, and Sheldon B. Weinberg. Queueing analysis of the IBM 2314 disk storage facility. Journal of the ACM, 15(4):577–589, October 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Antoy:2000:NNS

[AEH00] Sergio Antoy, Rachid Echahed, and Michael Hanus. A needed narrowing strategy. Journal of the ACM, 47(4):776–822, 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/citations/
journals/jacm/2000-47-4/p776-antoy/.

Avin:2012:SDC

[AEK^{+}12] Chen Avin, Yuval Emek, Erez Kantor, Zvi Lotker, David Peleg, and Liam Roditty. SINR diagrams: Convexity and its applications in wireless networks. Journal of the ACM, 59(4):18:1–18:34, August 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ascher:1958:SEU

[AF58] Marcia Ascher and George E. Forsythe. SWAC experiments on the use of orthogonal polynomials for data fitting. Journal of the ACM, 5(1):9–21, January 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Aanderaa:1967:SHP

[AF67] Stål Aanderaa and Patrick C. Fischer. The solvability of the halting problem for 2-state post machines. Journal of the ACM, 14(4):677–682, October 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Attar:1977:LFF

[AF77] R. Attar and A. S. Fraenkel. Local feedback in full-text retrieval systems. Journal of the ACM, 24(3):397–417, July 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Attiya:2003:AAP

[AF03] Hagit Attiya and Arie Fouren. Algorithms adapting to point contention. Journal of the ACM, 50(4):444–468, July 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Andrews:2005:SRS

[AFGZ05] Matthew Andrews, Antonio Fernández, Ashish Goel, and Lisa Zhang. Source routing and scheduling in packet networks. Journal of the ACM, 52(4):582–601, July 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Alur:1996:BRP

[AFH96] Rajeev Alur, Tomás Feder, and Thomas A. Henzinger. The benefits of relaxing punctuality. Journal of the ACM, 43(1):116– 146, January 1996. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/jacm/227602.html.

Arjomandi:1983:ESV

[AFL83] Eshrat Arjomandi, Michael J. Fischer, and Nancy A. Lynch. Efficiency of synchronous versus asynchronous distributed systems. Journal of the ACM, 30(3):449–456, July 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Alber:2004:PTD

[AFN04] Jochen Alber, Michael R. Fellows, and Rolf Niedermeier. Polynomial-time data reduction for dominating set. Journal of the ACM, 51(3):363–384, May 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ackerman:2019:CCP

[AFR19] Nathanael L. Ackerman, Cameron E. Freer, and Daniel M. Roy. On the computability of conditional probability. Journal of the ACM, 66(3):23:1–23:??, June 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=3321699.

Arora:1973:OSL

[AG73] S. R. Arora and A. Gallo. Optimization of static loading and sizing of multilevel memory systems. Journal of the ACM, 20 (2):307–319, April 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Abiteboul:1986:TSL

[AG86] Serge Abiteboul and Seymour Ginsburg. Tuple sequences and lexicographic indexes. Journal of the ACM, 33(3):409–422, July 1986. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/5926.html.

Ajtai:1987:MVP

[AG87] Miklos Ajtai and Yuri Gurevich. Monotone versus positive. Journal of the ACM, 34(4):1004–1015, October 1987. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
31852.html.

Agrawal:2017:NOR

[AG17] Shipra Agrawal and Navin Goyal. Near-optimal regret bounds for Thompson sampling. Journal of the ACM, 64(5):30:1–30:??, October 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Arden:1962:ATB

[AGG62] Bruce W. Arden, Bernard A. Galler, and Robert M. Graham. An algorithm for translating Boolean expressions. Journal of the ACM, 9(2):222–239, April 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Altman:2000:BSO

[AGH00] Eitan Altman, Bruno Gaujal, and Arie Hordijk. Balanced sequences and optimal routing. Journal of the ACM, 47(4): 752–775, July 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/citations/journals/jacm/2000-47-4/p752-altman/.

Ahmadi:2025:APC

[AGH^{+}25] Ali Ahmadi, Iman Gholami, MohammadTaghi Hajiaghayi, Peyman Jabbarzade, and Mohammad Mahdavi. 2-approximation for prize-collecting Steiner forest. Journal of the ACM, 72(2):17:1– 17:??, April 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Attiya:2009:COF

[AGHK09] Hagit Attiya, Rachid Guerraoui, Danny Hendler, and Petr Kuznetsov. The complexity of obstruction-free implementations. Journal of the ACM, 56(4):24:1–24:24, June 2009. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Attiya:2017:CTM

[AGHR17] Hagit Attiya, Alexey Gotsman, Sandeep Hans, and Noam Rinetzky. Characterizing transactional memory consistency conditions using observational refinement. Journal of the ACM, 65 (1):2:1–2:??, December 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ameloot:2017:PCT

[AGK^{+}17] Tom J. Ameloot, Gaetano Geck, Bas Ketsman, Frank Neven, and Thomas Schwentick. Parallel-correctness and transferability for conjunctive queries. Journal of the ACM, 64(5):36:1–36:??, October 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Albers:2000:MST

[AGL00] Susanne Albers, Naveen Garg, and Stefano Leonardi. Minimizing stall time in single and parallel disk systems. Journal of the ACM, 47(6):969–986, 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/citations/journals/jacm/2000-47-6/p969-albers/.

Afek:1995:CFS

[AGMT95] Yehuda Afek, David S. Greenberg, Michael Merritt, and Gadi Taubenfeld. Computing with faulty shared objects. Journal of the ACM, 42(6):1231–1274, November 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/227688.html.

Arden:1966:PAS

[AGOW66] B. W. Arden, B. A. Galler, T. C. O’Brien, and F. H. Westervelt. Program and addressing structure in a time-sharing environment. Journal of the ACM, 13(1):1–16, January 1966. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Andersen:2016:AOL

[AGPT16] Reid Andersen, Shayan Oveis Gharan, Yuval Peres, and Luca Trevisan. Almost optimal local graph clustering using evolving sets. Journal of the ACM, 63(2):15:1–15:??, May 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Awerbuch:1990:TBI

[AGPV90] Baruch Awerbuch, Oded Goldreich, David Peleg, and Ronen Vainish. A trade-off between information and communication in broadcast protocols. Journal of the ACM, 37(2):238–256, April 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/77618.html.

Awodey:2017:HIA

[AGS17] Steve Awodey, Nicola Gambino, and Kristina Sojakova. Homotopy-initial algebras in type theory. Journal of the ACM, 63(6):51:1–51:??, February 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Argue:2021:CCB

[AGTG21] C. J. Argue, Anupam Gupta, Ziye Tang, and Guru Guruganesh. Chasing convex bodies with linear competitive ratio. Journal of the ACM, 68(5):32:1–32:10, October 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3450349.

Aceto:1992:TDD

[AH92a] L. Aceto and M. Hennessy. Termination, deadlock, and divergence. Journal of the ACM, 39(1):147–187, January 1992. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
147527.html.

Allender:1992:LBL

[AH92b] Eric Allender and Lane A. Hemachandra. Lower bounds for the low hierarchy. Journal of the ACM, 39(1):234–251, January 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/147546.html.

Alur:1994:RTL

[AH94] Rajeev Alur and Thomas A. Henzinger. A really temporal logic. Journal of the ACM, 41(1):181–204, January 1994. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
174651.html.

Avni:2019:IDB

[AHC19] Guy Avni, Thomas A. Henzinger, and Ventsislav Chonev. Infinite-duration bidding games. Journal of the ACM, 66(4): 31:1–31:??, August 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3340295.

Angluin:1993:LRO

[AHK93] Dana Angluin, Lisa Hellerstein, and Marek Karpinski. Learning read-once formulas with queries. Journal of the ACM, 40(1):185– 210, January 1993. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/138061.html.

Alur:2002:ATT

[AHK02] Rajeev Alur, Thomas A. Henzinger, and Orna Kupferman. Alternating-time temporal logic. Journal of the ACM, 49(5): 672–713, September 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Aho:1968:IGE

[Aho68] Alfred V. Aho. Indexed grammars — an extension of contextfree grammars. Journal of the ACM, 15(4):647–671, October 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Aho:1969:NSA

[Aho69] Alfred V. Aho. Nested stack automata. Journal of the ACM, 16(3):383–406, July 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Agarwal:1997:ASP

[AHPSV97] Pankaj K. Agarwal, Sariel Har-Peled, Micha Sharir, and Kasturi R. Varadarajan. Approximating shortest paths on a convex polytope in three dimensions. Journal of the ACM, 44(4):567– 584, July 1997. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org:80/pubs/
citations/journals/jacm/1997-44-4/p567-agarwal/.

Agarwal:2004:AEM

[AHPV04] Pankaj K. Agarwal, Sariel Har-Peled, and Kasturi R. Varadarajan. Approximating extent measures of points. Journal of the ACM, 51(4):606–635, July 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Adamaszek:2019:ASI

[AHPW19] Anna Adamaszek, Sariel Har-Peled, and Andreas Wiese. Approximation schemes for independent set and sparse subsets of polygons. Journal of the ACM, 66(4):29:1–29:??, August 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=
3326122.

Aspnes:1994:CN

[AHS94] James Aspnes, Maurice Herlihy, and Nir Shavit. Counting networks. Journal of the ACM, 41(5):1020–1048, September 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/185815.html.

Aho:1976:BCL

[AHU76] A. V. Aho, D. S. Hirschberg, and J. D. Ullman. Bounds on the complexity of the longest common subsequence problem. Journal of the ACM, 23(1):1–12, January 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Adiri:1973:MQ

[AHY73] I. Adiri, M. Hofri, and M. Yadin. A multiprogramming queue. Journal of the ACM, 20(4):589–603, October 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Achlioptas:2016:RWF

[AI16] Dimitris Achlioptas and Fotis Iliopoulos. Random walks that find perfect objects and the Lovász local lemma. Journal of the ACM, 63(3):22:1–22:??, September 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Adachi:1984:SCG

[AIK84] Akeo Adachi, Shigeki Iwata, and Takumi Kasai. Some combinatorial game problems require Ω(n^{k}) time. Journal of the ACM, 31(2):361–376, April 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Aho:1976:OCG

[AJ76] A. V. Aho and S. C. Johnson. Optimal code generation for expression trees. Journal of the ACM, 23(3):488–501, July 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Also in Proc. ACM Symp. on Theory of Computing, pp. 207–217, 1975.

Aronson:1980:NFD

[AJM80] Alan R. Aronson, Barry E. Jacobs, and Jack Minker. A note on fuzzy deduction. Journal of the ACM, 27(4):599–603, October 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Aho:1977:CGE

[AJU77] A. V. Aho, S. C. Johnson, and J. D. Ullman. Code generation for expressions with common subexpressions. Journal of the ACM, 24(1):146–160, January 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Andrews:1970:KMC

[AK70] H. C. Andrews and J. Kane. Kronecker matrices, computer implementation, and generalized spectra. Journal of the ACM, 17 (2):260–268, April 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Atallah:1984:GPM

[AK84] Mikhail J. Atallah and S. Rao Kosaraju. Graph problems on a mesh-connected processor array. Journal of the ACM, 31(3): 649–667, July 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Abiteboul:1998:OIQ

[AK98] Serge Abiteboul and Paris C. Kanellakis. Object identity as a query language primitive. Journal of the ACM, 45(5):798–842, September 1998. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org:80/pubs/
citations/journals/jacm/1998-45-5/p1-abiteboul/.

Allender:2010:ALB

[AK10] Eric Allender and Michal Koucký. Amplifying lower bounds by means of self-reducibility. Journal of the ACM, 57(3):14:1–14:36, March 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Arora:2016:CPD

[AK16] Sanjeev Arora and Satyen Kale. A combinatorial, primal-dual approach to semidefinite programs. Journal of the ACM, 63 (2):12:1–12:??, May 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Atserias:2025:CRM

[AK25] Albert Atserias and Phokion G. Kolaitis. Consistency of relations over monoids. Journal of the ACM, 72(3):18:1–18:??, June 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Akima:1970:NMI

[Aki70] Hiroshi Akima. A new method of interpolation and smooth curve fitting based on local procedures. Journal of the ACM, 17(4):

589–602, October 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Asharov:2023:OOO

[AKL^{+}23] Gilad Asharov, Ilan Komargodski, Wei-Kai Lin, Kartik Nayak, Enoch Peserico, and Elaine Shi. OptORAMa: Optimal oblivious RAM. Journal of the ACM, 70(1):4:1–4:??, February 2023. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3566049.

Abrahamsen:2025:GEC

[AKM25] Mikkel Abrahamsen, Linda Kleist, and Tillmann Miltzow. Geometric embeddability of complexes is ∃R-complete. Journal of the ACM, 72(1):9:1–9:??, February 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3707201.

Aguilera:2011:DAS

[AKMS11] Marcos K. Aguilera, Idit Keidar, Dahlia Malkhi, and Alexander Shraer. Dynamic atomic storage without consensus. Journal of the ACM, 58(2):7:1–7:32, April 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Alon:2008:WNI

[AKN^{+}08] Noga Alon, Haim Kaplan, Gabriel Nivasch, Micha Sharir, and Shakhar Smorodinsky. Weak ϵ-nets and interval chains. Journal of the ACM, 55(6):28:1–28:32, December 2008. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Adil:2024:FAR

[AKPS24] Deeksha Adil, Rasmus Kyng, Richard Peng, and Sushant Sachdeva. Fast algorithms for ℓ_{p}-regression. Journal of the ACM, 71(5):34:1–34:??, October 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3686794.

Ambainis:2012:QLL

[AKS12] Andris Ambainis, Julia Kempe, and Or Sattath. A quantum Lovász local lemma. Journal of the ACM, 59(5):24:1–24:24, October 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

An:2015:ICA

[AKS15] Hyung-Chan An, Robert Kleinberg, and David B. Shmoys. Improving Christofides’ algorithm for the s-t path TSP. Journal of the ACM, 62(5):34:1–34:??, November 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Arenas:2005:ITA

[AL05] Marcelo Arenas and Leonid Libkin. An information-theoretic approach to normal forms for relational and XML data. Journal of the ACM, 52(2):246–283, March 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Arenas:2008:XDE

[AL08] Marcelo Arenas and Leonid Libkin. XML data exchange: Consistency and query answering. Journal of the ACM, 55(2):7:1– 7:72, May 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Allender:1989:UCC

[All89] Eric W. Allender. P -uniform circuit complexity. Journal of the ACM, 36(4):912–928, October 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/76370.html.

Arora:1998:PVH

[ALM^{+}98] Sanjeev Arora, Carsten Lund, Rajeev Motwani, Madhu Sudan, and Mario Szegedy. Proof verification and the hardness of approximation problems. Journal of the ACM, 45(3):501–555, May 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org:80/pubs/citations/
journals/jacm/1998-45-3/p501-arora/.

Alonso:1960:SMT

[Alo60] R. Alonso. A starting method for the three-point Adams predictor-corrector method. Journal of the ACM, 7(2):176–180, April 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Attiya:1994:WFA

[ALS94] Hagit Attiya, Nancy Lynch, and Nir Shavit. Are wait-free algorithms fast? Journal of the ACM, 41(4):725–763, July 1994. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
179902.html.

Arvestad:2009:GEM

[ALS09] Lars Arvestad, Jens Lagergren, and Bengt Sennblad. The gene evolution model and computing its associated probabilities. Journal of the ACM, 56(2):7:1–7:44, April 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Alt:1957:NNa

[Alt57a] F. L. Alt. News and notices. Journal of the ACM, 4(1):95– 96, January 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Alt:1957:NNb

[Alt57b] F. L. Alt. News and notices. Journal of the ACM, 4(2):221– 224, April 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Alt:1957:NNc

[Alt57c] F. L. Alt. News and notices. Journal of the ACM, 4(3):367– 370, July 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Alt:1957:NNd

[Alt57d] F. L. Alt. News and notices. Journal of the ACM, 4(4):534– 540, October 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Alt:1962:DPR

[Alt62] Franz L. Alt. Digital pattern recognition by moments. Journal of the ACM, 9(2):240–258, April 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Alt:1988:CCC

[Alt88] Helmut Alt. Comparing the combinational complexities of arithmetic functions. Journal of the ACM, 35(2):447–460, April 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/214084.html.

Alt:2003:JAB

[Alt03] Franz Alt. Journal of the ACM—the beginnings. Journal of the ACM, 50(1):10–11, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ashenhurst:1959:UFP

[AM59] Robert L. Ashenhurst and Nicholas Metropolis. Unnormalized floating point arithmetic. Journal of the ACM, 6(3):415–428, July 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Augustson:1970:ASG

[AM70] J. Gary Augustson and Mack Minker. An analysis of some graph theoretical cluster techniques. Journal of the ACM, 17(4):571– 588, October 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Allen:1978:SOB

[AM78a] Brian Allen and Ian Munro. Self-organizing binary search trees. Journal of the ACM, 25(4):526–535, October 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Arvillias:1978:PPC

[AM78b] A. C. Arvillias and D. G. Maritsas. Partitioning the period of a class of m-sequences and application to pseudorandom number generation. Journal of the ACM, 25(4):675–686, October 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Altenkamp:1980:CUP

[AM80] Doris Altenkamp and Kurt Mehlhorn. Codes: Unequal probabilities, unequal letter costs. Journal of the ACM, 27(3):412–427, July 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Arbib:1982:PCE

[AM82] Michael A. Arbib and Ernest G. Manes. The pattern-of-calls expansion is the canonical fixpoint for recursive definitions. Journal of the ACM, 29(2):577–602, April 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Adler:1985:SAW

[AM85] Ilan Adler and Nimrod Megiddo. A simplex algorithm whose average number of steps is bounded between two quadratic functions of the smaller dimension. Journal of the ACM, 32(4):871– 895, October 1985. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/4222.html.

Abadi:1990:NDF

[AM90] Martín Abadi and Zohar Manna. Nonclausal deduction in firstorder temporal logic. Journal of the ACM, 37(2):279–317, April 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/77617.html.

Alon:1994:PLP

[AM94] Noga Alon and Nimrod Megiddo. Parallel linear programming in fixed dimension almost surely in constant time. Journal of the ACM, 41(2):422–434, March 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/174661.html.

Achlioptas:2007:FCL

[AM07] Dimitris Achlioptas and Frank Mcsherry. Fast computation of low-rank matrix approximations. Journal of the ACM, 54(2):9:1– 9:19, April 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Alur:2009:ANS

[AM09] Rajeev Alur and P. Madhusudan. Adding nesting structure to words. Journal of the ACM, 56(3):16:1–16:43, May 2009. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Atserias:2020:ARN

[AM20] Albert Atserias and Moritz Müller. Automating resolution is NP-hard. Journal of the ACM, 67(5):31:1–31:17, October 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3409472.

Arya:2009:STT

[AMM09] Sunil Arya, Theocharis Malamatos, and David M. Mount. Spacetime tradeoffs for approximate nearest neighbor searching. Journal of the ACM, 57(1):1:1–1:54, November 2009. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Arya:1998:OAA

[AMN^{+}98] Sunil Arya, David M. Mount, Nathan S. Netanyahu, Ruth Silverman, and Angela Y. Wu. An optimal algorithm for approximate nearest neighbor searching fixed dimensions. Journal of the ACM, 45(6):891–923, November 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:/

/www.acm.org:80/pubs/citations/journals/jacm/1998-45-
6/p891-arya/.

Ahuja:1990:FAS

[AMOT90] Ravindra K. Ahuja, Kurt Mehlhorn, James B. Orlin, and Robert E. Tarjan. Faster algorithms for the shortest path problem. Journal of the ACM, 37(2):213–223, April 1990. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
77615.html.

Ashcroft:1973:DPM

[AMP73] Edward Ashcroft, Zohar Manna, and Amir Pnueli. Decidable properties of monadic functional schemas. Journal of the ACM, 20(3):489–499, July 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Arthur:2011:SAM

[AMR11] David Arthur, Bodo Manthey, and Heiko Röglin. Smoothed analysis of the k-means method. Journal of the ACM, 58(5): 19:1–19:??, October 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Aleksandrov:2005:DAS

[AMS05] L. Aleksandrov, A. Maheshwari, and J.-R. Sack. Determining approximate shortest paths on weighted polyhedral surfaces. Journal of the ACM, 52(1):25–53, January 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anders:1965:EBN

[And65a] Edward B. Anders. An error bound for a numerical filtering technique. Journal of the ACM, 12(1):136–140, January 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anderson:1965:IPN

[And65b] Donald G. Anderson. Iterative procedures for nonlinear integral equations. Journal of the ACM, 12(4):547–560, October 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anders:1966:ERI

[And66] Edward B. Anders. An extension of Romberg integration procedures to N-variables. Journal of the ACM, 13(4):505–510, October 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Andrews:1968:CCR

[And68a] Peter B. Andrews. A correction concerning resolution. Journal of the ACM, 15(4):720, October 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [And68b].

Andrews:1968:RM

[And68b] Peter B. Andrews. Resolution with merging. Journal of the ACM, 15(3):367–381, July 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [And68a].

Andrews:1981:TPM

[And81] Peter B. Andrews. Theorem proving by matings. Journal of the ACM, 28(2):193–214, April 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Angluin:1982:IRL

[Ang82] Dana Angluin. Inference of reversible languages. Journal of the ACM, 29(3):741–765, July 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1954:NNa

[Ano54a] Anonymous. News and notices. Journal of the ACM, 1(1):44, January 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1954:NNb

[Ano54b] Anonymous. News and notices. Journal of the ACM, 1(2):92, April 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1954:NNc

[Ano54c] Anonymous. News and notices. Journal of the ACM, 1(3):137– 138, July 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1954:NNd

[Ano54d] Anonymous. News and notices. Journal of the ACM, 1(4):192, October 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1955:NNa

[Ano55a] Anonymous. News and notices. Journal of the ACM, 2(1):50– 52, January 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1955:NNb

[Ano55b] Anonymous. News and notices. Journal of the ACM, 2(2):117– 118, April 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1955:NNc

[Ano55c] Anonymous. News and notices. Journal of the ACM, 2(3):208– 210, July 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1955:NNd

[Ano55d] Anonymous. News and notices. Journal of the ACM, 2(4):280– 282, October 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1956:DCNa

[Ano56a] Anonymous. Digital Computer Newsletter. Journal of the ACM, 3(1):44–64, January 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1956:DCNb

[Ano56b] Anonymous. Digital Computer Newsletter. Journal of the ACM, 3(2):114–128, April 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1956:DCNc

[Ano56c] Anonymous. Digital Computer Newsletter. Journal of the ACM, 3(3):244–263, July 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1956:DCNd

[Ano56d] Anonymous. Digital Computer Newsletter. Journal of the ACM, 3(4):383–403, October 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1956:EN

[Ano56e] Anonymous. Editor’s note. Journal of the ACM, 3(4):265, October 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1956:NNa

[Ano56f] Anonymous. News and notices. Journal of the ACM, 3(1):41– 43, January 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1956:NNb

[Ano56g] Anonymous. News and notices. Journal of the ACM, 3(2):110– 111, April 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1956:NNc

[Ano56h] Anonymous. News and notices. Journal of the ACM, 3(3):240– 243, July 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1956:NNd

[Ano56i] Anonymous. News and notices. Journal of the ACM, 3(4):379– 382, October 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1957:CMC

[Ano57] Anonymous. Conference on matrix computations. Journal of the ACM, 4(4):520–523, October 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1958:A

[Ano58a] Anonymous. Announcement. Journal of the ACM, 5(1):117, January 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1958:AI

[Ano58b] Anonymous. Author index, 1954–1958. Journal of the ACM, 5 (4):397–403, October 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1960:BR

[Ano60] Anonymous. Book reviews. Journal of the ACM, 7(2):189– 200, April 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). The Book Reviews contributions move to ACM Computing Reviews after this issue.

Anonymous:1968:CJA

[Ano68] Anonymous. Contributions to the Journal of the Association for Computing Machinery; ACM author instructions for manuscript documentation unit; categories of the computing sciences. Journal of the ACM, 15(4):490–492, October 1968. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1974:CJA

[Ano74] Anonymous. Contributions to the Journal of the Association of Computing Machinery. Journal of the ACM, 21(3):525–526, July 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1979:CPF

[Ano79] Anonymous. Corrigendum: “Papers from the Fourth ACM Symposium on Principles of Programming Languages”. Journal of the ACM, 26(2):382, April 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1995:AI

[Ano95a] Anonymous. Author index. Journal of the ACM, 42(6):1305– 1306, November 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1995:SI

[Ano95b] Anonymous. Subject index. Journal of the ACM, 42(6):1307– 1308, November 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1997:AI

[Ano97] Anonymous. Author index. Journal of the ACM, 44(6):892, November 1997. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org:80/pubs/
citations/journals/jacm/1997-44-6/p892-author_index/.

Anonymous:1998:AI

[Ano98a] Anonymous. Author index. Journal of the ACM, 45(6):1078– 1079, November 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1998:MPC

[Ano98b] Anonymous. In memoriam: Paris C. Kanellakis. Journal of the ACM, 45(5):751, September 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:1998:SI

[Ano98c] Anonymous. Subject index. Journal of the ACM, 45(6):1076– 1077, November 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Anonymous:2015:IAF

[Ano15] Anonymous. Invited articles foreword. Journal of the ACM, 62 (4):29:1–29:??, August 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Agarwal:2004:LAP

[ANP^{+}04] Pankaj K. Agarwal, Eran Nevo, János Pach, Rom Pinchasi, Micha Sharir, and Shakhar Smorodinsky. Lenses in arrangements of pseudo-circles and their applications. Journal of the ACM, 51(2):139–186, March 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Achlioptas:2007:MSR

[ANP07] Dimitris Achlioptas, Assaf Naor, and Yuval Peres. On the maximum satisfiability of random formulas. Journal of the ACM, 54 (2):10:1–10:21, April 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ambainis:2002:DQC

[ANTSV02] Andris Ambainis, Ashwin Nayak, Amnon Ta-Shma, and Umesh Vazirani. Dense quantum coding and quantum finite automata. Journal of the ACM, 49(4):496–511, July 2002. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ameloot:2013:RTD

[ANV13] Tom J. Ameloot, Frank Neven, and Jan Van Den Bussche. Relational transducers for declarative networking. Journal of the ACM, 60(2):15:1–15:38, April 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Apt:1986:CNR

[AP86] K. R. Apt and G. D. Plotkin. Countable nondeterminism and random assignment. Journal of the ACM, 33(4):724–767, October 1986. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/6494.html.

Afrati:1993:PCS

[AP93] Foto Afrati and Christos H. Papadimitriou. The parallel complexity of simple logic programs. Journal of the ACM, 40(4): 891–916, September 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/153752.html.

Awerbuch:1995:OTM

[AP95] Baruch Awerbuch and David Peleg. Online tracking of mobile users. Journal of the ACM, 42(5):1021–1058, September 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/210132.html.

Altschuler:2025:ASH

[AP25] Jason M. Altschuler and Pablo A. Parrilo. Acceleration by stepsize hedging: Multi-step descent and the silver stepsize schedule. Journal of the ACM, 72(2):12:1–12:??, April 2025. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Arenas:2013:DEB

[APR13] Marcelo Arenas, Jorge Pérez, and Juan Reutter. Data exchange beyond complete data. Journal of the ACM, 60(4):28:1–28:59, August 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Apt:1983:FJP

[Apt83] Krzysztof R. Apt. Formal justification of a proof system for communicating sequential processes. Journal of the ACM, 30 (1):197–216, January 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Arkin:1991:MCP

[APY91] E. M. Arkin, C. H. Papadimitriou, and M. Yannakakis. Modularity of cycles and paths in graphs. Journal of the ACM, 38(2):255– 274, April 1991. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/103517.html.

Aharonov:2005:LPN

[AR05] Dorit Aharonov and Oded Regev. Lattice problems in NP ∩ coNP. Journal of the ACM, 52(5):749–765, September 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Arbib:1961:TMF

[Arb61] Michael Arbib. Turing machines, finite automata and neural nets. Journal of the ACM, 8(4):467–475, October 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Arora:1998:PTA

[Aro98] Sanjeev Arora. Polynomial time approximation schemes for Euclidean traveling salesman and other geometric problems. Journal of the ACM, 45(5):753–782, September 1998. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic).  URL http://www.acm.org:80/pubs/citations/
journals/jacm/1998-45-5/p1-arora/.

Ackermann:2008:ICS

[ARV08] Heiner Ackermann, Heiko Röglin, and Berthold Vöcking. On the impact of combinatorial structure on congestion games. Journal of the ACM, 55(6):25:1–25:22, December 2008. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Arora:2009:EFG

[ARV09] Sanjeev Arora, Satish Rao, and Umesh Vazirani. Expander flows, geometric embeddings and graph partitioning. Journal of the ACM, 56(2):5:1–5:37, April 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Atta:1954:CGH

[AS54] Susie E. Atta and Ward C. Sangren. Calculation of generalized hypergeometric series. Journal of the ACM, 1(4):170–172, October 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Awerbuch:1997:MCD

[AS97] Baruch Awerbuch and Leonard J. Schulman. The maintenance of common data in a distributed system. Journal of the ACM, 44 (1):86–103, January 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Arora:1998:PCP

[AS98] Sanjeev Arora and Shmuel Safra. Probabilistic checking of proofs: a new characterization of NP. Journal of the ACM, 45 (1):70–122, January 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org:80/

pubs/citations/journals/jacm/1998-45-1/p70-arora/.

Aaronson:2004:QLB

[AS04] Scott Aaronson and Yaoyun Shi. Quantum lower bounds for the collision and the element distinctness problems. Journal of the ACM, 51(4):595–605, July 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Angelopoulos:2013:PLU

[AS13] Spyros Angelopoulos and Pascal Schweitzer. Paging and list update under bijective analysis. Journal of the ACM, 60(2):7:1– 7:18, April 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ashenhurst:1964:FEU

[Ash64] R. L. Ashenhurst. Function evaluation in unnormalized arithmetic. Journal of the ACM, 11(2):168–187, April 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Aspnes:1998:LBD

[Asp98] James Aspnes. Lower bounds for distributed coin-flipping and randomized consensus. Journal of the ACM, 45(3):415–450, May 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org:80/pubs/citations/
journals/jacm/1998-45-3/p415-aspnes/.

Ambainis:2006:CHM

[ASV06] Andris Ambainis, Leonard J. Schulman, and Umesh Vazirani. Computing with highly mixed states. Journal of the ACM, 53 (3):507–531, May 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Attiya:1988:CAR

[ASW88] Hagit Attiya, Marc Snir, and Manfred K. Warmuth. Computing on an anonymous ring. Journal of the ACM, 35(4):845– 875, October 1988. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/48247.html.

Alexander:1971:BCD

[AT71] J. C. Alexander and A. I. Thaler. The boundary count of digital pictures. Journal of the ACM, 18(1):105–112, January 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Andersson:2007:DOS

[AT07] Arne Andersson and Mikkel Thorup. Dynamic ordered sets with exponential search trees. Journal of the ACM, 54(3):13:1–13:40, June 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Altman:2010:AAP

[AT10] Alon Altman and Moshe Tennenholtz. An axiomatic approach to personalized ranking systems. Journal of the ACM, 57(4):26:1– 26:35, April 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Avron:2011:RAE

[AT11] Haim Avron and Sivan Toledo. Randomized algorithms for estimating the trace of an implicit symmetric positive semi-definite matrix. Journal of the ACM, 58(2):8:1–8:17, April 2011. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Atallah:1994:PAE

[Ata94] Mikhail J. Atallah. Parallel algorithms for evaluating sequences of set-manipulation operations. Journal of the ACM, 41(6): 1049–1088, November 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/195617.html.

Atserias:2004:SCU

[Ats04] Albert Atserias. On sufficient conditions for unsatisfiability of random formulas. Journal of the ACM, 51(2):281–311, March 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Armoni:2000:SAC

[ATSWZ00] Roy Armoni, Amnon Ta-Shma, Avi Widgerson, and Shiyu Zhou. An O(log(n)^{4/3}) space algorithm for (s, t) connectivity in undirected graphs. Journal of the ACM, 47(2): 294–311, March 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/citations/journals/jacm/2000-47-2/p294-armoni/.

Atta:1957:EPE

[Att57] Susie E. Atta. Effect of propagated error on inverse of Hilbert matrix. Journal of the ACM, 4(1):36–40, January 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Apt:1982:CTL

[AV82] Krzysztof R. Apt and M. H. Van Emden. Contributions to the theory of logic programming. Journal of the ACM, 29(3):841– 862, July 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Abiteboul:1988:EOR

[AV88] Serge Abiteboul and Victor Vianu. Equivalence and optimization of relational transactions. Journal of the ACM, 35(1):70– 120, January 1988. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/42271.html.

Abiteboul:1989:TBA

[AV89] Serge Abiteboul and Victor Vianu. A transaction-based approach to relational database specification. Journal of the ACM, 36(4):758–789, October 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/76363.html.

Anari:2020:PGP

[AV20] Nima Anari and Vijay V. Vazirani. Planar graph perfect matching is in NC. Journal of the ACM, 67(4):21:1–21:34, August 2020. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL https://dl.acm.org/doi/abs/
10.1145/3397504.

Abiteboul:1997:FLR

[AVV97] Serge Abiteboul, Moshe Y. Vardi, and Victor Vianu. Fixpoint logics, relational machines, and computational complexity. Journal of the ACM, 44(1):30–56, January 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Abadi:2008:SAC

[AW08] Martín Abadi and Bogdan Warinschi. Security analysis of cryptographically controlled access to XML documents. Journal of the ACM, 55(2):6:1–6:29, May 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Awerbuch:1985:CNS

[Awe85] Baruch Awerbuch. Complexity of network synchronization. Journal of the ACM, 32(4):804–823, October 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
4227.html.

Alon:2013:MSN

[AY13] Noga Alon and Raphael Yuster. Matrix sparsification and nested dissection over arbitrary fields. Journal of the ACM, 60(4):25:1– 25:18, August 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Alon:1995:CC

[AYZ95] Noga Alon, Raphael Yuster, and Uri Zwick. Color-coding. Journal of the ACM, 42(4):844–856, July 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/210337.html.

Andrews:2005:STV

[AZ05] Matthew Andrews and Lisa Zhang. Scheduling over a timevarying user-dependent channel with applications to high-speed wireless data. Journal of the ACM, 52(5):809–834, September 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Andrews:2006:LHU

[AZ06] Matthew Andrews and Lisa Zhang. Logarithmic hardness of the undirected edge-disjoint paths problem. Journal of the ACM, 53 (5):745–761, September 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Andrews:2008:ATH

[AZ08] Matthew Andrews and Lisa Zhang. Almost-tight hardness of directed congestion minimization. Journal of the ACM, 55(6): 27:1–27:20, December 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Baader:1993:UCT

[Baa93] Franz Baader. Unification in commutative theories, Hilbert’s basis theorem, and Gröbner bases. Journal of the ACM, 40(3): 477–503, July 1993. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/174133.html.

Baber:1963:TST

[Bab63] R. L. Baber. Tape searching techniques. Journal of the ACM, 10 (4):478–486, October 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Babad:1975:GME

[Bab75] Jair M. Babad. A generalized multi-entrance time-sharing priority queue. Journal of the ACM, 22(2):231–247, April 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Babichenko:2016:QCA

[Bab16] Yakov Babichenko. Query complexity of approximate Nash equilibria. Journal of the ACM, 63(4):36:1–36:??, November 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Backus:1954:ISS

[Bac54] J. W. Backus. The IBM 701 Speedcoding system. Journal of the ACM, 1(1):4–6, January 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:

//community.computerhistory.org/scc/projects/FORTRAN/
paper/p4-backus.pdf.

Ben-Amram:1992:PVA

[BAG92] Amir M. Ben-Amram and Zvi Galil. On pointers versus addresses. Journal of the ACM, 39(3):617–648, July 1992. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
146666.html.

Ben-Amram:2014:RFL

[BAG14] Amir M. Ben-Amram and Samir Genaim. Ranking functions for linear-constraint loops. Journal of the ACM, 61(4):26:1–26:??, July 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bailey:1966:GSE

[Bai66] John S. Bailey. Generalized single-ended counters. Journal of the ACM, 13(3):412–418, July 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Baker:1956:PCS

[Bak56] Charles L. Baker. The PACT I coding system for the IBM Type 701. Journal of the ACM, 3(4):272–278, October 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Baker:1962:IRB

[Bak62] Frank B. Baker. Information retrieval based upon latent class analysis. Journal of the ACM, 9(4):512–521, October 1962. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Baker:1977:ASF

[Bak77] Brenda S. Baker. An algorithm for structuring flowgraphs. Journal of the ACM, 24(1):98–120, January 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Baker:1994:AAN

[Bak94] Brenda S. Baker. Approximation algorithms for NP-complete problems on planar graphs. Journal of the ACM, 41(1):153– 180, January 1994. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/174650.html.

Bates:2021:DFR

[BAL^{+}21] Stephen Bates, Anastasios Angelopoulos, Lihua Lei, Jitendra Malik, and Michael Jordan. Distribution-free, risk-controlling prediction sets. Journal of the ACM, 68(6):43:1–43:34, December 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3478535.

Ben-Amram:2006:BSL

[BAP06] Amir M. Ben-Amram and Holger Petersen. Backing up in singly linked lists. Journal of the ACM, 53(4):681–705, July 2006. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bareiss:1960:RPM

[Bar60] Erwin H. Bareiss. Resultant procedure and the mechanization of the Graeffe process. Journal of the ACM, 7(4):346–386, October 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Barnes:1965:GAS

[Bar65] Bruce Barnes. Groups of automorphisms and sets of equivalence classes of input for automata. Journal of the ACM, 12(4):561– 565, October 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Barenboim:2016:DCS

[Bar16] Leonid Barenboim. Deterministic (Δ + 1)-coloring in sublinear (in Δ) time in static, dynamic, and faulty networks. Journal of the ACM, 63(5):47:1–47:??, December 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bashkow:1958:CPR

[Bas58] T. R. Bashkow. A “curve plotting” routine for the inverse Laplace transform of rational functions. Journal of the ACM, 5(1):52–56, January 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Basu:1999:NRQ

[Bas99] Saugata Basu. New results on quantifier elimination over real closed fields and applications to constraint databases. Journal of the ACM, 46(4):537–555, July 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/articles/journals/jacm/
1999-46-4/p537-basu/p537-basu.pdf;   http://www.acm.
org/pubs/citations/journals/jacm/1999-46-4/p537-basu/

.

Bauer:1956:ICS

[Bau56] Walter F. Bauer. An integrated computation system for the ERA-1103. Journal of the ACM, 3(3):181–185, July 1956. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bauer:1958:MMI

[Bau58a] F. L. Bauer. On modern matrix iteration processes of Bernoulli and Graeffe type. Journal of the ACM, 5(3):246–257, July 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Baumann:1958:HSR

[Bau58b] D. M. Baumann. A high-scanning-rate storage device for computer applications. Journal of the ACM, 5(1):76–88, January 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Baudet:1978:AIM

[Bau78] Gérard M. Baudet. Asynchronous iterative methods for multiprocessors. Journal of the ACM, 25(2):226–244, April 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bauer:1985:SCS

[Bau85] M. A. Bauer. Soundness and completeness of a synthesis algorithm based on example computations. Journal of the ACM, 32(2):249–279, April 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/3150.html.

Bavel:1968:STP

[Bav68] Zamir Bavel. Structure and transition-preserving functions of finite automata. Journal of the ACM, 15(1):135–158, January 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bayes:1972:MVS

[Bay72] A. J. Bayes. A minimum variance sampling technique for simulation models. Journal of the ACM, 19(4):734–741, October 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bazilevskii:1957:UED

[Baz57] IU. IA. Bazilevsk�ıı. The universal electronic digital machine (URAL) for engineering research. Journal of the ACM, 4(4): 511–519, October 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bofinger:1958:PPP

[BB58] Eve Bofinger and V. J. Bofinger. On a periodic property of pseudo-random sequences. Journal of the ACM, 5(3):261–265, July 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bratman:1959:SSS

[BB59] Harvey Bratman and Ira V. Boldt, Jr. The share 709 system: Supervisory control. Journal of the ACM, 6(2):152–155, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brigham:1961:GSP

[BB61] R. C. Brigham and P. D. Burgess. Generalized simulation of post office systems. Journal of the ACM, 8(2):252–259, April 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Borko:1963:ADC

[BB63] Harold Borko and Myrna Bernick. Automatic document classification. Journal of the ACM, 10:151–162, 1963. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [BB64].

Borko:1964:ADC

[BB64] Harold Borko and Myrna Bernick. Automatic document classification. Part II. Additional experiments. Journal of the ACM, 11(2):138–151, April 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [BB63].

Ballantyne:1977:APT

[BB77] A. Michael Ballantyne and W. W. Bledsoe. Automatic proofs of theorems in analysis using nonstandard techniques. Journal of the ACM, 24(3):353–374, July 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bender:1985:ESF

[BB85] Edward A. Bender and Jon T. Butler. Enumeration of structured flowcharts. Journal of the ACM, 32(3):537–548, July 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/3832.html.

Buntine:1994:SED

[BB94] Wray L. Buntine and Hans-Jürgen Bürckert. On solving equations and disequations. Journal of the ACM, 41(4):591–629, July 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/179813.html.

Bay:1995:DLR

[BB95] Paul Bay and Gianfranco Bilardi. Deterministic on-line routing on area-universal networks. Journal of the ACM, 42(3):614– 640, May 1995. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/210417.html.

Balcan:2010:DMS

[BB10] Maria-Florina Balcan and Avrim Blum. A discriminative model for semi-supervised learning. Journal of the ACM, 57(3):19:1– 19:46, March 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Beimel:2000:LFR

[BBB^{+}00] Amos Beimel, Francesco Bergadano, Nader H. Bshouty, Eyal Kushilevitz, and Stefano Varricchio. Learning functions represented as multiplicity automata. Journal of the ACM, 47 (3):506–530, May 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/citations/journals/jacm/2000-47-5/p506-beimel/.

Bonamy:2021:ESA

[BBB^{+}21] Marthe Bonamy, Édouard Bonnet, Nicolas Bousquet, Pierre Charbit, Panos Giannopoulos, Eun Jung Kim, Paweł Rzazewski, Florian Sikora, and Stéphan Thomassé. EPTAS and subexponential algorithm for maximum clique on disk and unit ball graphs. Journal of the ACM, 68(2):9:1–9:38, March 2021. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3433160.

Berger:1995:NOA

[BBBL95] Bonnie Berger, Martin Brady, Donna Brown, and Tom Leighton. Nearly optimal algorithms and bounds for multilayer channel routing. Journal of the ACM, 42(2):500–542, March 1995. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
201037.html.

Beals:2001:QLB

[BBC^{+}01] Robert Beals, Harry Buhrman, Richard Cleve, Michele Mosca, and Ronald de Wolf. Quantum lower bounds by polynomials. Journal of the ACM, 48(4):778–797, July 2001. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Beyersdorff:2020:FSQ

[BBCP20] Olaf Beyersdorff, Ilario Bonacina, Leroy Chew, and Jan Pich. Frege systems for quantified Boolean logic. Journal of the ACM, 67(2):9:1–9:36, May 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3381881.

Baruah:2015:PUS

[BBD^{+}15] Sanjoy Baruah, Vincenzo Bonifaci, Gianlorenzo D’Angelo, Haohan Li, Alberto Marchetti-Spaccamela, Suzanne Van Der Ster, and Leen Stougie. Preemptive uniprocessor scheduling of mixedcriticality sporadic task systems. Journal of the ACM, 62(2): 14:1–14:??, May 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Baer:1970:LOP

[BBE70] J. L. Baer, D. P. Bovet, and G. Estrin. Legality and other properties of graph models of computations. Journal of the ACM, 17(3):543–554, July 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bhattacharyya:2021:PIE

[BBE^{+}21] Arnab Bhattacharyya, Édouard Bonnet, László Egri, Suprovat Ghoshal, Karthik C. S., Bingkai Lin, Pasin Manurangsi, and Dániel Marx. Parameterized intractability of even set and shortest vector problem. Journal of the ACM, 68(3):16:1–16:40, May 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3444942.

Brandl:2018:PIE

[BBEG18] Florian Brandl, Felix Brandt, Manuel Eberl, and Christian Geist. Proving the incompatibility of efficiency and strategyproofness via SMT solving. Journal of the ACM, 65(2):6:1–6:??, March 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Baltag:2023:TMC

[BBFD23] Alexandru Baltag, Nick Bezhanishvili, and David Fernández-Duque. The topological mu-calculus: Completeness and decidability. Journal of the ACM, 70(5):33:1–33:??, October 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3623268.

Beeri:1989:MCN

[BBG89] Catriel Beeri, Philip A. Bernstein, and Nathan Goodman. A model for concurrency in nested transaction systems. Journal of the ACM, 36(2):230–269, April 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/62046.html.

Balcan:2013:CUA

[BBG13] Maria-Florina Balcan, Avrim Blum, and Anupam Gupta. Clustering under approximation stability. Journal of the ACM, 60 (2):8:1–8:34, April 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Blumer:1987:CIF

[BBH^{+}87] A. Blumer, J. Blumer, D. Haussler, R. McConnell, and A. Ehrenfeucht. Complete inverted files for efficient text retrieval and analysis. Journal of the ACM, 34(3):578–595, July 1987. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
28873.html.

Balliu:2021:LBM

[BBH^{+}21] Alkida Balliu, Sebastian Brandt, Juho Hirvonen, Dennis Olivetti, Mikaël Rabie, and Jukka Suomela. Lower bounds for maximal matchings and maximal independent sets. Journal of the ACM, 68(5):39:1–39:30, October 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3461458.

Baeten:1993:DBE

[BBK93] J. C. M. Baeten, J. A. Bergstra, and J. W. Klop. Decidability of bisimulation equivalence for processes generating context-free languages. Journal of the ACM, 40(3):653–682, July 1993. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
174141.html.

Barto:2021:AAP

[BBKO21] Libor Barto, Jakub Bulín, Andrei Krokhin, and Jakub Oprsal. Algebraic approach to promise constraint satisfaction. Journal of the ACM, 68(4):28:1–28:66, August 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3457606.

Bansal:2015:PCA

[BBMN15] Nikhil Bansal, Niv Buchbinder, Aleksander Madry, and Joseph (Seffi) Naor. A polylogarithmic-competitive algorithm for the k-server problem. Journal of the ACM, 62(5):40:1–40:??, November 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bansal:2012:PDR

[BBN12] Nikhil Bansal, Niv Buchbinder, and Joseph (Seffi) Naor. A primal-dual randomized algorithm for weighted paging. Journal of the ACM, 59(4):19:1–19:24, August 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Boyar:1995:SZK

[BBP95] Joan Boyar, Gilles Brassard, and René Peralta. Subquadratic zero-knowledge. Journal of the ACM, 42(6):1169–1193, November 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/227686.html.

Bashe:1954:IT

[BBR54] C. J. Bashe, W. Buchholz, and N. Rochester. The IBM Type 702. Journal of the ACM, 1(4):149–169, October 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Balcazar:1986:PTH

[BBS86] Jose L. Balcázar, Ronald V. Book, and Uwe Schöning. The polynomial-time hierarchy and sparse oracles. Journal of the ACM, 33(3):603–617, July 1986. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/5937.html.

Bollobas:1990:CDC

[BBS90] Béla Bollobás, Andrei Z. Broder, and Istvan Simon. The cost distribution of clustering in random probing. Journal of the ACM, 37(2):224–237, April 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/77619.html.

Bauer:1954:DHS

[BC54] Walter F. Bauer and John W. Carr III. On the demonstration of high-speed digital computers. Journal of the ACM, 1(4):177– 182, October 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Blum:1961:ABB

[BC61] E. K. Blum and P. C. Curtis, Jr. Asymptotic behavior of the best polynomial approximation. Journal of the ACM, 8(4):645– 647, October 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brzozowski:1969:DRE

[BC69] J. A. Brzozowski and Rina Cohen. On decompositions of regular events. Journal of the ACM, 16(1):132–144, January 1969. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Burnett:1973:CPR

[BC73] G. J. Burnett and E. G. Coffman, Jr. A combinatorial problem related to interleaved memory systems. Journal of the ACM, 20 (1):39–45, January 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Berman:1974:MLT

[BC74] G. Berman and A. W. Colijn. A modified list technique allowing binary search. Journal of the ACM, 21(2):227–232, April 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bernstein:1981:USJ

[BC81] Philip A. Bernstein and Dah-Ming W. Chiu. Using semi-joins to solve relational queries. Journal of the ACM, 28(1):25–40, January 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bhatt:1993:TRW

[BC93] Sandeep Bhatt and Jin-Yi Cai. Taking random walks to grow trees in hypercubes. Journal of the ACM, 40(3):741–764, July 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/174144.html.

Brinkman:2005:IDR

[BC05] Bo Brinkman and Moses Charikar. On the impossibility of dimension reduction in ℓ_{1}. Journal of the ACM, 52(5):766–788, September 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Baldan:2014:LTC

[BC14] Paolo Baldan and Silvia Crafa. A logic for true concurrency. Journal of the ACM, 61(4):24:1–24:??, July 2014. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bender:2023:IHO

[BCFC^{+}23] Michael A. Bender, Alex Conway, Martín Farach-Colton, William Kuszmaul, and Guido Tagliavini. Iceberg hashing: Optimizing many hash-table criteria at once. Journal of the ACM, 70 (6):40:1–40:??, December 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3625817.

Baeten:2007:CRE

[BCG07] J. C. M. Baeten, F. Corradini, and C. A. Grabmayer. A characterization of regular expressions under bisimulation. Journal of the ACM, 54(2):6:1–6:28, April 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bhatt:1996:OEB

[BCH^{+}96] Sandeep N. Bhatt, Fan R. K. Chung, Jia-Wei Hong, F. Thomson Leighton, Bojana Obrenić, Arnold L. Rosenberg, and Eric J. Schwabe. Optimal emulations by butterfly-like networks. Journal of the ACM, 43(2):293–330, March 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/jacm/226658.html.

Burgisser:2019:CHB

[BCL19] Peter Bürgisser, Felipe Cucker, and Pierre Lairez. Computing the homology of basic semialgebraic sets in weak exponential time. Journal of the ACM, 66(1):5:1–5:??, January 2019. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brown:1956:PPM

[BCLM56] J. H. Brown, John W. Carr III, Boyd Larrowe, and J. R. McReynolds. Prevention of propagation of machine error in long problems. Journal of the ACM, 3(4):348–354, October 1956. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brakerski:2021:CTQ

[BCM^{+}21] Zvika Brakerski, Paul Christiano, Urmila Mahadev, Umesh Vazirani, and Thomas Vidick. A cryptographic test of quantumness and certifiable randomness from a single quantum device. Journal of the ACM, 68(5):31:1–31:47, October 2021. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3441309.

Baskett:1975:OCM

[BCMP75] Forest Baskett, K. Mani Chandy, Richard R. Muntz, and Fernando G. Palacios. Open, closed and mixed networks of queues with different classes of customers. Journal of the ACM, 22 (2):248–260, April 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bangalore:2020:PSE

[BCP20] Laasya Bangalore, Ashish Choudhury, and Arpita Patra. The power of shunning: Efficient asynchronous Byzantine agreement revisited*. Journal of the ACM, 67(3):14:1–14:59, June 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3388788.

Balkovich:1974:CPG

[BCPW74] E. Balkovich, W. Chiu, L. Presser, and R. Wood. Comments on a paper by Gaver. Journal of the ACM, 21(2):314–315, April 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Billionnet:1992:EAT

[BCS92] A. Billionnet, M. C. Costa, and A. Sutter. An efficient algorithm for a task allocation problem. Journal of the ACM, 39(3):502– 518, July 1992. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/146646.html.

Baker:1985:ARC

[BCW85] Brenda S. Baker, Edward G. Coffman, Jr., and Dan E. Willard. Algorithms for resolving conflicts in dynamic storage allocation. Journal of the ACM, 32(2):327–343, April 1985. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Burstall:1977:TSD

[BD77] R. M. Burstall and John Darlington. A transformation system for developing recursive programs. Journal of the ACM, 24(1):44– 67, January 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Book:1987:EDG

[BD87] Ronald V. Book and Ding-Zhu Du. The existence and density of generalized complexity cores. Journal of the ACM, 34(3):718– 730, July 1987. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/28880.html.

Bachmair:1994:EIC

[BD94] Leo Bachmair and Nachum Dershowitz. Equational inference, canonical proofs, and proof orderings. Journal of the ACM, 41(2):236–276, March 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/174655.html.

Ben-David:2023:NMT

[BDB23] Shalev Ben-David and Eric Blais. A new minimax theorem for randomized algorithms. Journal of the ACM, 70(6):38:1–38:??, December 2023. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL https://dl.acm.org/doi/10.
1145/3626514.

Balcan:2024:HMD

[BDD^{+}24] Maria-Florina Balcan, Dan Deblasio, Travis Dick, Carl Kingsford, Tuomas Sandholm, and Ellen Vitercik. How much data is sufficient to learn high-performing algorithms? Journal of the ACM, 71(5):32:1–32:??, October 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3676278.

Bruno:1981:STE

[BDF81] J. Bruno, P. Downey, and G. N. Frederickson. Sequencing tasks with exponential service times to minimize the expected flow time or makespan. Journal of the ACM, 28(1):100–113, January 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bushkov:2019:PTT

[BDFG19] Victor Bushkov, Dmytro Dziuma, Panagiota Fatourou, and Rachid Guerraoui. The PCL theorem: Transactions cannot be parallel, consistent, and live. Journal of the ACM, 66(1):2:1– 2:??, January 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Beeri:1984:SAR

[BDFS84] Catriel Beeri, Martin Dowd, Ronald Fagin, and Richard Statman. On the structure of Armstrong relations for functional dependencies. Journal of the ACM, 31(1):30–46, January 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bulatov:2013:EFB

[BDG^{+}13] Andrei A. Bulatov, Martin Dyer, Leslie Ann Goldberg, Mark Jerrum, and Colin Mcquillan. The expressibility of functions on the Boolean domain, with applications to counting CSPs. Journal of the ACM, 60(5):32:1–32:36, October 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ballard:2012:GEC

[BDHS12] Grey Ballard, James Demmel, Olga Holtz, and Oded Schwartz. Graph expansion and communication costs of fast matrix multiplication. Journal of the ACM, 59(6):32:1–32:23, December 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Special issue text supplied by Joanne Pello.

Buchet:2024:SHO

[BDK24] Mickaël Buchet, Bianca B. Dornelas, and Michael Kerber. Sparse higher order Cech filtrations. Journal of the ACM, 71(4):28:1– 28:??, August 2024. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL https://dl.acm.org/doi/10.
1145/3666085.

Benedikt:1998:REP

[BDLW98] Michael Benedikt, Guozhu Dong, Leonid Libkin, and Limsoon Wong. Relational expressive power of constraint query languages. Journal of the ACM, 45(1):1–34, January 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org:80/pubs/citations/
journals/jacm/1998-45-1/p1-benedikt/.

Bloch:1987:WVA

[BDS87] Joshua J. Bloch, Dean S. Daniels, and Alfred Z. Spector. A weighted voting algorithm for replicated directories. Journal of the ACM, 34(4):859–909, October 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/31847.html.

Balcan:2024:LBG

[BDSV24] Maria-Florina Balcan, Travis Dick, Tuomas Sandholm, and Ellen Vitercik. Learning to branch: Generalization guarantees and limits of data-independent discretization. Journal of the ACM, 71(2):13:1–13:??, April 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3637840.

Bailey:1962:SFS

[BE62] John S. Bailey and George Epstein. Single function shifting counters. Journal of the ACM, 9(3):375–378, July 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bloom:1991:FHL

[BÉ91] Stephen L. Bloom and Zoltán Ésik. Floyd-Hoare logic in iteration theories. Journal of the ACM, 38(4):887–934, October 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/115352.html.

Bibel:1997:DTR

[BE97] W. Bibel and E. Eder. Decomposition of tautogies into regular formulas and strong completeness of connection-graph resolution. Journal of the ACM, 44(2):320–344, March 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Dedicated to J. A. Robinson. See erratum [SW01].

Barenboim:2011:DDV

[BE11] Leonid Barenboim and Michael Elkin. Deterministic distributed vertex coloring in polylogarithmic time. Journal of the ACM, 58 (5):23:1–23:??, October 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Beatty:1972:AAC

[Bea72] James C. Beatty. An axiomatic approach to code optimization for expressions. Journal of the ACM, 19(4):613–640, October 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [Bea73a, Bea73b].

Beatty:1973:EAAa

[Bea73a] James C. Beatty. Errata: “An axiomatic approach to code optimization for expressions”. Journal of the ACM, 20(1):188, January 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Bea72, Bea73b].

Beatty:1973:EAAb

[Bea73b] James C. Beatty. Errata: “An axiomatic approach to code optimization for expressions”. Journal of the ACM, 20(3):538, July 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Bea72, Bea73a].

Beatty:1982:RBL

[Bea82] John C. Beatty. On the relationship between the LL(1) and LR(1) grammars. Journal of the ACM, 29(4):1007–1022, October 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Blondin:2021:RPT

[BEF^{+}21] Michael Blondin, Matthias Englert, Alain Finkel, Stefan GÖller, Christoph Haase, Ranko Lazić, Pierre Mckenzie, and Patrick Totzke. The reachability problem for two-dimensional vector addition systems with states. Journal of the ACM, 68(5):34:1– 34:43, October 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/

doi/10.1145/3464794.

Boroujeni:2021:AED

[BEG^{+}21] Mahdi Boroujeni, Soheil Ehsani, Mohammad Ghodsi, Mohammadtaghi Hajiaghayi, and Saeed Seddighin. Approximating edit distance in truly subquadratic time: Quantum and MapReduce. Journal of the ACM, 68(3):19:1–19:41, May 2021. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3456807.

Barenboim:2022:LID

[BEG22] Leonid Barenboim, Michael Elkin, and Uri Goldenberg. Locallyiterative distributed (Δ + 1)-coloring and applications. Journal of the ACM, 69(1):5:1–5:26, February 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3486625.

Braverman:2017:CRC

[BEGH17] Mark Braverman, Klim Efremenko, Ran Gelles, and Bernhard Haeupler. Constant-rate coding for multiparty interactive communication is impossible. Journal of the ACM, 65(1):4:1–4:??, December 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Braverman:2022:OSC

[BEGY22] Mark Braverman, Klim Efremenko, Ran Gelles, and Michael Yitayew. Optimal short-circuit resilient formulas. Journal of the ACM, 69(4):26:1–26:??, August 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/doi/10.1145/3538390.

Blumer:1989:LVC

[BEHW89] Anselm Blumer, Andrzej Ehrenfeucht, David Haussler, and Manfred K. Warmuth. Learnability and the Vapnik–Chervonenkis dimension. Journal of the ACM, 36(4):929–965, October 1989. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
76371.html.

Ben-Eliezer:2022:FAR

[BEJWY22] Omri Ben-Eliezer, Rajesh Jayaram, David P. Woodruff, and Eylon Yogev. A framework for adversarially robust streaming algorithms. Journal of the ACM, 69(2):17:1–17:33, April 2022. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3498334.

Bellman:1960:SMA

[Bel60] Richard Bellman. Sequential machines, ambiguity, and dynamic programming. Journal of the ACM, 7(1):24–28, January 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bellman:1962:DPT

[Bel62] Richard Bellman. Dynamic programming treatment of the travelling salesman problem. Journal of the ACM, 9(1):61–63, January 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bubeck:2021:KBM

[BEL21] Sébastien Bubeck, Ronen Eldan, and Yin Tat Lee. Kernel-based methods for bandit convex optimization. Journal of the ACM, 68(4):25:1–25:35, August 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3453721.

Bender:1980:NFF

[Ben80] Edward A. Bender. The number of fanout-free functions with various gates. Journal of the ACM, 27(1):181–190, January 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bilardi:2009:AIR

[BEP09] Gianfranco Bilardi, Kattamuri Ekanadham, and Pratap Pattnaik. On approximating the ideal random access machine by physical machines. Journal of the ACM, 56(5):27:1–27:57, August 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Barenboim:2016:LDS

[BEPS16] Leonid Barenboim, Michael Elkin, Seth Pettie, and Johannes Schneider. The locality of distributed symmetry breaking. Journal of the ACM, 63(3):20:1–20:??, September 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bergman:1954:MSB

[Ber54] Stefan Bergman. A method of solving boundary value problems of mathematical physics on punch card machines. Journal of the ACM, 1(3):101–104, July 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Berman:1958:MTM

[Ber58] Martin F. Berman. A method for transposing a matrix. Journal of the ACM, 5(4):383–384, October 1958. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [WSD59].

Berlekamp:1963:PDD

[Ber63] Elwyn R. Berlekamp. Program for double-dummy bridge problems — a new strategy for mechanical game playing. Journal of the ACM, 10:357–364, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Berman:1969:LAM

[Ber69] Gerald Berman. Lattice approximations to the minima of functions of several variables. Journal of the ACM, 16(2):286–294, April 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Berztiss:1973:BPI

[Ber73] A. T. Berztiss. A backtrack procedure for isomorphism of directed graphs. Journal of the ACM, 20(3):365–377, July 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bertsch:1975:ORP

[Ber75] Eberhard Bertsch. An observation on relative parsing time. Journal of the ACM, 22(4):493–498, October 1975. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ban:2006:ISP

[BER06] Yih-En Andrew Ban, Herbert Edelsbrunner, and Johannes Rudolph. Interface surfaces for protein-protein complexes. Journal of the ACM, 53(3):361–378, May 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Beus:1970:UIS

[Beu70] H. Lynn Beus. The use of information in sorting. Journal of the ACM, 17(3):482–495, July 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bourne:1961:SMS

[BF61] Charles P. Bourne and Donald F. Ford. A study of methods for systematically abbreviating English words and names. Journal of the ACM, 8(4):538–552, October 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Barnes:1967:MEI

[BF67] Bruce H. Barnes and John M. Fitzgerald. Minimal experiments for input-independent machines. Journal of the ACM, 14(4):683– 686, October 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Boneh:2001:EGS

[BF01] Dan Boneh and Matthew Franklin. Efficient generation of shared RSA keys. Journal of the ACM, 48(4):702–722, July 2001. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Buhrman:1999:MS

[BFG^{+}99] Harry Buhrman, Matthew Franklin, Juan A. Garay, Jaap-Henk Hoepman, John Tromp, and Paul Vitányi. Mutual search. Journal of the ACM, 46(4):517–536, July 1999. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL http://www.acm.org/pubs/articles/journals/
jacm/1999-46-4/p517-buhrman/p517-buhrman.pdf;   http:
//www.acm.org/pubs/citations/journals/jacm/1999-46-4/
p517-buhrman/.

Buchsbaum:2003:ITC

[BFG03] Adam L. Buchsbaum, Glenn S. Fowler, and Raffaele Giancarlo. Improving table compression with combinatorial optimization. Journal of the ACM, 50(6):825–851, November 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Benedikt:2008:XSP

[BFG08] Michael Benedikt, Wenfei Fan, and Floris Geerts. XPath satisfiability in the presence of DTDs. Journal of the ACM, 55(2): 8:1–8:79, May 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bustoz:1979:ITD

[BFGL79] Joaquín Bustoz, Alan Feldstein, Richard Goodman, and Seppo Linnainmaa. Improved trailing digits estimates applied to optimal computer arithmetic. Journal of the ACM, 26(4):716–730, October 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Barcelo:2020:SOC

[BFGP20] Pablo Barceló, Diego Figueira, Georg Gottlob, and Andreas Pieris. Semantic optimization of conjunctive queries. Journal of the ACM, 67(6):34:1–34:60, November 2020. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3424908.

Bender:2019:SEB

[BFGY19] Michael A. Bender, Jeremy T. Fineman, Seth Gilbert, and Maxwell Young. Scaling exponential backoff: Constant throughput, polylogarithmic channel-access attempts, and robustness. Journal of the ACM, 66(1):6:1–6:??, January 2019. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Barvinok:2003:GMT

[BFJ^{+}03] Alexander Barvinok, Sándor P. Fekete, David S. Johnson, Arie Tamir, Gerhard J. Woeginger, and Russ Woodroofe. The geometric maximum traveling salesman problem. Journal of the ACM, 50(5):641–664, September 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bodlaender:2016:MK

[BFL^{+}16] Hans L. Bodlaender, Fedor V. Fomin, Daniel Lokshtanov, Eelko Penninkx, Saket Saurabh, and Dimitrios M. Thilikos. (meta) kernelization. Journal of the ACM, 63(5):44:1–44:??, December 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bishop:2019:ELR

[BFM^{+}19] Steve Bishop, Matthew Fairbairn, Hannes Mehnert, Michael Norrish, Tom Ridge, Peter Sewell, Michael Smith, and Keith Wansbrough. Engineering with logic: Rigorous test-oracle specification and validation for TCP/IP and the Sockets API. Journal of the ACM, 66(1):1:1–1:??, January 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Beeri:1983:DAD

[BFMY83] Catriel Beeri, Ronald Fagin, David Maier, and Mihalis Yannakakis. On the desirability of acyclic database schemes. Journal of the ACM, 30(3):479–513, July 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Batu:2013:TCD

[BFR^{+}13] Tugkan Batu, Lance Fortnow, Ronitt Rubinfeld, Warren D. Smith, and Patrick White. Testing closeness of discrete distributions. Journal of the ACM, 60(1):4:1–4:25, February 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bonakdarpour:2022:DAC

[BFR^{+}22] Borzoo Bonakdarpour, Pierre Fraigniaud, Sergio Rajsbaum, David Rosenblueth, and Corentin Travers. Decentralized asynchronous crash-resilient runtime verification. Journal of the ACM, 69(5):34:1–34:??, October 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/doi/10.1145/3550483.

Broder:2001:GAD

[BFU01] Andrei Z. Broder, Alan M. Frieze, and Eli Upfal. A general approach to dynamic packet routing with bounded buffers.

Journal of the ACM, 48(2):324–349, March 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/citations/journals/jacm/
2001-48-2/p324-broder/.

Boyle:1970:ASA

[BG70] J. M. Boyle and A. A. Grau. An algorithmic semantics for AL-GOL 60 identifier denotation. Journal of the ACM, 17(2):361– 382, April 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Boesch:1977:CPD

[BG77] F. T. Boesch and J. F. Gimpel. Covering the points of a digraph with point-disjoint paths and its application to code optimization. Journal of the ACM, 24(2):192–198, April 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bol:1996:MNP

[BG96] Roland Bol and Jan Friso Groote. The meaning of negative premises in transition system specifications. Journal of the ACM, 43(5):863–914, September 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/jacm/234756.html.

Bachmair:1998:OCC

[BG98] Leo Bachmair and Harald Ganzinger. Ordered chaining calculi for first-order theories of transitive relations. Journal of the ACM, 45(6):1007–1049, November 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:/

/www.acm.org:80/pubs/citations/journals/jacm/1998-45-
6/p1007-bachmair/.

Bartal:2000:HSA

[BG00] Yair Bartal and Eddie Grove. The harmonic k-server algorithm is competitive. Journal of the ACM, 47(1):1–15, January 2000. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/citations/journals/jacm/
2000-47-1/p1-bartal/.

Basin:2001:ACA

[BG01] David Basin and Harald Ganzinger. Automated complexity analysis based on ordered resolution. Journal of the ACM, 48(1):70– 109, January 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/pubs/

articles/journals/jacm/2001-48-1/p70-basin/p70-basin.
pdf;   http://www.acm.org/pubs/citations/journals/jacm/
2001-48-1/p70-basin/.

Bonacina:2015:FSC

[BG15] Ilario Bonacina and Nicola Galesi. A framework for space complexity in algebraic proof systems. Journal of the ACM, 62(3): 23:1–23:??, June 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Baier:2012:PA

[BGB12] Christel Baier, Marcus Grösser, and Nathalie Bertrand. Probabilistic ω-automata. Journal of the ACM, 59(1):1:1–1:52, February 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bonnet:2024:TWI

[BGdM^{+}24] Édouard Bonnet, Ugo Giocanti, Patrice Ossona de Mendez, Pierre Simon, Stéphan Thomassé, and Szymon Toruńczyk. Twin-width IV: Ordered graphs and matrices. Journal of the ACM, 71(3):21:1–21:??, June 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/doi/10.1145/3651151.

Bertrand:2017:QDD

[BGG17] Nathalie Bertrand, Blaise Genest, and Hugo Gimbert. Qualitative determinacy and decidability of stochastic games with signals. Journal of the ACM, 64(5):33:1–33:??, October 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bhargava:2024:FMM

[BGG^{+}24] Vishwas Bhargava, Sumanta Ghosh, Zeyu Guo, Mrinal Kumar, and Chris Umans. Fast multivariate multipoint evaluation over all finite fields. Journal of the ACM, 71(3):22:1–22:??, June 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3652025.

Bruni:2023:CIP

[BGGR23] Roberto Bruni, Roberto Giacobazzi, Roberta Gori, and Francesco Ranzato. A correctness and incorrectness program logic. Journal of the ACM, 70(2):15:1–15:??, April 2023. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3582267.

Birman:1995:OSP

[BGH^{+}95] Alexander Birman, H. Richard Gail, Sidney L. Hantler, Zvi Rosberg, and Moshe Sidi. An optimal service policy for buffer systems. Journal of the ACM, 42(3):641–657, May 1995. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
210422.html.

Barak:2012:IPO

[BGI^{+}12] Boaz Barak, Oded Goldreich, Russell Impagliazzo, Steven Rudich, Amit Sahai, Salil Vadhan, and Ke Yang. On the (im)possibility of obfuscating programs. Journal of the ACM, 59(2):6:1–6:48, April 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Blei:2010:NCR

[BGJ10] David M. Blei, Thomas L. Griffiths, and Michael I. Jordan. The nested Chinese restaurant process and Bayesian nonparametric inference of topic hierarchies. Journal of the ACM, 57(2):7:1– 7:30, January 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bonchi:2022:SDR

[BGK^{+}22] Filippo Bonchi, Fabio Gadducci, Aleks Kissinger, Pawel Sobocinski, and Fabio Zanasi. String diagram rewrite theory I: Rewriting with Frobenius structure. Journal of the ACM, 69(2):14:1–14:58, April 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3502719.

Bhargava:2023:FAM

[BGKM23] Vishwas Bhargava, Sumanta Ghosh, Mrinal Kumar, and Chandra Kanta Mohapatra. Fast, algebraic multivariate multipoint evaluation in small characteristic and applications. Journal of the ACM, 70(6):42:1–42:??, December 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3625226.

Bera:2022:CSD

[BGL^{+}22] Suman K. Bera, Lior Gishboliner, Yevgeny Levanzov, C. Seshadhri, and Asaf Shapira. Counting subgraphs in degenerate graphs. Journal of the ACM, 69(3):23:1–23:21, June 2022. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3520240.

Bshouty:1998:NTD

[BGM^{+}98] Nader H. Bshouty, Sally A. Goldman, H. David Mathias, Subhash Suri, and Hisao Tamaki. Noise-tolerant distribution-free learning of general geometric concepts. Journal of the ACM, 45(5):863–890, September 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org:80/pubs/citations/journals/jacm/1998-45-5/p1-
bshouty/.

Blelloch:1999:PES

[BGM99] Guy E. Blelloch, Phillip B. Gibbons, and Yossi Matias. Provably efficient scheduling for languages with fine-grained parallelism. Journal of the ACM, 46(2):281–321, March 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org:80/pubs/citations/
journals/jacm/1999-46-2/p281-blelloch/.

Blasiok:2022:GSP

[BGN^{+}22] Jarosław Błasiok, Venkatesan Guruswami, Preetum Nakkiran, Atri Rudra, and Madhu Sudan. General strong polarization. Journal of the ACM, 69(2):11:1–11:67, April 2022. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3491390.

Baccelli:1984:EEA

[BGP84] François Baccelli, Erol Gelenbe, and Brigitte Plateau. An end-to-end approach to the resequencing problem. Journal of the ACM, 31(3):474–485, July 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bournez:2017:PTC

[BGP17] Olivier Bournez, Daniel S. Graça, and Amaury Pouly. Polynomial time corresponds to solutions of polynomial ordinary differential equations of polynomial length. Journal of the ACM, 64 (6):38:1–38:??, November 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Byrka:2013:STA

[BGRS13] Jaroslaw Byrka, Fabrizio Grandoni, Thomas Rothvoss, and Laura Sanità. Steiner tree approximation via iterative randomized rounding. Journal of the ACM, 60(1):6:1–6:33, February 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Blelloch:2020:PRI

[BGSS20] Guy E. Blelloch, Yan Gu, Julian Shun, and Yihan Sun. Parallelism in randomized incremental algorithms. Journal of the ACM, 67(5):27:1–27:27, October 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/doi/10.1145/3402819.

Bramble:1965:ASM

[BH65] J. H. Bramble and B. E. Hubbard. Approximation of solutions of mixed boundary value problems for Poisson’s equation by finite differences. Journal of the ACM, 12(1):114–123, January 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bellmore:1974:TMP

[BH74] Mandell Bellmore and Saman Hong. Transformation of multisalesmen problem to the standard traveling salesman problem. Journal of the ACM, 21(3):500–504, July 1974. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Beame:1989:OBD

[BH89] Paul Beame and Johan Håstad. Optimal bounds for decision problems on the CRCW PRAM. Journal of the ACM, 36(3): 643–670, July 1989. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/65958.html.

Behnezhad:2023:EFM

[BHH23] Soheil Behnezhad, Mohammadtaghi Hajiaghayi, and David G. Harris. Exponentially faster massively parallel maximal matching. Journal of the ACM, 70(5):34:1–34:??, October 2023. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3617360.

Bellman:1959:ADP

[BHK59] Richard Bellman, John Holland, and Robert Kalaba. On an application of dynamic programming to the synthesis of logical systems. Journal of the ACM, 6(4):486–493, October 1959. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bergstra:1990:MA

[BHK90] J. A. Bergstra, J. Heering, and P. Klint. Module algebra. Journal of the ACM, 37(2):335–372, April 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/77621.html.

Bateni:2011:ASS

[BHM11] Mohammadhossein Bateni, Mohammadtaghi Hajiaghayi, and Dániel Marx. Approximation schemes for Steiner forest on planar graphs and graphs of bounded treewidth. Journal of the ACM, 58(5):21:1–21:??, October 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bloniarz:1984:ASH

[BHR84a] P. A. Bloniarz, H. B. Hunt, III, and D. J. Rosenkrantz. Algebraic structures with hard equivalence and minimization problems. Journal of the ACM, 31(4):879–904, October 1984. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brookes:1984:TCS

[BHR84b] S. D. Brookes, C. A. R. Hoare, and A. W. Roscoe. A theory of communicating sequential processes. Journal of the ACM, 31(3): 560–599, July 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bernstein:2019:OBM

[BHR19] Aaron Bernstein, Jacob Holm, and Eva Rotenberg. Online bipartite matching with amortized O(log^{2} n) replacements. Journal of the ACM, 66(5):37:1–37:??, September 2019. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/ft_gateway.cfm?id=3344999.

Berman:2018:CFC

[BHT18] Itay Berman, Iftach Haitner, and Aris Tentes. Coin flipping of any constant bias implies one-way functions. Journal of the ACM, 65(3):14:1–14:??, March 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bibel:1981:MC

[Bib81] Wolfgang Bibel. On matrices with connections. Journal of the ACM, 28(4):633–645, October 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Babaioff:2018:MSP

[BIKK18] Moshe Babaioff, Nicole Immorlica, David Kempe, and Robert Kleinberg. Matroid secretary problems. Journal of the ACM, 65 (6):35:1–35:??, November 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/ft_gateway.cfm?id=3212512.

Babaioff:2020:SAO

[BILW20] Moshe Babaioff, Nicole Immorlica, Brendan Lucier, and S. Matthew Weinberg. A simple and approximately optimal mechanism for an additive buyer. Journal of the ACM, 67(4): 24:1–24:40, August 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3398745.

Bloom:1995:BCT

[BIM95] Bard Bloom, Sorin Istrail, and Albert R. Meyer. Bisimulation can’t be traced. Journal of the ACM, 42(1):232–268, January 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/200876.html.

Ben-Israel:1963:EMC

[BIW63] A. Ben-Israel and S. J. Wersan. An elimination method for computing the generalized inverse of an arbitrary complex matrix. Journal of the ACM, 10:532–537, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bringmann:2018:ABP

[BIZ18] Karl Bringmann, Christian Ikenmeyer, and Jeroen Zuiddam. On algebraic branching programs of small width. Journal of the ACM, 65(5):32:1–32:??, September 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/ft_gateway.cfm?id=3209663.

Barlow:1966:MSR

[BJ66] C. A. Barlow, Jr. and E. L. Jones. A method for the solution of roots of a nonlinear equation and for solution of the general eigenvalue problem. Journal of the ACM, 13(1):135–142, January 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Burns:1982:DRI

[BJL^{+}82] James E. Burns, Paul Jackson, Nancy A. Lynch, Michael J. Fischer, and Gary L. Peterson. Data requirements for implementation of N-process mutual exclusion using a single shared variable. Journal of the ACM, 29(1):183–205, January 1982. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Beyer:1979:LAI

[BJM79] T. Beyer, W. Jones, and S. Mitchell. Linear algorithms for isomorphism of maximal outerplanar graphs. Journal of the ACM, 26(4):603–610, October 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bauer:1959:APP

[BJP59] Walter F. Bauer, Mario L. Juncosa, and Alan J. Perlis. ACM publications policies and plans. Journal of the ACM, 6(2):121– 122, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Burge:1971:AM

[BK71] William H. Burge and Alan G. Konheim. An accessing model. Journal of the ACM, 18(3):400–404, July 1971. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Beck:1972:CAR

[BK72] Robert E. Beck and Bernard Kolman. Computer approaches to the representation of Lie algebras. Journal of the ACM, 19(4): 577–589, October 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Blake:1977:BBB

[BK77a] Ian F. Blake and Alan G. Konheim. Big buckets are (are not) better! Journal of the ACM, 24(4):591–606, October 1977. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bookstein:1977:ORA

[BK77b] Abraham Bookstein and Don Kraft. Operations research applied to document indexing and retrieval decisions. Journal of the ACM, 24(3):418–427, July 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brent:1978:FAM

[BK78] R. P. Brent and H. T. Kung. Fast algorithms for manipulating formal power series. Journal of the ACM, 25(4):581–595, October 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Baker:1979:CMB

[BK79] Brenda S. Baker and S. Rao Kosaraju. A comparison of multilevel break and next statements. Journal of the ACM, 26(3): 555–566, July 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brent:1981:ATC

[BK81] R. P. Brent and H. T. Kung. The area-time complexity of binary multiplication. Journal of the ACM, 28(3):521–534, July 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [BK82].

Brent:1982:CAT

[BK82] R. P. Brent and H. T. Kung. Corrigendum: “The area-time complexity of binary multiplication”. Journal of the ACM, 29 (3):904, July 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [BK81].

Beeri:1986:EIA

[BK86] Catriel Beeri and Michael Kifer. Elimination of intersection anomalies from database schemes. Journal of the ACM, 33(3): 423–450, July 1986. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/5927.html.

Beeri:1987:TIA

[BK87] Catriel Beeri and Michael Kifer. A theory of intersection anomalies in relational database schemes. Journal of the ACM, 34(3): 544–577, July 1987. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/28872.html.

Blum:1995:DPC

[BK95] Manuel Blum and Sampath Kannan. Designing programs that check their work. Journal of the ACM, 42(1):269–291, January 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/200880.html.

Borradaile:2009:AMS

[BK09] Glencora Borradaile and Philip Klein. An O(n log n) algorithm for maximum st-flow in a directed planar graph. Journal of the ACM, 56(2):9:1–9:30, April 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bodirsky:2010:CTC

[BK10] Manuel Bodirsky and Jan Kára. The complexity of temporal constraint satisfaction problems. Journal of the ACM, 57(2): 9:1–9:41, January 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Barto:2014:CSP

[BK14a] Libor Barto and Marcin Kozik. Constraint satisfaction problems solvable by local consistency methods. Journal of the ACM, 61 (1):3:1–3:??, January 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brotherston:2014:UPS

[BK14b] James Brotherston and Max Kanovich. Undecidability of propositional separation logic and its neighbours. Journal of the ACM, 61(2):14:1–14:??, April 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Boxma:1984:PFS

[BKK84] O. J. Boxma, F. P. Kelly, and A. G. Konheim. The product form for sojourn time distributions in cyclic exponential queues. Journal of the ACM, 31(1):128–133, January 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brazdil:2014:EAP

[BKK14] Tomás Brázdil, Stefan Kiefer, and Antonín Kucera. Efficient analysis of probabilistic programs with an unbounded counter.

Journal of the ACM, 61(6):41:1–41:??, November 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bienvenu:2018:OMQ

[BKK^{+}18] Meghyn Bienvenu, Stanislav Kikot, Roman Kontchakov, Vladimir V. Podolskii, and Michael Zakharyaschev. Ontologymediated queries: Combined complexity and succinctness of rewritings via circuit complexity. Journal of the ACM, 65(5): 28:1–28:??, September 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/ft_gateway.cfm?id=3191832.

Berger:1999:RTD

[BKL99] Bonnie Berger, Jon Kleinberg, and Tom Leighton. Reconstructing a three-dimensional model with arbitrary errors. Journal of the ACM, 46(2):212–235, March 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:/

/www.acm.org:80/pubs/citations/journals/jacm/1999-46-
2/p212-berger/.

Benedikt:2006:CFO

[BKL^{+}06] Michael Benedikt, Bart Kuijpers, Christof Löding, Jan Van den Bussche, and Thomas Wilke. A characterization of first-order topological properties of planar spatial data. Journal of the ACM, 53(2):273–305, March 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Basin:2015:MMF

[BKMZ15] David Basin, Felix Klaedtke, Samuel Müller, and Eugen Zalinescu. Monitoring metric first-order temporal properties. Journal of the ACM, 62(2):15:1–15:??, May 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brakerski:2014:FIC

[BKN14] Zvika Brakerski, Yael Tauman Kalai, and Moni Naor. Fast interactive coding against adversarial noise. Journal of the ACM, 61 (6):35:1–35:??, November 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bansal:2007:SSM

[BKP07] Nikhil Bansal, Tracy Kimbrel, and Kirk Pruhs. Speed scaling to manage energy and temperature. Journal of the ACM, 54 (1):3:1–3:39, March 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Borodin:2001:AQT

[BKR^{+}01] Allan Borodin, Jon Kleinberg, Prabhakar Raghavan, Madhu Sudan, and David P. Williamson. Adversarial queuing theory. Journal of the ACM, 48(1):13–38, January 2001. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL http://www.acm.org/pubs/articles/journals/
jacm/2001-48-1/p13-borodin/p13-borodin.pdf;   http:/
/www.acm.org/pubs/citations/journals/jacm/2001-48-1/
p13-borodin/.

Barak:2010:SIN

[BKS^{+}10] B. Barak, G. Kindler, R. Shaltiel, B. Sudakov, and A. Wigderson. Simulating independence: New constructions of condensers, Ramsey graphs, dispersers, and extractors. Journal of the ACM, 57(4):20:1–20:52, April 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Babaioff:2015:TMI

[BKS15] Moshe Babaioff, Robert D. Kleinberg, and Aleksandrs Slivkins. Truthful mechanisms with implicit payment computation. Journal of the ACM, 62(2):10:1–10:??, May 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Beame:2017:CSP

[BKS17] Paul Beame, Paraschos Koutris, and Dan Suciu. Communication steps for parallel query processing. Journal of the ACM, 64(6): 40:1–40:??, November 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Badanidiyuru:2018:BK

[BKS18] Ashwinkumar Badanidiyuru, Robert Kleinberg, and Aleksandrs Slivkins. Bandits with knapsacks. Journal of the ACM, 65(3): 13:1–13:??, March 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bentley:1978:ANM

[BKST78] J. L. Bentley, H. T. Kung, M. Schkolnick, and C. D. Thompson. On the average number of maxima in a set of vectors and applications. Journal of the ACM, 25(4):536–543, October 1978.

CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bhattacharya:2024:DMB

[BKSW24] Sayan Bhattacharya, Peter Kiss, Thatchaphol Saranurak, and David Wajc. Dynamic matching with better-than-2 approximation in polylogarithmic update time. Journal of the ACM, 71 (5):33:1–33:??, October 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3679009.

Brush:1967:SOD

[BKT67] D. G. Brush, J. J. Kohfeld, and G. T. Thompson. Solution of ordinary differential equations using two “off-step” points. Journal of the ACM, 14(4):769–784, October 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bonnet:2022:TWT

[BKTW22] Édouard Bonnet, Eun Jung Kim, Stéphan Thomassé, and Rémi Watrigant. Twin-width I: Tractable FO model checking. Journal of the ACM, 69(1):3:1–3:46, February 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3486655.

Boyle:2025:MCR

[BKV25] Elette Boyle, Ilan Komargodski, and Neekon Vafa. Memory checking requires logarithmic overhead. Journal of the ACM, 72 (2):13:1–13:??, April 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Blum:2003:NTL

[BKW03] Avrim Blum, Adam Kalai, and Hal Wasserman. Noise-tolerant learning, the parity problem, and the statistical query model. Journal of the ACM, 50(4):506–519, July 2003. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Byrne:1966:PRK

[BL66] George D. Byrne and Robert J. Lambert. Pseudo-Runge–Kutta methods involving two points. Journal of the ACM, 13(1):114– 123, January 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bruno:1975:GOC

[BL75] J. L. Bruno and T. Lassagne. The generation of optimal code for stack machines. Journal of the ACM, 22(3):382–396, July 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Berry:1979:MOC

[BL79] Gérard Berry and Jean-Jacques Lévy. Minimal and optimal computations of recursive programs. Journal of the ACM, 26(1):148– 175, January 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Baccelli:1990:EPP

[BL90] François Baccelli and Zhen Liu. On the execution of parallel programs on multiprocessor systems — a queuing theory approach. Journal of the ACM, 37(2):373–414, April 1990. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
77622.html.

Blumofe:1999:SMC

[BL99] Robert D. Blumofe and Charles E. Leiserson. Scheduling multithreaded computations by work stealing. Journal of the ACM, 46(5):720–748, September 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:

//www.acm.org/pubs/articles/journals/jacm/1999-46-5/
p720-blumofe/p720-blumofe.pdf;   http://www.acm.org/
pubs/citations/journals/jacm/1999-46-5/p720-blumofe/.

Benedikt:2000:RQI

[BL00] Michael Benedikt and Leonid Libkin. Relational queries over interpreted structures. Journal of the ACM, 47(4):644–680, July 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/citations/
journals/jacm/2000-47-4/p644-benedikt/.

Becchetti:2004:NSM

[BL04a] Luca Becchetti and Stefano Leonardi. Nonclairvoyant scheduling to minimize the total flow time on single and parallel machines. Journal of the ACM, 51(4):517–539, July 2004. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Burgisser:2004:LBB

[BL04b] Peter Bürgisser and Martin Lotz. Lower bounds on the bounded coefficient complexity of bilinear maps. Journal of the ACM, 51(3):464–482, May 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Beame:2019:TVN

[BL19] Paul Beame and Vincent Liew. Toward verifying nonlinear integer arithmetic. Journal of the ACM, 66(3):22:1–22:??, June 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=
3319396.

Blanyer:1955:PMD

[Bla55] Carl G. Blanyer. Precision modulators and demodulators. Journal of the ACM, 2(4):229–242, October 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Blair:1959:CTM

[Bla59] Charles R. Blair. On computer transcription of manual Morse. Journal of the ACM, 6(3):429–442, July 1959. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Blanc:2025:SSA

[Bla25] Guy Blanc. Subsampling suffices for adaptive data analysis. Journal of the ACM, 72(1):2:1–2:??, February 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3698104.

Bledsoe:1966:SRM

[Ble66] W. W. Bledsoe. Some results on multicategory pattern recognition. Journal of the ACM, 13(2):304–316, April 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Blikle:1972:AUC

[Bli72] Andrzej Blikle. Addressless units for carrying out loop-free computations. Journal of the ACM, 19(1):136–157, January 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brafman:1997:ALK

[BLMS97] Ronen I. Brafman, Jean-Claude Latombe, Yoram Moses, and Yoav Shoham. Applications of a logic of knowledge to motion planning under uncertainty. Journal of the ACM, 44(5):633–668, September 1997. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org:80/pubs/
citations/journals/jacm/1997-44-5/p633-brafman/.

Babaioff:2009:SVC

[BLP09] Moshe Babaioff, Ron Lavi, and Elan Pavlov. Single-value combinatorial auctions and algorithmic implementation in undominated strategies. Journal of the ACM, 56(1):4:1–4:32, January 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Barcelo:2010:XII

[BLPS10] Pablo Barceló, Leonid Libkin, Antonella Poggi, and Cristina Sirangelo. XML with incomplete information. Journal of the ACM, 58(1):4:1–4:62, December 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Blanc:2022:PLD

[BLQT22] Guy Blanc, Jane Lange, Mingda Qiao, and Li-Yang Tan. Properly learning decision trees in almost polynomial time. Journal of the ACM, 69(6):39:1–39:??, December 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3561047.

Biswal:2010:EBS

[BLR10] Punyashloka Biswal, James R. Lee, and Satish Rao. Eigenvalue bounds, spectral partitioning, and metrical deformations via flows. Journal of the ACM, 57(3):13:1–13:23, March 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Blum:2013:LTA

[BLR13] Avrim Blum, Katrina Ligett, and Aaron Roth. A learning theory approach to noninteractive database privacy. Journal of the ACM, 60(2):12:1–12:25, April 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Barcelo:2014:QRG

[BLR14] Pablo Barceló, Leonid Libkin, and Juan L. Reutter. Querying regular graph patterns. Journal of the ACM, 61(1):8:1–8:??, January 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Borodin:1992:OLA

[BLS92] Allan Borodin, Nathan Linial, and Michael E. Saks. An optimal on-line algorithm for metrical task system. Journal of the ACM, 39(4):745–763, October 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/146588.html.

Benedikt:2003:DRF

[BLSS03] Michael Benedikt, Leonid Libkin, Thomas Schwentick, and Luc Segoufin. Definable relations and first-order query languages over strings. Journal of the ACM, 50(5):694–751, September 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Baccelli:1993:ESP

[BLT93] François Baccelli, Zhen Liu, and Don Towsley. Extremal scheduling of parallel processing with and without real-time constraints. Journal of the ACM, 40(5):1209–1237, November 1993. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
169745.html.

Blum:1994:LAS

[BLTY94] Avrim Blum, Ming Li, John Tromp, and Mihalis Yannakakis. Linear approximation of shortest superstrings. Journal of the ACM, 41(4):630–647, July 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/179818.html.

Blumenthal:1958:DMF

[Blu58] Sherman Blumenthal. A dual master file system for a tape processing computer. Journal of the ACM, 5(4):319–327, October 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Blum:1959:EDF

[Blu59] Marvin Blum. On exponential digital filters. Journal of the ACM, 6(2):283–304, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Blum:1966:FSD

[Blu66] E. K. Blum. A formal system of differentiation. Journal of the ACM, 13(4):495–504, October 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Blum:1967:MIT

[Blu67] Manuel Blum. A machine-independent theory of the complexity of recursive functions. Journal of the ACM, 14(2):322–336, April 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Blum:1971:EPS

[Blu71] Manuel Blum. On effective procedures for speeding up algorithms. Journal of the ACM, 18(2):290–305, April 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Blum:1994:NAA

[Blu94] Avrim Blum. New approximation algorithms for graph coloring. Journal of the ACM, 41(3):470–516, May 1994. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
176586.html.

Brooker:1962:GTP

[BM62] R. A. Brooker and D. Morris. A general translation program for phrase structure languages. Journal of the ACM, 9(1):1– 10, January 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bavel:1970:CRA

[BM70] Zamir Bavel and David E. Muller. Connectivity and reversibility in automata. Journal of the ACM, 17(2):231–240, April 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Boyer:1975:PTA

[BM75] Robert S. Boyer and J. Strother Moore. Proving theorems about LISP functions. Journal of the ACM, 22(1):129–144, January 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bader:1982:GOL

[BM82] Christopher Bader and Arnaldo Moura. A generalization of Ogden’s lemma. Journal of the ACM, 29(2):404–406, April 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bagchi:1983:SAU

[BM83] A. Bagchi and A. Mahanti. Search algorithms under different kinds of heuristics–A comparative study. Journal of the ACM, 30(1):1–21, January 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Boyer:1984:MPU

[BM84] Robert S. Boyer and J. Strother Moore. A mechanical proof of the unsolvability of the halting problem. Journal of the ACM, 31(3):441–458, July 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bagchi:1985:TAH

[BM85] A. Bagchi and A. Mahanti. Three approaches to heuristic search in networks. Journal of the ACM, 32(1):1–27, January 1985. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
2458.html.

Bellare:1992:HSG

[BM92] Mihir Bellare and Silvio Micali. How to sign given any trapdoor permutation. Journal of the ACM, 39(1):214–233, January 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/147537.html.

Busch:1996:CTB

[BM96] Costas Busch and Marios Mavronicolas. A combinatorial treatment of balancing networks. Journal of the ACM, 43(5):794–839, September 1996. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/jacm/234754.html.

Buchin:2011:DTT

[BM11] Kevin Buchin and Wolfgang Mulzer. Delaunay triangulations in O(sort(n)) time and more. Journal of the ACM, 58(2):6:1– 6:27, April 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bodirsky:2018:DTC

[BMM18] Manuel Bodirsky, Barnaby Martin, and Antoine Mottet. Discrete temporal constraint satisfaction problems. Journal of the ACM, 65(2):9:1–9:??, March 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Balas:1991:PSA

[BMPT91] Egon Balas, Donald Miller, Joseph Pekny, and Paolo Toth. A parallel shortest augmenting path algorithm for the assignment problem. Journal of the ACM, 38(4):985–1004, October 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/115349.html.

Bistarelli:1997:SBC

[BMR97] Stefano Bistarelli, Ugo Montanari, and Francesca Rossi. Semiring-based constraint satisfaction and optimization. Journal of the ACM, 44(2):201–236, March 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bojanczyk:2009:TVL

[BMSS09] Mikoaj Bojańczyk, Anca Muscholl, Thomas Schwentick, and Luc Segoufin. Two-variable logic on data trees and XML reasoning. Journal of the ACM, 56(3):13:1–13:48, May 2009. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Baccelli:1989:AFJ

[BMT89] François Baccelli, William A. Massey, and Don Towsley. Acyclic fork-join queueing networks. Journal of the ACM, 36(3):615– 642, July 1989. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/65957.html.

Beaudry:1992:MPA

[BMT92] M. Beaudry, P. McKenzie, and D. Thérien. The membership problem in aperiodic transformation monoids. Journal of the ACM, 39(3):599–616, July 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/146661.html.

Bose:1962:SP

[BN62] R. C. Bose and R. J. Nelson. A sorting problem. Journal of the ACM, 9(2):282–296, April 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bhat:1971:BPA

[BN71] U. Narayan Bhat and Richard E. Nance. Busy period analysis of a time-sharing system modeled as a semi-Markov process. Journal of the ACM, 18(2):221–238, April 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bhat:1979:ECE

[BN79] U. Narayan Bhat and Richard E. Nance. An evaluation of CPU efficiency under dynamic quantum allocation. Journal of the ACM, 26(4):761–778, October 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bazzi:2001:SFT

[BN01] Rida A. Bazzi and Gil Neiger. Simplifying fault-tolerance: providing the abstraction of crash failures. Journal of the ACM, 48 (3):499–554, May 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/citations/journals/jacm/2001-48-3/p499-bazzi/.

Banerjee:2005:OCE

[BN05] Anindya Banerjee and David A. Naumann. Ownership confinement ensures representation independence for object-oriented programs. Journal of the ACM, 52(6):894–960, November 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Banerjee:2013:LRGb

[BN13] Anindya Banerjee and David A. Naumann. Local reasoning for global invariants, Part II: Dynamic boundaries. Journal of the ACM, 60(3):19:1–19:73, June 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Berkholz:2023:NOL

[BN23] Christoph Berkholz and Jakob Nordström. Near-optimal lower bounds on quantifier depth and Weisfeiler–Leman refinement steps. Journal of the ACM, 70(5):32:1–32:??, October 2023. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3195257.

Bar-Noy:2001:UAA

[BNBYF^{+}01] Amotz Bar-Noy, Reuven Bar-Yehuda, Ari Freund, Joseph (Seffi) Naor, and Baruch Schieber. A unified approach to approximating resource allocation and scheduling. Journal of the ACM, 48(5): 1069–1090, September 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bell:1994:MIP

[BNNS94] Colin Bell, Anil Nerode, Raymond T. Ng, and V. S. Subrahmanian. Mixed integer programming methods for computing nonmonotonic deductive databases. Journal of the ACM, 41(6): 1178–1215, November 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/195637.html. A prelim-

inary version is available as a University of Maryland TR Nr. CS-TR-2801, Dec 1991.

Banerjee:2013:LRGa

[BNR13] Anindya Banerjee, David A. Naumann, and Stan Rosenberg. Local reasoning for global invariants, Part I: Region logic. Journal of the ACM, 60(3):18:1–18:56, June 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bernstein:2025:NWS

[BNWN25] Aaron Bernstein, Danupon Nanongkai, and Christian Wulff-Nilsen. Negative-weight single-source shortest paths in nearlinear time. Journal of the ACM, 72(4):23:1–23:??, August 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bracken:1956:GSH

[BO56] Robert H. Bracken and Bruce G. Oldfield. A general system for handling alphameric information on the IBM 701 computer. Journal of the ACM, 3(3):175–180, July 1956. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bhargavan:2002:FVS

[BOG02] Karthikeyan Bhargavan, Davor Obradovic, and Carl A. Gunter. Formal verification of standards for distance vector routing protocols. Journal of the ACM, 49(4):538–576, July 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bojanczyk:1984:OAN

[Boj84] A. Bojańczyk. Optimal asynchronous Newton method for the solution of nonlinear equations. Journal of the ACM, 31(4):792– 803, October 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bolour:1979:OPM

[Bol79] Azad Bolour. Optimality properties of multiple-key hashing functions. Journal of the ACM, 26(2):196–210, April 1979. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bonzon:1970:NSC

[Bon70] P. Bonzon. Necessary and sufficient conditions for dynamic programming of combinatorial type. Journal of the ACM, 17(4): 675–682, October 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Booker:1965:NES

[Boo65] Aaron Booker. Numerical evaluation of symmetric polynomials. Journal of the ACM, 12(1):90–94, January 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Book:1978:SRC

[Boo78] Ronald V. Book. Simple representations of certain classes of languages. Journal of the ACM, 25(1):23–31, January 1978. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Book:1982:COT

[Boo82] Ronald V. Book. Confluent and other types of Thue systems. Journal of the ACM, 29(1):171–182, January 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Borodin:1972:CCE

[Bor72a] A. Borodin. Computational complexity and the existence of complexity gaps. Journal of the ACM, 19(1):158–174, January 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [Bor72b].

Borodin:1972:CCC

[Bor72b] A. Borodin. Corrigendum: “Computational complexity and the existence of complexity gaps”. Journal of the ACM, 19(3):576, July 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Bor72a].

Bottenbruch:1962:SUA

[Bot62] H. Bottenbruch. Structure and use of ALGOL 60. Journal of the ACM, 9(2):161–221, April 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Boyell:1957:PMI

[Boy57] Roger L. Boyell. Programmed multiplication on the IBM 407. Journal of the ACM, 4(4):442–449, October 1957. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Boyar:1989:ISP

[Boy89] Joan Boyar. Inferring sequences produced by pseudo-random number generators. Journal of the ACM, 36(1):129–141, January 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/59305.html;   http://www.imada.sdu.dk/~joan/

.

Boyland:2005:RAG

[Boy05] John Tang Boyland. Remote attribute grammars. Journal of the ACM, 52(4):627–687, July 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bilardi:1989:STC

[BP89] G. Bilardi and F. P. Preparata. Size-time complexity of Boolean networks for prefix computations. Journal of the ACM, 36 (2):362–382, April 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/62052.html.

Bergstra:2001:RMB

[BP01] Jan A. Bergstra and Alban Ponse. Register-machine based processes. Journal of the ACM, 48(6):1207–1241, November 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Beal:2003:GSR

[BP03a] Marie-Pierre Béal and Dominique Perrin. On the generating sequences of regular languages on k symbols. Journal of the ACM, 50(6):955–980, November 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bilardi:2003:ACS

[BP03b] Gianfranco Bilardi and Keshav Pingali. Algorithms for computing the static single assignment form. Journal of the ACM, 50 (3):375–425, May 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bojanczyk:2011:XEL

[BP11] Mikołaj Bojańczyk and Paweł Parys. XPath evaluation in linear time. Journal of the ACM, 58(4):17:1–17:33, July 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bodirsky:2015:STG

[BP15] Manuel Bodirsky and Michael Pinsker. Schaefer’s theorem for graphs. Journal of the ACM, 62(3):19:1–19:??, June 2015. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bodwin:2023:RSP

[BP23] Greg Bodwin and Merav Parter. Restorable shortest path tiebreaking for edge-faulty graphs. Journal of the ACM, 70 (5):28:1–28:??, October 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3603542.

Badanidiyuru:2012:TBN

[BPC^{+}12] Ashwinkumar Badanidiyuru, Arpita Patra, Ashish Choudhury, Kannan Srinathan, and C. Pandu Rangan. On the trade-off between network connectivity, round complexity, and communication complexity of reliable message transmission. Journal of the ACM, 59(5):22:1–22:35, October 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bilardi:2016:NOA

[BPP^{+}16] Gianfranco Bilardi, Andrea Pietracaprina, Geppino Pucci, Michele Scquizzato, and Francesco Silvestri. Network-oblivious algorithms. Journal of the ACM, 63(1):3:1–3:??, March 2016.

CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brock:1954:PAT

[BR54] Paul Brock and Sybil Rock. Problems in acceptance testing in digital computers. Journal of the ACM, 1(2):82–87, April 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Benschop:1971:MSE

[BR71] N. F. Benschop and H. C. Ratz. A mean square estimate of the generated roundoff error in constant matrix iterative processes. Journal of the ACM, 18(1):48–62, January 1971. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Berger:1991:SWI

[BR91] Bonnie Berger and John Rompel. Simulating (log^{c} n)-wise independence in NC. Journal of the ACM, 38(4):1026–1046, October 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/115347.html.

Basu:1996:CAC

[BR96] Saugata Basu and Marie-Françoise Roy. On the combinatorial and algebraic complexity of quantifier elimination. Journal of the ACM, 43(6):1002–1045, November 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/jacm/235813.html.

Brunsch:2015:ISA

[BR15] Tobias Brunsch and Heiko Röglin. Improved smoothed analysis of multiobjective optimization. Journal of the ACM, 62(1):4:1– 4:??, February 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bodirsky:2023:DCT

[BR23] Manuel Bodirsky and Jakub Rydval. On the descriptive complexity of temporal constraint satisfaction problems. Journal of the ACM, 70(1):2:1–2:??, February 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3566051.

Brans:1967:CPN

[Bra67] Carl H. Brans. A computer program for the nonnumerical testing and reduction of sets of algebraic partial differential equations. Journal of the ACM, 14(1):45–62, January 1967. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brandwajn:1977:QMM

[Bra77] Alexandre Brandwajn. A queueing model of multiprogrammed computer systems under full load conditions. Journal of the ACM, 24(2):222–240, April 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brand:1978:PCP

[Bra78] Daniel Brand. Path calculus in program verification. Journal of the ACM, 25(4):630–651, October 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bracha:1987:ERR

[Bra87] Gabriel Bracha. An O(log n) expected rounds randomized Byzantine generals protocol. Journal of the ACM, 34(4):910– 920, October 1987. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/42229.html.

Braverman:2010:PIF

[Bra10] Mark Braverman. Polylogarithmic independence fools AC^{0} circuits. Journal of the ACM, 57(5):28:1–28:10, June 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Breuer:1970:SCP

[Bre70] M. A. Breuer. Simplification of the covering problem with application to Boolean expressions. Journal of the ACM, 17(1): 166–181, January 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brent:1974:PEG

[Bre74] Richard P. Brent. The parallel evaluation of general arithmetic expressions. Journal of the ACM, 21(2):201–206, April 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brent:1976:FMP

[Bre76] Richard P. Brent. Fast multiple-precision evaluation of elementary functions. Journal of the ACM, 23(2):242–251, April 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Briggs:1990:CTC

[Bri90] David A. Briggs. A correction of the termination conditions of the Henschen-Naqvi technique. Journal of the ACM, 37(4):712– 719, October 1990. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/96562.html.

Batni:1974:EAF

[BRK74] Ramachendra P. Batni, Jeffrey D. Russell, and Charles R. Kime. An efficient algorithm for finding an irredundant set cover. Journal of the ACM, 21(3):351–355, July 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brown:1958:LT

[Bro58] A. F. R. Brown. Language translation. Journal of the ACM, 5(1):1–8, January 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brown:1964:TSA

[Bro64] R. R. Brown. Tape sets and automata. Journal of the ACM, 11 (1):10–14, January 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Broome:1965:DOS

[Bro65] Paul W. Broome. Discrete orthogonal sequences. Journal of the ACM, 12(2):151–168, April 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brown:1971:EAC

[Bro71] W. S. Brown. On Euclid’s algorithm and the computation of polynomial greatest common divisors. Journal of the ACM, 18 (4):478–504, October 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brown:1982:DCR

[Bro82] Theodore Brown. Determination of the conditional response for quantum allocation algorithms. Journal of the ACM, 29(2):448– 460, April 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Broy:1997:CRI

[Bro97] Manfred Broy. Compositional refinement of interactive systems. Journal of the ACM, 44(6):850–891, November 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org:80/pubs/citations/
journals/jacm/1997-44-6/p850-broy/.

Brooks:2003:TGC

[Bro03] Frederick P. Brooks, Jr. Three great challenges for half-century-old computer science. Journal of the ACM, 50(1):25–26, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Balkanski:2022:LOS

[BRS22] Eric Balkanski, Aviad Rubinstein, and Yaron Singer. The limitations of optimization from samples. Journal of the ACM, 69(3):21:1–21:33, June 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3511018.

Borodini:1997:HMC

[BRSU97] Allan Borodini, Prabhakar Raghavan, Baruch Schieber, and Eli Upfal. How much can hardware help routing? Journal of the ACM, 44(5):726–741, September 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:/

/www.acm.org:80/pubs/citations/journals/jacm/1997-44-
5/p726-borodini/.

Bruno:1976:SJS

[Bru76] John L. Bruno. Sequencing jobs with stochastic task structures on a single machine. Journal of the ACM, 23(4):655–664, October 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bryant:1982:MPR

[Bry82] R. M. Bryant. Maximum processing rates of memory bound systems. Journal of the ACM, 29(2):461–477, April 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bryant:1991:MHV

[Bry91] Randal E. Bryant. A methodology for hardware verification based on logic simulation. Journal of the ACM, 38(2):299– 328, April 1991. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/103519.html.

Brzozowski:1964:DRE

[Brz64] Janusz A. Brzozowski. Derivatives of regular expressions. Journal of the ACM, 11(4):481–494, October 1964. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brzozowski:1967:RSE

[Brz67] J. A. Brzozowski. Roots of star events. Journal of the ACM, 14(3):466–477, July 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Boehm:1959:SSM

[BS59] E. M. Boehm and T. B. Steel, Jr. The share 709 system: Machine implementation of symbolic programming. Journal of the ACM, 6(2):134–140, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bruno:1972:EAC

[BS72] J. Bruno and K. Steiglitz. The expression of algorithms by charts. Journal of the ACM, 19(3):517–525, July 1972. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Blevins:1974:CED

[BS74] M. M. Blevins and G. W. Stewart. Calculating the eigenvectors of diagonally dominant matrices. Journal of the ACM, 21 (2):261–271, April 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bruno:1976:CGO

[BS76] John Bruno and Ravi Sethi. Code generation for a one-register machine. Journal of the ACM, 23(3):502–510, July 1976. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Buckley:1985:BTP

[BS85] G. N. Buckley and A. Silberschatz. Beyond two-phase locking. Journal of the ACM, 32(2):314–326, April 1985. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
3151.html.

Baltsan:1988:SPB

[BS88] Avikam Baltsan and Micha Sharir. On the shortest paths between two convex polyhedra. Journal of the ACM, 35(2):267– 287, April 1988. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/214094.html.

Brzozowski:1989:UFR

[BS89] J. A. Brzozowski and C.-J. Seger. A unified framework for race analysis of asynchronous networks. Journal of the ACM, 36(1):20–45, January 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/59301.html.

Bubeck:2023:ULR

[BS23] Sébastien Bubeck and Mark Sellke. A universal law of robustness via isoperimetry. Journal of the ACM, 70(2):10:1–10:??, April 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3578580.

Ben-Sasson:2023:PGR

[BSCI^{+}23] Eli Ben-Sasson, Dan Carmon, Yuval Ishai, Swastik Kopparty, and Shubhangi Saraf. Proximity gaps for Reed–Solomon codes. Journal of the ACM, 70(5):31:1–31:??, October 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3614423.

Bshouty:1993:CFR

[Bsh93] Nader H. Bshouty. On the complexity of functions for random access machines. Journal of the ACM, 40(2):211–223, April 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/151262.html.

Ben-Sasson:2016:CRP

[BSKK^{+}16] Eli Ben-Sasson, Yohay Kaplan, Swastik Kopparty, Or Meir, and Henning Stichtenoth. Constant rate PCPs for circuit-SAT with sublinear query complexity. Journal of the ACM, 63(4):32:1– 32:??, November 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ben-Sasson:2014:ACA

[BSLRZ14] Eli Ben-Sasson, Shachar Lovett, and Noga Ron-Zewi. An additive combinatorics approach relating rank to communication complexity. Journal of the ACM, 61(4):22:1–22:??, July 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Becker:1982:SAM

[BSP82] Ronald I. Becker, Stephen R. Schach, and Yehoshua Perl. A shifting algorithm for min-max tree partitioning. Journal of the ACM, 29(1):58–67, January 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Beame:2003:TST

[BSSV03] Paul Beame, Michael Saks, Xiaodong Sun, and Erik Vee. Timespace trade-off lower bounds for randomized computation of decision problems. Journal of the ACM, 50(2):154–195, March 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bhargava:2020:DFS

[BSV20] Vishwas Bhargava, Shubhangi Saraf, and Ilya Volkovich. Deterministic factorization of sparse polynomials with bounded individual degree. Journal of the ACM, 67(2):8:1–8:28, May 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3365667.

Ben-Sasson:2001:SPN

[BSW01] Eli Ben-Sasson and Avi Wigderson. Short proofs are narrow— resolution made simple. Journal of the ACM, 48(2):149–169, March 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/pubs/

citations/journals/jacm/2001-48-2/p149-ben-sasson/.

Berkovits:1965:MMA

[BSZ65] S. Berkovits, M. Schlessing, and N. Zierler. On mappings for modular arithmetic, II. Journal of the ACM, 12(4):545–546, October 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Blum:2006:OAM

[BSZ06] Avrim Blum, Tuomas Sandholm, and Martin Zinkevich. Online algorithms for market clearing. Journal of the ACM, 53(5): 845–879, September 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bracken:1957:ISC

[BT57] R. H. Bracken and H. E. Tillitt. Information searching with the 701 calculator. Journal of the ACM, 4(2):131–136, April 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brown:1969:MSC

[BT69] W. S. Brown and J. F. Traub. MERCURY: a system for the computer-aided distribution of technical reports. Journal of the ACM, 16(1):13–25, January 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brown:1971:EAT

[BT71] W. S. Brown and J. F. Traub. On Euclid’s algorithm and the theory of subresultants. Journal of the ACM, 18(4):505–514, October 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brown:1979:FMA

[BT79] Mark R. Brown and Robert E. Tarjan. A fast merging algorithm. Journal of the ACM, 26(2):211–226, April 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bracha:1985:ACB

[BT85] Gabriel Bracha and Sam Toueg. Asynchronous consensus and broadcast protocols. Journal of the ACM, 32(4):824–840, October 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/214134.html.

Bergstra:1995:ESC

[BT95] J. A. Bergstra and J. V. Tucker. Equational specifications, complete term rewriting systems, and computable and semicomputable algebras. Journal of the ACM, 42(6):1194–1230, November 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/227687.html.

Bshouty:1996:FSM

[BT96] Nader H. Bshouty and Christino Tamon. On the Fourier spectrum of monotone functions. Journal of the ACM, 43(4):747– 770, July 1996. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/jacm/234564.html.

Brafman:2000:ATT

[BT00] Ronen I. Brafman and Moshe Tennenholtz. An axiomatic treatment of three qualitative decision criteria. Journal of the ACM, 47(3):452–482, May 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/citations/journals/jacm/2000-47-5/p452-brafman/.

Bergstra:2007:RNA

[BT07] J. A. Bergstra and J. V. Tucker. The rational numbers as an abstract data type. Journal of the ACM, 54(2):7:1–7:25, April 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Barany:2015:GN

[BTS15] Vince Bárány, Balder Ten Cate, and Luc Segoufin. Guarded negation. Journal of the ACM, 62(3):22:1–22:??, June 2015. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bui:1977:ECP

[Bui77] T. D. Bui. Errata and comments on a paper by J. R. Cash. Journal of the ACM, 24(4):623, October 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Cas76].

Bui:1979:SSS

[Bui79] T. D. Bui. Some A-stable and L-stable methods for the numerical integration of stiff ordinary differential equations. Journal of the ACM, 26(3):483–493, July 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bulatov:2006:DTC

[Bul06] Andrei A. Bulatov. A dichotomy theorem for constraint satisfaction problems on a 3-element set. Journal of the ACM, 53 (1):66–120, January 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bulatov:2013:CCC

[Bul13] Andrei A. Bulatov. The complexity of the counting constraint satisfaction problem. Journal of the ACM, 60(5):34:1–34:41, October 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bunch:1971:ESM

[Bun71] James R. Bunch. Equilibration of symmetric matrices in the max-norm. Journal of the ACM, 18(4):566–572, October 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Burks:1970:ECA

[Bur70] Arthur W. Burks. Essays on Cellular Automata. University of Illinois Press, Urbana, IL, USA, 1970. xxvi + 375 pp. LCCN QA267.5.S4 B87. A classic collection of papers on cellular automata intended as a companion to von Neumann’s Theory of Self Reproducing Automata [von66]. Includes papers by Burks, Thatcher, Moore, Myhill, Ulam, and Holland. See [Hol62b].

Burge:1976:ABS

[Bur76] William H. Burge. An analysis of binary search trees formed from sequences of nondistinct keys. Journal of the ACM, 23(3): 451–454, July 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Butcher:1965:MMM

[But65] J. C. Butcher. A modified multistep method for the numerical integration of ordinary differential equations. Journal of the ACM, 12(1):124–135, January 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Butcher:1967:MGR

[But67] John C. Butcher. A multistep generalization of Runge–Kutta methods with four or five stages. Journal of the ACM, 14(1):84– 99, January 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Butler:1978:ADF

[But78] Jon T. Butler. Analysis and design of fanout-free networks of positive symmetric gates. Journal of the ACM, 25(3):481–498, July 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Butcher:1979:TIR

[But79] J. C. Butcher. A transformed implicit Runge–Kutta method. Journal of the ACM, 26(4):731–738, October 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Beeri:1984:PPD

[BV84] Catriel Beeri and Moshe Y. Vardi. A proof procedure for data dependencies. Journal of the ACM, 31(4):718–741, October 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bertsimas:2004:SCP

[BV04] Dimitris Bertsimas and Santosh Vempala. Solving convex programs by random walks. Journal of the ACM, 51(4):540–556, July 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bitansky:2018:IOF

[BV18] Nir Bitansky and Vinod Vaikuntanathan. Indistinguishability obfuscation from functional encryption. Journal of the ACM, 65 (6):39:1–39:??, November 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/ft_gateway.cfm?id=3234511.

Bauer:1957:SGP

[BW57a] Walter F. Bauer and George P. West. A system for generalpurpose analog-digital computation. Journal of the ACM, 4 (1):12–17, January 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Burks:1957:LAPa

[BW57b] Arthur W. Burks and Hao Wang. The logic of automata, Part I. Journal of the ACM, 4(2):193–218, April 1957. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Burks:1957:LAPb

[BW57c] Arthur W. Burks and Hao Wang. The logic of automata, Part II. Journal of the ACM, 4(3):279–297, July 1957. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bambos:1991:SPP

[BW91] Nicholas Bambos and Jean Walrand. On stability and performance of parallel processing systems. Journal of the ACM, 38 (2):429–452, April 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/103520.html.

Bell:1994:RBG

[BW94] Timothy C. Bell and Ian H. Witten. The relationship between greedy parsing and symbolwise text compression. Journal of the ACM, 41(4):708–724, July 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/179892.html.

Bass:1973:OHN

[BY73] Leonard Bass and Paul Young. Ordinal hierarchies and naming complexity classes. Journal of the ACM, 20(4):668–686, October 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Boyer:1996:APO

[BY96] Robert S. Boyer and Yuan Yu. Automated proofs of object code for a widely used microprocessor. Journal of the ACM, 43(1):166–192, January 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/jacm/227603.html.

Bar-Yehuda:2017:DAV

[BYCHS17] Reuven Bar-Yehuda, Keren Censor-Hillel, and Gregory Schwartzman. A distributed (2 + ϵ)-approximation for vertex cover in O(log Δ/ϵ log log Δ) rounds. Journal of the ACM, 64(3):23:1– 23:??, June 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Baeza-Yates:1996:FTS

[BYG96] Richardo A. Baeza-Yates and Gaston H. Gonnet. Fast text searching for regular expressions or automaton searching on tries. Journal of the ACM, 43(6):915–936, November 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
jacm/235810.html.

Bar-Yossef:2008:RSS

[BYG08] Ziv Bar-Yossef and Maxim Gurevich. Random sampling from a search engine’s index. Journal of the ACM, 55(5):24:1–24:74, October 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brand:1983:CFS

[BZ83] Daniel Brand and Pitro Zafiropulo. On communicating finitestate machines. Journal of the ACM, 30(2):323–342, April 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ben-Zvi:2014:BLH

[BZ14] Ido Ben-Zvi. Beyond Lamport’s happened-before: On time bounds and the ordering of events in distributed systems. Journal of the ACM, 61(2):13:1–13:??, April 2014. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Bauwens:2023:UAO

[BZ23] Bruno Bauwens and Marius Zimand. Universal almost optimal compression and Slepian–Wolf coding in probabilistic polynomial time. Journal of the ACM, 70(2):9:1–9:??, April 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3575807.

Cotten:1974:PTS

[CAA74] L. W. Cotten and A. M. Abd-Alla. Processing times for segmented jobs with I/O compute overlap. Journal of the ACM, 21 (1):18–30, January 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cohen-Addad:2024:FDT

[CADK^{+}24] Vincent Cohen-Addad, Debarati Das, Evangelos Kipouridis, Nikos Parotsidis, and Mikkel Thorup. Fitting distances by tree metrics minimizing the total error within a constant factor. Journal of the ACM, 71(2):10:1–10:??, April 2024. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3639453.

Cohen-Addad:2021:ATL

[CADMD21] Vincent Cohen-Addad, Éric Colin De Verdière, Dániel Marx, and Arnaud De Mesmay. Almost tight lower bounds for hard cutting problems in embedded graphs. Journal of the ACM, 68 (4):30:1–30:26, August 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3450704.

Cohen-Addad:2021:NLT

[CAFS21] Vincent Cohen-Addad, Andreas Emil Feldmann, and David Saulpic. Near-linear time approximation schemes for clustering in doubling metrics. Journal of the ACM, 68(6):44:1–44:34, December 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3477541.

Cohen-addad:2019:HCO

[CaKMTM19] Vincent Cohen-addad, Varun Kanade, Frederik Mallmann-Trenn, and Claire Mathieu. Hierarchical clustering: Objective functions and algorithms. Journal of the ACM, 66(4): 26:1–26:??, August 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3321386.

Caldwell:1959:NDM

[Cal59] George C. Caldwell. A note on the downhill method. Journal of the ACM, 6(2):223–225, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Calingaert:1961:TDP

[Cal61] Peter Calingaert. Two-dimensional parity checking. Journal of the ACM, 8(2):186–200, April 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cantor:1962:APB

[Can62] David G. Cantor. On the ambiguity problem of Backus systems. Journal of the ACM, 9(4):477–479, October 1962. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Canetti:2020:UCS

[Can20] Ran Canetti. Universally composable security. Journal of the ACM, 67(5):28:1–28:94, October 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/doi/10.1145/3402457.

Caplain:2002:CPS

[Cap02] Gilbert Caplain. Correctness properties in a shared-memory parallel language. Journal of the ACM, 49(6):785–827, November 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Carr:1957:IPA

[Car57] John W. Carr III. Inaugural presidential address. Journal of the ACM, 4(1):5–7, January 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Carr:1958:EBR

[Car58] John W. Carr III. Error bounds for the Runge–Kutta single-step integration process. Journal of the ACM, 5(1):39–44, January 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Case:1971:NDS

[Cas71] John Case. A note on degrees of self-describing Turing machines. Journal of the ACM, 18(3):329–338, July 1971. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cash:1975:CIR

[Cas75] J. R. Cash. A class of implicit Runge–Kutta methods for the numerical integration of stiff ordinary differential equations. Journal of the ACM, 22(4):504–511, October 1975. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cash:1976:SIR

[Cas76] J. R. Cash. Semi-implicit Runge–Kutta procedures with error estimates for the numerical integration of stiff systems of ordinary differential equations. Journal of the ACM, 23(3):455–460, July 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See comments and errata in [Bui77].

Causey:1958:SEB

[Cau58] Robert L. Causey. On some error bounds of Givens. Journal of the ACM, 5(2):127–131, April 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Caviness:1970:CFS

[Cav70] B. F. Caviness. On canonical forms and simplification. Journal of the ACM, 17(2):385–396, April 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Constable:1972:SPL

[CB72] Robert L. Constable and Allan B. Borodin. Subrecursive programming languages, Part I: Efficiency and program structure. Journal of the ACM, 19(3):526–568, July 1972. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cesa-Bianchi:1999:SES

[CBDF^{+}99] Nicolò Cesa-Bianchi, Eli Dichterman, Paul Fischer, Eli Shamir, and Hans Ulrich Simon. Sample-efficient strategies for learning in the presence of noise. Journal of the ACM, 46(5): 684–719, September 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/articles/journals/jacm/1999-46-5/p684-cesa-
bianchi/p684-cesa-bianchi.pdf;   http://www.acm.org/
pubs/citations/journals/jacm/1999-46-5/p684-cesa-bianchi/

.

Cesa-Bianchi:1997:HUE

[CBFH^{+}97] Nicolò Cesa-Bianchi, Yoav Freund, David Haussler, David P. Helmbold, Robert E. Schapire, and Manfred K. Warmuth. How to use expert advice. Journal of the ACM, 44(3):427–485, May 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/citations/
journals/jacm/1997-44-3/p427-cesa-bianchi/.

Campbell:1961:NCT

[CBHH61] Edwin S. Campbell, R. Buehler, J. O. Hirschfelder, and D. Hughes. Numerical construction of Taylor series approximations for a set of simultaneous first order differential equations. Journal of the ACM, 8(3):374–383, July 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chu:1967:EPD

[CC67] J. T. Chu and J. C. Chueh. Error probability in decision functions for character recognition. Journal of the ACM, 14(2):273– 280, April 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cody:1976:ERA

[CC76a] R. A. Cody and E. G. Coffman, Jr. Errata: “Record allocation for minimizing expected retrieval costs on drum-like storage devices”. Journal of the ACM, 23(4):572, October 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [CC76b].

Cody:1976:RAM

[CC76b] R. A. Cody and E. G. Coffman, Jr. Record allocation for minimizing expected retrieval costs on drum-like storage devices. Journal of the ACM, 23(1):103–115, January 1976. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [CC76a].

Conforti:1995:CLP

[CC95] Michele Conforti and Gérard Cornuéjols. A class of logic problems solvable by linear programming. Journal of the ACM, 42 (5):1107–1113, September 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/210137.html.

Chen:2004:DDN

[CC04] Chung-Min Chen and Christine T. Cheng. From discrepancy to declustering: Near-optimal multidimensional declustering strategies for range queries. Journal of the ACM, 51(1):46–73, January 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chekuri:2016:PBG

[CC16] Chandra Chekuri and Julia Chuzhoy. Polynomial bounds for the grid-minor theorem. Journal of the ACM, 63(5):40:1–40:??, December 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cai:2017:CCC

[CC17] Jin-Yi Cai and Xi Chen. Complexity of counting CSP with complex weights. Journal of the ACM, 64(3):19:1–19:??, June 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Calderbank:1984:OHS

[CCF84] A. R. Calderbank, E. G. Coffman, Jr., and L. Flatto. Optimum head separation in a disk system with two read/write heads. Journal of the ACM, 31(4):826–838, October 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Calderbank:1988:ODP

[CCF88] A. R. Calderbank, E. G. Coffman, Jr., and Leopold Flatto. Optimal directory placement on disk storage devices. Journal of the ACM, 35(2):433–446, April 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/42287.html.

Chia:2023:NLQ

[CCL23] Nai-Hui Chia, Kai-Min Chung, and Ching-Yi Lai. On the need for large quantum depth. Journal of the ACM, 70(1):6:1–6:??, February 2023. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL https://dl.acm.org/doi/10.
1145/3570637.

Cousot:2012:TSS

[CCM12] Patrick Cousot, Radhia Cousot, and Laurent Mauborgne. Theories, solvers and static analysis by abstract interpretation. Journal of the ACM, 59(6):31:1–31:56, December 2012. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). Special issue text supplied by Joanne Pello.

Chan:2022:FDO

[CCMS22] T.-H. Hubert Chan, Kai-Min Chung, Bruce Maggs, and Elaine Shi. Foundations of differentially oblivious algorithms. Journal of the ACM, 69(4):27:1–27:??, August 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3555984.

Cassaigne:2014:ATC

[CCSS14] Julien Cassaigne, James D. Currie, Luke Schaeffer, and Jeffrey Shallit. Avoiding three consecutive blocks of the same size and same sum. Journal of the ACM, 61(2):10:1–10:??, April 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Conte:1960:ADM

[CD60] Samuel D. Conte and Ralph T. Dames. On an alternating direction method for solving the plate problem with mixed boundary conditions. Journal of the ACM, 7(3):264–273, July 1960. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cabay:1987:SLE

[CD87a] Stanley Cabay and Bart Domzy. Systems of linear equations with dense univariate polynomial coefficients. Journal of the ACM, 34(3):646–660, July 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/28877.html.

Chazelle:1987:ICO

[CD87b] B. Chazelle and D. P. Dobkin. Intersection of convex objects in two and three dimensions. Journal of the ACM, 34(1):1– 27, January 1987. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/24036.html.

Cosnard:1994:OAP

[CD94] Michel Cosnard and El Mostafa Daoudi. Optimal algorithms for parallel Givens factorization on a coarse-grained PRAM. Journal of the ACM, 41(2):399–421, March 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/174660.html.

Chen:2008:MAB

[CD08] Xi Chen and Xiaotie Deng. Matching algorithmic bounds for finding a Brouwer fixed point. Journal of the ACM, 55(3):13:1– 13:23, July 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Czumaj:2021:EST

[CD21] Artur Czumaj and Peter Davies. Exploiting spontaneous transmissions for broadcasting and leader election in radio networks. Journal of the ACM, 68(2):13:1–13:22, March 2021. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3446383.

Chewi:2024:QLB

[CdDPL^{+}24] Sinho Chewi, Jaume de Dios Pont, Jerry Li, Chen Lu, and Shyam Narayanan. Query lower bounds for log-concave sampling. Journal of the ACM, 71(4):29:1–29:??, August 2024. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3673651.

Cheng:2000:SE

[CDE^{+}00] Siu-Wing Cheng, Tamal K. Dey, Herbert Edelsbrunner, Michael A. Facello, and Shang-Hua Teng. Silver exudation. Journal of the ACM, 47(5):883–904, 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:

//www.acm.org/pubs/citations/journals/jacm/2000-47-5/
p883-cheng/.

Chakraborty:2020:AED

[CDG^{+}20] Diptarka Chakraborty, Debarati Das, Elazar Goldenberg, Michal Koucký, and Michael Saks. Approximating edit distance within constant factor in truly sub-quadratic time. Journal of the ACM, 67(6):36:1–36:22, November 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/doi/10.1145/3422823.

Collins:1992:VDD

[CDMP92] Oliver Collins, Sam Dolinar, Robert McEliece, and Fabrizio Pollara. A VLSI decomposition of the deBruijn graph. Journal of the ACM, 39(4):931–948, October 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/146620.html.

Calcagno:2011:CSA

[CDOY11] Cristiano Calcagno, Dino Distefano, Peter W. O’Hearn, and Hongseok Yang. Compositional shape analysis by means of biabduction. Journal of the ACM, 58(6):26:1–26:66, December 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chaput:2014:AMP

[CDPP14] Philippe Chaput, Vincent Danos, Prakash Panangaden, and Gordon Plotkin. Approximating Markov processes by averaging. Journal of the ACM, 61(1):5:1–5:??, January 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cotumaccio:2023:CLO

[CDPP23] Nicola Cotumaccio, Giovanna D’Agostino, Alberto Policriti, and Nicola Prezza. Co-lexicographically ordering automata and regular languages — part I. Journal of the ACM, 70(4):27:1–27:??, August 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3607471.

Coppersmith:1993:RWW

[CDRS93] Don Coppersmith, Peter Doyle, Prabhakar Raghavan, and Marc Snir. Random walks on weighted graphs and applications to on-line algorithms. Journal of the ACM, 40(3):421–453, July 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/174131.html.

Chen:2023:RIC

[CDRS23] Ho-Lin Chen, David Doty, Wyatt Reeves, and David Soloveichik. Rate-independent computation in continuous chemical reaction networks. Journal of the ACM, 70(3):22:1–22:??, June 2023. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3590776.

Chen:2009:SCC

[CDT09] Xi Chen, Xiaotie Deng, and Shang-Hua Teng. Settling the complexity of computing two-player Nash equilibria. Journal of the ACM, 56(3):14:1–14:57, May 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cramer:2021:MSV

[CDW21] Ronald Cramer, Léo Ducas, and Benjamin Wesolowski. Mildly short vectors in cyclotomic ideal lattices in quantum polynomial time. Journal of the ACM, 68(2):8:1–8:26, March 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3431725.

Chen:1972:MRM

[CE72] Y. E. Chen and D. L. Epley. Memory requirements in a multiprocessing environment. Journal of the ACM, 19(1):57–69, January 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chazelle:1992:OAI

[CE92] Bernard Chazelle and Herbert Edelsbrunner. An optimal algorithm for intersecting line segments in the plane. Journal of the ACM, 39(1):1–54, January 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/147511.html.

Canny:2000:SBA

[CE00] John F. Canny and Ioannis Z. Emiris. A subdivision-based algorithm for the sparse resultant. Journal of the ACM, 47 (3):417–451, May 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/citations/journals/jacm/2000-47-3/p417-canny/.

Coleman:1986:PFS

[CEG86] Thomas F. Coleman, Anders Edenbrandt, and John R. Gilbert. Predicting fill for sparse orthogonal factorization. Journal of the ACM, 33(3):517–532, July 1986. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/5932.html. Cited

in Åke Björck’s bibliography on least squares, which is avail-

able by anonymous ftp from math.liu.se in pub/references.

Chaudhuri:2000:TBS

[CeLT00] Soma Chaudhuri, Maurice erlihy, Nancy A. Lynch, and Mark R. Tuttle. Tight bounds for k-set agreement. Journal of the ACM, 47(5):912–943, 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:

//www.acm.org/pubs/articles/journals/jacm/2000-47-5/
p912-chaudhuri/p912-chaudhuri.pdf;   http://www.acm.
org/pubs/citations/journals/jacm/2000-47-5/p912-chaudhuri/

.

Certaine:1958:SPR

[Cer58] J. Certaine. On sequences of pseudo-random numbers of maximal length. Journal of the ACM, 5(4):353–356, October 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Copi:1958:REL

[CEW58] Irving M. Copi, Calvin C. Elgot, and Jesse B. Wright. Realization of events by logical nets. Journal of the ACM, 5(2):181–196, April 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Curtis:1959:ADP

[CF59] Philip C. Curtis, Jr. and Werner L. Frank. An algorithm for the determination of the polynomial of best minimax approximation to a function defined on a finite point set. Journal of the ACM, 6(3):395–404, July 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chartres:1968:USD

[CF68] B. A. Chartres and J. J. Florentin. A universal syntax-directed top-down analyzer. Journal of the ACM, 15(3):447–464, July 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chen:2012:AOA

[CF12] Yijia Chen and Jörg Flum. From almost optimal algorithms to logics for complexity classes via listings and a halting problem. Journal of the ACM, 59(4):17:1–17:34, August 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chua:2020:SGF

[CF20] Jeroen Chua and Pedro F. Felzenszwalb. Scene grammars, factor graphs, and belief propagation. Journal of the ACM, 67(4):19:1– 19:41, August 2020. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL https://dl.acm.org/doi/abs/
10.1145/3396886.

Cygan:2017:TLB

[CFG^{+}17] Marek Cygan, Fedor V. Fomin, Alexander Golovnev, Alexander S. Kulikov, Ivan Mihajlin, Jakub Pachocki, and Arkadiusz Socala. Tight lower bounds on graph embedding problems. Journal of the ACM, 64(3):18:1–18:??, June 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chiesa:2022:SII

[CFGS22] Alessandro Chiesa, Michael A. Forbes, Tom Gur, and Nicholas Spooner. Spatial isolation implies zero knowledge even in a quantum world. Journal of the ACM, 69(2):15:1–15:44, April 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3511100.

Chebolu:2010:FMM

[CFM10] Prasad Chebolu, Alan Frieze, and P’all Melsted. Finding a maximum matching in a sparse random graph in O(n) expected time. Journal of the ACM, 57(4):24:1–24:27, April 2010. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chung:1957:TIC

[CG57] J. H. Chung and C. C. Gotlieb. Test of an inventory control system on FERUT. Journal of the ACM, 4(2):121–130, April 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chartres:1967:CEB

[CG67] Bruce A. Chartres and James C. Geuder. Computable error bounds for direct solution of linear equations. Journal of the ACM, 14(1):63–71, January 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chang:1970:ASC

[CG70a] S. K. Chang and A. Gill. Algorithmic solution of the changemaking problem. Journal of the ACM, 17(1):113–122, January 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Corneil:1970:EAG

[CG70b] D. G. Corneil and C. C. Gotlieb. An efficient algorithm for graph isomorphism. Journal of the ACM, 17(1):51–64, January 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Conway:1986:RNE

[CG86] A. E. Conway and N. D. Georganas. RECAL–A new efficient algorithm for the exact analysis of multiple-chain queuing networks. Journal of the ACM, 33(4):768–791, October 1986. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
6495.html.

Coffman:1993:PCP

[CG93] E. G. Coffman, Jr. and M. R. Garey. Proof of the 4/3 conjecture for preemptive vs. nonpreemptive two-processor scheduling. Journal of the ACM, 40(5):991–1018, November 1993. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
174148.html.

Cerone:2018:ASI

[CG18] Andrea Cerone and Alexey Gotsman. Analysing snapshot isolation. Journal of the ACM, 65(2):11:1–11:??, March 2018. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Clarke:1983:EAH

[CGH83] Edmund M. Clarke, Jr., Steven M. German, and Joseph Y. Halpern. Effective axiomatizations of Hoare logics. Journal of the ACM, 30(3):612–636, July 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Canetti:2004:ROM

[CGH04] Ran Canetti, Oded Goldreich, and Shai Halevi. The random oracle methodology, revisited. Journal of the ACM, 51(4):557– 594, July 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chuzhoy:2005:ACH

[CGH^{+}05] Julia Chuzhoy, Sudipto Guha, Eran Halperin, Sanjeev Khanna, Guy Kortsarz, Robert Krauthgamer, and Joseph (Seffi) Naor. Asymmetric k-center is log^{∗} n-hard to approximate. Journal of the ACM, 52(4):538–551, July 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Clarke:2003:CGA

[CGJ^{+}03] Edmund Clarke, Orna Grumberg, Somesh Jha, Yuan Lu, and Helmut Veith. Counterexample-guided abstraction refinement for symbolic model checking. Journal of the ACM, 50(5):752– 794, September 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chia:2022:SBS

[CGL^{+}22] Nai-Hui Chia, András Pal Gilyén, Tongyang Li, Han-Hsuan Lin, Ewin Tang, and Chunhao Wang. Sampling-based sublinear lowrank matrix arithmetic framework for dequantizing quantum machine learning. Journal of the ACM, 69(5):33:1–33:??, October 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3549524.

Charalampopoulos:2023:AOE

[CGL^{+}23] Panagiotis Charalampopoulos, Paweł Gawrychowski, Yaowei Long, Shay Mozes, Seth Pettie, Oren Weimann, and Christian Wulff-Nilsen. Almost optimal exact distance oracles for planar graphs. Journal of the ACM, 70(2):12:1–12:??, April 2023. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3580474.

Chierichetti:2018:RSC

[CGLP18] Flavio Chierichetti, George Giakkoupis, Silvio Lattanzi, and Alessandro Panconesi. Rumor spreading and conductance. Journal of the ACM, 65(4):17:1–17:??, August 2018. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chaudhury:2024:EET

[CGM24] Bhaskar Ray Chaudhury, Jugal Garg, and Kurt Mehlhorn. EFX exists for three agents. Journal of the ACM, 71(1):4:1–4:??, February 2024. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL https://dl.acm.org/doi/10.
1145/3616009.

Chazal:2013:PBC

[CGOS13] Frédéric Chazal, Leonidas J. Guibas, Steve Y. Oudot, and Primoz Skraba. Persistence-based clustering in Riemannian manifolds. Journal of the ACM, 60(6):41:1–41:??, November 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chen:2002:MG

[CGP02] Zhi-Zhong Chen, Michelangelo Grigni, and Christos H. Papadimitriou. Map graphs. Journal of the ACM, 49(2):127–138, March 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cygan:2015:AAB

[CGS15] Marek Cygan, Harold N. Gabow, and Piotr Sankowski. Algorithmic applications of Baur–Strassen’s Theorem: Shortest cycles, diameter, and matchings. Journal of the ACM, 62(4):28:1– 28:??, August 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chou:2024:SAA

[CGSV24] Chi-Ning Chou, Alexander Golovnev, Madhu Sudan, and Santhoshini Velusamy. Sketching approximability of all finite CSPs. Journal of the ACM, 71(2):15:1–15:??, April 2024. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3649435.

Chan:2010:ULD

[CGT10] T.-H. Hubert Chan, Anupam Gupta, and Kunal Talwar. Ultralow-dimensional embeddings for doubling metrics. Journal of the ACM, 57(4):21:1–21:26, April 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cohen:1979:TAD

[CH79] Jacques Cohen and Timothy Hickey. Two algorithms for determining volumes of convex polyhedra. Journal of the ACM, 26(3): 401–414, July 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Culik:1984:SEP

[CH84] Karel Culik, II and Tero Harju. The ω-sequence equivalence problem for D0L systems is decidable. Journal of the ACM, 31 (2):282–298, April 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Collings:1986:IGF

[CH86] Bruce Jay Collings and G. Barry Hembree. Initializing generalized feedback shift register pseudorandom number generators. Journal of the ACM, 33(4):706–711, October 1986. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/

6493.html. See also [CH88].

Collings:1988:AIG

[CH88] Bruce Jay Collings and G. Barry Hembree. Addendum to “initializing generalized feedback shift register pseudorandom number generators”. Journal of the ACM, 35(4):1001, October 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [CH86].

Castellani:1989:DB

[CH89] Ilaria Castellani and Matthew Hennessy. Distributed bisimulations. Journal of the ACM, 36(4):887–911, October 1989. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
76369.html.

Cockett:1990:DTR

[CH90] J. R. B. Cockett and J. A. Herrera. Decision tree reduction. Journal of the ACM, 37(4):815–842, October 1990. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
96576.html.

Chan:1991:IRD

[CH91a] Edward P. F. Chan and Héctor J. Hernández. Independencereducible database schemes. Journal of the ACM, 38(4):854– 886, October 1991. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/115362.html.

Chandru:1991:EHS

[CH91b] V. Chandru and J. N. Hooker. Extended Horn sets in propositional logic. Journal of the ACM, 38(1):205–221, January 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/102789.html.

Chatterjee:2014:EDA

[CH14] Krishnendu Chatterjee and Monika Henzinger. Efficient and dynamic algorithms for alternating Büchi games and maximal end-component decomposition. Journal of the ACM, 61(3):15:1– 15:??, May 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chase:1962:SPP

[Cha62] P. E. Chase. Stability properties of predictor-corrector methods for ordinary differential equations. Journal of the ACM, 9(4): 457–468, October 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chaitin:1966:LPC

[Cha66a] Gregory J. Chaitin. On the length of programs for computing finite binary sequences. Journal of the ACM, 13(4):547–569, October 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chartres:1966:ACP

[Cha66b] Bruce A. Chartres. Automatic controlled precision calculations. Journal of the ACM, 13(3):386–403, July 1966. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chaitin:1969:LPC

[Cha69a] Gregory J. Chaitin. On the length of programs for computing finite binary sequences: Statistical considerations. Journal of the ACM, 16(1):145–159, January 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chaitin:1969:SSP

[Cha69b] Gregory J. Chaitin. On the simplicity and speed of programs for computing infinite sets of natural numbers. Journal of the ACM, 16(3):407–422, July 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chang:1970:UPI

[Cha70] C. L. Chang. The unit proof and the input proof in theorem proving. Journal of the ACM, 17(4):698–707, October 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chang:1972:GMT

[Cha72] Shi-Kuo Chang. The generation of minimal trees with a Steiner topology. Journal of the ACM, 19(4):699–711, October 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chaitin:1974:ITL

[Cha74] Gregory J. Chaitin. Information-theoretic limitations of formal systems. Journal of the ACM, 21(3):403–424, July 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chaitin:1975:TPS

[Cha75] Gregory J. Chaitin. A theory of program size formally identical to information theory. Journal of the ACM, 22(3):329–340, July 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chazelle:1990:LBOa

[Cha90a] Bernard Chazelle. Lower bounds for orthogonal range searching: I. the reporting case. Journal of the ACM, 37(2):200–212, April 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/77614.html.

Chazelle:1990:LBOb

[Cha90b] Bernard Chazelle. Lower bounds for orthogonal range searching: II. the arithmetic model. Journal of the ACM, 37(3):439– 463, July 1990. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/79149.html.

Chazelle:2000:SHA

[Cha00a] Bernard Chazelle. The soft heap: an approximate priority queue with optimal error rate. Journal of the ACM, 47(6):1012–1027, 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/citations/
journals/jacm/2000-47-6/p1012-chazelle/.

Chazelve:2000:MST

[Cha00b] Bernard Chazelve. A minimum spanning tree algorithm with Inverse-Ackermann type complexity. Journal of the ACM, 47(6): 1028–1047, 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/pubs/

citations/journals/jacm/2000-47-6/p1028-chazelve/.

Chan:2001:DPC

[Cha01] Timothy M. Chan. Dynamic planar convex hull operations in near-logarithmic amortized time. Journal of the ACM, 48(1):1– 12, January 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/pubs/

articles/journals/jacm/2001-48-1/p1-chan/p1-chan.pdf;
 http://www.acm.org/pubs/citations/journals/jacm/2001-
48-1/p1-chan/.

Chan:2010:DDS

[Cha10] Timothy M. Chan. A dynamic data structure for 3-D convex hulls and 2-D nearest neighbor queries. Journal of the ACM, 57 (3):16:1–16:15, March 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chazelle:2014:CBF

[Cha14] Bernard Chazelle. The convergence of bird flocking. Journal of the ACM, 61(4):21:1–21:??, July 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chan:2016:ARP

[Cha16] Siu On Chan. Approximation resistance from pairwiseindependent subgroups. Journal of the ACM, 63(3):27:1–27:??, September 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chen:2017:TFG

[Che17] Hubie Chen. The tractability frontier of graph-like first-order query sets. Journal of the ACM, 64(4):26:1–26:??, September 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cheraghchi:2019:CUB

[Che19] Mahdi Cheraghchi. Capacity upper bounds for deletion-type channels. Journal of the ACM, 66(2):9:1–9:??, April 2019. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/ft_gateway.cfm?id=3281275.

Chin:1986:SPI

[Chi86] Francis Chin. Security problems on inference control for SUM, MAX, and MIN queries. Journal of the ACM, 33(3):451– 464, July 1986. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/5928.html.

Chatterjee:2015:MSS

[CHJS15] Krishnendu Chatterjee, Thomas A. Henzinger, Barbara Jobstmann, and Rohit Singh. Measuring and synthesizing systems in probabilistic environments. Journal of the ACM, 62(1):9:1–9:??, February 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cohen:1982:UBS

[CHK82] Jacques Cohen, Timothy Hickey, and Joel Katcoff. Upper bounds for speedup in parallel parsing. Journal of the ACM, 29(2):408–428, April 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Curien:1996:CPW

[CHL96] Pierre-Louis Curien, Thérèse Hardin, and Jean-Jacques Lévy. Confluence properties of weak and strong calculi of explicit substitutions. Journal of the ACM, 43(2):362–397, March 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
jacm/226675.html.

Chong:2001:CTO

[CHL01] Ka Wong Chong, Yijie Han, and Tak Wah Lam. Concurrent threads and optimal parallel minimum spanning trees algorithm. Journal of the ACM, 48(2):297–323, March 2001. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/citations/journals/jacm/
2001-48-2/p297-chong/.

Chow:1980:CTD

[Cho80] We-Min Chow. The cycle time distribution of exponential cyclic queues. Journal of the ACM, 27(2):281–286, April 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chen:2022:BNP

[CHO^{+}22] Lijie Chen, Shuichi Hirahara, Igor Carboni Oliveira, Ján Pich, Ninad Rajgopal, and Rahul Santhanam. Beyond natural proofs: Hardness magnification and locality. Journal of the ACM, 69 (4):25:1–25:??, August 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3538391.

Chandy:1977:PFL

[CHT77] K. Mani Chandy, John H. Howard, and Donald F. Towsley. Product form and local balance in queueing networks. Journal of the ACM, 24(2):250–263, April 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chandra:1996:WFD

[CHT96] Tushar Deepak Chandra, Vassos Hadzilacos, and Sam Toueg. The weakest failure detector for solving consensus. Journal of the ACM, 43(4):685–722, July 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/jacm/234549.html.

Chu:1961:SMS

[Chu61] J. T. Chu. Some methods for simplifying switching circuits using “don’t care” conditions. Journal of the ACM, 8(4):497–512, October 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chu:1965:ODF

[Chu65] J. T. Chu. Optimum decision functions for computer character recognition. Journal of the ACM, 12(2):213–226, April 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chu:1981:NME

[Chu81] C. K. Chu. A note on multiple error detection in ASCII numeric data communication. Journal of the ACM, 28(2):265–269, April 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Clarkson:2012:SOM

[CHW12] Kenneth L. Clarkson, Elad Hazan, and David P. Woodruff. Sublinear optimization for machine learning. Journal of the ACM, 59(5):23:1–23:49, October 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Crane:1974:SSSa

[CI74a] Michael A. Crane and Donald L. Iglehart. Simulating stable stochastic systems, I: General multiserver queues. Journal of the ACM, 21(1):103–113, January 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Crane:1974:SSSb

[CI74b] Michael A. Crane and Donald L. Iglehart. Simulating stable stochastic systems, II: Markov chains. Journal of the ACM, 21 (1):114–123, January 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chang:1988:POW

[CIV88] Jik H. Chang, Oscar H. Ibarra, and Anastasios Vergis. On the power of one-way communication. Journal of the ACM, 35(3):

697–726, July 1988. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/44493.html.

Cohen:2000:BTD

[CJJK00] David Cohen, Peter Jeavons, Peter Jonsson, and Manolis Koubarakis. Building tractable disjunctive constraints. Journal of the ACM, 47(5):826–853, 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:

//www.acm.org/pubs/citations/journals/jacm/2000-47-5/
p826-cohen/.

Csirik:2006:SSA

[CJK^{+}06] Janos Csirik, David S. Johnson, Claire Kenyon, James B. Orlin, Peter W. Shor, and Richard R. Weber. On the Sum-of-Squares algorithm for bin packing. Journal of the ACM, 53(1):1–65, January 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cheng:2022:DDE

[CJLW22] Kuan Cheng, Zhengzhong Jin, Xin Li, and Ke Wu. Deterministic document exchange protocols and almost optimal binary codes for edit errors. Journal of the ACM, 69(6):44:1–44:??, December 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3561046.

Crane:1965:PCA

[CK65] R. L. Crane and R. W. Klopfenstein. A predictor-corrector algorithm with an increased range of absolute stability. Journal of the ACM, 12(2):227–241, April 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Coffman:1968:FQM

[CK68] Edward G. Coffman and Leonard Kleinrock. Feedback queueing models for time-shared systems. Journal of the ACM, 15(4):549– 576, October 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chand:1970:ACP

[CK70] Donald R. Chand and Sham S. Kapur. An algorithm for convex polytopes. Journal of the ACM, 17(1):78–86, January 1970.

CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chang:1976:CCC

[CK76] Lena Chang and James F. Korsh. Canonical coin changing and greedy solutions. Journal of the ACM, 23(3):418–422, July 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cherniavsky:1979:CCH

[CK79] John C. Cherniavsky and Samuel N. Kamin. A complete and consistent Hoare axiomatics for a simple programming language. Journal of the ACM, 26(1):119–128, January 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Callahan:1995:DMP

[CK95a] Paul B. Callahan and S. Rao Kosaraju. A decomposition of multidimensional point sets with applications to k-nearest-neighbors and n-body potential fields. Journal of the ACM, 42(1):67– 90, January 1995. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/200853.html.

Chien:1995:PAR

[CK95b] Andrew A. Chien and Jae H. Kim. Planar-adaptive routing: Low-Cost adaptive networks for multiprocessors. Journal of the ACM, 42(1):91–123, January 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/200856.html.

Chuzhoy:2009:PFC

[CK09] Julia Chuzhoy and Sanjeev Khanna. Polynomial flow-cut gaps and hardness of directed cut problems. Journal of the ACM, 56 (2):6:1–6:28, April 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chierichetti:2015:LPF

[CK15] Flavio Chierichetti and Ravi Kumar. LSH-preserving functions and their applications. Journal of the ACM, 62(5):33:1–33:??, November 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chor:1998:PIR

[CKGS98] Benny Chor, Eyal Kushilevitz, Oded Goldreich, and Madhu Sudan. Private information retrieval. Journal of the ACM, 45(6): 965–981, November 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org:80/

pubs/citations/journals/jacm/1998-45-6/p965-chor/.

Chen:2007:ATB

[CKL^{+}07] Jiangzhuo Chen, Robert D. Kleinberg, László Lovász, Rajmohan Rajaraman, Ravi Sundaram, and Adrian Vetta. (almost) tight bounds and existence theorems for single-commodity confluent flows. Journal of the ACM, 54(4):16:1–16:32, July 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cheung:2013:FMR

[CKL13] Ho Yee Cheung, Tsz Chiu Kwok, and Lap Chi Lau. Fast matrix rank algorithms and applications. Journal of the ACM, 60(5): 31:1–31:25, October 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cormode:2023:RES

[CKL^{+}23] Graham Cormode, Zohar Karnin, Edo Liberty, Justin Thaler, and Pavel Veselý. Relative error streaming quantiles. Journal of the ACM, 70(5):30:1–30:??, October 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3617891.

Chen:2025:MFM

[CKL^{+}25] Li Chen, Rasmus Kyng, Yang Liu, Richard Peng, Maximilian Probst Gutenberg, and Sushant Sachdeva. Maximum flow and minimum-cost flow in almost-linear time. Journal of the ACM, 72(3):19:1–19:??, June 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cadek:2014:CAM

[CKM^{+}14] Martin Cadek, Marek Krcál, Jirí Matousek, Francis Sergeraert, Lukás Vokrínek, and Uli Wagner. Computing all maps into a sphere. Journal of the ACM, 61(3):17:1–17:??, May 2014. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cygan:2018:FHC

[CKN18] Marek Cygan, Stefan Kratsch, and Jesper Nederlof. Fast Hamiltonicity checking via bases of perfect matchings. Journal of the ACM, 65(3):12:1–12:??, March 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chandran:2014:PAA

[CKOR14] Nishanth Chandran, Bhavana Kanukurthi, Rafail Ostrovsky, and Leonid Reyzin. Privacy amplification with asymptotically optimal entropy loss. Journal of the ACM, 61(5):29:1–29:??, August 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cook:2017:VIE

[CKP17] Byron Cook, Heidy Khlaaf, and Nir Piterman. Verifying increasingly expressive temporal logics for infinite-state systems. Journal of the ACM, 64(2):15:1–15:??, June 2017. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chandra:1981:A

[CKS81] Ashok K. Chandra, Dexter C. Kozen, and Larry J. Stockmeyer. Alternation. Journal of the ACM, 28(1):114–133, January 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Christodoulou:2016:BCA

[CKS16] George Christodoulou, Annamária Kovács, and Michael Schapira. Bayesian combinatorial auctions. Journal of the ACM, 63(2): 11:1–11:??, May 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cosmadakis:1990:PTI

[CKV90] Stavros S. Cosmadakis, Paris C. Kanellakis, and Moshe Y. Vardi. Polynomial-time implication problems for unary inclusion dependencies. Journal of the ACM, 37(1):15–46, January 1990. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
78937.html.

Crane:1962:SGC

[CL62] Roger L. Crane and Robert J. Lambert. Stability of a generalized corrector formula. Journal of the ACM, 9(1):104–117, January 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cicalese:2011:CRE

[CL11] Ferdinando Cicalese and Eduardo Sany Laber. On the competitive ratio of evaluating priced functions. Journal of the ACM, 58(3):9:1–9:40, May 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chuzhoy:2016:PAA

[CL16] Julia Chuzhoy and Shi Li. A polylogarithmic approximation algorithm for edge-disjoint paths with congestion 2. Journal of the ACM, 63(5):45:1–45:??, December 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Clarke:1979:PLC

[Cla79] Edmund Melson Clarke, Jr. Programming language constructs for which it is impossible to obtain good Hoare axiom systems. Journal of the ACM, 26(1):129–147, January 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Clarkson:1995:VAL

[Cla95] Kenneth L. Clarkson. Las Vegas algorithms for linear and integer programming when the dimension is small. Journal of the ACM, 42(2):488–499, March 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/201036.html.

Chen:2008:FPA

[CLL^{+}08] Jianer Chen, Yang Liu, Songjian Lu, Barry O’Sullivan, and Igor Razgon. A fixed-parameter algorithm for the directed feedback vertex set problem. Journal of the ACM, 55(5):21:1–21:19, October 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Czerwinski:2021:RPP

[CLL^{+}21] Wojciech Czerwiński, Sławomir Lasota, Ranko Lazić, JÉrôme Leroux, and Filip Mazowiecki. The reachability problem for Petri nets is not elementary. Journal of the ACM, 68(1):7:1–7:28, February 2021. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL https://dl.acm.org/doi/10.
1145/3422822.

Candes:2011:RPC

[CLMW11] Emmanuel J. Candès, Xiaodong Li, Yi Ma, and John Wright. Robust principal component analysis? Journal of the ACM, 58 (3):11:1–11:37, May 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chor:1986:ANT

[CLRS86] Benny Chor, Charles E. Leiserson, Ronald L. Rivest, and James B. Shearer. An application of number theory to the organization of raster-graphics memory. Journal of the ACM, 33(1):86–104, January 1986. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/4800.html.

Chan:2016:ACS

[CLRS16] Siu On Chan, James R. Lee, Prasad Raghavendra, and David Steurer. Approximate constraint satisfaction requires large LP relaxations. Journal of the ACM, 63(4):34:1–34:??, November 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cohen:2021:SLP

[CLS21] Michael B. Cohen, Yin Tat Lee, and Zhao Song. Solving linear programs in the current matrix multiplication time. Journal of the ACM, 68(1):3:1–3:39, February 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3424305.

Chan:2018:SPH

[CLTZ18] T.-H. Hubert Chan, Anand Louis, Zhihao Gavin Tang, and Chenzi Zhang. Spectral properties of hypergraph Laplacian and approximation algorithms. Journal of the ACM, 65(3):15:1– 15:??, March 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Choudhury:1995:CNC

[CLW95] Gagan H. Choudhury, Kin K. Leung, and Ward Whitt. Calculating normalization constants of closed queuing networks by numerically inverting their generating functions. Journal of the ACM, 42(5):935–970, September 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/210122.html.

Chen:2025:CFP

[CLY25] Xi Chen, Yuhao Li, and Mihalis Yannakakis. Computing a fixed point of contraction maps in polynomial queries. Journal of the ACM, 72(4):25:1–25:??, August 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cordray:1957:RRP

[CM57] R. E. Cordray and Robert M. Mason. Remarks on a recent paper: “Arithmetical analysis of digital computing nets”. Journal of the ACM, 4(4):524–529, October 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Jef56].

Chow:1960:BCS

[CM60] Tse-Sun Chow and Harold Willis Milnes. Boundary contraction solution of Laplace’s differential equation II. Journal of the ACM, 7(1):37–45, January 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chow:1961:NSN

[CM61] Tse-Sun Chow and Harold Willis Milnes. Numerical solution of the Neumann and mixed boundary value problems by boundary contraction. Journal of the ACM, 8(3):336–358, July 1961. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cocke:1964:UTS

[CM64] John Cocke and Marvin Minsky. Universality of tag systems with P = 2. Journal of the ACM, 11(1):15–20, January 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Coveyou:1967:FAU

[CM67] R. R. Coveyou and R. D. MacPherson. Fourier analysis of uniform random number generators. Journal of the ACM, 14(1): 100–119, January 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chien:1974:DSM

[CM74] R. T. Chien and E. A. Mark. A document storage method based on polarized distance. Journal of the ACM, 21(2):233–245, April 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cooper:1978:FPU

[CM78] W. S. Cooper and M. E. Maron. Foundations of probabilistic and utility-theoretic indexing. Journal of the ACM, 25(1):67– 80, January 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chandy:1983:CPF

[CM83] K. M. Chandy and A. J. Martin. A characterization of productform queueing networks. Journal of the ACM, 30(2):286–299, April 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chazelle:1985:MCV

[CM85] Bernard Chazelle and Louis Monier. A model of computation for VLSI with related complexity results. Journal of the ACM, 32(3): 573–588, July 1985. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/3834.html.

Chan:1987:AQE

[CM87] Edward P. F. Chan and Alberto O. Mendelzon. Answering queries on embedded-complete database schemes. Journal of the ACM, 34(2):349–375, April 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/23007.html.

Cunto:1989:ACS

[CM89] Walter Cunto and J. Ian Munro. Average case selection. Journal of the ACM, 36(2):270–279, April 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/62047.html.

Cohen:1993:SPT

[CM93] Edith Cohen and Nimrod Megiddo. Strongly polynomial-time and NC algorithms for detecting cycles in periodic graphs. Journal of the ACM, 40(4):791–830, September 1993. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
153727.html.

Cheng:1996:BEI

[CM96] William C. Cheng and Richard R. Muntz. Bounding errors introduced by clustering of customers in closed product-form queuing networks. Journal of the ACM, 43(4):641–669, July 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
jacm/234539.html.

Crescenzi:2004:AIE

[CM04] Valter Crescenzi and Giansalvatore Mecca. Automatic information extraction from large websites. Journal of the ACM, 51(5): 731–779, September 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Czerwinski:2018:MTP

[CMNP18] Wojciech Czerwiński, Wim Martens, Matthias Niewerth, and Pawel Parys. Minimization of tree patterns. Journal of the ACM, 65(4):26:1–26:??, August 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chor:1989:SCT

[CMS89] Benny Chor, Michael Merritt, and David B. Shmoys. Simple constant-time consensus protocols in realistic failure models. Journal of the ACM, 36(3):591–614, July 1989. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/

65956.html. Review: Computing Reviews, June 1990.

Chattopadhyay:2020:LAR

[CMS20] Arkadev Chattopadhyay, Nikhil S. Mande, and Suhail Sherif. The log-approximate-rank conjecture is false. Journal of the ACM, 67(4):23:1–23:28, August 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3396695.

Coffman:1970:WTD

[CMT70] E. G. Coffman, Jr., R. R. Muntz, and H. Trotter. Waiting time distributions for processor-sharing systems. Journal of the ACM, 17(1):123–130, January 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ciaccia:2020:FCA

[CMT20] Paolo Ciaccia, Davide Martinenghi, and Riccardo Torlone. Foundations of context-aware preference propagation. Journal of the ACM, 67(1):4:1–4:43, February 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3375713.

Cormode:2012:CSD

[CMYZ12] Graham Cormode, S. Muthukrishnan, Ke Yi, and Qin Zhang. Continuous sampling from distributed streams. Journal of the ACM, 59(2):10:1–10:25, April 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chang:1995:BSE

[CN95] C. S. Chang and R. Nelson. Bounds on the speedup and efficiency of partial synchronization in parallel processing systems. Journal of the ACM, 42(1):204–231, January 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/200872.html.

Chuzhoy:2006:NHR

[CN06] Julia Chuzhoy and Joseph (Seffi) Naor. New hardness results for congestion minimization and machine scheduling. Journal of the ACM, 53(5):707–721, September 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cornuejols:1985:TSP

[CNP85] G. Cornuéjols, D. Naddef, and W. Pulleyblank. The traveling salesman problem in graphs with 3-edge cutsets. Journal of the ACM, 32(2):383–410, April 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/3154.html.

Cohen:2007:DEA

[CNS07] Sara Cohen, Werner Nutt, and Yehoshua Sagiv. Deciding equivalences among conjunctive aggregate queries. Journal of the ACM, 54(2):5:1–5:50, April 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Clenshaw:1984:BFP

[CO84] C. W. Clenshaw and F. W. J. Olver. Beyond floating point. Journal of the ACM, 31(2):319–328, April 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Coja-Oghlan:2017:BPG

[CO17] Amin Coja-Oghlan. Belief propagation guided decimation fails on random formulas. Journal of the ACM, 63(6):49:1–49:??, February 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Coffman:1968:ATT

[Cof68] E. G. Coffman, Jr. Analysis of two time-sharing algorithms designed for limited swapping. Journal of the ACM, 15(3):341–353, July 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Coffman:1969:ADI

[Cof69a] E. G. Coffman, Jr. Analysis of a drum input/output queue under scheduled operation in a paged computer system. Journal of the ACM, 16(1):73–90, January 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Coffman:1969:EAD

[Cof69b] E. G. Coffman, Jr. Errata: “Analysis of a drum input/output queue under scheduled operation in a paged computer system”. Journal of the ACM, 16(4):646, October 1969. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Coffman:2003:J

[Cof03] Ed Coffman. JACM 1976–1979. Journal of the ACM, 50(1): 16, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cohn:1955:SEI

[Coh55] Harvey Cohn. Some experiments in ideal factorization on the MIDAC. Journal of the ACM, 2(2):111–116, April 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cohn:1964:PLM

[Coh64] Martin Cohn. Properties of linear machines. Journal of the ACM, 11(3):296–301, July 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cohen:2000:PTN

[Coh00] Edith Cohen. Polylog-time and near-linear work approximation scheme for undirected shortest paths. Journal of the ACM, 47 (1):132–166, January 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/citations/journals/jacm/2000-47-1/p132-cohen/.

Cole:1958:NNS

[Col58] R. W. Cole. A note on numerical solution of certain linear boundary value problems. Journal of the ACM, 5(3):258–260, July 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Collins:1967:SRP

[Col67] George E. Collins. Subresultants and reduced polynomial remainder sequences. Journal of the ACM, 14(1):128–142, January 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Colmerauer:1970:TPR

[Col70] Alain Colmerauer. Total precedence relations. Journal of the ACM, 17(1):14–30, January 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cole:1971:DPS

[Col71a] Stephen N. Cole. Deterministic pushdown store machines and real-time computation. Journal of the ACM, 18(2):306–328, April 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Collins:1971:CMP

[Col71b] George E. Collins. The calculation of multivariate polynomial resultants. Journal of the ACM, 18(4):515–532, October 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cole:1987:SSN

[Col87] Richard Cole. Slowing down sorting networks to obtain faster sorting algorithms. Journal of the ACM, 34(1):200–208, January 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/7537.html.

Comba:1968:PDI

[Com68] Paul G. Comba. A procedure for detecting intersections of threedimensional objects. Journal of the ACM, 15(3):354–366, July 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Conte:1957:SIF

[Con57] S. D. Conte. A stable implicit finite difference approximation to a fourth order parabolic equation. Journal of the ACM, 4(1):18– 23, January 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Constable:1972:OG

[Con72] Robert L. Constable. The operator gap. Journal of the ACM, 19 (1):175–183, January 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Condon:1991:SBP

[Con91] Anne Condon. Space-bounded probabilistic game automata. Journal of the ACM, 38(2):472–494, April 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
128681.html.

Cooper:1964:FRD

[Coo64] William S. Cooper. Fact retrieval and deductive questionanswering information retrieval systems. Journal of the ACM, 11(2):117–137, April 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cook:1966:SDP

[Coo66] Stephen A. Cook. The solvability of the derivability problem for one-normal systems. Journal of the ACM, 13(2):223–225, April 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cook:1971:CPM

[Coo71] Stephen A. Cook. Characterizations of pushdown machines in terms of time-bounded computers. Journal of the ACM, 18(1): 4–18, January 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cook:2003:IPV

[Coo03] Stephen Cook. The importance of the P versus NP question. Journal of the ACM, 50(1):27–29, January 2003. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Corbato:1963:CJM

[Cor63] F. J. Corbató. On the coding of Jacobi’s method for computing eigenvalues and eigenvectors of symmetric matrices. Journal of the ACM, 10:123–125, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Corbato:2003:CSC

[Cor03] Fernando J. Corbató. On computer system challenges. Journal of the ACM, 50(1):30–31, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Coveyou:1960:SCG

[Cov60] R. R. Coveyou. Serial correlation in the generation of pseudorandom numbers. Journal of the ACM, 7(1):72–74, January 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chonev:2016:COP

[COW16] Ventsislav Chonev, Joël Ouaknine, and James Worrell. On the complexity of the orbit problem. Journal of the ACM, 63(3): 23:1–23:??, September 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chonev:2023:ZEP

[COW23] Ventsislav Chonev, Joel Ouaknine, and James Worrell. On the zeros of exponential polynomials. Journal of the ACM, 70(4): 26:1–26:??, August 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/

doi/10.1145/3603543.

Cosmadakis:1984:URV

[CP84] Stavros S. Cosmadakis and Christos H. Papadimitriou. Updates of relational views. Journal of the ACM, 31(4):742–760, October 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Crochemore:1991:TWS

[CP91] Maxime Crochemore and Dominique Perrin. Two-way string matching. Journal of the ACM, 38(3):651–675, July 1991. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
116845.html.

Chang:2021:NOD

[CPSZ21] Yi-Jun Chang, Seth Pettie, Thatchaphol Saranurak, and Hengjie Zhang. Near-optimal distributed triangle enumeration via expander decompositions. Journal of the ACM, 68(3):21:1–21:36, May 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3446330.

Conway:1994:EDM

[CPT94] Adrian E. Conway, Eugene Pinsky, and Srinivasan Tridandapani. Efficient decomposition methods for the analysis of multi-facility blocking models. Journal of the ACM, 41(4):648–675, July 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/179838.html.

Chen:2017:CNM

[CPY17] Xi Chen, Dimitris Paparas, and Mihalis Yannakakis. The complexity of non-monotone markets. Journal of the ACM, 64(3): 20:1–20:??, June 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Conte:1956:KTO

[CR56] S. D. Conte and R. F. Reeves. A Kutta third-order procedure for solving differential equations requiring minimal storage. Journal of the ACM, 3(1):22–25, January 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chaudhuri:1986:AGP

[CR86a] R. Chaudhuri and A. N. V. Rao. Approximating grammar probabilities: Solution of a conjecture. Journal of the ACM, 33(4):702– 705, October 1986. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/214099.html.

Cosnard:1986:CPQ

[CR86b] M. Cosnard and Y. Robert. Complexity of parallel QR factorization. Journal of the ACM, 33(4):712–723, October 1986. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
214102.html.

Compton:1995:EDT

[CR95] Kevin J. Compton and Chinya Ravishankar. Expected deadlock time in a multiprocessing system. Journal of the ACM, 42(3): 562–583, May 1995. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/210412.html.

Castaneda:2012:NCT

[CR12] Armando Castañeda and Sergio Rajsbaum. New combinatorial topology bounds for renaming: The upper bound. Journal of the ACM, 59(1):3:1–3:49, February 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Crandall:1954:NTF

[Cra54] Stephen H. Crandall. Numerical treatment of a fourth order parabolic partial differential equation. Journal of the ACM, 1 (3):111–117, July 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Crandall:1955:IVE

[Cra55] Stephen H. Crandall. Implicit vs. explicit recurrence formulas for the linear diffusion equation. Journal of the ACM, 2(1):42– 49, January 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Crandall:1957:ORF

[Cra57] Stephen H. Crandall. Optimum recurrence formulas for a fourth order parabolic partial differential equation. Journal of the ACM, 4(4):467–471, October 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Crespi-Reghizzi:1978:NCF

[CRGM78] S. Crespi-Reghizzi, G. Guida, and D. Mandrioli. Noncounting context-free languages. Journal of the ACM, 25(4):571–580, October 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Castaneda:2018:UCO

[CRR18] Armando Castañeda, Sergio Rajsbaum, and Michel Raynal. Unifying concurrent objects and distributed tasks: Intervallinearizability. Journal of the ACM, 65(6):45:1–45:??, November 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=
3266457.

Case:2001:LCA

[CRS01] John Case, Dayanand S. Rajan, and Anil M. Shende. Lattice computers for approximating Euclidean space. Journal of the ACM, 48(1):110–144, January 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/articles/journals/jacm/
2001-48-1/p110-case/p110-case.pdf;   http://www.acm.
org/pubs/citations/journals/jacm/2001-48-1/p110-case/

.

Chen:2023:ETH

[CRTY23] Lijie Chen, Ron D. Rothblum, Roei Tell, and Eylon Yogev. On exponential-time hypotheses, derandomization, and circuit lower bounds. Journal of the ACM, 70(4):25:1–25:??, August 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3593581.

Crump:1976:NIL

[Cru76] Kenny S. Crump. Numerical inversion of Laplace transforms using a Fourier series approximation. Journal of the ACM, 23 (1):89–96, January 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Charikar:2024:BMV

[CRWW24] Moses Charikar, Prasanna Ramakrishnan, Kangning Wang, and Hongxun Wu. Breaking the metric voting distortion barrier. Journal of the ACM, 71(6):42:1–42:??, December 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3689625.

Cryer:1967:NSQ

[Cry67] C. W. Cryer. On the numerical solution of a quasi-linear elliptic equation. Journal of the ACM, 14(2):363–375, April 1967. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cryer:1970:ASF

[Cry70] C. W. Cryer. On the approximate solution of free boundary problems using finite differences. Journal of the ACM, 17(3): 397–411, July 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cryer:1971:TPA

[Cry71] Colin W. Cryer. Topological problems arising when solving boundary value problems for elliptic partial differential equations by the method of finite differences. Journal of the ACM, 18(1):63–74, January 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cudia:1968:DUF

[CS68] Dennis F. Cudia and Wilson E. Singletary. Degrees of unsolvability in formal grammars. Journal of the ACM, 15(4):680–692, October 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chang:1971:CLR

[CS71] C. L. Chang and J. R. Slagle. Completeness of linear refutation for theories with equality. Journal of the ACM, 18(1):126–136, January 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Comer:1977:CTI

[CS77] Douglas Comer and Ravi Sethi. The complexity of trie index construction. Journal of the ACM, 24(3):428–440, July 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Coffman:1983:ISE

[CS83] E. G. Coffman, Jr. and Ravi Sethi. Instruction sets for evaluating arithmetic expressions. Journal of the ACM, 30(3):457–478, July 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Courtois:1984:BPE

[CS84] P.-J. Courtois and P. Semal. Bounds for the positive eigenvectors of nonnegative matrices and for their approximations by decomposition. Journal of the ACM, 31(4):804–825, October 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cao:1985:IAD

[CS85] Wei-Lu Cao and William J. Stewart. Iterative aggregation/ disaggregation techniques for nearly uncoupled Markov chains. Journal of the ACM, 32(3):702–719, July 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
214137.html.

Chvatal:1988:MHE

[CS88a] Vašek Chvátal and Endre Szemerédi. Many hard examples for resolution. Journal of the ACM, 35(4):759–768, October 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/48016.html.

Cole:1988:OVC

[CS88b] Richard Cole and Alan Siegel. Optimal VLSI circuits for sorting. Journal of the ACM, 35(4):777–809, October 1988. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
48017.html.

Cucker:1999:CED

[CS99] Felipe Cucker and Steve Smale. Complexity estimates depending on condition and round-off error. Journal of the ACM, 46(1): 113–184, January 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org:80/

pubs/citations/journals/jacm/1999-46-1/p113-cucker/.

Chazelle:2011:OGR

[CS11] Bernard Chazelle and C. Seshadhri. Online geometric reconstruction. Journal of the ACM, 58(4):14:1–14:32, July 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Conitzer:2007:WEF

[CSL07] Vincent Conitzer, Tuomas Sandholm, and Jérôme Lang. When are elections with few candidates hard to manipulate? Journal of the ACM, 54(3):14:1–14:33, June 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chudnovsky:2020:DOH

[CSSS20] Maria Chudnovsky, Alex Scott, Paul Seymour, and Sophie Spirkl. Detecting an odd hole. Journal of the ACM, 67(1):5:1– 5:12, February 2020. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL https://dl.acm.org/doi/abs/
10.1145/3375720.

Chen:2018:SQC

[CST^{+}18] Xi Chen, Rocco A. Servedio, Li-Yang Tan, Erik Waingarten, and Jinyu Xie. Settling the query complexity of non-adaptive junta testing. Journal of the ACM, 65(6):40:1–40:??, November 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cheung:2007:TSP

[CSV07] Ling Cheung, Mariëlle Stoelinga, and Frits Vaandrager. A testing scenario for probabilistic processes. Journal of the ACM, 54(6): 29:1–29:45, December 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chadha:2009:ECR

[CSV09] Rohit Chadha, A. Prasad Sistla, and Mahesh Viswanathan. On the expressiveness and complexity of randomization in finite state monitors. Journal of the ACM, 56(5):26:1–26:44, August 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chin:1981:MMF

[CT81] Francis Y. Chin and Long Lieh Tsai. On J-maximal and Jminimal flow-shop schedules. Journal of the ACM, 28(3):462– 476, July 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chandra:1996:UFD

[CT96] Tushar Deepak Chandra and Sam Toueg. Unreliable failure detectors for reliable distributed systems. Journal of the ACM, 43(2):225–267, March 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/jacm/226647.html.

Chor:2006:FML

[CT06] Benny Chor and Tamir Tuller. Finding a maximum likelihood tree is hard. Journal of the ACM, 53(5):722–744, September 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Corbato:1959:GSB

[CU59] Fernando J. Corbató and Jack L. Uretsky. Generation of spherical Bessel functions in digital computers. Journal of the ACM, 6(3):366–375, July 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cudia:1970:GPF

[Cud70] Dennis F. Cudia. General problems of formal grammars. Journal of the ACM, 17(1):31–43, January 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Culik:1979:PHC

[Cul79] K. Culik, II. A purely homomorphic characterization of recursively enumerable sets. Journal of the ACM, 26(2):345–350, April 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cunningham:1985:OAR

[Cun85] William H. Cunningham. Optimal attack and reinforcement of a network. Journal of the ACM, 32(3):549–561, July 1985. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
3829.html.

Curtis:1959:FCF

[Cur59a] H. Allen Curtis. A functional canonical form. Journal of the ACM, 6(2):245–258, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Curtis:1959:MCF

[Cur59b] H. Allen Curtis. Multifunctional circuits in functional canonical form. Journal of the ACM, 6(4):538–547, October 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Curtis:1961:GTC

[Cur61] H. Allen Curtis. A generalized tree circuit. Journal of the ACM, 8(4):484–496, October 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Curtis:1962:MRV

[Cur62] H. Allen Curtis. Multiple reduction of variable dependency of sequential machines. Journal of the ACM, 9(3):324–344, July 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Curtis:1963:GTC

[Cur63a] H. Allen Curtis. Generalized tree circuit—the basic building block of an extended decomposition theory. Journal of the ACM, 10:562–581, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Curtis:1963:UDT

[Cur63b] H. Allen Curtis. Use of decomposition theory in the solution of the state assignment problem of sequential machines. Journal of the ACM, 10:386–411, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Curtis:1965:TMS

[Cur65] M. W. Curtis. A Turing machine simulator. Journal of the ACM, 12(1):1–13, January 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cuthill:1959:MNB

[CV59] Elizabeth H. Cuthill and Richard S. Varga. A method of normalized block iteration. Journal of the ACM, 6(2):236–244, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chatterjee:2017:CMP

[CV17] Krishnendu Chatterjee and Yaron Velner. The complexity of mean-payoff pushdown games. Journal of the ACM, 64(5): 34:1–34:??, October 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Conn:1965:ODS

[CvH65] Richard W. Conn and Richard E. von Holdt. An online display for the study of approximating functions. Journal of the ACM, 12(3):326–349, July 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chang:1965:ART

[CW65] Wei Chang and Donald J. Wong. Analysis of real time multiprogramming. Journal of the ACM, 12(4):581–588, October 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Carroll:1967:APP

[CW67] A. B. Carroll and R. T. Wetherald. Applications of parallel processing to numerical weather prediction. Journal of the ACM, 14(3):591–614, July 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chen:1996:TED

[CW96] Weidong Chen and David S. Warren. Tabled evaluation with delaying for general logic programs. Journal of the ACM, 43(1):20– 74, January 1996. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/jacm/227597.html.

Clarkson:2017:LRA

[CW17] Kenneth L. Clarkson and David P. Woodruff. Low-rank approximation and regression in input sparsity time. Journal of the ACM, 63(6):54:1–54:??, February 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chow:1982:CFQ

[CY82] D. Chow and C. T. Yu. On the construction of feedback queries. Journal of the ACM, 29(1):127–151, January 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Courcoubetis:1995:CPV

[CY95] Costas Courcoubetis and Mihalis Yannakakis. The complexity of probabilistic verification. Journal of the ACM, 42(4):857– 907, July 1995. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/210339.html.

Chen:2014:CFT

[CYZG14] Binbin Chen, Haifeng Yu, Yuda Zhao, and Phillip B. Gibbons. The cost of fault tolerance in multi-party communication complexity. Journal of the ACM, 61(3):19:1–19:??, May 2014. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dubner:1968:NIL

[DA68] H. Dubner and J. Abate. Numerical inversion of Laplace transforms by relating them to the finite Fourier cosine transform. Journal of the ACM, 15(1):115–123, January 1968. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Daduna:1988:BPS

[Dad88] Hans Daduna. Busy periods for subnetworks in stochastic networks: Mean value analysis. Journal of the ACM, 35(3):668– 674, July 1988. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/44495.html.

Daley:1973:EIC

[Dal73] Robert P. Daley. An example of information and computation resource trade-off. Journal of the ACM, 20(4):687–695, October 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Darwiche:2001:DNN

[Dar01] Adnan Darwiche. Decomposable negation normal form. Journal of the ACM, 48(4):608–647, July 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Darwiche:2003:DAI

[Dar03] Adnan Darwiche. A differential approach to inference in Bayesian networks. Journal of the ACM, 50(3):280–305, May 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dass:1975:SBC

[Das75] Bal Kishan Dass. A sufficient bound for codes correcting bursts with weight constraint. Journal of the ACM, 22(4):501–503, October 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Das:1977:MRF

[Das77] Shawpawn Kumar Das. A machine representation of finite T_{0} topologies. Journal of the ACM, 24(4):676–692, October 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Davis:1977:PCT

[Dav77] Philip J. Davis. Proof, completeness, transcendentals, and sampling. Journal of the ACM, 24(2):298–310, April 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Davis:1990:CER

[Dav90] Henry W. Davis. Cost-error relationships in A* tree-searching. Journal of the ACM, 37(2):195–199, April 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
77595.html.

Davies:2017:TLA

[Dav17] Rowan Davies. A temporal logic approach to binding-time analysis. Journal of the ACM, 64(1):1:1–1:??, March 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dowdy:1992:SCB

[DCKT92] Lawrence W. Dowdy, Brian M. Carlson, Alan T. Krantz, and Satish K. Tripathi. Single-class bounds of multi-class queuing networks. Journal of the ACM, 39(1):188–213, January 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/147530.html.

deChampeaux:1977:IBH

[dCS77] Dennis de Champeaux and Lenie Sint. An improved bidirectional heuristic search algorithm. Journal of the ACM, 24(2):177–191, April 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

De:2014:NOS

[DDFS14] Anindya De, Ilias Diakonikolas, Vitaly Feldman, and Rocco A. Servedio. Nearly optimal solutions for the Chow parameters problem and low-weight approximation of halfspaces. Journal of the ACM, 61(2):11:1–11:??, April 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dolev:1987:MSN

[DDS87] Danny Dolev, Cynthia Dwork, and Larry Stockmeyer. On the minimal synchronism needed for distributed consensus. Journal of the ACM, 34(1):77–97, January 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/7533.html.

Doerr:2016:PMM

[DDST16] Benjamin Doerr, Carola Doerr, Reto Spöhel, and Henning Thomas. Playing Mastermind with many colors. Journal of the ACM, 63(5):42:1–42:??, December 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dolev:1993:PSM

[DDWY93] Danny Dolev, Cynthia Dwork, Orli Waarts, and Moti Yung. Perfectly secure message transmission. Journal of the ACM, 40(1):17–47, January 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/138036.html.

deChampeaux:1983:BHS

[de 83] Dennis de Champeaux. Bidirectional heuristic search again. Journal of the ACM, 30(1):22–32, January 1983. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

deChampeaux:1986:SFI

[de 86] Dennis de Champeaux. Subproblem finder and instance checker, two cooperating modules for theorem provers. Journal of the ACM, 33(4):633–657, October 1986. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/6491.html.

Dean:1989:UTH

[Dea89] Thomas Dean. Using temporal hierarchies to efficiently maintain large temporal databases. Journal of the ACM, 36(4):687– 718, October 1989. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/76360.html.

Debray:1992:EDA

[Deb92] Saumya K. Debray. Efficient dataflow analysis of logic programs. Journal of the ACM, 39(4):949–984, October 1992. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
146624.html.

Dujmovic:2021:ALP

[DEG^{+}21] Vida Dujmović, Louis Esperet, Cyril Gavoille, Gwenaël Joret, Piotr Micek, and Pat Morin. Adjacency labelling for planar graphs (and beyond). Journal of the ACM, 68(6):42:1–42:33, December 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3477542.

Delbrouck:1970:FQS

[Del70] L. E. N. Delbrouck. A feedback queueing system with batch arrivals, bulk service, and queue-dependent service time. Journal of the ACM, 17(2):314–323, April 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

DeMillo:1980:STT

[DEL80] Richard A. DeMillo, Stanley C. Eisenstat, and Richard J. Lipton. Space-time trade-offs in structured programming: An improved combinatorial embedding theorem. Journal of the ACM, 27(1): 123–127, January 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Demolombe:1992:SCS

[Dem92] Robert Demolombe. Syntactical characterization of a subset of domain-independent formulas. Journal of the ACM, 39(1):71– 94, January 1992. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/147520.html.

Dennis:1958:HSC

[Den58] Jack B. Dennis. A high-speed computer technique for the transportation problem. Journal of the ACM, 5(2):132–153, April 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Denman:1961:CGO

[Den61] Harry H. Denman. Computer generation of optimized subroutines. Journal of the ACM, 8(1):104–116, January 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [GG61].

Dennis:1965:SDM

[Den65] Jack B. Dennis. Segmentation and the design of multiprogrammed computer systems. Journal of the ACM, 12(4):589– 602, October 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Deo:1976:NHT

[Deo76] Narsingh Deo. Note on Hopcroft and Tarjan’s planarity algorithm. Journal of the ACM, 23(1):74–75, January 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [HT74].

Dolev:2010:RBC

[DEP10] Shlomi Dolev, Yuval Elovici, and Rami Puzis. Routing betweenness centrality. Journal of the ACM, 57(4):25:1–25:27, April 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Devroye:1986:NHB

[Dev86] Luc Devroye. A note on the height of binary search trees. Journal of the ACM, 33(3):489–498, July 1986. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/5930.html.

Dinitz:2024:SAI

[DFGN24] Michael Dinitz, Jeremy Fineman, Seth Gilbert, and Calvin Newport. Smoothed analysis of information spreading in dynamic networks. Journal of the ACM, 71(3):17:1–17:??, June 2024. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3661831.

Deligkas:2024:PCT

[DFHM24] Argyrios Deligkas, John Fearnley, Alexandros Hollender, and Themistoklis Melissourgos. Pure-circuit: Tight inapproximability for PPAD. Journal of the ACM, 71(5):31:1–31:??, October 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3678166.

Demaine:2005:SPA

[DFHT05] Erik D. Demaine, Fedor V. Fomin, Mohammadtaghi Hajiaghayi, and Dimitrios M. Thilikos. Subexponential parameterized algorithms on bounded-genus graphs and H-minor-free graphs. Journal of the ACM, 52(6):866–893, November 2005. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dyer:1991:RPT

[DFK91] Martin Dyer, Alan Frieze, and Ravi Kannan. A random polynomial time algorithm for approximating the volume of convex bodies. Journal of the ACM, 38(1):1–17, January 1991. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/

102783.html. A constant time oracle is assumed for determining if a point in space is inside or outside a convex body in n-dimensional Euclidean space. The algorithm runs in time bounded by a polynomial in n, the dimension of the body, and 1/ϵ, where ϵ is the relative error bound. With probability 3/4, it finds an approximation satisfying the error bound.

Dutting:2024:OAT

[DFN^{+}24] Paul Dütting, Zhe Feng, Harikrishna Narasimhan, David C. Parkes, and Sai Srivatsa Ravindranath. Optimal auctions through deep learning: Advances in differentiable economics. Journal of the ACM, 71(1):5:1–5:??, February 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3630749.

Dubois:2002:QDT

[DFPP02] Didier Dubois, Hélène Fargier, Henri Prade, and Patrice Perny. Qualitative decision theory: from Savage’s axioms to nonmonotonic reasoning. Journal of the ACM, 49(4):455–495, July 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dolev:2014:FTA

[DFSL14] Danny Dolev, Matthias Függer, Ulrich Schmid, and Christoph Lenzen. Fault-tolerant algorithms for tick-generation in asynchronous logic: Robust pulse generation. Journal of the ACM, 61(5):30:1–30:??, August 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dwyer:1957:MRM

[DG57] Paul S. Dwyer and Bernard A. Galler. The method of reduced matrices for a general transportation problem. Journal of the ACM, 4(3):308–313, July 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Douglas:1962:ADM

[DG62] Jim Douglas, Jr. and James E. Gunn. Alternating direction methods for parabolic systems in m space variables. Journal of the ACM, 9(4):450–456, October 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dill:1971:GSS

[DG71] C. Dill and C. W. Gear. A graphical search for stiffly stable methods for ordinary differential equations. Journal of the ACM, 18(1):75–79, January 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Datta:1984:SCD

[DG84] Ajoy Datta and S. Ghosh. Synthesis of a class of deadlock-free Petri nets. Journal of the ACM, 31(3):486–506, July 1984. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

deSouzaeSilva:1989:CAP

[dG89] Edmundo de Souza e Silva and H. Richard Gail. Calculating availability and performability measures of repairable computer systems using randomization. Journal of the ACM, 36(1):171– 193, January 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dey:1998:CHG

[DG98] Tamal K. Dey and Sumanta Guha. Computing homology groups of simplicial complexes in R^{3}. Journal of the ACM, 45(2):266– 287, March 1998. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org:80/pubs/
citations/journals/jacm/1998-45-2/p266-dey/.

Dvir:2016:SPS

[DG16] Zeev Dvir and Sivakanth Gopi. 2-server PIR with subpolynomial communication. Journal of the ACM, 63(4):39:1–39:??, November 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dottling:2021:IBE

[DG21] Nico Döttling and Sanjam Garg. Identity-based encryption from the Diffie–Hellman assumption. Journal of the ACM, 68(3): 14:1–14:46, May 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/

doi/10.1145/3422370.

Dong:2025:NDM

[DGG^{+}25] Sally Dong, Yu Gao, Gramoz Goranci, Yin Tat Lee, Sushant Sachdeva, Richard Peng, and Guanghao Ye. Nested dissection meets IPMs: Planar min-cost flow in nearly-linear time. Journal of the ACM, 72(4):27:1–27:??, August 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dubiner:1994:FTP

[DGM94] Moshe Dubiner, Zvi Galil, and Edith Magen. Faster tree pattern matching. Journal of the ACM, 41(2):205–213, March 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/174653.html.

Dyer:2007:CHD

[DGP07] Martin Dyer, Leslie Ann Goldberg, and Mike Paterson. On counting homomorphisms to directed acyclic graphs. Journal of the ACM, 54(6):27:1–27:23, December 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dooly:2001:LAT

[DGS01] Daniel R. Dooly, Sally A. Goldman, and Stephen D. Scott. On-line analysis of the TCP acknowledgment delay problem. Journal of the ACM, 48(2):243–273, March 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/citations/journals/jacm/
2001-48-2/p243-dooly/.

Downing:1956:SIC

[DH56] A. C. Downing, Jr. and A. S. Householder. Some inverse characteristic value problems. Journal of the ACM, 3(3):203–207, July 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

DiDonato:1959:NFC

[DH59] A. R. DiDonato and A. V. Hershey. New formulas for computing incomplete elliptic integrals of the first and second kind. Journal of the ACM, 6(4):515–526, October 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Digricoli:1986:EBB

[DH86] Vincent J. Digricoli and Malcolm C. Harrison. Equality-based binary resolution. Journal of the ACM, 33(2):253–289, April 1986. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/5389.html.

Drusinsky:1994:PBC

[DH94] Doron Drusinsky and David Harel. On the power of bounded concurrency I: Finite automata. Journal of the ACM, 41(3): 517–539, May 1994. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/176587.html.

Danna:2017:UMM

[DHK^{+}17] Emilie Danna, Avinatan Hassidim, Haim Kaplan, Alok Kumar, Yishay Mansour, Danny Raz, and Michal Segalov. Upward maxmin fairness. Journal of the ACM, 64(1):2:1–2:??, March 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Disser:2019:TBU

[DHK19] Yann Disser, Jan Hackfeld, and Max Klimm. Tight bounds for undirected graph exploration with pebbles and multiple agents. Journal of the ACM, 66(6):40:1–40:??, October 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/ft_gateway.cfm?id=3356883.

Dughmi:2021:BFB

[DHKN21] Shaddin Dughmi, Jason Hartline, Robert D. Kleinberg, and Rad Niazadeh. Bernoulli factories and black-box reductions in mechanism design. Journal of the ACM, 68(2):10:1–10:30, March 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3440988.

Dudik:2020:OEO

[DHL^{+}20] Miroslav Dudík, Nika Haghtalab, Haipeng Luo, Robert E. Schapire, Vasilis Syrgkanis, and Jennifer Wortman Vaughan. Oracle-efficient online learning and auction design. Journal of the ACM, 67(5):26:1–26:57, October 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3402203.

Dantzig:1965:UPF

[DHM65] G. B. Dantzig, R. P. Harvey, and R. D. McKnight. Updating the product form of the inverse for the revised simplex method (A summary). Journal of the ACM, 12(4):603, October 1965. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

The complete paper is published in Proceedings of the 20th National Conference of the Association for Computing Machinery, Cleveland, August 1965.

Dolev:1995:DFT

[DHSS95] Danny Dolev, Joseph Y. Halpern, Barbara Simons, and Ray Strong. Dynamic fault-tolerant clock synchronization. Journal of the ACM, 42(1):143–185, January 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/200870.html.

Dwork:1997:CSM

[DHW97] Cynthia Dwork, Maurice Herlihy, and Orli Waarts. Contention in shared memory algorithms. Journal of the ACM, 44(6):779–805, November 1997. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org:80/pubs/
citations/journals/jacm/1997-44-6/p779-dwork/.

DiPaola:1969:RSS

[Di 69a] Robert A. Di Paola. Random sets in subrecursive hierarchies. Journal of the ACM, 16(4):621–630, October 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

DiPaola:1969:RUD

[Di 69b] Robert A. Di Paola. The recursive unsolvability of the decision problem for the class of definite formulas. Journal of the ACM, 16(2):324–327, April 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

DiPaola:1973:SDP

[Di 73] Robert A. Di Paola. The solvability of the decision problem for classes of proper formulas and related results. Journal of the ACM, 20(1):112–126, January 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Donatiello:1987:ACP

[DI87] Lorenzo Donatiello and Balakrishna R. Iyer. Analysis of a composite performance reliability measure for fault-tolerant systems. Journal of the ACM, 34(1):179–199, January 1987. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
7536.html.

Demetrescu:2004:NAD

[DI04] Camil Demetrescu and Giuseppe F. Italiano. A new approach to dynamic all pairs shortest paths. Journal of the ACM, 51(6): 968–992, November 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Demetrescu:2005:TOF

[DI05] Camil Demetrescu and Giuseppe F. Italiano. Trade-offs for fully dynamic transitive closure on DAGs: breaking through the O(n^{2}) barrier. Journal of the ACM, 52(2):147–156, March 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dill:1989:CES

[Dil89] Jens M. Dill. A counter-example for “A simpler construction for showing the intrinsically exponential complexity of the circularity problem for attribute grammars”. Journal of the ACM, 36 (1):92–96, January 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/77393.html. See [Jaz81].

Dinur:2007:PTG

[Din07] Irit Dinur. The PCP theorem by gap amplification. Journal of the ACM, 54(3):12:1–12:44, June 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dixon:1973:RTP

[Dix73] John K. Dixon. Z-resolution: Theorem-Proving with compiled axioms. Journal of the ACM, 20(1):127–147, January 1973. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Davis:1981:AST

[DJ81] Ernest Davis and Jeffrey M. Jaffe. Algorithms for scheduling tasks on unrelated processors. Journal of the ACM, 28(4):721– 736, October 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Deineko:2008:AMC

[DJKK08] Vladimir Deineko, Peter Jonsson, Mikael Klasson, and Andrei Krokhin. The approximability of MAX CSP with fixed-value constraints. Journal of the ACM, 55(4):16:1–16:37, September 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dyer:2017:SMC

[DJM17] Martin Dyer, Mark Jerrum, and Haiko Müller. On the switch Markov chain for perfect matchings. Journal of the ACM, 64 (2):12:1–12:??, June 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dujmovic:2020:PGB

[DJM^{+}20] Vida Dujmović, Gwenaël Joret, Piotr Micek, Pat Morin, Torsten Ueckerdt, and David R. Wood. Planar graphs have bounded queue-number. Journal of the ACM, 67(4):22:1–22:38, August 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3385731.

Devanur:2019:NOO

[DJSW19] Nikhil R. Devanur, Kamal Jain, Balasubramanian Sivan, and Christopher A. Wilkens. Near optimal online algorithms and fast approximation algorithms for resource allocation problems. Journal of the ACM, 66(1):7:1–7:??, January 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Duchi:2014:PAL

[DJW14] John C. Duchi, Michael I. Jordan, and Martin J. Wainwright. Privacy aware learning. Journal of the ACM, 61(6):38:1–38:??, November 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

DiGri:1959:SSI

[DK59] Vincent J. DiGri and Jane E. King. The share 709 system: Input-Output translation. Journal of the ACM, 6(2):141–144, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dixon:1965:COS

[DK65] W. J. Dixon and R. A. Kronmal. The choice of origin and scale for graphs. Journal of the ACM, 12(2):259–261, April 1965. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dinur:2024:FGC

[DKK24] Itai Dinur, Nathan Keller, and Ohad Klein. Fine-grained cryptanalysis: Tight conditional bounds for dense k-SUM and k-XOR. Journal of the ACM, 71(3):23:1–23:??, June 2024. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3653014.

Datta:2018:RD

[DKM^{+}18] Samir Datta, Raghav Kulkarni, Anish Mukherjee, Thomas Schwentick, and Thomas Zeume. Reachability is in DynFO. Journal of the ACM, 65(5):33:1–33:??, September 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/ft_gateway.cfm?id=3212685.

Die:1998:HLU

[DKP98] Xiaotie Die, Tiko Kameda, and Christos Papadimitriou. How to learn an unknown environment. I: the rectilinear case. Journal of the ACM, 45(2):215–245, March 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:/

/www.acm.org:80/pubs/citations/journals/jacm/1998-45-
2/p215-die/.

Danos:2007:MC

[DKP07] Vincent Danos, Elham Kashefi, and Prakash Panangaden. The measurement calculus. Journal of the ACM, 54(2):8:1–8:45, April 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Diekert:2015:RLC

[DKRW15] Volker Diekert, Manfred Kufleitner, Klaus Reinhardt, and Tobias Walter. Regular languages are Church–Rosser congruential. Journal of the ACM, 62(5):39:1–39:??, November 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dvorak:2013:TFO

[DKT13] Zdenek Dvorák, Daniel Král, and Robin Thomas. Testing firstorder properties for subclasses of sparse graphs. Journal of the ACM, 60(5):36:1–36:24, October 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Derr:1956:SAA

[DL56] J. I. Derr and R. C. Luke. Semi-automatic allocation of data storage for PACT I. Journal of the ACM, 3(4):299–308, October 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

deSouzaeSilva:1989:CJQ

[dL89] Edmundo de Souza e Silva and S. S. Lavenberg. Calculating joint queue-length distributions in product-form queuing networks. Journal of the ACM, 36(1):194–207, January 1989. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

DeVerdiere:2007:OPD

[DL07] Éric Colin De Verdière and Francis Lazarus. Optimal pants decompositions and shortest homotopic cycles on an orientable surface. Journal of the ACM, 54(4):18:1–18:27, July 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Despre:2019:CGI

[DL19] Vincent Despré and Francis Lazarus. Computing the geometric intersection number of curves. Journal of the ACM, 66(6): 45:1–45:??, December 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3363367.

Dalphin:1973:BLS

[DLN73] John F. Dalphin and Victor Lovass-Nagy. Best least squares solutions to finite difference equations using the generalized inverse and tensor product methods. Journal of the ACM, 20(2):279– 289, April 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dolev:1986:RAA

[DLP^{+}86] Danny Dolev, Nancy A. Lynch, Shlomit S. Pinter, Eugene W. Stark, and William E. Weihl. Reaching approximate agreement in the presence of faults. Journal of the ACM, 33(3):499– 516, July 1986. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/5931.html.

Dwork:1988:CPP

[DLS88] Cynthia Dwork, Nancy Lynch, and Larry Stockmeyer. Consensus in the presence of partial synchrony. Journal of the ACM, 35 (2):288–323, April 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/42283.html.

Dallery:1994:ERS

[DLT94] Yves Dallery, Zhen Liu, and Don Townsley. Equivalence, reversibility, symmetry and concavity properties in fork-join queueing networks with blocking. Journal of the ACM, 41(5):903–942, September 1994. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/185776.html.

Duffield:2007:PSE

[DLT07] Nick Duffield, Carsten Lund, and Mikkel Thorup. Priority sampling for estimation of arbitrary subset sums. Journal of the ACM, 54(6):32:1–32:37, December 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dalirrooyfard:2025:HAD

[DLW25] Mina Dalirrooyfard, Ray Li, and Virginia Vassilevska Williams. Hardness of approximate diameter: Now for undirected graphs. Journal of the ACM, 72(1):6:1–6:??, February 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3704631.

Dulmage:1962:MAH

[DM62] A. L. Dulmage and N. S. Mendelsohn. Matrices associated with the Hitchcock problem. Journal of the ACM, 9(4):409–418, October 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dobkin:1981:OTM

[DM81] David Dobkin and J. Ian Munro. Optimal time minimal space selection algorithms. Journal of the ACM, 28(3):454–461, July 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Degano:1987:MDS

[DM87] Pierpaolo Degano and Ugo Montanari. A model for distributed systems based on graph rewriting. Journal of the ACM, 34 (2):411–449, April 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/24038.html.

Dhoked:2025:AFT

[DM25] Sahil Dhoked and Neeraj Mittal. Adaptive and fair transformation for recoverable mutual exclusion. Journal of the ACM, 72 (4):24:1–24:??, August 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Doron:2022:NOP

[DMOZ22] Dean Doron, Dana Moshkovitz, Justin Oh, and David Zuckerman. Nearly optimal pseudorandomness from hardness. Journal of the ACM, 69(6):43:1–43:??, December 2022. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3555307.

Dinitz:2008:BCB

[DMR08] Yefim Dinitz, Shlomo Moran, and Sergio Rajsbaum. Bit complexity of breaking and achieving symmetry in chains and rings. Journal of the ACM, 55(1):3:1–3:28, February 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

deMesmay:2020:ERN

[dMRST20] Arnaud de Mesmay, Yo’av Rieck, Eric Sedgwick, and Martin Tancer. Embeddability in R^{3} is NP-hard. Journal of the ACM, 67(4):20:1–20:29, August 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3396593.

Dwork:2003:MFM

[DNRS03] Cynthia Dwork, Moni Naor, Omer Reingold, and Larry Stockmeyer. Magic functions: In memoriam: Bernard M. Dwork 1923– 1998. Journal of the ACM, 50(6):852–921, November 2003. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dwork:2004:CZK

[DNS04] Cynthia Dwork, Moni Naor, and Amit Sahai. Concurrent zeroknowledge. Journal of the ACM, 51(6):851–898, November 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Detlefs:2005:STP

[DNS05] David Detlefs, Greg Nelson, and James B. Saxe. Simplify: a theorem prover for program checking. Journal of the ACM, 52 (3):365–473, May 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Douglas:1959:REN

[Dou59] Jim Douglas, Jr. Round-off error in the numerical solution of the heat equation. Journal of the ACM, 6(1):48–58, January 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Doyle:1961:SRM

[Doy61] Lauren B. Doyle. Semantic road maps for literature searchers. Journal of the ACM, 8(4):553–578, October 1961. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Doyle:1962:OUS

[Doy62] W. Doyle. Operations useful for similarity-invariant pattern recognition. Journal of the ACM, 9(2):259–267, April 1962. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Doyle:1965:ACR

[Doy65] Lauren B. Doyle. Is automatic classification a reasonable application of statistical analysis of text? Journal of the ACM, 12 (4):473–489, October 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Doyen:2023:SGS

[Doy23] Laurent Doyen. Stochastic games with synchronization objectives. Journal of the ACM, 70(3):23:1–23:??, June 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3588866.

Davis:1960:CPQ

[DP60] Martin Davis and Hilary Putman. A computing procedure for quantification theory. Journal of the ACM, 7(3):201–215, July 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dressler:1975:PPS

[DP75] Robert E. Dressler and S. Thomas Parker. Primes with a prime subscript. Journal of the ACM, 22(3):380–381, July 1975. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dechter:1985:GBF

[DP85] Rina Dechter and Judea Pearl. Generalized best-first search strategies and the optimality of A*. Journal of the ACM, 32(3): 505–536, July 1985. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/3830.html.

Davies:2001:MAS

[DP01] Rowan Davies and Frank Pfenning. A modal analysis of staged computation. Journal of the ACM, 48(3):555–604, May 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/citations/
journals/jacm/2001-48-3/p555-davies/.

Duan:2014:LTA

[DP14] Ran Duan and Seth Pettie. Linear-time approximation for maximum weight matching. Journal of the ACM, 61(1):1:1–1:??, January 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dory:2022:EFS

[DP22] Michal Dory and Merav Parter. Exponentially faster shortest paths in the congested clique. Journal of the ACM, 69 (4):29:1–29:??, August 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3527213.

Dowd:1989:PBS

[DPRS89] Martin Dowd, Yehoshua Perl, Larry Rudolph, and Michael Saks. The periodic balanced sorting network. Journal of the ACM, 36(4):738–757, October 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/76362.html.

Devanur:2008:MEP

[DPSV08] Nikhil R. Devanur, Christos H. Papadimitriou, Amin Saberi, and Vijay V. Vazirani. Market equilibrium via a primal–dual algorithm for a convex program. Journal of the ACM, 55(5): 22:1–22:18, October 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Delgrande:2018:GBR

[DPW18a] James P. Delgrande, Pavlos Peppas, and Stefan Woltran. General belief revision. Journal of the ACM, 65(5):29:1– 29:??, September 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3203409.

Dziembowski:2018:NMC

[DPW18b] Stefan Dziembowski, Krzysztof Pietrzak, and Daniel Wichs. Non-malleable codes. Journal of the ACM, 65(4):20:1–20:??, August 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Davis:1954:MPO

[DR54] Philip Davis and Philip Rabinowitz. A multiple purpose orthonormalizing code and its uses. Journal of the ACM, 1(4): 183–191, October 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dolev:1985:BIE

[DR85] Danny Dolev and Rüdiger Reischuk. Bounds on information exchange for Byzantine agreement. Journal of the ACM, 32(1):191– 204, January 1985. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/214112.html.

Dymond:2000:PRO

[DR00] Patrick W. Dymond and Walter L. Ruzzo. Parallel RAMs with owned global memory and deterministic context-free language recognition. Journal of the ACM, 47(1):16–45, January 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/citations/
journals/jacm/2000-47-1/p16-dymond/.

Dechter:2003:MBG

[DR03] Rina Dechter and Irina Rish. Mini-buckets: a general scheme for bounded inference. Journal of the ACM, 50(2):107–153, March 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Drmota:2003:AAH

[Drm03] Michael Drmota. An analytic approach to the height of binary search trees II. Journal of the ACM, 50(3):333–374, May 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dolev:1990:ESB

[DRS90] Danny Dolev, Ruediger Reischuk, and H. Raymond Strong. Early stopping in Byzantine agreement. Journal of the ACM, 37(4):720–741, October 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/96565.html.

Dughmi:2016:OMC

[DRY16] Shaddin Dughmi, Tim Roughgarden, and Qiqi Yan. Optimal mechanisms for combinatorial auctions and combinatorial public projects via convex rounding. Journal of the ACM, 63(4): 30:1–30:??, November 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

denBroeder:1958:PSD

[dS58] George G. den Broeder, Jr. and Harry J. Smith. A property of semi-definite Hermitian matrices. Journal of the ACM, 5(3): 244–245, July 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Downey:1978:ACA

[DS78] Peter J. Downey and Ravi Sethi. Assignment commands with array references. Journal of the ACM, 25(4):652–666, October 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dobkin:1991:MGE

[DS91] David Dobkin and Subhash Suri. Maintenance of geometric extrema. Journal of the ACM, 38(2):275–298, April 1991. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
103518.html.

Dwork:1992:FSVa

[DS92a] Cynthia Dwork and Larry Stockmeyer. Finite state verifiers I: The power of interaction. Journal of the ACM, 39(4):800– 828, October 1992. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/146599.html.

Dwork:1992:FSVb

[DS92b] Cynthia Dwork and Larry Stockmeyer. Finite state verifiers II: Zero knowledge. Journal of the ACM, 39(4):829–858, October 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/146601.html.

Dalvi:2012:DPI

[DS12] Nilesh Dalvi and Dan Suciu. The dichotomy of probabilistic inference for unions of conjunctive queries. Journal of the ACM, 59 (6):30:1–30:87, December 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Special issue text supplied by Joanne Pello.

Drmota:2013:MTD

[DS13] Michael Drmota and Wojciech Szpankowski. A master theorem for discrete divide and conquer recurrences. Journal of the ACM, 60(3):16:1–16:49, June 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dutta:2022:DRU

[DSS22] Pranjal Dutta, Nitin Saxena, and Amit Sinhababu. Discovering the roots: Uniform closure results for algebraic classes under factoring. Journal of the ACM, 69(3):18:1–18:39, June 2022. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3510359.

Downey:1980:VCS

[DST80] Peter J. Downey, Ravi Sethi, and Robert Endre Tarjan. Variations on the common subexpression problem. Journal of the ACM, 27(4):758–771, October 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dillencourt:1992:CGA

[DST92a] Michael B. Dillencourt, Hanan Samet, and Markku Tamminen. Corrigenda: “A general approach to connected-component labelling for arbitrary image representations”. Journal of the ACM, 39(4):985–986, October 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [DST92b].

Dillencourt:1992:GAC

[DST92b] Michael B. Dillencourt, Hanan Samet, and Markku Tamminen. A general approach to connected-component labelling for arbitrary image representations. Journal of the ACM, 39(2):253– 280, April 1992. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/128750.html. See also [DST92a].

Driscoll:1994:FPL

[DST94] James R. Driscoll, Daniel D. K. Sleator, and Robert E. Tarjan. Fully persistent lists with catenation. Journal of the ACM, 41 (5):943–959, September 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/185791.html.

Dubey:1966:SDC

[Dub66] Satya D. Dubey. Statistical determination of certain mathematical constants and functions using computers. Journal of the ACM, 13(4):511–525, October 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dunham:1965:CPM

[Dun65] Charles B. Dunham. Convergence problems in Maehly’s second method. Journal of the ACM, 12(2):181–186, April 1965. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dunham:1966:CPM

[Dun66] Charles B. Dunham. Convergence problems in Maehly’s second method: Part II. Journal of the ACM, 13(1):108–113, January 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dunham:1974:ECA

[Dun74] Charles B. Dunham. Efficiency of Chebyshev approximation on finite subsets. Journal of the ACM, 21(2):311–313, April 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

deVogelaere:1959:RPT

[dV59] René de Vogelaere. Remarks on the paper “Tchebysheff approximations for power series”. Journal of the ACM, 6(1):111–114, January 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Min57].

deVries:1971:MSD

[dV71] Ronald C. de Vries. Minimal sets of distinct literals for a logically passive function. Journal of the ACM, 18(3):431–443, July 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

DeNicola:1995:TLB

[DV95] Rocco De Nicola and Frits Vaandrager. Three logics for branching bisimulation. Journal of the ACM, 42(2):458–487, March 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/201032.html.

Dell:2014:SAN

[DV14] Holger Dell and Dieter Van Melkebeek. Satisfiability allows no nontrivial sparsification unless the polynomial-time hierarchy collapses. Journal of the ACM, 61(4):23:1–23:??, July 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dobzinski:2016:IRT

[DV16] Shahar Dobzinski and Jan Vondrák. Impossibility results for truthful combinatorial auctions with submodular valuations. Journal of the ACM, 63(1):5:1–5:??, March 2016. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

DeMillo:1977:SSM

[DVSC77] R. A. DeMillo, K. Vairavan, and E. Sycara-Cyranski. A study of schedules as models of synchronous parallel computation. Journal of the ACM, 24(4):544–565, October 1977. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dressler:1964:ERT

[DW64] R. F. Dressler and W. Werner. Error rates for two methods of statistical pattern recognition. Journal of the ACM, 11(4):471– 480, October 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dwork:1999:SEB

[DW99] Cynthia Dwork and Orli Waarts. Simple and efficient bounded concurrent timestamping and the traceable use abstraction. Journal of the ACM, 46(5):633–666, September 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dolev:2004:SSC

[DW04] Shlomi Dolev and Jennifer L. Welch. Self-stabilizing clock synchronization in the presence of Byzantine faults. Journal of the ACM, 51(5):780–799, September 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

deWitte:1958:EII

[dWF58] Leendeert de Witte and Kenneth P. Fournier. Evaluation of integrals involving combinations of Bessel functions and circular functions. Journal of the ACM, 5(2):119–126, April 1958. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Donald:1993:KMP

[DXCR93] Bruce Donald, Patrick Xavier, John Canny, and John Reif. Kinodynamic motion planning. Journal of the ACM, 40(5):1048–1066, November 1993. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/174150.html.

Dyer:1968:GMM

[Dye68] James Dyer. Generalized multistep methods in satellite orbit computation. Journal of the ACM, 15(4):712–719, October 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Earnest:1974:STC

[Ear74] Christopher Earnest. Some topics in code optimization. Journal of the ACM, 21(1):76–102, January 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Earnest:1972:AGO

[EBA72] C. P. Earnest, K. G. Balke, and J. Anderson. Analysis of graphs by ordering of nodes. Journal of the ACM, 19(1):23–42, January 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Edwards:1964:CPP

[EC64] A. Wood Edwards and Robert L. Chambers. Can A priori probabilities help in character recognition? Journal of the ACM, 11 (4):465–470, October 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Echols:1968:SIL

[EC68] Robert E. Echols and Leon Cooper. Solution of integer linear programming problems by direct search. Journal of the ACM, 15(1):75–84, January 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Edalat:2022:SAL

[Eda22] Abbas Edalat. Smooth approximation of Lipschitz maps and their subgradients. Journal of the ACM, 69(1):8:1–8:32, February 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3481805.

Edmundson:1969:NMA

[Edm69] H. P. Edmundson. New methods in automatic extracting. Journal of the ACM, 16(2):264–285, April 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Edwards:1954:SAM

[Edw54] C. M. Edwards. Survey of analog multiplication schemes. Journal of the ACM, 1(1):27–35, January 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Elias:1975:CSS

[EF75] Peter Elias and Richard A. Flower. The complexity of some simple retrieval problems. Journal of the ACM, 22(3):367–379, July 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Engelfriet:1989:PSV

[EF89] Joost Engelfriet and Gilberto Filé. Passes, sweeps, and visits in attribute grammars. Journal of the ACM, 36(4):841–869, October 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/76366.html.

Ephraim:2022:SSP

[EFKP22] Naomi Ephraim, Cody Freitag, Ilan Komargodski, and Rafael Pass. SPARKs: Succinct parallelizable arguments of knowledge. Journal of the ACM, 69(5):31:1–31:??, October 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3549523.

Ellen:2007:TLB

[EFR07] Faith Ellen, Panagiota Fatourou, and Eric Ruppert. Time lower bounds for implementations of multi-writer snapshots. Journal of the ACM, 54(6):30:1–30:34, December 2007. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Evfimievski:2010:EP

[EFW10] Alexandre Evfimievski, Ronald Fagin, and David Woodruff. Epistemic privacy. Journal of the ACM, 58(1):2:1–2:45, December 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ernst:1982:MDC

[EG82] George W. Ernst and Michael M. Goldstein. Mechanical discovery of classes of problem-solving strategies. Journal of the ACM, 29(1):1–23, January 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Eiter:1995:CLB

[EG95] Thomas Eiter and Georg Gottlob. The complexity of logicbased abduction. Journal of the ACM, 42(1):3–42, January 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/200838.html.

Eiter:2000:ESO

[EGG00] Thomas Eiter, Yuri Gurevich, and Georg Gottlob. Existential second-order logic over strings. Journal of the ACM, 47 (1):77–131, January 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/citations/journals/jacm/2000-47-1/p77-eiter/.

Eppstein:1992:SDPa

[EGGI92a] David Eppstein, Zvi Galil, Raffaele Giancarlo, and Giuseppe F. Italiano. Sparse dynamic programming I: Linear cost functions. Journal of the ACM, 39(3):519–545, July 1992. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
146650.html.

Eppstein:1992:SDPb

[EGGI92b] David Eppstein, Zvi Galil, Raffaele Giancarlo, and Giuseppe F. Italiano. Sparse dynamic programming II: Convex and concave cost functions. Journal of the ACM, 39(3):546–567, July 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/146656.html.

Estein:1997:STS

[EGIN97] David Estein, Zvi Galil, Giusee F. Italiano, and Amnon Nissenzweig. Sparsification — a technique for speeding up dynamic graph algorithms. Journal of the ACM, 44(5):669–696, September 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org:80/pubs/citations/
journals/jacm/1997-44-5/p669-eppstein/.

Esparza:2016:PVA

[EGM16] Javier Esparza, Pierre Ganty, and Rupak Majumdar. Parameterized verification of asynchronous shared-memory systems. Journal of the ACM, 63(1):10:1–10:??, March 2016. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Emerson:1986:SNR

[EH86] E. Allen Emerson and Joseph Y. Halpern. “sometimes” and “not never” revisited: On branching versus linear time temporal logic. Journal of the ACM, 33(1):151–178, January 1986. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
4999.html.

Ehrlich:1958:NMS

[Ehr58] L. W. Ehrlich. A numerical method of solving a heat flow problem with moving boundary. Journal of the ACM, 5(2):161–176, April 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ehrlich:1959:MCS

[Ehr59] Louis W. Ehrlich. Monte Carlo solutions of boundary value problems involving the difference analogue of ∂^{2}u/∂x^{2} + ∂^{2}u/∂y^{2} + (K/y)(∂u/∂y) = 0. Journal of the ACM, 6(2):204–218, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ehrlich:1973:LAG

[Ehr73] Gideon Ehrlich. Loopless algorithms for generating permutations, combinations, and other combinatorial configurations. Journal of the ACM, 20(3):500–513, July 1973. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ehrich:1982:TSI

[Ehr82] H.-D. Ehrich. On the theory of specification, implementation, and parametrization of abstract data types. Journal of the ACM, 29(1):206–227, January 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Eades:1984:SHP

[EHR84] Peter Eades, Michael Hickey, and Ronald C. Read. Some Hamilton paths and a minimal change algorithm. Journal of the ACM, 31(1):19–29, January 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Edmonds:1972:TIA

[EK72] Jack Edmonds and Richard M. Karp. Theoretical improvements in algorithmic efficiency for network flow problems. Journal of the ACM, 19(2):248–264, April 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Esparza:2010:NPA

[EKL10] Javier Esparza, Stefan Kiefer, and Michael Luttenberger. Newtonian program analysis. Journal of the ACM, 57(6):33:1–33:47, October 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Emek:2020:AGN

[EKLS20] Yuval Emek, Shay Kutten, Ron Lavi, and Yangguang Shi. Approximating generalized network design under (dis)economies of scale with applications to energy efficiency. Journal of the ACM, 67(1):7:1–7:33, February 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3377387.

Esparza:2020:UTL

[EKS20] Javier Esparza, Jan Kretínský, and Salomon Sickert. A unified translation of linear temporal logic to ω-automata. Journal of the ACM, 67(6):33:1–33:61, November 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3417995.

Eldred:1959:TRB

[Eld59] Richard D. Eldred. Test routines based on symbolic logical statements. Journal of the ACM, 6(1):33–36, January 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Elgot:1954:SVT

[Elg54] Calvin C. Elgot. On single vs. triple address computing machines. Journal of the ACM, 1(3):118–123, July 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Elias:1974:ESR

[Eli74] Peter Elias. Efficient storage and retrieval by content and address of static files. Journal of the ACM, 21(2):246–260, April 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Elkin:2020:SDD

[Elk20a] M. Elkin. A simple deterministic distributed MST algorithm with near-optimal time and message complexities. Journal of the ACM, 67(2):13:1–13:15, May 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3380546.

Elkin:2020:DES

[Elk20b] Michael Elkin. Distributed exact shortest paths in sublinear time. Journal of the ACM, 67(3):15:1–15:36, June 2020. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3387161.

Ellis:1972:HPP

[Ell72] Clarence A. Ellis. The halting problem for probabilistic contextfree generators. Journal of the ACM, 19(3):396–399, July 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Eiter:2007:CAA

[EM07] Thomas Eiter and Kazuhisa Makino. On computing all abductive explanations from a propositional Horn theory. Journal of the ACM, 54(5):24:1–24:54, October 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Emanuel:1963:WSC

[Ema63] George Emanuel. The Wilf stability criterion for numerical integration. Journal of the ACM, 10:557–561, 1963. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Even:2000:DCA

[ENRS00] Guy Even, Joseph (Seffi) Naor, Satish Rao, and Baruch Schieber. Divide-and-conquer approximation algorithms via spreading metrics. Journal of the ACM, 47(4):585–616, July 2000. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/citations/journals/jacm/
2000-47-4/p585-even/.

Even:1972:PGT

[EPL72] S. Even, A. Pnueli, and A. Lempel. Permutation graphs and transitive graphs. Journal of the ACM, 19(3):400–410, July 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ehrhard:2018:FAP

[EPT18] Thomas Ehrhard, Michele Pagani, and Christine Tasson. Full abstraction for probabilistic PCF. Journal of the ACM, 65(4): 23:1–23:??, August 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Elgot:1964:RMA

[ER64a] C. C. Elgot and J. D. Rutledge. RS-machines with almost blank tape. Journal of the ACM, 11(3):313–337, July 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Elgot:1964:RAS

[ER64b] Calvin C. Elgot and Abraham Robinson. Random-access stored program machines, an approach to programming languages. Journal of the ACM, 11(4):365–399, October 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ehrenfeucht:1980:SEP

[ER80a] A. Ehrenfeucht and G. Rozenberg. The sequence equivalence problem is decidable for 0S systems. Journal of the ACM, 27 (4):656–663, October 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Engelfriet:1980:FPL

[ER80b] J. Engelfriet and G. Rozenberg. Fixed point languages, equality languages, and representation of recursively enumerable languages. Journal of the ACM, 27(3):499–518, July 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ercoli:1960:LEE

[Erc60] Paolo Ercoli. Letters to the Editor: Errors due to overflow in arithmetic operations. Communications of the Association for Computing Machinery, 3(12):A9, December 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [EV57].

Erdelyi:1966:ROL

[Erd66] Ivan Erdelyi. On the “reverse order law” related to the generalized inverse of matrix products. Journal of the ACM, 13(3): 439–443, July 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ernst:1969:SCS

[Ern69] George W. Ernst. Sufficient conditions for the success of GPS. Journal of the ACM, 16(4):517–533, October 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ehrenfeucht:1981:MRC

[ERR81] A. Ehrenfeucht, G. Rozenberg, and K. Ruohonen. A morphic representation of complements of recursively enumerable sets. Journal of the ACM, 28(4):706–714, October 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Esparza:2024:ENL

[ERS24] Javier Esparza, Rubén Rubio, and Salomon Sickert. Efficient normalization of linear temporal logic. Journal of the ACM, 71(2):16:1–16:??, April 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3651152.

Eisner:1976:MTE

[ES76] Mark J. Eisner and Dennis G. Severance. Mathematical techniques for efficient record segmentation in large shared databases. Journal of the ACM, 23(4):619–635, October 1976. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Even:1981:LED

[ES81] Shimon Even and Yossi Shiloach. An on-line edge-deletion problem. Journal of the ACM, 28(1):1–4, January 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Eager:1986:BHM

[ES86] Derek L. Eager and Kenneth C. Sevcik. Bound hierarchies for multiple-class queuing networks. Journal of the ACM, 33(1):179– 206, January 1986. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/4992.html.

Ezra:2009:UFT

[ES09] Esther Ezra and Micha Sharir. On the union of fat tetrahedra in three dimensions. Journal of the ACM, 57(1):2:1–2:23, November 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Elkin:2015:OES

[ES15] Michael Elkin and Shay Solomon. Optimal Euclidean spanners: Really short, thin, and lanky. Journal of the ACM, 62(5):35:1– 35:??, November 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

ElDin:2017:NOA

[ES17] Mohab Safey El Din and Éric Schost. A nearly optimal algorithm for deciding connectivity queries in smooth and bounded real algebraic sets. Journal of the ACM, 63(6):48:1–48:??, February 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Esch:1960:NSC

[Esc60] Robin E. Esch. A necessary and sufficient condition for stability of partial difference equation problems. Journal of the ACM, 7 (2):163–175, April 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Engelfriet:1980:SMC

[ESvL80] Joost Engelfriet, Erik Meineche Schmidt, and Jan van Leeuwen. Stack machines and classes of nonnested macro languages. Journal of the ACM, 27(1):96–117, January 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Even:1985:HCT

[ESY85] Shimon Even, Alan L. Selman, and Yacov Yacobi. Hard-core theorems for complexity classes. Journal of the ACM, 32(1):205– 217, January 1985. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/214111.html.

Even:1976:CPW

[ET76] S. Even and R. E. Tarjan. A combinatorial problem which is complete in polynomial space. Journal of the ACM, 23(4):710– 719, October 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ercoli:1957:EDO

[EV57] Paolo Ercoli and Roberto Vacca. Errors due to overflow in arithmetic operations particularly as regards FINAC electronic computer. Journal of the ACM, 4(4):450–455, October 1957. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See letter [Erc60].

Estrin:1962:CAO

[EV62a] G. Estrin and C. R. Viswanathan. Correction and addendum: “Organization of a ‘fixed-plus-variable’ structure computer for computation of eigenvalues and eigenvectors of real symmetric matrices”. Journal of the ACM, 9(4):522, October 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [EV62b].

Estrin:1962:OFP

[EV62b] G. Estrin and C. R. Viswanathan. Organization of a “fixed-plus-variable” structure computer for computation of eigenvalues and eigenvectors of real symmetric matrices. Journal of the ACM, 9 (1):41–60, January 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [EV62a].

Easton:1975:ECC

[EW75] M. C. Easton and C. K. Wong. The effect of a capacity constraint on the minimal cost of a partition. Journal of the ACM, 22(4): 441–449, October 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Etessami:2009:RMC

[EY09] Kousha Etessami and Mihalis Yannakakis. Recursive Markov chains, stochastic grammars, and monotone systems of nonlinear equations. Journal of the ACM, 56(1):1:1–1:66, January 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Etessami:2015:RMD

[EY15] Kousha Etessami and Mihalis Yannakakis. Recursive Markov decision processes and recursive stochastic games. Journal of the ACM, 62(2):11:1–11:??, May 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Feldman:1981:MBF

[FACP81] R. M. Feldman, G. W. Adkins, G. L. Curry, and U. W. Pooch. Measurement bias in feedback queues. Journal of the ACM, 28 (2):351–357, April 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fagin:1982:HCD

[Fag82] Ronald Fagin. Horn clauses and database dependencies. Journal of the ACM, 29(4):952–985, October 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fagin:1983:DAH

[Fag83] Ronald Fagin. Degrees of acyclicity for hypergraphs and relational database schemes. Journal of the ACM, 30(3):514–550, July 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fairthorne:1967:MIF

[Fai67] Robert A. Fairthorne. Morphology of “information flow”. Journal of the ACM, 14(4):710–719, October 1967. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Faigle:1985:OLO

[Fai85] Ulrich Faigle. On ordered languages and the optimization of linear functions by greedy algorithms. Journal of the ACM, 32 (4):861–870, October 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/4998.html.

Falkoff:1962:APS

[Fal62] A. D. Falkoff. Algorithms for parallel-search memories. Journal of the ACM, 9(4):488–511, October 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Farr:1963:LPS

[Far63] Edwin H. Farr. Lattice properties of sequential machines. Journal of the ACM, 10:365–385, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Frazer:1972:BOM

[FB72] W. D. Frazer and B. T. Bennett. Bounds on optimal merge performance, and a strategy for optimality. Journal of the ACM, 19(4):641–648, October 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fuller:1975:ADS

[FB75] Samuel H. Fuller and Forest Baskett. An analysis of drum storage units. Journal of the ACM, 22(1):83–105, January 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Also published in/as: Stanford, DSL, SEL TR.26,.

Frisch:2008:SSD

[FCB08] Alain Frisch, Giuseppe Castagna, and Véronique Benzaken. Semantic subtyping: Dealing set-theoretically with function, union, intersection, and negation types. Journal of the ACM, 55(4): 19:1–19:64, September 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Farach-Colton:2000:SCS

[FCFM00] Martin Farach-Colton, Paolo Ferragina, and S. Muthukrishnan. On the sorting-complexity of suffix tree construction. Journal of the ACM, 47(6):987–1011, 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/citations/journals/jacm/
2000-47-6/p987-farach-colton/.

Farina:2022:SUN

[FCMG22] Gabriele Farina, Andrea Celli, Alberto Marchesi, and Nicola Gatti. Simple uncoupled no-regret learning dynamics for extensive-form correlated equilibrium. Journal of the ACM, 69 (6):41:1–41:??, December 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3563772.

Fagin:1976:IMR

[FE76] Ronald Fagin and Malcolm C. Easton. The independence of miss ratio on page size. Journal of the ACM, 23(1):128–146, January 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Feit:1984:FAT

[Fei84] Sidnie Dresher Feit. A fast algorithm for the two-variable integer programming problem. Journal of the ACM, 31(1):99–113, January 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Feige:1998:TAS

[Fei98] Uriel Feige. A threshold of ln n for approximating set cover. Journal of the ACM, 45(4):634–652, July 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:/

/www.acm.org:80/pubs/citations/journals/jacm/1998-45-
4/p634-feige/.

Feigenbaum:2003:SCG

[Fei03] Edward A. Feigenbaum. Some challenges and grand challenges for computational intelligence. Journal of the ACM, 50(1):32– 40, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ferguson:1960:IOB

[Fer60] David E. Ferguson. Input-output buffering and Fortran. Journal of the ACM, 7(1):1–9, January 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ferguson:1964:MCI

[Fer64] James Ferguson. Multivariable curve interpolation. Journal of the ACM, 11(2):221–228, April 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Flum:2002:QET

[FFG02] Jörg Flum, Markus Frick, and Martin Grohe. Query evaluation via tree-decompositions. Journal of the ACM, 49(6):716–752, November 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Flesca:2008:MXQ

[FFM08] S. Flesca, F. Furfaro, and E. Masciari. On the minimization of XPath queries. Journal of the ACM, 55(1):2:1–2:46, February 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Feldstein:1976:CED

[FG76] Alan Feldstein and Richard Goodman. Convergence estimates for the distribution of trailing digits. Journal of the ACM, 23 (2):287–297, April 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fussenegger:1979:CAL

[FG79] Frank Fussenegger and Harold N. Gabow. A counting approach to lower bounds for selection problems. Journal of the ACM, 26 (2):227–238, April 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ferragina:1999:SBT

[FG99] Paolo Ferragina and Roberto Grossi. The string B-tree: a new data structure for string search in external memory and its applications. Journal of the ACM, 46(2):236–280, March 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org:80/pubs/citations/
journals/jacm/1999-46-2/p236-ferragina/.

Frick:2001:DFO

[FG01] Markus Frick and Martin Grohe. Deciding first-order properties of locally tree-decomposable structures. Journal of the ACM, 48 (6):1184–1206, November 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Franceschini:2005:PSC

[FG05] Gianni Franceschini and Viliam Geffert. An in-place sorting with O(n log n) comparisons and O(n) moves. Journal of the ACM, 52(4):515–537, July 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fauconnier:2010:TFD

[FG10] Carole Delporte-Gallet Hugues Fauconnier and Rachid Guerraoui. Tight failure detection bounds on atomic object implementations. Journal of the ACM, 57(4):22:1–22:32, April 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fearnley:2023:CGD

[FGHS23] John Fearnley, Paul Goldberg, Alexandros Hollender, and Rahul Savani. The complexity of gradient descent: CLS = PPAD ∩ PLS. Journal of the ACM, 70(1):7:1–7:??, February 2023. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3568163.

Fearnley:2025:CCK

[FGHS25] John Fearnley, Paul Goldberg, Alexandros Hollender, and Rahul Savani. The complexity of computing KKT solutions of quadratic programs. Journal of the ACM, 72(5):31:1–31:48, October 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fomin:2009:MCA

[FGK09] Fedor V. Fomin, Fabrizio Grandoni, and Dieter Kratsch. A measure & conquer approach for the analysis of exact algorithms. Journal of the ACM, 56(5):25:1–25:32, August 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fayolle:1977:SOC

[FGL77] Guy Fayolle, Erol Gelenbe, and Jacques Labetoulle. Stability and optimal control of the packet switching broadcast channel. Journal of the ACM, 24(3):375–386, July 1977. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Feige:1996:IPH

[FGL^{+}96] Uriel Feige, Shafi Goldwasser, Laszlo Lovász, Shmuel Safra, and Mario Szegedy. Interactive proofs and the hardness of approximating cliques. Journal of the ACM, 43(2):268–292, March 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
jacm/226652.html.

Fomin:2019:EAM

[FGLS19] Fedor V. Fomin, Serge Gaspers, Daniel Lokshtanov, and Saket Saurabh. Exact algorithms via monotone local search. Journal of the ACM, 66(2):8:1–8:??, April 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=3284176.

Furst:1988:FMG

[FGM88] Merrick L. Furst, Jonathan L. Gross, and Lyle A. McGeoch. Finding a maximum-genus graph imbedding. Journal of the ACM, 35(3):523–534, July 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/44485.html.

Ferragina:2005:BTC

[FGMS05] Paolo Ferragina, Raffaele Giancarlo, Giovanni Manzini, and Marinella Sciortino. Boosting textual compression in optimal linear time. Journal of the ACM, 52(4):688–713, July 2005. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Farber:1964:SSM

[FGP64] D. J. Farber, R. E. Griswold, and I. P. Polonsky. SNOBOL, a string manipulation language. Journal of the ACM, 11(1):21– 30, January 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Feldman:2017:SAL

[FGR^{+}17] Vitaly Feldman, Elena Grigorescu, Lev Reyzin, Santosh S. Vempala, and Ying Xiao. Statistical algorithms and a lower bound for detecting planted cliques. Journal of the ACM, 64(2):8:1– 8:??, June 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Franklin:2000:EGG

[FGY00] Matthew Franklin, Zvi Galil, and Moti Yung. Eavesdropping games: a graph-theoretic approach to privacy in distributed systems. Journal of the ACM, 47(2):225–243, March 2000. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/citations/journals/jacm/
2000-47-2/p225-franklin/.

Feng:2021:FSC

[FGYZ21] Weiming Feng, Heng Guo, Yitong Yin, and Chihao Zhang. Fast sampling and counting k-SAT solutions in the local lemma regime. Journal of the ACM, 68(6):40:1–40:42, December 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3469832.

Fisher:1980:DLC

[FH80] Marshall L. Fisher and Dorit S. Hochbaum. Database location in computer networks. Journal of the ACM, 27(4):718–735, October 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fagin:1994:RAK

[FH94] Ronald Fagin and Joseph Y. Halpern. Reasoning about knowledge and probability. Journal of the ACM, 41(2):340–367, March 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/

0004-5411/174658.html. See corrigendum [FH98].

Fagin:1998:CRA

[FH98] Ronald Fagin and Joseph Y. Halpern. Corrigendum: “Reasoning about Knowledge and Probability”. Journal of the ACM, 45(1): 214, January 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [FH94].

Friedman:2001:PMD

[FH01] Nir Friedman and Joseph Y. Halpern. Plausibility measures and default reasoning. Journal of the ACM, 48(4):648–685, July 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fleck:1972:CSA

[FHO72] A. C. Fleck, S. T. Hedetniemi, and R. H. Oehmke. S-semigroups of automata. Journal of the ACM, 19(1):3–10, January 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fich:1998:SCR

[FHS98] Faith Fich, Maurice Herlihy, and Nir Shavit. On the space complexity of randomized synchronization. Journal of the ACM, 45(5):843–862, September 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org:80/pubs/citations/journals/jacm/1998-45-5/p1-
fich/.

Fawzi:2013:LDN

[FHS13] Omar Fawzi, Patrick Hayden, and Pranab Sen. From lowdistortion norm embeddings to explicit uncertainty relations and efficient information locking. Journal of the ACM, 60(6):44:1– 44:??, November 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fahrbach:2022:EWO

[FHTZ22] Matthew Fahrbach, Zhiyi Huang, Runzhou Tao, and Morteza Zadimoghaddam. Edge-weighted online bipartite matching. Journal of the ACM, 69(6):45:1–45:??, December 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3556971.

Fagin:1991:MTA

[FHV91] Ronald Fagin, Joseph Y. Halpern, and Moshe Y. Vardi. A modeltheoretic analysis of knowledge. Journal of the ACM, 38(2):382– 428, April 1991. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/128680.html. A preliminary version ap-

peared in Proc. 25th IEEE Symposium on Foundations of Computer Science, 1984.

Fagin:1992:WCM

[FHV92] Ronald Fagin, Joseph Y. Halpern, and Moshe Y. Vardi. What can machines know? On the properties of knowledge in distributed systems. Journal of the ACM, 39(2):328–376, April 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/150945.html.

Fike:1959:NPC

[Fik59] C. T. Fike. Note on the practical computation of proper values. Journal of the ACM, 6(3):360–362, July 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fisher:1956:HOD

[Fis56] Michael E. Fisher. Higher order differences in the analogue solution of partial differential equations. Journal of the ACM, 3 (4):325–347, October 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fisher:1958:PMA

[Fis58] Michael E. Fisher. Proposed methods for the analog solution of Fredholm’s integral equation. Journal of the ACM, 5(4):357– 369, October 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fischer:1965:GPO

[Fis65a] Patrick C. Fischer. Generation of primes by a one-dimensional real-time iterative array. Journal of the ACM, 12(3):388–394, July 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fischer:1965:FTM

[Fis65b] Patrick C. Fischer. On formalisms for Turing machines. Journal of the ACM, 12(4):570–580, October 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fischer:2003:J

[Fis03] Michael J. Fischer. JACM 1983–1986. Journal of the ACM, 50(1):17, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fitzpatrick:1960:SBR

[Fit60] G. B. Fitzpatrick. Synthesis of binary ring counters of given periods. Journal of the ACM, 7(3):287–297, July 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Freiberger:1960:CFF

[FJ60] Walter F. Freiberger and Richard H. Jones. Computation of the frequency function of a quadratic form in random normal variables. Journal of the ACM, 7(3):245–250, July 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fiorini:2025:IPB

[FJWY25] Samuel Fiorini, Gwenaël Joret, Stefan Weltge, and Yelena Yuditsky. Integer programs with bounded subdeterminants and two nonzeros per row. Journal of the ACM, 72(1):3:1–3:??, February 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3695985.

Fuchs:1985:MCT

[FK85] Ken Fuchs and Dennis Kafura. Memory-constrained task scheduling on a network of dual processors. Journal of the ACM, 32(1):102–129, January 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/2456.html.

Farach:1999:EAI

[FK99] Martin Farach and Sampath Kannan. Efficient algorithms for inverting evolution. Journal of the ACM, 46(4):437–449, July 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/citations/
journals/jacm/1999-46-4/p437-farach/.

Franek:2015:RSS

[FK15] Peter Franek and Marek Krcál. Robust satisfiability of systems of equations. Journal of the ACM, 62(4):26:1–26:??, August 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fraigniaud:2016:OAL

[FK16] Pierre Fraigniaud and Amos Korman. An optimal ancestry labeling scheme with applications to XML trees and universal posets. Journal of the ACM, 63(1):6:1–6:??, March 2016. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fagin:2011:PDE

[FKK11] Ronald Fagin, Benny Kimelfeld, and Phokion G. Kolaitis. Probabilistic data exchange. Journal of the ACM, 58(4):15:1–15:55, July 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fekete:1998:ISC

[FKL98] Alan Fekete, M. Frans Kaashoek, and Nancy Lynch. Implementing sequentially consistent shared objects using broadcast and point-to-point communication. Journal of the ACM, 45(1):35– 69, January 1998. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org:80/pubs/
citations/journals/jacm/1998-45-1/p35-fekete/.

Fraigniaud:2013:TCT

[FKP13] Pierre Fraigniaud, Amos Korman, and David Peleg. Towards a complexity theory for local distributed computing. Journal of the ACM, 60(5):35:1–35:26, October 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fraigniaud:2019:PBS

[FKR19] Pierre Fraigniaud, Amos Korman, and Yoav Rodeh. Parallel Bayesian search with no coordination. Journal of the ACM, 66 (3):17:1–17:??, June 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3304111.

Fagin:2015:DSF

[FKRV15] Ronald Fagin, Benny Kimelfeld, Frederick Reiss, and Stijn Vansummeren. Document spanners: a formal approach to information extraction. Journal of the ACM, 62(2):12:1–12:??, May 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fredman:1984:SST

[FKS84] Michael L. Fredman, János Komlós, and Endre Szemerédi. Storing a sparse table with O(1) worst case access time. Journal of the ACM, 31(3):538–544, July 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Extends the work of Tarjan and Yao [TY79], using a two-level data structure, the first containing pointers to the second, and the second containing blocks accessible by a perfect hashing function.

Freivalds:1995:IFL

[FKS95] Rūsiņš Freivalds, Efim Kinber, and Carl H. Smith. On the impact of forgetting on learning machines. Journal of the ACM, 42(6): 1146–1168, November 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/227685.html.

Frieze:2004:FMC

[FKV04] Alan Frieze, Ravi Kannan, and Santosh Vempala. Fast Monte-Carlo algorithms for finding low-rank approximations. Journal of the ACM, 51(6):1025–1041, November 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fenner:1974:SNB

[FL74] T. I. Fenner and G. Loizou. Some new bounds on the condition numbers of optimally scaled matrices. Journal of the ACM, 21 (3):514–524, July 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Frederickson:1987:ELS

[FL87] Greg N. Frederickson and Nancy A. Lynch. Electing a leader in a synchronous ring. Journal of the ACM, 34(1):98–115, January 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/7919.html.

Fellows:1988:NTP

[FL88] Michael R. Fellows and Michael A. Langston. Nonconstructive tools for proving polynomial-time decidability. Journal of the ACM, 35(3):727–739, July 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/44491.html.

Fan:2002:XIC

[FL02] Wenfei Fan and Leonid Libkin. On XML integrity constraints in the presence of DTDs. Journal of the ACM, 49(3):368–406, May 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Friedrichs:2018:PMT

[FL18] Stephan Friedrichs and Christoph Lenzen. Parallel metric tree embedding based on an algebraic view on Moore-Bellman-Ford. Journal of the ACM, 65(6):43:1–43:??, November 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fleck:1962:IGA

[Fle62] A. C. Fleck. Isomorphism groups of automata. Journal of the ACM, 9(4):469–476, October 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fleck:1965:AGA

[Fle65] A. C. Fleck. On the automorphism group of an automaton. Journal of the ACM, 12(4):566–569, October 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Flinn:1960:MFM

[Fli60] E. A. Flinn. A modification of Filon’s method of numerical integration. Journal of the ACM, 7(2):181–184, April 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ferragina:2009:CIL

[FLMM09] Paolo Ferragina, Fabrizio Luccio, Giovanni Manzini, and S. Muthukrishnan. Compressing and indexing labeled trees, with applications. Journal of the ACM, 57(1):4:1–4:33, November 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fekete:1993:IIR

[FLMS93] Alan Fekete, Nancy Lynch, Yishay Mansour, and John Spinelli. The impossibility of implementing reliable communication in the face of crashes. Journal of the ACM, 40(5):1087–1107, November 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/169676.html.

Flores:1960:CTA

[Flo60] Ivan Flores. Computer time for address calculation sorting. Journal of the ACM, 7(4):389–409, October 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Flores:1961:AIC

[Flo61a] Ivan Flores. Analysis of internal computer sorting. Journal of the ACM, 8(1):41–80, January 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [Goe61].

Floyd:1961:DLS

[Flo61b] Robert W. Floyd. A descriptive language for symbol manipulation. Journal of the ACM, 8(4):579–584, October 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Floyd:1963:SAO

[Flo63] Robert W. Floyd. Syntactic analysis and operator precedence. Journal of the ACM, 10:316–333, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Flores:1964:DWT

[Flo64] Ivan Flores. Derivation of a waiting-time factor for a multiplebank memory. Journal of the ACM, 11(3):265–282, July 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Floyd:1967:NA

[Flo67] Robert W. Floyd. Nondeterministic algorithms. Journal of the ACM, 14(4):636–644, October 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fortune:1983:ESS

[FLO83] Steven Fortune, Daniel Leivant, and Michael O’Donnell. The expressiveness of simple and second-order type structures. Journal of the ACM, 30(1):151–185, January 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fijalkow:2025:MPA

[FLO^{+}25] Nathanael Fijalkow, Engel Lefaucheux, Pierre Ohlmann, Joël Ouaknine, Amaury Pouly, and James Worrell. On the monniaux problem in abstract interpretation. Journal of the ACM, 72(2): 11:1–11:??, April 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fischer:1985:IDC

[FLP85] Michael J. Fischer, Nancy A. Lynch, and Michael S. Paterson. Impossibility of distributed consensus with one faulty process. Journal of the ACM, 32(2):374–382, April 1985. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/

214121.html. This paper proves that every completely asynchronous, deterministic algorithm for Byzantine agreement has the possibility of nontermination, even with only one faulty processor. This impossibility result does not hold in the synchronous case. For completely asynchronous probabilistic algorithms, the problem is avoided since termination is only required with probability 1. See Section xxx for an example of such a probabilistic algorithm for asynchronous Byzantine agreement.

Fomin:2016:ECR

[FLPS16] Fedor V. Fomin, Daniel Lokshtanov, Fahad Panolan, and Saket Saurabh. Efficient computation of representative families with applications in parameterized and exact algorithms. Journal of the ACM, 63(4):29:1–29:??, November 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fomin:2018:EGM

[FLS18] Fedorr V. Fomin, Daniel Lokshtanov, and Saket Saurabh. Excluded grid minors and efficient polynomial-time approximation schemes. Journal of the ACM, 65(2):10:1–10:??, March 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fleisig:1974:IME

[FLSY74] S. Fleisig, D. Loveland, A. K. Smiley III, and D. L. Yarmush. An implementation of the model elimination proof procedure. Journal of the ACM, 21(1):124–139, January 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fortnow:2005:TSL

[FLvMV05] Lance Fortnow, Richard Lipton, Dieter van Melkebeek, and Anastasios Viglas. Time-space lower bounds for satisfiability. Journal of the ACM, 52(6):835–865, November 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Frazer:1970:SSA

[FM70] W. D. Frazer and A. C. McKellar. Samplesort: a sampling approach to minimal storage tree sorting. Journal of the ACM, 17(3):496–507, July 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fishman:1979:EMC

[FM79] George S. Fishman and Louis R. Moore. Estimating the mean of a correlated binary sequence with an application to discrete event simulation. Journal of the ACM, 26(1):82–94, January 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ferragina:2005:ICT

[FM05] Paolo Ferragina and Giovanni Manzini. Indexing compressed text. Journal of the ACM, 52(4):552–581, July 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fayolle:1980:SPA

[FMI80] G. Fayolle, I. Mitrani, and R. Iasnogorodski. Sharing a processor among many job classes. Journal of the ACM, 27(3):519–532, July 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ferrere:2019:RTL

[FMNP19] Thomas Ferrère, Oded Maler, Dejan Nicković, and Amir Pnueli. From real-time logic to timed automata. Journal of the ACM, 66 (3):19:1–19:??, June 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3286976.

Fiorini:2015:ELB

[FMP^{+}15] Samuel Fiorini, Serge Massar, Sebastian Pokutta, Hans Raj Tiwary, and Ronald De Wolf. Exponential lower bounds for polytopes in combinatorial optimization. Journal of the ACM, 62 (2):17:1–17:??, May 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fischer:1972:RTS

[FMR72] Patrick C. Fischer, Albert R. Meyer, and Arnold L. Rosenberg. Real-time simulation of multihead tape units. Journal of the ACM, 19(4):590–607, October 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Frandsen:1997:DWP

[FMS97] Gudmund Skovbjerg Frandsen, Peter Bro Miltersen, and Sven Skyum. Dynamic word problems. Journal of the ACM, 44 (2):257–271, March 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fisher:2010:NDD

[FMW10] Kathleen Fisher, Yitzhak Mandelbaum, and David Walker. The next 700 data description languages. Journal of the ACM, 57(2): 10:1–10:51, January 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fonlupt:1993:DPG

[FN93] Jean Fonlupt and Armand Nachef. Dynamic programming and the graphical traveling salesman problem. Journal of the ACM, 40(5):1165–1187, November 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/169803.html.

Fagin:2010:SIS

[FN10] Ronald Fagin and Alan Nash. The structure of inverses in schema mappings. Journal of the ACM, 57(6):31:1–31:57, October 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Feldman:2023:OWC

[FNFSZ23] Moran Feldman, Ashkan Norouzi-Fard, Ola Svensson, and Rico Zenklusen. The one-way communication complexity of submodular maximization with applications to streaming and robustness. Journal of the ACM, 70(4):24:1–24:??, August 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3588564.

Fugger:2021:TBA

[FNS21] Matthias Függer, Thomas Nowak, and Manfred Schwarz. Tight bounds for asymptotic and approximate consensus. Journal of the ACM, 68(6):46:1–46:35, December 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3485242.

Fiat:1992:NH

[FNSS92] Amos Fiat, Moni Naor, Jeanette P. Schmidt, and Alan Siegel. Nonoblivious hashing. Journal of the ACM, 39(4):764–782, October 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/146591.html.

Foreword:2016:IPF

[For16] Foreword. Invited paper foreword. Journal of the ACM, 63(4): 38:1–38:??, November 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Faenza:2014:SWS

[FOS14] Yuri Faenza, Gianpaolo Oriolo, and Gautier Stauffer. Solving the weighted stable set problem in claw-free graphs via decomposition. Journal of the ACM, 61(4):20:1–20:??, July 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fox:1970:ALP

[Fox70] B. L. Fox. Accelerating list processing in discrete programming. Journal of the ACM, 17(2):383–384, April 1970. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Flajolet:1986:PMR

[FP86] Philippe Flajolet and Claude Puech. Partial match retrieval of multidimensional data. Journal of the ACM, 33(2):371– 407, April 1986. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/5453.html.

Fischer:1994:FPQ

[FP94] Michael J. Fischer and Michael S. Paterson. Fishspear: a priority queue algorithm. Journal of the ACM, 41(1):3–30, January 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/174645.html.

Fleming:2022:RTL

[FPPR22] Noah Fleming, Denis Pankratov, Toniann Pitassi, and Robert Robere. Random Θ(log n)-CNFs are hard for cutting planes. Journal of the ACM, 69(3):19:1–19:32, June 2022. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3486680.

Fernandes:1998:EDV

[FPS98] Paulo Fernandes, Brigitte Plateau, and William J. Stewart. Efficient descriptor-vector multiplications in stochastic automata networks. Journal of the ACM, 45(3):381–414, May 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org:80/pubs/citations/
journals/jacm/1998-45-3/p381-fernandes/.

Freudenstein:1963:NSS

[FR63] Ferdinand Freudenstein and Bernard Roth. Numerical solution of systems of nonlinear equations. Journal of the ACM, 10:550– 556, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Frank:1958:FZA

[Fra58a] Werner L. Frank. Finding zeros of arbitrary functions. Journal of the ACM, 5(2):154–160, April 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Franklin:1958:NSC

[Fra58b] J. N. Franklin. On the numerical solution of characteristic equations in flutter analysis. Journal of the ACM, 5(1):45–51, January 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Frank:1960:SLS

[Fra60] Werner L. Frank. Solution of linear systems by Richardson’s method. Journal of the ACM, 7(3):274–286, July 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fraenkel:1961:UIC

[Fra61] Aviezri S. Fraenkel. The user of index calculus and Mersenne primes for the design of a high-speed digital multiplier. Journal of the ACM, 8(1):87–96, January 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fraser:1965:SMC

[Fra65] W. Fraser. A survey of methods for computing minimax and near-minimax polynomial approximations for functions of a single independent variable. Journal of the ACM, 12(3):295–314, July 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Frank:1969:AOD

[Fra69] H. Frank. Analysis and optimization of disk storage devices for time-sharing systems. Journal of the ACM, 16(4):602–620, October 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fraser:1976:API

[Fra76] Donald Fraser. Array permutation by index-digit permutation. Journal of the ACM, 23(2):298–309, April 1976. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Frederickson:1979:AAS

[Fre79] Greg N. Frederickson. Approximation algorithms for some postman problems. Journal of the ACM, 26(3):538–554, July 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fredman:1981:LBC

[Fre81] Michael L. Fredman. A lower bound on the complexity of orthogonal range queries. Journal of the ACM, 28(4):696–705, October 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fredman:1982:CMA

[Fre82a] Michael L. Fredman. The complexity of maintaining an array and computing its partial sums. Journal of the ACM, 29(1): 250–260, January 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Freuder:1982:SCB

[Fre82b] Eugene C. Freuder. A sufficient condition of backtrack-free search. Journal of the ACM, 29(1):24–32, January 1982. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Frederickson:1983:IDS

[Fre83] Greg N. Frederickson. Implicit data structures for the dictionary problem. Journal of the ACM, 30(1):80–94, January 1983. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Freuder:1985:SCB

[Fre85] Eugene C. Freuder. A sufficient condition for backtrack-bounded search. Journal of the ACM, 32(4):755–761, October 1985. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
4225.html.

Frederickson:1991:PGD

[Fre91] Greg N. Frederickson. Planar graph decomposition and all pairs shortest paths. Journal of the ACM, 38(1):162–204, January 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fredman:1999:EPH

[Fre99] Michael L. Fredman. On the efficiency of pairing heaps and related data structures. Journal of the ACM, 46(4):473–501, July 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL http://www.acm.org/pubs/articles/
journals/jacm/1999-46-4/p473-fredman/p473-fredman.pdf;
 http://www.acm.org/pubs/citations/journals/jacm/1999-
46-4/p473-fredman/.

Friend:1956:SEC

[Fri56] Edward Harry Friend. Sorting on electronic computer systems. Journal of the ACM, 3(3):134–168, July 1956. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Friedman:1962:DPC

[Fri62] Joyce Friedman. A decision procedure for computations of finite automata. Journal of the ACM, 9(3):315–323, July 1962. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Friedman:1963:CPS

[Fri63a] Joyce Friedman. A computer program for a solvable case of the decision problem. Journal of the ACM, 10:348–356, 1963. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Friedman:1963:SDP

[Fri63b] Joyce Friedman. A semi-decision procedure for the functional calculus. Journal of the ACM, 10(1):1–24, January 1963. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Friedli:1973:OCA

[Fri73] Armin Friedli. Optimal covering algorithms in methods of search for solving polynomial equations. Journal of the ACM, 20 (2):290–300, April 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Froese:1961:ERK

[Fro61] Charlotte Froese. An evaluation of Runge–Kutta type methods for higher order differential equations. Journal of the ACM, 8 (4):637–644, October 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Flajolet:2006:HWS

[FSV06] Philippe Flajolet, Wojciech Szpankowski, and Brigitte Vallée. Hidden word statistics. Journal of the ACM, 53(1):147–183, January 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fredman:1987:FHT

[FT87] Michael L. Fredman and Robert Endre Tarjan. Fibonacci heaps and their uses in improved network optimization algorithms. Journal of the ACM, 34(3):596–615, July 1987. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
28874.html.

Fich:1988:PCE

[FT88] Faith E. Fich and Martin Tompa. The parallel complexity of exponentiating polynomials over finite fields. Journal of the ACM, 35(3):651–667, July 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/44496.html.

Fulek:2022:AEC

[FT22] Radoslav Fulek and Csaba D. Tóth. Atomic embeddability, clustered planarity, and thickenability. Journal of the ACM, 69(2): 13:1–13:34, April 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/

doi/10.1145/3502264.

Floyd:1982:CRE

[FU82] Robert W. Floyd and Jeffrey D. Ullman. The compilation of regular expressions into integrated circuits. Journal of the ACM, 29(3):603–622, July 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fushimi:1990:RNG

[Fus90] Masanori Fushimi. Random number generation with the recursion X_{t} = X_{t−3p} ⊕ X_{t−3q}. Journal of Computational and Applied Mathematics, 31(1):105–118, July 24, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/037704279090341V.

Frumkin:2002:TBC

[FV02] Michael A. Frumkin and Rob F. Van der Wijngaart. Tight bounds on cache use for stencil operations on rectangular grids. Journal of the ACM, 49(3):434–453, May 2002. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Franaszek:1974:SDF

[FW74] P. A. Franaszek and T. J. Wagner. Some distribution-free aspects of paging algorithm performance. Journal of the ACM, 21(1):31– 39, January 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Fiduccia:1977:AHL

[FZ77] C. M. Fiduccia and Y. Zalcstein. Algebras having linear multiplicative complexities. Journal of the ACM, 24(2):311–331, April 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gabow:1976:EIE

[Gab76] Harold N. Gabow. An efficient implementation of Edmonds’ algorithm for maximum matching in graphs. Journal of the ACM, 23(2):221–234, April 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gabow:1982:ALA

[Gab82] Harold N. Gabow. An almost-linear algorithm for two-processor scheduling. Journal of the ACM, 29(3):766–780, July 1982. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gabow:2006:UEG

[Gab06] Harold N. Gabow. Using expander graphs to find vertex connectivity. Journal of the ACM, 53(5):800–844, September 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gabbay:2016:SCN

[Gab16] Murdoch J. Gabbay. Semantics out of context: Nominal absolute denotations for first-order logic and computation. Journal of the ACM, 63(3):25:1–25:??, September 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gaina:2020:FCH

[Gai20] Daniel Gaina. Forcing and calculi for hybrid logics. Journal of the ACM, 67(4):25:1–25:55, August 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3400294.

Galil:1981:SMR

[Gal81] Z. Galil. String matching in real time. Journal of the ACM, 28 (1):134–149, January 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Galil:1982:ALT

[Gal82] Zvi Galil. An almost linear-time algorithm for computing a dependency basis in a relational database. Journal of the ACM, 29 (1):96–102, January 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Galil:1995:CTO

[Gal95] Zvi Galil. A constant-time optimal parallel string-matching algorithm. Journal of the ACM, 42(4):908–918, July 1995. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
210341.html.

Gati:1983:CSP

[Gat83] Georg Gati. The complexity of solving polynomial equations by quadrature. Journal of the ACM, 30(3):637–640, July 1983. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gauss:1959:CMO

[Gau59] E. J. Gauss. A comparison of machine organizations by their performance of the iteration solution of linear equations. Journal of the ACM, 6(4):476–485, October 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [Gau60].

Gauss:1960:CCM

[Gau60] E. J. Gauss. Corrigendum: “A comparison of machine organizations by their performance of the iteration solution of linear equations”. Journal of the ACM, 7(2):188, April 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Gau59].

Gauss:1961:LLW

[Gau61a] E. J. Gauss. Locating the largest word in a file using a modified memory. Journal of the ACM, 8(3):418–425, July 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gautschi:1961:RCC

[Gau61b] Walter Gautschi. Recursive computation of certain integrals. Journal of the ACM, 8(1):21–40, January 1961. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gauss:1964:EPS

[Gau64] E. J. Gauss. Estimation of power spectral density by filters. Journal of the ACM, 11(1):98–103, January 1964. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gaver:1967:PMM

[Gav67] D. P. Gaver, Jr. Probability models for multiprogramming computer systems. Journal of the ACM, 14(3):423–438, July 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gaver:1971:ART

[Gav71] Donald P. Gaver. Analysis of remote terminal backlogs under heavy demand conditions. Journal of the ACM, 18(3):405–415, July 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gavish:1983:FAC

[Gav83] Bezalel Gavish. Formulations and algorithms for the capacitated minimal directed tree problem. Journal of the ACM, 30(1):118– 132, January 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Golomb:1965:BP

[GB65] Solomon W. Golomb and Leonard D. Baumert. Backtrack programming. Journal of the ACM, 12(4):516–524, October 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Grzymala-Busse:1969:APA

[GB69a] Jerzy W. Grzymala-Busse. Automorphisms of polyadic automata. Journal of the ACM, 16(2):208–219, April 1969. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [GB69b].

Grzymala-Busse:1969:EAP

[GB69b] Jerzy W. Grzymala-Busse. Errata: “Automorphisms of polyadic automata”. Journal of the ACM, 16(4):646, October 1969. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [GB69a].

Grzymala-Busse:1969:PRR

[GB69c] Jerzy W. Grzymala-Busse. On the periodic representations and the reducibility of periodic automata. Journal of the ACM, 16 (3):432–441, July 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [GB69c, GB70].

Grzymala-Busse:1970:EPR

[GB70] Jerzy W. Grzymala-Busse. Errata: “On the periodic representations and the reducibility of periodic automata”. Journal of the ACM, 17(4):739, October 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [GB69c].

Gill:1974:AEC

[GB74] John Gill and Manuel Blum. On almost everywhere complex recursive functions. Journal of the ACM, 21(3):425–435, July 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goguen:1992:IAM

[GB92] Joseph A. Goguen and Rod M. Burstall. Institutions: Abstract model theory for specification and programming. Journal of the ACM, 39(1):95–146, January 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/

147524.html. Draft, as Report ECS-LFCS-90-106, Computer Science Department, University of Edinburgh, January 1990; an ancestor is “Introducing Institutions,” in Proceedings, Logics of Programming Workshop, Edward Clarke and Dexter Kozen, editors, Springer Lecture Notes in Computer Science, Volume 164, pages 221–256, 1984.

Glass:1965:SSM

[GC65] H. Glass and L. Cooper. Sequential search: a method for solving constrained optimized problems. Journal of the ACM, 12(1):71– 82, January 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Greene:1984:CVA

[GE84] Jonathan W. Greene and Abbas El Gamal. Configuration of VLSI arrays in the presence of defects. Journal of the ACM, 31 (4):694–717, October 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gal:2001:MPU

[GE01] Avigdor Gal and Jonathan Eckstein. Managing periodically updated data in relational databases: a stochastic modeling approach. Journal of the ACM, 48(6):1141–1183, November 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gelenbe:1975:ACS

[Gel75] Erol Gelenbe. On approximate computer system models. Journal of the ACM, 22(2):261–269, April 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gelenbe:1979:OCI

[Gel79] Erol Gelenbe. On the optimum checkpoint interval. Journal of the ACM, 26(2):259–270, April 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gentleman:1978:SCR

[Gen78] W. Morven Gentleman. Some complexity results for matrix computations on parallel processors. Journal of the ACM, 25(1):112– 115, January 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gordon:1960:NMC

[GF60] N. L. Gordon and A. H. Flasterstein. A note on a method of computing the gamma function. Journal of the ACM, 7(4):387– 388, October 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gill:1967:PDS

[GF67] Arthur Gill and J. Robert Flexer. Periodic decomposition of sequential machines. Journal of the ACM, 14(4):666–676, October 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Greibach:1980:SPS

[GF80] S. A. Greibach and E. P. Friedman. Superdeterministic PDAs: a subclass with a decidable inclusion problem. Journal of the ACM, 27(4):675–700, October 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Greenberg:1987:EMC

[GFL87] Albert G. Greenberg, Philippe Flajolet, and Richard E. Ladner. Estimating the multiplicities of conflicts to speed their resolution in multiple access channels. Journal of the ACM, 34(2):289– 325, April 1987. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/23006.html. Reviews: Computing Re-

views, Vol. 30, No. 7.

Gardner:1961:LEO

[GG61] R. J. Gardner and T. H. Gosling. Letter to the editor: “Optimized subroutines”. Journal of the ACM, 8(4):648–649, October 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Den61].

Greibach:1972:MA

[GG72] Sheila Greibach and Seymour Ginsburg. Multitape AFA. Journal of the ACM, 19(2):193–221, April 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gabrielian:1974:GS

[GG74] Armen Gabrielian and Seymour Ginsburg. Grammar schemata. Journal of the ACM, 21(2):213–226, April 1974. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Garfinkel:1978:BTS

[GG78] R. S. Garfinkel and K. C. Gilbert. The bottleneck traveling salesman problem: Algorithms and probabilistic analysis. Journal of the ACM, 25(3):435–448, July 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Godoy:2013:HPD

[GG13a] Guillem Godoy and Omer Giménez. The HOM problem is decidable. Journal of the ACM, 60(4):23:1–23:44, August 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gottlob:2013:DCA

[GG13b] Georg Gottlob and Gianluigi Greco. Decomposing combinatorial auctions and set packing problems. Journal of the ACM, 60(4): 24:1–24:39, August 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Galanis:2017:AMP

[GGG^{+}17] Andreas Galanis, Andreas Göbel, Leslie Ann Goldberg, John Lapinskas, and David Richerby. Amplifiers for the Moran process. Journal of the ACM, 64(1):5:1–5:??, March 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Georgiou:2013:AG

[GGGK13] Chryssis Georgiou, Seth Gilbert, Rachid Guerraoui, and Dariusz R. Kowalski. Asynchronous gossip. Journal of the ACM, 60 (2):11:1–11:42, April 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ginsburg:1967:OWS

[GGH67a] Seymour Ginsburg, Sheila A. Greibach, and Michael A. Harrison. One-way stack automata. Journal of the ACM, 14(2):389–418, April 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ginsburg:1967:SAC

[GGH67b] Seymour Ginsburg, Sheila A. Greibach, and Michael A. Harrison. Stack automata and compiling. Journal of the ACM, 14(1):172– 201, January 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goldreich:1986:HCR

[GGM86] Oded Goldreich, Shafi Goldwasser, and Silvio Micali. How to construct random functions. Journal of the ACM, 33(4):792– 807, October 1986. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/6503.html. A computational complex-

ity measure of the randomness of functions is introduced, and, assuming the existence of one-way functions, a pseudo-random function generator is presented.

Goodman:1988:SBE

[GGMM88] Jonathan Goodman, Albert G. Greenberg, Neal Madras, and Peter March. Stability of binary exponential backoff. Journal of the ACM, 35(3):579–602, July 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/44488.html.

Goldreich:1998:PTC

[GGR98] Oded Goldreich, Shari Goldwasser, and Dana Ron. Property testing and its connection to learning and approximation. Journal of the ACM, 45(4):653–750, July 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:/

/www.acm.org:80/pubs/citations/journals/jacm/1998-45-
4/p653-goldreich/.

Gabow:1989:EIG

[GGS89] Harold N. Gabow, Zvi Galil, and Thomas H. Spencer. Efficient implementation of graph algorithms using contraction. Journal of the ACM, 36(3):540–572, July 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/65954.html.

Georges:2024:CPV

[GGV^{+}24] Aïna Linn Georges, Armaël Guéneau, Thomas Van Strydonck, Amin Timany, Alix Trieu, Dominique Devriese, and Lars Birkedal. Cerise: Program verification on a capability machine in the presence of untrusted code. Journal of the ACM, 71 (1):3:1–3:??, February 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3623510.

Goldstine:1959:PDN

[GH59] H. H. Goldstine and L. P. Horowitz. A procedure for the diagonalization of normal matrices. Journal of the ACM, 6(2):176– 195, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ginsburg:1964:SMM

[GH64] Seymour Ginsburg and Thomas N. Hibbard. Solvability of machine mappings of regular sets to regular sets. Journal of the ACM, 11(3):302–312, July 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ginsburg:1968:OWN

[GH68] Seymour Ginsburg and Michael A. Harrison. One-way nondeterministic real-time list-storage languages. Journal of the ACM, 15(3):428–446, July 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ginsburg:1970:TWB

[GH70] Seymour Ginsburg and John Hopcroft. Two-way balloon automata and AFL. Journal of the ACM, 17(1):3–13, January 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gray:1972:CRP

[GH72] James N. Gray and Michael A. Harrison. On the covering and reduction problems for context-free grammars. Journal of the ACM, 19(4):675–698, October 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gray:1973:CPS

[GH73] James N. Gray and Michael A. Harrison. Canonical precedence schemes. Journal of the ACM, 20(2):214–234, April 1973. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ginsburg:1986:SSR

[GH86] Seymour Ginsburg and Richard Hull. Sort sets in the relational model. Journal of the ACM, 33(3):465–488, July 1986. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
5929.html.

Geffner:2023:LBI

[GH23] Ivan Geffner and Joseph Y. Halpern. Lower bounds on implementing mediators in asynchronous systems with rational and malicious agents. Journal of the ACM, 70(2):13:1–13:??, April 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3578579.

Gelernter:1960:FCL

[GHG60] Herbert Gelernter, J. R. Hansen, and C. L. Gerberich. A Fortrancompiled list-processing language. Journal of the ACM, 7(2): 87–101, April 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goos:2024:SPC

[GHJ^{+}24] Mika Göös, Alexandros Hollender, Siddhartha Jain, Gilbert Maystre, William Pires, Robert Robere, and Ran Tao. Separations in proof complexity and TFNP. Journal of the ACM, 71(4):26:1–26:??, August 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3663758.

Gordon:2011:CFS

[GHKL11] S. Dov Gordon, Carmit Hazay, Jonathan Katz, and Yehuda Lindell. Complete fairness in secure Two-Party computation. Journal of the ACM, 58(6):24:1–24:37, December 2011. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gottlob:2021:SMS

[GHKL21] Georg Gottlob, André Hernich, Clemens Kupke, and Thomas Lukasiewicz. Stable model semantics for guarded existential rules and description logics: Decidability and complexity. Journal of the ACM, 68(5):35:1–35:87, October 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3447508.

Galil:1987:DVI

[GHL^{+}87] Zvi Galil, Christoph M. Hoffmann, Eugene M. Luks, Claus P. Schnorr, and Andreas Weber. An O(n^{3} log n) deterministic and an O(n^{3}) Las Vegas isomorphism test for trivalent graphs. Journal of the ACM, 34(3):513–531, July 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/28870.html.

Ghosh:2025:OFL

[GHL25] Arka Ghosh, Piotr Hofman, and Sławomir Lasota. Orbit-finite linear programming. Journal of the ACM, 72(1):1:1–1:??, February 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3703909.

Gupta:2022:OBC

[GHLL22] Anupam Gupta, David G. Harris, Euiwoong Lee, and Jason Li. Optimal bounds for the k-cut problem. Journal of the ACM, 69(1):2:1–2:18, February 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3478018.

Goodman:1975:AHC

[GHS75] S. E. Goodman, S. T. Hedetniemi, and P. J. Slater. Advances on the Hamiltonian completion problem. Journal of the ACM, 22(3):352–360, July 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Grohe:2009:LBP

[GHS09] Martin Grohe, André Hernich, and Nicole Schweikardt. Lower bounds for processing data with few random accesses to external memory. Journal of the ACM, 56(3):12:1–12:58, May 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goos:2013:LBL

[GHS13] Mika Göös, Juho Hirvonen, and Jukka Suomela. Lower bounds for local approximation. Journal of the ACM, 60(5):39:1–39:23, October 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gurari:1979:NCN

[GI79] Eitan M. Gurari and Oscar H. Ibarra. An NP-complete numbertheoretic problem. Journal of the ACM, 26(3):567–581, July 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gurari:1981:CEP

[GI81] Eitan M. Gurari and Oscar H. Ibarra. The complexity of the equivalence problem for simple programs. Journal of the ACM, 28(3):535–560, July 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gurari:1982:TWC

[GI82] Eitan M. Gurari and Oscar H. Ibarra. Two-way counter machines and Diophantine equations. Journal of the ACM, 29(3):863–873, July 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Giammo:1964:MMA

[Gia64] T. Giammo. A mathematical model for the automatic scaling of a function. Journal of the ACM, 11(1):79–83, January 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gibbs:1969:CGA

[Gib69] Norman E. Gibbs. A cycle generation algorithm for finite undirected linear graphs. Journal of the ACM, 16(4):564–568, October 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gill:1960:ANN

[Gil60] Arthur Gill. Analysis of nets by numerical methods. Journal of the ACM, 7(3):251–254, July 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gill:1963:WDP

[Gil63] Arthur Gill. On a weight distribution problem, with application to the design of stochastic generators. Journal of the ACM, 10 (1):110–122, January 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gill:1965:ASS

[Gil65] Arthur Gill. Analysis and synthesis of stable linear sequential circuits. Journal of the ACM, 12(1):141–149, January 1965. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gilbert:1966:SAL

[Gil66a] Philip Gilbert. On the syntax of algorithmic languages. Journal of the ACM, 13(1):90–107, January 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gill:1966:RIO

[Gil66b] A. Gill. Realization of input-output relations by sequential machines. Journal of the ACM, 13(1):33–42, January 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gilmore:1968:SPA

[Gil68] P. A. Gilmore. Structuring of parallel algorithms. Journal of the ACM, 15(2):176–192, April 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ginsburg:1958:LSU

[Gin58] Seymour Ginsburg. On the length of the smallest uniform experiment which distinguishes the terminal states of a machine. Journal of the ACM, 5(3):266–280, July 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ginsburg:1959:RSS

[Gin59] Seymour Ginsburg. On the reduction of superfluous states in a sequential machine. Journal of the ACM, 6(2):259–282, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ginsburg:1960:CPP

[Gin60] Seymour Ginsburg. Connective properties preserved in minimal state machines. Journal of the ACM, 7(4):311–325, October 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ginsburg:1961:CSI

[Gin61a] Seymour Ginsburg. Compatibility of states in input-independent machines. Journal of the ACM, 8(3):400–403, July 1961. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ginsburg:1961:STA

[Gin61b] Seymour Ginsburg. Sets of tapes accepted by different types of automata. Journal of the ACM, 8(1):81–86, January 1961. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ginzburg:1967:PCE

[Gin67] A. Ginzburg. A procedure for checking equality of regular expressions. Journal of the ACM, 14(2):355–362, April 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Galil:1999:FDP

[GIS99] Zvi Galil, Giuseppe F. Italiano, and Neil Sarnak. Fully dynamic planarity testing with applications. Journal of the ACM, 46 (1):28–91, January 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org:80/

pubs/citations/journals/jacm/1999-46-1/p28-galil/.

Givens:1957:CVV

[Giv57] Wallace Givens. The characteristic value-vector problem. Journal of the ACM, 4(3):298–307, July 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Givens:1958:CMC

[Giv58] Wallace Givens. Conference on matrix computations. Journal of the ACM, 5(1):100–115, January 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Garey:1976:CNO

[GJ76a] M. R. Garey and D. S. Johnson. The complexity of near-optimal graph coloring. Journal of the ACM, 23(1):43–49, January 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Garey:1976:STN

[GJ76b] M. R. Garey and D. S. Johnson. Scheduling tasks with nonuniform deadlines on two processors. Journal of the ACM, 23(3): 461–467, July 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Garey:1978:SNC

[GJ78] M. R. Garey and D. S. Johnson. “strong” NP-completeness results: Motivation, examples, and implications. Journal of the ACM, 25(3):499–508, July 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gonzalez:1980:NAP

[GJ80] Teofilo F. Gonzalez and Donald B. Johnson. A new algorithm for preemptive scheduling of trees. Journal of the ACM, 27(2):287– 312, April 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Grant:1982:FGD

[GJ82] John Grant and Barry E. Jacobs. On the family of generalized dependency constraints. Journal of the ACM, 29(4):986–997, October 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goldberg:2012:APF

[GJ12] Leslie Ann Goldberg and Mark Jerrum. Approximating the partition function of the ferromagnetic Potts model. Journal of the ACM, 59(5):25:1–25:31, October 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Guo:2019:UST

[GJL19] Heng Guo, Mark Jerrum, and Jingcheng Liu. Uniform sampling through the Lovász local lemma. Journal of the ACM, 66 (3):18:1–18:??, June 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3310131.

Ganardi:2021:BSL

[GJL21] Moses Ganardi, Artur Jez, and Markus Lohrey. Balancing straight-line programs. Journal of the ACM, 68(4):27:1–27:40, August 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3457389.

Goldreich:2012:TGO

[GJS12] Oded Goldreich, Brendan Juba, and Madhu Sudan. A theory of goal-oriented communication. Journal of the ACM, 59(2):8:1– 8:65, April 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goodman:1958:COL

[GK58] N. R. Goodman and S. Katz. Calculating open loop transfer functions from closed loop measurements. Journal of the ACM, 5(3):289–297, July 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Greenwald:1959:SSP

[GK59] Irwin D. Greenwald and Maureen Kane. The share 709 system: Programming and modification. Journal of the ACM, 6(2):128– 133, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gotlieb:1968:SCI

[GK68] C. C. Gotlieb and S. Kumar. Semantic clustering of index terms. Journal of the ACM, 15(4):493–513, October 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gottlieb:1984:CRP

[GK84] Allan Gottlieb and Clyde P. Kruskal. Complexity results for permuting data and other computations on parallel processors. Journal of the ACM, 31(2):193–209, April 1984. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Grigoriadis:1988:NCH

[GK88] M. D. Grigoriadis and B. Kalantari. A new class of heuristic algorithms for weighted perfect matching. Journal of the ACM, 35(4):769–776, October 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/48015.html.

Gao:1994:CRM

[GK94] Shaodi Gao and Michael Kaufmann. Channel routing of multiterminal nets. Journal of the ACM, 41(4):791–818, July 1994. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
179927.html.

Goodrich:1996:SPP

[GK96] Michael T. Goodrich and S. Rao Kosaraju. Sorting on a parallel pointer machine with applications to set expression evaluation. Journal of the ACM, 43(2):331–361, March 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
jacm/226670.html.

Goldwasser:1999:PTU

[GK99] Shafi Goldwasser and Joe Kilian. Primality testing using elliptic curves. Journal of the ACM, 46(4):450–472, July 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL http://www.acm.org/pubs/articles/journals/
jacm/1999-46-4/p450-goldwasser/p450-goldwasser.pdf;
http://www.acm.org/pubs/citations/journals/jacm/1999-
46-4/p450-goldwasser/.

Gottlob:2004:MDE

[GK04] Georg Gottlob and Christoph Koch. Monadic datalog and the expressive power of languages for Web information extraction. Journal of the ACM, 51(1):74–113, January 2004. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gupta:2014:ACD

[GKKS14] Ankit Gupta, Pritish Kamath, Neeraj Kayal, and Ramprasad Saptharishi. Approaching the chasm at depth four. Journal of the ACM, 61(6):33:1–33:??, November 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Garay:2024:BBP

[GKL24] Juan Garay, Aggelos Kiayias, and Nikos Leonardos. The bitcoin backbone protocol: Analysis and applications. Journal of the ACM, 71(4):25:1–25:??, August 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/doi/10.1145/3653445.

Grohe:2022:GDC

[GKpKL22] Martin Grohe, Benjamin Lucien Kaminski, Joost pieter Katoen, and Peter Lindner. Generative Datalog with continuous distributions. Journal of the ACM, 69(6):46:1–46:??, December 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3559102.

Gupta:2007:ACS

[GKPR07] Anupam Gupta, Amit Kumar, Martin P´ al, and Tim Roughgarden. Approximation via cost sharing: Simpler and better approximation algorithms for network design. Journal of the ACM, 54(3):11:1–11:38, June 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gottlob:2005:CXQ

[GKPS05] Georg Gottlob, Christoph Koch, Reinhard Pichler, and Luc Segoufin. The complexity of XPath query evaluation and XML typing. Journal of the ACM, 52(2):284–335, March 2005. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gandhi:2006:DRA

[GKPS06] Rajiv Gandhi, Samir Khuller, Srinivasan Parthasarathy, and Aravind Srinivasan. Dependent rounding and its applications to approximation algorithms. Journal of the ACM, 53(3):324–360, May 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gorman:1955:ACI

[GKR55] T. P. Gorman, R. G. Kelly, and R. B. Reddy. Automatic coding for the IBM 701. Journal of the ACM, 2(4):253–261, October 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goldwasser:2015:DCI

[GKR15] Shafi Goldwasser, Yael Tauman Kalai, and Guy N. Rothblum. Delegating computation: Interactive proofs for muggles. Journal of the ACM, 62(4):27:1–27:??, August 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ganor:2016:ESI

[GKR16] Anat Ganor, Gillat Kol, and Ran Raz. Exponential separation of information and communication for Boolean functions. Journal of the ACM, 63(5):46:1–46:??, December 2016. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gottlob:2004:ESO

[GKS04] Georg Gottlob, Phokion G. Kolaitis, and Thomas Schwentick. Existential second-order logic over graphs: Charting the tractability frontier. Journal of the ACM, 51(2):312–362, March 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gottlob:2006:CQT

[GKS06] Georg Gottlob, Christoph Koch, and Klaus U. Schulz. Conjunctive queries over trees. Journal of the ACM, 53(2):238–272, March 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Grohe:2017:DFO

[GKS17] Martin Grohe, Stephan Kreutzer, and Sebastian Siebertz. Deciding first-order properties of nowhere dense graphs. Journal of the ACM, 64(3):17:1–17:??, June 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gershinsky:1964:AHI

[GL64] Morris Gershinsky and David A. Levine. Aitken-Hermite interpolation. Journal of the ACM, 11(3):352–356, July 1964. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Good:1970:CIA

[GL70] Donald I. Good and Ralph L. London. Computer interval arithmetic: Definition and proof of correct implementation. Journal of the ACM, 17(4):603–612, October 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gaver:1971:PMB

[GL71] Donald P. Gaver, Jr. and Peter A. W. Lewis. Probability models for buffer storage allocation problems. Journal of the ACM, 18 (1):186–198, January 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ginsburg:1976:SCC

[GL76] Seymour Ginsburg and Nancy Lynch. Size complexity in contextfree grammar forms. Journal of the ACM, 23(4):582–598, October 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gottlob:1985:ESA

[GL85] G. Gottlob and A. Leitsch. On the efficiency of subsumption algorithms. Journal of the ACM, 32(2):280–295, April 1985. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
214118.html.

Gonnet:1988:EHL

[GL88a] Gaston H. Gonnet and Per-Åke Larson. External hashing with limited internal storage. Journal of the ACM, 35(1):161–184, January 1988. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/42274.html.

Gonzalez:1988:LTA

[GL88b] Teofilo F. Gonzalez and Sing-Ling Lee. A linear time algorithm for optimal routing around a rectangle. Journal of the ACM, 35(4):810–831, October 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/48018.html.

Grohe:2022:IIP

[GL22] Martin Grohe and Peter Lindner. Independence in infinite probabilistic databases. Journal of the ACM, 69(5):37:1–37:??, October 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3549525.

Glantz:1956:NM

[Gla56] Herbert T. Glantz. A note on microprogramming. Journal of the ACM, 3(2):77–84, April 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Glantz:1957:RID

[Gla57] Herbert T. Glantz. On the recognition of information with a digital computer. Journal of the ACM, 4(2):178–188, April 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Glauz:1965:NSO

[Gla65] R. D. Glauz. On the numerical solution of ordinary and partial differential equations using integral relations. Journal of the ACM, 12(2):175–180, April 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Glasserman:1991:SCP

[Gla91] Paul Glasserman. Structural conditions for perturbation analysis of queuing systems. Journal of the ACM, 38(4):1005–1025, October 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/115348.html.

Goldstein:1957:BLA

[GLH57] Allen A. Goldstein, Norman Levine, and James B. Hereshoff. On the “best” and “least Qth” approximation of an overdetermined system of linear equations. Journal of the ACM, 4(3):341–347, July 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Glicksman:1965:CME

[Gli65] Stephen Glicksman. Concerning the merging of equal length tape files. Journal of the ACM, 12(2):254–258, April 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goubault-Larrecq:2023:DTA

[GLJT23] Jean Goubault-Larrecq, Xiaodong Jia, and Clément Théron. A domain-theoretic approach to statistical programming languages. Journal of the ACM, 70(5):35:1–35:??, October 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3611660.

Grohe:2025:CCG

[GLNS25] Martin Grohe, Moritz Lichter, Daniel Neuen, and Pascal Schweitzer. Compressing CFI graphs and lower bounds for the Weisfeiler–Leman refinements. Journal of the ACM, 72(3):21:1– 21:??, June 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gottlob:2021:CAG

[GLPR21] Georg Gottlob, Matthias Lanzinger, Reinhard Pichler, and Igor Razgon. Complexity analysis of generalized and fractional hypertree decompositions. Journal of the ACM, 68(5):38:1–38:50, October 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3457374.

Guruswami:2025:PIH

[GLR^{+}25] Venkatesan Guruswami, Bingkai Lin, Xuandi Ren, Yican Sun, and Kewen Wu. Parameterized inapproximability hypothesis under ETH. Journal of the ACM, 72(5):32:1–32:40, October 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gottlob:2001:CAC

[GLS01] Georg Gottlob, Nicola Leone, and Francesco Scarcello. The complexity of acyclic conjunctive queries. Journal of the ACM, 48 (3):431–498, May 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gluss:1963:MOS

[Glu63] Brian Gluss. A method for obtaining suboptimal group-testing policies using dynamic programming and information theory. Journal of the ACM, 10:89–96, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gottlob:2012:STB

[GLVV12] Georg Gottlob, Stephanie Tien Lee, Gregory Valiant, and Paul Valiant. Size and treewidth bounds for conjunctive queries. Journal of the ACM, 59(3):16:1–16:35, June 2012. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gustavson:1970:SGO

[GLW70] F. G. Gustavson, W. Liniger, and R. Willoughby. Symbolic generation of an optimal Crout algorithm for sparse systems of linear equations. Journal of the ACM, 17(1):87–109, January 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Greenwald:1956:CAU

[GM56] I. D. Greenwald and H. G. Martin. Conclusions after using the PACT I advanced coding technique. Journal of the ACM, 3(4): 309–313, October 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gurk:1961:DSI

[GM61] Herbert M. Gurk and Jack Minker. The design and simulation of an information processing system. Journal of the ACM, 8 (2):260–270, April 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gurk:1970:SRI

[GM70] Herbert M. Gurk and Jack Minker. Storage requirements for information handling centers. Journal of the ACM, 17(1):65– 77, January 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gotlieb:1973:PMH

[GM73] C. C. Gotlieb and G. H. MacEwen. Performance of movablehead disk storage devices. Journal of the ACM, 20(4):604–623, October 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [GM75].

Gotlieb:1975:EPM

[GM75] C. C. Gotlieb and G. H. MacEwen. Errata: “Performance of movable-head disk storage devices”. Journal of the ACM, 22(2): 172, April 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [GM73].

Galil:1979:FSA

[GM79] Zvi Galil and Nimrod Megiddo. A fast selection algorithm and the problem of optimum distribution of effort. Journal of the ACM, 26(1):58–64, January 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ghezzi:1980:APS

[GM80] Carlo Ghezzi and Dino Mandrioli. Augmenting parsers to support incrementality. Journal of the ACM, 27(3):564–579, July 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Greenberg:1992:HFF

[GM92] Albert G. Greenberg and Neal Madras. How fair is fair queuing? Journal of the ACM, 39(3):568–598, July 1992. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
146658.html.

Goren:2020:S

[GM20] Guy Goren and Yoram Moses. Silence. Journal of the ACM, 67(1):3:1–3:26, February 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3377883.

Garcia-Molina:1985:HAV

[GMB85] Hector Garcia-Molina and Daniel Barbara. How to assign votes in a distributed system. Journal of the ACM, 32(4):841–860, October 1985. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/4223.html.

Goel:2015:PCA

[GML15] Gagan Goel, Vahab Mirrokni, and Renato Paes Leme. Polyhedral clinching auctions and the AdWords polytope. Journal of the ACM, 62(3):18:1–18:??, June 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gnesi:1981:DPG

[GMM81] Stefania Gnesi, Ugo Montanari, and Alberto Martelli. Dynamic programming as graph searching: an algebraic approach. Journal of the ACM, 28(4):737–751, October 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goldberg:2000:CRC

[GMPS00] Leslie Ann Goldberg, Philip D. Mackenzie, Mike Paterson, and Aravind Srinivasan. Contention resolution with constant expected delay. Journal of the ACM, 47(6):1048–1096, 2000. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/citations/journals/jacm/
2000-47-6/p1048-goldberg/.

Gal:2017:WFR

[GMPZ17] Ya’akov (Kobi) Gal, Moshe Mash, Ariel D. Procaccia, and Yair Zick. Which is the fairest (rent division) of them all? Journal of the ACM, 64(6):39:1–39:??, November 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gottlob:2009:GHD

[GMS09] Georg Gottlob, Zoltán Miklós, and Thomas Schwentick. Generalized hypertree decompositions: NP-hardness and tractable variants. Journal of the ACM, 56(6):30:1–30:32, September 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Guha:2010:AAR

[GMS10] Sudipto Guha, Kamesh Munagala, and Peng Shi. Approximation algorithms for restless bandit problems. Journal of the ACM, 58 (1):3:1–3:50, December 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gaifman:1993:UOP

[GMSV93] Haim Gaifman, Harry Mairson, Yehoshua Sagiv, and Moshe Y. Vardi. Undecidable optimization problems for database logic programs. Journal of the ACM, 40(3):683–713, July 1993. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
174142.html.

Graham:1986:NDS

[GMV86] Marc H. Graham, Alberto O. Mendelzon, and Moshe Y. Vardi. Notions of dependency satisfaction. Journal of the ACM, 33(1): 105–129, January 1986. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/4798.html. Also published

in/as: Stanford Un., 1984.

Goldstine:1959:JMR

[GMvN59] H. H. Goldstine, F. J. Murray, and J. von Neumann. The Jacobi method for real symmetric matrices. Journal of the ACM, 6 (1):59–96, January 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goldreich:1991:PYN

[GMW91] Oded Goldreich, Silvio Micali, and Avi Wigderson. Proofs that yield nothing but their validity or all languages in NP have zero-knowledge proof systems. Journal of the ACM, 38(3):691– 729, July 1991. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/116852.html. They show that for a lan-

guage L in NP and a string w in L, there exists a probabilistic interactive proof that efficiently demonstrates membership of x in L without conveying additional information. Previously, zeroknowledge proofs were known only for some problems that were in both NP and co-NP. A preliminary version of this paper appeared in Proc. 27th Ann. IEEE Symp. on Foundations of Computer Science, 1986, under the title “Proofs that yield nothing but their validity and a methodology of cryptographic protocol design.”.

Glass:1994:TMA

[GN94] Christopher J. Glass and Lionel M. Ni. The turn model for adaptive routing. Journal of the ACM, 41(5):874–902, September 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/185682.html.

Gottlob:2008:ECC

[GN08] Georg Gottlob and Alan Nash. Efficient core computation in data exchange. Journal of the ACM, 55(2):9:1–9:49, May 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gallier:1993:AFC

[GNP^{+}93] Jean Gallier, Paliath Narendran, David Plaisted, Stan Raatz, and Wayne Snyder. An algorithm for finding canonical sets of ground rewrite rules in polynomial time. Journal of the ACM, 40(1):1–16, January 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/138032.html.

Gagie:2020:FFS

[GNP20] Travis Gagie, Gonzalo Navarro, and Nicola Prezza. Fully functional suffix trees and optimal text searching in BWT-runs bounded space. Journal of the ACM, 67(1):2:1–2:54, February 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3375890.

Gallier:1992:TPU

[GNRS92] Jean Gallier, Paliath Narendran, Stan Raatz, and Wayne Snyder. Theorem proving using equational matings and rigid E– unification. Journal of the ACM, 39(2):377–429, April 1992. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
128754.html.

Goldreich:1996:SPS

[GO96] Oded Goldreich and Rafail Ostrovsky. Software protection and simulation on oblivious RAMs. Journal of the ACM, 43(3):431– 473, May 1996. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/jacm/233553.html.

Guard:1969:SAM

[GOBS69] J. R. Guard, F. C. Oglesby, J. H. Bennett, and L. G. Settle. Semi-automated mathematics. Journal of the ACM, 16(1):49– 62, January 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goetz:1961:LEI

[Goe61] Martin Goetz. Letter to the editor: “Internal sorting and external merging”. Journal of the ACM, 8(4):649–650, October 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Nag59, Liv60, Flo61a, Nag61].

Goffman:1971:MMA

[Gof71] William Goffman. A mathematical method for analyzing the growth of a scientific discipline. Journal of the ACM, 18(1):173– 185, January 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goldstein:1957:DCNa

[Gol57a] Gordon D. Goldstein. Digital Computer Newsletter. Journal of the ACM, 4(1):97–120, January 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goldstein:1957:DCNb

[Gol57b] Gordon D. Goldstein. Digital Computer Newsletter. Journal of the ACM, 4(2):225–244, April 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goldstein:1957:DCNc

[Gol57c] Gordon D. Goldstein. Digital Computer Newsletter. Journal of the ACM, 4(3):371–391, July 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goldstein:1957:DCNd

[Gol57d] Gordon D. Goldstein. Digital Computer Newsletter. Journal of the ACM, 4(4):541–558, October 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goldstine:1960:FRP

[Gol60] H. H. Goldstine. Footnote to a recent paper. Journal of the ACM, 7(1):78–79, January 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goldschlager:1982:UIP

[Gol82] Leslie M. Goldschlager. A universal interconnection pattern for parallel computers. Journal of the ACM, 29(4):1073–1086, October 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gonnet:1981:ELL

[Gon81] Gaston H. Gonnet. Expected length of the longest probe sequence in hash code searching. Journal of the ACM, 28(2):289– 304, April 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goodrich:2011:RSS

[Goo11] Michael T. Goodrich. Randomized Shellsort: a simple dataoblivious sorting algorithm. Journal of the ACM, 58(6):27:1– 27:26, December 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gordon:1956:OPI

[Gor56] Barry Gordon. An optimizing program for the IBM 650. Journal of the ACM, 3(1):3–5, January 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gorn:1957:SPM

[Gor57] Saul Gorn. Standardized programming methods and universal coding. Journal of the ACM, 4(3):254–273, July 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gorn:1963:DGA

[Gor63] Saul Gorn. Detection of generative ambiguities in context-free mechanical languages. Journal of the ACM, 10:196–208, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Groth:2012:NTN

[GOS12] Jens Groth, Rafail Ostrovsky, and Amit Sahai. New techniques for noninteractive zero-knowledge. Journal of the ACM, 59(3): 11:1–11:35, June 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goyal:2013:VCT

[GOS13] Navin Goyal, Neil Olver, and F. Bruce Shepherd. The VPN conjecture is true. Journal of the ACM, 60(3):17:1–17:17, June 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gotlieb:1954:RCE

[Got54] C. C. Gotlieb. Running a computer efficiently. Journal of the ACM, 1(3):124–127, July 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gottlob:1995:NTC

[Got95a] Georg Gottlob. NP trees and Carnap’s modal logic. Journal of the ACM, 42(2):421–457, March 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/201031.html.

Gottlob:1995:TDL

[Got95b] Georg Gottlob. Translating default logic into standard autoepistemic logic. Journal of the ACM, 42(4):711–740, July 1995. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
210334.html.

Gotlieb:2003:GAJ

[Got03] C. C. Gotlieb. A Golden Age for JACM. Journal of the ACM, 50(1):14, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Galil:1983:EGP

[GP83] Zvi Galil and Wolfgang J. Pauli. An efficient general-purpose parallel computer. Journal of the ACM, 30(2):360–387, April 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Grochow:2018:CCP

[GP18a] Joshua A. Grochow and Toniann Pitassi. Circuit complexity, proof complexity, and polynomial identity testing: The ideal proof system. Journal of the ACM, 65(6):37:1–37:??, November 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gronlund:2018:TDL

[GP18b] Allan Grønlund and Seth Pettie. Threesomes, degenerates, and love triangles. Journal of the ACM, 65(4):22:1–22:??, August 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goaoc:2019:SNC

[GPP^{+}19] Xavier Goaoc, Pavel Paták, Zuzana Patáková, Martin Tancer, and Uli Wagner. Shellability is NP-complete. Journal of the ACM, 66(3):21:1–21:??, June 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=3314024.

Gyssens:1989:UAT

[GPV89] Marc Gyssens, Jan Paredaens, and Dirk Van Gucht. A uniform approach toward handling atomic and structured information in the nested relational database model. Journal of the ACM, 36(4):790–825, October 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/76364.html.

Galler:1960:GTC

[GR60] B. A. Galler and D. P. Rozenburg. A generalization of a theorem of Carr on error bounds for Runge–Kutta procedures. Journal of the ACM, 7(1):57–60, January 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ginsburg:1962:TFL

[GR62] Seymour Ginsburg and H. Gordon Rice. Two families of languages related to ALGOL. Journal of the ACM, 9(3):350–371, July 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ginsburg:1963:OWP

[GR63a] Seymour Ginsburg and G. F. Rose. Operations which preserve definability in languages. Journal of the ACM, 10:175–195, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ginsburg:1963:SRU

[GR63b] Seymour Ginsburg and Gene F. Rose. Some recursively unsolvable problems in ALGOL-like languages. Journal of the ACM, 10(1):29–47, January 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goldstein:1973:MP

[GR73] A. J. Goldstein and P. L. Richman. A midpoint phenomenon. Journal of the ACM, 20(2):301–304, April 1973. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gordon:1974:BBM

[GR74] William J. Gordon and Richard F. Riesenfeld. Bernstein-Bézier methods for the computer-aided design of free form curves and surfaces. Journal of the ACM, 21(2):293–310, April 1974. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gross:1979:LTP

[GR79] Jonathan L. Gross and Ronald H. Rosen. A linear time planarity algorithm for 2-complexes. Journal of the ACM, 26(4):611–617, October 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goldberg:1998:BFD

[GR98] Andrew V. Goldberg and Satish Rao. Beyond the flow decomposition barrier. Journal of the ACM, 45(5):783–797, September 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org:80/pubs/citations/
journals/jacm/1998-45-5/p1-goldberg/.

Goldreich:2017:LTD

[GR17] Oded Goldreich and Dana Ron. On learning and testing dynamic environments. Journal of the ACM, 64(3):21:1–21:??, June 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goemans:2020:PPC

[GR20] Michel X. Goemans and Thomas Rothvoss. Polynomiality for bin packing with a constant number of item types. Journal of the ACM, 67(6):38:1–38:21, November 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3421750.

Graney:1955:MAT

[Gra55a] Edward P. Graney. Maintenance and acceptance tests used on the MIDAC. Journal of the ACM, 2(2):95–98, April 1955. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gray:1955:PTE

[Gra55b] H. J. Gray, Jr. Propagation of truncation errors in the numerical solution of ordinary differential equations by repeated closures. Journal of the ACM, 2(1):6–17, January 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Grau:1962:TOS

[Gra62] A. A. Grau. A translator-oriented symbolic programming language. Journal of the ACM, 9(4):480–487, October 1962. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Grau:1963:RNR

[Gra63] A. A. Grau. On the reduction of number range in the use of the Graeffe process. Journal of the ACM, 10:538–544, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Graham:1977:PTA

[Gra77] Susan L. Graham. Papers from Third ACM Symposium on Principles of Programming Languages. Journal of the ACM, 24(1):96–97, January 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gray:2003:WND

[Gra03] Jim Gray. What next?: a dozen information-technology research goals. Journal of the ACM, 50(1):41–57, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Greenberger:1961:NNP

[Gre61] Martin Greenberger. Notes on a new pseudo-random number generator. Journal of the ACM, 8(2):163–167, April 1961. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Greibach:1965:NNF

[Gre65] Sheila A. Greibach. A new normal-form theorem for context-free phrase structure grammars. Journal of the ACM, 12(1):42–52, January 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Greibach:1966:URL

[Gre66] Sheila A. Greibach. The unsolvability of the recognition of linear context-free languages. Journal of the ACM, 13(4):582–587, October 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Greibach:1969:IHC

[Gre69] Sheila A. Greibach. An infinite hierarchy of context-free languages. Journal of the ACM, 16(1):91–106, January 1969. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Green:1973:PEF

[Gre73] Christopher D. Green. A path entropy function for rooted trees. Journal of the ACM, 20(3):378–384, July 1973. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Griffiths:1968:UEP

[Gri68] T. V. Griffiths. The unsolvability of the equivalence problem for Λ-free nondeterministic generalized machines. Journal of the ACM, 15(3):409–413, July 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Griffith:1973:MMA

[Gri73] Arnold K. Griffith. Mathematical models for automatic line detection. Journal of the ACM, 20(1):62–80, January 1973. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Grimson:1986:CLC

[Gri86] W. Eric L. Grimson. The combinatorics of local constraints in model-based recognition and localization from sparse data. Journal of the ACM, 33(4):658–686, October 1986. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
6492.html.

Grohe:2007:CHC

[Gro07] Martin Grohe. The complexity of homomorphism and constraint satisfaction problems seen from the other side. Journal of the ACM, 54(1):1:1–1:24, March 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Grohe:2012:FPD

[Gro12] Martin Grohe. Fixed-point definability and polynomial time on graphs with excluded minors. Journal of the ACM, 59(5): 27:1–27:64, October 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Giacobazzi:2000:MAI

[GRS00] Roberto Giacobazzi, Francesco Ranzato, and Francesca Scozzari. Making abstract interpretations complete. Journal of the ACM, 47(2):361–416, March 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:

//www.acm.org/pubs/citations/journals/jacm/2000-47-2/
p361-giacobazzi/.

Gruttke:1970:PRK

[Gru70] William B. Gruttke. Pseudo-Runge–Kutta methods of the fifth order. Journal of the ACM, 17(4):613–628, October 1970. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gerberich:1957:CCM

[GS57] C. L. Gerberich and W. C. Sangren. Codes for the classical membrane problem. Journal of the ACM, 4(4):477–486, October 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ginsburg:1963:QCF

[GS63] Seymour Ginsburg and Edwin H. Spanier. Quotients of contextfree languages. Journal of the ACM, 10:487–492, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gragg:1964:GMP

[GS64] William B. Gragg and Hans J. Stetter. Generalized multistep predictor-corrector methods. Journal of the ACM, 11(2):188– 209, April 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ginsburg:1965:MLT

[GS65] Seymour Ginsburg and Edwin H. Spanier. Mappings of languages by two-tape devices. Journal of the ACM, 12(3):423–434, July 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ghosh:1969:FOS

[GS69] S. P. Ghosh and M. E. Senko. File organization: On the selection of random access index points for sequential files. Journal of the ACM, 16(4):569–579, October 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Grossman:1973:PRS

[GS73] David D. Grossman and Harvey F. Silverman. Placement of records on a secondary storage device to minimize access time. Journal of the ACM, 20(3):429–438, July 1973. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gonzalez:1976:OSS

[GS76] Teofilo Gonzalez and Sartaj Sahni. Open shop scheduling to minimize finish time. Journal of the ACM, 23(4):665–679, October 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Galil:1978:LTL

[GS78a] Zvi Galil and Joel Seiferas. A linear-time on-line recognition algorithm for “palstar”. Journal of the ACM, 25(1):102–111, January 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gonzalez:1978:PSU

[GS78b] Teofilo Gonzalez and Sartaj Sahni. Preemptive scheduling of uniform processor systems. Journal of the ACM, 25(1):92–101, January 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Goodman:1983:SCT

[GS83] Nathan Goodman and Oded Shmueli. Syntactic characterization of tree database schemas. Journal of the ACM, 30(4):767–786, October 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

GonzalezSmith:1985:PAD

[GS85] M. E. Gonzalez Smith and J. A. Storer. Parallel algorithms for data compression. Journal of the ACM, 32(2):344–373, April 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/3152.html.

Gurevich:1990:NLT

[GS90] Yuri Gurevich and Saharon Shelah. Nondeterministic linear-time tasks may require substantially nonlinear deterministic time in the case of sublinear work space. Journal of the ACM, 37(3): 674–687, July 1990. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/214070.html.

Gasarch:1992:LQ

[GS92a] William I. Gasarch and Carl H. Smith. Learning via queries. Journal of the ACM, 39(3):649–674, July 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/

146670.html. A shorter version is in 29th FOCS conference, 1988, pp. 130-137.

German:1992:RAS

[GS92b] Steven M. German and A. Prasad Sistla. Reasoning about systems with many processes. Journal of the ACM, 39(3):675– 735, July 1992. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/146681.html.

Golumbic:1993:CAR

[GS93] Martin Charles Golumbic and Ron Shamir. Complexity and algorithms for reasoning about time: A graph-theoretic approach. Journal of the ACM, 40(5):1108–1133, November 1993. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
169675.html.

Goldreich:2006:LTC

[GS06] Oded Goldreich and Madhu Sudan. Locally testable codes and PCPs of almost-linear length. Journal of the ACM, 53(4):558– 655, July 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gilpin:2007:LAI

[GS07] Andrew Gilpin and Tuomas Sandholm. Lossless abstraction of imperfect information games. Journal of the ACM, 54(5): 25:1–25:30, October 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gottlob:2010:SMD

[GS10] Georg Gottlob and Pierre Senellart. Schema mapping discovery from data instances. Journal of the ACM, 57(2):6:1–6:37, January 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Guruswami:2016:ORC

[GS16] Venkatesan Guruswami and Adam Smith. Optimal rate code constructions for computationally simple channels. Journal of the ACM, 63(4):35:1–35:??, November 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gasieniec:2021:EPC

[GS21] Leszek Gasieniec and Grzegorz Stachowiak. Enhanced phase clocks, population protocols, and fast space optimal leader election. Journal of the ACM, 68(1):2:1–2:21, February 2021. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3424659.

Garg:2022:TRM

[GS22] Sanjam Garg and Akshayaram Srinivasan. Two-round multiparty secure computation from minimal assumptions. Journal of the ACM, 69(5):36:1–36:??, October 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3566048.

Gabor:1989:RCG

[GSH89] Csaba P. Gabor, Kenneth J. Supowit, and Wen-Lian Hsu. Recognizing circle graphs in polynomial time. Journal of the ACM, 36(3):435–473, July 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/65951.html.

Green:1959:ETA

[GSK59] Bert F. Green, Jr., J. E. Keith Smith, and Laura Klem. Empirical tests of an additive random number generator. Journal of the ACM, 6(4):527–537, October 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gebauer:2016:LLA

[GST16] Heidi Gebauer, Tibor Szabó, and Gábor Tardos. The local lemma is asymptotically tight for SAT. Journal of the ACM, 63 (5):43:1–43:??, December 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gurevich:1984:SNH

[GSV84] Yuri Gurevich, Larry Stockmeyer, and Uzi Vishkin. Solving NPhard problems on graphs that are almost trees and an application to facility location problems. Journal of the ACM, 31(3):459– 473, July 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Galanis:2015:IAS

[GSV15] Andreas Galanis, Daniel Stefankovic, and Eric Vigoda. Inapproximability for antiferromagnetic spin systems in the tree nonuniqueness region. Journal of the ACM, 62(6):50:1–50:??, December 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Galil:1988:MCF

[GT88a] Zvi Galil and Éva Tardos. An O(n^{2}(m + n log n) log n) mincost flow algorithm. Journal of the ACM, 35(2):374–386, April 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/

0004-5411/214090.html. Preliminary version in Proc. 27th Annual IEEE Symposium on the Foundations of Computer Science, pages 1–9, 1986.

Goldberg:1988:NAM

[GT88b] Andrew V. Goldberg and Robert E. Tarjan. A new approach to the maximum-flow problem. Journal of the ACM, 35(4):921– 940, October 1988. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/61051.html. Preliminary version in

Proc. 18th Annual ACM Symposium on the Theory of Computing, pages 136–146, 1986.

Goldberg:1989:FMC

[GT89] Andrew V. Goldberg and Robert E. Tarjan. Finding minimumcost circulations by canceling negative cycles. Journal of the ACM, 36(4):873–886, October 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/76368.html. Pre-

liminary version in Proc. 20th Annual ACM Symposium on the Theory of Computing, pages 388–397, 1987.

Gabow:1991:FSA

[GT91] Harold N. Gabow and Robert E. Tarjan. Faster scaling algorithms for general graph-matching problems. Journal of the ACM, 38(4):815–853, October 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/115366.html.

Gaitonde:2023:PAS

[GT23] Jason Gaitonde and Éva Tardos. The price of anarchy of strategic queuing systems. Journal of the ACM, 70(3):20:1–20:??, June 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3587250.

Goguen:1977:IAS

[GTWW77] J. A. Goguen, J. W. Thatcher, E. G. Wagner, and J. B. Wright. Initial algebra semantics and continuous algebras. Journal of the ACM, 24(1):68–95, January 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). An early version is “Initial Algebra Semantics”, with James Thatcher, IBM T. J. Watson Research Center, Report RC 4865, May 1974.

Ginsburg:1966:ACF

[GU66a] Seymour Ginsburg and Joseph Ullian. Ambiguity in context free languages. Journal of the ACM, 13(1):62–89, January 1966. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ginsburg:1966:PUI

[GU66b] Seymour Ginsburg and Joseph Ullian. Preservation of unambiguity and inherent ambiquity in context-free languages. Journal of the ACM, 13(3):364–368, July 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Guibas:1978:AHT

[Gui78] Leo J. Guibas. The analysis of hashing techniques that exhibit k-ary clustering. Journal of the ACM, 25(4):544–555, October 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Geerts:2016:ASC

[GUK^{+}16] Floris Geerts, Thomas Unger, Grigoris Karvounarakis, Irini Fundulaki, and Vassilis Christophides. Algebraic structures for capturing the provenance of SPARQL queries. Journal of the ACM, 63(1):7:1–7:??, March 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gupta:1979:BSC

[Gup79] Udaiprakash Gupta. Bounds on storage for consecutive retrieval. Journal of the ACM, 26(1):28–36, January 1979. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gurari:1985:DPP

[Gur85] Eitan M. Gurari. Decidable problems for powerful programs. Journal of the ACM, 32(2):466–483, April 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
3157.html.

Gusfield:1983:PCC

[Gus83] Dan Gusfield. Parametric combinatorial computing and a problem of program module distribution. Journal of the ACM, 30(3): 551–563, July 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Guruswami:2009:UER

[GUV09] Venkatesan Guruswami, Christopher Umans, and Salil Vadhan. Unbalanced expanders and randomness extractors from Parvaresh–Vardy codes. Journal of the ACM, 56(4):20:1–20:34, June 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gutenmakher:1961:PUI

[GV61] L. I. Gutenmakher and G. E. Vleduts. The prospects for the utilization of informational-logical machines in chemistry (USSR). Journal of the ACM, 8(2):240–251, April 1961. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Gorbunov:2015:ABE

[GVW15] Sergey Gorbunov, Vinod Vaikuntanathan, and Hoeteck Wee. Attribute-based encryption for circuits. Journal of the ACM, 62 (6):45:1–45:??, December 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Graham:1976:FUL

[GW76] Susan L. Graham and Mark Wegman. A fast and usually linear algorithm for global flow analysis. Journal of the ACM, 23(1): 172–202, January 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Greenberg:1985:LBT

[GW85] Albert G. Greenberg and Shmuel Winograd. A lower bound on the time needed in the worst case to resolve conflicts deterministically in multiple access channels. Journal of the ACM, 32(3): 589–596, July 1985. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/214125.html.

Graham:1990:EES

[GW90] Marc H. Graham and Ke Wang. On the equivalence of an Egd to a set of Fd’s. Journal of the ACM, 37(3):474–490, July 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/79151.html.

Goemans:1995:IAA

[GW95] Michel X. Goemans and David P. Williamson. Improved approximation algorithms for maximum cut and satisfiability problems using semidefinite programming. Journal of the ACM, 42(6): 1115–1145, November 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/227684.html.

Gonczarowski:2021:SCM

[GW21] Yannai A. Gonczarowski and S. Matthew Weinberg. The sample complexity of up-to-ϵ multi-dimensional revenue maximization. Journal of the ACM, 68(3):15:1–15:28, May 2021. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3439722.

Goaoc:2023:CHR

[GW23] Xavier Goaoc and Emo Welzl. Convex hulls of random order types. Journal of the ACM, 70(1):8:1–8:??, February 2023. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3570636.

Guruswami:2022:ORL

[GX22] Venkatesan Guruswami and Chaoping Xing. Optimal rate list decoding over bounded alphabets using algebraic-geometric codes. Journal of the ACM, 69(2):10:1–10:48, April 2022. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3506668.

Graham:1980:IBW

[GYY80] Ronald L. Graham, Andrew C. Yao, and F. Frances Yao. Information bounds are weak in the shortest distance problem. Journal of the ACM, 27(3):428–444, July 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ginsburg:1982:PFD

[GZ82] Seymour Ginsburg and Sami Mohammed Zaiddan. Properties of functional-dependency families. Journal of the ACM, 29(3): 678–698, July 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Heizer:1956:TFS

[HA56] L. E. Heizer and S. J. Abraham. Transfer function simulation by means of amplifiers and potentiometers. Journal of the ACM, 3(3):186–198, July 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Haber:1970:SNA

[Hab70] Seymour Haber. Sequences of numbers that are approximately completely equidistributed. Journal of the ACM, 17(2):269–272, April 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hadzilacos:1988:TRD

[Had88] Vassos Hadzilacos. A theory of reliability in database systems. Journal of the ACM, 35(1):121–145, January 1988. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
42272.html.

Haeupler:2015:SFD

[Hae15] Bernhard Haeupler. Simple, fast and deterministic gossip and rumor spreading. Journal of the ACM, 62(6):47:1–47:??, December 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Haeupler:2016:ANC

[Hae16] Bernhard Haeupler. Analyzing network coding (gossip) made easy. Journal of the ACM, 63(3):26:1–26:??, September 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hajek:1983:PFT

[Haj83] Bruce Hajek. The proof of a folk theorem on queuing delay with applications to routing in networks. Journal of the ACM, 30 (4):834–851, October 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Halpern:1997:BEC

[Hal97] Joseph Halpern. On becoming Editor-in-Chief of JACM. Journal of the ACM, 44(3):363–365, May 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Halpern:1998:TPP

[Hal98] Joseph Y. Halpern. Time to publication: a progress report. Journal of the ACM, 45(3):379–380, May 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:/

/www.acm.org:80/pubs/citations/journals/jacm/1998-45-
3/p379-halpern/.

Halpern:1999:CET

[Hal99a] J. Halpern. Corrigendum: Editorial: Taking stock. Journal of the ACM, 46(4):576, March 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Hal99b].

Halpern:1999:ETS

[Hal99b] J. Halpern. Editorial: Taking stock. Journal of the ACM, 46 (3):323–324, March 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See corrigendum [Hal99a].

Halpern:2000:EBR

[Hal00] Joe Halpern. Editorial: a bill of rights and responsibilities. Journal of the ACM, 47(5):823–825, 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/citations/journals/jacm/
2000-47-5/p823-halpern/.

Halpern:2002:UTP

[Hal02] Joe Halpern. Update: Time to publication statistics. Journal of the ACM, 49(6):715, November 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Halpern:2003:JA

[Hal03] Joseph Y. Halpern. JACM’s 50th anniversary. Journal of the ACM, 50(1):3–7, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hallgren:2007:PTQ

[Hal07] Sean Hallgren. Polynomial-time quantum algorithms for Pell’s equation and the principal ideal problem. Journal of the ACM, 54(1):4:1–4:19, March 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hammersley:1956:CMC

[Ham56a] J. M. Hammersley. Conditional Monte Carlo. Journal of the ACM, 3(2):73–76, April 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1956:BRa

[Ham56b] R. W. Hamming. Book reviews. Journal of the ACM, 3(2):112– 113, April 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1956:BRb

[Ham56c] R. W. Hamming. Book reviews. Journal of the ACM, 3(3):239, July 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1956:BRc

[Ham56d] R. W. Hamming. Book reviews. Journal of the ACM, 3(4):376– 378, October 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1957:BRa

[Ham57a] R. W. Hamming. Book reviews. Journal of the ACM, 4(1):93– 94, January 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1957:BRb

[Ham57b] R. W. Hamming. Book reviews. Journal of the ACM, 4(2):219– 220, April 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1957:BRc

[Ham57c] R. W. Hamming. Book reviews. Journal of the ACM, 4(3):362– 366, July 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1957:BRd

[Ham57d] R. W. Hamming. Book reviews. Journal of the ACM, 4(4):530– 533, October 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1958:BRa

[Ham58a] R. W. Hamming. Book reviews. Journal of the ACM, 5(1): 116, January 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1958:BRb

[Ham58b] R. W. Hamming. Book reviews. Journal of the ACM, 5(2):197– 203, April 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1958:BRc

[Ham58c] R. W. Hamming. Book reviews. Journal of the ACM, 5(3):298– 308, July 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1958:BRd

[Ham58d] R. W. Hamming. Book reviews. Journal of the ACM, 5(4):389– 396, October 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1959:BRa

[Ham59a] R. W. Hamming. Book reviews. Journal of the ACM, 6(1):115– 120, January 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1959:BRb

[Ham59b] R. W. Hamming. Book reviews. Journal of the ACM, 6(2):306– 312, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1959:BRc

[Ham59c] R. W. Hamming. Book reviews. Journal of the ACM, 6(3):443– 458, July 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1959:BRd

[Ham59d] R. W. Hamming. Book reviews. Journal of the ACM, 6(4):548– 556, October 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1959:SPC

[Ham59e] R. W. Hamming. Stable predictor-corrector methods for ordinary differential equations. Journal of the ACM, 6(1):37–47, January 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1960:BR

[Ham60] R. W. Hamming. Book reviews. Journal of the ACM, 7(1):80– 86, January 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1969:OMV

[Ham69] R. W. Hamming. One man’s view of computer science. Journal of the ACM, 16(1):3–12, January 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hansen:1962:QJM

[Han62] Eldon R. Hansen. On quasicyclic Jacobi methods. Journal of the ACM, 9(1):118–135, January 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hansen:1963:DM

[Han63] Eldon R. Hansen. On the Danilewski method. Journal of the ACM, 10:102–109, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Harlow:1957:HPI

[Har57] Francis H. Harlow. Hydrodynamic problems involving large fluid distortions. Journal of the ACM, 4(2):137–142, April 1957. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Harary:1960:CPM

[Har60] Frank Harary. On the consistency of precedence matrices. Journal of the ACM, 7(3):255–259, July 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Harrison:1963:NCI

[Har63a] Michael A. Harrison. The number of classes of invertible Boolean functions. Journal of the ACM, 10(1):25–28, January 1963. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hartmanis:1963:FRS

[Har63b] J. Hartmanis. Further results on the structure of sequential machines. Journal of the ACM, 10:78–88, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hartt:1964:SAP

[Har64] Kenneth Hartt. Some analytical procedures for computers and their applications to a class of multidimensional integrals. Journal of the ACM, 11(4):416–421, October 1964. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Harrison:1966:AES

[Har66] Michael A. Harrison. On asymptotic estimates in switching and automata theory. Journal of the ACM, 13(1):151–157, January 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hartmanis:1967:MRC

[Har67] J. Hartmanis. On memory requirements for context-free language recognition. Journal of the ACM, 14(4):663–665, October 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hartmanis:1968:CCO

[Har68] J. Hartmanis. Computational complexity of one-tape Turing machine computations. Journal of the ACM, 15(2):325–339, April 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hartmanis:1969:CUP

[Har69] J. Hartmanis. On the complexity of undecidable problems in automata theory. Journal of the ACM, 16(1):160–167, January 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Haralick:1974:DRD

[Har74] Robert M. Haralick. The Diclique representation and decomposition of binary relations. Journal of the ACM, 21(3):356–366, July 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Harris:1976:CTP

[Har76] Britton Harris. A code for the transportation problem of linear programming. Journal of the ACM, 23(1):155–157, January 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Harel:1986:ETI

[Har86] David Harel. Effective transformations on infinite trees, with applications to high undecidability, dominoes, and fairness. Journal of the ACM, 33(1):224–248, January 1986. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/4993.html.

Hartmanis:2003:SCC

[Har03] Juris Hartmanis. Separation of complexity classes. Journal of the ACM, 50(1):58–62, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hashimoto:1970:MSL

[Has70] Masahiro Hashimoto. A method for solving large matrix equations reduced from Fredholm integral equations of the second kind. Journal of the ACM, 17(4):629–636, October 1970. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Haastad:2001:SOI

[Hås01] Johan Håstad. Some optimal inapproximability results. Journal of the ACM, 48(4):798–859, July 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Haastad:2021:SDF

[Hås21] Johan Håstad. On small-depth Frege proofs for Tseitin for grids. Journal of the ACM, 68(1):1:1–1:31, February 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3425606.

Hayes:1975:FSS

[Hay75] John P. Hayes. The fanout structure of switching functions. Journal of the ACM, 22(4):551–571, October 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hayes:1976:EFF

[Hay76] John P. Hayes. Enumeration of fanout-free Boolean functions. Journal of the ACM, 23(4):700–709, October 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Harchol-Balter:2002:TAU

[HB02] Mor Harchol-Balter. Task assignment with unknown duration. Journal of the ACM, 49(2):260–288, March 2002. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Harrison:2021:RTD

[HB21] P. G. Harrison and J. Bor. Response time distribution in a tandem pair of queues with batch processing. Journal of the ACM, 68(4):22:1–22:41, August 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3448973.

Hull:1963:EPC

[HC63] T. E. Hull and A. L. Creemer. Efficiency of predictor-corrector procedures. Journal of the ACM, 10:291–301, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Huang:1972:IPM

[HC72] Nancy M. Huang and Randall E. Cline. Inversion of persymmetric matrices having Toeplitz inverses. Journal of the ACM, 19(3):437–444, July 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hickey:1988:APA

[HC88] Timothy Hickey and Jacques Cohen. Automating program analysis. Journal of the ACM, 35(1):185–220, January 1988. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
42275.html.

Hull:1964:MCR

[HD64] T. E. Hull and A. R. Dobell. Mixed congruential random number generators for binary machines. Journal of the ACM, 11(1):31– 40, January 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hwang:1973:CMA

[HD73] F. K. Hwang and D. N. Deutsch. A class of merging algorithms. Journal of the ACM, 20(1):148–159, January 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Holm:2001:PLD

[HdLT01] Jacob Holm, Kristian de Lichtenberg, and Mikkel Thorup. Polylogarithmic deterministic fully-dynamic algorithms for connectivity, minimum spanning tree, 2-edge, and biconnectivity. Journal of the ACM, 48(4):723–760, July 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hehner:1977:ICP

[Heh77] Eric C. R. Hehner. Information content of programs and operation encoding. Journal of the ACM, 24(2):290–297, April 1977.

CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Heindel:1971:IAA

[Hei71] Lee E. Heindel. Integer arithmetic algorithms for polynomial real zero determination. Journal of the ACM, 18(4):533–548, October 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Heller:1961:SAM

[Hel61] J. Heller. Sequencing aspects of multiprogramming. Journal of the ACM, 8(3):426–439, July 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Helman:1989:CSD

[Hel89] Paul Helman. A common schema for dynamic programming and branch and bound algorithms. Journal of the ACM, 36(1):97– 128, January 1989. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/59304.html.

Hemmerle:1964:ASS

[Hem64] W. J. Hemmerle. Algebraic specification of statistical models for analysis of variance computations. Journal of the ACM, 11 (2):234–239, April 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Henrici:1956:ACP

[Hen56a] Peter Henrici. Automatic computations with power series. Journal of the ACM, 3(1):10–15, January 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Henrici:1956:SCR

[Hen56b] Peter Henrici. A subroutine for computations with rational numbers. Journal of the ACM, 3(1):6–9, January 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Henrici:1970:MSS

[Hen70] Peter Henrici. Methods of search for solving polynomial equations. Journal of the ACM, 17(2):273–283, April 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Henschen:1979:TPC

[Hen79] L. J. Henschen. Theorem proving by covering expressions. Journal of the ACM, 26(3):385–400, July 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hennessy:1985:AT

[Hen85] M. Hennessy. Acceptance trees. Journal of the ACM, 32(4):896– 928, October 1985. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/4249.html.

Herr:1974:SMS

[Her74] David G. Herr. On a statistical model of Strand and Westwater for the numerical solution of a Fredholm integral equation of the first kind. Journal of the ACM, 21(1):1–5, January 1974. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Herlihy:1990:CAD

[Her90] Maurice Herlihy. Concurrency and availability as dual properties of replicated atomic data. Journal of the ACM, 37(2):257– 278, April 1990. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/77616.html.

Halton:1957:MIE

[HH57] J. H. Halton and D. C. Hanscomb. A method of increasing the efficiency of Monte Carlo integration. Journal of the ACM, 4(3): 329–340, July 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hartley:1963:MCC

[HH63] H. O. Hartley and D. L. Harris. Monte Carlo computations in normal correlation problems. Journal of the ACM, 10:302–306, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hartmanis:1971:OTC

[HH71] J. Hartmanis and J. E. Hopcroft. An overview of the theory of computational complexity. Journal of the ACM, 18(3):444–476, July 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Harrison:1974:PDL

[HH74] Michael A. Harrison and Ivan M. Havel. On the parsing of deterministic languages. Journal of the ACM, 21(4):525–548, October 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hirst:1994:PBC

[HH94] Tirza Hirst and David Harel. On the power of bounded concurrency II: Pushdown automata. Journal of the ACM, 41(3): 540–554, May 1994. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/176588.html.

Helmert:2014:MSA

[HHHN14] Malte Helmert, Patrik Haslum, Jörg Hoffmann, and Raz Nissim. Merge-and-shrink abstraction: a method for generating lower bounds in factored state spaces. Journal of the ACM, 61(3):16:1– 16:??, May 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Harper:1993:FDL

[HHP93] Robert Harper, Furio Honsell, and Gordon Plotkin. A framework for defining logics. Journal of the ACM, 40(1):143–184, January 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/

0004-5411/138060.html. Preliminary version appeared in Proc. 2nd IEEE Symposium on Logic in Computer Science, 1987, 194– 204.

Helmi:2014:SCL

[HHPW14] Maryam Helmi, Lisa Higham, Eduardo Pacheco, and Philipp Woelfel. The space complexity of long-lived and one-shot timestamp implementations. Journal of the ACM, 61(1):7:1–7:??, January 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hemaspaandra:1997:EAD

[HHR97] Edith Hemaspaandra, Lane A. Hemaspaandra, and Jörg Rothe. Exact analysis of Dodgson elections: Lewis Carroll’s 1876 voting system is complete for parallel access to NP. Journal of the ACM, 44(6):806–825, November 1997. CODEN JACOAH.

ISSN 0004-5411 (print), 1557-735X (electronic). URL http:/

/www.acm.org:80/pubs/citations/journals/jacm/1997-44-
6/p806-hemaspaandra/.

Hildebrandt:1959:REI

[HI59] Paul Hildebrandt and Harold Isbitz. Radix exchange—an internal sorting method for digital computers. Journal of the ACM, 6(2):156–163, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Heath:1992:PGG

[HI92] Lenwood S. Heath and Sorin Istrail. The pagenumber of genus g graphs is O(g). Journal of the ACM, 39(3):479–501, July 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/146643.html.

Hibbard:1961:LUB

[Hib61] Thomas N. Hibbard. Least upper bounds on minimal terminal state experiments for two classes of sequential machines. Journal of the ACM, 8(4):601–612, October 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hibbard:1962:SCP

[Hib62] Thomas N. Hibbard. Some combinatorial properties of certain trees with applications to sorting and searching. Journal of the ACM, 9(1):13–28, January 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hibbard:1963:SSA

[Hib63] Thomas N. Hibbard. A simple sorting algorithm. Journal of the ACM, 10:142–150, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hibbard:1974:CLG

[Hib74] Thomas N. Hibbard. Context-limited grammars. Journal of the ACM, 21(3):446–453, July 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hirschhorn:1958:SCB

[Hir58] Edwin Hirschhorn. Simplification of a class of Boolean functions. Journal of the ACM, 5(1):67–75, January 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hirschberg:1977:ALC

[Hir77] Daniel S. Hirschberg. Algorithms for the longest common subsequence problem. Journal of the ACM, 24(4):664–675, October 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hooke:1961:DSS

[HJ61] Robert Hooke and T. A. Jeeves. “direct search” solution of numerical and statistical problems. Journal of the ACM, 8(2):212– 229, April 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hickey:2001:IAP

[HJV01] T. Hickey, Q. Ju, and M. H. Van Emden. Interval arithmetic: From principles to implementation. Journal of the ACM, 48 (5):1038–1068, September 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://csr.

uvic.ca/~vanemden/Publications/hickeyJuvE.ps;   http:
//www.cs.brandeis.edu/~tim/Papers/jacm01.pdf;   http:
//www.cs.brandeis.edu/~tim/Papers/jacm01.ps;   http:/
/www.cs.brandeis.edu/~tim/Papers/jacm01.ps.gz.

Hamilton:1954:IMD

[HK54] F. E. Hamilton and E. C. Kubie. The IBM Magnetic Drum Calculator Type 650. Journal of the ACM, 1(1):13–20, January 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamer:1955:TOA

[HK55] Howard Hamer and Jerome D. Kennedy. Testing of operational amplifiers. Journal of the ACM, 2(2):92–94, April 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hyafil:1977:CPE

[HK77] L. Hyafil and H. T. Kung. The complexity of parallel evaluation of linear recurrences. Journal of the ACM, 24(3):513–521, July 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Henzinger:1999:RFD

[HK99] Monika R. Henzinger and Valerie King. Randomized fully dynamic graph algorithms with polylogarithmic time per operation. Journal of the ACM, 46(4):502–516, July 1999. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/citations/journals/jacm/
1999-46-4/p502-henzinger/.

Hellerstein:2002:MIB

[HKM^{+}02] Joseph M. Hellerstein, Elias Koutsoupias, Daniel P. Miranker, Christos H. Papadimitriou, and Vasilis Samoladas. On a model of indexability and its bounds for range queries. Journal of the ACM, 49(1):35–55, January 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Harsha:2012:IPP

[HKM12] Prahladh Harsha, Adam Klivans, and Raghu Meka. An invariance principle for polytopes. Journal of the ACM, 59(6):29:1– 29:25, December 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Special issue text supplied by Joanne Pello.

Hassidim:2022:ARS

[HKM^{+}22] Avinatan Hassidim, Haim Kaplan, Yishay Mansour, Yossi Matias, and Uri Stemmer. Adversarially robust streaming algorithms via differential privacy. Journal of the ACM, 69(6): 42:1–42:??, December 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3556972.

Horwitz:1966:IRA

[HKMW66] L. P. Horwitz, R. M. Karp, R. E. Miller, and S. Winograd. Index register allocation. Journal of the ACM, 13(1):43–61, January 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Henzinger:2018:DSS

[HKN18] Monika Henzinger, Sebastian Krinninger, and Danupon Nanongkai. Decremental single-source shortest paths on undirected graphs in near-linear total update time. Journal of the ACM, 65(6): 36:1–36:??, November 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3218657.

Hoover:1984:BFL

[HKP84] H. J. Hoover, M. M. Klawe, and N. J. Pippenger. Bounding fan-out in logical networks. Journal of the ACM, 31(1):13–18, January 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Huang:2020:FOM

[HKT^{+}20] Zhiyi Huang, Ning Kang, Zhihao Gavin Tang, Xiaowei Wu, Yuhao Zhang, and Xue Zhu. Fully online matching. Journal of the ACM, 67(3):17:1–17:25, June 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3390890.

Holtz:1966:SRD

[HL66] Howard Holtz and C. T. Leondes. The synthesis of recursive digital filters. Journal of the ACM, 13(2):262–280, April 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hoffmann:1976:CTS

[HL76] Christoph M. Hoffmann and Lawrence H. Landweber. A completeness theorem for straight-line programs with structured variables. Journal of the ACM, 23(1):203–220, January 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hirschberg:1987:NAF

[HL87] D. S. Hirschberg and L. L. Larmore. New applications of failure functions. Journal of the ACM, 34(3):616–625, July 1987. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
28875.html.

Hillar:2013:MTP

[HL13] Christopher J. Hillar and Lek-Heng Lim. Most tensor problems are NP-Hard. Journal of the ACM, 60(6):45:1–45:??, November 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Herlihy:1992:COM

[HLMW92] Maurice Herlihy, Nancy Lynch, Michael Merritt, and William Weihl. On the correctness of orphan management algorithms.

Journal of the ACM, 39(4):881–930, October 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
146616.html.

Hella:2001:LAO

[HLNW01] Lauri Hella, Leonid Libkin, Juha Nurmonen, and Limsoon Wong. Logics with aggregate operators. Journal of the ACM, 48(4):880– 907, July 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hass:1999:CCK

[HLP99] Joel Hass, Jeffrey C. Lagarias, and Nicholas Pippenger. The computational complexity of knot and link problems. Journal of the ACM, 46(2):185–211, March 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:/

/www.acm.org:80/pubs/citations/journals/jacm/1999-46-
2/p185-hass/.

Horvath:1977:LAP

[HLS77] Edward C. Horvath, Shui Lam, and Ravi Sethi. A level algorithm for preemptive scheduling. Journal of the ACM, 24(1):32–43, January 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Harris:1972:NFQ

[HM72] Carl M. Harris and Paul G. Marlin. A note on feedback queues with bulk service. Journal of the ACM, 19(4):727–733, October 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hennessy:1985:ALN

[HM85a] Matthew Hennessy and Robin Milner. Algebraic laws for nondeterminism and concurrency. Journal of the ACM, 32(1):137– 161, January 1985. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/2460.html.

Hochbaum:1985:ASC

[HM85b] Dorit S. Hochbaum and Wolfgang Maass. Approximation schemes for covering and packing problems in image processing and VLSI. Journal of the ACM, 32(1):130–136, January 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/214106.html.

Halpern:1990:KCK

[HM90] Joseph Y. Halpern and Yoram Moses. Knowledge and common knowledge in a distributed environment. Journal of the ACM, 37(3):549–587, July 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/79161.html. A preliminary

version appeared in Proc. 3rd ACM Symposium on Principles of Distributed Computing, 1984.

Harrow:2013:TPS

[HM13] Aram W. Harrow and Ashley Montanaro. Testing product states, quantum Merlin–Arthur games and tensor optimization. Journal of the ACM, 60(1):3:1–3:43, February 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hong:1983:CTO

[HMR83] Jia-Wei Hong, Kurt Mehlhorn, and Arnold L. Rosenberg. Cost trade-offs in graph embeddings, with applications. Journal of the ACM, 30(4):709–728, October 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hallgren:2010:LQC

[HMR^{+}10] Sean Hallgren, Cristopher Moore, Martin Rötteler, Alexander Russell, and Pranab Sen. Limitations of quantum coset states for graph isomorphism. Journal of the ACM, 57(6):34:1–34:33, October 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Halevy:2001:SAD

[HMSS01] Alon Y. Halevy, Inderpal Singh Mumick, Yehoshua Sagiv, and Oded Shmueli. Static analysis in datalog extensions. Journal of the ACM, 48(5):971–1012, September 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hansen:2013:SIS

[HMZ13] Thomas Dueholm Hansen, Peter Bro Miltersen, and Uri Zwick. Strategy iteration is strongly polynomial for 2-player turn-based stochastic games with a constant discount factor. Journal of the ACM, 60(1):1:1–1:16, February 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Henschen:1984:CQR

[HN84] Lawrence J. Henschen and Shamim A. Naqvi. On compiling queries in recursive first-order databases. Journal of the ACM, 31(1):47–85, January 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Haastad:2004:SAR

[HN04] Johan Håstad and Mats Nåslund. The security of all RSA and discrete log bits. Journal of the ACM, 51(2):187–230, March 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Harju:2007:PUW

[HN07] Tero Harju and Dirk Nowotka. Periodicity and unbordered words: a proof of the extended Duval conjecture. Journal of the ACM, 54(4):20:1–20:20, July 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hoffmann:1982:PMT

[HO82] Christoph M. Hoffmann and Michael J. O’Donnell. Pattern matching in trees. Journal of the ACM, 29(1):68–95, January 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hoare:1978:SPP

[Hoa78] C. A. R. Hoare. Some properties of predicate transformers. Journal of the ACM, 25(3):461–480, July 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hoare:2003:VCG

[Hoa03] Tony Hoare. The verifying compiler: a grand challenge for computing research. Journal of the ACM, 50(1):63–69, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hobby:1989:RCC

[Hob89] John D. Hobby. Rasterizing curves of constant width. Journal of the ACM, 36(2):209–229, April 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/62045.html.

Hobby:1993:GAT

[Hob93] John D. Hobby. Generating automatically tuned bitmaps from outlines. Journal of the ACM, 40(1):48–94, January 1993. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
138040.html.

Hockney:1965:FDS

[Hoc65] R. W. Hockney. A fast direct solution of Poisson’s equation using Fourier analysis. Journal of the ACM, 12(1):95–113, January 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hochbaum:2001:EAI

[Hoc01] Dorit S. Hochbaum. An efficient algorithm for image segmentation, Markov random fields and related problems. Journal of the ACM, 48(4):686–701, July 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hodes:1970:LCG

[Hod70] Louis Hodes. The logical complexity of geometric properties in the plane. Journal of the ACM, 17(2):339–347, April 1970. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hofri:1977:COB

[Hof77] Micha Hofri. On certain output-buffer management techniques — a stochastic model. Journal of the ACM, 24(2):241–249, April 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hofri:1984:AIS

[Hof84] Micha Hofri. Analysis of interleaved storage via a constantservice queuing system with Markov-chain-driven input. Journal of the ACM, 31(3):628–648, July 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hogger:1981:DLP

[Hog81] C. J. Hogger. Derivation of logic programs. Journal of the ACM, 28(2):372–392, April 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Holdiman:1962:MTR

[Hol62a] Thomas A. Holdiman. Management techniques for real time computer programming. Journal of the ACM, 9(3):387–404, July 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Holland:1962:OLT

[Hol62b] John H. Holland. Outline for a logical theory of adaptive systems. Journal of the ACM, 9(3):297–314, July 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Early work on genetic algorithms. Reprinted in [Bur70].

Homer:1987:MDP

[Hom87] Steven Homer. Minimal degrees for polynomial reducibilities. Journal of the ACM, 34(2):480–491, April 1987. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
23009.html.

Honeyman:1982:TSF

[Hon82] Peter Honeyman. Testing satisfaction of functional dependencies. Journal of the ACM, 29(3):668–677, July 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hooper:1966:IPP

[Hoo66a] Philip K. Hooper. The immortality problem for post normal systems. Journal of the ACM, 13(4):594–599, October 1966. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hooper:1966:MPN

[Hoo66b] Philip K. Hooper. Monogenic post normal systems of arbitrary degree. Journal of the ACM, 13(3):359–363, July 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hrushovski:2023:SAP

[HOPW23] Ehud Hrushovski, Joël Ouaknine, Amaury Pouly, and James Worrell. On strongest algebraic program invariants. Journal of the ACM, 70(5):29:1–29:??, October 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3614319.

Horowitz:1975:SAP

[Hor75] Ellis Horowitz. A sorting algorithm for polynomial multiplication. Journal of the ACM, 22(4):450–462, October 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Horvath:1978:SSA

[Hor78] Edward C. Horvath. Stable sorting in asymptotically optimal time and extra space. Journal of the ACM, 25(2):177–199, April 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Householder:1956:BNA

[Hou56a] Alston S. Householder. Bibliography on numerical analysis. Journal of the ACM, 3(2):85–100, April 1956. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Householder:1956:CMI

[Hou56b] Alston S. Householder. On the convergence of matrix iterations. Journal of the ACM, 3(4):314–324, October 1956. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Householder:1956:PAA

[Hou56c] Alston S. Householder. Presidential Address to the ACM, Philadelphia, September 14, 1955. Journal of the ACM, 3(1): 1–2, January 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Householder:1957:RPA

[Hou57] Alston S. Householder. Retiring presidential address. Journal of the ACM, 4(1):1–4, January 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Householder:1958:ASM

[Hou58a] A. S. Householder. The approximate solution of matrix problems. Journal of the ACM, 5(3):205–243, July 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Householder:1958:GER

[Hou58b] Alston S. Householder. Generated error in rotational tridiagonalization. Journal of the ACM, 5(4):335–338, October 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Householder:1958:UTN

[Hou58c] Alston S. Householder. Unitary triangularization of a nonsymmetric matrix. Journal of the ACM, 5(4):339–342, October 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hoffman:1959:MSB

[HP59] Walter Hoffman and Richard Pavley. A method for the solution of the Nth best path problem. Journal of the ACM, 6(4):506– 514, October 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hamming:1966:CIF

[HP66] R. W. Hamming and R. S. Pinkham. A class of integration formulas. Journal of the ACM, 13(3):430–438, July 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hammer:1972:MPB

[HP72a] Peter L. Hammer and Uri N. Peled. On the maximization of a pseudo-Boolean function. Journal of the ACM, 19(2):265–282, April 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Heacox:1972:ATT

[HP72b] Harry C. Heacox, Jr. and Paul W. Purdom, Jr. Analysis of two time-sharing queueing models. Journal of the ACM, 19(1):70– 91, January 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Horowitz:1976:PST

[HP76] Steven L. Horowitz and Theodosios Pavlidis. Picture segmentation by a tree traversal algorithm. Journal of the ACM, 23 (2):368–388, April 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Harrison:1990:RMI

[HP90] Peter G. Harrison and Naresh M. Patel. The representation of multistage interconnection networks in queuing models of parallel systems. Journal of the ACM, 37(4):863–898, October 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/96599.html.

Hannenhalli:1999:TCT

[HP99] Sridhar Hannenhalli and Pavel A. Pevzner. Transforming cabbage into turnip: polynomial algorithm for sorting signed permutations by reversals. Journal of the ACM, 46(1):1–27, January 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org:80/pubs/citations/
journals/jacm/1999-46-1/p1-hannenhalli/.

Halpern:2007:CRA

[HP07] Joseph Y. Halpern and Riccardo Pucella. Characterizing and reasoning about probabilistic and non-probabilistic expectation. Journal of the ACM, 54(3):15:1–15:49, June 2007. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Har-Peled:2015:NPL

[HPR15] Sariel Har-Peled and Benjamin Raichel. Net and prune: a linear time algorithm for Euclidean distance problems. Journal of the ACM, 62(6):44:1–44:??, December 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hellerstein:1996:HMQ

[HPRW96] Lisa Hellerstein, Krishnan Pillaipakkamnatt, Vijay Raghavan, and Dawn Wilkins. How many queries are needed to learn? Journal of the ACM, 43(5):840–862, September 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
jacm/234755.html.

Hopcroft:1977:TVS

[HPV77] John Hopcroft, Wolfgang Paul, and Leslie Valiant. On time versus space. Journal of the ACM, 24(2):332–337, April 1977. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Huang:2024:BAO

[HPZ24] Shang-En Huang, Seth Pettie, and Leqi Zhu. Byzantine agreement with optimal resilience via statistical fraud detection. Journal of the ACM, 71(2):12:1–12:??, April 2024. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3639454.

House:1963:ECP

[HR63a] R. W. House and T. Rado. Erratum: “On a computer program for obtaining irreducible representations for two-level multiple input-output logical systems”. Journal of the ACM, 10: 256, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [HR63b].

House:1963:CPO

[HR63b] R. W. House and T. Rado. On a computer program for obtaining irreducible representations for two-level multiple inputoutput logical systems. Journal of the ACM, 10(1):48–77, January 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [HR63a].

Hunt:1977:ECP

[HR77] Harry B. Hunt III and Daniel J. Rosenkrantz. On equivalence and containment problems for formal languages. Journal of the ACM, 24(3):387–396, July 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Harrison:1978:AGR

[HR78] Harrison and Rubin. Another generalization of resolution. Journal of the ACM, 25(2):341–351, April 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hsiang:1991:PRC

[HR91] Jieh Hsiang and Michaël Rusinowitch. Proving refutational completeness of theorem-proving strategies: The transfinite semantic tree method. Journal of the ACM, 38(3):559–587, July 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/116833.html.

Hollender:2025:EFC

[HR25] Alexandros Hollender and Aviad Rubinstein. Envy-free cakecutting for four agents. Journal of the ACM, 72(5):34:1–34:54, October 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Huson:2002:GPM

[HRM02] Daniel H. Huson, Knut Reinert, and Eugene W. Myers. The greedy path-merging algorithm for contig scaffolding. Journal of the ACM, 49(5):603–615, September 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Heath:1988:PMP

[HRS88] Lenwood S. Heath, Arnold L. Rosenberg, and Bruce T. Smith. The physical mapping problem for parallel architectures. Journal of the ACM, 35(3):603–634, July 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/44489.html.

Haghtalab:2024:SAA

[HRS24] Nika Haghtalab, Tim Roughgarden, and Abhishek Shetty. Smoothed analysis with adaptive adversaries. Journal of the ACM, 71(3):19:1–19:??, June 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/doi/10.1145/3656638.

Haastad:2017:ACD

[HRST17] Johan Håstad, Benjamin Rossman, Rocco A. Servedio, and Li-Yang Tan. An average-case depth hierarchy theorem for Boolean circuits. Journal of the ACM, 64(5):35:1–35:??, October 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Haneman:1955:CCA

[HS55] V. S. Haneman and J. W. Senders. Correlation computation on analog devices. Journal of the ACM, 2(4):267–279, October 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hempstead:1956:PLE

[HS56] Gus Hempstead and Jules I. Schwartz. PACT loop expansion. Journal of the ACM, 3(4):292–298, October 1956. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hennie:1966:TTS

[HS66] F. C. Hennie and R. E. Stearns. Two-tape simulations of multitape Turing machines. Journal of the ACM, 13(4):533–546, October 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hartmanis:1968:RPA

[HS68] J. Hartmanis and H. Shank. On the recognition of primes by automata. Journal of the ACM, 15(3):382–389, July 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Harrison:1971:GCO

[HS71] Michael A. Harrison and Mario Schkolnick. A grammatical characterization of one-way nondeterministic stack languages. Journal of the ACM, 18(1):148–172, January 1971. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Horowitz:1974:CPA

[HS74] Ellis Horowitz and Sartaj Sahni. Computing partitions with applications to the knapsack problem. Journal of the ACM, 21 (2):277–292, April 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Horowitz:1975:CED

[HS75] E. Horowitz and S. Sahni. On computing the exact determinant of matrices with polynomial entries. Journal of the ACM, 22 (1):38–50, January 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Horowitz:1976:EAA

[HS76] Ellis Horowitz and Sartaj Sahni. Exact and approximate algorithms for scheduling nonidentical processors. Journal of the ACM, 23(2):317–327, April 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hunt:1978:CLB

[HS78a] H. B. Hunt, III and T. G. Szymanski. Corrigendum: “Lower bounds and reductions between grammar problems”. Journal of the ACM, 26(4):687–688, October 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [HS78b].

Hunt:1978:LBR

[HS78b] Harry B. Hunt, III and Thomas G. Szymanski. Lower bounds and reductions between grammar problems. Journal of the ACM, 25(1):32–51, January 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [HS78a].

Hochbaum:1986:UAA

[HS86] Dorit S. Hochbaum and David B. Shmoys. A unified approach to approximation algorithms for bottleneck problems. Journal of the ACM, 33(3):533–550, July 1986. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/5933.html.

Hochbaum:1987:UDA

[HS87] Dorit S. Hochbaum and David B. Shmoys. Using dual approximation algorithms for scheduling problems: Theoretical and practical results. Journal of the ACM, 34(1):144–162, January 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/7535.html.

Hochbaum:1990:CSO

[HS90] Dorit S. Hochbaum and J. George Shanthikumar. Convex separable optimization is not much harder than linear optimization. Journal of the ACM, 37(4):843–862, October 1990. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
96597.html.

Halpern:1991:PML

[HS91] Joseph Y. Halpern and Yoav Shoham. A propositional modal logic of time intervals. Journal of the ACM, 38(4):935–962, October 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/115351.html.

Herlihy:1999:TSA

[HS99] Maurice Herlihy and Nir Shavit. The topological structure of asynchronous computability. Journal of the ACM, 46(6):858–923, November 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:

//www.acm.org/pubs/articles/journals/jacm/1999-46-6/
p858-herlihy/p858-herlihy.pdf.

Harris:2019:MTF

[HS19] David G. Harris and Aravind Srinivasan. The Moser–Tardos framework with partial resampling. Journal of the ACM, 66(5): 36:1–36:??, September 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3342222.

Haeupler:2021:SSC

[HS21] Bernhard Haeupler and Amirbehshad Shahrasbi. Synchronization strings: Codes for insertions and deletions approaching the singleton bound. Journal of the ACM, 68(5):36:1–36:39, October 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3468265.

Haeupler:2011:NCA

[HSS11] Bernhard Haeupler, Barna Saha, and Aravind Srinivasan. New constructive aspects of the Lovász Local Lemma. Journal of the ACM, 58(6):28:1–28:28, December 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Harris:2018:DCS

[HSS18] David G. Harris, Johannes Schneider, and Hsin-Hao Su. Distributed (Δ + 1)-coloring in sublogarithmic rounds. Journal of the ACM, 65(4):19:1–19:??, August 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Heller:1976:AIM

[HST76] D. E. Heller, D. K. Stevenson, and J. F. Traub. Accelerated iterative methods for the solution of tridiagonal systems on parallel computers. Journal of the ACM, 23(4):636–654, October 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hon:2014:SEF

[HSTV14] Wing-Kai Hon, Rahul Shah, Sharma V. Thankachan, and Jeffrey Scott Vitter. Space-efficient frameworks for top-k string retrieval. Journal of the ACM, 61(2):9:1–9:??, April 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hsu:1975:AFM

[Hsu75] Harry T. Hsu. An algorithm for finding a minimal equivalent graph of a digraph. Journal of the ACM, 22(1):11–16, January 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hsu:1987:RPP

[Hsu87] Wen-Lian Hsu. Recognizing planar perfect graphs. Journal of the ACM, 34(2):255–288, April 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/31330.html.

Hsu:1988:CMI

[Hsu88] Wen-Lian Hsu. The coloring and maximum independent set problems on planar perfect graphs. Journal of the ACM, 35(3): 535–563, July 1988. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/44486.html.

Hopcroft:1974:EPT

[HT74] John Hopcroft and Robert Tarjan. Efficient planarity testing. Journal of the ACM, 21(4):549–568, October 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See comments and errata in [Deo76].

Halpern:1993:KPA

[HT93] Joseph Y. Halpern and Mark R. Tuttle. Knowledge, probability, and adversaries. Journal of the ACM, 40(4):917–962, September 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/153770.html.

Hu:2024:PAJ

[HT24] Xiao Hu and Yufei Tao. Parallel acyclic joins: Optimal algorithms and cyclicity separation. Journal of the ACM, 71(1): 6:1–6:??, February 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/

doi/10.1145/3633512.

Hibbard:1966:IIA

[HU66] Thomas N. Hibbard and Joseph Ullian. The independence of inherent ambiguity from complementedness among context-free languages. Journal of the ACM, 13(4):588–593, October 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hopcroft:1968:DUQ

[HU68a] J. E. Hopcroft and J. D. Ullman. Decidable and undecidable questions about automata. Journal of the ACM, 15(2):317–324, April 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hopcroft:1968:RBT

[HU68b] John E. Hopcroft and Jeffrey D. Ullman. Relations between time and tape complexities. Journal of the ACM, 15(3):414–427, July 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hopcroft:1969:SRT

[HU69] J. E. Hopcroft and J. D. Ullman. Some results on tape-bounded Turing machines. Journal of the ACM, 16(1):168–177, January 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hecht:1974:CRF

[HU74] M. S. Hecht and J. D. Ullman. Characterizations of reducible flow graphs. Journal of the ACM, 21(3):367–375, July 1974. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hueckel:1971:OWL

[Hue71] Manfred H. Hueckel. An operator which locates edges in digitized pictures. Journal of the ACM, 18(1):113–125, January 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hueckel:1973:LVO

[Hue73] Manfred H. Hueckel. A local visual operator which recognizes edges and lines. Journal of the ACM, 20(4):634–647, October 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [Hue74].

Hueckel:1974:ELV

[Hue74] Manfred H. Hueckel. Erratum: “A local visual operator which recognizes edges and lines”. Journal of the ACM, 21(2):350, April 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Hue73].

Huet:1980:CRA

[Hue80] Gérard Huet. Confluent reductions: Abstract properties and applications to term rewriting systems. Journal of the ACM, 27 (4):797–821, October 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Preliminary version in Proceedings, 18th IEEE Symposium on Foundations of Computer Science, IEEE, 1977, pages 30–45.

Huffman:1957:DUH

[Huf57] David A. Huffman. The design and use of hazard-free switching networks. Journal of the ACM, 4(1):47–62, January 1957. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hull:1984:FSI

[Hul84] Richard Hull. Finitely specifiable implicational dependency families. Journal of the ACM, 31(2):210–226, April 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hunt:1982:CFL

[Hun82a] H. B. Hunt, III. On the complexity of flowchart and loop program schemes and programming languages. Journal of the ACM, 29 (1):228–249, January 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hunt:1982:DGP

[Hun82b] H. B. Hunt, III. On the decidability of grammar problems. Journal of the ACM, 29(2):429–447, April 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hunt:1984:TTM

[Hun84] H. B. Hunt III. Terminating Turing machine computations and the complexity and/or decidability of correspondence problems, grammars, and program schemes. Journal of the ACM, 31 (2):299–318, April 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Haldar:1995:CWM

[HV95] S. Haldar and K. Vidyasankar. Constructing 1-writer multiread multivalued atomic variables from regular variables. Journal of the ACM, 42(1):186–203, January 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/200871.html.

Haldar:2002:BCT

[HV02] Sibsankar Haldar and Paul Vitányi. Bounded concurrent timestamp systems using vector clocks. Journal of the ACM, 49(1):

101–126, January 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Haldar:2007:SRW

[HV07] Sibsankar Haldar and K. Vidyasankar. On specification of Read/ Write shared variables. Journal of the ACM, 54(6):31:1–31:19, December 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Harvey:2022:PMF

[HvdH22] David Harvey and Joris van der Hoeven. Polynomial multiplication over finite fields in time O(n log n). Journal of the ACM, 69(2):12:1–12:40, April 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3505584.

Harvey:2017:FPM

[HVL17] David Harvey, Joris Van Der Hoeven, and Grégoire Lecerf. Faster polynomial multiplication over finite fields. Journal of the ACM, 63(6):52:1–52:??, February 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hume:1955:TSA

[HW55] J. N. P. Hume and Beatrice H. Worsley. Transcode: a system of automatic coding for FERUT. Journal of the ACM, 2(4):243– 252, October 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hicks:1967:NSP

[HW67] J. S. Hicks and J. Wei. Numerical solution of parabolic partial differential equations with two-point boundary conditions by use of the method of lines. Journal of the ACM, 14(3):549–562, July 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Henschen:1974:URH

[HW74] L. Henschen and L. Wos. Unit refutations and Horn sets. Journal of the ACM, 21(4):590–605, October 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hirschberg:1976:PTA

[HW76] D. S. Hirschberg and C. K. Wong. A polynomial-time algorithm for the knapsack problem with two variables. Journal of the ACM, 23(1):147–154, January 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Han:1990:EFP

[HW90] Yijie Han and Robert A. Wagner. An efficient and fast parallelconnected component algorithm. Journal of the ACM, 37(3): 626–642, July 1990. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/214077.html.

Hochstein:2004:EDR

[HW04] Jan M. Hochstein and Karsten Weihe. Edge-disjoint routing in plane switch graphs in linear time. Journal of the ACM, 51(4): 636–670, July 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Halpern:2008:FFX

[HW08] Joseph Y. Halpern and Vicky Weissman. A formal foundation for XrML. Journal of the ACM, 55(1):4:1–4:42, February 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hwang:1979:ARM

[Hwa79] F. K. Hwang. An O(n log n) algorithm for rectilinear minimal spanning tree. Journal of the ACM, 26(2):177–182, April 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hoover:1956:WTD

[HWB56] William R. Hoover, John J. Wedel, and Joseph R. Bruman. Wind tunnel data reduction using paper-tape storage media. Journal of the ACM, 3(2):101–109, April 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Halpern:1990:CRR

[HWW90] Joseph Y. Halpern, John H. Williams, and Edward L. Wimmers. Completeness of rewrite rules and rewrite strategies for FP. Journal of the ACM, 37(1):86–143, January 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/78939.html.

Hofri:1975:PSD

[HY75] Micha Hofri and Micha Yadin. A processor in series with demand-interrupting devices — a stochastic model. Journal of the ACM, 22(2):270–290, April 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hwang:1977:OBS

[HY77] K. Hwang and S. B. Yao. Optimal batched searching of tree structured files in multiprocessor computer systems. Journal of the ACM, 24(3):441–454, July 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hull:1984:FMT

[HY84] Richard Hull and Chee K. Yap. The format model: a theory of database organization. Journal of the ACM, 31(3):518–537, July 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Also published in/as: ACM SIGACT-SIGMOD Symp. on Principles of Database Systems, 1982, pp.205–211.

Honda:2016:MAS

[HYC16] Kohei Honda, Nobuko Yoshida, and Marco Carbone. Multiparty asynchronous session types. Journal of the ACM, 63(1):9:1–9:??, March 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Henderson:1980:SPS

[HZ80] Peter B. Henderson and Yechezkel Zalcstein. Synchronization problems solvable by generalized PV systems. Journal of the ACM, 27(1):60–71, January 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Halpern:1992:LKG

[HZ92] Joseph Y. Halpern and Lenore D. Zuck. A little knowledge goes a long way: Knowledge-Based derivations and correctness proofs for a family of protocols. Journal of the ACM, 39(3):449– 478, July 1992. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/146638.html.

Herrmann:2016:CCQ

[HZ16] Christian Herrmann and Martin Ziegler. Computational complexity of quantum satisfiability. Journal of the ACM, 63(2):

19:1–19:??, May 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ibaraki:1983:DMC

[IAWK83] Toshihide Ibaraki, Hussein M. Abdel-Wahab, and Tiko Kameda. Design of minimum-cost deadlock-free systems. Journal of the ACM, 30(4):736–751, October 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ibarra:1972:NCN

[Iba72] Oscar H. Ibarra. A note concerning nondeterministic tape complexities. Journal of the ACM, 19(4):608–612, October 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ibaraki:1977:PDR

[Iba77] Toshihide Ibaraki. The power of dominance relations in branchand-bound algorithms. Journal of the ACM, 24(2):264–279, April 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ibarra:1978:RBM

[Iba78] O. H. Ibarra. Reversal-bounded multicounter machines and their decision problems. Journal of the ACM, 25(1):116–133, January 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Inselberg:1987:CAP

[ICR87] Alfred Inselberg, Tuval Chomut, and Mordechai Reif. Convexity algorithms in parallel coordinates. Journal of the ACM, 34 (4):765–801, October 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/32221.html.

Iwata:2001:CSP

[IFF01] Satoru Iwata, Lisa Fleischer, and Satoru Fujishige. A combinatorial strongly polynomial algorithm for minimizing submodular functions. Journal of the ACM, 48(4):761–777, July 2001. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Iglehart:1976:SSS

[Igl76] Donald L. Iglehart. Simulating stable stochastic systems, VI: Quantile estimation. Journal of the ACM, 23(2):347–360, April 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ingargiola:1974:FOD

[IK74] Giorgio Ingargiola and James F. Korsh. Finding optimal demand paging algorithms. Journal of the ACM, 21(1):40–53, January 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ibarra:1975:FAA

[IK75] Oscar H. Ibarra and Chul E. Kim. Fast approximation algorithms for the knapsack and sum of subset problems. Journal of the ACM, 22(4):463–468, October 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ibarra:1977:HAS

[IK77] Oscar H. Ibarra and Chul E. Kim. Heuristic algorithms for scheduling independent tasks on nonidentical processors. Journal of the ACM, 24(2):280–289, April 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ikenmeyer:2019:CHF

[IKL^{+}19] Christian Ikenmeyer, Balagopal Komarath, Christoph Lenzen, Vladimir Lysikov, Andrey Mokhov, and Karteek Sreenivasaiah. On the complexity of hazard-free circuits. Journal of the ACM, 66(4):25:1–25:??, August 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/ft_gateway.cfm?id=3320123.

Im:2017:CAC

[IKM17] Sungjin Im, Janardhan Kulkarni, and Kamesh Munagala. Competitive algorithms from competitive equilibria: Non-clairvoyant scheduling under polyhedral constraints. Journal of the ACM, 65(1):3:1–3:??, December 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Iglehart:1979:RSI

[IL79] Donald L. Iglehart and Peter A. W. Lewis. Regenerative simulation with internal controls. Journal of the ACM, 26(2):271–282, April 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ibarra:1983:SEP

[IL83] Oscar H. Ibarra and Brian S. Leininger. On the simplification and equivalence problems for straight-line programs. Journal of the ACM, 30(3):641–656, July 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Imielinski:1984:IIR

[IL84] Tomasz Imieliński and Witold Lipski, Jr. Incomplete information in relational databases. Journal of the ACM, 31(4):761–791, October 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Irving:1987:EAO

[ILG87] Robert W. Irving, Paul Leather, and Dan Gusfield. An efficient algorithm for the “optimal” stable marriage. Journal of the ACM, 34(3):532–543, July 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/28871.html.

Ishii:1997:OTP

[ILP97] Alexander T. Ishii, Charles E. Leiserson, and Marios C. Papaefthymiou. Optimizing two-phase, level-clocked circuitry. Journal of the ACM, 44(1):148–199, January 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ibarra:1983:PAD

[IM83] Oscar H. Ibarra and Shlomo Moran. Probabilistic algorithms for deciding equivalence of straight-line programs. Journal of the ACM, 30(1):217–228, January 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). They study the complexity of deciding the equivalence of straight-line programs, i.e., those in which there are no loops, and only statements of the form x := y, x := y + z, x := y − z, and x := y ∗ z are permitted. Given two such programs P and Q, Ibarra and Moran ask the question: Is P = Q? If the domain of the variables is an infinite field such as the integers, then they show that there exists a polynomial-time probabilistic algorithm to solve this problem. If the domain is a finite field, the problem is shown to be NP-hard.

Imielinski:1991:AQP

[Imi91] Tomasz Imielinski. Abstraction in query processing. Journal of the ACM, 38(3):534–558, July 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/116832.html.

Impagliazzo:2019:PS

[IMZ19] Russell Impagliazzo, Raghu Meka, and David Zuckerman. Pseudorandomness from shrinkage. Journal of the ACM, 66(2): 11:1–11:??, April 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3230630.

Indyk:2006:SDP

[Ind06] Piotr Indyk. Stable distributions, pseudorandom generators, embeddings, and data stream computation. Journal of the ACM, 53(3):307–323, May 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Iwata:2019:IRD

[IOT19] Satoru Iwata, Taihei Oki, and Mizuyo Takamatsu. Index reduction for differential-algebraic equations with mixed matrices. Journal of the ACM, 66(5):35:1–35:??, September 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/ft_gateway.cfm?id=3341499.

Itai:1978:SMP

[IRT78] Alon Itai, Michael Rodeh, and Steven L. Tanimoto. Some matching problems for bipartite graphs. Journal of the ACM, 25(4): 517–525, October 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Isaac:1956:SAC

[IS56] E. J. Isaac and R. C. Singleton. Sorting by address calculation. Journal of the ACM, 3(3):169–174, July 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Iglehart:1978:RSR

[IS78] Donald L. Iglehart and Gerald S. Shedler. Regenerative simulation of response times in networks of queues. Journal of the ACM, 25(3):449–460, July 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Israel:1957:STT

[Isr57] David R. Israel. Simulation techniques for the test and evaluation of real-time computer programs. Journal of the ACM, 4(3):354– 361, July 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Immorlica:2022:ABK

[ISSS22] Nicole Immorlica, Karthik Sankararaman, Robert Schapire, and Aleksandrs Slivkins. Adversarial bandits with knapsacks. Journal of the ACM, 69(6):40:1–40:??, December 2022. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3557045.

Itai:1978:TCF

[Ita78] Alon Itai. Two-commodity flow. Journal of the ACM, 25(4):596– 611, October 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Irani:1971:NLC

[IW71] K. B. Irani and V. L. Wallace. On network linguistics and the conversational design of queueing networks. Journal of the ACM, 18(4):616–629, October 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ioannidis:1991:TAT

[IW91] Yannis E. Ioannidis and Eugene Wong. Towards an algebraic theory of recursion. Journal of the ACM, 38(2):329–381, April 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/103521.html.

JaJa:1980:CBF

[Ja’80] Joseph Ja’Ja’. Computation of bilinear forms over finite fields. Journal of the ACM, 27(4):822–830, October 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

JaJa:1983:TST

[Ja’83] Joseph Ja’Ja’. Time-space trade-offs for some algebraic problems. Journal of the ACM, 30(3):657–667, July 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

JaJa:1984:VCS

[Ja’84] Joseph Ja’Ja’. The VLSI complexity of selected graph problems. Journal of the ACM, 31(2):377–391, April 1984. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jacobs:1954:ERA

[Jac54] H. Jacobs, Jr. Equipment reliability as applied to analogue computers. Journal of the ACM, 1(1):21–26, January 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jacobs:1982:DL

[Jac82] Barry E. Jacobs. On database logic. Journal of the ACM, 29 (2):310–332, April 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jaffe:1983:DSR

[Jaf83] Jeffrey M. Jaffe. Decentralized simulation of resource managers. Journal of the ACM, 30(2):300–322, April 1983. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jaffar:1990:MCW

[Jaf90] Joxan Jaffar. Minimal and complete word unification. Journal of the ACM, 37(1):47–85, January 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/78938.html.

Jain:2015:NSD

[Jai15] Rahul Jain. New strong direct product results in communication complexity. Journal of the ACM, 62(3):20:1–20:??, June 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jayanti:1997:RWF

[Jay97] Prasad Jayanti. Robust wait-free hierarchies. Journal of the ACM, 44(4):592–614, July 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:/

/www.acm.org:80/pubs/citations/journals/jacm/1997-44-
4/p592-jayanti/.

Jazayeri:1981:SCS

[Jaz81] Mehdi Jazayeri. A simpler construction for showing the intrinsically exponential complexity of the circularity problem for attribute grammars. Journal of the ACM, 28(4):715–720, October 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [Dil89].

Jeavons:1997:CPC

[JCG97] Peter Jeavons, David Cohen, and Marc Gyssens. Closure properties of constraints. Journal of the ACM, 44(4):527–548, July 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org:80/pubs/citations/
journals/jacm/1997-44-4/p527-jeavons/.

Jayanti:1998:FTW

[JCT98] Prasad Jayanti, Tushar Deepak Chandra, and Sam Toueg. Faulttolerant wait-free shared objects. Journal of the ACM, 45(3): 451–500, May 1998. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org:80/pubs/
citations/journals/jacm/1998-45-3/p451-jayanti/.

Jeffrey:1956:AAD

[Jef56] Richard C. Jeffrey. Arithmetical analysis of digital computing nets. Journal of the ACM, 3(4):360–375, October 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [CM57].

Jenkins:1973:BMI

[Jen73] M. A. Jenkins. Bernouilli’s method with implicit shifting. Journal of the ACM, 20(3):539–544, July 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jez:2016:RSP

[Jez16] Artur Jez. Recompression: a simple and powerful technique for word equations. Journal of the ACM, 63(1):4:1–4:??, March 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jez:2019:DCU

[Jez19] Artur Jez. Deciding context unification. Journal of the ACM, 66(6):39:1–39:??, October 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/ft_gateway.cfm?id=3356904.

Jiang:2023:MCF

[Jia23] Haotian Jiang. Minimizing convex functions with rational minimizers. Journal of the ACM, 70(1):5:1–5:??, February 2023. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3566050.

Jain:2011:QP

[JJUW11] Rahul Jain, Zhengfeng Ji, Sarvagya Upadhyay, and John Watrous. QIP = PSPACE. Journal of the ACM, 58(6):30:1–30:27, December 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Johnson:1978:LBS

[JK78] Donald B. Johnson and Samuel D. Kashdan. Lower bounds for selection in X + Y and other multisets. Journal of the ACM, 25 (4):556–570, October 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jin:2023:FPA

[JL23] Yaonan Jin and Pinyan Lu. First price auction is 1 − 1/e^{2} efficient. Journal of the ACM, 70(5):36:1–36:??, October 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3617902.

Jiang:2000:LBA

[JLV00] Tao Jiang, Ming Li, and Paul Vitányi. A lower bound on the average-case complexity of Shellsort. Journal of the ACM, 47(5):905–911, 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/citations/journals/jacm/2000-47-5/p905-jiang/.

Juncosa:1960:ICR

[JM60] M. L. Juncosa and T. W. Mullikin. On the increase of convergence rates of relaxation procedures for elliptic partial differential equations. Journal of the ACM, 7(1):29–36, January 1960. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jones:1977:ESP

[JM77] N. D. Jones and S. S. Muchnick. Even simple programs are hard to analyze. Journal of the ACM, 24(2):338–350, April 1977. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jones:1978:CFM

[JM78] Neil D. Jones and Steven S. Muchnick. The complexity of finite memory programs with recursion. Journal of the ACM, 25 (2):312–321, April 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jing:2024:EPT

[JM24] Yifan Jing and Bojan Mohar. Efficient polynomial-time approximation scheme for the genus of dense graphs. Journal of the ACM, 71(6):38:1–38:??, December 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https:

//dl.acm.org/doi/10.1145/3690821.

Jean-Marie:1993:PQR

[JMG93] Alain Jean-Marie and Levent Gün. Parallel queues with resequencing. Journal of the ACM, 40(5):1188–1208, November 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/169748.html.

Jain:2003:GFL

[JMM^{+}03] Kamal Jain, Mohammad Mahdian, Evangelos Markakis, Amin Saberi, and Vijay V. Vazirani. Greedy facility location algorithms analyzed using dual fitting with factor-revealing LP. Journal of the ACM, 50(6):795–824, November 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Johnson:1979:RAF

[JMMW79] Donald B. Johnson, Webb Miller, Brian Minnihan, and Celia Wrathall. Reducibility among floating-point graphs. Journal of the ACM, 26(4):739–760, October 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jin:2021:NOM

[JNG^{+}21] Chi Jin, Praneeth Netrapalli, Rong Ge, Sham M. Kakade, and Michael I. Jordan. On nonconvex optimization for machine learning: Gradients, stochasticity, and saddle points. Journal of the ACM, 68(2):11:1–11:29, March 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/doi/10.1145/3418526.

Joachim:1959:ME

[Joa59] Gertrud S. Joachim. Memory efficiency. Journal of the ACM, 6(2):172–175, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Johansen:1961:MGM

[Joh61] Donald E. Johansen. A modified Givens method for the eigenvalue evaluation of large matrices. Journal of the ACM, 8(3):

331–335, July 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Johnson:1971:PRZ

[Joh71] S. C. Johnson. On the problem of recognizing zero. Journal of the ACM, 18(4):559–565, October 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Johnson:1973:NDS

[Joh73] Donald B. Johnson. A note on Dijkstra’s shortest path algorithm. Journal of the ACM, 20(3):385–388, July 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Johnson:1977:EAS

[Joh77] Donald B. Johnson. Efficient algorithms for shortest paths in sparse networks. Journal of the ACM, 24(1):1–13, January 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Johnson:1987:PAM

[Joh87] Donald B. Johnson. Parallel algorithms for minimum cuts and maximum flows in planar networks. Journal of the ACM, 34 (4):950–967, October 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/31849.html. Prelimi-

nary version in Proc. 23rd Annual IEEE Symposium on the Foundations of Computer Science, pages 244–254, 1982.

Johnson:1993:FCC

[Joh93] C. A. Johnson. Factorization and circuit in the connection method. Journal of the ACM, 40(3):536–557, July 1993. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
174136.html.

Jones:1962:ANA

[Jon62] Terence G. Jones. An algorithm for the numerical application of a linear operator. Journal of the ACM, 9(4):440–449, October 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Staff:2003:JEC

[Jou03] Journal of the ACM staff. JACM editors-in-chief. Journal of the ACM, 50(1):9, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Joyner:1976:RSD

[Joy76] William H. Joyner, Jr. Resolution strategies as decision procedures. Journal of the ACM, 23(3):398–417, July 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

JaJa:1984:ITD

[JP84] J. Ja’Ja’ and V. K. Prasanna Kumar. Information transfer in distributed computing with applications to VLSI. Journal of the ACM, 31(1):150–162, January 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jouannaud:2007:PHO

[JR07] Jean-Pierre Jouannaud and Albert Rubio. Polymorphic higherorder recursive path orderings. Journal of the ACM, 54(1):2:1– 2:48, March 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jain:2009:PQR

[JRS09] Rahul Jain, Jaikumar Radhakrishnan, and Pranab Sen. A property of quantum relative entropy with an application to privacy in quantum communication. Journal of the ACM, 56(6):33:1– 33:32, September 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jacob:2014:SSS

[JRS^{+}14] Riko Jacob, Andrea Richa, Christian Scheideler, Stefan Schmid, and Hanjo Täubig. SKIP +: a self-stabilizing skip graph. Journal of the ACM, 61(6):36:1–36:??, November 2014. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jerrum:1982:SEC

[JS82] Mark Jerrum and Marc Snir. Some exact complexity results for straight-line computations over semirings. Journal of the ACM, 29(3):874–897, July 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Joung:1996:CSC

[JS96] Yuh-Jzer Joung and Scott A. Smolka. A comprehensive study of the complexity of multiparty interaction. Journal of the ACM, 43(1):75–115, January 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/jacm/227601.html.

Jancar:2008:UBD

[JS08] Petr Jančar and Jivří Srba. Undecidability of bisimilarity by defender’s forcing. Journal of the ACM, 55(1):5:1–5:26, February 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jayasimha:2003:FDD

[JSMM03] D. N. Jayasimha, Loren Schwiebert, D. Manivannan, and Jeff A. May. A foundation for designing deadlock-free routing algorithms in wormhole networks. Journal of the ACM, 50(2):250– 275, March 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jansson:2016:IAC

[JSS16] Jesper Jansson, Chuanqi Shen, and Wing-Kin Sung. Improved algorithms for constructing consensus trees. Journal of the ACM, 63(3):28:1–28:??, September 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jiang:1997:THB

[JSV97] Tao Jiang, Joel I. Seiferas, and Paul M. B. Vitányi. Two heads are better than two tapes. Journal of the ACM, 44(2):237–256, March 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jerrum:2004:PTA

[JSV04] Mark Jerrum, Alistair Sinclair, and Eric Vigoda. A polynomialtime approximation algorithm for the permanent of a matrix with nonnegative entries. Journal of the ACM, 51(4):671–697, July 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jump:1975:IAC

[JT75] J. Robert Jump and P. S. Thiagarajan. On the interconnection of asynchronous control structures. Journal of the ACM, 22(4):

596–612, October 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Juncosa:2003:T

[Jun03] Mario L. Juncosa. Transitions 1959–1964. Journal of the ACM, 50(1):12–13, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jain:2001:AAM

[JV01] Kamal Jain and Vijay V. Vazirani. Approximation algorithms for metric facility location and k -Median problems using the primaldual schema and Lagrangian relaxation. Journal of the ACM, 48 (2):274–296, March 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/citations/journals/jacm/2001-48-2/p274-jain/.

Jagadeesan:2023:LEM

[JWW^{+}23] Meena Jagadeesan, Alexander Wei, Yixin Wang, Michael I. Jordan, and Jacob Steinhardt. Learning equilibria in matching markets with bandit feedback. Journal of the ACM, 70(3):19:1–19:??, June 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3583681.

Kosaraju:1988:OSB

[KA88] S. Rao Kosaraju and Mikhail J. Atallah. Optimal simulations between mesh-connected arrays of processors. Journal of the ACM, 35(3):635–650, July 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/44494.html.

Kacewicz:1979:IKS

[Kac79] Boleslaw Kacewicz. Integrals with a kernel in the solution of nonlinear equations in N dimensions. Journal of the ACM, 26 (2):239–249, April 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kahale:1995:EER

[Kah95] Nabil Kahale. Eigenvalues and expansion of regular graphs. Journal of the ACM, 42(5):1091–1106, September 1995. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
210136.html.

Kaltofen:1988:GCD

[Kal88] Erich Kaltofen. Greatest common divisors of polynomials given by straight-line programs. Journal of the ACM, 35(1):231–264, January 1988. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/45069.html.

Kameda:1980:TDF

[Kam80] Tiko Kameda. Testing deadlock-freedom of computer systems. Journal of the ACM, 27(2):270–280, April 1980. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kameda:1982:FSQ

[Kam82] Hisao Kameda. A finite-source queue with different customers. Journal of the ACM, 29(2):478–491, April 1982. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kaminski:1987:LTA

[Kam87] Michael Kaminski. A linear time algorithm for residue computation and a fast algorithm for division with a sparse divisor. Journal of the ACM, 34(4):968–984, October 1987. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
31850.html.

Kaminski:1989:NPV

[Kam89] Michael Kaminski. A note on probabilistically verifying integer and polynomial products. Journal of the ACM, 36(1): 142–149, January 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/214082.html. The author

describes probabilistic algorithms for verifying the product of two n-bit integers in O(n) bit operations, and for verifying the product of two polynomials of degree n over integral domains in 4n + o(n) algebraic operations. The error probability is o(\frac{1}{n^{1−ϵ}}) for any ϵ > 0.

Kanner:1965:NBC

[Kan65] Herbert Kanner. Number base conversion in significant digit arithmetic. Journal of the ACM, 12(2):242–246, April 1965. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kannan:1980:PAT

[Kan80] Ravindran Kannan. A polynomial algorithm for the two-variable integer programming problem. Journal of the ACM, 27(1):118– 122, January 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kannan:1983:PTA

[Kan83] Ravindran Kannan. Polynomial-time aggregation of integer programming problems. Journal of the ACM, 30(1):133–145, January 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kaplan:1968:SCR

[Kap68] Donald M. Kaplan. Some completeness results in the mathematical theory of computation. Journal of the ACM, 15(1):124–134, January 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Karp:1967:SBS

[Kar67] Richard M. Karp. Some bounds on the storage requirements of sequential machines and Turing machines. Journal of the ACM, 14(3):478–489, July 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Karim:1968:TSG

[Kar68] Abbas I. Abdel Karim. A theorem for the stability of general predictor-corrector methods for the solution of systems of differential equations. Journal of the ACM, 15(4):706–711, October 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Karr:1981:SFT

[Kar81] Michael Karr. Summation in finite terms. Journal of the ACM, 28(2):305–350, April 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Karp:1994:PRR

[Kar94] Richard M. Karp. Probabilistic recurrence relations. Journal of the ACM, 41(6):1136–1150, November 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/195632.html.

Karger:2000:MCN

[Kar00] David R. Karger. Minimum cuts in near-linear time. Journal of the ACM, 47(1):46–76, January 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:

//www.acm.org/pubs/citations/journals/jacm/2000-47-1/
p46-karger/.

Kautz:1971:ACA

[Kau71] William H. Kautz. An augmented content-addressed memory array for implementation with large-scale integration. Journal of the ACM, 18(1):19–33, January 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Knuth:1967:PLA

[KB67] Donald E. Knuth and Richard H. Bigelow. Programming languages for automata. Journal of the ACM, 14(4):615–635, October 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kumar:1981:PAB

[KB81] Sarangan Krishna Kumar and Melvin A. Breuer. Probabilistic aspects of Boolean switching functions via a new transform. Journal of the ACM, 28(3):502–520, July 1981. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kaminski:1989:MCP

[KB89] Michael Kaminski and Nader H. Bshouty. Multiplicative complexity of polynomial multiplication over finite fields. Journal of the ACM, 36(1):150–170, January 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/59306.html.

Karbyshev:2017:PDI

[KBI^{+}17] Aleksandr Karbyshev, Nikolaj Bjørner, Shachar Itzhaky, Noam Rinetzky, and Sharon Shoham. Property-directed inference of universal invariants or proving their absence. Journal of the ACM, 64(1):7:1–7:??, March 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kearns:1998:ENT

[Kea98] Michael Kearns. Efficient noise-tolerant learning from statistical queries. Journal of the ACM, 45(6):983–1006, November 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org:80/pubs/citations/
journals/jacm/1998-45-6/p983-kearns/.

Kedem:1979:CDR

[Ked79] Zvi M. Kedem. Combining dimensionality and rate of growth arguments for establishing lower bounds on the number of multiplications and divisions. Journal of the ACM, 26(3):582–601, July 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Keitel:1956:EMT

[Kei56] Glenn H. Keitel. An extension of Milne’s three-point method. Journal of the ACM, 3(3):212–222, July 1956. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Keller:1961:FAP

[Kel61] Herbert B. Keller. Finite automata, pattern recognition and perceptrons. Journal of the ACM, 8(1):1–20, January 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Keller:1973:PPSa

[Kel73a] Robert M. Keller. Parallel program schemata and maximal parallelism I. Fundamental results. Journal of the ACM, 20(3):514– 537, July 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Keller:1973:PPSb

[Kel73b] Robert M. Keller. Parallel program schemata and maximal parallelism II: Construction of closures. Journal of the ACM, 20 (4):696–710, October 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kernighan:1971:OSP

[Ker71] Brian W. Kernighan. Optimal sequential partitions of graphs. Journal of the ACM, 18(1):34–40, January 1971. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kritzinger:1974:TTC

[KG74] P. S. Kritzinger and J. W. Graham. A theorem in the theory of compromise merge methods. Journal of the ACM, 21(1):157– 160, January 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Klingman:1964:NPA

[KH64] W. R. Klingman and D. M. Himmelblau. Nonlinear programming with the aid of a multiple-gradient summation technique. Journal of the ACM, 11(4):400–415, October 1964. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Khot:2005:HAS

[Kho05] Subhash Khot. Hardness of approximating the shortest vector problem in lattices. Journal of the ACM, 52(5):789–808, September 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kifer:1998:DAQ

[Kif98] Michael Kifer. On the decidability and axiomatization of query finiteness in deductive databases. Journal of the ACM, 45(4): 588–633, July 1998. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org:80/pubs/
citations/journals/jacm/1998-45-4/p588-kifer/.

Katoh:1981:ABS

[KIM81] N. Katoh, T. Ibaraki, and H. Mine. An algorithm for the K best solutions of the resource allocation problem. Journal of the ACM, 28(4):752–764, October 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kirkpatrick:1981:ULB

[Kir81] David G. Kirkpatrick. A unified lower bound for selection and set partitioning problems. Journal of the ACM, 28(1):150–165, January 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kissin:1991:ULB

[Kis91] Gloria Kissin. Upper and lower bounds on switching energy in VLSI. Journal of the ACM, 38(1):222–254, January 1991. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
102790.html.

Krokhin:2003:RAT

[KJJ03] Andrei Krokhin, Peter Jeavons, and Peter Jonsson. Reasoning about temporal relations: The tractable subalgebras of Allen’s interval algebra. Journal of the ACM, 50(5):591–640, September 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kagiwada:1970:IVT

[KK70] H. H. Kagiwada and R. Kalaba. An initial-value theory for Fredholm integral equations with semidegenerate kernels. Journal of the ACM, 17(2):412–419, April 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Katz:1974:PDL

[KK74] Steven Katz and Alan G. Konheim. Priority disciplines in a loop system. Journal of the ACM, 21(2):340–349, April 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kolaitis:2013:RGP

[KK13] Phokion G. Kolaitis and Swastik Kopparty. Random graphs and the parity quantifier. Journal of the ACM, 60(5):37:1–37:34, October 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kempe:2004:SGR

[KKD04] David Kempe, Jon Kleinberg, and Alan Demers. Spatial gossip and resource location protocols. Journal of the ACM, 51(6): 943–967, November 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kaminski:2022:CEP

[KKG^{+}22] Mark Kaminski, Egor V. Kostylev, Bernardo Cuenca Grau, Boris Motik, and Ian Horrocks. The complexity and expressive power of limit Datalog. Journal of the ACM, 69(1):6:1–6:83, February 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3495009.

Kaminski:2018:WPR

[KKMO18] Benjamin Lucien Kaminski, Joost-Pieter Katoen, Christoph Matheja, and Federico Olmedo. Weakest precondition reasoning for expected runtimes of randomized algorithms. Journal of the ACM, 65(5):30:1–30:??, September 2018. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/ft_gateway.cfm?id=3208102.

Kim:2025:FAD

[KKPW25] Eun Jung Kim, Stefan Kratsch, Marcin Pilipczuk, and Magnus Wahlström. Flow-augmentation I: Directed graphs. Journal of the ACM, 72(1):5:1–5:??, February 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3706103.

Karger:1995:RLT

[KKT95] David R. Karger, Philip N. Klein, and Robert E. Tarjan. A randomized linear-time algorithm to find minimum spanning trees. Journal of the ACM, 42(2):321–328, March 1995. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
201022.html.

Kaneko:1970:ARE

[KL70] Toyohisa Kaneko and Bede Liu. Accumulation of round-off error in fast Fourier transforms. Journal of the ACM, 17(4):637–654, October 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kaneko:1973:LRE

[KL73] Toyohisa Kaneko and Bede Liu. On local roundoff errors in floating-point arithmetic. Journal of the ACM, 20(3):391–398, July 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kannan:1986:PTA

[KL86] R. Kannan and R. J. Lipton. Polynomial-time algorithm for the orbit problem. Journal of the ACM, 33(4):808–821, October 1986. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/6496.html.

Korilis:1995:EEN

[KL95] Yannis A. Korilis and Aurel A. Lazar. On the existence of equilibria in noncooperative optimal flow control. Journal of the ACM, 42(3):584–613, May 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/210415.html.

Klawe:1985:TBB

[Kla85] Maria M. Klawe. A tight bound for black and white pebbles on the pyramid. Journal of the ACM, 32(1):218–228, January 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/214115.html.

Kleinfeld:1960:TEV

[Kle60] Erwin Kleinfeld. Techniques for enumerating Veblen-Wedderburn systems. Journal of the ACM, 7(4):330–337, October 1960. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kleinrock:1966:SPM

[Kle66] Leonard Kleinrock. Sequential processing machines (S.P.M.) analyzed with a queueing theory model. Journal of the ACM, 13 (2):179–193, April 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Klee:1967:MCC

[Kle67a] Victor Klee. A method for constructing circuit codes. Journal of the ACM, 14(3):520–528, July 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kleinrock:1967:TSS

[Kle67b] Leonard Kleinrock. Time-shared systems: a theoretical treatment. Journal of the ACM, 14(2):242–261, April 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kleinberg:1999:ASH

[Kle99] Jon M. Kleinberg. Authoritative sources in a hyperlinked environment. Journal of the ACM, 46(5):604–632, September 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/pubs/

articles/journals/jacm/1999-46-5/p604-kleinberg/p604-
kleinberg.pdf;   http://www.acm.org/pubs/citations/journals/
jacm/1999-46-5/p604-kleinberg/.

Koch:1997:WPE

[KLM^{+}97] Richard R. Koch, F. T. Leighton, Bruce M. Maggs, Satish B. Rao, Arnold L. Rosenberg, and Eric J. Schwabe. Workpreserving emulations of fixed-connection networks. Journal of the ACM, 44(1):104–147, January 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kane:2019:NOL

[KLM19] Daniel M. Kane, Shachar Lovett, and Shay Moran. Near-optimal linear decision trees for k-SUM and related problems. Journal of the ACM, 66(3):16:1–16:??, June 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/ft_gateway.cfm?id=3285953.

Klopfenstein:1961:ZNF

[Klo61] R. W. Klopfenstein. Zeros of nonlinear functions. Journal of the ACM, 8(3):366–373, July 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kutylowski:2000:PSC

[KLOW00] Mirosław Kutyłowski, Krzysztof Loryś, Brigitte Oesterdiekhoff, and Rolf Wanka. Periodification scheme: constructing sorting networks with constant period. Journal of the ACM, 47(5): 944–967, 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/pubs/

citations/journals/jacm/2000-47-5/p944-kutylowski/.

Kung:1975:FMS

[KLP75] H. T. Kung, F. Luccio, and F. P. Preparata. On finding the maxima of a set of vectors. Journal of the ACM, 22(4):469– 476, October 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kantor:2015:TWC

[KLPP15] Erez Kantor, Zvi Lotker, Merav Parter, and David Peleg. The topology of wireless communication. Journal of the ACM, 62(5): 37:1–37:??, November 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kaplan:2005:AAA

[KLSS05] Haim Kaplan, Moshe Lewenstein, Nira Shafrir, and Maxim Sviridenko. Approximation algorithms for asymmetric TSP by decomposing directed regular multigraphs. Journal of the ACM, 52(4):602–626, July 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Klug:1982:ERA

[Klu82] Anthony Klug. Equivalence of relational algebra and relational calculus query languages having aggregate functions. Journal of the ACM, 29(3):699–717, July 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Klug:1983:LEI

[Klu83] Anthony Klug. Locking expressions for increased database concurrency. Journal of the ACM, 30(1):36–54, January 1983. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Klug:1988:CQC

[Klu88] Anthony Klug. On conjunctive queries containing inequalities. Journal of the ACM, 35(1):146–160, January 1988. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
42273.html.

Kearns:1994:LBF

[KLV94] Michael Kearns, Ming Li, and Leslie Valiant. Learning Boolean formulas. Journal of the ACM, 41(6):1298–1328, November 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/195656.html.

Kifer:1995:LFO

[KLW95] Michael Kifer, Georg Lausen, and James Wu. Logical foundations of object-oriented and frame-based languages. Journal of the ACM, 42(4):741–843, July 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/210335.html.

Kleinrock:1972:PSQ

[KM72a] L. Kleinrock and R. R. Muntz. Processor sharing queueing models of mixed scheduling disciplines for time shared systems. Journal of the ACM, 19(3):464–482, July 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Konheim:1972:SLS

[KM72b] Alan G. Konheim and Bernd Meister. Service in a loop system. Journal of the ACM, 19(1):92–108, January 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Konheim:1974:WLT

[KM74] Alan G. Konheim and Bernd Meister. Waiting lines and times in a system with polling. Journal of the ACM, 21(3):470–490, July 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kowalski:2020:PCA

[KM20] Dariusz R. Kowalski and Miguel A. Mosteiro. Polynomial counting in anonymous dynamic networks with applications to anonymous dynamic algebraic computations. Journal of the ACM, 67(2):11:1–11:17, May 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3385075.

Karhumaki:2000:ELR

[KMP00] Juhani Karhumäki, Filippo Mignosi, and Wojciech Plandowski. The expressibility of languages and relations by word equations. Journal of the ACM, 47(3):483–505, May 2000. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/citations/journals/jacm/
2000-47-5/p483-karhumaki/.

Kirsch:2012:ERA

[KMP^{+}12] Adam Kirsch, Michael Mitzenmacher, Andrea Pietracaprina, Geppino Pucci, Eli Upfal, and Fabio Vandin. An efficient rigorous approach for identifying statistically significant frequent itemsets. Journal of the ACM, 59(3):12:1–12:22, June 2012. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kumar:2009:UAS

[KMPS09] V. S. Anil Kumar, Madhav V. Marathe, Srinivasan Parthasarathy, and Aravind Srinivasan. A unified approach to scheduling on unrelated parallel machines. Journal of the ACM, 56(5):28:1–28:31, August 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kurtz:1995:ICF

[KMR95] Stuart A. Kurtz, Stephen R. Mahaney, and James S. Royer. The isomorphism conjecture fails relative to a random oracle. Journal of the ACM, 42(2):401–420, March 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/201030.html.

Kopparty:2017:HRL

[KMRZS17] Swastik Kopparty, Or Meir, Noga Ron-Zewi, and Shubhangi Saraf. High-rate locally correctable and locally testable codes with sub-polynomial query complexity. Journal of the ACM, 64 (2):11:1–11:??, June 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Karger:1998:AGC

[KMS98] David Karger, Rajeev Motwani, and Madhu Sudan. Approximate graph coloring by semidefinite programming. Journal of the ACM, 45(2):246–265, March 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:/

/www.acm.org:80/pubs/citations/journals/jacm/1998-45-
2/p246-karger/.

Kunnemann:2025:CVR

[KMS^{+}25] Marvin Künnemann, Filip Mazowiecki, Lia Schütze, Henry Sinclair-Banks, and Karol Wegrzycki. Coverability in VASS revisited: Improving Rackoff’s bounds to obtain conditional optimality. Journal of the ACM, 72(5):33:1–33:27, October 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Karp:1967:OCU

[KMW67] Richard M. Karp, Raymond E. Miller, and Shmuel Winograd. The organization of computations for uniform recurrence relations. Journal of the ACM, 14(3):563–590, July 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kuhn:2016:LCL

[KMW16] Fabian Kuhn, Thomas Moscibroda, and Roger Wattenhofer. Local computation: Lower and upper bounds. Journal of the ACM, 63(2):17:1–17:??, May 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kleinrock:1981:OSA

[KN81] Leonard Kleinrock and Arne Nilsson. On optimal scheduling algorithms for time-shared systems. Journal of the ACM, 28(3): 477–486, July 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kane:2014:SJL

[KN14] Daniel M. Kane and Jelani Nelson. Sparser Johnson– Lindenstrauss transforms. Journal of the ACM, 61(1):4:1–4:??, January 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Knauer:1975:SPC

[Kna75] Bernd Knauer. A simple planarity criterion. Journal of the ACM, 22(2):226–230, April 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Knessl:1993:STD

[Kne93] Charles Knessl. On the sojourn time distribution in a finite capacity processor shared queue. Journal of the ACM, 40(5): 1238–1301, November 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/169736.html.

Khamis:2024:CDP

[KNP^{+}24] Mahmoud Abo Khamis, Hung Q. Ngo, Reinhard Pichler, Dan Suciu, and Yisu Remy Wang. Convergence of datalog over (pre-) semirings. Journal of the ACM, 71(2):8:1–8:??, April 2024. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3643027.

Knuth:1961:MDL

[Knu61] Donald E. Knuth. Minimizing drum latency time. Journal of the ACM, 8(2):119–150, April 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Komargodski:2019:WBV

[KNY19] Ilan Komargodski, Moni Naor, and Eylon Yogev. White-box vs. black-box complexity of search problems: Ramsey and graph property testing. Journal of the ACM, 66(5):34:1–34:??, September 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=
3341106.

Ko:1990:SCR

[Ko90] Ker-I Ko. Separating and collapsing results on the relativized probabilistic polynomial-time hierarchy. Journal of the ACM, 37(2):415–438, April 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/77623.html.

Kobayashi:1974:ADAa

[Kob74a] Hisashi Kobayashi. Application of the diffusion approximation to queueing networks I: Equilibrium queue distributions. Journal of the ACM, 21(2):316–328, April 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kobayashi:1974:ADAb

[Kob74b] Hisashi Kobayashi. Application of the diffusion approximation to queueing networks II: Nonequilibrium distributions and applications to computer modeling. Journal of the ACM, 21(3): 459–469, July 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kobayashi:2013:MCH

[Kob13] Naoki Kobayashi. Model checking higher-order programs. Journal of the ACM, 60(3):20:1–20:62, June 2013. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kochen:1969:AQA

[Koc69] Manfred Kochen. Automatic question-answering of English-like questions about simple diagrams. Journal of the ACM, 16(1):26– 48, January 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kock:2023:WGP

[Koc23] Joachim Kock. Whole-grain Petri nets and processes. Journal of the ACM, 70(1):1:1–1:??, February 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3559103.

Koltun:2004:ATU

[Kol04] Vladlen Koltun. Almost tight upper bounds for vertical decompositions in four dimensions. Journal of the ACM, 51(5):699–730, September 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kong:2024:DTP

[Kon24] Yuqing Kong. Dominantly truthful peer prediction mechanisms with a finite number of tasks. Journal of the ACM, 71(2):9:1– 9:??, April 2024. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL https://dl.acm.org/doi/10.
1145/3638239.

Korolev:1958:CCC

[Kor58] L. N. Korolev. Coding and code compression. Journal of the ACM, 5(4):328–330, October 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Translated by Morris D. Friedman.

Korth:1983:LPD

[Kor83] Henry F. Korth. Locking primitives in a database system. Journal of the ACM, 30(1):55–79, January 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kortelainen:1989:CFC

[Kor89] Juha Kortelainen. The conjecture of Fliess on commutative context-free languages. Journal of the ACM, 36(4):870–872, October 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Korf:2008:LTD

[Kor08] Richard E. Korf. Linear-time disk-based implicit graph search. Journal of the ACM, 55(6):26:1–26:40, December 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kou:1977:LDA

[Kou77] Lawrence T. Kou. On live-dead analysis for global data flow problems. Journal of the ACM, 24(3):473–483, July 1977. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kowalski:1975:PPU

[Kow75] Robert Kowalski. A proof procedure using connection graphs. Journal of the ACM, 22(4):572–595, October 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kowaltowski:1979:DSC

[Kow79] Tomasz Kowaltowski. Data structures and correctness of programs. Journal of the ACM, 26(2):283–301, April 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Katz:2009:ESA

[KOY09] Jonathan Katz, Rafail Ostrovsky, and Moti Yung. Efficient and secure authenticated key exchange using weak passwords. Journal of the ACM, 57(1):3:1–3:39, November 2009. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kaplan:1971:NQI

[KP71] M. A. Kaplan and R. A. Papetti. A note on quadrilateral interpolation. Journal of the ACM, 18(4):576–585, October 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Krishnaswamy:1980:CSS

[KP80] Ramachandran Krishnaswamy and Arthur B. Pyster. On the correctness of semantic-syntax-directed translations. Journal of the ACM, 27(2):338–355, April 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kolaitis:1990:SCA

[KP90] Phokion G. Kolaitis and Christos H. Papadimitriou. Some computational aspects of circumscription. Journal of the ACM, 37 (1):1–14, January 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/78936.html.

Koutsoupias:1995:SC

[KP95] Elias Koutsoupias and Christos H. Papadimitriou. On the kserver conjecture. Journal of the ACM, 42(5):971–983, September 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/210128.html.

Koscielski:1996:CMA

[KP96] Antoni Kościelski and Leszek Pacholski. Complexity of Makanin’s algorithm. Journal of the ACM, 43(4):670–684, July 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
jacm/234543.html.

Kalyanasundaram:2000:SPC

[KP00] Bala Kalyanasundaram and Kirk Prauhs. Speed is as powerful as clairvoyance. Journal of the ACM, 47(4):617–643, July 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/citations/
journals/jacm/2000-47-4/p617-kalyanasundaram/.

Kameda:2002:PDD

[KP02] Hisao Kameda and Odile Pourtallier. Paradoxes in distributed decisions on optimal load balancing for networks of homogeneous computers. Journal of the ACM, 49(3):407–433, May 2002. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kalyanasundaram:2003:MFT

[KP03] Bala Kalyanasundaram and Kirk R. Pruhs. Minimizing flow time nonclairvoyantly. Journal of the ACM, 50(4):551–567, July 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kutten:2015:CUL

[KPP^{+}15] Shay Kutten, Gopal Pandurangan, David Peleg, Peter Robinson, and Amitabh Trehan. On the complexity of universal leader election. Journal of the ACM, 62(1):7:1–7:??, February 2015. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kleinberg:2004:SP

[KPR04] Jon Kleinberg, Christos Papadimitriou, and Prabhakar Raghavan. Segmentation problems. Journal of the ACM, 51(2):263– 280, March 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kiefer:2019:WLD

[KPS19] Sandra Kiefer, Ilia Ponomarenko, and Pascal Schweitzer. The Weisfeiler–Leman dimension of planar graphs is at most 3. Journal of the ACM, 66(6):44:1–44:??, December 2019. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/ft_gateway.cfm?id=3333003.

Kurokawa:2018:FEG

[KPW18] David Kurokawa, Ariel D. Procaccia, and Junxing Wang. Fair enough: Guaranteeing approximate maximin shares. Journal of the ACM, 65(2):8:1–8:??, March 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Konheim:1976:QMF

[KR76] Alan G. Konheim and Martin Reiser. A queueing model with finite waiting room and blocking. Journal of the ACM, 23(2):328– 341, April 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kinariwala:1977:FSA

[KR77] Bharat Kinariwala and A. G. Rao. Flow switching approach to the maximum flow problem: I. Journal of the ACM, 24(4):630– 645, October 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kapoor:1989:OLB

[KR89] Sanjiv Kapoor and Edward M. Reingold. Optimum lopsided binary trees. Journal of the ACM, 36(3):573–590, July 1989. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
65955.html.

Katajainen:1992:ALM

[KR92] Jyrki Katajainen and Timo Raita. An analysis of the longest match and the greedy heuristics in text encoding. Journal of the ACM, 39(2):281–294, April 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/128751.html.

Karloff:1993:RAP

[KR93] Howard J. Karloff and Prabhakar Raghavan. Randomized algorithms and pseudorandom numbers. Journal of the ACM, 40(3): 454–476, July 1993. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/174132.html.

Khardon:1997:LR

[KR97] Roni Khardon and Dan Roth. Learning to reason. Journal of the ACM, 44(5):697–725, September 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:/

/www.acm.org:80/pubs/citations/journals/jacm/1997-44-
5/p697-khardon/.

Kimelfeld:2014:TMS

[KR14] Benny Kimelfeld and Christopher Ré. Transducing Markov sequences. Journal of the ACM, 61(5):32:1–32:??, August 2014.

CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kramer:1962:NSC

[Kra62] Henry P. Kramer. A note on the self-consistency definitions of generalization and inductive inference. Journal of the ACM, 9 (2):280–281, April 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Krider:1964:FAA

[Kri64] Lee Krider. A flow analysis algorithm. Journal of the ACM, 11 (4):429–436, October 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kronmal:1964:EPN

[Kro64] Richard Kronmal. Evaluation of a pseudorandom normal number generator. Journal of the ACM, 11(3):357–363, July 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Krogh:1966:PCM

[Kro66] Fred T. Krogh. Predictor-corrector methods of high order with improved stability characteristics. Journal of the ACM, 13(3): 374–385, July 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Krogh:1967:TIN

[Kro67] Fred T. Krogh. A test for instability in the numerical solution of ordinary differential equations. Journal of the ACM, 14(2):351– 354, April 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Krogh:1973:TSN

[Kro73] Fred T. Krogh. On testing a subroutine for the numerical integration of ordinary differential equations. Journal of the ACM, 20(4):545–562, October 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kalai:2022:HDC

[KRR22] Yael Tauman Kalai, Ran Raz, and Ron D. Rothblum. How to delegate computations: The power of no-signaling proofs. Journal of the ACM, 69(1):1:1–1:82, February 2022. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3456867.

Khandekar:2009:GPU

[KRV09] Rohit Khandekar, Satish Rao, and Umesh Vazirani. Graph partitioning using single commodity flows. Journal of the ACM, 56 (4):19:1–19:15, June 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Klein:1963:CAG

[KS63] Sheldon Klein and Robert F. Simmons. A computational approach to grammatical coding of English words. Journal of the ACM, 10:334–347, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kobayashi:1966:CPD

[KS66] Kojiro Kobayashi and Shigeru Sekiguchi. On the class of predicates decidable by two-way multitape finite automata. Journal of the ACM, 13(2):236–261, April 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kohler:1974:CTC

[KS74] Walter H. Kohler and Kenneth Steiglitz. Characterization and theoretical comparison of branch-and-bound algorithms for permutation problems. Journal of the ACM, 21(1):140–156, January 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kohler:1975:EAG

[KS75] Walter H. Kohler and Kenneth Steiglitz. Exact, approximate, and guaranteed accuracy algorithms for the flow-shop problem n/2/F/\overline{f}. Journal of the ACM, 22(1):106–114, January 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kafura:1978:ADM

[KS78] D. G. Kafura and V. Y. Shen. An algorithm to design the memory configuration of a computer network. Journal of the ACM, 25(3):365–377, July 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kedem:1983:LPE

[KS83] Zvi M. Kedem and Abraham Silberschatz. Locking protocols: From exclusive to shared locks. Journal of the ACM, 30(4):787– 804, October 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Karger:1996:NAM

[KS96] David R. Karger and Clifford Stein. A new approach to the minimum cut problem. Journal of the ACM, 43(4):601–640, July 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
jacm/234534.html.

King:2011:BBB

[KS11] Valerie King and Jared Saia. Breaking the O(n^{2}) bit barrier: Scalable Byzantine agreement with an adaptive adversary. Journal of the ACM, 58(4):18:1–18:24, July 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

King:2016:BAE

[KS16] Valerie King and Jared Saia. Byzantine agreement in expected polynomial time. Journal of the ACM, 63(2):13:1–13:??, May 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Karstadt:2020:MML

[KS20] Elaye Karstadt and Oded Schwartz. Matrix multiplication, a little faster. Journal of the ACM, 67(1):1:1–1:31, February 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3364504.

Karkkainen:2006:LWS

[KSB06] Juha Kärkkäinen, Peter Sanders, and Stefan Burkhardt. Linear work suffix array construction. Journal of the ACM, 53(6): 918–936, November 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kubert:1968:PRF

[KSG68] B. Kubert, J. Szabo, and S. Guilieri. The perspective representation of functions of two variables. Journal of the ACM, 15 (2):193–204, April 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Krause:1975:AST

[KSS75] K. L. Krause, V. Y. Shen, and H. D. Schwetman. Analysis of several task-scheduling algorithms for a model of multiprogramming computer systems. Journal of the ACM, 22(4):522–550, October 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [KSS77].

Krause:1977:EAS

[KSS77] K. L. Krause, V. Y. Shen, and H. D. Schwetman. Errata: “Analysis of several task-scheduling algorithms for a model of multiprogramming computer systems”. Journal of the ACM, 24(3): 527, July 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [KSS75].

Kumar:2010:LTA

[KSS10] Amit Kumar, Yogish Sabharwal, and Sandeep Sen. Linear-time approximation schemes for clustering problems in any dimensions. Journal of the ACM, 57(2):5:1–5:32, January 2010. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kister:1957:EC

[KSU^{+}57] J. Kister, P. Stein, S. Ulam, W. Walden, and M. Wells. Experiments in chess. Journal of the ACM, 4(2):174–177, April 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kleinberg:2019:BEM

[KSU19] Robert Kleinberg, Aleksandrs Slivkins, and Eli Upfal. Bandits and experts in metric spaces. Journal of the ACM, 66(4): 30:1–30:??, August 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3299873.

Kleinberg:2009:TEU

[KSW09] Jon Kleinberg, Aleksandrs Slivkins, and Tom Wexler. Triangulation and embedding using small sets of beacons. Journal of the ACM, 56(6):32:1–32:37, September 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kopparty:2014:HRC

[KSY14] Swastik Kopparty, Shubhangi Saraf, and Sergey Yekhanin. Highrate codes with sublinear-time decoding. Journal of the ACM, 61(5):28:1–28:??, August 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kohfeld:1967:MMM

[KT67] J. J. Kohfeld and G. T. Thompson. Multistep methods with modified predictors and correctors. Journal of the ACM, 14(1): 155–166, January 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kohfeld:1968:MNM

[KT68] J. J. Kohfeld and G. T. Thompson. A modification of Nordsieck’s method using an “off-step” point. Journal of the ACM, 15(3): 390–401, July 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kasami:1969:SAP

[KT69] T. Kasami and K. Torii. A syntax-analysis procedure for unambiguous context-free grammars. Journal of the ACM, 16(3): 423–431, July 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kung:1974:OOO

[KT74] H. T. Kung and J. F. Traub. Optimal order of one-point and multipoint iteration. Journal of the ACM, 21(4):643–651, October 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kung:1978:AAF

[KT78] H. T. Kung and J. F. Traub. All algebraic functions can be computed fast. Journal of the ACM, 25(2):245–260, April 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Knessl:1990:AEL

[KT90] Charles Knessl and Charles Tier. Asymptotic expansions for large closed queueing networks. Journal of the ACM, 37(1):144– 174, January 1990. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/78940.html.

Kaplan:1999:PFR

[KT99] Haim Kaplan and Robert E. Tarjan. Purely functional, real-time deques with catenation. Journal of the ACM, 46(5):577–603, September 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kleinberg:2002:AAC

[KT02] Jon Kleinberg and Éva Tardos. Approximation algorithms for classification problems with pairwise relationships: metric labeling and Markov random fields. Journal of the ACM, 49(5): 616–639, September 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kawarabayashi:2017:CCG

[KT17] Ken-Ichi Kawarabayashi and Mikkel Thorup. Coloring 3colorable graphs with less than n^{1/5} colors. Journal of the ACM, 64(1):4:1–4:??, March 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kawarabayashi:2019:DEC

[KT19] Ken-Ichi Kawarabayashi and Mikkel Thorup. Deterministic edge connectivity in near-linear time. Journal of the ACM, 66(1):4:1– 4:??, January 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kfoury:1994:AMT

[KTU94] A. J. Kfoury, J. Tiuryn, and P. Urzyczyn. An analysis of ML typability. Journal of the ACM, 41(2):368–398, March 1994. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
174659.html.

Kam:1976:GDF

[KU76] John B. Kam and Jeffrey D. Ullman. Global data flow analysis and iterative algorithms. Journal of the ACM, 23(1):158–171, January 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kaplan:1980:SAI

[KU80] Marc A. Kaplan and Jeffrey D. Ullman. A scheme for the automatic inference of variable types. Journal of the ACM, 27(1):

128–145, January 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Karlin:1988:PHE

[KU88] Anna R. Karlin and Eli Upfal. Parallel hashing: an efficient implementation of shared memory. Journal of the ACM, 35(4): 876–892, October 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Koutris:2015:QBD

[KUB^{+}15] Paraschos Koutris, Prasang Upadhyaya, Magdalena Balazinska, Bill Howe, and Dan Suciu. Query-based data pricing. Journal of the ACM, 62(5):43:1–43:??, November 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kung:1976:NAL

[Kun76] H. T. Kung. New algorithms and lower bounds for the parallel evaluation of certain rational expressions and recurrences. Journal of the ACM, 23(2):252–261, April 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kurtzberg:1962:AMA

[Kur62] Jerome M. Kurtzberg. On approximation methods for the assignment problem. Journal of the ACM, 9(4):419–439, October 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kearns:1994:CLL

[KV94a] Michael Kearns and Leslie Valiant. Cryptographic limitations on learning Boolean formulae and finite automata. Journal of the ACM, 41(1):67–95, January 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/174647.html.

Khuller:1994:BAG

[KV94b] Samir Khuller and Uzi Vishkin. Biconnectivity approximations and graph carvings. Journal of the ACM, 41(2):214–235, March 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/174654.html.

Kolaitis:2004:F

[KV04] Phokion Kolaitis and Victor Vianu. Foreword. Journal of the ACM, 51(1):1, January 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Khot:2015:UGC

[KV15] Subhash A. Khot and Nisheeth K. Vishnoi. The unique games conjecture, integrality gap for cut problems and embeddability of negative-type metrics into l_{1}. Journal of the ACM, 62(1): 8:1–8:??, February 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kannan:2004:CGB

[KVV04] Ravi Kannan, Santosh Vempala, and Adrian Vetta. On clusterings: Good, bad and spectral. Journal of the ACM, 51(3): 497–515, May 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kupferman:2000:ATA

[KVW00] Orna Kupferman, Moshe Y. Vardi, and Pierre Wolper. An automata-theoretic approach to branching-time model checking. Journal of the ACM, 47(2):312–360, March 2000. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/citations/journals/jacm/
2000-47-2/p312-kupferman/.

Krishnamoorthi:1966:TSO

[KW66] B. Krishnamoorthi and Roger C. Wood. Time-shared operations with both interarrival and service times exponential. Journal of the ACM, 13(3):317–338, July 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Karp:1985:FPA

[KW85] Richard M. Karp and Avi Wigderson. A fast parallel algorithm for the maximal independent set problem. Journal of the ACM, 32(4):762–773, October 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/4226.html. This impor-

tant paper showed that the maximal independent set problem for graphs can be solved in polylogarithmic time using a polynomial number of processes on a PRAM in which concurrent reads and writes are disallowed. They derive their algorithm from a randomized one using a technique that has become known as derandomization via k-wise independence.

Klein:1989:PLB

[KW89] Rolf Klein and Derick Wood. On the path length of binary trees. Journal of the ACM, 36(2):280–289, April 1989. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
62048.html.

Kratsch:2020:RSI

[KW20] Stefan Kratsch and Magnus WahlströM. Representative sets and irrelevant vertices: New tools for kernelization. Journal of the ACM, 67(3):16:1–16:50, June 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3390887.

Krevner:1983:ITS

[KY83] Yael Krevner and Amiram Yehudai. An iteration theorem for simple precedence languages. Journal of the ACM, 30(4):820– 833, October 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Karp:1993:RPA

[KZ93] Richard M. Karp and Yanjun Zhang. Randomized parallel algorithms for backtrack search and branch-and-bound computation. Journal of the ACM, 40(3):765–789, July 1993. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
174145.html.

Karp:1998:PEG

[KZ98] Richard M. Karp and Yangun Zhang. On parallel evaluation of game trees. Journal of the ACM, 45(6):1050–1075, November 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org:80/pubs/citations/
journals/jacm/1998-45-6/p1050-karp/.

Kolmogorov:2013:CCV

[KZ13] Vladimir Kolmogorov and Stanislav Zivný. The complexity of conservative valued CSPs. Journal of the ACM, 60(2):10:1–10:38, April 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Korf:2005:FS

[KZTH05] Richard E. Korf, Weixiong Zhang, Ignacio Thayer, and Heath Hohwald. Frontier search. Journal of the ACM, 52(5):715–748, September 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

LaBudde:1964:TNC

[La 64] C. Donald La Budde. Two new classes of algorithms for finding the eigenvalues and eigenvectors of real symmetric matrices. Journal of the ACM, 11(1):53–58, January 1964. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Laasonen:1958:SPD

[Laa58a] Pentti Laasonen. On the solution of Poisson’s difference equation. Journal of the ACM, 5(4):370–382, October 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Laasonen:1958:TED

[Laa58b] Pentti Laasonen. On the truncation error of discrete approximations to the solutions of Dirichlet problems in a domain with corners. Journal of the ACM, 5(1):32–38, January 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ladner:1975:SPT

[Lad75] Richard E. Ladner. On the structure of polynomial time reducibility. Journal of the ACM, 22(1):155–171, January 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Laird:2013:GSP

[Lai13] J. Laird. Game semantics for a polymorphic programming language. Journal of the ACM, 60(4):29:1–29:27, August 2013. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lam:1982:DSG

[Lam82] Simon S. Lam. Dynamic scaling and growth behavior of queuing network normalization constants. Journal of the ACM, 29 (2):492–513, April 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lamport:1983:WBG

[Lam83] L. Lamport. The weak Byzantine generals problem. Journal of the ACM, 30(3):668–676, July 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lamport:1986:MEPa

[Lam86a] Leslie Lamport. The mutual exclusion problem: Part I — the theory of interprocess communication. Journal of the ACM, 33 (2):313–326, April 1986. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/5384.html.

Lamport:1986:MEPb

[Lam86b] Leslie Lamport. The mutual exclusion problem: Part II — statement and solutions. Journal of the ACM, 33(2):327–348, April 1986. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/5385.html.

Lampson:2003:GCU

[Lam03] Butler Lampson. Getting computers to understand. Journal of the ACM, 50(1):70–72, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lance:1959:SAT

[Lan59] G. N. Lance. Solution of algebraic and transcendental equations on an automatic digital computer. Journal of the ACM, 6(1):97– 101, January 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

LaPaugh:1993:RDH

[LaP93] Andrea S. LaPaugh. Recontamination does not help to search a graph. Journal of the ACM, 40(2):224–245, April 1993. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
151263.html.

Larson:1976:ECC

[Lar76] Richard G. Larson. Efficiency of computation of Cayley tables of 2-groups. Journal of the ACM, 23(2):235–241, April 1976. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Larson:1983:AUH

[Lar83] Per-Åke Larson. Analysis of uniform hashing. Journal of the ACM, 30(4):805–819, October 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Latouche:1981:AAM

[Lat81] Guy Latouche. Algorithmic analysis of a multiprogrammingmultiprocessor computer system. Journal of the ACM, 28(4): 662–679, October 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lawler:1964:AMB

[Law64] Eugene L. Lawler. An approach to multilevel Boolean minimization. Journal of the ACM, 11(3):283–294, July 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lazar:1984:OFC

[Laz84] Aurel A. Lazar. Optimal flow control of an M/M/m queue. Journal of the ACM, 31(1):86–98, January 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lueker:1979:LTA

[LB79] George S. Lueker and Kellogg S. Booth. A linear time algorithm for deciding interval graph isomorphism. Journal of the ACM, 26(2):183–195, April 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lewis:1967:SDS

[LBB67] P. A. W. Lewis, P. B. Baxendale, and J. L. Bennett. Statistical discrimination of the synonymy/antonymy relationship between words. Journal of the ACM, 14(1):20–44, January 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Leyton-Brown:2009:EHM

[LBNS09] Kevin Leyton-Brown, Eugene Nudelman, and Yoav Shoham. Empirical hardness models: Methodology and a case study on combinatorial auctions. Journal of the ACM, 56(4):22:1–22:52, June 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lebedev:1956:HSE

[Leb56] S. A. Lebedev. The high-speed electronic calculating machine of the Academy of Sciences of the U.S.S.R. Journal of the ACM, 3(3):129–133, July 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Translated by Curtis D. Benster.

Lipton:1976:STH

[LED76] R. J. Lipton, S. C. Eisenstat, and R. A. DeMillo. Space and time hierarchies for classes of control structures and data structures. Journal of the ACM, 23(4):720–732, October 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lee:1961:CAM

[Lee61] C. Y. Lee. Categorizing automata by W -machine programs. Journal of the ACM, 8(3):384–399, July 1961. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lee:1972:FLR

[Lee72] Richard C. T. Lee. Fuzzy logic and the resolution principle. Journal of the ACM, 19(1):109–119, January 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lee:1980:TDV

[Lee80] D. T. Lee. Two-dimensional Voronoi diagrams in the L_{p}-metric. Journal of the ACM, 27(4):604–618, October 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lee:2002:FCF

[Lee02] Lillian Lee. Fast context-free grammar parsing requires fast Boolean matrix multiplication. Journal of the ACM, 49(1):1– 15, January 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lehmer:1957:SCR

[Leh57] D. H. Lehmer. Sorting cards with respect to a modulus. Journal of the ACM, 4(1):41–46, January 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lehmer:1961:MMS

[Leh61] D. H. Lehmer. A machine method for solving polynomial equations. Journal of the ACM, 8(2):151–162, April 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lehot:1974:OAD

[Leh74] Philippe G. H. Lehot. An optimal algorithm to detect a line graph and output its root graph. Journal of the ACM, 21(4): 569–575, October 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Leiner:1954:SSD

[Lei54] Alan L. Leiner. System specifications for the DYSEAC. Journal of the ACM, 1(2):57–81, April 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Leichner:1957:DCC

[Lei57] Gene H. Leichner. Designing computer circuits with a computer. Journal of the ACM, 4(2):143–147, April 1957. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Leighton:1991:LE

[Lei91] Tom Leighton. Letter from the editor. Journal of the ACM, 38 (3):515, July 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Leighton:2003:J

[Lei03] Tom Leighton. JACM 1991–1997. Journal of the ACM, 50(1): 19, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lengauer:1989:HPT

[Len89] Thomas Lengauer. Hierarchical planarity testing algorithms. Journal of the ACM, 36(3):474–509, July 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
65952.html.

Lesh:1958:MSD

[Les58] F. Lesh. Methods of simulating a differential analyzer on a digital computer. Journal of the ACM, 5(3):281–288, July 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Leung:1993:ESS

[Leu93] Kin K. Leung. An execution/sleep scheduling policy for serving an additional job in priority queueing systems. Journal of the ACM, 40(2):394–417, April 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/151269.html.

Levin:2013:FI

[Lev13] Leonid A. Levin. Forbidden information. Journal of the ACM, 60(2):9:1–9:9, April 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lewis:1978:RSC

[Lew78] Harry R. Lewis. Renaming a set of clauses as a Horn set. Journal of the ACM, 25(1):134–135, January 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ladner:1980:PPC

[LF80] Richard E. Ladner and Michael J. Fischer. Parallel prefix computation. Journal of the ACM, 27(4):831–838, October 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Leivant:1988:MRF

[LF88] Daniel Leivant and Tim Fernando. Meager and replete failures of relative completeness. Journal of the ACM, 35(4):953–964, October 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/63139.html.

Lund:1992:AMI

[LFKN92] Carsten Lund, Lance Fortnow, Howard Karloff, and Noam Nisan. Algebraic methods for interactive proof systems. Journal of the ACM, 39(4):859–868, October 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/146605.html.

Lee:2014:MSP

[LGT14] James R. Lee, Shayan Oveis Gharan, and Luca Trevisan. Multiway spectral partitioning and higher-order Cheeger inequalities. Journal of the ACM, 61(6):37:1–37:??, November 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Larmore:1990:FAO

[LH90] Lawrence L. Larmore and Daniel S. Hirschberg. A fast algorithm for optimal length-limited Huffman codes. Journal of the ACM, 37(3):464–473, July 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/79150.html.

Li:1959:ODC

[Li59] Shu-T’ien Li. Origin and development of the Chinese abacus. Journal of the ACM, 6(1):102–110, January 1959. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Liberatore:2001:MRR

[Lib01] Paolo Liberatore. Monotonic reductions, representative equivalence, and compilation of intractable problems. Journal of the ACM, 48(6):1091–1125, November 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lichter:2023:SRL

[Lic23] Moritz Lichter. Separating rank logic from polynomial time. Journal of the ACM, 70(2):14:1–14:??, April 2023. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3572918.

Lieblein:1959:GAV

[Lie59] Julius Lieblein. A general analysis of variance scheme applicable to a computer with a very large memory. Journal of the ACM, 6 (4):469–475, October 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lien:1982:EDM

[Lie82] Y. Edmund Lien. On the equivalence of database models. Journal of the ACM, 29(2):333–362, April 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Linz:1969:LMM

[Lin69] Peter Linz. Linear multistep methods for Volterra integrodifferential equations. Journal of the ACM, 16(2):295–301, April 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lin:2018:PCB

[Lin18] Bingkai Lin. The parameterized complexity of the k-biclique problem. Journal of the ACM, 65(5):34:1–34:??, September 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=
3212622.

Lions:1971:SRC

[Lio71] John Lions. Some results concerning the reduction of binary matrices. Journal of the ACM, 18(3):424–430, July 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lipton:1957:TPT

[Lip57] S. Lipton. Two programming techniques for one-plus-one address computers. Journal of the ACM, 4(3):274–278, July 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lipski:1981:DII

[Lip81] Witold Lipski, Jr. On databases with incomplete information. Journal of the ACM, 28(1):41–70, January 1981. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Liu:1963:SVA

[Liu63] C. N. Liu. A state variable assignment method for asynchronous sequential switching circuits. Journal of the ACM, 10:209–216, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Liu:1969:LFP

[Liu69] C. L. Liu. Lattice functions, pair algebras, and finite-state machines. Journal of the ACM, 16(3):442–454, July 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lively:1960:LER

[Liv60] J. A. Lively. Letter to the editor: Remarks on “Amphisbaenic sorting”. Journal of the ACM, 7(2):187, April 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Nag59, Nag61, Goe61].

Lord:1983:SLA

[LKK83] R. E. Lord, J. S. Kowalik, and S. P. Kumar. Solving linear algebraic equations on an MIMD computer. Journal of the ACM, 30(1):103–117, January 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Liu:1973:SAM

[LL73] C. L. Liu and James W. Layland. Scheduling algorithms for multiprogramming in a hard-real-time environment. Journal of the ACM, 20(1):46–61, January 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lawler:1978:PSU

[LL78] E. L. Lawler and J. Labetoulle. On preemptive scheduling of unrelated parallel processors by linear programming. Journal of the ACM, 25(4):612–619, October 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lee:1985:SLP

[LL85] C. C. Lee and D. T. Lee. A simple on-line bin-packing algorithm. Journal of the ACM, 32(3):562–572, July 1985. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
3833.html.

Lloyd:1982:CPS

[Llo82] Errol L. Lloyd. Critical path scheduling with resource and processor constraints. Journal of the ACM, 29(3):781–811, July 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lindell:2006:CAB

[LLR06] Yehuda Lindell, Anna Lysyanskaya, and Tal Rabin. On the composition of authenticated Byzantine Agreement. Journal of the ACM, 53(6):881–917, November 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lenzen:2010:TBC

[LLW10] Christoph Lenzen, Thomas Locher, and Roger Wattenhofer. Tight bounds for clock synchronization. Journal of the ACM, 57(2):8:1–8:42, January 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lubachevsky:1986:CAA

[LM86] Boris Lubachevsky and Debasis Mitra. A chaotic asynchronous algorithm for computing the fixed point of a nonnegative matrix of unit spectral radius. Journal of the ACM, 33(1):130– 150, January 1986. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/4801.html.

Ladkin:1994:BCP

[LM94a] Peter B. Ladkin and Roger D. Maddux. On binary constraint problems. Journal of the ACM, 41(3):435–469, May 1994. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
176585.html.

Lui:1994:CBS

[LM94b] John C. S. Lui and Richard R. Muntz. Computing bounds on steady state availability of repairable computer systems. Journal of the ACM, 41(4):676–707, July 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/179848.html.

Liu:2023:RLG

[LM23] Allen Liu and Ankur Moitra. Robustly learning general mixtures of Gaussians. Journal of the ACM, 70(3):21:1–21:??, June 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3583680.

Linial:1993:CDC

[LMN93] Nathan Linial, Yishay Mansour, and Noam Nisan. Constant depth circuits, Fourier transform, and learnability. Journal of the ACM, 40(3):607–620, July 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/174138.html.

Leonardi:2003:BDQ

[LMNM03] Emilio Leonardi, Marco Mellia, Fabio Neri, and Marco Ajmone Marsan. Bounds on delays and queue lengths in input-queued cell switches. Journal of the ACM, 50(4):520–550, July 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lamport:1985:SCP

[LMS85] Leslie Lamport and P. M. Melliar-Smith. Synchronizing clocks in the presence of faults. Journal of the ACM, 32(1):52–78, January 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/2457.html.

Libkin:2016:QGD

[LMV16] Leonid Libkin, Wim Martens, and Domagoj Vrgoc. Querying graphs with data. Journal of the ACM, 63(2):14:1–14:??, May 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Li:2002:CSS

[LMW02] Ming Li, Bin Ma, and Lusheng Wang. On the closest string and substring problems. Journal of the ACM, 49(2):157–171, March 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Li:2005:BPC

[LMW05] Ninghui Li, John C. Mitchell, and William H. Winsborough. Beyond proof-of-compliance: security analysis in trust management. Journal of the ACM, 52(3):474–514, May 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Li:2025:VCP

[LNP^{+}25] Jason Li, Danupon Nanongkai, Debmalya Panigrahi, Thatchaphol Saranurak, and Sorrachai Yingchareonthawornchai. Vertex connectivity in poly-logarithmic max-flows. Journal of the ACM, 72(4):28:1–28:??, August 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Leiner:1959:PNM

[LNSW59] A. L. Leiner, W. A. Notz, J. L. Smith, and A. Weinberger. PI-LOT — a new multiple computer system. Journal of the ACM, 6(3):313–335, July 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Liu:1997:EBA

[LNT97] Zhen Liu, Philippe Nain, and Don Towsley. Exponential bounds with applications to call admission. Journal of the ACM, 44 (3):366–394, May 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/citations/journals/jacm/1997-44-3/p366-liu/.

Lagarias:1985:SLD

[LO85] J. C. Lagarias and A. M. Odlyzko. Solving low-density subset sum problems. Journal of the ACM, 32(1):229–246, January 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/2461.html.

Logrippo:1978:REM

[Log78] Luigi Logrippo. Renamings and economy of memory in program schemata. Journal of the ACM, 25(1):10–22, January 1978. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Logrippo:1979:RMP

[Log79] Luigi Logrippo. Renamings, maximal parallelism, and spacetime tradeoff in program schemata. Journal of the ACM, 26(4): 819–833, October 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Loizou:1969:NJC

[Loi69] Georghios Loizou. Nonnormality and Jordan condition numbers of matrices. Journal of the ACM, 16(4):580–584, October 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lombardi:1961:SHF

[Lom61] Lionello Lombardi. System handling of functional operators. Journal of the ACM, 8(2):168–185, April 1961. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lombardi:1962:MSN

[Lom62] Lionello Lombardi. Mathematical structure of nonarithmetic data processing procedures. Journal of the ACM, 9(1):136–159, January 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lomet:1973:FTD

[Lom73] David Bruce Lomet. A formalization of transition diagram systems. Journal of the ACM, 20(2):235–257, April 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lehmann:2002:TRA

[LOS02] Daniel Lehmann, Liadan Ita O’Callaghan, and Yoav Shoham. Truth revelation in approximately efficient combinatorial auctions. Journal of the ACM, 49(5):577–602, September 2002. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lotkin:1956:NMM

[Lot56] Mark Lotkin. A note on the midpoint method of integration. Journal of the ACM, 3(3):208–211, July 1956. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Loveland:1968:MTP

[Lov68] Donald W. Loveland. Mechanical theorem-proving by model elimination. Journal of the ACM, 15(2):236–251, April 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Lov69b].

Loveland:1969:SFM

[Lov69a] D. W. Loveland. A simplified format for the model elimination theorem-proving procedure. Journal of the ACM, 16(3):349–363, July 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Loveland:1969:EMT

[Lov69b] Donald W. Loveland. Errata: “Mechanical theorem-proving by model elimination”. Journal of the ACM, 16(4):646, October 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Lov68].

Loveland:1972:UVS

[Lov72] D. W. Loveland. A unifying view of some linear Herbrand procedures. Journal of the ACM, 19(2):366–384, April 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Loveman:1977:PIS

[Lov77] David B. Loveman. Program improvement by source-to-source transformation. Journal of the ACM, 24(1):121–145, January 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lovett:2016:CBR

[Lov16] Shachar Lovett. Communication is bounded by root of rank. Journal of the ACM, 63(1):1:1–1:??, March 2016. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lowe:1968:IDB

[Low68] Thomas C. Lowe. The influence of data base characteristics and usage on direct access file organization. Journal of the ACM, 15 (4):535–548, October 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lewis:1973:GFS

[LP73] T. G. Lewis and W. G. Payne. Generalized feedback shift register pseudorandom number algorithm. Journal of the ACM, 20(3): 456–468, July 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See important errata, and algorithm and code improvements, in [Fus90].

Lee:1979:OAF

[LP79] D. T. Lee and F. P. Preparata. An optimal algorithm for finding the kernel of a polygon. Journal of the ACM, 26(3):415–421, July 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Losch:2014:DSN

[LP14] Steffen Lösch and Andrew M. Pitts. Denotational semantics with nominal Scott domains. Journal of the ACM, 61(4):27:1–27:??, July 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lin:2015:CRN

[LP15] Huijia Lin and Rafael Pass. Constant-round nonmalleable commitments from any one-way function. Journal of the ACM, 62 (1):5:1–5:??, February 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Li:2025:DMC

[LP25] Jason Li and Debmalya Panigrahi. Deterministic minimum cut in poly-logarithmic maximum flows. Journal of the ACM, 72(4): 29:1–29:??, August 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lyubashevsky:2013:ILL

[LPR13] Vadim Lyubashevsky, Chris Peikert, and Oded Regev. On ideal lattices and learning with errors over rings. Journal of the ACM, 60(6):43:1–43:??, November 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lotker:2015:IDA

[LPSP15] Zvi Lotker, Boaz Patt-Shamir, and Seth Pettie. Improved distributed approximate matching. Journal of the ACM, 62(5):

38:1–38:??, November 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Low:2002:UTV

[LPW02] Steven H. Low, Larry L. Peterson, and Limin Wang. Understanding TCP Vegas: a duality model. Journal of the ACM, 49 (2):207–235, March 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lin:1965:CST

[LR65] Shen Lin and Tibor Rado. Computer studies of Turing machine problems. Journal of the ACM, 12(2):196–212, April 1965. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Landweber:1972:RPA

[LR72] L. H. Landweber and E. L. Robertson. Recursive properties of abstract complexity classes. Journal of the ACM, 19(2):296–308, April 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lin:1977:DPG

[LR77] Benjamin W. Y. Lin and Ronald L. Rardin. Development of a parametric generating procedure for integer programming test problems. Journal of the ACM, 24(3):465–472, July 1977. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Landweber:1978:PCF

[LR78] L. H. Landweber and E. L. Robertson. Properties of conflict-free and persistent Petri nets. Journal of the ACM, 25(2):352–364, April 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lemme:1979:SPA

[LR79] James M. Lemme and John R. Rice. Speedup in parallel algorithms for adaptive quadrature. Journal of the ACM, 26(1):65– 71, January 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Loveland:1981:DRG

[LR81] D. W. Loveland and C. R. Reddy. Deleting repeated goals in the problem reduction format. Journal of the ACM, 28(4):646– 661, October 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Leighton:1999:MMF

[LR99] Tom Leighton and Satish Rao. Multicommodity max-flow mincut theorems and their use in designing approximation algorithms. Journal of the ACM, 46(6):787–832, November 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL http://www.acm.org/pubs/articles/journals/
jacm/1999-46-6/p787-leighton/p787-leighton.pdf.

Lenzen:2019:SSB

[LR19] Christoph Lenzen and Joel Rybicki. Self-stabilising Byzantine clock synchronisation is almost as easy as consensus. Journal of the ACM, 66(5):32:1–32:??, September 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/ft_gateway.cfm?id=3339471.

Lipton:1980:EHS

[LRY80] Richard J. Lipton, Arnold L. Rosenberg, and Andrew C. Yao. External hashing schemes for collections of data structures. Journal of the ACM, 27(1):81–95, January 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lewis:1968:SDT

[LS68] P. M. Lewis II and R. E. Stearns. Syntax-directed transduction. Journal of the ACM, 15(3):465–488, July 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lewis:1971:CQM

[LS71] P. A. W. Lewis and G. S. Shedler. A cyclic-queue model of system overhead in multiprogrammed computer systems. Journal of the ACM, 18(1):199–220, January 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lipton:1977:LTA

[LS77] R. J. Lipton and L. Snyder. A linear time algorithm for deciding subject security. Journal of the ACM, 24(3):455–464, July 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lichtenstein:1980:GPS

[LS80] David Lichtenstein and Michael Sipser. GO is polynomial-space hard. Journal of the ACM, 27(2):393–401, April 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Leong:1981:NRT

[LS81a] Benton L. Leong and Joel I. Seiferas. New real-time simulations of multihead tape units. Journal of the ACM, 28(1):166–180, January 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lieberherr:1981:CPS

[LS81b] K. J. Lieberherr and E. Specker. Complexity of partial satisfaction. Journal of the ACM, 28(2):411–421, April 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Long:1986:RCC

[LS86] Timothy J. Long and Alan L. Selman. Relativizing complexity classes with sparse oracles. Journal of the ACM, 33(3):618– 627, July 1986. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/5938.html.

Lee:1987:ORS

[LS87] Yann-Hang Lee and Kang G. Shin. Optimal reconfiguration strategy for a degradable multimodule computing system. Journal of the ACM, 34(2):326–348, April 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/23016.html.

Lee:1988:ODU

[LS88] Yann-Hang Lee and Kang G. Shin. Optimal design and use of retry in fault-tolerant computer systems. Journal of the ACM, 35(1):45–69, January 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/42269.html.

Lin:1995:PCT

[LS95] Fangzhen Lin and Yoav Shoham. Provably correct theories of action. Journal of the ACM, 42(2):293–320, March 1995. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
201021.html.

Lavi:2011:TNO

[LS11] Ron Lavi and Chaitanya Swamy. Truthful and near-optimal mechanism design via linear programming. Journal of the ACM, 58(6):25:1–25:24, December 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Los:2023:BAC

[LS23] Dimitrios Los and Thomas Sauerwald. Balanced allocations with the choice of noise. Journal of the ACM, 70(6):37:1–37:??, December 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3625386.

Lichter:2024:CPT

[LS24] Moritz Lichter and Pascal Schweitzer. Choiceless polynomial time with witnessed symmetric choice. Journal of the ACM, 71(2):7:1–7:??, April 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3648104.

Lamport:1982:BGP

[LSP82] Leslie Lamport, Robert Shostak, and Marshall Pease. The Byzantine generals problem. ACM Transactions on Programming Languages and Systems, 4(3):382–401, July 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Limaye:2025:SLB

[LST25] Nutan Limaye, Srikanth Srinivasan, and Sébastien Tavenas. Superpolynomial lower bounds against low-depth algebraic circuits. Journal of the ACM, 72(4):26:1–26:??, August 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lengauer:1982:ATB

[LT82] Thomas Lengauer and Robert E. Tarjan. Asymptotically tight bounds on time-space trade-offs in a pebble game. Journal of the ACM, 29(4):1087–1130, October 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Luo:1993:CCD

[LT93] Zhi-Quan Luo and John N. Tsitsiklis. On the communication complexity of distributed algebraic computation. Journal of the ACM, 40(5):1019–1047, November 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/174149.html.

Labrada:2022:GSF

[LTT22] Elizabeth Labrada, Matías Toro, and Éric Tanter. Gradual system F. Journal of the ACM, 69(5):38:1–38:??, October 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3555986.

Li:1996:HSC

[LTV96] Ming Li, John Tromp, and Paul M. B. Vitányi. How to share concurrent wait-free variables. Journal of the ACM, 43(4):723– 746, July 1996. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/jacm/234556.html.

Lueker:2009:IBA

[Lue09] George S. Lueker. Improved bounds on the average length of longest common subsequences. Journal of the ACM, 56(3):17:1– 17:38, May 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Luke:1957:RAE

[Luk57] Yudell L. Luke. Rational approximations to the exponential function. Journal of the ACM, 4(1):24–29, January 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lehmer:1957:CNS

[LV57] Emma Lehmer and H. S. Vandiver. On the computation of the number of solutions of certain trinomial congruences. Journal of the ACM, 4(4):505–510, October 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Libkin:2009:IPS

[LV09] Leonid Libkin and Victor Vianu. Introduction to PODS 2007 special section. Journal of the ACM, 56(6):29:1, September 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lykouris:2021:CCM

[LV21] Thodoris Lykouris and Sergei Vassilvitskii. Competitive caching with machine learned advice. Journal of the ACM, 68(4):24:1– 24:25, August 2021. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL https://dl.acm.org/doi/10.
1145/3447579.

Loberman:1957:FPC

[LW57] H. Loberman and A. Weinberger. Formal procedures for connecting terminals with a minimum total wire length. Journal of the ACM, 4(4):428–437, October 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lee-Whiting:1963:EFC

[LW63a] G. E. Lee-Whiting. Erratum: “Formulas for computing incomplete elliptic integrals of the first and second kinds”. Journal of the ACM, 10:412, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [LW63b].

Lee-Whiting:1963:FCI

[LW63b] G. E. Lee-Whiting. Formulas for computing incomplete elliptic integrals of the first and second kinds. Journal of the ACM, 10 (2):126–130, April 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [LW63a].

Lowrance:1975:ESS

[LW75] Roy Lowrance and Robert A. Wagner. An extension of the stringto-string correction problem. Journal of the ACM, 22(2):177– 183, April 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lengauer:1993:EDP

[LW93] Thomas Lengauer and Egon Wanke. Efficient decision procedures for graph properties on centext-free graph languages. Journal of the ACM, 40(2):368–393, April 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/151268.html.

Li:2008:BSD

[LW08] Ninghui Li and Qihua Wang. Beyond separation of duty: an algebra for specifying high-level security policies. Journal of the ACM, 55(3):12:1–12:56, July 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Logan:1955:ASE

[LWS55] Benjamin F. Logan, George R. Welti, and George C. Sponsler. Analogue study of electron trajectories. Journal of the ACM, 2 (1):28–41, January 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lovett:2021:DLC

[LWZ21] Shachar Lovett, Kewen Wu, and Jiapeng Zhang. Decision list compression by mild random restrictions. Journal of the ACM, 68(6):45:1–45:17, December 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/doi/10.1145/3485007.

Leiner:1961:SGP

[LY61] A. L. Leiner and W. W. Youden. A system for generating “pronounceable” names using a computer. Journal of the ACM, 8 (1):97–103, January 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Li:1989:NLB

[LY89] Ming Li and Yaacov Yesha. New lower bounds for parallel computation. Journal of the ACM, 36(3):671–680, July 1989. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
65959.html.

Lund:1994:HAM

[LY94] Carsten Lund and Mihalis Yannakakis. On the hardness of approximating minimization problems. Journal of the ACM, 41(5): 960–981, September 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Liu:2025:PSS

[LY25] Hongyang Liu and Yitong Yin. Parallelize single-site dynamics up to Dobrushin criterion. Journal of the ACM, 72(1):7:1–7:??, February 2025. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL https://dl.acm.org/doi/10.
1145/3708558.

Lyness:1969:NAS

[Lyn69] J. N. Lyness. Notes on the adaptive Simpson quadrature routine. Journal of the ACM, 16(3):483–495, July 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lynch:1975:RCS

[Lyn75] Nancy Lynch. On reducibility to complex or sparse sets. Journal of the ACM, 22(3):341–345, July 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lynch:1977:LSR

[Lyn77] Nancy Lynch. Log space recognition and translation of parenthesis languages. Journal of the ACM, 24(4):583–590, October 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lipton:1977:WPS

[LZ77] Richard J. Lipton and Yechezkel Zalcstein. Word problems solvable in logspace. Journal of the ACM, 24(3):522–526, July 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lasserre:2001:LTA

[LZ01] Jean B. Lasserre and Eduardo S. Zeron. A Laplace transform algorithm for the volume of a convex polytope. Journal of the ACM, 48(6):1126–1140, November 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lim:2025:QCC

[LZ25] Donghyun Lim and Martin Ziegler. Quantitative coding and complexity theory of continuous data: Part I: Motivation, definition, consequences. Journal of the ACM, 72(1):4:1–4:??, February 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3705609.

Liu:2025:QWP

[LZBY25] Junyi Liu, Li Zhou, Gilles Barthe, and Mingsheng Ying. Quantum weakest preconditions for reasoning about expected runtimes of quantum programs. Journal of the ACM, 72(3):20:1– 20:??, June 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Macon:1955:CEH

[Mac55] Nathaniel Macon. On the computation of exponential and hyperbolic functions using continued fractions. Journal of the ACM, 2 (4):262–266, October 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Macon:1957:CLP

[Mac57] Nathaniel Macon. Condensation and look-up procedures for double entry tables. Journal of the ACM, 4(4):456–458, October 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

MacLaren:1966:ISR

[Mac66] M. Donald MacLaren. Internal sorting by radix plus sifting. Journal of the ACM, 13(3):404–411, July 1966. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Maddison:1966:PNL

[Mad66] R. N. Maddison. A procedure for nonlinear least squares refinement in adverse practical conditions. Journal of the ACM, 13 (1):124–134, January 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Maehly:1960:MFR

[Mae60] Hans J. Maehly. Methods for fitting rational approximations, Part I: Telescoping procedures for continued fractions. Journal of the ACM, 7(2):150–162, April 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Maehly:1963:MFR

[Mae63] Hans J. Maehly. Methods for fitting rational approximations, parts II and III. Journal of the ACM, 10:257–278, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Maekawa:1977:QMC

[Mae77] Mamoru Maekawa. Queueing models for computer systems connected by a communication line. Journal of the ACM, 24(4): 566–582, October 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Maier:1978:CSP

[Mai78] David Maier. The complexity of some problems on subsequences and supersequences. Journal of the ACM, 25(2):322–336, April 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Maier:1980:MCR

[Mai80] David Maier. Minimum covers in the relational database model. Journal of the ACM, 27(4):664–674, October 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mamelak:1966:PCL

[Mam66] J. S. Mamelak. The placement of computer logic modules. Journal of the ACM, 13(4):615–629, October 1966. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Manacher:1967:PSR

[Man67a] G. K. Manacher. Production and stabilization of real-time task schedules. Journal of the ACM, 14(3):439–465, July 1967. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mancino:1967:RIS

[Man67b] O. G. Mancino. Resolution by iteration of some nonlinear systems. Journal of the ACM, 14(2):341–350, April 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Manna:1969:PPF

[Man69] Zohar Manna. Properties of programs and the first-order predicate calculus. Journal of the ACM, 16(2):244–255, April 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Manacher:1975:NLT

[Man75] Glenn Manacher. A new linear-time “on-line” algorithm for finding the smallest initial palindrome of a string. Journal of the ACM, 22(3):346–351, July 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Manacher:1979:FJS

[Man79a] Glenn K. Manacher. The Ford-Johnson sorting algorithm is not optimal. Journal of the ACM, 26(3):441–456, July 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Manacher:1979:SIH

[Man79b] Glenn K. Manacher. Significant improvements to the Hwang-Lin merging algorithm. Journal of the ACM, 26(3):434–440, July 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Manzini:2001:ABW

[Man01] Giovanni Manzini. An analysis of the Burrows — Wheeler transform. Journal of the ACM, 48(3):407–430, May 2001. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/citations/journals/jacm/
2001-48-3/p407-manzini/.

Mandelbaum:2003:RH

[Man03] Mark Mandelbaum. A report from Headquarters. Journal of the ACM, 50(1):21–23, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Markov:1958:ICS

[Mar58] A. A. Markov. On the inversion complexity of a system of functions. Journal of the ACM, 5(4):331–334, October 1958. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Translated by Morris D. Friedman.

Marimont:1959:NMC

[Mar59] Rosalind B. Marimont. A new method for checking the consistency of precedence matrices. Journal of the ACM, 6(2):164–171, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Maron:1961:AIE

[Mar61] M. E. Maron. Automatic indexing: an experiment inquiry. Journal of the ACM, 8(3):404–417, July 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Martin:1971:DEA

[Mar71] William A. Martin. Determining the equivalence of algebraic expressions by hash coding. Journal of the ACM, 18(4):549– 558, October 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Martelli:1976:GEA

[Mar76] Alberto Martelli. A Gaussian elimination algorithm for the enumeration of cut sets in a graph. Journal of the ACM, 23(1):58– 73, January 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Martel:1982:PSR

[Mar82] Charles Martel. Preemptive scheduling with release times, deadlines, and due times. Journal of the ACM, 29(3):812–829, July 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Marx:2013:THP

[Mar13] Dániel Marx. Tractable hypergraph properties for constraint satisfaction and conjunctive queries. Journal of the ACM, 60(6): 42:1–42:??, November 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mason:1956:DAC

[Mas56] Robert M. Mason. The digital approximation of contours. Journal of the ACM, 3(4):355–359, October 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mateti:1981:DPC

[Mat81] Prabhaker Mateti. A decision procedure for the correctness of a class of programs. Journal of the ACM, 28(2):215–222, April 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Matula:1982:BDS

[Mat82] David W. Matula. Basic digit sets for radix representation. Journal of the ACM, 29(4):1131–1143, October 1982. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mate:1990:NPT

[Mát90] Atilla Máté. Nondeterministic polynomial-time computations and models of arithmetic. Journal of the ACM, 37(1):175–193, January 1990. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/78941.html.

Maurer:1966:TCI

[Mau66] Ward Douglas Maurer. A theory of computer instructions. Journal of the ACM, 13(2):226–235, April 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Maurer:1969:DPI

[Mau69] Herman A. Maurer. A direct proof of the inherent ambiguity of a simple context-free language. Journal of the ACM, 16(2):256– 260, April 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mayers:2001:USQ

[May01] Dominic Mayers. Unconditional security in quantum cryptography. Journal of the ACM, 48(3):351–406, May 2001. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/citations/journals/jacm/
2001-48-3/p351-mayers/.

Macon:1956:GED

[MB56] Nathaniel Macon and Margaret Baskervill. On the generation of errors in the digital evaluation of continued fractions. Journal of the ACM, 3(3):199–202, July 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Matula:1983:SLO

[MB83] David W. Matula and Leland L. Beck. Smallest-last ordering and clustering and graph coloring algorithms. Journal of the ACM, 30(3):417–427, July 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mahanti:1985:GHS

[MB85] A. Mahanti and A. Bagchi. AND/OR graph heuristic search methods. Journal of the ACM, 32(1):28–51, January 1985. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
2459.html.

Manacher:1989:OCS

[MBM89] G. K. Manacher, T. D. Bui, and T. Mai. Optimum combinations of sorting and merging. Journal of the ACM, 36(2):290– 334, April 1989. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/62049.html.

Muntz:1970:PSR

[MC70] R. R. Muntz and E. G. Coffman, Jr. Preemptive scheduling of real-time tasks on multiprocessor systems. Journal of the ACM, 17(2):324–338, April 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mulligan:1972:CBA

[MC72] Gordon D. Mulligan and D. G. Corneil. Corrections to Bierstone’s algorithm for generating cliques. Journal of the ACM, 19(2):244–247, April 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Michel:1974:SFQ

[MC74] J. A. Michel and E. G. Coffman, Jr. Synthesis of a feedback queueing discipline for computer operation. Journal of the ACM, 21(2):329–339, April 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mitra:1987:RPC

[MC87] Debasis Mitra and Randall A. Cieslak. Randomized parallel communications on an extension of the Omega network. Journal of the ACM, 34(4):802–824, October 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/

42226.html. This is an extension of Valiant and Aleliunas’ algorithm to eliminate the need for scheduling. This algorithm also works on networks of fixed degree nodes.

McAllester:1993:ART

[McA93] David A. McAllester. Automatic recognition of tractability in inference relations. Journal of the ACM, 40(2):284–303, April 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/151265.html.

McAllester:2002:CAS

[McA02] David McAllester. On the complexity analysis of static analyses. Journal of the ACM, 49(4):512–537, July 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

McClellan:1973:ESS

[McC73] Michael T. McClellan. The exact solution of systems of linear equations with polynomial coefficients. Journal of the ACM, 20 (4):563–588, October 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

McCreight:1976:SES

[McC76] Edward M. McCreight. A space-economical suffix tree construction algorithm. Journal of the ACM, 23(2):262–272, April 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

McCarthy:2003:PPC

[McC03] John McCarthy. Problems and projections in CS for the next 49 years. Journal of the ACM, 50(1):73–79, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

McDermott:1982:NLI

[McD82] Drew McDermott. Nonmonotonic logic II: Nonmonotonic modal theories. Journal of the ACM, 29(1):33–57, January 1982. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

McGee:1959:GKS

[McG59] W. C. McGee. Generalization: Key to successful electronic data processing. Journal of the ACM, 6(1):1–23, January 1959. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

McKie:1978:ESF

[McK78] William McKie. An example of a skewing function. Journal of the ACM, 25(2):261–265, April 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

McKenna:1987:AES

[McK87] James McKenna. Asymptotic expansions of the sojourn time distribution functions of jobs in closed, product-form queuing networks. Journal of the ACM, 34(4):985–1003, October 1987. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
31851.html.

McLure:1965:PLS

[McL65] Robert M. McLure. A programming language for simulating digital systems. Journal of the ACM, 12(1):14–22, January 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

McLean:1984:FMA

[McL84] John McLean. A formal method for the abstract specification of software. Journal of the ACM, 31(3):600–627, July 1984. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

McNaughton:1967:PG

[McN67] Robert McNaughton. Parenthesis grammars. Journal of the ACM, 14(3):490–500, July 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Martin:1967:MCS

[ME67] David Martin and Gerald Estrin. Models of computations and systems—evaluation of vertex probabilities in graph models of computations. Journal of the ACM, 14(2):281–299, April 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Megiddo:1983:APC

[Meg83] Nimrod Megiddo. Applying parallel computation algorithms in the design of serial algorithms. Journal of the ACM, 30(4):852– 865, October 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Megiddo:1984:LPL

[Meg84] Nimrod Megiddo. Linear programming in linear time when the dimension is fixed. Journal of the ACM, 31(1):114–127, January 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Melahn:1956:DCV

[Mel56] Wesley S. Melahn. A description of a cooperative venture in the production of an automatic coding system. Journal of the ACM, 3(4):266–271, October 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Menon:1965:PCC

[Men65] M. V. Menon. On a problem concerning a central storage device served by multiple terminals. Journal of the ACM, 12(3):350– 355, July 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mendelzon:1979:AMD

[Men79] Alberto O. Mendelzon. On axiomatizing multivalued dependencies in relational databases. Journal of the ACM, 26(1):37–44, January 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mendelson:2015:LC

[Men15] Shahar Mendelson. Learning without concentration. Journal of the ACM, 62(3):21:1–21:??, June 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mendelson:2019:ULP

[Men19] Shahar Mendelson. An unrestricted learning procedure. Journal of the ACM, 66(6):42:1–42:??, December 2019. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/ft_gateway.cfm?id=3361699.

Mercer:1957:MP

[Mer57] Robert J. Mercer. Micro-programming. Journal of the ACM, 4(2):157–171, April 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Meyers:1965:MCN

[Mey65] Leroy F. Meyers. Morphological classification in the National Bureau of Standards mechanical translation system. Journal of the ACM, 12(4):437–472, October 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Meyer:1969:NSF

[Mey69] Albert R. Meyer. A note on star-free events. Journal of the ACM, 16(2):220–225, April 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

MeyeraufderHeide:1984:PLS

[Mey84] Friedhelm Meyer auf der Heide. A polynomial linear search algorithm for the n-dimensional knapsack problem. Journal of the ACM, 31(3):668–676, July 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

MeyerAufDerHeide:1985:LBS

[Mey85] Friedhelm Meyer Auf Der Heide. Lower bounds for solving linear Diophantine equations on random access machines. Journal of the ACM, 32(4):929–937, October 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/4250.html.

MeyeraufderHeide:1988:FAD

[Mey88] Friedhelm Meyer auf der Heide. Fast algorithms for Ndimensional restrictions of hard problems. Journal of the ACM, 35(3):740–747, July 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/44490.html.

Mattson:1963:FWC

[MF63] R. L. Mattson and O. Firschein. Feature word construction for use with pattern recognition algorithms: an experimental study. Journal of the ACM, 10:458–477, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

McAllester:1993:TSF

[MG93] David McAllester and Robert Givan. Taxonomic syntax for first order inference. Journal of the ACM, 40(2):246–283, April 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/151264.html.

Mundhenk:2000:CFH

[MGLA00] Martin Mundhenk, Judy Goldsmith, Christopher Lusena, and Eric Allender. Complexity of finite-horizon Markov decision process problems. Journal of the ACM, 47(4):681–720, July 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/citations/
journals/jacm/2000-47-4/p681-mundhenk/.

Meyer:1982:ADP

[MH82] Albert R. Meyer and Joseph Y. Halpern. Axiomatic definitions of programming languages: a theoretical assessment. Journal of the ACM, 29(2):555–576, April 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

McCune:1989:MSC

[MH89] William W. McCune and Lawrence J. Henschen. Maintaining state constraints in relational databases: A proof theoretic basis. Journal of the ACM, 36(1):46–68, January 1989. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/

59302.html. An early version appears as Argonne National Laboratory Tech. Memo ANL/MCS-TM-94, 1987.

Megiddo:1988:CSG

[MHG^{+}88] N. Megiddo, S. L. Hakimi, M. R. Garey, D. S. Johnson, and C. H. Papadimitriou. The complexity of searching a graph. Journal of the ACM, 35(1):18–44, January 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/42268.html.

Mickunas:1976:CCP

[Mic76] M. Dennis Mickunas. On the complete covering problem for LR(k) grammars. Journal of the ACM, 23(1):17–30, January 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Miehle:1957:BTF

[Mie57] William Miehle. Burroughs truth function evaluator. Journal of the ACM, 4(2):189–192, April 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Miller:1973:TAN

[Mil73] Webb Miller. Toward abstract numerical analysis. Journal of the ACM, 20(3):399–408, July 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Miller:1975:CSN

[Mil75] Webb Miller. Computer search for numerical instability. Journal of the ACM, 22(4):512–521, October 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Milner:1979:FFA

[Mil79] Robin Milner. Flowgraphs and flow algebras. Journal of the ACM, 26(4):794–818, October 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Miller:2003:J

[Mil03] Ray Miller. JACM 1973–1975. Journal of the ACM, 50(1):15, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Minnick:1957:TAP

[Min57] Robert C. Minnick. Tshebysheff approximations for power series. Journal of the ACM, 4(4):487–504, October 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [dV59].

Minnick:1967:SMR

[Min67] Robert C. Minnick. Survey of microcellular research. Journal of the ACM, 14(2):203–241, April 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Minsky:1970:FCC

[Min70] Marvin L. Minsky. Form and content in computer science. Journal of the ACM, 17(2):197–215, April 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Minoura:1982:DAR

[Min82] Toshimi Minoura. Deadlock avoidance revisited. Journal of the ACM, 29(4):1023–1048, October 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mitrani:1972:NMS

[Mit72] I. Mitrani. Nonpriority multiprogramming systems under heavy demand conditions — customers’ viewpoint. Journal of the ACM, 19(3):445–452, July 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mitra:1974:SAH

[Mit74] Debasis Mitra. Some aspects of hierarchical memory systems. Journal of the ACM, 21(1):54–65, January 1974. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Maron:1960:RPI

[MK60] M. E. Maron and J. L. Kuhns. On relevance, probabilistic indexing and information retrieval. Journal of the ACM, 7(3):216–244, July 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mickunas:1976:TLG

[MLS76] M. D. Mickunas, R. L. Lancaster, and V. B. Schneider. Transforming LR(k) grammars to LR(1), SLR(1), and (1,1) bounded right-context grammars. Journal of the ACM, 23(3):511–533, July 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

MacLaren:1965:URN

[MM65] M. Donald MacLaren and George Marsaglia. Uniform random number generators. Journal of the ACM, 12(1):83–89, January 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Micchelli:1975:HOS

[MM75] C. A. Micchelli and W. L. Miranker. High order search methods for finding roots. Journal of the ACM, 22(1):51–60, January 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Milne:1979:CPT

[MM79] George Milne and Robin Milner. Concurrent processes and their syntax. Journal of the ACM, 26(2):302–321, April 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mayr:1981:CFC

[MM81] Ernst W. Mayr and Albert R. Meyer. The complexity of the finite containment problem for Petri nets. Journal of the ACM, 28(3):561–576, July 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

McKenna:1984:AEI

[MM84] J. McKenna and Debasis Mitra. Asymptotic expansions and integral representations of moments of queue lengths in closed Markovian networks. Journal of the ACM, 31(2):346–360, April 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mitra:1986:AEC

[MM86] Debasis S. Mitra and J. McKenna. Asymptotic expansions for closed Markovian networks with state-dependent service rates. Journal of the ACM, 33(3):568–592, July 1986. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
5935.html.

Matsumoto:2024:BIH

[MM24] Namiko Matsumoto and Arya Mazumdar. Binary iterative hard thresholding converges with optimal number of measurements for 1-bit compressed sensing. Journal of the ACM, 71(5):35:1–35:??, October 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3680542.

Mohring:1997:MRB

[MMHW97] Rolf H. Möhring, Matthias Müller-Hannemann, and Karsten Weihe. Mesh refinement via bidirected flows: modeling, complexity, and computational results. Journal of the ACM, 44 (3):395–426, May 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/citations/journals/jacm/1997-44-3/p395-hohring/.

Mostefaoui:2015:SFA

[MMR15] Achour Mostéfaoui, Hamouma Moumen, and Michel Raynal. Signature-free asynchronous binary Byzantine consensus with t < n/3, O(n^{2}) messages, and O(1) expected time. Journal of the ACM, 62(4):31:1–31:??, August 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Maneva:2007:NLS

[MMW07] Elitza Maneva, Elchanan Mossel, and Martin J. Wainwright. A new look at survey propagation and its generalizations. Journal of the ACM, 54(4):17:1–17:41, July 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Merro:2005:BTM

[MN05] Massimo Merro and Francesco Zappa Nardelli. Behavioral theory for mobile ambients. Journal of the ACM, 52(6):961–1023, November 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Miksa:2024:GMP

[MN24] Mladen Miksa and Jakob Nordström. A generalized method for proving polynomial calculus degree lower bounds. Journal of the ACM, 71(6):37:1–37:??, December 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3675668.

McNaughton:1988:CRT

[MNO88] Robert McNaughton, Paliath Narendran, and Friedrich Otto. Church–Rosser Thue systems and formal languages. Journal of the ACM, 35(2):324–344, April 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/42284.html.

Miller:1956:PCI

[MO56] Robert C. Miller, Jr. and Bruce G. Oldfield. Producing computer instructions for the PACT I compiler. Journal of the ACM, 3 (4):288–291, October 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Morris:1968:CSM

[MO68] Gerald L. Morris and Patrick L. Odell. Common solutions for n matrix equations with applications. Journal of the ACM, 15 (2):272–274, April 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mock:1956:LOP

[Moc56] Owen R. Mock. Logical organization of the PACT I compiler. Journal of the ACM, 3(4):279–287, October 1956. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mock:1971:NAN

[Moc71] M. S. Mock. Numerical analysis of a nonlinear diffusion problem. Journal of the ACM, 18(3):373–380, July 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Moitra:2019:ACL

[Moi19] Ankur Moitra. Approximate counting, the Lovász local lemma, and inference in graphical models. Journal of the ACM, 66(2): 10:1–10:??, April 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3268930.

Moler:1967:IRF

[Mol67] Cleve B. Moler. Iterative refinement in floating point. Journal of the ACM, 14(2):316–321, April 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Molloy:2018:FTC

[Mol18] Michael Molloy. The freezing threshold for k-colourings of a random graph. Journal of the ACM, 65(2):7:1–7:??, March 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Montanari:1968:MOS

[Mon68] U. Montanari. A method for obtaining skeletons using a quasi-Euclidean distance. Journal of the ACM, 15(4):600–624, October 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Montanari:1969:CSD

[Mon69] Ugo Montanari. Continuous skeletons from digitized images. Journal of the ACM, 16(4):534–549, October 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Montanari:1970:LPD

[Mon70] G. Ugo Montanari. On limit properties in digitization schemes. Journal of the ACM, 17(2):348–360, April 1970. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Moore:1967:CAS

[Moo67] J. B. Moore. A convergent algorithm for solving polynomial equations. Journal of the ACM, 14(2):311–315, April 1967. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Morrison:1959:NQM

[Mor59] David Morrison. Numerical quadrature in many dimensions. Journal of the ACM, 6(2):219–222, April 1959. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Morrison:1960:RUT

[Mor60] David D. Morrison. Remarks on the unitary triangularization of a nonsymmetric matrix. Journal of the ACM, 7(2):185–186, April 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Morrison:1962:OMS

[Mor62] D. Morrison. Optimal mesh size in the numerical integration of an ordinary differential equation. Journal of the ACM, 9(1):98– 103, January 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Morrison:1968:PPA

[Mor68a] Donald R. Morrison. PATRICIA–practical algorithm to retrieve information coded in alphanumeric. Journal of the ACM, 15(4): 514–534, October 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Morse:1968:MMA

[Mor68b] Stephen P. Morse. A mathematical model for the analysis of contour-line data. Journal of the ACM, 15(2):205–220, April 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Morgenstern:1973:NLB

[Mor73] Jacques Morgenstern. Note on a lower bound of the linear complexity of the fast Fourier transform. Journal of the ACM, 20 (2):305–306, April 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Morgenstern:1975:LCC

[Mor75] Jacques Morgenstern. The linear complexity of computation. Journal of the ACM, 22(2):184–194, April 1975. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Morishita:1997:ACP

[Mor97] Shinichi Morishita. Avoiding Cartesian products for multiple joins. Journal of the ACM, 44(1):57–85, January 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Moshman:1954:GPR

[Mos54] Jack Moshman. The generation of pseudo-random numbers on a decimal calculator. Journal of the ACM, 1(2):88–91, April 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Moshos:1955:AIA

[Mos55] George J. Moshos. Analog interpolator for automatic control. Journal of the ACM, 2(2):83–91, April 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Moshman:1958:ASE

[Mos58] Jack Moshman. The application of sequential estimation to computer simulation and Monte Carlo procedures. Journal of the ACM, 5(4):343–352, October 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Motwani:1994:ACA

[Mot94] Rajeev Motwani. Average-case analysis of algorithms for matching and related problems. Journal of the ACM, 41(6):1329–1356, November 1994. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/195663.html.

Moursund:1965:CSL

[Mou65a] David Moursund. Chebyshev solution of n + 1 linear equations in n unknowns. Journal of the ACM, 12(3):383–387, July 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Moursund:1965:EMR

[Mou65b] David G. Moursund. Examination of multiple roots and root clusters of a polynomial using the Bernoulli procedure. Journal of the ACM, 12(2):169–174, April 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mowle:1967:AGS

[Mow67] Frederic J. Mowle. An algorithm for generating stable feedback shift registers of order n. Journal of the ACM, 14(3):529–542, July 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mowle:1970:CNS

[Mow70] Frederic J. Mowle. Controllability of nonlinear sequential networks. Journal of the ACM, 17(3):518–524, July 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Milnes:1959:BCS

[MP59] Harold W. Milnes and Renfrey B. Potts. Boundary contraction solution of Laplace’s differential equation. Journal of the ACM, 6(2):226–235, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mileto:1965:SCA

[MP65] Franco Mileto and Gianfranco Putzolu. Statistical complexity of algorithms for Boolean function minimization. Journal of the ACM, 12(3):364–375, July 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [MP66a].

Mileto:1966:CSC

[MP66a] Franco Mileto and Gianfranco Potzolu. Corrigenda: “Statistical complexity of algorithms for Boolean function minimization”. Journal of the ACM, 13(3):469, July 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [MP65].

Minsky:1966:USN

[MP66b] Marvin Minsky and Seymour Papert. Unrecognizable sets of numbers. Journal of the ACM, 13(2):281–286, April 1966. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Manna:1970:FPF

[MP70a] Zohar Manna and Amir Pnueli. Formalization of properties of functional programs. Journal of the ACM, 17(3):555–569, July 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Meyer:1970:NLS

[MP70b] C. D. Meyer and R. J. Painter. Note on a least squares inverse for a matrix. Journal of the ACM, 17(1):110–112, January 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mylopoulos:1971:TPQa

[MP71a] J. P. Mylopoulos and T. Pavlidis. On the topological properties of quantized spaces. I. the notion of dimension. Journal of the ACM, 18(2):239–246, April 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mylopoulos:1971:TPQb

[MP71b] J. P. Mylopoulos and T. Pavlidis. On the topological properties of quantized spaces. II. connectivity and order of connectivity. Journal of the ACM, 18(2):247–254, April 1971. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

McAfee:1972:ADS

[MP72] J. McAfee and L. Presser. An algorithm for the design of simple precedence grammars. Journal of the ACM, 19(3):385–395, July 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Muller:1975:BCN

[MP75] David E. Muller and Franco P. Preparata. Bounds to complexities of networks for sorting and for switching. Journal of the ACM, 22(2):195–201, April 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Muller:1976:RAE

[MP76] David E. Muller and Franco P. Preparata. Restructuring of arithmetic expressions for parallel evaluation. Journal of the ACM, 23(3):534–543, July 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mehlhorn:1986:RTR

[MP86] K. Mehlhorn and F. P. Preparata. Routing through a rectangle. Journal of the ACM, 33(1):60–85, January 1986. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
4994.html.

Mitchell:1991:WRP

[MP91] Joseph S. B. Mitchell and Christos H. Papadimitriou. The weighted region problem: Finding shortest paths through a weighted planar subdivision. Journal of the ACM, 38(1):18– 73, January 1991. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/102784.html.

Mandow:2010:MSC

[MP10] Lawrence Mandow and José Luis Pérez De La Cruz. Multiobjective A^{∗} search with consistent heuristics. Journal of the ACM, 57 (5):27:1–27:25, June 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mottet:2024:SAA

[MP24] Antoine Mottet and Michael Pinsker. Smooth approximations: an algebraic approach to CSPs over finitely bounded homogeneous structures. Journal of the ACM, 71(5):36:1–36:??, October 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3689207.

MacKenzie:1998:CRP

[MPR98] P. D. MacKenzie, C. G. Plaxton, and R. Rajaraman. On contention resolution protocols and associated probabilistic phenomena. Journal of the ACM, 45(2):324–378, March 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org:80/pubs/citations/
journals/jacm/1998-45-2/p324-mackenzie/.

Miltersen:1996:ACM

[MPT96] Peter Bro Miltersen, Mike Paterson, and Jun Tarui. The asymptotic complexity of merging networks. Journal of the ACM, 43 (1):147–165, January 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/jacm/227693.html.

Milne:1959:SNS

[MR59] W. E. Milne and R. R. Reynolds. Stability of a numerical solution of differential equations. Journal of the ACM, 6(2):196–203, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Milne:1960:SNS

[MR60] W. E. Milne and R. R. Reynolds. Stability of a numerical solution of differential equations—Part II. Journal of the ACM, 7(1):46– 56, January 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Milne:1962:FOM

[MR62] W. E. Milne and R. R. Reynolds. Fifth-order methods for the numerical solution of ordinary differential equations. Journal of the ACM, 9(1):64–70, January 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Murray:1987:IPR

[MR87] Neil V. Murray and Erik Rosenthal. Inference with path resolution and semantic graphs. Journal of the ACM, 34(2):225– 254, April 1987. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/23716.html.

Murray:1993:DMP

[MR93] Neil V. Murray and Erik Rosenthal. Dissolution: Making paths vanish. Journal of the ACM, 40(3):504–535, July 1993. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
174135.html.

Martinez:1998:RBS

[MR98] Conrado Martínez and Salvador Roura. Randomized binary search trees. Journal of the ACM, 45(2):288–323, March 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org:80/pubs/citations/
journals/jacm/1998-45-2/p288-martinez/.

Mulzer:2008:MWT

[MR08] Wolfgang Mulzer and Günter Rote. Minimum-weight triangulation is NP-hard. Journal of the ACM, 55(2):11:1–11:29, May 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Moshkovitz:2010:TQP

[MR10a] Dana Moshkovitz and Ran Raz. Two-query PCP with subconstant error. Journal of the ACM, 57(5):29:1–29:29, June 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Motik:2010:RDL

[MR10b] Boris Motik and Riccardo Rosati. Reconciling description logics and rules. Journal of the ACM, 57(5):30:1–30:62, June 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mostefaoui:2003:CIV

[MRR03] Achour Mostefaoui, Sergio Rajsbaum, and Michel Raynal. Conditions on input vectors for consensus solvability in asynchronous distributed systems. Journal of the ACM, 50(6):922–954, November 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mock:1959:SSP

[MS59] Owen Mock and Charles J. Swift. The share 709 system: Programmed input-output buffering. Journal of the ACM, 6(2):145– 151, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Michie:1974:SBD

[MS74] D. Michie and E. E. Sibert. Some binary derivation systems. Journal of the ACM, 21(2):175–190, April 1974. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Muller:1989:IMD

[MS89] John H. Muller and Jeremy Spinrad. Incremental modular decomposition. Journal of the ACM, 36(1):1–19, January 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/59300.html.

Mansour:1992:IBP

[MS92] Yishay Mansour and Baruch Schieber. The intractability of bounded protocols for on-line sequence transmission over non-FIFO channels. Journal of the ACM, 39(4):783–799, October 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/146596.html.

Marcus:1996:FMD

[MS96] Sherry Marcus and V. S. Subrahmanian. Foundations of multimedia database systems. Journal of the ACM, 43(3):463– 523, May 1996. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/jacm/233554.html.

Miklau:2004:CEF

[MS04] Gerome Miklau and Dan Suciu. Containment and equivalence for a fragment of XPath. Journal of the ACM, 51(1):2–45, January 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mastrolilli:2011:HAF

[MS11] Monaldo Mastrolilli and Ola Svensson. Hardness of approximating flow and job shop scheduling problems. Journal of the ACM, 58(5):20:1–20:??, October 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Miller:2016:RPS

[MS16a] Carl A. Miller and Yaoyun Shi. Robust protocols for securely expanding randomness and distributing keys using untrusted quantum devices. Journal of the ACM, 63(4):33:1–33:??, November 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Momke:2016:RAE

[MS16b] Tobias Mömke and Ola Svensson. Removing and adding edges for the traveling salesman problem. Journal of the ACM, 63 (1):2:1–2:??, March 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Makarychev:2018:SOP

[MS18] Konstantin Makarychev and Maxim Sviridenko. Solving optimization problems with diseconomies of scale via decoupling. Journal of the ACM, 65(6):42:1–42:??, November 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/ft_gateway.cfm?id=3266140.

Malajovich:2019:TNC

[MS19] Gregorio Malajovich and Mike Shub. A theory of NPcompleteness and ill-conditioning for approximate real computations. Journal of the ACM, 66(4):27:1–27:??, August 2019. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/ft_gateway.cfm?id=3321479.

Mika:2021:FFU

[MS21] Maksymilian Mika and Marek Szykuła. The Frobenius and factor universality problems of the Kleene star of a finite set of words. Journal of the ACM, 68(3):18:1–18:22, May 2021. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3447237.

Moran:1985:ART

[MSM85] Shlomo Moran, Marc Snir, and Udi Manber. Applications of Ramsey’s theorem to decision tree complexity. Journal of the ACM, 32(4):938–949, October 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/4259.html.

Meghini:2001:MMI

[MSS01] Carlo Meghini, Fabrizio Sebastiani, and Umberto Straccia. A model of multimedia information retrieval. Journal of the ACM, 48(5):909–970, September 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Marsden:2024:ECO

[MSSV24] Annie Marsden, Vatsal Sharan, Aaron Sidford, and Gregory Valiant. Efficient convex optimization requires superlinear memory. Journal of the ACM, 71(6):41:1–41:??, December 2024. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3689208.

Mansour:1991:LBI

[MST91] Yishay Mansour, Baruch Schieber, and Prasoon Tiwari. A lower bound for integer greatest common divisor computations. Journal of the ACM, 38(2):453–471, April 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/103522.html.

Marek:1993:MNL

[MST93] V. Wiktor Marek, Grigori F. Shvarts, and Mirosław Truszczyński. Modal nonmonotonic logics: Ranges, characterization, computation. Journal of the ACM, 40(4):963–990, September 1993. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
153773.html.

Mohring:1999:ASS

[MSU99] Rolf H. Möhring, Andreas S. Schulz, and Marc Uetz. Approximation in stochastic scheduling: the power of LP-based priority policies. Journal of the ACM, 46(6):924–942, November 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL http://www.acm.org/pubs/articles/journals/
jacm/1999-46-6/p924-mohring/p924-mohring.pdf.

Mehta:2007:AGO

[MSVV07] Aranyak Mehta, Amin Saberi, Umesh Vazirani, and Vijay Vazirani. AdWords and generalized online matching. Journal of the ACM, 54(5):22:1–22:19, October 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Maurer:1982:DHG

[MSW82] H. A. Maurer, A. Salomaa, and D. Wood. Dense hierarchies of grammatical families. Journal of the ACM, 29(1):118–126, January 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Maurer:1983:SFT

[MSW83] H. A. Maurer, A. Salomaa, and D. Wood. A supernormal-form theorem for context-free grammars. Journal of the ACM, 30 (1):95–102, January 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Maier:1981:CTI

[MSY81] David Maier, Yehoshua Sagiv, and Mihalis Yannakakis. On the complexity of testing implications of functional and join dependencies. Journal of the ACM, 28(4):680–695, October 1981. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Moyles:1969:AFM

[MT69] Dennis M. Moyles and Gerald L. Thompson. An algorithm for finding a minimum equivalent graph of a digraph. Journal of the ACM, 16(3):455–460, July 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Manber:1985:CPP

[MT85] Udi Manber and Martin Tompa. Complexity of problems on probabilistic, nondeterministic, and alternating decision trees. Journal of the ACM, 32(3):720–732, July 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/

3838.html. This paper compares lower bounds on the running times of algorithms that allow probabilistic, non-deterministic and alternating control on decision trees. Decision trees that allow internal randomization at the expense of a small probability of error are shown to run no faster asymptotically than ordinary decision trees for a collection of problems. An earlier version of this publication appeared in Proc. 14th Ann. ACM Symp. on Theory of Computing, 1982, pp. 234–244.

MixBarrington:1988:FMF

[MT88] David A. Mix Barrington and Denis Thérien. Finite monoids and the fine structure of NC^{1}. Journal of the ACM, 35(4):941– 952, October 1988. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/63138.html.

Marek:1991:AL

[MT91] Wiktor Marek and Miroslaw Truszczynski. Autoepistemic logic. Journal of the ACM, 38(3):588–619, July 1991. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
116836.html.

Mehlhorn:1993:DIS

[MT93] Kurt Mehlhorn and Athanasios Tsakalidis. Dynamic interpolation search. Journal of the ACM, 40(3):621–634, July 1993. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
174139.html.

Moser:2010:CPG

[MT10] Robin A. Moser and Gábor Tardos. A constructive proof of the general Lovász local lemma. Journal of the ACM, 57(2): 11:1–11:15, January 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Monniaux:2019:CCA

[MT19] David Monniaux and Valentin Touzeau. On the complexity of cache analysis for different replacement policies. Journal of the ACM, 66(6):41:1–41:??, December 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https:

//dl.acm.org/ft_gateway.cfm?id=3366018.

Murawski:2021:GSI

[MT21] Andrzej S. Murawski and Nikos Tzevelekos. Game semantics for interface middleweight Java. Journal of the ACM, 68(1): 4:1–4:51, February 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/

doi/10.1145/3428676.

Miller:1997:SSP

[MTTV97] Gary L. Miller, Shang-Hua Teng, William Thurston, and Stephen A. Vavasis. Separators for sphere-packings and nearest neighbor graphs. Journal of the ACM, 44(1):1–29, January 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Miller:1960:IPF

[MTZ60] C. E. Miller, A. W. Tucker, and R. A. Zemlin. Integer programming formulation of traveling salesman problems. Journal of the ACM, 7(4):326–329, October 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mukhopadhyay:1968:REN

[Muk68] Amar Mukhopadhyay. Representation of events in the von Neumann cellular model. Journal of the ACM, 15(4):693–705, October 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Muller:1959:CMG

[Mul59] Mervin E. Muller. A comparison of methods for generating normal deviates on digital computers. Journal of the ACM, 6(3): 376–383, July 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mulvey:1978:PSP

[Mul78] John M. Mulvey. Pivot strategies for primal-simplex network codes. Journal of the ACM, 25(2):266–270, April 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mulmuley:1991:FPP

[Mul91] Ketan Mulmuley. A fast planar partition algorithm, II. Journal of the ACM, 38(1):74–103, January 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/102785.html.

Mulmuley:2011:PVN

[Mul11] Ketan D. Mulmuley. On P vs. NP and geometric complexity theory: Dedicated to Sri Ramakrishna. Journal of the ACM, 58 (2):5:1–5:26, April 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Murray:1955:MR

[Mur55] F. J. Murray. Mechanisms and robots. Journal of the ACM, 2(2): 61–82, April 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Musser:1975:MPF

[Mus75] David R. Musser. Multivariate polynomial factorization. Journal of the ACM, 22(2):291–308, April 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Musser:1978:EPI

[Mus78] David R. Musser. On the efficiency of a polynomial irreducibility test. Journal of the ACM, 25(2):271–282, April 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Milo:2005:I

[MV05] Tova Milo and Victor Vianu. Introduction. Journal of the ACM, 52(2):245, March 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Miles:2015:SPN

[MV15] Eric Miles and Emanuele Viola. Substitution-permutation networks, pseudorandom functions, and natural proofs. Journal of the ACM, 62(6):46:1–46:??, December 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mangoubi:2025:PLR

[MV25] Oren Mangoubi and Nisheeth K. Vishnoi. Private low-rank approximation for covariance matrices, Dyson Brownian motion, and eigenvalue-gap bounds for Gaussian perturbations. Journal of the ACM, 72(2):14:1–14:??, April 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

McKellar:1978:DPR

[MW78] A. C. McKellar and C. K. Wong. Dynamic placement of records in linear storage. Journal of the ACM, 25(3):421–434, July 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mitra:1984:PMD

[MW84] Debasis Mitra and P. J. Weinberger. Probabilistic models of database locking: Solutions, computational algorithms, and asymptotics. Journal of the ACM, 31(4):855–878, October 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Manna:1986:SRA

[MW86] Zohar Manna and Richard Waldinger. Special relations in automated deduction. Journal of the ACM, 33(1):1–59, January 1986. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/4905.html.

Mendelson:1979:PMO

[MY79] Haim Mendelson and Uri Yechiali. Performance measures for ordered lists in random-access files. Journal of the ACM, 26(4):

654–667, October 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Mendelson:1980:NAA

[MY80] Haim Mendelson and Uri Yechiali. A new approach to the analysis of linear probing schemes. Journal of the ACM, 27(3):474– 483, July 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Melamed:1984:NCS

[MY84] Benjamin Melamed and Micha Yadin. Numerical computation of sojourn-time distributions in queuing networks. Journal of the ACM, 31(4):839–854, October 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Moran:2016:SCS

[MY16] Shay Moran and Amir Yehudayoff. Sample compression schemes for VC classes. Journal of the ACM, 63(3):21:1–21:??, September 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Myers:1992:FRA

[Mye92] Gene Myers. A four Russians algorithm for regular expression pattern matching. Journal of the ACM, 39(2):430–448, April 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/128755.html.

Myers:1999:FBV

[Mye99] Gene Myers. A fast bit-vector algorithm for approximate string matching based on dynamic programming. Journal of the ACM, 46(3):395–415, May 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org:80/

pubs/citations/journals/jacm/1999-46-3/p395-myers/.

Nagler:1959:AS

[Nag59] H. Nagler. Amphisbaenic sorting. Journal of the ACM, 6(4): 459–468, October 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [Liv60, Nag61, Goe61].

Nagler:1961:LEA

[Nag61] H. Nagler. Letter to the editor: an answer to Mr. J. A. Lively’s remarks on the paper “Amphisbaenic sorting”. Journal of the ACM, 8(1):117, January 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Liv60, Nag59, Goe61].

Natapoff:1967:ITC

[Nat67] Alan Natapoff. Irreducible topological components of an arbitrary Boolean truth function and generation of their minimal coverings. Journal of the ACM, 14(2):376–381, April 1967. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Nau:1983:DQF

[Nau83] Dana S. Nau. Decision quality as a function of search depth on game trees. Journal of the ACM, 30(4):687–708, October 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Naughton:1989:MFF

[Nau89] Jeffrey F. Naughton. Minimizing function-free recursive inference rules. Journal of the ACM, 36(1):69–91, January 1989. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
59303.html.

Nebel:1995:RAT

[NB95] Bernhard Nebel and Hans-Jürgen Bürckert. Reasoning about temporal relations: a maximal tractable subclass of Allen’s interval algebra. Journal of the ACM, 42(1):43–66, January 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/200848.html.

Nederhof:1996:LTS

[NB96] Mark-Jan Nederhof and Eberhard Bertsch. Linear-time suffix parsing for deterministic languages. Journal of the ACM, 43(3): 524–554, May 1996. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/jacm/233555.html.

Nance:1972:BPA

[NBC72] Richard E. Nance, U. Narayan Bhat, and Billy G. Claybrook. Busy period analysis of a time-sharing system: Transform inversion. Journal of the ACM, 19(3):453–463, July 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Negron:1966:DOT

[Neg66] C. D. Negron. Digital one-third octave spectral analysis. Journal of the ACM, 13(4):605–614, October 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Nelson:1987:SCT

[Nel87] Randolph Nelson. Stochastic catastrophe theory in computer performance modeling. Journal of the ACM, 34(3):661–685, July 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/28878.html.

Netherwood:1959:LMT

[Net59] Douglas B. Netherwood. Logic matrices and the truth function problem. Journal of the ACM, 6(3):405–414, July 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Neumann:1954:SODa

[Neu54a] A. J. Neumann. Supplement: ONR Digital Computer Newsletter. Journal of the ACM, 1(1):45–55, January 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Neumann:1954:SODb

[Neu54b] A. J. Neumann. Supplement: ONR Digital Computer Newsletter. Journal of the ACM, 1(2):93–100, April 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Neumann:1954:SODc

[Neu54c] A. J. Neumann. Supplement: ONR Digital Computer Newsletter. Journal of the ACM, 1(3):139–148, July 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Neumann:1954:SODd

[Neu54d] A. J. Neumann. Supplement: ONR Digital Computer Newsletter. Journal of the ACM, 1(4):193–200, October 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Neumann:1955:SODa

[Neu55a] A. J. Neumann. Supplement: ONR Digital Computer Newsletter. Journal of the ACM, 2(1):53–60, January 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Neumann:1955:SODb

[Neu55b] A. J. Neumann. Supplement: ONR Digital Computer Newsletter. Journal of the ACM, 2(2):119–136, April 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Neumann:1955:SODc

[Neu55c] A. J. Neumann. Supplement: ONR Digital Computer Newsletter. Journal of the ACM, 2(3):211–228, July 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Neumann:1955:SODd

[Neu55d] A. J. Neumann. Supplement: ONR Digital Computer Newsletter. Journal of the ACM, 2(4):283–298, October 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Nevins:1970:PLA

[Nev70] Arthur J. Nevins. A programming language with automatic goal generation and selection. Journal of the ACM, 17(2):216–230, April 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Nevins:1974:HOL

[Nev74] Arthur J. Nevins. A human oriented logic for automatic theoremproving. Journal of the ACM, 21(4):606–621, October 1974. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Newman:1965:LMU

[New65] D. J. Newman. Location of the maximum on unimodal surfaces. Journal of the ACM, 12(3):395–398, July 1965. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Nicol:1993:CCS

[Nic93] David M. Nicol. The cost of conservative synchronization in parallel discrete event simulations. Journal of the ACM, 40(2):304– 333, April 1993. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/151266.html.

Nipkow:1990:UPA

[Nip90] Tobias Nipkow. Unification in primal algebras, their powers and their varieties. Journal of the ACM, 37(4):742–776, October 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/96569.html.

Nivasch:2010:IBN

[Niv10] Gabriel Nivasch. Improved bounds and new techniques for Davenport–Schinzel sequences and their generalizations. Journal of the ACM, 57(3):17:1–17:44, March 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Nelson:1957:SCS

[NJ57] R. T. Nelson and J. R. Jackson. SWAC computations for some m × n scheduling problems. Journal of the ACM, 4(4):438–441, October 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Nadathur:1990:HOH

[NM90] Gopalan Nadathur and Dale Miller. Higher-order Horn clauses. Journal of the ACM, 37(4):777–814, October 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
96570.html.

Nelson:1980:FDP

[NO80] Greg Nelson and Derek C. Oppen. Fast decision procedures based on congruence closure. Journal of the ACM, 27(2):356– 364, April 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Noetzel:1979:GQD

[Noe79] Andrew S. Noetzel. A generalized queueing discipline for product form network solutions. Journal of the ACM, 26(4):779–793, October 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Nieuwenhuis:2006:SSS

[NOT06] Robert Nieuwenhuis, Albert Oliveras, and Cesare Tinelli. Solving SAT and SAT Modulo Theories: From an abstract Davis– Putnam–Logemann–Loveland procedure to DPLL(T). Journal of the ACM, 53(6):937–977, November 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Nourani:1983:AIT

[Nou83] C. Farshid Nourani. Abstract implementations and their correctness proofs. Journal of the ACM, 30(2):343–359, April 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ngo:2018:WCO

[NPRR18] Hung Q. Ngo, Ely Porat, Christopher Ré, and Atri Rudra. Worst-case optimal join algorithms. Journal of the ACM, 65(3): 16:1–16:??, March 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Naughton:1994:HFP

[NR94] Jeffrey F. Naughton and Raghu Ramakrishnan. How to forget the past without repeating it. Journal of the ACM, 41(6): 1151–1177, November 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/195634.html.

Naor:2004:NTC

[NR04] Moni Naor and Omer Reingold. Number-theoretic constructions of efficient pseudo-random functions. Journal of the ACM, 51 (2):231–262, March 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Naor:2009:COM

[NR09] Moni Naor and Guy N. Rothblum. The complexity of online memory checking. Journal of the ACM, 56(1):2:1–2:46, January 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Nassimi:1980:ORA

[NS80] David Nassimi and Sartaj Sahni. An optimal routing algorithm for mesh-connected parallel computers. Journal of the ACM, 27 (1):6–29, January 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Nassimi:1982:PPS

[NS82] David Nassimi and Sartaj Sahni. Parallel permutation and sorting algorithms and a new generalized connection network. Journal of the ACM, 29(3):642–667, July 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Nayak:2006:LAQ

[NS06a] Ashwin Nayak and Julia Salzman. Limits on the ability of quantum states to convey classical messages. Journal of the ACM, 53 (1):184–206, January 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Nederhof:2006:PPS

[NS06b] Mark-Jan Nederhof and Giorgio Satta. Probabilistic parsing strategies. Journal of the ACM, 53(3):406–436, May 2006. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Neiger:2024:FMC

[NSSV24] Vincent Neiger, Bruno Salvy, Éric Schost, and Gilles Villard. Faster modular composition. Journal of the ACM, 71(2):11:1– 11:??, April 2024. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL https://dl.acm.org/doi/10.
1145/3638349.

Nowak:2024:TCC

[NSW24] Thomas Nowak, Ulrich Schmid, and Kyrill Winkler. Topological characterization of consensus in distributed systems. Journal of the ACM, 71(6):39:1–39:??, December 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3687302.

Neiger:1993:SSC

[NT93a] Gil Neiger and Sam Toueg. Simulating synchronized clocks and common knowledge in distributed systems. Journal of the ACM, 40(2):334–367, April 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/151267.html.

Nelson:1993:PES

[NT93b] Randolph Nelson and Donald Towsley. A performance evaluation of several priority policies for parallel processing systems. Journal of the ACM, 40(3):714–740, July 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/174143.html.

Nodine:1995:GSO

[NV95] Mark H. Nodine and Jeffrey Scott Vitter. Greed sort: Optimal deterministic sorting on parallel disks. Journal of the ACM, 42 (4):919–933, July 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/210343.html.

Neven:2002:ESD

[NV02] Frank Neven and Jan Van Den Bussche. Expressiveness of structured document query languages based on attribute grammars. Journal of the ACM, 49(1):56–100, January 2002. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Nievergelt:1973:UBT

[NW73] J. Nievergelt and C. K. Wong. Upper bounds for the total path length of binary trees. Journal of the ACM, 20(1):1–6, January 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

OConnor:1964:MIM

[O’C64] John O’Connor. Mechanized indexing methods and their testing. Journal of the ACM, 11(4):437–449, October 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

OConnor:1965:ASR

[O’C65] John O’Connor. Automatic subject recognition in scientific papers: An empirical study. Journal of the ACM, 12(4):490–515, October 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Odell:1967:CFP

[OD67] P. L. Odell and H. P. Decell. On computing the fixed-point probability vector of ergodic transition matrices. Journal of the ACM, 14(4):765–768, October 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Oehmke:1963:SAI

[Oeh63] Robert H. Oehmke. On the structures of an automaton and its input semigroup. Journal of the ACM, 10:521–525, 1963. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Oettinger:1957:AIA

[Oet57] Anthony G. Oettinger. Account identification for automatic data processing. Journal of the ACM, 4(3):245–253, July 1957. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ott:1961:DSM

[OF61] Gene Ott and Neil H. Feinstein. Design of sequential machines from their regular expressions. Journal of the ACM, 8(4):585– 600, October 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Omlin:1996:CDF

[OG96] Christian W. Omlin and C. Lee Giles. Constructing deterministic finite-state automata in recurrent neural networks. Journal of the ACM, 43(6):937–972, November 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/jacm/235811.html.

Orponen:1994:IC

[OKSW94] Pekka Orponen, Ker-I Ko, Uwe Schöning, and Osamu Watanabe. Instance complexity. Journal of the ACM, 41(1):96–121, January 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/174648.html.

Oliver:1967:AFE

[Oli67] I. Oliver. Analysis of factorial experiments using generalized matrix operations. Journal of the ACM, 14(3):508–519, July 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Omahen:1978:AAD

[OM78] K. Omahen and V. Marathe. Analysis and applications of the delay cycle for the M/M/C queueing system. Journal of the ACM, 25(2):283–303, April 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Omahen:1977:CBM

[Oma77] Kenneth J. Omahen. Capacity bounds for multiresource queues. Journal of the ACM, 24(4):646–663, October 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Orden:1968:BSA

[ON68] A. Orden and V. Nalbandian. A bidirectional simplex algorithm. Journal of the ACM, 15(2):221–235, April 1968. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Oney:1975:QAS

[One75] Walter C. Oney. Queueing analysis of the scan policy for movinghead disks. Journal of the ACM, 22(3):397–412, July 1975. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

ONeil:1999:OPL

[OOW99] Elizabeth J. O’Neil, Patrick E. O’Neil, and Gerhard Weikum. An optimality proof of the LRU-K page replacement algorithm. Journal of the ACM, 46(1):92–112, January 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org:80/pubs/citations/
journals/jacm/1999-46-1/p92-o_neil/.

Oppen:1980:RAR

[Opp80] Derek C. Oppen. Reasoning about recursively defined data structures. Journal of the ACM, 27(3):403–411, July 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ouaknine:2019:DMM

[OPSPW19] Joël Ouaknine, Amaury Pouly, João Sousa-Pinto, and James Worrell. On the decidability of membership in matrixexponential semigroups. Journal of the ACM, 66(3):15:1–15:??, June 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/ft_gateway.
cfm?id=3286487.

Orda:1990:SPM

[OR90] Ariel Orda and Raphael Rom. Shortest-path and minimum delay algorithms in networks with time-dependent edge-length. Journal of the ACM, 37(3):607–625, July 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/214078.html.

OHearn:2000:APL

[OR00] Peter W. O’Hearn and John C. Reynolds. From Algol to polymorphic linear lambda-calculus. Journal of the ACM, 47(1): 167–223, January 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/citations/journals/jacm/2000-47-1/p167-o_hearn/.

Ostrovsky:2002:PTA

[OR02] Rafail Ostrovsky and Yuval Rabani. Polynomial-time approximation schemes for geometric min-sum median clustering. Journal of the ACM, 49(2):139–156, March 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ostrovsky:2007:LDE

[OR07] Rafail Ostrovsky and Yuval Rabani. Low distortion embeddings for edit distance. Journal of the ACM, 54(5):23:1–23:16, October 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ormsby:1961:DNF

[Orm61] Joseph F. A. Ormsby. Design of numerical filters with applications to missile data processing. Journal of the ACM, 8(3): 440–466, July 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ostrovsky:2012:ELT

[ORSS12] Rafail Ostrovsky, Yuval Rabani, Leonard J. Schulman, and Chaitanya Swamy. The effectiveness of Lloyd-type methods for the k-means problem. Journal of the ACM, 59(6):28:1–28:22, December 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Special issue text supplied by Joanne Pello.

Ortega:1960:SST

[Ort60] J. M. Ortega. On Sturm sequences for tridiagonal matrices. Journal of the ACM, 7(3):260–263, July 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Osborne:1960:PCM

[Osb60] E. E. Osborne. On pre-conditioning of matrices. Journal of the ACM, 7(4):338–345, October 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Osborne:1961:LSS

[Osb61] E. E. Osborne. On least squares solutions of linear equations. Journal of the ACM, 8(4):628–636, October 1961. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). Cited in Åke Björck’s bibliography on least squares, which is available

by anonymous ftp from math.liu.se in pub/references.

ODonnell:2022:FP

[OST22] Ryan O’Donnell, Rocco A. Servedio, and Li-Yang Tan. Fooling polytopes. Journal of the ACM, 69(2):9:1–9:37, April 2022. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3460532.

OHearn:1995:PLV

[OT95] P. W. O’Hearn and R. D. Tennent. Parametricity and local variables. Journal of the ACM, 42(3):658–709, May 1995. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
210425.html.

Otto:2012:HAG

[Ott12] Martin Otto. Highly acyclic groups, hypergraph covers, and the guarded fragment. Journal of the ACM, 59(1):5:1–5:40, February 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Olver:2020:SFS

[OV20] Neil Olver and László A. Végh. A simpler and faster strongly polynomial algorithm for generalized flow maximization. Journal of the ACM, 67(2):10:1–10:26, May 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3383454.

Orlin:2023:DSP

[OV23] James B. Orlin and László Végh. Directed shortest paths via approximate cost balancing. Journal of the ACM, 70(1):3:1–3:??, February 2023. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL https://dl.acm.org/doi/10.
1145/3565019.

Overheu:1966:AMS

[Ove66] D. L. Overheu. An abstract machine for symbolic computation. Journal of the ACM, 13(3):444–468, July 1966. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Overbeek:1974:NCA

[Ove74] Ross A. Overbeek. A new class of automated theorem-proving algorithms. Journal of the ACM, 21(2):191–200, April 1974. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ovens:2024:SCC

[Ove24] Sean Ovens. The space complexity of consensus from swap. Journal of the ACM, 71(1):1:1–1:??, February 2024. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3631390.

Ozveren:1991:SSD

[ÖWA91] Cüneyt M. Özveren, Alan S. Willsky, and Panos J. Antsaklis. Stability and stabilizability of discrete event dynamic systems. Journal of the ACM, 38(3):730–752, July 1991. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
116855.html.

Oyamaguchi:1987:EPR

[Oya87] Michio Oyamaguchi. The equivalence problem for real-time DP-DAs. Journal of the ACM, 34(3):731–760, July 1987. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
28881.html.

Pacholski:2010:SCP

[Pac10] Witold Charatonik Leszek Pacholski. Set constraints with projections. Journal of the ACM, 57(4):23:1–23:37, April 2010. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pager:1970:EA

[Pag70] David Pager. On the efficiency of algorithms. Journal of the ACM, 17(4):708–714, October 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pager:1974:FRP

[Pag74] David Pager. Further results on the problem of finding minimal length programs for decision tables. Journal of the ACM, 21 (2):207–212, April 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pagano:1976:LCC

[Pag76] Marcello Pagano. On the linear convergence of a covariance factorization algorithm. Journal of the ACM, 23(2):310–316, April 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Papadimitriou:1976:CET

[Pap76] Christos H. Papadimitriou. On the complexity of edge traversing. Journal of the ACM, 23(3):544–554, July 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Papadimitriou:1979:OFF

[Pap79a] Christos H. Papadimitriou. Optimality of the fast Fourier transform. Journal of the ACM, 26(1):95–102, January 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Papadimitriou:1979:SCD

[Pap79b] Christos H. Papadimitriou. The serializability of concurrent database updates. Journal of the ACM, 26(4):631–653, October 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Papadimitriou:1981:CIP

[Pap81] Christos H. Papadimitriou. On the complexity of integer programming. Journal of the ACM, 28(4):765–768, October 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Papadimitriou:1982:TDC

[Pap82] Christos H. Papadimitriou. A theorem in database concurrency control. Journal of the ACM, 29(4):998–1006, October 1982. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [Pap85].

Papadimitriou:1984:CUS

[Pap84] Christos H. Papadimitriou. On the complexity of unique solutions. Journal of the ACM, 31(2):392–400, April 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Papadimitriou:1985:CTD

[Pap85] Christos H. Papadimitriou. Correction to “A theorem in database concurrency control”. Journal of the ACM, 32(3):750, July 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Pap82].

Papachristou:1987:ATL

[Pap87] Christos A. Papachristou. Associative table lookup processing for multioperand residue arithmetic. Journal of the ACM, 34 (2):376–396, April 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/23017.html.

Parsons:1955:SDC

[Par55] Frances L. Parsons. A simple desk-calculator method for checking binary results of digital computer arithmetic operations. Journal of the ACM, 2(3):205–207, July 1955. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Parter:1961:SCR

[Par61] Seymour V. Parter. Some computational results on “two-line” iterative methods for the biharmonic difference equation. Journal of the ACM, 8(3):359–365, July 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Parikh:1966:CFL

[Par66] Rohit J. Parikh. On context-free languages. Journal of the ACM, 13(4):570–581, October 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Parchmann:1979:CSM

[Par79] Rainer Parchmann. Control system model for critically timed sources. Journal of the ACM, 26(1):1–5, January 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [Par80].

Parchmann:1980:CCS

[Par80] R. Parchmann. Corrigendum: “Control system model for critically timed sources”. Journal of the ACM, 27(2):402, April 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Par79].

Pavlidis:1972:LCF

[Pav72] T. Pavlidis. Linear and context-free graph grammars. Journal of the ACM, 19(1):11–22, January 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pawlikowski:1980:MWT

[Paw80] Krzysztof Pawlikowski. Message waiting time in a packet switching system. Journal of the ACM, 27(1):30–41, January 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Platzman:1989:SCP

[PB89] Loren K. Platzman and John J. Bartholdi, III. Spacefilling curves and the planar travelling salesman problem. Journal of the ACM, 36(4):719–737, October 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/76361.html.

Pong:1998:FVC

[PD98] Fong Pong and Michel Dubois. Formal verification of complex coherence protocols using symbolic state models. Journal of the ACM, 45(4):557–587, July 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:/

/www.acm.org:80/pubs/citations/journals/jacm/1998-45-
4/p557-pong/.

Pease:1967:MIU

[Pea67] Marshall C. Pease. Matrix inversion using parallel processing. Journal of the ACM, 14(4):757–764, October 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pease:1968:AFF

[Pea68] Marshall C. Pease. An adaptation of the fast Fourier transform for parallel processing. Journal of the ACM, 15(2):252–264, April 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pease:1969:IMP

[Pea69a] Marshall C. Pease. Inversion of matrices by partitioning. Journal of the ACM, 16(2):302–314, April 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pease:1969:OLS

[Pea69b] Marshall C. Pease. Organization of large scale Fourier processors. Journal of the ACM, 16(3):474–482, July 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Peleg:1987:CDL

[Pel87] David Peleg. Concurrent dynamic logic. Journal of the ACM, 34 (2):450–479, April 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/23008.html.

Pellegrini:1998:EFS

[Pel98] Marco Pellegrini. Electrostatic fields without singularities: theory, algorithms and error analysis. Journal of the ACM, 45(6):924–964, November 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org:80/pubs/citations/journals/jacm/1998-45-

6/p924-pellegrini/. The author compares his methods against fast multipole methods for point-to-volume integrals.

Penney:1965:BSC

[Pen65] Walter Penney. A “binary” system for complex numbers. Journal of the ACM, 12(2):247–248, April 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Perley:1954:ASG

[Per54] Richmond Perley. Automatic strain-gage and thermocouple recording on punched cards. Journal of the ACM, 1(1):36–43, January 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Perkins:1956:EPC

[Per56] Robert Perkins. EASIAC, a pseudo-computer. Journal of the ACM, 3(2):65–72, April 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Perlis:1967:SAS

[Per67] Alan J. Perlis. The synthesis of algorithmic systems. Journal of the ACM, 14(1):1–9, January 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Peterka:1962:MOS

[Pet62] James J. Peterka. A method for obtaining specific values of compiling-parameter functions. Journal of the ACM, 9(3):379– 386, July 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Peterson:1976:TPL

[Pet76] G. E. Peterson. Theorem proving with lemmas. Journal of the ACM, 23(4):573–581, October 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pettie:2015:SBD

[Pet15] Seth Pettie. Sharp bounds on Davenport–Schinzel sequences of every order. Journal of the ACM, 62(5):36:1–36:??, November 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pippenger:1979:RAC

[PF79] Nicholas Pippenger and Michael J. Fischer. Relations among complexity measures. Journal of the ACM, 26(2):361–381, April 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pfaltz:1972:GS

[Pfa72] John L. Pfaltz. Graph structures. Journal of the ACM, 19(3): 411–422, July 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pfaltz:1975:RGK

[Pfa75] John L. Pfaltz. Representing graphs by Knuth trees. Journal of the ACM, 22(3):361–366, July 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Perl:1975:EGO

[PGE75] Y. Perl, M. R. Garey, and S. Even. Efficient generation of optimal prefix code: Equiprobable words using unequal cost letters. Journal of the ACM, 22(2):202–214, April 1975. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Potier:1976:AAC

[PGL76] D. Potier, E. Gelenbe, and J. Lenfant. Adaptive allocation of central processing unit quanta. Journal of the ACM, 23(1):97– 102, January 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Phillips:1962:TNS

[Phi62] David L. Phillips. A technique for the numerical solution of certain integral equations of the first kind. Journal of the ACM, 9(1):84–97, January 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Phuong:1976:SIP

[Phu76] Ta Huu Phuong. Solution of integer programs with a quadratic objective function. Journal of the ACM, 23(3):468–474, July 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pickard:1964:TSS

[Pic64a] William F. Pickard. Tables for the step-by-step integration of ordinary differential equations of the first order. Journal of the ACM, 11(2):229–233, April 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pickard:1964:TGS

[Pic64b] William F. Pickard. Tables of the generalized Stirling numbers of the first kind. Journal of the ACM, 11(1):70–78, January 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pickett:1967:NCA

[Pic67] H. E. Pickett. Note concerning the algebraic theory of automata. Journal of the ACM, 14(2):382–388, April 1967. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pietrzykowski:1973:CMS

[Pie73] Tomasz Pietrzykowski. A complete mechanization of secondorder type theory. Journal of the ACM, 20(2):333–364, April 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pippenger:1978:TST

[Pip78] Nicholas Pippenger. A time-space trade-off. Journal of the ACM, 25(3):509–515, July 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pippenger:1989:ICM

[Pip89] Nicholas Pippenger. Invariance of complexity measures for networks with unreliable gates. Journal of the ACM, 36(3):531– 539, July 1989. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/77248.html.

Pitt:1989:PII

[Pit89] L. Pitt. Probabilistic inductive inference. Journal of the ACM, 36(2):383–433, April 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/62053.html. Inductive

inference machines construct total recursive functions ϕ(x) given examples of the input and output of ϕ. Probabilistic inductive inference machines are permitted coin tosses while constructing ϕ, and are only required to construct ϕ with probability p, 0 < p < 1. This paper shows a discrete hierarchy of inferability parameterized by p, for p ≤ 1/2. Any machine that can be constructed by probabilistic inference with p > 1/2 can also be constructed deterministically.

Pitts:2006:ASR

[Pit06] Andrew M. Pitts. Alpha-structural recursion and induction. Journal of the ACM, 53(3):459–506, May 2006. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Papadimitriou:1980:FSL

[PK80] Christos H. Papadimitriou and Paris C. Kanellakis. Flowshop scheduling with limited temporary storage. Journal of the ACM, 27(3):533–549, July 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pflug:1987:LPN

[PK87] Georg Ch. Pflug and Hans W. Kessler. Linear probing with a nonuniform address distribution. Journal of the ACM, 34 (2):397–410, April 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/42225.html.

Pachl:1984:LBD

[PKR84] J. Pachl, E. Korach, and D. Rotem. Lower bounds for distributed maximum-finding algorithms. Journal of the ACM, 31(4):905– 918, October 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pattipati:1990:AMV

[PKT90] K. R. Pattipati, M. M. Kostreva, and J. L. Teele. Approximate mean value analysis algorithms for queueing networks: Existence, uniqueness, and convergence results. Journal of the ACM, 37(3):643–673, July 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/214074.html.

Plandowski:2004:SWE

[Pla04] Wojciech Plandowski. Satisfiability of word equations with constants is in PSPACE. Journal of the ACM, 51(3):483–496, May 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Plemmons:1974:LLS

[Ple74] Robert J. Plemmons. Linear least squares by elimination and MGS. Journal of the ACM, 21(4):581–585, October 1974. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Philippatos:1971:ECI

[PM71] G. C. Philippatos and D. R. Moscato. Effects of constrained information on player decisions in experimental business simulation: Some empirical evidence. Journal of the ACM, 18(1):94– 104, January 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

PucciDeFarias:2006:CEA

[PM06] Daniela Pucci De Farias and Nimrod Megiddo. Combining expert advice in reactive environments. Journal of the ACM, 53(5): 762–799, September 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Puschel:2009:PSD

[PMH09] Markus Püschel, Peter A. Milder, and James C. Hoe. Permuting streaming data using RAMs. Journal of the ACM, 56(2):10:1– 10:34, April 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pollack:1964:MRC

[Pol64] Maurice Pollack. Message route control in a large teletype network. Journal of the ACM, 11(1):104–116, January 1964. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Porter:1966:UMG

[Por66] Sigmund N. Porter. Use of multiwrite for general programmability of search memories. Journal of the ACM, 13(3):369–373, July 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Potechin:2017:BMS

[Pot17] Aaron Potechin. Bounds on monotone switching networks for directed connectivity. Journal of the ACM, 64(4):29:1–29:??, September 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Paz:1965:UDS

[PP65] A. Paz and B. Peleg. Ultimate-definite and symmetric-definite events and automata. Journal of the ACM, 12(3):399–410, July 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Paturi:2005:IET

[PPSZ05] Ramamohan Paturi, Pavel Pudlák, Michael E. Saks, and Francis Zane. An improved exponential-time algorithm for k-SAT. Journal of the ACM, 52(3):337–364, May 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Prawitz:1960:MPP

[PPV60] Dag Prawitz, Håkan Prawitz, and Neri Voghera. A mechanical proof procedure and its realization in an electronic computer. Journal of the ACM, 7(2):102–128, April 1960. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pirri:1999:SCM

[PR99] Fiora Pirri and Ray Reiter. Some contributions to the metatheory of the situation calculus. Journal of the ACM, 46(3):325– 361, May 1999. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org:80/pubs/
citations/journals/jacm/1999-46-3/p325-pirri/.

Pettie:2002:OMS

[PR02] Seth Pettie and Vijaya Ramachandran. An optimal minimum spanning tree algorithm. Journal of the ACM, 49(1):16–34, January 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Papadimitriou:2008:CCE

[PR08] Christos H. Papadimitriou and Tim Roughgarden. Computing correlated equilibria in multi-player games. Journal of the ACM, 55(3):14:1–14:29, July 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Prather:1960:CAD

[Pra60] Ronald Prather. Computational aids for determining the minimal form of a truth function. Journal of the ACM, 7(4):299–310, October 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Price:1976:NEC

[Pri76] Thomas G. Price. A note on the effect of the central processor service time distribution on processor utilization in multiprogrammed computer systems. Journal of the ACM, 23(2):342– 346, April 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Proskurowski:1980:GBT

[Pro80] Andrzej Proskurowski. On the generation of binary trees. Journal of the ACM, 27(1):1–2, January 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Palmer:1974:EBC

[PRR74] E. M. Palmer, M. A. Rahimi, and R. W. Robinson. Efficiency of a binary comparison storage technique. Journal of the ACM, 21(3):376–384, July 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Purdom:1970:SPB

[PS70] Paul W. Purdom, Jr. and Stephen M. Stigler. Statistical properties of the buddy system. Journal of the ACM, 17(4):683–697, October 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Perl:1978:FTD

[PS78] Y. Perl and Y. Shiloach. Finding two disjoint paths between two pairs of vertices in a graph. Journal of the ACM, 25(1): 1–9, January 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Prabhala:1980:ECE

[PS80] Bhaskaram Prabhala and Ravi Sethi. Efficient computation of expressions with common subexpressions. Journal of the ACM, 27(1):146–163, January 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Perl:1981:MMT

[PS81a] Yehoshua Perl and Stephen R. Schach. Max-min tree partitioning. Journal of the ACM, 28(1):5–15, January 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Peterson:1981:CSR

[PS81b] Gerald E. Peterson and Mark E. Stickel. Complete sets of reductions for some equational theories. Journal of the ACM, 28 (2):223–264, April 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pierce:2000:BEP

[PS00] Benjamin C. Pierce and Davide Sangiorgi. Behavioral equivalence in the polymorphic pi-calculus. Journal of the ACM, 47 (3):531–584, May 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/citations/journals/jacm/2000-47-5/p531-pierce/.

Pease:1980:RAP

[PSL80] M. Pease, R. Shostak, and L. Lamport. Reaching agreements in the presence of faults. Journal of the ACM, 27(2):228–234, April 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). This paper is similar to their 1982 publication [LSP82], but contains a rigorous proof of the impossibility of Byzantine agreement for the case n = 3, t = 1. As usual, n is the total number of processes and t is the number of faulty processes.

Peres:2013:APS

[PSSZ13] Yuval Peres, Dmitry Sotnikov, Benny Sudakov, and Uri Zwick. All-pairs shortest paths in O(n^{2}) time with high probability. Journal of the ACM, 60(4):26:1–26:25, August 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pope:1957:MFR

[PT57] David A. Pope and C. Tompkins. Maximizing functions of rotations—experiments concerning speed of diagonalization of symmetric matrices using Jacobi’s method. Journal of the ACM, 4(4):459–466, October 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Patrascu:2012:PST

[PT12] Mihai Pătraşcu and Mikkel Thorup. The power of simple tabulation hashing. Journal of the ACM, 59(3):14:1–14:50, June 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Platzer:2020:DEI

[PT20] André Platzer and Yong Kiam Tan. Differential equation invariance axiomatization. Journal of the ACM, 67(1):6:1–6:66, April 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3380825.

Panwar:1988:OSP

[PTW88] Shivendra S. Panwar, Don Towsley, and Jack K. Wolf. Optimal scheduling policies for a class of queues with customer deadlines to the beginning of service. Journal of the ACM, 35(4):832– 844, October 1988. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/48019.html.

Peleg:1989:TBS

[PU89] David Peleg and Eli Upfal. A trade-off between space and efficiency for routing tables. Journal of the ACM, 36(3):510– 530, July 1989. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/65953.html.

Pulvari:1955:MMU

[Pul55] Charles F. Pulvari. Memory matrix using ferroelectric condensers as bistable elements. Journal of the ACM, 2(3):169–185, July 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pippenger:1976:SGT

[PV76] Nicholas Pippenger and Leslie G. Valiant. Shifting graphs and their applications. Journal of the ACM, 23(3):423–432, July 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pitt:1988:CLL

[PV88] Leonard Pitt and Leslie G. Valiant. Computational limitations on learning from examples. Journal of the ACM, 35(4):965– 984, October 1988. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/63140.html.

Pitt:1993:MCD

[PW93] Leonard Pitt and Manfred K. Warmuth. The minimum consistent DFA problem cannot be approximated within any polynomial. Journal of the ACM, 40(1):95–142, January 1993. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
138042.html.

Papadimitriou:1982:CRS

[PY82] Christos H. Papadimitriou and Mihalis Yannakakis. The complexity of restricted spanning tree problems. Journal of the ACM, 29(2):285–309, April 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Pyle:1964:GIC

[Pyl64] L. Duane Pyle. Generalized inverse computations using the gradient projection method. Journal of the ACM, 11(4):422–428, October 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Place:2019:GHF

[PZ19] Thomas Place and Marc Zeitoun. Going higher in first-order quantifier alternation hierarchies on words. Journal of the ACM, 66(2):12:1–12:??, April 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/ft_gateway.cfm?id=3303991.

Quinlan:1968:FDP

[QH68] J. R. Quinlan and E. B. Hunt. A formal deductive problemsolving system. Journal of the ACM, 15(4):625–646, October 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rosberg:1976:MQE

[RA76] Z. Rosberg and I. Adiri. Multilevel queues with external priorities. Journal of the ACM, 23(4):680–690, October 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Raghvendra:2020:NLT

[RA20] Sharath Raghvendra and Pankaj K. Agarwal. A near-linear time ϵ-approximation algorithm for geometric bipartite matching. Journal of the ACM, 67(3):18:1–18:19, June 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3393694.

Rabinowitz:1966:NEC

[Rab66] Philip Rabinowitz. Numerical experiments in conformal mapping by the method of orthonormal polynomials. Journal of the ACM, 13(2):296–303, April 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rabin:1989:EDI

[Rab89] Michael O. Rabin. Efficient dispersal of information for security, load balancing, and fault tolerance. Journal of the ACM, 36(2):335–348, April 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/62050.html.

Rabin:1994:RSS

[Rab94] Tal Rabin. Robust sharing of secrets when the dealer is honest or cheating. Journal of the ACM, 41(6):1089–1109, November 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/195621.html.

Rackoff:1982:RQI

[Rac82] Charles Rackoff. Relativized questions involving probabilistic algorithms. Journal of the ACM, 29(1):261–268, January 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Rackoff attempts to prove R ≠ P by assuming P ≠ NP and relativization (i.e., for a class of languages C, C^{A} is the same as C except that one can answer questions concerning membership in A in constant time). Interestingly, he proves that for some oracle A, P^{A} ≠ NP^{A} and R^{A} ≠ P^{A}, and at the same time, for some other oracle B, P^{B} ≠ NP^{B} and R^{B} ≠ P^{B}. An earlier version of this paper appeared in Proc. 10th Ann. ACM Symp. on Theory of Computing, 1978, pp. 338–342.

Raghavan:2003:EPE

[Rag03] Prabhakar Raghavan. Editorial: Preserving excellence through change. Journal of the ACM, 50(4):427–428, July 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rahme:1969:NLN

[Rah69] H. S. Rahme. A new look at the numerical integration of ordinary differential equations. Journal of the ACM, 16(3):496–506, July 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rahme:1970:SAN

[Rah70] H. S. Rahme. Stability analysis of a new algorithm used for integrating a system of ordinary differential equations. Journal of the ACM, 17(2):284–293, April 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ralston:1959:FQF

[Ral59] A. Ralston. A family of quadrature formulas which achieve high accuracy in composite rules. Journal of the ACM, 6(3):384–394, July 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ralston:1963:ERF

[Ral63] Anthony Ralston. On economization of rational functions. Journal of the ACM, 10:279–282, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ramamoorthy:1966:AGC

[Ram66] C. V. Ramamoorthy. Analysis of graphs by connectivity considerations. Journal of the ACM, 13(2):211–222, April 1966. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ramakrishnan:1980:STC

[Ram80] K. G. Ramakrishnan. Solving two-commodity transportation problems with coupling constraints. Journal of the ACM, 27(4): 736–757, October 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ramachandran:1986:DML

[Ram86] Vijaya Ramachandran. On driving many long wires in a VLSI layout. Journal of the ACM, 33(4):687–701, October 1986. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
6784.html.

Ramshaw:1988:EGT

[Ram88] Lyle Ramshaw. Eliminating go to’s while preserving program structure. Journal of the ACM, 35(4):893–920, October 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/48021.html.

Raney:1958:SF

[Ran58] George N. Raney. Sequential functions. Journal of the ACM, 5(2):177–180, April 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rao:1978:PAC

[Rao78] Gururaj S. Rao. Performance analysis of cache memories. Journal of the ACM, 25(3):378–395, July 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rasch:1970:QTS

[Ras70] Philip J. Rasch. A queueing theory study of round-robin scheduling of time-shared computer systems. Journal of the ACM, 17 (1):131–145, January 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Razborov:2003:PPC

[Raz03] Alexander Razborov. Propositional proof complexity. Journal of the ACM, 50(1):80–82, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Raz:2004:RLB

[Raz04] Ran Raz. Resolution lower bounds for the weak pigeonhole principle. Journal of the ACM, 51(2):115–138, March 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Raz:2009:MLF

[Raz09] Ran Raz. Multi-linear formulas for permanent and determinant are of super-polynomial size. Journal of the ACM, 56(2):8:1– 8:17, April 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Raz:2013:TRL

[Raz13] Ran Raz. Tensor-rank and lower bounds for arithmetic formulas. Journal of the ACM, 60(6):40:1–40:??, November 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Razborov:2016:NKT

[Raz16] Alexander Razborov. A new kind of tradeoffs in propositional proof complexity. Journal of the ACM, 63(2):16:1–16:??, May 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Raz:2019:FLR

[Raz19] Ran Raz. Fast learning requires good memory: a time-space lower bound for parity learning. Journal of the ACM, 66(1):3:1– 3:??, January 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rahli:2019:BIC

[RBCC19] Vincent Rahli, Mark Bickford, Liron Cohen, and Robert L. Constable. Bar induction is compatible with constructive type theory. Journal of the ACM, 66(2):13:1–13:??, April 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/ft_gateway.cfm?id=3305261.

Reynolds:1969:SGR

[RC69] B. G. Reynolds and W. F. Cutlip. Synchronization and general repetitive machines, with applications to ultimate definite automata. Journal of the ACM, 16(2):226–234, April 1969. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ramamoorthy:1970:OMH

[RC70] C. V. Ramamoorthy and K. M. Chandy. Optimization of memory hierarchies in multiprogrammed systems. Journal of the ACM, 17(3):426–445, July 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rowland:1986:SSA

[RC86] John H. Rowland and John R. Cowles. Small sample algorithms for the identification of polynomials. Journal of the ACM, 33 (4):822–829, October 1986. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rowland:1981:UTP

[RD81] John H. Rowland and Philip J. Davis. On the use of transcendentals for program testing. Journal of the ACM, 28(1):181–190, January 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Reddy:2003:TOP

[Red03] Raj Reddy. Three open problems in AI. Journal of the ACM, 50(1):83–86, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Reed:2003:HRB

[Ree03] Bruce Reed. The height of a random binary search tree. Journal of the ACM, 50(3):306–332, May 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Regev:2004:NLB

[Reg04] Oded Regev. New lattice-based cryptographic constructions. Journal of the ACM, 51(6):899–942, November 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Regev:2009:LLE

[Reg09] Oded Regev. On lattices, learning with errors, random linear codes, and cryptography. Journal of the ACM, 56(6):34:1–34:40, September 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Regev:2025:EQF

[Reg25] Oded Regev. An efficient quantum factoring algorithm. Journal of the ACM, 72(1):10:1–10:??, February 2025. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3708471.

Reiter:1968:SPC

[Rei68] Raymond Reiter. Scheduling parallel computations. Journal of the ACM, 15(4):590–599, October 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Reiter:1971:TRO

[Rei71] Raymond Reiter. Two results on ordering for resolution with merging and linear format. Journal of the ACM, 18(4):630–646, October 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Reingold:1972:OSS

[Rei72] Edward M. Reingold. On the optimality of some set algorithms. Journal of the ACM, 19(4):649–659, October 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Reiss:1979:SDC

[Rei79] Steven P. Reiss. Security in databases; a combinatorial study. Journal of the ACM, 26(1):45–57, January 1979. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Reischuk:1980:IBP

[Rei80a] Rüdiger Reischuk. Improved bounds on the problem of timespace trade-off in the pebble game. Journal of the ACM, 27(4): 839–849, October 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Reiter:1980:EDC

[Rei80b] Raymond Reiter. Equality and domain closure for first-order databases. Journal of the ACM, 27(2):235–249, April 1980. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Reif:1984:SC

[Rei84] John H. Reif. Symmetric complementation. Journal of the ACM, 31(2):401–421, April 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Reiter:1986:SSC

[Rei86] Raymond Reiter. A sound and sometimes complete query evaluation algorithm for relational databases with null values. Journal of the ACM, 33(2):349–370, April 1986. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/5388.html.

Reischuk:1987:SWP

[Rei87] Rüdiger Reischuk. Simultaneous WRITES of parallel random access machines do not help to compute simple arithmetic functions. Journal of the ACM, 34(1):163–178, January 1987. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
22944.html.

Reingold:2008:UCL

[Rei08] Omer Reingold. Undirected connectivity in log-space. Journal of the ACM, 55(4):17:1–17:24, September 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rem:1981:CSP

[Rem81] Martin Rem. The closure statement: a programming language construct allowing ultraconcurrent execution. Journal of the ACM, 28(2):393–410, April 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Reynolds:1977:SDF

[Rey77] John C. Reynolds. Semantics of the domain of flow diagrams. Journal of the ACM, 24(3):484–503, July 1977. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rigal:1967:CGS

[RG67] J. L. Rigal and J. Gaches. On the compatibility of a given solution with the data of a linear system. Journal of the ACM, 14(3):543–548, July 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Richards:1973:EESb

[Ric73a] Donald L. Richards. Efficient exercising of switching elements in combinational nets. Journal of the ACM, 20(2):320–332, April 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Richards:1973:EESa

[Ric73b] Donald L. Richards. Efficient exercising of switching elements in nets of identical gates. Journal of the ACM, 20(1):88–111, January 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Richman:1974:CSI

[Ric74] Paul L. Richman. Computing a subinterval of the image. Journal of the ACM, 21(3):454–458, July 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rice:1975:MAQ

[Ric75] John R. Rice. A metalgorithm for adaptive quadrature. Journal of the ACM, 22(1):61–82, January 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rider:1976:SAA

[Rid76] Kenneth Lloyd Rider. A simple approximation to the average queue size in the time-dependent M/M/1 queue. Journal of the ACM, 23(2):361–367, April 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ritchie:1963:FAS

[Rit63] Robert W. Ritchie. Finite automata and the set of squares. Journal of the ACM, 10:528–531, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rivest:1978:OAK

[Riv78] Ronald L. Rivest. Optimal arrangement of keys in a hash table. Journal of the ACM, 25(2):200–209, April 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rinberg:2023:IVL

[RK23] Arik Rinberg and Idit Keidar. Intermediate value linearizability: a quantitative correctness criterion. Journal of the ACM, 70(2):17:1–17:??, April 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3584699.

Ross:1975:CSE

[RKN75] G. Terry Ross, D. Klingman, and A. Napier. A computational study of the effects of problem dimensions on solution times for transportation problems. Journal of the ACM, 22(3):413–424, July 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Reiser:1980:MVA

[RL80] M. Reiser and S. S. Lavenberg. Mean-value analysis of closed multichain queuing networks. Journal of the ACM, 27(2):313– 322, April 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [RL81].

Reiser:1981:CMV

[RL81] M. Reiser and S. S. Lavenberg. Corrigendum: “Mean-Value analysis of closed multichain queuing networks”. Journal of the ACM, 28(3):629, July 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [RL80].

Rodrigue:1979:OER

[RMK79] Garry H. Rodrigue, Niel K. Madsen, and Jack I. Karush. Oddeven reductions for banded linear equations. Journal of the ACM, 26(1):72–81, January 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Robinson:1963:TPC

[Rob63] J. A. Robinson. Theorem-proving on the computer. Journal of the ACM, 10:163–174, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Robinson:1965:MOL

[Rob65] J. A. Robinson. A machine-oriented logic based on the resolution principle. Journal of the ACM, 12(1):23–41, January 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Robson:1971:ESS

[Rob71] J. M. Robson. An estimate of the store size necessary for dynamic storage allocation. Journal of the ACM, 18(3):416–423, July 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Related to online graph coloring.

Robson:1974:BSF

[Rob74] J. M. Robson. Bounds for some functions concerning dynamic storage allocation. Journal of the ACM, 21(3):491–499, July 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rock:1984:TMN

[Röc84] Hans Röck. The three-machine no-wait flow shop is NPcomplete. Journal of the ACM, 31(2):336–345, April 1984. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rockmore:1994:ECF

[Roc94] Daniel N. Rockmore. Efficient computation of Fourier inversion for finite groups. Journal of the ACM, 41(1):31–66, January 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/174646.html.

Rosenberg:1967:RTD

[Ros67a] Arnold L. Rosenberg. Real-time definable languages. Journal of the ACM, 14(4):645–662, October 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rosenkrantz:1967:MEN

[Ros67b] Daniel J. Rosenkrantz. Matrix equations and normal forms for context-free grammars. Journal of the ACM, 14(3):501–507, July 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rosenberg:1968:IRT

[Ros68] Arnold L. Rosenberg. On the independence of real-time definability and certain structural properties of context-free languages. Journal of the ACM, 15(4):672–679, October 1968. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rosenkrantz:1969:PGC

[Ros69] Daniel J. Rosenkrantz. Programmed grammars and classes of formal languages. Journal of the ACM, 16(1):107–131, January 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rosenberg:1970:NAC

[Ros70a] Arnold L. Rosenberg. A note on ambiguity of context-free languages and presentations of semilinear sets. Journal of the ACM, 17(1):44–50, January 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rosenfeld:1970:CDP

[Ros70b] Azriel Rosenfeld. Connectivity in digital pictures. Journal of the ACM, 17(1):146–160, January 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rose:1971:NCO

[Ros71] Donald J. Rose. A note on consistent ordering and zero circulation. Journal of the ACM, 18(4):573–575, October 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rosenberg:1972:ADG

[Ros72] Arnold L. Rosenberg. Addressable data graphs. Journal of the ACM, 19(2):309–340, April 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rosen:1973:TMS

[Ros73a] Barry K. Rosen. Tree-manipulation systems and Church–Rosser theorems. Journal of the ACM, 20(1):160–187, January 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rosenfeld:1973:ACD

[Ros73b] Azriel Rosenfeld. Arcs and curves in digital pictures. Journal of the ACM, 20(1):81–87, January 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rosenberg:1974:ASE

[Ros74] Arnold L. Rosenberg. Allocating storage for extendible arrays. Journal of the ACM, 21(4):652–670, October 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [Ros75].

Rosenberg:1975:CAS

[Ros75] Arnold L. Rosenberg. Corrigendum: “Allocating storage for extendible arrays”. Journal of the ACM, 22(2):308, April 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Ros74].

Rosen:1979:DFA

[Ros79a] Barry K. Rosen. Data flow analysis for procedural languages. Journal of the ACM, 26(2):322–344, April 1979. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rosenberg:1979:EDS

[Ros79b] Arnold L. Rosenberg. Encoding data structures in trees. Journal of the ACM, 26(4):668–689, October 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rosenberg:1983:TDV

[Ros83] Arnold L. Rosenberg. Three-dimensional VLSI: a case study. Journal of the ACM, 30(3):397–416, July 1983. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ross:1994:MSM

[Ros94] Kenneth A. Ross. Modular stratification and magic sets for datalog programs with negation. Journal of the ACM, 41(6): 1216–1266, November 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/195646.html.

Rosenkrantz:2003:J

[Ros03] Daniel J. Rosenkrantz. JACM 1986–1990. Journal of the ACM, 50(1):18, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rossman:2008:HPT

[Ros08] Benjamin Rossman. Homomorphism preservation theorems. Journal of the ACM, 55(3):15:1–15:53, July 2008. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rotenburg:1960:NPR

[Rot60] A. Rotenburg. A new pseudo-random number generator. Journal of the ACM, 7(1):75–77, January 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Roth:1970:ASL

[Rot70] Richard H. Roth. An approach to solving linear discrete optimization problems. Journal of the ACM, 17(2):303–313, April 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rothvoss:2017:MPE

[Rot17] Thomas Rothvoss. The matching polytope has exponential extension complexity. Journal of the ACM, 64(6):41:1–41:??, November 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Roura:2001:IMT

[Rou01] Salvador Roura. Improved master theorems for divide-and-conquer recurrences. Journal of the ACM, 48(2):170–205, March 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/citations/
journals/jacm/2001-48-2/p170-roura/.

Roughgarden:2015:IRP

[Rou15] Tim Roughgarden. Intrinsic robustness of the price of anarchy. Journal of the ACM, 62(5):32:1–32:??, November 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Roughgarden:2024:TFM

[Rou24] Tim Roughgarden. Transaction fee mechanism design. Journal of the ACM, 71(4):30:1–30:??, August 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3674143.

Rowan:1957:PTS

[Row57] T. C. Rowan. Psychological tests and selection of computer programmers. Journal of the ACM, 4(3):348–353, July 1957. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rosenfeld:1966:SOD

[RP66] Azriel Rosenfeld and John L. Pfaltz. Sequential operations in digital picture processing. Journal of the ACM, 13(4):471–494, October 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rosenthal:1989:GAC

[RP89] Arnon Rosenthal and José A. Pino. A generalized algorithm for centrality problems on trees. Journal of the ACM, 36(2):349– 361, April 1989. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/62051.html.

Rodeh:1981:LAD

[RPE81] Michael Rodeh, Vaughan R. Pratt, and Shimon Even. Linear algorithm for data compression via string matching. Journal of the ACM, 28(1):16–24, January 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Randell:1964:SST

[RR64] B. Randell and L. J. Russell. Single-scan techniques for the translation of arithmetic expressions in ALGOL 60. Journal of the ACM, 11(2):159–167, April 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ramesh:1992:NPM

[RR92] R. Ramesh and I. V. Ramakrishnan. Nonlinear pattern matching in trees. Journal of the ACM, 39(2):295–316, April 1992. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
128752.html.

Rich:1963:MFA

[RS63] Robert P. Rich and Harry Shaw. A method for finding all the zeros of f(z). Journal of the ACM, 10:545–549, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Reinwald:1966:CLE

[RS66] Lewis T. Reinwald and Richard M. Soland. Conversion of limited-entry decision tables to optimal computer programs I: Minimum average processing time. Journal of the ACM, 13(3): 339–358, July 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Reinwald:1967:CLE

[RS67] Lewis T. Reinwald and Richard M. Soland. Conversion of limited-entry decision tables to optimal computer programs II: Minimum storage requirement. Journal of the ACM, 14(4):742– 756, October 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rosenberg:1977:HSE

[RS77] Arnold L. Rosenberg and Larry J. Stockmeyer. Hashing schemes for extendible arrays. Journal of the ACM, 24(2):199–221, April 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Raoult:1983:PNC

[RS83] Jean-Claude Raoult and Ravi Sethi. Properties of a notation for combining functions. Journal of the ACM, 30(3):595–611, July 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Reif:1994:MPP

[RS94a] John Reif and Micha Sharir. Motion planning in the presence of moving obstacles. Journal of the ACM, 41(4):764–790, July 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/179911.html.

Reif:1994:SEU

[RS94b] John H. Reif and James A. Storer. A single-exponential upper bound for finding shortest paths in three dimensions. Journal of the ACM, 41(5):1013–1019, September 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/185811.html.

Rivest:1994:DBI

[RS94c] Ronald L. Rivest and Robert E. Schapire. Diversity-based inference of finite automata. Journal of the ACM, 41(3):555–589, May 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/176589.html.

Rodeh:1999:FCA

[RS99] Michael Rodeh and Mooly Sagiv. Finding circular attributes in attribute grammars. Journal of the ACM, 46(4):556–575, July 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL http://www.acm.org/pubs/articles/
journals/jacm/1999-46-4/p556-rodeh/p556-rodeh.pdf;
http://www.acm.org/pubs/citations/journals/jacm/1999-
46-4/p556-rodeh/.

Rawling:2000:OQL

[RS00] J. P. Rawling and S. A. Selesnick. Orthologic and quantum logic: models and computational elements. Journal of the ACM, 47 (4):721–751, July 2000. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/citations/journals/jacm/2000-47-4/p721-rawling/.

Remy:2009:QPT

[RS09a] Jan Remy and Angelika Steger. A quasi-polynomial time approximation scheme for minimum weight triangulation. Journal of the ACM, 56(3):15:1–15:47, May 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Roughgarden:2009:QIC

[RS09b] Tim Roughgarden and Mukund Sundararajan. Quantifying inefficiency in cost-sharing mechanisms. Journal of the ACM, 56 (4):23:1–23:33, June 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ross:2005:AOP

[RSG05] Robert Ross, V. S. Subrahmanian, and John Grant. Aggregate operators in probabilistic databases. Journal of the ACM, 52 (1):54–101, January 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Reed:1969:GSR

[RT69] I. S. Reed and Rein Turn. A generalization of shift-register sequence generators. Journal of the ACM, 16(3):461–473, July 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rhee:1988:SDA

[RT88] Wansoo T. Rhee and Michel Talagrand. Some distributions that allow perfect packing. Journal of the ACM, 35(3):564–578, July 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/44487.html.

Ryu:1990:ADP

[RT90] In Kyung Ryu and Alexander Thomasian. Analysis of database performance with dynamic locking. Journal of the ACM, 37(3): 491–523, July 1990. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/79152.html.

Roughgarden:2002:HBS

[RT02] Tim Roughgarden and Éva Tardos. How bad is selfish routing? Journal of the ACM, 49(2):236–259, March 2002. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Raz:2022:OSB

[RT22] Ran Raz and Avishay Tal. Oracle separation of BQP and PH. Journal of the ACM, 69(4):30:1–30:??, August 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3530258.

Ross:1994:MCS

[RTW94] Keith W. Ross, Danny H. K. Tsang, and Jie Wang. Monte Carlo summation and integration applied to multiclass queuing networks. Journal of the ACM, 41(6):1110–1135, November 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/195630.html.

Rubinoff:1955:DCR

[Rub55] Morris Rubinoff. Digital computers for real-time simulation. Journal of the ACM, 2(3):186–204, July 1955. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rubin:1974:SPH

[Rub74] Frank Rubin. A search procedure for Hamilton paths and circuits. Journal of the ACM, 21(4):576–580, October 1974. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rubin:2015:KDT

[Rub15] Natan Rubin. On kinetic Delaunay triangulations: a nearquadratic bound for unit speed motions. Journal of the ACM, 62 (3):25:1–25:??, June 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rubin:2022:IBW

[Rub22] Natan Rubin. An improved bound for weak epsilon-nets in the plane. Journal of the ACM, 69(5):32:1–32:??, October 2022. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3555985.

Ruschitzka:1982:PJC

[Rus82] Manfred Ruschitzka. The performance of job classes with distinct policy functions. Journal of the ACM, 29(2):514–526, April 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rutishauser:1955:SPT

[Rut55] Heinz Rutishauser. Some programming techniques for the ER-METH. Journal of the ACM, 2(1):1–4, January 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rutledge:1964:IPS

[Rut64] J. D. Rutledge. On Ianov’s program schemata. Journal of the ACM, 11(1):1–9, January 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rotem:1978:GBT

[RV78] Doron Rotem and Y. L. Varol. Generation of binary trees from ballot sequences. Journal of the ACM, 25(3):396–404, July 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Raoult:1980:OSE

[RV80] Jean-Claude Raoult and Jean Vuillemin. Operational and semantic equivalence between recursive programs. Journal of the ACM, 27(4):772–796, October 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Reif:1987:LTS

[RV87] John H. Reif and Leslie G. Valiant. A logarithmic time sort for linear size networks. Journal of the ACM, 34(1):60–76, January 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/7532.html.

Rudelson:2007:SLM

[RV07] Mark Rudelson and Roman Vershynin. Sampling from large matrices: an approach through geometric functional analysis. Journal of the ACM, 54(4):21:1–21:19, July 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Roughgarden:2018:SCL

[RVW18] Tim Roughgarden, Sergei Vassilvitskii, and Joshua R. Wang. Shuffles and circuits (on lower bounds for modern parallel computation). Journal of the ACM, 65(6):41:1–41:??, November 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rabin:1963:WHT

[RW63] Michael O. Rabin and Hao Wang. Words in the history of a Turing machine with a fixed input. Journal of the ACM, 10: 526–527, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rohlicek:1988:RPM

[RW88] Jan Robin Rohlicek and Alan S. Willsky. The reduction of perturbed Markov generators: An algorithm exposing the role of transient states. Journal of the ACM, 35(3):675–696, July 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/44497.html.

Raz:1992:MCM

[RW92] Ran Raz and Avi Wigderson. Monotone circuits for matching require linear depth. Journal of the ACM, 39(3):736–744, July 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/146684.html.

Rutschmann:2023:CKC

[RW23] Daniel Rutschmann and Manuel Wettstein. Chains, Koch chains, and point sets with many triangulations. Journal of the ACM, 70(3):18:1–18:??, June 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3585535.

Rathmann:1994:CHS

[RWM94] Peter K. Rathmann, Marianne Winslett, and Mark Manasse. Circumscription with homomorphisms: Solving the equality and counterexample problems. Journal of the ACM, 41(5):819–873, September 1994. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/185678.html.

Romero:2023:PAM

[RWZ23] Miguel Romero, Marcin Wrochna, and Stanislav Zivný. Pliability and approximating Max-CSPs. Journal of the ACM, 70 (6):41:1–41:??, December 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3626515.

Ross:1991:OLB

[RY91] Keith W. Ross and David D. Yao. Optimal load balancing and scheduling in a distributed computer system. Journal of the ACM, 38(3):676–690, July 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/116847.html.

Rao:1965:MMA

[RZ65] T. R. N. Rao and N. Zierler. On mappings for modular arithmetic, I. Journal of the ACM, 12(4):542–544, October 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Romashchenko:2019:OCM

[RZ19] Andrei Romashchenko and Marius Zimand. An operational characterization of mutual information in algorithmic information theory. Journal of the ACM, 66(5):38:1–38:??, September 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=
3356867.

Ron-Zewi:2024:LPA

[RZR24] Noga Ron-Zewi and Ron Rothblum. Local proofs approaching the witness length. Journal of the ACM, 71(3):18:1–18:??, June 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3661483.

Ron-Zewi:2025:PFC

[RZR25] Noga Ron-Zewi and Ron Rothblum. Proving as fast as computing: Succinct arguments with constant prover overhead. Journal of the ACM, 72(2):15:1–15:??, April 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sacca:1985:CDH

[Sac85] Domenico Saccà. Closures of database hypergraphs. Journal of the ACM, 32(4):774–803, October 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/4997.html.

Sagiv:1980:AIM

[Sag80] Yehoshua Sagiv. An algorithm for inferring multivalued dependencies with an application to propositional logic. Journal of the ACM, 27(2):250–262, April 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sagiv:1991:EQI

[Sag91] Yehoshua Sagiv. Evaluation of queries in independent database schemes. Journal of the ACM, 38(1):120–161, January 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/102787.html.

Singh:1994:EAR

[SAG94] Ambuj K. Singh, James H. Anderson, and Mohamed G. Gouda. The elusive atomic register. Journal of the ACM, 41(2):311– 339, March 1994. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/174657.html.

Sahni:1975:AAK

[Sah75] Sartaj Sahni. Approximate algorithms for the 0/1 knapsack problem. Journal of the ACM, 22(1):115–124, January 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sahni:1976:ASI

[Sah76] Sartaj K. Sahni. Algorithms for scheduling independent tasks. Journal of the ACM, 23(1):116–127, January 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Smith:1964:DCS

[SAKS64] K. U. Smith, S. D. Ansell, J. Koehler, and G. H. Servos. Digital computer system for dynamic analysis of speech and sound feedback mechanisms. Journal of the ACM, 11(2):240–251, April 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Salton:1960:NMP

[Sal60] Gerard Salton. A new method for the payment of bills and the transfer of credit. Journal of the ACM, 7(2):140–149, April 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Salton:1963:ADR

[Sal63] Gerard Salton. Associative document retrieval techniques using bibliographic information. Journal of the ACM, 10:440–457, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Salomaa:1966:TCA

[Sal66] Arto Salomaa. Two complete axiom systems for the algebra of regular events. Journal of the ACM, 13(1):158–169, January 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Salton:1969:PJ

[Sal69] G. Salton. A policy for JACM. Journal of the ACM, 16(1):1– 2, January 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Salton:1970:RAJ

[Sal70] G. Salton. On the role of the ACM Journal. Journal of the ACM, 17(1):1–2, January 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Salton:1971:EST

[Sal71a] G. Salton. Editorial: Some thoughts on scientific information dissemination. Journal of the ACM, 18(1):1–3, January 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Salton:1971:I

[Sal71b] G. Salton. Introduction. Journal of the ACM, 18(4):477, October 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Salton:1972:EWC

[Sal72] G. Salton. Editorial: What is computer science? Journal of the ACM, 19(1):1–2, January 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Salton:1973:RSA

[Sal73] G. Salton. Recent studies in automatic text analysis and document retrieval. Journal of the ACM, 20(2):258–278, April 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Samet:1981:CCL

[Sam81] Hanan Samet. Connected component labeling using quadtrees. Journal of the ACM, 28(3):487–501, July 1981. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sanders:1965:DAC

[San65] Jerry Sanders. Document association and classification based on L-languages. Journal of the ACM, 12(2):249–253, April 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sandewall:1969:PPS

[San69] Erik J. Sandewall. A planning problem solver based on lookahead in stochastic game trees. Journal of the ACM, 16(3):364– 382, July 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Santos:1972:NBG

[San72] Eugene S. Santos. A note on bracketed grammars. Journal of the ACM, 19(2):222–224, April 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sandhu:1988:SPM

[San88] Ravinderpal Singh Sandhu. The schematic protection model: Its definition and analysis for acyclic attenuation schemes. Journal of the ACM, 35(2):404–432, April 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/42286.html.

Santos:1996:LPF

[San96] Eugene Santos, Jr. On linear potential functions for approximating Bayesian computations. Journal of the ACM, 43(3):399– 430, May 1996. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/jacm/233552.html.

Sarwate:1983:NNM

[Sar83] D. V. Sarwate. A note on “A note on multiple error detection in ASCII numeric data communication”. Journal of the ACM, 30(1):33–35, January 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Savage:1972:CWT

[Sav72] J. E. Savage. Computational work and time on finite machines. Journal of the ACM, 19(4):660–674, October 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Savage:1984:STT

[Sav84] John E. Savage. Space-time trade-offs for banded matrix problems. Journal of the ACM, 31(2):422–437, April 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Slagle:1968:EMT

[SB68] James R. Slagle and Philip Bursky. Experiments with a multipurpose, theorem-proving heuristic program. Journal of the ACM, 15(1):85–99, January 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sloate:1973:SCM

[SB73] Harry M. Sloate and Theodore A. Bickart. A-stable composite multistep methods. Journal of the ACM, 20(1):7–26, January 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Samelson:1982:NPF

[SB82] Christopher L. Samelson and William G. Bulgren. A note on product-form solution for queuing networks with Poisson arrivals and general service-time distributions with finite means. Journal of the ACM, 29(3):830–840, July 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sahni:1980:SIT

[SC80] Sartaj Sahni and Yookun Cho. Scheduling independent tasks with due times on a uniform processor system. Journal of the ACM, 27(3):550–563, July 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sistla:1985:CPL

[SC85] A. P. Sistla and E. M. Clarke. The complexity of propositional linear temporal logics. Journal of the ACM, 32(3):733–749, July 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/3837.html.

Sen:2002:TTC

[SCD02] Sandeep Sen, Siddhartha Chatterjee, and Neeraj Dumir. Towards a theory of cache-efficient algorithms. Journal of the ACM, 49(6):828–858, November 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Schwartz:1961:ASP

[Sch61] Eugene S. Schwartz. An automatic sequencing procedure with application to parallel programming. Journal of the ACM, 8(4): 513–537, October 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Schwartz:1963:DMR

[Sch63] Eugene S. Schwartz. A dictionary for minimum redundancy encoding. Journal of the ACM, 10(4):413–439, October 1963. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Schroeder:1964:IDS

[Sch64] R. Schroeder. Input data source limitations for real-time operation of digital computers. Journal of the ACM, 11(2):152–158, April 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Schwartz:1966:LPC

[Sch66] J. Schwartz. Large parallel computers. Journal of the ACM, 13 (1):25–32, January 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Schutzenberger:1968:RAS

[Sch68] Marcel Paul Schützenberger. A remark on acceptable sets of numbers. Journal of the ACM, 15(2):300–303, April 1968. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Schultz:1972:SMM

[Sch72] Gary D. Schultz. A stochastic model for message assembly buffering with a comparison of block assignment strategies. Journal of the ACM, 19(3):483–495, July 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Schubert:1974:ILR

[Sch74] L. K. Schubert. Iterated limiting recursion and the program minimization problem. Journal of the ACM, 21(3):436–445, July 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Schnorr:1978:SQC

[Sch78] C. P. Schnorr. Satisfiability is quasilinear complete in NQL. Journal of the ACM, 25(1):136–145, January 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Schwartz:1980:FPA

[Sch80] J. T. Schwartz. Fast probabilistic algorithms for verification of polynomial identities. Journal of the ACM, 27(4):701–717, October 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Schonhage:1988:NLB

[Sch88] Arnold Schönhage. A nonlinear lower bound for random-access machines under logarithmic cost. Journal of the ACM, 35(3): 748–754, July 1988. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/44492.html.

Sciore:1982:CAF

[Sci82] Edward Sciore. A complete axiomatization of full join dependencies. Journal of the ACM, 29(2):373–393, April 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Scott:1958:APF

[Sco58] A. E. Scott. Automatic preparation of flow chart listings. Journal of the ACM, 5(1):57–66, January 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Smith:1962:MAN

[SD62] D. R. Smith and C. H. Davidson. Maintained activity in neural nets. Journal of the ACM, 9(2):268–279, April 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Slagle:1969:ESP

[SD69] James H. Slagle and John K. Dixon. Experiments with some programs that search game trees. Journal of the ACM, 16(2):189– 207, April 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Schmidt:1976:FBA

[SD76] Douglas C. Schmidt and Larry E. Druffel. A fast backtracking algorithm to test directed graphs for isomorphism using distance matrices. Journal of the ACM, 23(3):433–445, July 1976. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Schassberger:1983:TRT

[SD83] R. Schassberger and H. Daduna. The time for a round trip in a cycle of exponential queues. Journal of the ACM, 30(1):146– 150, January 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sagiv:1981:EBR

[SDPF81] Yehoshua Sagiv, Claude Delobel, D. Stott Parker, Jr., and Ronald Fagin. An equivalence between relational database dependencies and a fragment of propositional logic. Journal of the ACM, 28(3):435–453, July 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [SDPF87].

Sagiv:1987:CEB

[SDPF87] Y. Sagiv, C. Delobel, D. S. Parker, Jr., and Ronald Fagin. Correction to “an equivalence between relational database dependencies and a fragment of propositional logic”. Journal of the ACM, 34(4):1016–1018, October 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/31853.html. See

[SDPF81].

Stefanov:2018:POE

[SDS^{+}18] Emil Stefanov, Marten Van Dijk, Elaine Shi, T.-H. Hubert Chan, Christopher Fletcher, Ling Ren, Xiangyao Yu, and Srinivas Devadas. Path ORAM: an extremely simple oblivious RAM protocol. Journal of the ACM, 65(4):18:1–18:??, August 2018. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Seiden:2002:OPP

[Sei02] Steven S. Seiden. On the online bin packing problem. Journal of the ACM, 49(5):640–671, September 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Selkow:1972:OPC

[Sel72] Stanley M. Selkow. One-pass complexity of digital picture properties. Journal of the ACM, 19(2):283–295, April 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sethi:1974:TCR

[Set74] Ravi Sethi. Testing for the Church–Rosser property. Journal of the ACM, 21(4):671–679, October 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [Set75].

Sethi:1975:ETC

[Set75] Ravi Sethi. Errata: “Testing for the Church–Rosser property”. Journal of the ACM, 22(3):424, July 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Set74].

Sethi:1978:CEE

[Set78] Ravi Sethi. Conditional expressions with equality tests. Journal of the ACM, 25(4):667–674, October 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sethi:1982:UAM

[Set82] Ravi Sethi. Useless actions make a difference: Strict serializability of database updates. Journal of the ACM, 29(2):394–403, April 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sevcik:1974:SMT

[Sev74] Kenneth C. Sevcik. Scheduling for minimum total loss using service time distributions. Journal of the ACM, 21(1):66–75, January 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stewart:2015:UBN

[SEY15] Alistair Stewart, Kousha Etessami, and Mihalis Yannakakis. Upper bounds for Newton’s method on monotone polynomial systems, and P-time model checking of probabilistic one-counter automata. Journal of the ACM, 62(4):30:1–30:??, August 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Solomon:1980:NEB

[SF80] Marvin Solomon and Raphael A. Finkel. A note on enumerating binary trees. Journal of the ACM, 27(1):3–5, January 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sozeau:2025:CCT

[SFLB^{+}25] Matthieu Sozeau, Yannick Forster, Meven Lennon-Bertrand, Jakob Nielsen, Nicolas Tabareau, and Théo Winterhalter. Correct and complete type checking and certified erasure for Coq, in Coq. Journal of the ACM, 72(1):8:1–8:??, February 2025. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3706056.

Seiferas:1978:SNT

[SFM78] Joel I. Seiferas, Michael J. Fischer, and Albert R. Meyer. Separating nondeterministic time complexity classes. Journal of the ACM, 25(1):146–167, January 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sklansky:1968:FPT

[SFR68] J. Sklansky, M. Finkelstein, and E. C. Russell. A formalism for program translation. Journal of the ACM, 15(2):165–175, April 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sahni:1976:CAP

[SG76] Sartaj Sahni and Teofilo Gonzalez. P -complete approximation problems. Journal of the ACM, 23(3):555–565, July 1976. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sarma:2011:EPG

[SGP11] Atish Das Sarma, Sreenivas Gollapudi, and Rina Panigrahy. Estimating PageRank on graph streams. Journal of the ACM, 58 (3):13:1–13:19, May 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Schank:1975:IPC

[SGRR75] Roger C. Schank, Neil M. Goldman, Charles J. Rieger III, and Christopher K. Riesbeck. Inference and paraphrase by computer. Journal of the ACM, 22(3):309–328, July 1975. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sasaki:1988:TCM

[SH88] Galen H. Sasaki and Bruce Hajek. The time complexity of maximum matching by simulated annealing. Journal of the ACM, 35 (2):387–403, April 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/46160.html.

Sterling:2021:LRT

[SH21] Jonathan Sterling and Robert Harper. Logical relations as types: Proof-relevant parametricity for program modules. Journal of the ACM, 68(6):41:1–41:47, December 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3474834.

Soloviev:2022:IAU

[SH22] Matvey Soloviev and Joseph Y. Halpern. Information acquisition under resource limitations in a noisy environment. Journal of the ACM, 69(3):20:1–20:37, June 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/doi/10.1145/3510024.

Shaw:1966:DAC

[Sha66] H. Shaw. Discrete analogs for continuous filters. Journal of the ACM, 13(4):600–604, October 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shapiro:1967:ARP

[Sha67a] Marvin B. Shapiro. An algorithm for reconstructing protein and RNA sequences. Journal of the ACM, 14(4):720–731, October 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shaw:1967:MMM

[Sha67b] Brian Shaw. Modified multistep methods based on a nonpolynomial interpolant. Journal of the ACM, 14(1):143–154, January 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shampine:1970:EPN

[Sha70a] L. F. Shampine. Efficiency of a procedure for near-minimax approximation. Journal of the ACM, 17(4):655–660, October 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shaw:1970:PGR

[Sha70b] Alan C. Shaw. Parsing of graph-representable pictures. Journal of the ACM, 17(3):453–481, July 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shankar:1988:MPC

[Sha88] N. Shankar. A mechanical proof of the Church–Rosser theorem. Journal of the ACM, 35(3):475–522, July 1988. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
44484.html.

Shamir:1992:IP

[Sha92] Adi Shamir. IP = PSPACE. Journal of the ACM, 39(4):869– 877, October 1992. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/146609.html. This paper shows that

the set of problems for which interactive protocols exist is precisely the set of problems which are solvable within polynomial space on a Turing machine.

Shenitzer:1957:CAC

[She57a] A. Shenitzer. Chebyshev approximation of a continuous function by a class of functions. Journal of the ACM, 4(1):30–35, January 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sherman:1957:DTP

[She57b] Bernard Sherman. Determination of three percentiles of the ω_{n} distribution function. Journal of the ACM, 4(4):472–476, October 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sheldon:1959:SNS

[She59a] J. W. Sheldon. On the spectral norms of several iterative processes. Journal of the ACM, 6(4):494–505, October 1959. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shell:1959:SSC

[She59b] Donald L. Shell. The share 709 system: a cooperative effort. Journal of the ACM, 6(2):123–127, April 1959. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sheng:1965:CTL

[She65a] C. L. Sheng. Correction: “Threshold logic elements used as a probability transformer”. Journal of the ACM, 12(3):435, July 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [She65b].

Sheng:1965:TLE

[She65b] C. L. Sheng. Threshold logic elements used as a probability transformer. Journal of the ACM, 12(2):262–276, April 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [She65a].

Shemer:1967:SMC

[She67] Jack E. Shemer. Some mathematical considerations of timesharing scheduling algorithms. Journal of the ACM, 14(2):262– 272, April 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shen:1992:IPS

[She92] A. Shen. IP = PSPACE: Simplified proof. Journal of the ACM, 39(4):878–880, October 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/146613.html.

Sherstov:2014:CLB

[She14] Alexander A. Sherstov. Communication lower bounds using directional derivatives. Journal of the ACM, 61(6):34:1–34:??, November 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shimizu:1969:SAM

[Shi69] Tamio Shimizu. A stochastic approximation method for optimization problems. Journal of the ACM, 16(3):511–516, July 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shiloach:1980:PSU

[Shi80] Yossi Shiloach. A polynomial solution in the undirected two paths problem. Journal of the ACM, 27(3):445–456, July 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shibuya:2010:GST

[Shi10] Tetsuo Shibuya. Geometric suffix tree: Indexing protein 3-D structures. Journal of the ACM, 57(3):15:1–15:17, March 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shostak:1977:SIM

[Sho77] Robert E. Shostak. On the SUP-INF method for proving Presburger formulas. Journal of the ACM, 24(4):529–543, October 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shostak:1979:PDP

[Sho79] Robert E. Shostak. A practical decision procedure for arithmetic with function symbols. Journal of the ACM, 26(2):351–360, April 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shostak:1981:DLI

[Sho81] Robert Shostak. Deciding linear inequalities by computing loop residues. Journal of the ACM, 28(4):769–779, October 1981. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shostak:1984:DCT

[Sho84] Robert E. Shostak. Deciding combinations of theories. Journal of the ACM, 31(1):1–12, January 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shor:2003:WHM

[Sho03] Peter W. Shor. Why haven’t more quantum algorithms been found? Journal of the ACM, 50(1):87–90, January 2003. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shrager:1970:NRL

[Shr70] Richard I. Shrager. Nonlinear regression with linear constraints: An extension of the magnified diagonal method. Journal of the ACM, 17(3):446–452, July 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sigillito:1967:CMA

[Sig67] V. G. Sigillito. On a continuous method of approximating solutions of the heat equation. Journal of the ACM, 14(4):732–741, October 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Simon:1963:EHC

[Sim63] Herbert A. Simon. Experiments with a heuristic compiler. Journal of the ACM, 10:493–506, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

SimoesPereira:1965:BME

[Sim65] J. M. S. Simões Pereira. On the Boolean matrix equation M′ = ∨_{i=1}M^{i}. Journal of the ACM, 12(3):376–382, July 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [Sim67].

SimoesPereira:1967:CBM

[Sim67] J. M. S. Simões Pereira. Corrigendum: “On the Boolean Matrix Equation M′ = ∨_{i=1}M^{i}”. Journal of the ACM, 14(2):419–420, April 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Sim65].

Simon:1984:PQF

[Sim84] B. Simon. Priority queues with feedback. Journal of the ACM, 31 (1):134–149, January 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Singh:1986:IMS

[Sin86] Samar Singh. Improved methods for storing and updating information in the out-of-kilter algorithm. Journal of the ACM, 33(3): 551–567, July 1986. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/5934.html.

Suzuki:1980:VDP

[SJ80] Norihisa Suzuki and David Jefferson. Verification decidability of Pressburger array programs. Journal of the ACM, 27(1):191– 205, January 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Salkin:1973:SCA

[SK73] Harvey M. Salkin and Ronald D. Koncal. Set covering by an all integer algorithm: Computational experience. Journal of the ACM, 20(2):189–193, April 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sahasrabudhe:1977:SFI

[SK77] S. E. Sahasrabudhe and A. D. Kulkarni. On solving Fredholm integral equations of the first kind. Journal of the ACM, 24(4): 624–629, October 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sameh:1978:SPL

[SK78] A. H. Sameh and D. J. Kuck. On stable parallel linear system solvers. Journal of the ACM, 25(1):81–91, January 1978. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Silberschatz:1980:CHD

[SK80] Abraham Silberschatz and Zvi Kedem. Consistency in hierarchical database systems. Journal of the ACM, 27(1):72–80, January 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Selman:1996:KCT

[SK96] Bart Selman and Henry Kautz. Knowledge compilation and theory approximation. Journal of the ACM, 43(2):193–224, March 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
jacm/226644.html.

Skeel:1979:SNS

[Ske79] Robert D. Skeel. Scaling for numerical stability in Gaussian elimination. Journal of the ACM, 26(3):494–526, July 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sklansky:1970:TCO

[Skl70] J. Sklansky. Thresholded convolution operations. Journal of the ACM, 17(1):161–165, January 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Schreiber:2018:OMW

[SKM18] Ethan L. Schreiber, Richard E. Korf, and Michael D. Moffitt. Optimal multi-way number partitioning. Journal of the ACM, 65(4):24:1–24:??, August 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Skutella:2001:CQS

[Sku01] Martin Skutella. Convex quadratic and semidefinite programming relaxations in scheduling. Journal of the ACM, 48(2): 206–242, March 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/citations/journals/jacm/2001-48-2/p206-skutella/

.

Salton:1968:CEI

[SL68] G. Salton and M. E. Lesk. Computer evaluation of indexing and text processing. Journal of the ACM, 15(1):8–36, January 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Simon:1971:OSS

[SL71] Richard Simon and Richard C. T. Lee. On the optimal solutions to and/or series-parallel graphs. Journal of the ACM, 18(3): 354–372, July 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Singh:2015:AMB

[SL15] Mohit Singh and Lap Chi Lau. Approximating minimum bounded degree spanning trees to within one of optimal. Journal of the ACM, 62(1):1:1–1:??, February 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Slagle:1963:HPS

[Sla63] James R. Slagle. A heuristic program that solves symbolic integration problems in freshman calculus. Journal of the ACM, 10:507–520, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Slagle:1964:EAF

[Sla64] James R. Slagle. An efficient algorithm for finding certain minimum-cost procedures for making binary decisions. Journal of the ACM, 11(3):253–264, July 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Slagle:1967:ATP

[Sla67] James R. Slagle. Automatic theorem proving with renamable and semantic resolution. Journal of the ACM, 14(4):687–697, October 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Slagle:1970:ITR

[Sla70] James R. Slagle. Interpolation theorems for resolution in lower predicate calculus. Journal of the ACM, 17(3):535–542, July 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Slagle:1972:AFL

[Sla72a] James R. Slagle. An approach for finding C-linear complete inference systems. Journal of the ACM, 19(3):496–516, July 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Slagle:1972:ATP

[Sla72b] James R. Slagle. Automatic theorem proving with built-in theories including equality, partial ordering, and sets. Journal of the ACM, 19(1):120–135, January 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Slagle:1974:ATP

[Sla74] James R. Slagle. Automated theorem-proving for theories with simplifiers, commutativity, and associativity. Journal of the ACM, 21(4):622–642, October 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Slater:1976:DG

[Sla76] Peter J. Slater. R-domination in graphs. Journal of the ACM, 23(3):446–450, July 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

S:2019:PCA

[SLM19] Karthik C. S., Bundit Laekhanukit, and Pasin Manurangsi. On the parameterized complexity of approximating dominating set. Journal of the ACM, 66(5):33:1–33:??, September 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/ft_gateway.cfm?id=3325116.

Sackman:1964:ICO

[SM64] Harold Sackman and J. B. Munson. Investigation of computer operating time and system capacity for man-machine digital systems. Journal of the ACM, 11(4):450–464, October 1964. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sevcik:1981:DQN

[SM81] K. C. Sevcik and I. Mitrani. The distribution of queuing network states at input and output instants. Journal of the ACM, 28 (2):358–371, April 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shahrokhi:1990:MCF

[SM90] Farhad Shahrokhi and D. W. Matula. The maximum concurrent flow problem. Journal of the ACM, 37(2):318–334, April 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/77620.html.

Stockmeyer:2002:CLB

[SM02] Larry Stockmeyer and Albert R. Meyer. Cosmological lower bound on the circuit complexity of a small problem in logic. Journal of the ACM, 49(6):753–784, November 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Strong:1979:SWP

[SMC79] H. R. Strong, G. Markowsky, and A. K. Chandra. Search within a page. Journal of the ACM, 26(3):457–482, July 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Smith:1970:EBZ

[Smi70] Brian T. Smith. Error bounds for zeros of a polynomial based upon Gerschgorin’s theorems. Journal of the ACM, 17(4):661– 674, October 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Smith:1971:MPR

[Smi71a] C. S. Smith. Multiplicative pseudo-random number generators with prime modulus. Journal of the ACM, 18(4):586–593, October 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Smith:1971:SCU

[Smi71b] Alvy Ray Smith III. Simple computation-universal cellular spaces. Journal of the ACM, 18(3):339–353, July 1971. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Proof that 1D cellular automata are capable of supporting universal computation.

Smith:1979:CSP

[Smi79] Alan Jay Smith. Characterizing the storage process and its effect on the update of main memory by write through. Journal of the ACM, 26(1):6–27, January 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Smith:1982:PPA

[Smi82] Carl H. Smith. The power of pluralism for automatic program synthesis. Journal of the ACM, 29(4):1144–1165, October 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Smith:1984:RTA

[Smi84] Douglas R. Smith. Random trees and the analysis of branch and bound procedures. Journal of the ACM, 31(1):163–188, January 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Seidl:2018:EDT

[SMK18] Helmut Seidl, Sebastian Maneth, and Gregor Kemper. Equivalence of deterministic top-down tree-to-string transducers is decidable. Journal of the ACM, 65(4):21:1–21:??, August 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Staples:1985:FSN

[SN85] John Staples and V. L. Nguyen. A fixpoint semantics for nondeterministic data flow. Journal of the ACM, 32(2):411–444, April 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/3155.html.

Steinke:2004:UTS

[SN04] Robert C. Steinke and Gary J. Nutt. A unified theory of shared memory consistency. Journal of the ACM, 51(5):800–849, September 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sarma:2013:DRW

[SNPT13] Atish Das Sarma, Danupon Nanongkai, Gopal Pandurangan, and Prasad Tetali. Distributed random walks. Journal of the ACM, 60(1):2:1–2:31, February 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sobel:1962:OSN

[Sob62] Sheldon Sobel. Oscillating sort — a new sort merging technique. Journal of the ACM, 9(3):372–374, July 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sumii:2007:BTA

[SP07] Eijiro Sumii and Benjamin C. Pierce. A bisimulation for type abstraction and recursion. Journal of the ACM, 54(5):26:1–26:43, October 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Spencer:1997:TWT

[Spe97] Thomas H. Spencer. Time-work tradeoffs for parallel algorithms. Journal of the ACM, 44(5):742–778, September 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org:80/pubs/citations/
journals/jacm/1997-44-5/p742-spencer/.

Spielberg:1961:RPS

[Spi61] Kurt Spielberg. Representation of power series in terms of polynomials, rational approximations and continued fractions. Journal of the ACM, 8(4):613–627, October 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Spira:1969:TRG

[Spi69] Philip M. Spira. The time required for group multiplication. Journal of the ACM, 16(2):235–243, April 1969. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Seshadhri:2013:DAS

[SPK13] C. Seshadhri, Ali Pinar, and Tamara G. Kolda. An in-depth analysis of stochastic Kronecker graphs. Journal of the ACM, 60 (2):13:1–13:32, April 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sprugnoli:1992:GBT

[Spr92] Renzo Sprugnoli. The generation of binary trees as a numerical problem. Journal of the ACM, 39(2):317–327, April 1992. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
128753.html.

Stein:1960:CAD

[SR60] Marvin L. Stein and Jack Rose. Changing from analog to digital programming by digital techniques. Journal of the ACM, 7 (1):10–23, January 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Steffanelli:1971:SPT

[SR71] R. Steffanelli and A. Rosenfeld. Some parallel thinning algorithms for digital pictures. Journal of the ACM, 18(2):255–264, April 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Storer:1994:SPP

[SR94] James A. Storer and John H. Reif. Shortest paths in the plane with polygonal obstacles. Journal of the ACM, 41(5):982–1012, September 1994. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/185795.html.

Srinivasan:1965:ISC

[Sri65] A. V. Srinivasan. An investigation of some computational aspects of integer programming. Journal of the ACM, 12(4):525–535, October 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Srinivasan:1976:LPC

[Sri76] V. Srinivasan. Linear programming computational procedures for ordinal regression. Journal of the ACM, 23(3):475–487, July 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sekar:1997:PLS

[SRM97] R. Sekar, I. V. Ramakrishnan, and P. Mishra. On the power and limitations of strictness analysis. Journal of the ACM, 44 (3):505–525, May 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/citations/journals/jacm/1997-44-3/p505-sekar/.

Spitzbart:1958:CFC

[SS58] A. Spitzbart and D. L. Shell. A Chebycheff fitting criterion. Journal of the ACM, 5(1):22–31, January 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shepherdson:1963:CRF

[SS63] J. C. Shepherdson and H. E. Sturgis. Computability of recursive functions. Journal of the ACM, 10(2):217–255, April 1963. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Savitch:1979:TBR

[SS79] Walter J. Savitch and Michael J. Stimson. Time bounded random access machines with parallel processing. Journal of the ACM, 26(1):103–118, January 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Storer:1982:DCT

[SS82] James A. Storer and Thomas G. Szymanski. Data compression via textual substitution. Journal of the ACM, 29(4):928–951, October 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shalev:2006:SOL

[SS06a] Ori Shalev and Nir Shavit. Split-ordered lists: Lock-free extensible hash tables. Journal of the ACM, 53(3):379–405, May 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shmoys:2006:ASS

[SS06b] David B. Shmoys and Chaitanya Swamy. An approximation scheme for stochastic linear programming and its application to stochastic integer programs. Journal of the ACM, 53(6):978– 1012, November 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sitters:2006:GTS

[SS06c] René A. Sitters and Leen Stougie. The generalized two-server problem. Journal of the ACM, 53(3):437–458, May 2006. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Saxena:2013:SGC

[SS13] Nitin Saxena and C. Seshadhri. From Sylvester–Gallai configurations to rank bounds: Improved blackbox identity test for depth-3 circuits. Journal of the ACM, 60(5):33:1–33:33, October 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Schulman:2017:ACM

[SS17] Leonard J. Schulman and Alistair Sinclair. Analysis of a classical matrix preconditioning algorithm. Journal of the ACM, 64(2): 9:1–9:23, June 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stein:2024:PPE

[SS24] Dario Stein and Sam Staton. Probabilistic programming with exact conditions. Journal of the ACM, 71(1):2:1–2:??, February 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/doi/10.1145/
3632170.

Sippu:1996:AMS

[SSS96] Seppo Sippu and Eljas Soisalon-Soininen. An analysis of magic sets and related optimization strategies for logic queries. Journal of the ACM, 43(6):1046–1088, November 1996. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
jacm/235814.html.

Scharbrodt:2006:NAC

[SSS06] Mark Scharbrodt, Thomas Schickinger, and Angelika Steger. A new average case analysis for completion time scheduling. Journal of the ACM, 53(1):121–146, January 2006. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sippu:1983:CLK

[SSSU83] Seppo Sippu, Eljas Soisalon-Soininen, and Esko Ukkonen. The complexity of LALR(k) testing. Journal of the ACM, 30(2):259– 270, April 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Schweikardt:2022:EFQ

[SSV22] Nicole Schweikardt, Luc Segoufin, and Alexandre Vigny. Enumeration for FO queries over nowhere dense graphs. Journal of the ACM, 69(3):22:1–22:37, June 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3517035.

Saks:1998:EDP

[SSZ98] Michael Saks, Aravind Srinivasan, and Shiyu Zhou. Explicit ORdispersers with polylogarithmic degree. Journal of the ACM, 45 (1):123–154, January 1998. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org:80/

pubs/citations/journals/jacm/1998-45-1/p123-saks/.

Srinivasan:1972:AAL

[ST72] V. Srinivasan and G. L. Thompson. Accelerated algorithms for labeling and relabeling of trees, with applications to distribution problems. Journal of the ACM, 19(4):712–726, October 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Srinivasan:1973:BCA

[ST73] V. Srinivasan and G. L. Thompson. Benefit-cost analysis of coding techniques for the primal transportation algorithm. Journal of the ACM, 20(2):194–213, April 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shaw:1974:NME

[ST74] Mary Shaw and J. F. Traub. On the number of multiplications for the evaluation of a polynomial and some of its derivatives. Journal of the ACM, 21(1):161–167, January 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sethi:1980:CCV

[ST80] Ravi Sethi and Adrian Tang. Constructing call-by-value continuation semantics. Journal of the ACM, 27(3):580–597, July 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sleator:1985:SAB

[ST85] Daniel Dominic Sleator and Robert Endre Tarjan. Self-adjusting binary search trees. Journal of the ACM, 32(3):652–686, July 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/3835.html.

Srikanth:1987:OCS

[ST87] T. K. Srikanth and Sam Toueg. Optimal clock synchronization. Journal of the ACM, 34(3):626–645, July 1987. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
28876.html.

Spielman:2004:SAA

[ST04] Daniel A. Spielman and Shang-Hua Teng. Smoothed analysis of algorithms: Why the simplex algorithm usually takes polynomial time. Journal of the ACM, 51(3):385–463, May 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stark:1956:RCN

[Sta56] R. H. Stark. Rates of convergence in numerical solution of the diffusion equation. Journal of the ACM, 3(1):29–40, January 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stanfel:1970:TSO

[Sta70] Larry E. Stanfel. Tree structures for optimal searching. Journal of the ACM, 17(3):508–517, July 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stanfel:1972:PAD

[Sta72] L. E. Stanfel. Practical aspects of doubly chained trees for retrieval. Journal of the ACM, 19(3):425–436, July 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stark:1982:SPS

[Sta82] Eugene W. Stark. Semaphore primitives and starvation-free mutual exclusion. Journal of the ACM, 29(4):1049–1072, October 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Steel:1957:PI

[Ste57] T. B. Steel, Jr. Pact IA. Journal of the ACM, 4(1):8–11, January 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stewart:1967:MDM

[Ste67] G. W. Stewart III. A modification of Davidon’s minimization method to accept difference approximations to derivatives. Journal of the ACM, 14(1):72–83, January 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stewart:1983:CEB

[Ste83] G. W. Stewart. Computable error bounds for aggregated Markov chains. Journal of the ACM, 30(2):271–285, April 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stearns:2003:DVN

[Ste03] Richard E. Stearns. Deterministic versus nondeterministic time and lower bound problems. Journal of the ACM, 50(1):91–95, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stiles:1961:AFI

[Sti61] H. Edmund Stiles. The association factor in information retrieval. Journal of the ACM, 8(2):271–279, April 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stineman:1965:DTD

[Sti65] R. W. Stineman. Digital time-domain analysis of systems with widely separated poles. Journal of the ACM, 12(2):286–294, April 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stillman:1973:CWS

[Sti73] Rona B. Stillman. The concept of weak substitution in theoremproving. Journal of the ACM, 20(4):648–667, October 1973. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stickel:1981:UAA

[Sti81] Mark E. Stickel. A unification algorithm for associativecommutative functions. Journal of the ACM, 28(2):423–434, April 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stockmal:1964:CPR

[Sto64a] Frank Stockmal. Calculations with pseudo-random numbers. Journal of the ACM, 11(1):41–52, January 1964. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stoer:1964:DMC

[Sto64b] Josef Stoer. A direct method for Chebyshev approximation by rational functions. Journal of the ACM, 11(1):59–69, January 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stone:1970:AMP

[Sto70] Harold S. Stone. An algorithm for modular partitioning. Journal of the ACM, 17(1):182–195, January 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stone:1973:EPA

[Sto73] Harold S. Stone. An efficient parallel algorithm for the solution of a tridiagonal linear system of equations. Journal of the ACM, 20(1):27–38, January 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Strong:1983:VEF

[Str83] H. R. Strong. Vector execution of flow graphs. Journal of the ACM, 30(1):186–196, January 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Svensson:2020:CFA

[STV20] Ola Svensson, Jakub Tarnawski, and László A. Végh. A constant-factor approximation algorithm for the asymmetric traveling salesman problem. Journal of the ACM, 67(6):37:1– 37:53, November 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/

doi/10.1145/3424306.

Sethi:1970:GOC

[SU70] Ravi Sethi and J. D. Ullman. The generation of optimal code for arithmetic expressions. Journal of the ACM, 17(4):715–728, October 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Reprinted as pp. 229–247 in Compiler Techniques, B. W. Pollack, ed., Auerbach, Princeton NJ (1972).

Sadri:1982:TDL

[SU82] Fereidoon Sadri and Jeffrey D. Ullman. Template dependencies: a large class of dependencies in relational databases and its complete axiomatization. Journal of the ACM, 29(2):363–372, April 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Shaltiel:2005:SEA

[SU05] Ronen Shaltiel and Christopher Umans. Simple extractors for all min-entropies and a new pseudorandom generator. Journal of the ACM, 52(2):172–216, March 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sudborough:1975:NTB

[Sud75] I. H. Sudborough. A note on tape-bounded complexity classes and linear context-free languages. Journal of the ACM, 22(4): 499–500, October 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sudborough:1978:TCD

[Sud78] I. H. Sudborough. On the tape complexity of deterministic context-free languages. Journal of the ACM, 25(3):405–414, July 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Summers:1977:MLP

[Sum77] Phillip D. Summers. A methodology for LISP program construction from examples. Journal of the ACM, 24(1):161–175, January 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Supowit:1983:RNG

[Sup83] Kenneth J. Supowit. The relative neighborhood graph, with an application to minimum spanning trees. Journal of the ACM, 30(3):428–448, July 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Suri:1983:RQN

[Sur83] Rajan Suri. Robustness of queuing network formulas. Journal of the ACM, 30(3):564–594, July 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Suri:1987:IPA

[Sur87] Rajan Suri. Infinitesimal perturbation analysis for general discrete event systems. Journal of the ACM, 34(3):686–717, July 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/28879.html.

Skyum:1985:CTB

[SV85] S. Skyum and L. G. Valiant. A complexity theory based on Boolean algebra. Journal of the ACM, 32(2):484–502, April 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/3158.html.

Seiferas:1988:CE

[SV88] Joel I. Seiferas and Paul M. B. Vitányi. Counting is easy. Journal of the ACM, 35(4):985–1000, October 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/63141.html.

Sahai:2003:CPS

[SV03] Amit Sahai and Salil Vadhan. A complete problem for statistical zero knowledge. Journal of the ACM, 50(2):196–249, March 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Suciu:2006:I

[SV06] Dan Suciu and Victor Vianu. Introduction. Journal of the ACM, 53(2):207, March 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sebo:2019:SIP

[SV19] András Sebő and Anke Van Zuylen. The salesman’s improved paths through forests. Journal of the ACM, 66(4):28:1–28:??, August 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-

735X (electronic). URL https://dl.acm.org/ft_gateway.
cfm?id=3326123.

Sevcik:2013:CVC

[SVN^{+}13] Jaroslav Sevcík, Viktor Vafeiadis, Francesco Zappa Nardelli, Suresh Jagannathan, and Peter Sewell. CompCertTSO: a verified compiler for relaxed-memory concurrency. Journal of the ACM, 60(3):22:1–22:50, June 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Svoboda:1975:CTE

[Svo75] Antonin Svoboda. The concept of term exclusiveness and its effect on the theory of Boolean functions. Journal of the ACM, 22(3):425–440, July 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stefankovic:2009:ASA

[ŠVV09] Daniel Štefankovič, Santosh Vempala, and Eric Vigoda. Adaptive simulated annealing: a near-optimal connection between sampling and counting. Journal of the ACM, 56(3):18:1–18:36, May 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Strand:1968:SEN

[SW68] Otto Neall Strand and Ed R. Westwater. Statistical estimation of the numerical solution of a Fredholm integral equation of the first kind. Journal of the ACM, 15(1):100–114, January 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sagiv:1982:SDC

[SW82] Yehoshua Sagiv and Scott F. Walecka. Subset dependencies and a completeness result for a subclass of embedded multivalued dependencies. Journal of the ACM, 29(1):103–117, January 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Stewart:1991:M

[SW91] Bradley S. Stewart and Chelsea C. White, III. Multiobjective A*. Journal of the ACM, 38(4):775–814, October 1991. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
115368.html.

Stoer:1997:SMC

[SW97] Mechthild Stoer and Frank Wagner. A simple min-cut algorithm. Journal of the ACM, 44(4):585–591, July 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org:80/pubs/citations/
journals/jacm/1997-44-4/p585-stoer/.

Sima:1998:TN

[ŠW98] Jiří Šíma and Jiří Wiedermann. Theory of neuromata. Journal of the ACM, 45(1):155–178, January 1998. CODEN JACOAH.

ISSN 0004-5411 (print), 1557-735X (electronic). URL http:/

/www.acm.org:80/pubs/citations/journals/jacm/1998-45-
1/p155-scaronima/.

Siekmann:2001:ECW

[SW01] Jörg H. Siekmann and Graham Wrightson. Erratum: a counterexample to W. Bibel’s and E. Eder’s strong completeness result for connection graph resolution. Journal of the ACM, 48(1):145–147, January 2001. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/articles/journals/jacm/
2001-48-1/p145-siekmann/p145-siekmann.pdf;   http://
www.acm.org/pubs/citations/journals/jacm/2001-48-1/p145-

siekmann/. See [BE97].

Swartz:1980:PLS

[Swa80] G. Boyd Swartz. Polling in a loop system. Journal of the ACM, 27(1):42–59, January 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Swift:1957:MFM

[Swi57] Charles J. Swift. Machine features for a more automatic monitoring system on digital computers. Journal of the ACM, 4 (2):172–173, April 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sagiv:1980:EAR

[SY80] Yehoshua Sagiv and Mihalis Yannakakis. Equivalences among relational expressions with the union and difference operators. Journal of the ACM, 27(4):633–655, October 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Szpankowski:1990:PTA

[Szp90] Wojciech Szpankowski. Patricia tries again revisited. Journal of the ACM, 37(4):691–711, October 1990. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/214080.html.

Tai:1979:TTC

[Tai79] Kuo-Chung Tai. The tree-to-tree correction problem. Journal of the ACM, 26(3):422–433, July 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tal:1969:MDO

[Tal69] A. Tal. On monotone decomposable operators. Journal of the ACM, 16(3):507–510, July 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tarjan:1972:SUN

[Tar72] Robert Tarjan. Sorting using networks of queues and stacks. Journal of the ACM, 19(2):341–346, April 1972. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tarjan:1975:EGL

[Tar75] Robert Endre Tarjan. Efficiency of a good but not linear set union algorithm. Journal of the ACM, 22(2):215–225, April 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tarjan:1979:APC

[Tar79] Robert Endre Tarjan. Applications of path compression on balanced trees. Journal of the ACM, 26(4):690–715, October 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tarjan:1981:FAS

[Tar81a] Robert Endre Tarjan. Fast algorithms for solving path problems. Journal of the ACM, 28(3):594–614, July 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tarjan:1981:UAP

[Tar81b] Robert Endre Tarjan. A unified approach to path problems. Journal of the ACM, 28(3):577–593, July 1981. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tarsi:1983:OSS

[Tar83] Michael Tarsi. Optimal search on some game trees. Journal of the ACM, 30(3):389–396, July 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Also as Tech. Report UCLA-ENG-CSL-8108.

Tardos:2008:OPF

[Tar08] Gábor Tardos. Optimal probabilistic fingerprint codes. Journal of the ACM, 55(2):10:1–10:24, May 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tardos:2016:IAFa

[Tar16a] Eva Tardos. Invited articles foreword. Journal of the ACM, 63 (1):8:1–8:??, March 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tardos:2016:IAFb

[Tar16b] Eva Tardos. Invited articles foreword. Journal of the ACM, 63 (2):18:1–18:??, May 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tardos:2016:IAFc

[Tar16c] Eva Tardos. Invited articles foreword. Journal of the ACM, 63 (3):27:1–27:??, September 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tardos:2016:IAFd

[Tar16d] Eva Tardos. Invited articles foreword. Journal of the ACM, 63 (5):45:1–45:??, December 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tardos:2017:IAFa

[Tar17a] Eva Tardos. Invited article foreword. Journal of the ACM, 63(6): 53:1–53:??, February 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tardos:2017:IAFb

[Tar17b] Eva Tardos. Invited article foreword. Journal of the ACM, 64 (1):6:1–6:??, March 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tardos:2017:IAFd

[Tar17c] Eva Tardos. Invited article foreword. Journal of the ACM, 64 (3):22:1–22:??, June 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tardos:2017:IAFe

[Tar17d] Eva Tardos. Invited article foreword. Journal of the ACM, 64 (4):29:1–29:??, September 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tardos:2017:IAFh

[Tar17e] Eva Tardos. Invited article foreword. Journal of the ACM, 65(1): 5:1–5:??, December 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tardos:2017:IAFc

[Tar17f] Eva Tardos. Invited articles foreword. Journal of the ACM, 64 (2):13:1–13:??, June 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tardos:2017:IAFf

[Tar17g] Eva Tardos. Invited articles foreword. Journal of the ACM, 64 (5):35:1–35:??, October 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tardos:2017:IAFg

[Tar17h] Eva Tardos. Invited articles foreword. Journal of the ACM, 64 (6):40:1–40:??, November 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tardos:2018:IAFa

[Tar18a] Eva Tardos. Invited article foreword. Journal of the ACM, 65 (2):11:1–11:??, March 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tardos:2018:IAFb

[Tar18b] Eva Tardos. Invited article foreword. Journal of the ACM, 65 (3):16:1–16:??, March 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tardos:2018:IAFc

[Tar18c] Eva Tardos. Invited article foreword. Journal of the ACM, 65(4): 25:1–25:??, August 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tardos:2018:IAFd

[Tar18d] Eva Tardos. Invited article foreword. Journal of the ACM, 65(6): 44:1–44:??, November 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tardos:2019:IAFa

[Tar19a] Eva Tardos. Invited articles foreword. Journal of the ACM, 66 (3):20:1–20:??, June 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3328536.

Tardos:2019:IAFb

[Tar19b] Eva Tardos. Invited articles foreword. Journal of the ACM, 66 (6):43:1–43:??, December 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/ft_gateway.cfm?id=3371337.

Tardos:2020:IAFa

[Tar20a] Eva Tardos. Invited article foreword. Journal of the ACM, 67 (2):12:1, May 2020. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL https://dl.acm.org/doi/abs/
10.1145/3391411.

Tardos:2020:IAFb

[Tar20b] Eva Tardos. Invited articles foreword. Journal of the ACM, 67 (5):29:1, October 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/

doi/10.1145/3418066.

Tardos:2020:IAF

[Tar20c] Eva Tardos. Invited articles foreword. Journal of the ACM, 67(6):35:1, November 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3429262.

Tardos:2021:IAFa

[Tar21a] Eva Tardos. Invited article foreword. Journal of the ACM, 68 (1):6:1, February 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/

doi/10.1145/3442687.

Tardos:2021:IAFc

[Tar21b] Eva Tardos. Invited article foreword. Journal of the ACM, 68 (3):20:1, May 2021. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL https://dl.acm.org/doi/10.
1145/3448013.

Tardos:2021:IAFd

[Tar21c] Eva Tardos. Invited article foreword. Journal of the ACM, 68(4):29:1, August 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/

doi/10.1145/3464772.

Tardos:2021:IAFf

[Tar21d] Eva Tardos. Invited article foreword. Journal of the ACM, 68(6):47:1, December 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3490689.

Tardos:2021:IAFb

[Tar21e] Eva Tardos. Invited articles foreword. Journal of the ACM, 68(2):12e:1, March 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/

doi/10.1145/3456290.

Tardos:2021:IAFe

[Tar21f] Eva Tardos. Invited articles foreword. Journal of the ACM, 68 (5):37:1, October 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/

doi/10.1145/3485044.

Tardos:2022:IAFa

[Tar22a] Eva Tardos. Invited article foreword. Journal of the ACM, 69 (1):7:1, February 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/

doi/10.1145/3505279.

Tardos:2022:IAFb

[Tar22b] Eva Tardos. Invited article foreword. Journal of the ACM, 69(2):16:1, April 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/

doi/10.1145/3519019.

Taubenfeld:2022:ASM

[Tau22] Gadi Taubenfeld. Anonymous shared memory. Journal of the ACM, 69(4):24:1–24:??, August 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/doi/10.1145/3529752.

Tay:1993:OSM

[Tay93] Y. C. Tay. On the optimality of strategies for multiple joins. Journal of the ACM, 40(5):1067–1086, November 1993. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
174151.html.

Tzameret:2021:UIF

[TC21] Iddo Tzameret and Stephen A. Cook. Uniform, integral, and feasible proofs for the determinant identities. Journal of the ACM, 68(2):12:1–12:80, March 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/doi/10.1145/3431922.

tenCate:2009:CQC

[tCL09] Balder ten Cate and Carsten Lutz. The complexity of query containment in expressive fragments of XPath 2.0. Journal of the ACM, 56(6):31:1–31:48, September 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tezuka:1987:DGP

[Tez87] Shu Tezuka. On the discrepancy of GFSR pseudorandom numbers. Journal of the ACM, 34(4):939–949, October 1987. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
31848.html.

Tzelnic:1982:APB

[TG82] Percy Tzelnic and Izidor Gertner. An approach to program behavior modeling and optimal memory control. Journal of the ACM, 29(2):525–554, April 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Thompson:1957:BMS

[Tho57] Gene Thomas Thompson. On Bateman’s method for solving linear integral equations. Journal of the ACM, 4(3):314–328, July 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Thorup:1999:USS

[Tho99] Mikkel Thorup. Undirected single-source shortest paths with positive integer weights in linear time. Journal of the ACM, 46 (3):362–394, May 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org:80/

pubs/citations/journals/jacm/1999-46-3/p362-thorup/.

Thorup:2004:COR

[Tho04] Mikkel Thorup. Compact oracles for reachability and approximate distances in planar digraphs. Journal of the ACM, 51(6): 993–1024, November 2004. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Thorup:2007:EBP

[Tho07] Mikkel Thorup. Equivalence between priority queues and sorting. Journal of the ACM, 54(6):28:1–28:27, December 2007. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tsuda:1970:NIM

[TI70] Takao Tsuda and Kozo Ichida. Nonlinear interpolation of multivariable functions by the Monte Carlo method. Journal of the ACM, 17(2):420–425, April 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tillitt:1958:CPY

[Til58] Harley Tillitt. Computer programming for young students. Journal of the ACM, 5(4):309–318, October 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Titus:1955:GCP

[Tit55] C. K. Titus. A general card-program for the evaluation of the inverse Laplace transform. Journal of the ACM, 2(1):18–27, January 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tiwari:1987:LBC

[Tiw87] Prasoon Tiwari. Lower bounds on communication complexity in distributed computer networks. Journal of the ACM, 34(4):921– 938, October 1987. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/32978.html. Reviews: Computing Re-

views, Vol. 29, No. 9, September 1988.

Tiomkin:1991:NDM

[TK91] Michael Tiomkin and Michael Kaminski. Nonmonotonic default modal logics. Journal of the ACM, 38(4):963–984, October 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/115350.html.

Toyama:1995:TDS

[TKB95] Y. Toyama, J. W. Klop, and H. P. Barendregt. Termination for direct sums of left-linear complete term rewriting systems. Journal of the ACM, 42(6):1275–1304, November 1995. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
227689.html.

Timany:2024:LAT

[TKDB24] Amin Timany, Robbert Krebbers, Derek Dreyer, and Lars Birkedal. A logical approach to type soundness. Journal of the ACM, 71(6):40:1–40:??, December 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3676954.

Tempero:1995:RST

[TL95] Ewan D. Tempero and Richard E. Ladner. Recoverable sequence transmission protocols. Journal of the ACM, 42(5):1059–1090, September 1995. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/210133.html.

Tsuda:1966:NLE

[TM66] Takao Tsuda and Hiroshi Matsumoto. A note on linear extrapolation of multivariable functions by the Monte Carlo method. Journal of the ACM, 13(1):143–150, January 1966. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tourlakis:1973:SRC

[TM73] G. Tourlakis and J. Mylopoulos. Some results in computational topology. Journal of the ACM, 20(3):439–455, July 1973. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Takamizawa:1982:LTC

[TNS82] K. Takamizawa, T. Nishizeki, and N. Saito. Linear-time computability of combinatorial problems in series-parallel graphs. Journal of the ACM, 29(3):623–641, July 1982. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tobagi:1982:DPD

[Tob82] Fouad A. Tobagi. Distributions of packet delay and interdeparture time in slotted ALOHA and carrier sense multiple access. Journal of the ACM, 29(3):907–927, July 1982. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tornheim:1964:CMI

[Tor64] Leonard Tornheim. Convergence of multipoint iterative methods. Journal of the ACM, 11(2):210–220, April 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Toran:1991:CCD

[Tor91] Jacobo Torán. Complexity classes defined by counting quantifiers. Journal of the ACM, 38(3):753–774, July 1991. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
116858.html.

Towsley:1980:QNM

[Tow80] Don Towsley. Queuing network models with state-dependent routing. Journal of the ACM, 27(2):323–337, April 1980. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Toyama:1987:CRP

[Toy87] Yoshihito Toyama. On the Church–Rosser property for the direct sum of term rewriting systems. Journal of the ACM, 34(1):128– 143, January 1987. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/7534.html.

Tsitsiklis:1986:PPB

[TPH86] John N. Tsitsiklis, Christos H. Papadimitriou, and Pierre Humblet. The performance of a precedence-based queueing discipline. Journal of the ACM, 33(3):593–602, July 1986. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
5936.html.

Trainiter:1963:ARA

[Tra63] M. Trainiter. Addressing for random-access storage with multiple bucket capabilities. Journal of the ACM, 10(3):307–315, July 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Trauth:1966:GTA

[Tra66] Charles A. Trauth, Jr. Group-type automata. Journal of the ACM, 13(1):170–175, January 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Trauboth:1969:RFE

[Tra69] H. H. Trauboth. Recursive formulas for the evaluation of the convolution integral. Journal of the ACM, 16(1):63–72, January 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tootill:1971:RPT

[TRA71] J. P. R. Tootill, W. D. Robinson, and A. G. Adams. The runs up-and-down performance of Tausworthe pseudo-random number generators. Journal of the ACM, 18(3):381–399, July 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Trexler:1954:PUC

[Tre54] George F. Trexler. Public utility customer accounting on the Type 650 Magnetic Drum Data Processing Machine. Journal of the ACM, 1(4):173–176, October 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tootill:1973:ART

[TRE73] J. P. R. Tootill, W. D. Robinson, and D. J. Eagle. An asymptotically random Tausworthe sequence. Journal of the ACM, 20(3): 469–481, July 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Trevisan:2001:EPG

[Tre01] Luca Trevisan. Extractors and pseudorandom generators. Journal of the ACM, 48(4):860–879, July 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Trojan:1984:LBF

[Tro84] George M. Trojan. Lower bounds and fast algorithms for sequence acceleration. Journal of the ACM, 31(2):329–335, April 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Trotter:1974:ABV

[TS74] L. E. Trotter, Jr. and C. M. Shetty. An algorithm for the bounded variable integer programming problem. Journal of the ACM, 21(3):505–513, July 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Trivedi:1981:ODL

[TS81] Kishor S. Trivedi and Timothy M. Sigmon. Optimal design of linear storage hierarchies. Journal of the ACM, 28(2):270–288, April 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tsitsiklis:1995:ACC

[TS95] John N. Tsitsiklis and George D. Stamoulis. On the average communication complexity of asynchronous distributed algorithms. Journal of the ACM, 42(2):382–400, March 1995. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
201029.html.

TenCate:2010:TCL

[TS10] Balder Ten Cate and Luc Segoufin. Transitive closure logic, and nested tree walking automata, and Xpath. Journal of the ACM, 57(3):18:1–18:41, March 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tsao:1974:SPE

[Tsa74] Nai-Kuan Tsao. Some a posteriori error bounds in floating-point computations. Journal of the ACM, 21(1):6–17, January 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tay:1985:MVP

[TSG85] Y. C. Tay, Rajan Suri, and Nathan Goodman. A mean value performance model for locking in databases: The no-waiting case. Journal of the ACM, 32(3):618–651, July 1985. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
3831.html.

Tsichritzis:1970:EPS

[Tsi70] D. Tsichritzis. The equivalence problem of simple programs. Journal of the ACM, 17(4):729–738, October 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tsukiyama:1980:AEA

[TSOA80] S. Tsukiyama, I. Shirakawa, H. Ozaki, and H. Ariyoshi. An algorithm to enumerate all cutsets of a graph in linear time per cutset. Journal of the ACM, 27(4):619–632, October 1980. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tantawi:1985:OSL

[TT85] Asser N. Tantawi and Don Towsley. Optimal static load balancing in distributed computer systems. Journal of the ACM, 32 (2):445–465, April 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/3156.html.

Toyoda:1966:AAA

[TTK66] Junichi Toyoda, Yoshikazu Tezuka, and Yoshiro Kasahara. Analysis of the address assignment problem for clustered keys. Journal of the ACM, 13(4):526–532, October 1966. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tabareau:2021:MUP

[TTS21] Nicolas Tabareau, Éric Tanter, and Matthieu Sozeau. The marriage of univalence and parametricity. Journal of the ACM, 68 (1):5:1–5:44, February 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.

org/doi/10.1145/3429979.

Tarjan:1984:WCA

[Tv84] Robert E. Tarjan and Jan van Leeuwen. Worst-case analysis of set union algorithms. Journal of the ACM, 31(2):245–281, April 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Traub:2019:APT

[TV19] Vera Traub and Jens Vygen. Approaching 3/2 for the s–t-path TSP. Journal of the ACM, 66(2):14:1–14:??, April 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/ft_gateway.cfm?id=3309715.

Totschek:1961:IRT

[TW61] R. Totschek and R. C. Wood. An investigation of real-time solution of the transportation problem. Journal of the ACM, 8 (2):230–239, April 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Traub:1979:CCN

[TW79] J. F. Traub and H. Woźniakowski. Convergence and complexity of Newton iteration for operator equations. Journal of the ACM, 26(2):250–258, April 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Traub:1984:OSL

[TW84] J. F. Traub and H. Woźniakowski. On the optimal solution of large linear systems. Journal of the ACM, 31(3):545–559, July 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tripathi:1988:VAT

[TW88] Satish K. Tripathi and C. Murray Woodside. A vertex-allocation theorem for resources in queuing networks. Journal of the ACM, 35(1):221–230, January 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/45068.html.

Thilikos:2024:KV

[TW24] Dimitrios M. Thilikos and Sebastian Wiederrecht. Killing a vortex. Journal of the ACM, 71(4):27:1–27:??, August 2024. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL https://dl.acm.org/doi/10.1145/3664648.

Twomey:1963:NSF

[Two63] S. Twomey. On the numerical solution of Fredholm integral equations of the first kind by the inversion of the linear system produced by quadrature. Journal of the ACM, 10:97–101, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Trivedi:1980:OSC

[TWS80] Kishor S. Trivedi, Robert A. Wagner, and Timothy M. Sigmon. Optimal selection of CPU speed, device capacities, and file assignments. Journal of the ACM, 27(3):457–473, July 1980. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tang:2023:TBU

[TWZ23] Zhihao Gavin Tang, Xiaowei Wu, and Yuhao Zhang. Toward a better understanding of randomized greedy matching. Journal of the ACM, 70(6):39:1–39:??, December 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3614318.

Tarjan:1979:SST

[TY79] Robert Endre Tarjan and Andrew Chi-Chih Yao. Storing a sparse table. Communications of the Association for Computing Machinery, 22(11):606–611, November 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [FKS84].

Thorup:2005:ADO

[TZ05] Mikkel Thorup and Uri Zwick. Approximate distance oracles. Journal of the ACM, 52(1):1–24, January 2005. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Thapper:2016:CFV

[TZ16] Johan Thapper and Stanislav Zivný. The complexity of finitevalued CSPs. Journal of the ACM, 63(4):37:1–37:??, November 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Traub:2025:BTA

[TZ25] Vera Traub and Rico Zenklusen. Better-than-2 approximations for weighted tree augmentation and applications to Steiner tree. Journal of the ACM, 72(2):16:1–16:??, April 2025. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Uresin:1990:PAA

[ÜD90] Aydin Üresin and Michel Dubois. Parallel asynchronous algorithms for discrete data. Journal of the ACM, 37(3):588– 606, July 1990. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/79162.html.

Ukkonen:1982:EPS

[Ukk82] Esko Ukkonen. The equivalence problem for some non-real-time deterministic pushdown automata. Journal of the ACM, 29(4): 1166–1181, October 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ullman:1969:HSA

[Ull69] J. D. Ullman. Halting stack automata. Journal of the ACM, 16 (4):550–563, October 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ullman:1972:NEH

[Ull72] J. D. Ullman. A note on the efficiency of hashing functions. Journal of the ACM, 19(3):569–575, July 1972. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). Early work on the problem of finding optimal hash functions for open addressing.

Ullmann:1976:ASI

[Ull76] J. R. Ullmann. An algorithm for subgraph isomorphism. Journal of the ACM, 23(1):31–42, January 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Unruh:2015:RQT

[Unr15] Dominique Unruh. Revocable quantum timed-release encryption. Journal of the ACM, 62(6):49:1–49:??, December 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Upfal:1984:ESP

[Upf84] Eli Upfal. Efficient schemes for parallel communication. Journal of the ACM, 31(3):507–517, July 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Upfal:1992:DPR

[Upf92] Eli Upfal. An O(log N) deterministic packet-routing scheme. Journal of the ACM, 39(1):55–70, January 1992. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
147517.html.

Urquhart:1987:HER

[Urq87] Alasdair Urquhart. Hard examples for resolution. Journal of the ACM, 34(1):209–219, January 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/8928.html.

Usmani:1966:BVT

[Usm66] Riaz A. Usmani. Boundary value techniques for the numerical solution of certain initial value problems in ordinary differential equations. Journal of the ACM, 13(2):287–295, April 1966. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ullman:1988:ETT

[UV88] Jeffrey D. Ullman and Allen Van Gelder. Efficient tests for top-down termination of logical rules. Journal of the ACM, 35 (2):345–373, April 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/42285.html.

Upfal:1987:HSM

[UW87] Eli Upfal and Avi Wigderson. How to share memory in a distributed system. Journal of the ACM, 34(1):116–127, January 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/7926.html.

Valiant:1975:RRP

[Val75] Leslie G. Valiant. Regularity and related problems for deterministic pushdown automata. Journal of the ACM, 22(1):1–10, January 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Valiant:2000:NAC

[Val00] Leslie G. Valiant. A neuroidal architecture for cognitive computation. Journal of the ACM, 47(5):854–882, 2000. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/citations/journals/jacm/
2000-47-5/p854-valiant/.

Valiant:2003:TPC

[Val03] Leslie G. Valiant. Three problems in computer science. Journal of the ACM, 50(1):96–99, January 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Valiant:2009:E

[Val09] Leslie G. Valiant. Evolvability. Journal of the ACM, 56(1):3:1– 3:21, January 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Valiant:2015:FCS

[Val15] Gregory Valiant. Finding correlations in subquadratic time, with applications to learning parities and the closest pair problem. Journal of the ACM, 62(2):13:1–13:??, May 2015. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

VanGelder:1967:SNR

[Van67] A. Van Gelder. Some new results in pseudo-random number generation. Journal of the ACM, 14(4):785–792, October 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

vanWestrhenen:1972:SST

[van72] S. C. van Westrhenen. Statistical studies of theoremhood in classical propositional and first order predicate calculus. Journal of the ACM, 19(2):347–365, April 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vantilborgh:1978:EAE

[Van78] Hendrik Vantilborgh. Exact aggregation in exponential queueing networks. Journal of the ACM, 25(4):620–629, October 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vantilborgh:1985:AE

[Van85] Hendrik Vantilborgh. Aggregation with an error of O(ϵ^{2}). Journal of the ACM, 32(1):162–190, January 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
214107.html.

Vansummeren:2007:DWD

[Van07] Stijn Vansummeren. On deciding well-definedness for query languages on trees. Journal of the ACM, 54(4):19:1–19:37, July 2007. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vazirani:2012:NRC

[Vaz12] Vijay V. Vazirani. The notion of a rational convex program, and an algorithm for the Arrow–Debreu Nash bargaining game. Journal of the ACM, 59(2):7:1–7:36, April 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Veanes:2017:MD

[VBNB17] Margus Veanes, Nikolaj Bjørner, Lev Nachmanson, and Sergey Bereg. Monadic decomposition. Journal of the ACM, 64(2):14:1– 14:??, June 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

vanBeek:1995:MGC

[vD95] Peter van Beek and Rina Dechter. On the minimality and global consistency of row-convex constraint networks. Journal of the ACM, 42(3):543–561, May 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/210347.html.

vanBeek:1997:CTL

[vD97] Peter van Beek and Rina Dechter. Constraint tightness and looseness versus local and global consistency. Journal of the ACM, 44(4):549–566, July 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:/

/www.acm.org:80/pubs/citations/journals/jacm/1997-44-

4/p549-van_beek/. See erratum [ZY03].

vandeLiefvoort:1986:MAS

[vdLL86] Appie van de Liefvoort and Lester Lipsky. A matrix-algebraic solution to two K_{m} servers in a loop. Journal of the ACM, 33(1):207–223, January 1986. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/5391.html.

vandeSnepscheut:1989:DSS

[vdSS89] Jan L. A. van de Snepscheut and Johan B. Swenker. On the design of some systolic algorithms. Journal of the ACM, 36(4):826– 840, October 1989. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/76365.html.

Vempala:2010:RSB

[Vem10] Santosh S. Vempala. A random-sampling-based algorithm for learning intersections of halfspaces. Journal of the ACM, 57(6): 32:1–32:14, October 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Venkataraman:1987:DPE

[Ven87] K. N. Venkataraman. Decidability of the purely existential fragment of the theory of term algebras. Journal of the ACM, 34 (2):492–510, April 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/24037.html.

Verma:1995:TUH

[Ver95] Rakesh M. Verma. A theory of using history for equational systems with applications. Journal of the ACM, 42(5):984–1020, September 1995. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/210130.html.

Vitter:2006:DSR

[VH06] Jeffrey Scott Vitter and David Alexander Hutchinson. Distribution sort with randomized cycling. Journal of the ACM, 53(4): 656–680, July 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:1987:DFD

[Via87] Victor Vianu. Dynamic functional dependencies and database aging. Journal of the ACM, 34(1):28–59, January 1987. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
7918.html.

Vianu:2010:EJS

[Via10a] Victor Vianu. Editorial: JACM at the start of a new decade. Journal of the ACM, 57(3):12:1–12:3, March 2010. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2010:IAS

[Via10b] Victor Vianu. Invited articles section foreword. Journal of the ACM, 58(1):1:1, December 2010. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2011:IJI

[Via11a] Victor Vianu. Introduction to JACM invited article. Journal of the ACM, 58(3):12:1, May 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2011:IAF

[Via11b] Victor Vianu. Invited article foreword. Journal of the ACM, 58 (6):29:1, December 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2011:IAFa

[Via11c] Victor Vianu. Invited articles foreword. Journal of the ACM, 58 (4):16:1, July 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2011:IAFb

[Via11d] Victor Vianu. Invited articles foreword. Journal of the ACM, 58 (5):22:1–22:??, October 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2012:IAFa

[Via12a] Victor Vianu. Invited article foreword. Journal of the ACM, 59(1):4:1, February 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2012:IAFb

[Via12b] Victor Vianu. Invited article foreword. Journal of the ACM, 59 (2):9:1, April 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2012:IAFc

[Via12c] Victor Vianu. Invited article foreword. Journal of the ACM, 59 (3):15:1, June 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2012:IAFd

[Via12d] Victor Vianu. Invited article foreword. Journal of the ACM, 59(4):20:1, August 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2012:IAF

[Via12e] Victor Vianu. Invited article foreword. Journal of the ACM, 59(5):26:1, October 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2013:EJR

[Via13a] Victor Vianu. Editorial: JACM redux. Journal of the ACM, 60(5):30:1–30:2, October 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2013:IAFa

[Via13b] Victor Vianu. Invited article foreword. Journal of the ACM, 60(1):5:1, February 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2013:IAFb

[Via13c] Victor Vianu. Invited article foreword. Journal of the ACM, 60 (2):14:1, April 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2013:IAFc

[Via13d] Victor Vianu. Invited article foreword. Journal of the ACM, 60 (3):21:1, June 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2013:IAFe

[Via13e] Victor Vianu. Invited article foreword. Journal of the ACM, 60(5):38:1, October 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2013:IAFd

[Via13f] Victor Vianu. Invited articles foreword. Journal of the ACM, 60(4):27:1, August 2013. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2014:FIA

[Via14a] Victor Vianu. Foreword to invited articles section. Journal of the ACM, 61(1):6:1–6:??, January 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2014:IAFb

[Via14b] Victor Vianu. Invited article foreword. Journal of the ACM, 61 (5):31:1–31:??, August 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2014:IAF

[Via14c] Victor Vianu. Invited article foreword. Journal of the ACM, 61 (6):40:1–40:??, November 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2014:IAFa

[Via14d] Victor Vianu. Invited articles foreword. Journal of the ACM, 61 (4):25:1–25:??, July 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2015:IAFb

[Via15a] Victor Vianu. Invited article foreword. Journal of the ACM, 62 (2):16:1–16:??, May 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2015:IAF

[Via15b] Victor Vianu. Invited article foreword. Journal of the ACM, 62 (3):24:1–24:??, June 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2015:IAFa

[Via15c] Victor Vianu. Invited articles foreword. Journal of the ACM, 62 (1):6:1–6:??, February 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vitter:1983:ASP

[Vit83] Jeffrey Scott Vitter. Analysis of the search performance of coalesced hashing. Journal of the ACM, 30(2):231–258, April 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vitter:1987:DAD

[Vit87] Jeffrey Scott Vitter. Design and analysis of dynamic Huffman codes. Journal of the ACM, 34(4):825–845, October 1987. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/toc/Abstracts/0004-5411/
42227.html.

Varghese:2000:FSC

[VJ00] George Varghese and Mahesh Jayaram. The fault span of crash failures. Journal of the ACM, 47(2):244–293, March 2000. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL http://www.acm.org/pubs/citations/journals/jacm/
2000-47-2/p244-varghese/.

VanEmden:1976:SPL

[VK76] M. H. Van Emden and R. A. Kowalski. The semantics of predicate logic as a programming language. Journal of the ACM, 23 (4):733–742, October 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vitter:1996:OPD

[VK96] Jeffrey Scott Vitter and P. Krishnan. Optimal prefetching via data compression. Journal of the ACM, 43(5):771–793, September 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
jacm/234753.html.

vanKreveld:1993:UCS

[vO93] Marc J. van Kreveld and Mark H. Overmars. Union-copy structures and dynamic segment trees. Journal of the ACM, 40(3): 635–652, July 1993. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/174140.html.

Vocking:2003:HAH

[Vöc03] Berthold Vöcking. How asymmetry helps load balancing. Journal of the ACM, 50(4):568–589, July 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

vonHoldt:1956:IPC

[von56] Richard Elton von Holdt. An iterative procedure for the calculation of the eigenvalues and eigenvectors of a real symmetric matrix. Journal of the ACM, 3(3):223–238, July 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

vonHoldt:1962:ITD

[von62] Richard E. von Holdt. Inversion of triple-diagonal compound matrices. Journal of the ACM, 9(1):71–83, January 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

vonNeumann:1966:TSR

[von66] John von Neumann. Theory of Self-Reproducing Automata. University of Illinois Press, Urbana, IL, USA, 1966. xix + 388 pp. LCCN QA267 .V55. Von Neumann’s work on self-reproducing automata, completed and edited after his death by Arthur Burks. Also includes transcripts of von Neumann’s 1949 University of Illinois lectures on the “Theory and Organization of Complicated Automata”. See [Bur70].

vanGelder:1991:WFS

[vRS91] Allen van Gelder, Kenneth Ross, and John S. Schlipf. The wellfounded semantics for general logic programs. Journal of the ACM, 38(3):620–650, July 1991. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/116838.html. Prelimi-

nary abstract appeared in Seventh ACM Symposium on Principles of Database Systems, March 1988, pp. 221–230.

Vale:2024:CTL

[VSC24] Arthur Oliveira Vale, Zhong Shao, and Yixuan Chen. A compositional theory of linearizability. Journal of the ACM, 71(2): 14:1–14:??, April 2024. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/

doi/10.1145/3643668.

Valentine:1963:AMP

[VV63] Charles W. Valentine and C. Peter Van Dine. An algorithm for mimimax polynomial curve-fitting of discrete data. Journal of the ACM, 10:283–290, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Vianu:2009:IPS

[VV09] Victor Vianu and Jan Van den Bussche. Introduction to PODS 2006 special section. Journal of the ACM, 56(3):11:1, May 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Volzer:2012:DFR

[VV12] Hagen Völzer and Daniele Varacca. Defining fairness in reactive and concurrent systems. Journal of the ACM, 59(3):13:1–13:37, June 2012. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Valiant:2017:EUI

[VV17] Gregory Valiant and Paul Valiant. Estimating the unseen: Improved estimators for entropy and other properties. Journal of the ACM, 64(6):37:1–37:??, November 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

VandenBussche:1997:COC

[VVAG97] Jan Van den Bussche, Dirk Van Gucht, Marc Andries, and Marc Gyssens. On the completeness of object-creating database transformation languages. Journal of the ACM, 44(2):272–319, March 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

vanGlabbeek:1996:BTA

[vW96] Rob J. van Glabbeek and W. Peter Weijland. Branching time and abstraction in bisimulation semantics. Journal of the ACM, 43(3):555–600, May 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/jacm/233556.html.

Vazirani:2011:MEU

[VY11] Vijay V. Vazirani and Mihalis Yannakakis. Market equilibrium under separable, piecewise-linear, concave utilities. Journal of the ACM, 58(3):10:1–10:25, May 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wadel:1954:EDA

[Wad54] Louis B. Wadel. An electronic differential analyzer as a difference analyzer. Journal of the ACM, 1(3):128–136, July 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wadel:1956:SDF

[Wad56] Louis B. Wadel. Simulation of digital filters on an electronic analog computer. Journal of the ACM, 3(1):16–21, January 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wadey:1960:FPA

[Wad60] W. G. Wadey. Floating-point arithmetics. Journal of the ACM, 7(2):129–139, April 1960. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wagner:1976:SPA

[Wag76] Robert A. Wagner. A shortest path algorithm for edge-sparse graphs. Journal of the ACM, 23(1):50–57, January 1976. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wagner:1997:EUE

[Wag97] Robert A. Wagner. Evaluating uniform expressions within two steps of minimum parallel time. Journal of the ACM, 44(2):345– 361, March 1997. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Waite:1967:PDM

[Wai67] William M. Waite. Path detection in multidimensional iterative arrays. Journal of the ACM, 14(2):300–310, April 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Waksman:1968:CPN

[Wak68a] Abraham Waksman. Corrigendum: “A permutation network”. Journal of the ACM, 15(2):340, April 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Wak68b].

Waksman:1968:PN

[Wak68b] Abraham Waksman. A permutation network. Journal of the ACM, 15(1):159–163, January 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [Wak68a].

Waksman:1969:MR

[Wak69] Abraham Waksman. A model of replication. Journal of the ACM, 16(1):178–188, January 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Walsh:1961:CFM

[Wal61] John E. Walsh. Computer-feasible method for handling incomplete data in regression analysis. Journal of the ACM, 8(2):201– 211, April 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Walther:1988:MSU

[Wal88] Christoph Walther. Many-sorted unification. Journal of the ACM, 35(1):1–17, January 1988. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/45071.html.

Wang:1957:VTT

[Wan57] Hao Wang. A variant to Turing’s theory of computing machines. Journal of the ACM, 4(1):63–92, January 1957. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wang:1974:ACN

[Wan74a] Chung C. Wang. An algorithm for the chromatic number of a graph. Journal of the ACM, 21(3):385–391, July 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wang:1974:UEZ

[Wan74b] Paul S. Wang. The undecidability of the existence of zeros of real elementary functions. Journal of the ACM, 21(4):586–589, October 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wand:1978:NIR

[Wan78] Mitchell Wand. A new incompleteness result for Hoare’s system. Journal of the ACM, 25(1):168–175, January 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wand:1980:CBP

[Wan80] Mitchell Wand. Continuation-based program transformation strategies. Journal of the ACM, 27(1):164–180, January 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wang:1993:MRE

[Wan93] Tie-Cheng Wang. Z-module reasoning: an equality-oriented proving method with built-in ring axioms. Journal of the ACM, 40(3):558–606, July 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/174137.html.

Wang:2023:NAE

[Wan23] Haitao Wang. A new algorithm for Euclidean shortest paths in the plane. Journal of the ACM, 70(2):11:1–11:??, April 2023. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3580475.

Ward:1957:HMS

[War57] James A. Ward. The down-hill method of solving f(z) = 0. Journal of the ACM, 4(2):148–150, April 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Warshall:1962:TBM

[War62] Stephen Warshall. A theorem on Boolean matrices. Journal of the ACM, 9(1):11–12, January 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wasel:1954:MDP

[Was54] A. D. Wasel. A method of determining plate bending by use of a punched-card machine. Journal of the ACM, 1(3):105–110, July 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wasserstrom:1971:SBV

[Was71] E. Wasserstrom. Solving boundary-value problems by imbedding. Journal of the ACM, 18(4):594–602, October 1971. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wasilkowski:1980:CSI

[Was80] G. W. Wasilkowski. Can any stationary iteration using linear information be globally convergent? Journal of the ACM, 27 (2):263–269, April 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wasilkowski:1981:ECM

[Was81] G. W. Wasilkowski. n-evaluation conjecture for multipoint iterations for the solution of scalar nonlinear equations. Journal of the ACM, 28(1):71–80, January 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Watanabe:1961:SSS

[Wat61] Shigere Watanabe. 5-symbol 8-state and 5-symbol 6-state universal Turing machines. Journal of the ACM, 8(4):476–483, October 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Watson:1973:AIB

[Wat73] G. A. Watson. An algorithm for the inversion of block matrices of Toeplitz form. Journal of the ACM, 20(3):409–415, July 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wasserman:1997:SRR

[WB97] Hal Wasserman and Manuel Blum. Software reliability via run-time result-checking. Journal of the ACM, 44(6):826–849, November 1997. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org:80/pubs/
citations/journals/jacm/1997-44-6/p826-wasserman/.

Wong:1974:CPR

[WC74] C. K. Wong and Don Coppersmith. A combinatorial problem related to multimodule memory organizations. Journal of the ACM, 21(3):392–402, July 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wong:1976:BSE

[WC76] C. K. Wong and Ashok K. Chandra. Bounds for the string editing problem. Journal of the ACM, 23(1):13–16, January 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Weeg:1960:TEG

[Wee60] Gerard P. Weeg. Truncation error in the Graeffe root-squaring method. Journal of the ACM, 7(1):69–71, January 1960. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Weeg:1962:SAO

[Wee62] G. P. Weeg. The structure of an automaton and its operationpreserving transformation group. Journal of the ACM, 9(3):345– 349, July 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Weeg:1965:AGD

[Wee65] G. P. Weeg. The automorphism group of the direct product of strongly related automata. Journal of the ACM, 12(2):187–195, April 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [Wee67].

Weeks:1966:NIL

[Wee66] William T. Weeks. Numerical inversion of Laplace transforms using Laguerre functions. Journal of the ACM, 13(3):419–429, July 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Weeg:1967:CAG

[Wee67] G. P. Weeg. Corrigendum: “The Automorphism Group of the Direct Product of Strongly Related Automata”. Journal of the ACM, 14(2):421, April 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Wee65].

Wegner:1969:TNF

[Weg69] Peter Wegner. Translation networks and function composition. Journal of the ACM, 16(2):261–263, April 1969. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wegbreit:1976:VPP

[Weg76] Ben Wegbreit. Verifying program performance. Journal of the ACM, 23(4):691–699, October 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wegbreit:1977:CSI

[Weg77] Ben Wegbreit. Complexity of synthesizing inductive assertions. Journal of the ACM, 24(3):504–512, July 1977. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wegener:1988:CBP

[Weg88] Ingo Wegener. On the complexity of branching programs and decision trees for clique functions. Journal of the ACM, 35(2):461– 471, April 1988. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/46161.html.

Weil:1955:RRM

[Wei55] Herschel Weil. Reduction of runs in multiparameter computations. Journal of the ACM, 2(2):99–110, April 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Weintraub:1962:CBP

[Wei62] Sol Weintraub. Cumulative binomial probabilities. Journal of the ACM, 9(3):405–407, July 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Weiss:1966:VPI

[Wei66] Ruth A. Weiss. BE VISION, a package of IBM 7090 FORTRAN programs to draw orthographic views of combinations of plane and quadric surfaces. Journal of the ACM, 13(2):194–204, April 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Weinblatt:1972:NSA

[Wei72] Herbert Weinblatt. A new search algorithm for finding the simple cycles of a finite directed graph. Journal of the ACM, 19(1):43– 56, January 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Welch:1966:MAC

[Wel66] John T. Welch, Jr. A mechanical analysis of the cyclic structure of undirected linear graphs. Journal of the ACM, 13(2):205–210, April 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wein:2025:KHT

[WEM25] Alexander Wein, Ahmed El Alaoui, and Cristopher Moore. The Kikuchi hierarchy and tensor PCA. Journal of the ACM, 72(5): 35:1–35:40, October 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Werschulz:1979:MOO

[Wer79] Arthur G. Werschulz. Maximal order and order of information for numerical quadrature. Journal of the ACM, 26(3):527–537, July 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Westlake:1967:URN

[Wes67] W. J. Westlake. A uniform random number generator based on the combination of two congruential generators. Journal of the ACM, 14(2):337–340, April 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wagner:1974:SSC

[WF74] Robert A. Wagner and Michael J. Fischer. The string-to-string correction problem. Journal of the ACM, 21(1):168–173, January 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Worrell:1973:EOS

[WH73] R. B. Worrell and B. L. Hulme. Efficient ordering of set expressions for symbolic expansion. Journal of the ACM, 20(3): 482–488, July 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ward:1980:STM

[WH80] Stephen A. Ward and Robert H. Halstead, Jr. A syntactic theory of message passing. Journal of the ACM, 27(2):365–383, April 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wigderson:1983:IPG

[Wig83] Avi Wigderson. Improving the performance guarantee for approximate graph coloring. Journal of the ACM, 30(4):729–735, October 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Williams:1954:ACM

[Wil54] S. B. Williams. The Association for Computing Machinery. Journal of the ACM, 1(1):1–3, January 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wilf:1959:SCN

[Wil59a] Herbert S. Wilf. A stability criterion for numerical integration. Journal of the ACM, 6(3):363–365, July 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wilkinson:1959:SRM

[Wil59b] J. H. Wilkinson. Stability of the reduction of a matrix to almost triangular and triangular forms by elementary similarity transformations. Journal of the ACM, 6(3):336–359, July 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wilkinson:1961:EAD

[Wil61] J. H. Wilkinson. Error analysis of direct methods of matrix inversion. Journal of the ACM, 8(3):281–330, July 1961. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Williams:1962:APS

[Wil62] Leland H. Williams. Algebra of polynomials in several variables for a digital computer. Journal of the ACM, 9(1):29–40, January 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wilkes:1968:CTN

[Wil68] Maurice V. Wilkes. Computers then and now. Journal of the ACM, 15(1):1–7, January 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Willis:1970:CCP

[Wil70] David G. Willis. Computational complexity and probability constructions. Journal of the ACM, 17(2):241–259, April 1970. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wilkinson:1971:SCN

[Wil71] J. H. Wilkinson. Some comments from a numerical analyst. Journal of the ACM, 18(2):137–147, April 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wilhelm:1977:GMP

[Wil77] Neil C. Wilhelm. A general model for the performance of disk systems. Journal of the ACM, 24(1):14–31, January 1977. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wilf:1978:GBA

[Wil78] Herbert S. Wilf. A global bisection algorithm for computing the zeros of polynomials in the complex plane. Journal of the ACM, 25(3):415–420, July 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Williamson:1984:DFS

[Wil84] S. G. Williamson. Depth-first search and Kuratowski subgraphs. Journal of the ACM, 31(4):681–693, October 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Willard:1987:MST

[Wil87] Dan E. Willard. Multidimensional search trees that provide new types of memory reductions. Journal of the ACM, 34(4):846– 858, October 1987. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/42228.html.

Willard:1991:OSC

[Wil91] Dan E. Willard. Optimal sample cost residues for differential database batch query problems. Journal of the ACM, 38(1):104– 119, January 1991. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/102786.html.

Williams:2014:NAC

[Wil14] Ryan Williams. Nonuniform ACC circuit lower bounds. Journal of the ACM, 61(1):2:1–2:??, January 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Winograd:1964:IEL

[Win64] S. Winograd. Input-error-limiting automata. Journal of the ACM, 11(3):338–351, July 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Winograd:1965:TRP

[Win65a] S. Winograd. On the time required to perform addition. Journal of the ACM, 12(2):277–285, April 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Winters:1965:MML

[Win65b] William K. Winters. A modified method of latent class analysis for file organization in information retrieval. Journal of the ACM, 12(3):356–363, July 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Winograd:1967:TRP

[Win67] S. Winograd. On the time required to perform multiplication. Journal of the ACM, 14(4):793–802, October 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Winder:1971:CPT

[Win71] Robert O. Winder. Chow parameters in threshold logic. Journal of the ACM, 18(2):265–289, April 1971. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Winograd:1975:PEC

[Win75] S. Winograd. On the parallel evaluation of certain arithmetic expressions. Journal of the ACM, 22(4):477–492, October 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Winker:1982:GVF

[Win82] Steve Winker. Generation and verification of finite models and counterexamples using an automated theorem prover answering two open questions. Journal of the ACM, 29(2):273–284, April 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wirth:1968:CPP

[Wir68a] Niklaus Wirth. Corrigendum: “PL360, a programming language for the 360 computers”. Journal of the ACM, 15(4):489, October 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Wir68b].

Wirth:1968:PPL

[Wir68b] Niklaus Wirth. PL360, a programming language for the 360 computers. Journal of the ACM, 15(1):37–74, January 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [Wir68a].

Wiseman:1954:LIP

[Wis54] R. T. Wiseman. Life insurance premium billing and combined operations by electronic equipment. Journal of the ACM, 1(1): 7–12, January 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Williams:1957:CFE

[WJR57] Theodore J. Williams, R. Curtis Johnson, and Arthur Rose. Computations in the field of engineering chemistry. Journal of the ACM, 4(4):393–419, October 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wigderson:2015:IAF

[WK15] Avi Wigderson and Phokion Kolaitis. Invited articles foreword. Journal of the ACM, 62(5):41:1–41:??, November 2015. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Weinberger:1957:SDE

[WL57] A. Weinberger and H. Loberman. Symbolic designations for electrical connections. Journal of the ACM, 4(4):420–427, October 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Willard:1985:ARR

[WL85] Dan E. Willard and George S. Lueker. Adding range restriction capability to dynamic data structures. Journal of the ACM, 32 (3):597–617, July 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0004-5411/3839.html.

Wang:1985:FVS

[WLS85] Ching-Chy Wang, Errol L. Lloyd, and Mary Lou Soffa. Feedback vertex sets and cyclically reducible graphs. Journal of the ACM, 32(2):296–313, April 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/3159.html.

Wojcik:1974:ASR

[WM74] Anthony S. Wojcik and Gernot Metze. An analysis of some relationships between post and Boolean algebras. Journal of the ACM, 21(4):680–696, October 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wolman:1965:FOC

[Wol65] Eric Wolman. A fixed optimum cell-size for records of various lengths. Journal of the ACM, 12(1):53–70, January 1965. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wu:2003:PDC

[WR03] Wei Biao Wu and Chinya V. Ravishankar. The performance of difference coding for sets and relational tables. Journal of the ACM, 50(5):665–693, September 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wos:1965:ECS

[WRC65] Lawrence Wos, George A. Robinson, and Daniel F. Carson. Efficiency and completeness of the set of support strategy in theorem proving. Journal of the ACM, 12(4):536–541, October 1965. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wos:1967:CDT

[WRCS67] Lawrence Wos, George A. Robinson, Daniel F. Carson, and Leon Shalla. The concept of demodulation in theorem proving. Journal of the ACM, 14(4):698–709, October 1967. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wright:1975:CMP

[Wri75] J. W. Wright. The change-making problem. Journal of the ACM, 22(1):125–128, January 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wegbreit:1976:PPC

[WS76] Ben Wegbreit and Jay M. Spitzen. Proving properties of complex data structures. Journal of the ACM, 23(2):389–396, April 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wheeler:1959:LEM

[WSD59] D. J. Wheeler and H. P. F. Swinnerton-Dyer. Letter to the editor: “A method for transposing a matrix”. Journal of the ACM, 6 (2):305, April 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [Ber58].

Wirth:2019:NPD

[WSY^{+}19] Fabian R. Wirth, Sonja Stüdli, Jia Yuan Yu, Martin Corless, and Robert Shorten. Nonhomogeneous place-dependent Markov chains, unsynchronised AIMD, and optimisation. Journal of the ACM, 66(4):24:1–24:??, August 2019. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=3312741.

Wunderlich:1967:SPD

[Wun67] Marvin C. Wunderlich. Sieving procedures on a digital computer. Journal of the ACM, 14(1):10–19, January 1967. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Williams:1973:NYA

[WW73] Thomas A. Williams and Gregory P. White. A note on Yen’s algorithm for finding the length of all shortest paths in N-node nonnegative-distance networks. Journal of the ACM, 20(3):389– 390, July 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Williams:2018:SEB

[WW18] Virginia Vassilevska Williams and R. Ryan Williams. Subcubic equivalences between path, matrix, and triangle problems. Journal of the ACM, 65(5):27:1–27:??, September 2018. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/ft_gateway.cfm?id=3186893.

Wan:2025:ECR

[WXSD25] Jun Wan, Hanshen Xiao, Elaine Shi, and Srinivas Devadas. Expected constant round Byzantine broadcast under dishonest majority. Journal of the ACM, 72(4):30:1–30:??, August 2025. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wang:2014:NAI

[WYCF14] Xiaojing Wang, Qizhao Yuan, Hongliang Cai, and Jiajia Fang. A new approach to image sharing with high-security threshold structure. Journal of the ACM, 61(6):39:1–39:??, November 2014. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wang:1995:DMR

[WZC95] Ke Wang, Weining Zhang, and Siu-Cheung Chau. Decomposition of magic rewriting. Journal of the ACM, 42(2):329– 381, March 1995. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/201027.html.

Xu:2006:FAA

[XB06] Jinbo Xu and Bonnie Berger. Fast and accurate algorithms for protein side-chain packing. Journal of the ACM, 53(4):533–557, July 2006. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Xu:1983:RNR

[XDB83] Mei-Rui Xu, John E. Doner, and Ronald V. Book. Refining nondeterminism in relativizations of complexity classes. Journal of the ACM, 30(3):677–685, July 1983. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Xu:2025:BDP

[XZ25] Yunbei Xu and Assaf Zeevi. Bayesian design principles for frequentist sequential learning. Journal of the ACM, 72(5):37:1– 37:65, October 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yakimovsky:1976:BOD

[Yak76] Yoram Yakimovsky. Boundary and object detection in real world images. Journal of the ACM, 23(4):599–618, October 1976. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yannakakis:1979:ECR

[Yan79] Mihalis Yannakakis. The effect of a connectivity requirement on the complexity of maximum subgraph problems. Journal of the ACM, 26(4):618–630, October 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yannakakis:1982:TSL

[Yan82] Mihalis Yannakakis. A theory of safe locking policies in database systems. Journal of the ACM, 29(3):718–740, July 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yannakakis:1984:SL

[Yan84] Mihalis Yannakakis. Serializability by locking. Journal of the ACM, 31(2):227–244, April 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yannakakis:1985:PAM

[Yan85] Mihalis Yannakakis. A polynomial algorithm for the min-cut linear arrangement of trees. Journal of the ACM, 32(4):950– 988, October 1985. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0004-5411/4228.html.

Yao:1980:NAP

[Yao80] Andrew Chi-Chih Yao. New algorithms for bin packing. Journal of the ACM, 27(2):207–227, April 1980. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yao:1981:LBF

[Yao81a] Andrew Chi-Chih Yao. A lower bound to finding convex hulls. Journal of the ACM, 28(4):780–787, October 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yao:1981:STS

[Yao81b] Andrew Chi Chih Yao. Should tables be sorted? Journal of the ACM, 28(3):615–628, July 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yao:1982:PCK

[Yao82] Andrew Chi Chih Yao. On parallel computation for the knapsack problem. Journal of the ACM, 29(3):898–903, July 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yao:1985:UHO

[Yao85] Andrew C. Yao. Uniform hashing is optimal. Journal of the ACM, 32(3):687–693, July 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/3836.html. Also pub-

lished in/as: Stanford Un., CSD, TR-CS-85-1038, Jan. 1985.

Yao:2003:CPC

[Yao03] Andrew Chi-Chih Yao. Classical physics and the Church–Turing Thesis. Journal of the ACM, 50(1):100–105, January 2003. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yasuhara:1967:RPN

[Yas67] Ann Yasuhara. A remark on post normal systems. Journal of the ACM, 14(1):167–171, January 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Young:1999:GUS

[YCW99] F. Y. Young, Chris C. N. Chu, and D. F. Wong. Generation of universal series-parallel Boolean functions. Journal of the ACM, 46(3):416–435, May 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org:80/

pubs/citations/journals/jacm/1999-46-3/p416-young/.

Yu:1993:AMD

[YDL93] Philip S. Yu, Daniel M. Dias, and Stephen S. Lavenberg. On the analytical modeling of database concurrency control. Journal of the ACM, 40(4):831–872, September 1993. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0004-5411/153733.html.

Yeh:1968:GPA

[Yeh68] Raymond T. Yeh. Generalized pair algebra with applications to automata theory. Journal of the ACM, 15(2):304–316, April 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yekhanin:2008:TQL

[Yek08] Sergey Yekhanin. Towards 3-query locally decodable codes of subexponential length. Journal of the ACM, 55(1):1:1–1:16, February 2008. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yelowitz:1975:DPC

[Yel75] Lawrence Yelowitz. Derivation of a path-connectivity matrix for tagged flowcharts. Journal of the ACM, 22(1):145–154, January 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yen:1972:FLA

[Yen72] Jin Y. Yen. Finding the lengths of all shortest paths in N-node nonnegative-distance complete networks using 1/2N^{3} additions and N^{3} comparisons. Journal of the ACM, 19(3):423–424, July 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yershov:1965:OVM

[Yer65] A. P. Yershóv. One view of man-machine interaction (translated by Nicholas Zvegintzov). Journal of the ACM, 12(3):315–325, July 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yershov:1966:AAP

[Yer66] A. P. Yershóv. ALPHA—an automatic programming system of high efficiency. Journal of the ACM, 13(1):17–24, January 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yi:2012:DIO

[Yi12] Ke Yi. Dynamic indexability and the optimality of B-trees. Journal of the ACM, 59(4):21:1–21:19, August 2012. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yates:1957:APR

[YL57] F. Yates and S. Lipton. An automatic programming routine for the Elliott 401. Journal of the ACM, 4(2):151–156, April 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yu:1977:AER

[YL77] C. T. Yu and W. S. Luk. Analysis of effectiveness of retrieval in clustered files. Journal of the ACM, 24(4):607–622, October 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yu:1976:SMR

[YLC76] C. T. Yu, W. S. Luk, and T. Y. Cheung. A statistical model for relevance feedback in information retrieval. Journal of the ACM, 23(2):273–286, April 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yu:1982:TWI

[YLS82] C. T. Yu, K. Lam, and G. Salton. Term weighting in information retrieval using the term precision model. Journal of the ACM, 29 (1):152–170, January 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yoeli:1965:GCD

[Yoe65] Michael Yoeli. Generalized cascade decompositions of automata. Journal of the ACM, 12(3):411–422, July 1965. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Young:1955:OSD

[You55] David M. Young. ORDVAC solutions of the Dirichlet problem. Journal of the ACM, 2(3):137–161, July 1955. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Young:1956:REA

[You56] Robert L. Young. Report on experiments in approximating the solution of a differential equation. Journal of the ACM, 3(1):26– 28, January 1956. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Young:1958:ASR

[You58] Frederick H. Young. Analysis of shift register counters. Journal of the ACM, 5(4):385–388, October 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Youden:1963:IV

[You63] W. W. Youden. Index, volumes 1–10 (1954–1963). Journal of the ACM, 10:583–646, 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Young:1967:BWO

[You67] Tzay Y. Young. Binomial-weighted orthogonal polynomials. Journal of the ACM, 14(1):120–127, January 1967. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Young:1969:TTE

[You69] Paul R. Young. Toward a theory of enumerations. Journal of the ACM, 16(2):328–348, April 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Young:1977:OAP

[You77] Paul Young. Optimization among provably equivalent programs. Journal of the ACM, 24(4):693–700, October 1977. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yoeli:1964:ATA

[YR64] Michael Yoeli and Shlomo Rinon. Application of ternary algebra to the study of static hazards. Journal of the ACM, 11(1):84– 97, January 1964. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yao:1978:HBT

[YR78] Andrew C. Yao and Ronald L. Rivest. k + 1 heads are better than k. Journal of the ACM, 25(2):337–340, April 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yu:1976:PWE

[YS76] C. T. Yu and G. Salton. Precision weighting — an effective automatic indexing method. Journal of the ACM, 23(1):76–88, January 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yu:1978:EAI

[YSS78] C. T. Yu, G. Salton, and M. K. Siu. Effective automatic indexing using term addition and deletion. Journal of the ACM, 25 (2):210–225, April 1978. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yu:1975:FCT

[Yu75] Clement T. Yu. A formal construction of term classes. Journal of the ACM, 22(1):17–37, January 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yamauchi:2017:PFS

[YUKY17] Yukiko Yamauchi, Taichi Uehara, Shuji Kijima, and Masafumi Yamashita. Plane formation by synchronous mobile robots in the three-dimensional Euclidean space. Journal of the ACM, 64 (3):16:1–16:??, June 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yue:1973:OPR

[YW73] P. C. Yue and C. K. Wong. On the optimality of the probability ranking scheme in storage applications. Journal of the ACM, 20 (4):624–633, October 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yao:1976:LBM

[YY76] Andrew Chi-Chih Yao and Foong Frances Yao. Lower bounds on merging networks. Journal of the ACM, 23(3):566–571, July 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Yamakawa:2024:VQA

[YZ24] Takashi Yamakawa and Mark Zhandry. Verifiable quantum advantage without structure. Journal of the ACM, 71(3):20:1– 20:??, June 2024. CODEN JACOAH. ISSN 0004-5411 (print),

1557-735X (electronic). URL https://dl.acm.org/doi/10.
1145/3658665.

Yu:2018:CUD

[YZJ18] Haifeng Yu, Yuda Zhao, and Irvan Jahja. The cost of unknown diameter in dynamic networks. Journal of the ACM, 65(5):31:1– 31:??, September 2018. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3209665.

Zadeh:1972:TEE

[Zad72] Norman Zadeh. Theoretical efficiency of the Edmonds-Karp algorithm for computing maximal flows. Journal of the ACM, 19 (1):184–192, January 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Zafarullah:1969:FDS

[Zaf69] A. Zafarullah. Finite difference scheme for a third boundary value problem. Journal of the ACM, 16(4):585–591, October 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Zafarullah:1970:AML

[Zaf70] A. Zafarullah. Application of the method of lines to parabolic partial differential equations with error estimates. Journal of the ACM, 17(2):294–302, April 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Zarechnak:1959:TLL

[Zar59] Michael Zarechnak. Three levels of linguistic analysis in machine translation. Journal of the ACM, 6(1):24–32, January 1959. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Zhang:2025:SSC

[ZCLD25] Zihan Zhang, Yuxin Chen, Jason Lee, and Simon S. Du. Settling the sample complexity of online reinforcement learning. Journal of the ACM, 72(3):22:1–22:??, June 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Zeigler:1972:TFT

[Zei72] Bernard P. Zeigler. Towards a formal theory of modeling and simulation: Structure preserving morphisms. Journal of the ACM, 19(4):742–764, October 1972. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Zeleznik:1968:QNM

[Zel68] Frank J. Zeleznik. Quasi-Newton methods for nonlinear equations. Journal of the ACM, 15(2):265–271, April 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Zerling:1985:GBT

[Zer85] David Zerling. Generating binary trees using rotations. Journal of the ACM, 32(3):694–701, July 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/214141.html.

Zhandry:2021:HCQ

[Zha21] Mark Zhandry. How to construct quantum random functions. Journal of the ACM, 68(5):33:1–33:43, October 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3450745.

Zhuk:2020:PCD

[Zhu20] Dmitriy Zhuk. A proof of the CSP dichotomy conjecture. Journal of the ACM, 67(5):30:1–30:78, October 2020. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://

dl.acm.org/doi/10.1145/3402029.

Zielke:1968:IMS

[Zie68] Gerhard Zielke. Inversion of modified symmetric matrices. Journal of the ACM, 15(3):402–408, July 1968. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Zaroodny:1958:ASP

[ZL58] Serge J. Zaroodny and Tadeusz Leser. AYDAR, special purpose analog machine for yaw data reduction. Journal of the ACM, 5 (1):89–99, January 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Zhuk:2022:QMD

[ZM22] Dmitriy Zhuk and Barnaby Martin. QCSP monsters and the demise of the Chen Conjecture. Journal of the ACM, 69(5): 35:1–35:??, October 2022. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL https://dl.acm.org/

doi/10.1145/3563820.

Zohar:1969:TMI

[Zoh69] Shalhav Zohar. Toeplitz matrix inversion: The algorithm of W. F. Trench. Journal of the ACM, 16(4):592–601, October 1969. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Zohar:1974:STS

[Zoh74] Shalhav Zohar. The solution of a Toeplitz set of linear equations. Journal of the ACM, 21(2):272–276, April 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Zoltners:1978:DDB

[Zol78] Andris A. Zoltners. A direct descent binary knapsack algorithm. Journal of the ACM, 25(2):304–311, April 1978. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Zhou:1999:ABB

[ZS99] Yunhong Zhou and Subhash Suri. Analysis of a bounding box heuristic for object intersection. Journal of the ACM, 46(6): 833–857, November 1999. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/

pubs/articles/journals/jacm/1999-46-6/p833-zhou/p833-
zhou.pdf.

Zeng:2021:FJQ

[ZTX21] Yun Zeng, Jian Tan, and Cathy H. Xia. Fork and join queueing networks with heavy tails: Scaling dimension and throughput limit. Journal of the ACM, 68(3):17:1–17:30, May 2021. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL

https://dl.acm.org/doi/10.1145/3448213.

Zhong:2003:CTG

[ZW03] Ning Zhong and Klaus Weihrauch. Computability theory of generalized functions. Journal of the ACM, 50(4):469–505, July 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Zwick:2002:APS

[Zwi02] Uri Zwick. All pairs shortest paths using bridging sets and rectangular matrix multiplication. Journal of the ACM, 49(3):289– 317, May 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Zhang:2003:EPV

[ZY03] Yuanlin Zhang and Roland H. C. Yap. Erratum: P. van Beek and R. Dechter’s theorem on constraint looseness and local consistency. Journal of the ACM, 50(3):277–279, May 2003. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [vD97].

Zheng:2017:RNG

[ZZ17] Yu Zheng and Louxin Zhang. Reconciliation with nonbinary gene trees revisited. Journal of the ACM, 64(4):24:1–24:??, September 2017. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Zhang:2025:OMD

[ZZC^{+}25] Zihan Zhang, Wenhao Zhan, Yuxin Chen, Simon S. Du, and Jason Lee. Optimal multi-distribution learning. Journal of the ACM, 72(5):36:1–36:71, October 2025. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).